A Complete Bibliography of Lecture Notes in Mathematics (1975–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

29 November 2023 Version 1.09

Title word cross-reference

#1 [Rau75a]. #2 [Rau75b]. #3 [Rau75c].

(0, ∞, ∞) [EH79a]. (1, −1) [Sri78]. (2k + 1) [Law75a]. (A(U(∞))) [SV75c]. (¯∂_{b}) [Roc75a]. (G, τ) [KN78]. (≤ ∞) [Cos75]. (H) [Gue78]. (n + 1) [Kos79]. (p, r, s) [Mic78j]. (p ≥ 2) [Man76b]. (r, λ) [Mul78b]. (r, s) [Was76]. (X_{t−}) [Get75b]. (X_{t−}^{∗}) [Get75b]. 0 [Dem75, Het79, Ver77]. 1 [Dem75, DS78f, Gui79a, Rhi75, VVT77, Wei76h]. 1/2 [SS77b]. 10 [Hog75]. 1 < p ≤ ∞ [FJ75n]. 1: 2: 1 [vdAS79]. 2 [Dek79, FP78b, FP78d, GH78b, HM79b, Lac76, RT77, Smy78b]. 2 − (2n + 1, n, n − 1) [Bre79a]. 24 [Rob77b]. 2^{q} [Sri78]. 2 × 2 [Sha79c]. 3 [FR79a, Hem75, IŠ79, Joh79c, Joh79d, Joh79e, Joh79g, Joh79h, Joh79l, Joh79m, Mar79c, Sta77a]. 4 [Del77h, Mar79c, Moi77a, dS79]. 41/2 [Del77h]. 5 [Del77h]. 5/3 + B [Man76a]. 6 [Hog75]. 8 [Tak77a]. [0, 1] [Hof79a, Mis79b]. [0, ∞] [Bla76i]. [A] [Lad76a]. [G] [Lad76a]. ^{∗} [Fel77c]. ^{n} [Don76, Tou76]. ^{nd} [BLR75e]. ^{o} [Ste75c]. _{1} [Che76b, Ply79, Stø78b]. _{2} [Lod76]. _{∞} [SV75e]. _{o} [DO77]. A [Wil75f, Haz79a]. A(K) [Bri76]. A(M × S^{1}) [BLR75g]. A[X, σ, δ] [Les78]. A^{∞} [Cho76]. A_{1} [Blo79a]. A_{2} [Dub76a]. a_{3} [Tam78a, Tam78b]. a_{4} [Tam78a]. A_{∞} [May78a]. A_{p} [RR75e]. Ady = Bdx [Van79]. af [SV75b]. A[leftnormalfactorsemidirectproduct]F [FGR75f]. ℵ [Che76b, Oso79]. α [Low76, Maa78, Mol77e, Mol77l, Mol77o, Mol77n]. B [DK79, Phi79]. ¯∂ [DT77, Øvr76b, Rab78b, Sko76]. ¯∂f = F [Rab77]. β [Dub76a, Maa78, SNWC77a]. Bγ [Ser79a, SS78a]. mod2 [Gon78]. modp [Har78a, McC79]. C [EH76a, Gil78d, Raj77, Bin75c, GM79b]. C(S) [BJM78c, BJM78f, Gin76, Hei78, JW77, Mar76a]. C(X) [Bin75d, Bin75i, Csá77, HK79b]. C∗ [Ara78, Ars79, Dou77a, Mis79a, PS77, Sak78, Tak78a, Tak78b, AK78, FM79]. C[0, 1] [For79b]. C^{∗} [SV75a]. C^{1} [Rob75b]. C^{2} [Mok78a]. C^{∞} [EL76, Poé76b, Poé76e, Raï75, Tou76]. C^{r} [HPS77a]. C_{n} [McA76]. C_{n+1} [Coh76a]. C_{o}(M, X) [Beh79j]. c_{P^{n}} [ACH77g]. c_{w} [Car75]. cos πλ [Ess75b, Ess75a, Ess75c]. CP^{n} ↪ R^{4n−2α(n)} [Era78]. ∪ [HS79a]. d′′ [Kre75a, Uen75g]. d = 2, 4 [Tak77a]. D[0, 1] [Tay78c]. d′′ [Cat75]. D_{k} [Neu76]. δ [CL75a, Čud77, Ros78h]. δ^{2} [Cor79b]. Δ_{3}^{1} [Sol78a]. Δu + c(x, y)u = 0 [Hee76]. E [Bra79t, Mah76f, Mah76g]. E^{3} [Kui79]. E^{n} [Dav75b]. E_{2} [Mor78a]. E_{∞} [May76a, May76b, May77k, May77c, May77b, May77a]. ℓ [Del77g, Sri79a, SD77]. ℓ^{∞} [Tem76b]. ℓ^{n} [GD77a]. ℓ_{p}^{(n)} [DE77]. ϵ [Bre77k, Čud77, Gil78b, Hel79b, Mil78a]. ϵ_{5} [Pau78]. η_{j} [Mah79]. Ext(X) [Bro77]. F [Kra78a, Rob77a, Rud77b, Sha77c, Ste76b, Ste76a, BWK77i, Hel78b, Rhi75]. f(p_{n}) [Rhi75]. f(x, y) = 0 [Sch76k]. F_{2} [Ger79c]. F_{T} [Sim75a]. F_{t+}^{o} [YM76]. F ϕ = 0 [BR76a]. G [Dej78a, HM79b, Las79, MP78b, Mol77p, Mol77-28, Pet78b, Poé76a, Sch77b, tD79c, tD79h, KT75c, PPW78]. G′ [Lad76a]. G^{X} [Gui77a]. G_{n} [Kuk76]. Γ [BDPR76, BF78, DO77, Mol77t, SNWC77b]. γ_{0}(N) [Klu77]. Γ \ N [AT75e, AT75d]. GF(2) [Fra78a]. gl(2) [Gér79b]. GL(n, k) [Mul75a]. GL(n, C/R) [Clo79]. GL2 [LT76a]. GL_{2} [LT76a]. GL_{n}(F_{q}) [Spr75]. H [Gon78, Har78a, Hub78a, Hub78b, Kan78a, McC79, Rat78h, Wag77, Bra79i, Wag79]. H(X, X) [Gro79g]. H − R [AB75]. H^{1} [BM77b, DMY78, Dou77b, Mey75b]. H^{1}(R^{v}) [Mey77b]. H^{∞} [Sar76a]. H_{2}(Su_{n}) [Alp76]. Hol(V, P^{1}) [Nam79g]. HP^{∞} [CT78c]. I [Joh79i, Pet78b]. i: V → M [HPS77e]. IP^{1}(c) [Dek79]. IP_{n}(c) [Kre75a]. J [Jou77c, Kna79b, tD79j]. j(n, m) = j([n/2], m) + j([n + 1/2], m) + j(n, m − 1) [PR77b]. K [Ati77, Bla76e, Bre75a, DM78a, Dou77a, FM79, FI77, FP78j, Geo76, Gra76k, HL79a, Hau79, HR76, Hod75, HS75b, JS77, KS78a, Kar79, KV79b, dSLM75a, LS77, May77e, May78a, McL79, Mis79a, MS77, Rat78d, Rob76a, Seg77, Sna75, Sta79b, Ste76c, Tho79b, Wag76, Wor79, BF79b, Bra76a, Bro78c, GM77e, Jan79h, Law75a, Mul75a, Rei78a, Ste79]. K^{2} = p_{g} = 1 [Cat79]. K_{0} [Ell79]. K_{2} [Gre76, Rob79a, vdK76, Len76b]. K_{2}(Z[Z/5]) [Sha79c]. K_{m,n} [Chi78a]. K_{n} [Kuk76]. K_{O} [Goo79b]. κ [Mek77]. kO [May77h]. L [Beh79g, Buh78e, Car79c, CM79, GD77a, Hou77, Jou77b, Mar75, Mor79a, Ran79a]. L^{1} [FJ75a]. L^{2} [Lum78, Mos79, Shi75]. L^{2}(dm) [McD78]. L^{2}(R^{n}) [Cor79g]. L^{2}[0, ∞] [EWL78]. L^{∞} [Bai77b, CV77h, DN77, FJ75b, FJ75q, Sar76a]. L^{P} [Fer76b, FJ75n, BDE^{+}77d, BDE^{+}77a, BDE^{+}77c, BDE^{+}77b, BDE^{+}77i, Cos75, Don76]. L^{↑} [CV77h]. L_{3}^{h}(Z(G)) [HM79b]. L_{∞} [Nit77]. L_{∞_{ω}} [MR77j]. L_{p} [Akc79, BTW75d, Man76b]. L_{p}(2 ≤ p < ∞) [GMZ79]. Λ [JS76, Hof79b, Las77, Maa75, tD79a]. ⟨g, ϕ⟩ [BR76a]. (π_{i}0 → π_{i}^{S} → K_{i}Z) [Qui76a]. ≤ [Zee76a]. ≤ 0′ [Eps79g, Eps79l]. Lx = Nx [GM77h]. M [BK77k, Beh79h, Beh79g, Beh79j, Beh79i, Beh79k, Kös76b, MBY79, Rey78c, Fla76, Hal79b, Hej76d]. M^{μ} [Jam78r]. C [Erm76, Le 75, Suz78b]. C(t) [Coh79a]. C∗ [Suz78b]. C^{2} [Yam78b]. C^{n} [Blo77, GW79g, Roo78b]. CIP^{3} [DM79a]. CP^{3} [Moi77d]. CP^{n} [Dej78a]. E [Kol78]. N × N [Cai79]. P^{3} [Pie78]. P_{m}(C) [Kla79]. P_{n}(C) [BE78]. R [Cop78g]. R^{2} [Jol78]. R^{3} [Mor79c, NV79e]. R^{m} [Rot76b]. R^{N} [Sim78b, Tou76]. R_{+} [For78]. Z/p [Bro78c]. B_{∈} [dG75e]. B_{∋} [dG75a]. F(D, ℓ_{∞}, ℓ_{∈}, P, Q) [Sch75h, Sch75n]. I [May77d]. L [Cho77c]. ⇒^{∗} D [Eps79b, Eps79a, Eps79e, Eps79i, Eps79m]. μ [Gla76a, NR78a]. N [Bra79p, DP77, EGS75, FJ75p, Hit78, Poh75, Sto75b, AV79, Ber76a, Fel79, Kos79, MZ78, Moi77g, Pop77d, Sal78, Wag79]. n + 2 [Mol77j]. N = 1 [Tam78k]. N = 3 [Tam78j]. n > 0 [Mol77j]. n ≥ 0 [Coh76a]. n ≤ g [Nam79a]. O^{#} [Mat78a]. O_{β}^{N} [SNWC77m]. ω [Lod76]. ω∗ [Fra77b]. Ω^{∞}S^{∞}A [Kah78]. ω′_{X} [Bar78b]. ω_{ω} [Sol78a]. P [EGS75, Lub79a, dlH79b, AF78, BDE^{+}77e, BBL77, Bor75a, Bru79b, But75, Car77b, Con75a, Cos75, Fef75a, Gel76e, Gér75b, GD77a, Jam78m, Kat77, KOS76a, Kre76b, Laz75f, Lie77, Mat78a, Mul75a, Mul79, Rei78b, Rog77b, She75, Str79c, Voi77k, Wag76]. P (m) [Fra77b]. P (C^{4n}) [IP76c]. p > 0 [Fer79]. P_{n}C [Ree78b]. ∂ [Bol75, Rap76]. Π_{2}^{1} [Ric75]. ∝^{n} [Wil75d]. PSL(2, R) [Hej76a]. Q [Ber79e, Buh78h, DR75f, DR75h, Eic77, Wil76b]. Q(\sqrt{q}) [Eic77]. Q_{α} [BW77a]. R [Bou78d, Her75a, Her75b, Kra75a, Kra75f, RR78, War76l]. R(Σ_{k}) [Hof79e]. R(Σ_{k}⟨Γ⟩) [Hof79f]. R − μ [Bin75a]. R^{2} [Cam75, CY78, Dud79, dlR75]. R^{3} [Dud79, Jen77g]. R^{d} [Bal79, GKR77d]. R^{n} [Ahl77, dG75c, dG76b]. R_{n}(V) [Nam79a]. ρ [Heg78]. rmSGA41/2 [Del77b]. S [BJM78f, Hof79h, Iag75b, Pha75a, Raj77, Jur78a]. S(b) [Tam78e, Tam78n]. S^{1} [IP76d, IP76c, PE75, Bin76]. S^{1} × S^{3} [PE75]. S^{2} × T^{2} [Bra79b]. S^{3} [McG78, PE75]. Sℓ_{2}(F_{p}) [LM79]. SF(n + 1) [Coh76b]. SGA41/2 [Del77b]. Σ [Dic77, AEK78a, HK78a, Mey76b]. Σ_{2}^{1} [Ric75]. SK_{1} [Day76]. SL(2, R) [Nic75a]. sl(3, C) [Dix75]. SL(n, C) [Duf79a]. SL_{2}(k) [Spr77a]. SL_{2}(Z) [Zag77a]. SL_{n}(k) [Rub75, Rub75]. SL_{n}(R) [Rub79]. SO(3)R^{3} [Fal79]. SO_{o}(1, q) [Har75b]. Sp(n, R) [CO79a, Sch75b]. spin_{0}(1, d)Spin(d − 1) [Tak77a]. SU(2) [Cle75, ER75]. SU(n, n) [CO79b]. SU(p, q) [CO79a]. T [Gol78a, Osi78, Raţ79]. T^{3} [Her77a]. T^{f} [Sim75a]. τ [Hof79g, Hof79h, Hof79c, Hof79d]. J̃(X) [SS78b]. U [Get75p]. U(G) [Aub79]. U(∞) [SV75d, SV75f, SV75e]. \underline{P ic}_{X̃/k}^{#} [Bou78d]. UV [PE75]. V [Cou79, Woo78a]. V = L [Dev77a]. [vDash] [DM75c, MOP75]. W [Cor79b, Jos79]. W (X) [Pfe75]. W ∗ [Tak78c]. W_{∗}(Q, C_{n}) [ACH77a]. ℘ [Gér75c]. X [Bar78b, Bou78d]. x_{0}(N) [MS76c]. X_{1} [Ste78f]. X_{2} [Buk76]. X_{o}(N) [Ser79b]. y^{d} = f(x) [Sch76n]. Y′′ = Q(t)Y [Bor79c]. y^{q} − y = f(x) [Sch76n]. Z/pZ [AF78]. ZF + L^{1} ≠ HOD [Szc77a]. ZFC + V ≠ L [Bel76].

* [Bar79c, Fel77b, Fel77d]. *-algebraic [Fel77b, Fel77d]. *-Autonomous [Bar79c].

-actions [Cam75, dlR75]. -Adic [Gér75b, Gel76e, Mul79, Sri79a, SD77, Wag76, Jou77b, Lub79a]. -Adique [Mul75a, Bor75a, Del77g, Mul79, Jou77b]. -adiques [Car77b, Gér75b, Gér75c, Jou77c]. -algebra [BDE^{+}77e, SV75a]. -Algebraic [Fel77c]. -Algebras [AK78, SV75b, Mis79a, Dou77a, DR75f, PS77, Sak78, Tak78a, Tak78b]. -Algebren [Jan79h]. -algèbres [Ars79, FM79]. -Algorithme [Gil78b, Bre77k]. -Analytiques [Her75a, Her75b, Le 75]. -Anneaux [BWK77i, Las77, Ber79e, Cou79]. -Barrelled [HK78a]. -bases [Wag79]. -Bedingung [RR75e]. -bundles [Joh79i]. -calculus [Maa75]. -categorical [Che76b]. -closed [Kre75a]. -cobordisms [Wag77]. -cohomologie [Cat75, Gui79a]. -cohomology [Mos79]. -conjecture [Rat78h]. -connected [Wor79]. -contact [Bla76e]. -convergence [Nit77]. -convex [DK79]. -convexes [VVT77]. -cycle [HR76]. -DG-algebras [KT75c]. -dimensional [Ber76a]. -dimensions [Uen75g]. -Embedded [Bin75c]. -equation [Øvr76b, Sko76]. -Equilateral [Law75a]. -equivalence [GH78b]. -estimators [MBY79, Rey78c]. -extensions [Kre76b, LT76a, KN78]. -fermées [Kre75a]. -fields [Mey76b]. -flots [dSLM75a]. -flows [Fel79]. -folds [IŠ79]. -free [Mek77]. -functional [JS77]. -functions [Mah76f, Mah76g, Mor79a, Tam78n]. -Galerkin [Dou77b]. -gons [Law75a]. -group [Fef75a, HM79b]. -groups [But75, Con75a, Lie77, Rei78b, She75, Car79c, CM79, War76l]. -harmonic [Gla76a]. -homogeneous [Sch77b]. -homology [Dou77a, KS78a, Seg77]. -homomorphism [Kna79b, tD79j]. -Ideals [Beh79h]. -independence [Hel79b]. -indexed [Woo78a]. -Intégrables [Bin75a]. -invariant [Heg78, NR78a]. -invariantes [Poé76a]. -Kalkül [Maa75]. -lattice [CL75a]. -Liealgebra [Voi77k]. -Local [Bra79i]. -locally [AEK78a]. -manifold [Joh79g, Sta77a]. -Manifolds [Joh79h, FR79a, Hem75, Joh79c, Joh79d, Joh79e, Joh79l, Joh79m, Mar79c, Moi77a, dS79]. -map [Raţ79]. -matrix [Iag75b, Wil76b]. -measures [PPW78]. -models [Dub76a]. -module [Jos79]. -modules [Haz79a, AF78, HK79b]. -Moduln [Kra75f]. -Moduls [Kra75a]. -monogenic [Bra76a]. -movability [DS78f]. -nd [RT77]. -nets [BBL77]. -normal [Sal78, Bra79t]. -operators [BDE^{+}77i]. -oriented [May77h]. -ple [Hit78]. -point [Mat78a]. -points [Fra77b]. -primitive [BF79b]. -problem [Ess75a]. -projections [BDE^{+}77a, Beh79g]. -quasi-cohérent [Ste76b].

-quasi-cohérents [Ste76a]. -recursion [Low76, Maa78]. -regular [Jam78m]. -representations [DR75h]. -resonance [vdAS79]. -ring [Hof79h]. -Rings [Hof79c, Hof79b, Hal79b, Hof79g, tD79a, Hof79d]. -saturèe [Dic77]. -Schéma [Bou78d]. -selfdecomposable [Jur78a]. -semi-stable [Jur78a]. -series [Buh78e, Cho77c]. -set [GM77e, Osi78]. -sets [tD79c]. -smoothing [Las79]. -space [AV79, Rob77a]. -spaces [Fla76, Hel78b, AEK78i, Coh76a, BDE^{+}77d, BF78, Get75p, Gon78, Har78a, Hub78a, Hub78b, Kan78a, Kra78a, McC79, Raj77, Rud77b, Sha77c]. -spheres [Kos79, Mar79c]. -Splines [JS76, Kös76b]. -stable [Was76]. -stars [Hog75]. -structure [Pop77d, BDE^{+}77b, Beh79i, Beh79k, BDE^{+}77c, Beh79j]. -Structures [Mol77-28, Mol77t, Mol77p]. -systems [Mul78b, Geo76]. -table [GM79b]. -tensor [Mic78j]. -term [Mor78a]. -th [MZ78, DP77]. -theorem [Ess75b, Hod75]. -théorie [Bre75a, FM79, Kar79]. -Theory [MS77, Ste76c, Bro78c, Rei78a, Ste79, HS75b, Ati77, DM78a, FI77, FP78j, Gra76k, HL79a, Hau79, LS77, May77e, May78a, McL79, Mis79a, Ran79a, Rob76a, Seg77, Sna75, Sta79b, Tho79b, Wag76]. -times [Poh75]. -transversality [Dej78a]. -trees [Rog77b, McA76]. -tuple [Kos79]. -types [KV79b]. -unions [Oso79]. -valued [Mar76a, Phi79]. -variation [Bru79b, Str79c]. -vector [tD79h]. -versel [Poé76a]. -well-posed [Shi75]. -wertigen [Bin76]. -width [FJ75p].

/base [RR78]. /socle [RR78].

0 [RR75a, RR75g].

1 [Mol77z]. 17ème [Boc78]. 18th [NR78b]. 1967/68 [Ano79]. 1972 [BK75a]. 1972/73 [Ste75c]. 1973 [CK76, KZ76, Lel75b, Man75d, Pha75b]. 1973/1974 [CK76]. 1973/74 [Bou75, KZ76, Lel75b, Man75d]. 1974 [BOS75, CDV75, Cha75c, CK76, DM75c, DG75j, Eve75a, FCG75, GR75b, Gol75, HM76a, JMR75, Kel75c, Lel76, MOP75, Ode75, Pin75a, Rau75e, Ros75b, SW75d]. 1974/75 [Ano76, Lel76]. 1975 [BØS76, BK76a, BH76b, BR77c, BMS76, BCN77, BR76d, CW76a, CUW75, CS76d, GM77n, GN76, GDMW76, Hil76, MSZ76, MS77, Nor75, Tem76c, Wat76, Zel76b]. 1975/76 [Ano77, Lel77a]. 1976 [AL78, AHW77, BKK76, BCD77, Bar77b, BR77b, BGM77, BS77, BGS78, CV77a, CCJS79, DMW78, DKT78, ES76, Foa77a, GR76, HJLN77, Lel77a, LS78, Lit77a, MWW76, Nat77a, PdC77, SS76a, SZ77a, Sch78g, SZ77b, SZ77c, Ste76c]. 1976/77 [BDV^{+}78, BR77b, DMW78, LS78]. 1977 [Ano79, AT78, AK78, AD78b, AD78a, BH78, BM78a, BM78b, BR78a, BRP78, Cha78a, Cop78i, DCHRH78, DCV78, DMW79, Fáb79, FW79, Fen79, FMS79, GLL79, Hau78a, HH78b, HPS78, HS78, KK78a, Lai78, MMP78, Mil78b, MS78d, Nas79b, Nor78c, Ols78b, Wat78b, Wer78]. 1977/78 [Ano79, Ano79, DMW79]. 1978 [DJ79, DM79b, GM79d, HL79b, HP79, Hey79, HS79b, HW79b, LLS79, LS79a, Løn79b, NR78b, PW79, dlH79c]. 1979 [GR79b, Nat79, Wuy79b].

2 [Mol77-27]. 27th [HP79]. 29ème [Mal77]. 29th [Hey79]. 2nd [AHW77, HP79].

30ème [Mal78a]. 31ème [Mal79c]. 3R [Mal79d].

452 [Bou75]. 470 [Ano76]. 488 [Ano77].

506 [BDV^{+}78].

65th [DM75c].

73 [Par75a]. 74 [KZ76, Lel75b, Man75d]. 75 [ETFS75, Lel76]. 77 [BR77b, DMW78, KM78b, LS78]. 78 [Ano79, DMW79, ETFS79].

= [Rog76b].

A. [Bor78c, Cho75a, Ess75j, Ess75g, Fel77c, Lan78, Le 75, Poi77]. a.e [Bru78e]. Aarhus [DM79b]. Abbildungen [GMK75c, Küh79, RR75f]. Abel [Buc77]. abelian [But75, Byc79, Che76e, DR75i, Dus79, Ekl77, Ish76h, Kap78, KN78, MP79b, Mek77, Pop77i, Rog77b, Sch78f, War76b, War77b, Acc75d, AHW77, AT75a, BW77a, Fef75a, Fis77, FGR75i, JR77, Kre76b, She75]. abelien [Ber76c, LL76]. abeliennes [Gro77f]. abeliens [Nic77]. abhängige [Cré79]. Abhängigkeit [Bro76a]. Abhyankar [Cor79i]. Abolition [Sus75]. Abraham [SW75b]. Abschätzungen [Don76, Ess76]. absence [Dav75a, Eas76, GW79j]. absolument [Cou79]. Absolute [Bak76, Bak78a, Jac79g, Kun76, Mañ75a, Arm75, AH76, Bak79a, Cha78b, KLŠ76]. Absolutely [Sch76k, Sch76l, Sch78e, Won79g]. absorbing [AEK78r]. absorption [Sai79c, Wil75g]. Abstract [BK77a, Cor79e, Cuy79, Kea75, Mol77a, Mol77b, PS78, Ree76a, Aiz79, BK77b, BK77c, BK77f, Clé79, DSS76, Fel77a, Gua78, KMS75, Kuo75d, Mil79, Pet79, Roc75b]. Abstrahlung [Fri76b]. abstrait [BG78a]. abstraite [BCH78]. accélération [GB79, Bre77f]. access [BW75]. Accessibility [Ste75a]. accompanying [Kru76]. according [Alo79, Jan78]. accroissements [dSLM75c, Sid79]. acoustiques [Lév77]. Actes [CDV75, CV77a, CV79, CK76, Foa77a, Rau75e]. acting [Kaw75]. Action [NT79a, Raï75, Ker75a]. Actions [ACH77f, Dav78b, Dav78e, Mol75, Pag79, Ren79a, ACH77g, AH79, Ass79, Bot78, Cam75, CN77, Dav78c, Dav78i, EH76d, Fer78b, Got78a, GL75, Hil75, Hil78, IP76c, Jon79, Liu78a, Liu78b, Liu79b, NT79f, War76g, dlR75, tDP79]. activity [Kry78]. acyclic [Mar79c, Rob77c]. Ad [Kec78a, Sol78b]. Adams [Bru78s, Hof79b, MP78b]. adaptés [Fey78b, Fd79, de 78]. addenda [Lad76a]. Addendum [Sch78a]. Additif [Ram78b, Ram78a]. additifs [Yoe79]. Addition [Dov79, Uen75a, FR79e]. additional [FR79i, FR79j]. Additive [Gri79a, Gri79b, Wag79, EN79, Par76, Sha79a, Tur78]. additivity [Pop77a]. Adelaide [CW76a]. adéquation [Tua78]. Adic [Gér75b, Gel76e, Jou77b, Lub79a, Mul79, Sri79a, SD77, Wag76]. Adique [Mul75a, Bor75a, Del77g, Jou77b, Mul79]. adiques [Car77b, Gér75b, Gér75c, Jou77c]. Adjoined [Zie79a]. Adjoint [Far75h, Yan77a, AdS75b, AdS77, Far75g, Kal78, MMS78, PS78, Wol78]. adjoints [GV79c]. admitting [Hof76a, Hof79l, War76i]. Advanced [Bar77b]. Advances [Gau76, Bor78c, KZ76]. AF [SV75d]. AF-Algebras [SV75d]. Affine [Mat77c, Che79, Coh79b, DeM76b, Éli79a, FJ79, Jen77f, Mat77g, Pop77k, Rob76a]. affines [Mat77b, Mat77f]. Afterword [Bre77a]. agrégatif [Jol78]. Ahlfors [Ess75e]. al [Tit78]. Alberta [GM79d]. aléatoire [Éli79b, LL76]. Aléatoires [GKR77b, GKR77e, Art75, CG79a, Cré76, CL76b, ER75, Fer75, GKR77d, GKR77f, GKR77g, GKR77c, Hen79b, Jac79k, Kôn76, PS79, Roy75, Roy78]. Alfred [dBR77]. Algebra [Che76j, FGR75i, FMS79, HS79a, NR78b, Nov77a, SW75b, vO75c, Ama76b, AdS75a, BK77c, BK77f, BW77a, BDE^{+}77e, Blo79a, CM77, Che76k, DR75e, EH76e, Gui77b, Hof76b, Jos77, Jos79, Kad78, Lan79a, MN75b, Ply79, Ser75a, Sha79b, Sib76a, Sim78d, Spr77b, SV75a, Var77b, Cro75b, Voi77c]. Algebraic [ACH77b, Bor79c, Bro75a, Bro75d, Bro78c, Bru78s, CBR76, Del77b, DS78a, Fel77c, Han78b, Hau79, Ish76a, Ish76d, JW78, Kas78, Lau79a, May77e, Moi77a, Pet76b, Pop77g, Rog77b, Sto78a, Uen75b, Uen75f, Wei76a, Yan77b, Yan77c, BR75, EH76a, Ent78, Fel77b, Fel77d, FJ75s, FM78, GT77, Gra76k, Gro78o, IP76a, IP76c, Joh75a, Kaw79a, LS77, Lub79a, Mah76e, Mah76d, Mah76f, Mah76g, Mas75d, May78a, McK79, Moi77d, Moi77g, Moi78, Pop77b, Pop77e, Pop77f, Pop77h, Pop77j, Pop77k, Pop78, SW75a, Sar78, Sch76r, Seg77, Sna78, Sri79i, Ste79, Tak78b, Tho79b, Uen75c, Uen75e, Van78b, Cor79a, DM79b, HPS78, HS79b, Jou79d, Løn79b, Mil78b, Ols78b, Ste76c]. Algebraic-topological [Han78b]. algebraically [Bro78d]. algebraische [Fel76, Kra75d]. algebraischen [Voi77i, Voi77n]. algebraischer [Voi77s, Voi77h, Voi77f]. Algebras [AK78, AT75a, BK77a, DG75j, Jac75d, Kup75, Mis79a, MS77, NT79c, Sar76a, SV75d, Yan77t, BV77, Bar79a, BK77h, BK77i, BK77g, Bar78d, Bin75f, Bin75h, Bor78c, BP78, CP77a, CM77, Cho78a, Cic77, Cor79g, Cum76, Dal76, Day76, DG76c, Ded78, DW79b, Dou77a, Dou78, DR75f, Fla77, FGR75b, FGR75h, Gei77, Goo76c, Goo76f, Goo79c, Gui77b, Haa79, Hae78a, HL79a, Her78a, HW75d, Irv79, Jac75b, Jac78a, Joh77, Joh79p, Kai76, KT75c, KT75j, La 76, Lad76b, Lau79e, Mac75b, Mah77, Ne’77, Ols78a, Orz76, PS77, Peł76, Rin75, Rob77d, Roc77, Ros78a, Ros78b, Ros78c, Sak78, Sło78, Smi76f, Smi79, Sni79, SV75b, Sze76, Tak78a, Tak78b, Tay77b, TY78, Var76, Wei76a, Wei76d, Wei76f, Wei76p, Wei76r, Wil78b, Yan77j, Yan77o, vO75d, Jac75e]. algebras [Sai78]. Algèbre [And79b, Gue75, Ser75a, Art75, Boh75, CB75, Dix75, Gap76, Her75a, Mok78b, Mok78a, Ren79b, Yam79, dlH79a]. Algebren [Jan79h]. Algèbres [dlH79c, Ars79, Aup79a, Aup79b, Aup79c, Bar78c, FM79, Kar79, Nac78, Ren77, Vie78b, Vie78d, dlH79b, Aup79f, Vie78a, dlH79c].

Algébrique [Del77b, Gil78a, Ill77b, Cor79a, FM79, Gro77e, Jou79b, Kar79, Le 77, Ram78c, Sab76, Ser77]. algébriques [BF75d, BF75e, BE78, Jou79b, Jou79e, Jou79d, Jou79g, Tit78, Ver76, Ver79a, Voi79]. Algorithm [Tat75, FJ75p, GM77p, Geo78a, Kur78, MJ75, Mor78c, Pow78a, Sch79j, War77a, Wei76b, Wei76c, Wei76n, vdLT79]. algorithm-theoretic [Kur78]. Algorithme [CL76a, Gil78b, Lau75c, Bre77k]. Algorithmes [Bre77g, Bre77i]. algorithmic [Bör75b, Ras75]. algorithmiques [Lal77a]. Algorithms [Spe75b, AG75, Bul79, Gal76a, Geo79a, Law76, Sim78d, Sim78e]. allé [Dar77]. alliance [Ben76a]. Almost [Aus75, AR75, BJM78a, Bla76a, CP77a, Glu75, PT76, Phi76, Phi79, BD79, Bra79e, Bra79k, Cop78g, DR75i, Kas79, Sha77b]. along [Gre79, Nør78d, vdE79]. Alternanten [Som76]. Alternative [GM77a, Loo75a, Loo75b, BB75, Chu77a, Dev76, Kri76, Soc76, Soc77]. alternatives [Neu76]. amalgames [Lal77b]. amarts [Bel78]. amenability [Gla76b]. American [FW79, Nas79b, PdC77]. among [Sha79c]. Ampère [Sib77]. amphibia [Zee76d]. amples [Ray78b]. amusing [Doh75]. analogue [Lam78d]. analogues [Hof79i, Hol76, Koo77, dM75a]. analogy [ADB75]. Analycité [Dlo78a]. Analyse [BR76b, Bre77f, Gil78h, Kré78a, Lel75b, Lel76, Lel77a, LS78, Ler78, Mat75, Mat77b, Mat77c, Cle75, Kré78b, Lel78b, Mat77h, RS78, Tak79, Tem76a, ETFS75, ETFS79]. Analysis [AT75a, Bre77d, BS77, CDV75, CV77a, CV79, CV77c, CCJS79, FJ75l, FMS79, GN76, Goo76d, GM79d, KK78a, LLS79, Nas79b, Nov77a, Var77a, Ver79c, Wat76, Wat78b, BH76a, Bec79b, Ben76c, Cap77, Cop78b, Dah76, Eps79c, FH79, Gal75, GH77, Hid78, Hla79, HK79b, HC78, JL77, Kec78b, Ler78, Mal75b, Mor76, Oli76, Rus77, Sar77, Sij79, Var77c, Var77b, YS79, DV75b]. Analytic [BCD77, BK77a, BK77d, BØS76, Bib79a, Bib79c, Dwy78b, Fis76d, KS78a, Mor77a, Uen75c, Ber76a, Bin75g, CM77, Cow78, Ede76, Fer76b, Gar77b, Hir75b, HW75k, Ike79, KRZ77e, Mag79a, Mah76d, Mat77a, Mil76, Muj78, Nie77, Pop77e, Sha76a, Sha77c, Sib76a, Sre78, Ste76a, Sto75b, Tog76, Tra76, Tra78a, Uen75a, Wag77]. Analytical [BDNY77, Zie79b, Ben76a, Džu76, Kre75a, Pin79]. Analyticité [Sam76, Ava77]. analyticity [Iag75b]. Analytique [Băn78, AN78, Bar75, Bar78b, BR76b, Cho77b, Fou75, Hen75b, Maz76, Ram78c, Ste76a, Tei75, VR75]. Analytiques [Her75a, Her75b, HH78b, Lel77a, Ali75, Bar75, CDI78, Dlo78a, Dlo78b, Erm76, Gér78, Gru77, Hen75c, Hen75b, Jou79f, Kan75a, Kan75b, Kil78, Kre75a, Le 75, Maz77a, Ste75b, Suz78a, Thi77]. analytische [Dri79]. Anelastic [Blo79b]. Anfangswertproblemen [Mic76c]. Angeles [AK78]. angeströmten [Fri76b]. Angles [Val75]. angular [Vuo79]. Anhang [GMK75a, Kra75a, KPP^{+}78a]. Anhang. [Haz77a]. anharmonic [AHK76e]. Anisotropic [Dev78]. anneau [Ari79, Fer77, Les78, RR78, RR79a, Tou76]. Anneaux [BWK77a, BWK77b, BKW77, BWK77i, GP79, And79b, Bas75, Ber79e, Cou79, Las77, Mar79a, Pau78, Que77, Ren79a]. Année [Lel75b, Lel76, Lel77a, LS78, Mal77, Mal78a, Mal79c]. Annihilator [Sch78f]. annulus [Ław79]. Anosov [Bow75b, Mañ75b, Man75c]. ANSR’s [DS78c]. anti [Ład75]. anti-commutation [Ład75]. antiautomorphisme [Ren77]. Antiautomorphisms [Pet76a]. anticommutative [Kré78a]. Antisimple [Hau78b]. antisimples [Hau78b]. antisymmetric [SV75f]. Antwerp [BK75a, Wuy79b]. Anwendung [HJSW75, JLR76]. Anwendungen [Col76c, Vek76]. août [And77a]. Apercu [Spe75a]. aperiodicity [Lor79]. Aplicada [PdC77]. appearing [Śni77]. Appendix [Bar79c, Fel77c, Fis76a, Iag75a, Lan77, Olv79, Raţ77, Web78b, GMK75a]. Application [BM79a, BDR78, Ben77b, Das75, DKT78, FM79, FJ75c, FP78l, HS75a, Lel78a, Miy77, Mor76, Sch76d, Sch75f, Sch75e, Sch75g, Śni78, Tam78a, Ver79c, vO75a, Ber79c, Buc77, Car78c, Dra79, Egu79, Geo79a, GM77p, Gue75, Heg78, Ker75a, Ker75b, LeP78a, Maz76, Ne’77, Nix78, Pel78b, Ren77, Rog79a, Ros78i, Rum79, SV79, Saw75, Sch78b, Sco76, Sim75b, Sko76]. Applications [AL78, AT78, AK78, AD78b, AD78a, BF75a, BM78a, BM78b, Bis78a, Bis78b, Bou78b, BTW75b, CV77h, Cha78a, Cun75, Del77a, Dev77a, Dup78a, EH76c, FJ75l, FGR75i, FMS79, Gil78h, Goo76b, Goo76d, Gre78a, Gro78c, Hal79c, Hey78, Hil76, HW75a, HM77, Jac75b, JPR^{+}78, KK78a, KRZ77a, Lau75a, Lie78, Mah76a, May79, Miu76a, Mol77q, NT79b, Pop77a, Pop77b, Ros75b, Sat79b, Sco79, Sha78a, Sha78b, Sjö75, Tem76a, Wuy79b, AB75, Ama76a, AG75, Ars79, Bak76, Ban79c, BG78b, BG78d, Bur76, CO79a, Cam78b, CV77e, Cha76c, CMN79, Col75, Com77, CT76, EH76e, Eva77, Frö79a, Fuk76, GS76, Gal76j, Ger79a, Gra79a, Hac78b, IP76c, JS77, Kat75, Kue78, Lan79a, Lau79g, Led76, Lel78b, Lum76a, Lum77, Mar76d, Mém78, Mor77d, Mul78b, MRN78, Oli76, Pop77i, Rab78a]. applications [Reb78, Rob76e, Rog76a, Rog76b, Ron75, Ruc78b, STW78, Sen78, Sib76b, Sna78, Sto75b, Tak78a, Tay78g, dBS76, BF75a, FdlP76, Fey78a, Fd79, FMS79, GN76, Yor78b]. Applied [GLL79, Ale75, BH76a, Bar79h, Cot79, McC78b]. applying [EW78]. appraisal [BP76]. appreciation [Ray78a]. Approach [GK78d, Sch76p, Ahl77, AL79, Cur75c, Czy78, Dyn76, Fen75, FR76, Fri79, Gar76a, Gla76k, Gri76a, Joh79r, Kri76, McC79, Nag76, Neu76, Par79c, Rob76a, Sco77, dF75]. approaches [Bul79, LY79]. approximable [Pol79]. Approximant [Gil78g]. Approximants [CC79, Gil78b, Gil78c, Gil78a, Gil78d, Gil78h, Wuy79a, Chi78b, GM79c, Pin79, Cuy79]. Approximate [Dei77a, DW79a, DW79b, Mar79d, AL79, Bru78b, Bru76e, DW79d]. Approximating [Goo79c, Han75b, HW75b]. Approximation [Bai77a, Ban79c, Col76a, Cro76, DG76a, DDC79, Dub79a, FJ75d, FM76b, GM77b, GR79e, PW79, Rot76a, Sch77b, Sjö76a, Som76, Wuy79b, Bar78e, Bar79h, BK76g, Bru76a, Bru75, Cox78, DS76b, Fer76b, For79b, FS79b, GH79, GP77, Han78b, HP77, Her78a, Het76, Jer76a, Jer76b, JL77, JS76, Köh75, KO79, Lyc76, May79, OW78, Rob77a, Ryc78, Sar76b, Spe75b, Woo78b, Bra76b, SS76a, Str76, vG76]. Approximationen [HK76]. Approximations [Col76c, SV79, CG79b, CB76, DN77, DNR79, DDP78, Erd75b, Kik79, MW79, Nic75b, Nit77, OW75, OL76, Red79, vH79a]. Approximations-theorie [Col76c]. Approximationseigenschaft [KPP^{+}78c, KPP^{+}78d, KPP^{+}78e, KPP^{+}78i]. Approximationssatz [HJSW75, KPP^{+}78g]. Approximationssätze [KPP^{+}78b]. approximierbarkeit [Bla76i]. April [AK78, CS76d, ES76, GR76, GR79b, MWW76, MS77, MS78d, SZ77a]. Apropos [Sza78]. Äquivalenzsatz [Jur78r]. arbitraire [Dam78]. arbitraires [Geo78b]. arbitrary [DS76b, Geo78b, KT76, Wil75e]. Archimedean [BWK77c, Gel76d]. Archimediennes [BWK77c]. Archimediens [BWK77d, BWK77f, BWK77i]. areas [Dej78b, Roa75]. arguments [Edr77, Kue75a, Red79]. arising [BPFS76, GW76a, LS79b, Wei76d]. Arithmetic [HW75j, BS75, HW75a, Hir75c, HW75q, Lei75, MZ78]. arithmétiques [Des79b, Gra78b, Tho79c, Wal76, Wal77b]. Armidale [HW79b]. Arnol’d [Dem75]. Arrêt [CG79a, El 78]. arrangements [Rob76d]. arrays [EH78, Ład75, LW75, SM79a]. Art [Häm76, Nic78, Mei78]. artificial [vdLT79]. Artin [Gel77, Mül75b, Pop77j]. Artinian [Jat76, Len76a]. Asano [Rob76e]. ascendantes [Nic77]. ascending [Hal79h]. Aschbacher [Jan78]. Aspects [Eff78, GM77n, Tho79a, Azu79, Bab79, Bea76, Bel78, BP79b, BEGS75, Ede75, Fin76, Gor76, Gor78a, Roq75, Smi77, Sta77b, Trn77]. assisted [CS75]. Associated [FP78f, Ghi79a, Sil76a, And77k, Bro78d, Cod79, CH78d, CL75f, DS78a, Fed78, FP78g, FP78h, Han78c, Har75b, McD78, Pas79a, Run76, SV75a, Wol78]. Associative [Ros78b, Gon78]. associé [Éli79b, Nan75, Sec75]. associée [Fed78, GD77b, Rit76]. associées [CO79b, Har75b, Le 76, Let79]. Associés [CH78d, Gri77, LL76, Sch76f, Sch76g]. assorted [BG78b]. Assoziierte [Jan79b]. assumptions [Lan76a]. Asymptotic [Bih79, BK76f, Dur76b, EL77a, GV79b, Gre75b, Gro78d, KT76, Lei79, Lio76, Sai79a, Ver79c, Vuo79, Wei78a, Wil75b, vH79a, Bra79a, Bro75g, DV75c, EWW77, Har76b, Lev79a, Roc75b, SM79b, Wil75i, vdB79, Gat77]. asymptotically [Hes79]. Asymptotics [BLP79, Ell75, Roc75c, Lon79a, Trè75]. asymptotique [Gat77, Kôn76, Lép78c, Lum76a, Roy76a]. asymptotiques [Ler75, dV75a]. asymptotische [Jur78q, Jur78b]. Athens [MS77]. Atiyah [Sha78c, Sha78d]. Atlases [Aki78a]. atomic [Gra76c, HvW79]. atomique [BM77b]. Attached [Yan77t, Rib77, Yan77o]. Attaching [Joh79a]. attaining [JW77]. Attempting [Kea78]. attempts [Gla77, Mar77]. attraction [Gin79, Hah79]. Attractive [Pre76b]. attractor [Lan77, Par79b, Rue76]. Attractors [Man77, MMP78, HP76a, Kol77, Wil77a]. Aubin [Bou79b]. Aufgaben [Nic78]. auflösbaren [Voi77i]. Augmentation [Pas79d, BLP77, Pas79f, Pas79i, Rog75b]. August [Bar77b, BK75a, CW76a, Cop78i, DM79b, Fáb79, FW79, HP79, HPS78, HS78, HW79b, LLS79, Lit77a, LS79a, Løn79b, Pin75a]. Austin [Ode75]. Australia [Cop78i, HW79b, Cro75b]. Australian [BCN77, CW76a, Cro75b, HW79b, Lit77a, NR78b, SW75d]. Authors [Kue75b, LMW75]. automatic [Geo78a, Lau76]. Automatique [Tik79]. automorphes [Bor76]. Automorphic [Gel76a, Gel77, Ply79, Cow78, Hej76d, Kud77, Mor77a, Mor77b, Pom76]. automorphism [AdS75b, AdS77, Ber77a, Cho78a, Rob75a]. Automorphisms [Acc75d, BLR75a, BLR75d, Fre75, LT76a, Rob79c, Bro78e, Kad78, KK79, Lie78, Lin79b, Liu78b]. Autonomous [Bar79c, Alt79a, Bih79, Nus79, Wan79]. aux [Bon78b, Cam78b, CCP75, Cor79a, Del77a, FM79, FdlP76, Fey78a, Gue75, GKR77b, Hec76, Kôn76, Krá76b, Kré75b, Lai78, Las78, Mey76a, Sch76d, Sch76g, Tar78, Tho79a]. avenues [Rey78e]. averages [DGS76r, MP79a, Peł75]. Averaging [Man79, BM79a]. avoidance [Rat78o]. avril [Foa77a]. AW* [Sai78, Tak78c]. AW*-algebras [Sai78]. AW*-factor [Tak78c]. AW*-factors [Sai78]. ax [ET79, Che76c]. Axiom [Bow75a, Bow75g, Dev77c, Kle77i, Bow75c, Dev76, Dev77b, Kle77h, Sus75]. Axiomatic [GT75, Fen75]. Axiomatizations [HW75c]. Axioms [Fla75, MR77a, Boo75, Fef75b, Fri78d, KM78a, KP77]. axis [FJ75j]. ayant [Sab76]. Azumaya [La 76, Sze76].

B [Ste78e, ET79]. B. [Car78c, Ram79b]. B.P. [Gro78j]. Back [Kue75a, Ser75b, Ser75c, Cun75, vH79b]. Back-and-forth [Kue75a, Cun75]. Background [EH76b, Gel76b, Jam78a, FR79f, HH77e, Sta78b]. Bäcklund [Che76a, FM76a, Lam76, Miu76a, New76a, Rog76a, Run76, Sco76, Wah76]. Baernstein [Ess75j, Ess75g]. Baire [Bru78d, Cic77, Fro79b, Gol78a, KS76]. Balanced [AH79, SM79a, Mor79b, Mul75c]. Balancing [MP79a]. balayage [Fey78b]. balls [Top76, TDHJ76]. Banach [Bec76b, Bec79a, AEK78b, Aki78h, Ama76a, Aro76, Aup79a, Aup79b, Aup79c, Aup79f, Bad76, BCD77, Ban79a, Ban79b, Bec76a, BW76, BDE^{+}77b, Beh79a, Beh79b, Beh79f, Beh79i, Beh79k, Bor79a, BS76b, BM79b, CP77b, Car79d, Cha79b, CG79b, CBR76, Cig76, ČT78a, DC75a, Dal76, DJ77, Dei77h, DMY78, Die75c, Die75e, Die75f, Die75g, DW79a, DW79c, Dwy78b, Ede75, Elw78, EL77f, Fel77b, Fel77c, Fel77d, Fey78b, Fd79, Gar76b, Gie76, Gin79, Hei79, HJ77, Hof77, HK79b, Ist78, Jai76, Joh77, Kai76, Kré76a, Kue76b, Kue76a, Kue79, Kuo75b, Kuo75c, Kuo76, Kuo78, Lac77, Lak79a, Man78a, Mic78g, Nas79a, Ørn77, PT76, Peł76, Pis76, Rap76, Rob77a, Roc77, Ros76a, Sat76a, Sat79a, Sch77b, Sch78e, Sco79, Sjö76a, SS77c, Ste77a, Szu78, Tay77b, TW78b, TDHJ76, Var76, Wer75]. Banach [Woy75, dA79, de 78]. Banach-space [Sch78e]. Banachmoduln [Gro79b]. band [BG78c]. Bang [FJ75e, KK78b]. Baouendi [Ali75]. Barbara [Mil78b]. Barrelled [AEK78c, AEK78d, AEK78e, HK78a, HK78c]. Barrelledness [HK78b]. barrier [Chu77b, Yam76]. Barry [Ser79b]. bars [Chi76c]. base [Erm76, Gér79b, Nov75, Pou79, RR78, Clo79]. based [GK78a]. Bases [Mar75, Pro78, Vie78a, Vie78b, EN79, Nat77b, Peł75, Sta77c, Wag79, dG75i]. Basic [Bra79c, Che76d, DR75a, Fei78a, Fis76b, Geo79b, Bre77e, Cla79f, GK78g, Gra76d, Rat78e, SW78f, Wil77b, Jur78i]. basiques [Vie78d]. basis [Jam78s, dG75a, dG75e, dG75f]. Battelle [Hil76]. Baumert [LW75]. Baxter [Tho77]. Bazilevic [Fek79]. be [Her76, KRZ77b, MJ75, Mor79c]. Beach [BH78]. Bearbeitung [JLR76]. Becker [Lem77]. Bedingung [RR75e]. Begriffe [Mét77d]. Behandlung [Häm76]. Behavior [Bru78a, Blo79b, Dur76b, GH77, Gre75b, Ike79, KY79a, Lei79, Lev79a, Sai79a, SM79b, Wei78a]. behaviour [Ghi79b, Lio76]. beim [Dic75]. Belgium [Wuy79b]. Beltrami [Kru79, SNWC77j]. Bemerkungen [Jur78s, Mic76c, Sca75]. Bendixson [Cam75, Lie79]. Benveniste [Mey77c]. Bergman [GW79a, Sib76b]. Berkeley [BR77b]. Berlin [HP79]. Bernoulli [DGS76b, Led76, SW78c, Vas78]. Bernoullicity [Geo76]. Bernstein [Bon76, Haz77g, Mal75a, Sch77c, SS76d, Jan79b]. Berquier [Mic76b]. Besancon [BR78a]. Beschraenkter [RR75c]. beschränkter [RR75d]. Besov [BTW75a, BTW75b]. Bessel [Gro78e, Gro78f]. Best [Bar78e, JS76]. Bestapproximation [Hei76a]. besten [DG76a]. Betti [Bar78f, RR78, Roo79]. between [Ada78, And77d, Aro76, Che76a, Cho78a, Čud77, DO77, Fri75, Han75b, Hei79, Kud77, Mar76b, Mic75, New76a, RR78, Roc77, Roo79, Sar76a, Sto75a, Ver79b, tM76]. Bewegliche [Gro76]. Beweis [Eif76]. Beweisen [Fel75]. Bex [BR76d, Hau78a, dlH79c]. Bezeichnungen [Gro79a]. Bezeichnungssystems [Pfe75]. Beziehungen [GMK75b]. Bézout [Bra79e, Bra79q]. bezüglich [Mét77b]. BG [Dup78j]. Bi [Kna79b, Cha79a, Raï78, Rub75]. Bi-invariant [Rub75, Cha79a]. bi-invariantes [Rub75]. bi-invariants [Raï78]. Bi-stable [Kna79b]. Bias [MBY79]. Bias- [MBY79]. bibliography [Wuy79a]. bicharacteristics [Dui75]. bidimensionnel [Mét78b]. Bidualräume [Gro79b, Gro79d]. Bidualräumen [Gro79c]. Biennial [GM79d, Wat78b]. Bierutowice [MSZ76]. bifunctors [Mic78d]. bifurcating [Jos76]. Bifurcation [Col76b, Sat79c, Sat79h, Sij79, Ale79, CY78, GM77g, Ioo76, Kel79, Kik79, PP79, Raţ79, Sat79f, Sat79g]. Bifurcations [HM78b, Raţ77, Wan79]. Big [Hoc79a]. bigebras [Yan77r]. Biharmonic [SNWC77b, SNWC77a, Ede76, GP77, SNWC77c, SNWC77d, SNWC77e, SNWC77f, SNWC77i, SNWC77n, Smy78a]. biharmoniques [Smy78a]. biholomorphes [Sib76b]. biholomorphic [GW79g]. biinfinite [dB76]. Bilateral [AG75]. bimesures [Hor77]. binary [ST77b, SM79a]. binomial [Sve77]. Biology [DKT78, Had76, Had79]. bipartite [Qui79]. bipermutative [May77b]. Biplanes [Hug78]. biprefixes [Per77]. Birch [Lan78]. birds [Zee76d]. Biregular [IŠ79]. Birkhoff [Lut79]. Birthday [DM75c]. Biserial [Ful79]. bishops [Rob76d]. bisimplicial [BF78]. Bismut [Mok78c]. Bissections [Vie78c]. bit [Cos75]. bitopology [Brü79a]. bits [Sto79]. bivariate [SS77a]. Bivariational [BR76a]. Blaschke [McD78]. Blattner [Voi77m, Voi77a]. Blending [BC77a, DP77]. blocks [CS75, Mic75]. BMO [RR75g, DMY78, Mey75b]. BMO-Funktionen [RR75g]. Bochner [BDE^{+}77c, Man78a]. bodies [Blo79c, Cha75a]. body [BK76g]. Bogomolov [Des79a]. Bohr [Gla76d, Gla76e]. Bois [Del77b, Ill77b, Del77b]. Boltzmann [Ell75]. Bonn [BR77c, BRP78, PW79, SS76a, SZ77b, SZ77c]. Boolean [BV77, Cic77, Fla77, Mah77, MR77b]. bord [Ali75, And75, Dar77, Lum76a, Rap76]. Bordered [Pea76]. Bordism [Con79a, Ecc78, Hau76]. Borel [Bon75, KS76, Ram79a]. boreliennes [Mok76]. boréliens [Gri77]. Borell [Ess75c]. Borho [Bor78c]. borne [Del77j, Des79a, Jac77a]. Bornées [CL76a, Bam79]. Bornological [AEK78f]. bornologie [Coe77]. Bornologiques [Mat77c, Mat77g]. boson [AHK76c]. Boston [Par75a]. Bott [Gar78]. bound [Bom76, FR79h, FR79g, Mah76f, Mar79c, Mas75e, Poh75]. Boundaries [Joh79h, Goo78]. Boundary [BK77e, BGM77, Cho76, GM77c, Ham75e, Koß78, Neu79, Sil76b, Sko76, Wal78a, dR79, Ama76a, BM79a, BL75, BF76b, Boh78, Bri76, Col78, Cor79d, Cox78, FR76, Föl75, Fuč79, Fuj79, GM77m, Gla76c, Gri76a, Gri76b, Gus75, Hes79, Hla79, Ike78, Ike79, KV79a, Ław79, Leh79, Lio75b, LP79a, McC78b, Mil76, Moe79, Nør78d, Ole78, Per78a, Red79, Rei76a, Sch78b, Sch75e, Sch75q, Sij79, Sil76a, Stu76a, Tam78d, Ver79b, VGK79]. Boundary-value [GM77c, Koß78, Gus75]. Bounded [Fai79a, Fer76b, GW79b, SNWC77c, SNWC77d, Tam78c, Won76d, BK77h, BF76a, Bea79a, BL75, CM77, Die75e, Gar77b, Hen79a, Lam75, Sch75q, Sha77a, Tam78d, War79b, Wid78, Wie79, Won76b]. bounded-measurable [BK77h]. Boundedly [AEK78g]. Boundedness [Szu78, Tur78, War76c]. Bounds [Dup79, Gil76, BR76a, BK76e, BC77a, FR79i, FR79j, SS77a]. Bourbaki [Ano76, Ano77, Ano79, BDV^{+}78]. Bp [Dav79, Gia78, Gro78o, Gro78d, Gro78g, Gro78h, Gro78p, Gro78r, Kan78a]. Bp-operations [Dav79]. bracelet [Zee76f]. Brain [Zee76b]. branch [Chi78b, Mac79]. Branched [Gér75b, HPS77b, Kau78]. branching [Ang79, BLP79, Cap79, Jam78b, Kaw76, Saw75]. Brauer [Cha76a, Chi76b, DeM76b, KOS76a, Orz76, Par76, Rum79, Sni79, Spr75, Zel76b, Zel76a, vO75a]. Brelot [Kor76]. Brick [HR78]. brief [Far76e, Ros75a]. Brillouin [Chu76]. British [HPS78]. Brown [Boa79, Lan79a]. Brownian [Cau75, Cau76e, Cau76f, Cie75a, Goo78, Sil76c, Wil76a]. Brownien [Cau76e, Cau76f, Cau76g]. Bruxelles [GDMW76, HL75]. Bucharest [CCJS79]. buckling [Zee76c]. Bucur [Ill77b]. Budan [dBS76]. buds [Laz75c]. Bundle [Sen78, AdS75b, AdS77, Bla76h, Dek79, Dun78, May77h]. Bundles [Fel77c, Hof77, HK79b, KT75g, sC77, Dav78f, Dup78a, Dup78e, Dup78k, Dup79, Fel77b, Fel77d, Fre75, Gli78, Gre77b, Hir75a, Joh79i, KT75a, KT75b, KT75e, KT75f, KT75i, Liu78b, Mic78a, Ree78b, Tra78b, Uen75a, tD79h]. Burger [Bar78a]. Burnside [tD79b, tD79c, tD79f]. Burnside-ring [tD79b]. Butzer [Sjö76a].

C [Bor78c, Dok78, Ill77b]. C. [Dok78, Har79a, Sib77]. C.N.R.S [Foa77a]. cône [Gol78a]. Cônes [FdlP76, Han76b, de 78]. Cabal [KM78b]. cacti [McA75]. Calabi [Bou79b]. Calcul [BT77, Gil78d, Jac79b, Las77, Lau75c, And77a, Bad76, Bre79f, FP78l, Mey79, Pau78, Roc79, Rum79, Tem76a]. Calculability [Han75a]. Calculation [Far76l, Mol77c, GM79c, Mül78a]. Calculations [FP78a, FP78b, Pow78a]. Calculs [Ill77a, Kré76a]. Calculus [GDMW76, Ros75b, Bal76, BF76b, Fis76f, HC78, Kuo78, Lau75a, Maa75, Mil75a, Ros75a, Soc77, Sta77b, Ste78h, Str75, dS79, Ded77]. Calgary [GM79d]. California [BH78]. Calkin [dlH79a]. Caltech [KM78b]. can [MJ75]. Canada [HL79b, HS79b]. Canadian [GM79d, HS79b]. Canberra [BCN77, Cop78i, HS78, NR78b]. Cancellation [Bec76a, Sta79a, Smi76a, Sus79]. Cancellative [Gri79a, Gri79c]. Canonical [Geo79i, Mos77b, FM76a, Gib76, Hud78, Ład75]. canoniques [Dab79, Pes76b]. canonization [Wei76c, Wei76n]. Cantelli [Kri76]. Capacitance [Wid78]. capacities [Car79a]. capacity [BBDB^{+}78]. Car [Kad78]. caractère [Jol78]. Caractères [Rog75a, Bes75b, Car77b, Rod75, Rum79]. Caractérisation [Ali75, GKR77a, Jol78, Bin75a, Par78, Aup79a, Aup79b, Aup79c, Sab76]. caractéristique [And79b, Szp76]. caractéristiques [Gue75, HL75, Mal79b, Wag78]. Carbondale [Nat79]. Cardinal [dBS76, Boo75, KM78a, Pét77, tM76]. Cardinals [Ště78d, BG76a, Dev75, McA78, Sre78]. cardinaux [Ste78e]. Carol [Kue75b]. carré [CW77b, Duf79b, Hir78a, Mey76f]. carrée [Loh75]. carriers [MBY79]. cartes [Fou79]. cartesian [Bou76a, Nel76]. cas [CW77b, Cho75b, Cho78b, Fed78, Ste76b, Yam78a]. cascade [Fuj76]. cascades [Abr76b]. case [Ale78, Bra79c, Fed78, Gal76d, Ish76b, Ish76i, Mos79, Roc75b, Tam78j, Wei76f]. cases [DR75h, Pet76i, Tru76]. Catalan [Fin76]. catastrophe [Pos76, Tho76a, Was76, Zee76f]. Catastrophes [Hil76, Zee76a]. Categorial [Trn77]. Categorical [Dyc76, MR77h, Che76b, FK79e, MR77f, BH76b, HP79]. Categoricity [Fri78b, Low76]. Categories [Bar79c, Bro75d, Ekl75, FGR75i, JPR^{+}78, MR77h, Mic78g, Pet76b, Ver77, Aar79, Bou76a, BM79b, Cig76, FP78k, HM79a, Joh75b, Joh79r, Laz75e, Laz75g, MR77a, MR77j, MR77l, Mar76c, May77b, Nel76, Rob76e, Rob77d, Vol75a, Vol75b, Woo78a, Car79e, Gro77f]. Category [Hof76b, Mah77, Bar79d, DS76c, DS78d, DS78h, Fla76, Hof77, Par76, Pop77e, Pop77l, War76d]. Catenary [Rat78c, Rat78a, Rat78k]. Católica [Sch78g]. Cauchy [Vek76, BG78a, Hof76a, Hör75, KS78b, Lum76a, Lum76b, Wag78]. Cauchy-completions [Hof76a]. caustic [Dui78]. Cayley [Bab78, Gro78a, Wat78a]. CE [Han75b, Sch78c]. Cebysev [Ørn77, Zie79c]. Cech [Bra79p, EH76c, Tay77b]. celestial [Sze75]. Cell [DS78b, Can75a, Kom78]. Cell-like [DS78b, Can75a]. celle [RR78]. cells [BBDB^{+}78, Dav75a, Gla75a, Wei76i, Wei76o, Wei76q, Wei76m]. Cellular [Wei76d, Gla75a, Wei76a, Wei76f, Wei76p, Wei76r]. census [CCMS75]. Center [Hil76, HPS77c]. Central [Jai76, Man76b, Smi76a, CR79, Gar76b, Gin76, Hei78, Hei79, Hel78b, Mar76a, MW78, Ryc78, War76b, War76k, War76m, dA79]. central-limit [Hei79]. central-limite [CR79, Hei78]. Centrale [BGM77]. Centrality [Smi76b]. Centralizer [Beh79d, Beh79c, Gro79f]. Centralizer-norming [Beh79d]. Centralizers [Yan77d, Pal78b]. centre [Tua78]. century [Lam76]. cercle [Del77e, Ros77]. Certain [Jat76, Tol75e, Acc79, AR79, AS79, Bab76, BP76, Bör79b, Che78a, CSW76, Dav75a, DV75c, FP78e, Gal76b, Gat77, Gér77, Gui77b, Hel78b, Her79b, Kre75a, Meh78a, Mis78, Mul78b, PR77c, Sri78]. Certaines [GKR77a, And77i, CG79a, Dam78, DMY78, Gat77, Kar79, Kil78, Kôn76, Sec75, Yor78b]. Certains [DC78, KV77, MT77c, Tho79a, Bas75, Bor75a, Bre79f, CO79b, CB75, Gér78, GS79b, Gos78, Gui79a, Kre75a, Lép78b, Mat75, Raï75, Rap76, Sjö78, Ste75b, Suz78a]. Chain [Cun75, Eas78, Rat78c, CF79, Gus76a, Hal79h, Loo75a, Rat78a, Rat78b, Rat78j, Rat78k]. chaines [Geo78b, Let76, Nic77, WK79]. Chains [Rat78c, Geo78b, HW75b, Kle75, Par75b, iS78]. chaleur [dV75a]. Champ [GR79c, Hir78a, Mey76f, Sam76, Sec75]. champs [Her75a, Kré78b, Pes76b]. Chandra [Car75, Sri79f]. change [Clo79, Eis77a]. Changement [Gér79b, EW77]. Changements [Jac79c, Jac79d, Jac79e, EM77, Str78]. Changing [Buk76, BDNY77, CW77a]. chaos [CY78]. Chaotic [KY79a]. Chapitre [CH78b, CH78a]. Chapter [Kle77a, Kle77b, Kle77c, Kle77d, Kle77e, Kle77f, Kle77g, Uen75d, CH78a, CH78b]. Character [Hir79a, Sch76m, Jam78c, Sch77e]. Characteristic [Dup78f, Dup78b, Joh79b, KT75a, KT75b, KT75g, May78b, Qui76b, AF78, Bot78, Dup79, Fer79, Fes78, Heg79, Joh79k, Joh79o, Joh79n, Jon78b, Jos77, KT75e, KT75i, Kat77, KOS76a, Laz75f, Sha76b]. characteristics [Mal79b]. Characterization [Get75a, Gra76c, Joh79p, Man78a, BD79, Bro78e, BH76c, McA78]. Characterizations [CW79, DS78c, GK78a, GK78c, GK78d, GK78b, GK78e, FJ75p, Won76b]. characterize [Oma75]. Characterizing [BHP78a, Bro77, Bru78l]. Characters [Ker75c, Sri79b, How78, Rog75a, Sax77, Sch75b]. Charaktere [Kle77j]. Charaktergruppen [Bin76]. Charakterisierung [DG76a]. Charakteristik [KPP^{+}78g]. Charges [Lap79]. Chase [Gar76a]. Chebychev [CC79]. Chebyshev [Gau76, Het76]. chemical [Gud76]. chemically [AG76]. Chen [Des79b]. Cherenack [Uen75f]. Chern [Bou79b, Dup78c, Dup78d, Dup78j, Jou77a, Kos75b, Ree78b]. Chevalley [Gér75c, Hum76c, Mul79, Sri79h]. child [Hig75]. choice [Fla75, SB77]. chomologie [Del79]. choux [HL75]. Chow [Lie78]. Christer [Ess75c]. Chromatic [Bab78, BC78b, CW78, Gup78]. chromial [Tut75]. Chu [Bar79c]. Church [Maa75, Sta75]. cinquième [Pau78, RR79a]. Circle [sC77, Fre75, Gro78n, KRZ77i]. circuits [Lit75, Lit77b, PR77b]. circulants [Ead77]. circulation [dR79]. City [GR75b]. clamped [SNWC77g]. Class [Sta77d, Bak78a, BP79a, Bit76, BD78, Bru78d, But76, Cor79i, Dah76, Fla75, Gei77, GW79b, Gus76b, Guz77, Han78a, Het79, Hum79, Iva79b, Joh79l, KTR79, Kov78, LS79b, Mac78, MN79, MZ78, Mas79, MS79c, Rei75, Rei78a, Rog77c, Sko76, Sun79, Tam78e, Var76, Wei78b]. Classe [Gil78e, BM77b, BLC78, CDI78, GM77o, GD77b, Kor76, LR78a, Mal78b, Mar79a, Mat78b]. Classes [Bro75b, Cou79, Dup78f, GKR77a, HW75d, KT75g, Web77, Ave78, Beh79k, Bot78, Bré75e, Bru78p, Bru78q, BC77b, Cor75, Dav78c, Dud79, Dup78b, Ecc78, Ewi79, FF78, Fes78, FSW75, Gat77, GT75, Gro77f, Gue75, HL75, Har76a, Heg79, HK78i, Huš79, Ist78, Jon78b, Jou77a, Jur78a, KT75a, KT75b, KT75e, KT75i, Kec78c, Kol78, Kos75b, Kra76a, Kra78a, Mah79, MM75b, May78b, Qui76b, Ree78b, RB75, SNWC77m, Sch75a, Sch75q, Sch78f, Sch79i, Wag78, Wei76r]. Classical [Buh78a, FP78f, FJ75l, BDE^{+}77d, Car78a, Con77, Die75a, Dup78b, FP76, FP78g, FP78k, Fri76a, Gar77a, GK76a, GR79h, Gol78b, Gor78a, Gro78r, Lam78a, Pra75]. Classically [Fri78c]. classifiant [Cat75]. Classification [ACH77f, Bau77e, Car78a, Dav76, Dem75, KZ77, Man75c, Pop77g, SNWC77e, Sri79c, Sri79d, Uen75e, Uen75f, But75, HK79a, Joh75b, KRZ77h, Kaw79a, Lar79, Mar76d, Pop77h, Pop78, Sal78, Sun75, Van78b, VVT77, Zee76a]. classifications [Fef75c, LM79]. classify [Kea78]. Classifying [MR77c, Wor79, Ada78, Dup78k, KS79a, Lad77, Sna78]. classique [Gra77]. classiques [Jou77a, Mey76d]. Clean [Gui75]. Cleichungen [Vek76]. Clenshaw [Rei77c]. Clifford [Mic78a]. clique [BC78b]. clos [Que77]. close [Sch75q]. close-to-convex [Sch75q]. Closed [Fit75, Mat77a, AEK78c, AEK78h, Bou76a, Dei77j, Die75e, Glu75, Gre79, Kni78, Kre75a, Nel76, Ste78f]. Closing [Rob78]. CLT [ČT78a, Hah76]. club [Per79b]. Cluster [Sch77a, Fuj79]. cnoidal [Wah76]. CNPg [PdC77]. Co [LV75, Lie77, NT79a, NT79f, Wol78]. co-action [NT79a]. co-actions [NT79f]. co-adjoint [Wol78]. Co-topologie [LV75]. co-valuations [Lie77]. Coalgebras [Yan77e]. coarse [Oor77]. Coates [Lan78, Mas75a]. Cobordism [BC78a, Dol78, Mor78b, RT78, Sna78]. cobordisms [Wag77]. Cobordismus [Vog78]. Cocycles [Roo78a]. Codes [ST77b, Cum76, Per77, Rod79]. Codimension [Ser75d, Ser75e, Dav75b, Ecc78, Kol77, Lam75, Sch75u, Zee76a]. coding [AM78, Par79a, Sol78a]. coefficient [BDNY77, Pfl76, Tam78a, Tam78b]. Coefficients [And77g, And77f, Bro78c, Dur76b, Har76a, KRZ77b, Kwa76, Leg75, Lio76, Mag79b, Sve77, SD77, Zag77b, Cow79]. Coequalizers [Ros78a]. cofibrations [Bau77b, Bau77g, Bau77h]. cofinality [Buk76]. cofini [Cou78]. coforcing [Sac75]. Cohen [Hoc79a, Yam79]. Cohenschemata [Kra75b]. Coherence [FGR75a, Sta77a]. Coherent [Fis76c, Gru75c, Ste76a, Tra76, Tra78a, Ste76b]. cohérents [Thi77, Ste76a]. Cohomologie [Bar78c, Bor75a, DB77, Del77b, Ill76, Jou77b, Jou77a, Meb78, Szp76, BE78, Bou76b, Cat75, Del77g, GD77b, Gro77d, Gui79a, Hae75]. Cohomologies [HL75]. cohomologique [Jou77a]. Cohomology [AT75b, Bro76b, Ell78, Hae78a, KT75c, Mic78b, Sch78g, AL77, CT78b, Del77b, Dup78g, Dus79, GH78a, Har76a, Hir78b, Ill76, Jou77b, Kre76b, Lau79b, Lau77, Liu78b, Mos79, Sha76b, Śni77, Sri79a, Tay77b, Wag76, tD79d]. Coincidence [GM77d, GM77f, GM77e, GM77g]. collaboration [Ill77b, Mol77-28, Uen75f]. Collection [Kue75b, LMW75, Wal78b]. Collège [Ser75a, Ser75a]. Collocation [FU78]. Colloque [CDV75, CV77a, CV79, Cha75c, HH78b, Rau75e]. Colloquium [DM75c, JMR75, LLS79, MOP75, Par75a, SS76a]. colocal [Yan77r]. color [Ber78b]. coloriage [Fou79]. colour [MS76a]. coloured [Sax77]. colouring [Hei76b]. Columbia [HPS78]. columns [Sri78]. combination [Iva79a, Fin75]. combinations [EN76]. Combinatoire [Foa77a]. combinatoires [Alo79]. Combinatorial [CN75a, HS78, HW79b, Pel77, Erd77, Hal78b, Hin79, Mac78, McA78, Rob75a, Rob76a, Rog77a, Sco77, Sha79a, Sta77b, Wil77b, Alo79, CW76a, HS78, HW79b, Lit77a, SW75d]. Combinatorics [Kle77i, Mac79, Ent78, Ker75a, Kle77h, Nat77b]. combinatory [Cur75b]. combustion [AW75]. Comma [Cum76]. Comma-free [Cum76]. comme [CM75, Rot78, Sak77]. comment [Mol77h]. Commented [Wuy79a]. Comments [Pra75, Rat78d, FM76a, SM79b]. commutant [NT79g]. Commutants [Cow78]. commutateurs [Lép78b]. Commutation [Ral79b, Cha78c, Ład75, Pes76b, Van78a]. Commutative [Bra79d, CDV75, CV77a, DR75g, Hoc79b, Laz75a, Sło78, vO75c, Con79f, Dal76, Day76, Eff78, Fai79b, Gil75a, Gil75b, Lan79a, Mac75b, Ros78b, SW75a, Sha79b, Yam79, CDV75, CV77a, CV79]. commutatives [Aup79a]. commutativity [Bre75c, Wei78b]. Commutator [Wal78b]. Commutators [Yan77f]. commuting [CP77a, Lub78]. compacité [Sch76e]. Compact [Aro76, BJM78a, Bre77b, Bre77d, DGS76d, Hey78, Mil75b, Nam79e, Poé76b, Raj77, Sch75h, Sch75u, Sch75v, Uen75f, AEK78i, BK77i, BK77g, Ber76c, BJM78e, Bra79k, Bri76, Bro78e, CV77j, CM76, DGS76k, Duf75b, DR75i, Dup78d, Edw75a, Eps75, ER77, Far76g, FJ75o, FJK79, Hej76e, HPS77j, Hol78a, Huš79, KR79, Kre75a, Lie79, LR79, Lub79a, MP79b, Nam79c, Nam79d, Nam79f, Pop77h, Sch75n, SV75a, Tak79, tD79b]. compact-continuous [BK77i]. compacte [Le 77, dV75a]. Compactes [HH78b, Bou79b, GS79a, Kil78, Sub75]. Compactification [Pop77c, Bra79p, Get75i, Gla76d, Gla76e, Ike78]. Compactifications [Kor79, BJM78f, Hal79e]. Compactly [Frö76, Die75g, Het79]. Compactness [CV77b, Dei77b, Hal79c, Lie78, BK77k, Lou78, Mus76]. compactological [CM76]. compacts [Bar75, Bou79a, CL76b, Kre75a, Le 75, Loe75, Mok79]. companions [HW75n]. Comparaison [Bre77h]. Comparing [Dob75]. Comparison [Bom77, Get75b, Get75c, RF79, Cor79g, GW79e, Lak79a]. Complément [Ber77b, Mey75b]. complémentaires [CO79b]. complementarity [Cot75, Cot79, Man76d]. complementary [Eav79, MS78e]. Complemented [Dub79b]. Complements [Fey79a, Jac79f, Bou78c]. Complete [Bro79a, Che76e, HRW78, Mic78c, BG78b, Bof75, Bra79n, Che76f, Che76g, Che76h, Fer79, FSW75, HWH78, HW75a, HW75d, HW75e, HW75f, HW75g, Hog75, Kaw79a, Lie77, Lub79a, Qui79, Sch75a, War76a]. completely [DR75d]. Complètement [BWK77e, Dix75, Fey78a, Que77, Ren77, Sch76h, Sec75]. Completeness [AEK78h, BH79, MR77d, Vol75a, Kre75c, LEV75, MR77e, Smi76f, Win75]. completion [AEK78q, But78]. Completions [Mor78b, War76a, Bör79b, Her79a, HW75n, Hof76a, War76l]. Complets [BWK77d]. Complex [CCJS79, Fis76d, Gre77b, KZ76, LLS79, Moi77b, RT78, Uen75f, Bar78e, Coh79a, Dav79, Duf75b, Dup78i, Fis76a, Frö79a, Jen77e, Jen77d, Mal75b, Mas75c, Mas75h, Mas75g, MS75, Mor78b, Mor77c, Nam79c, Nam79f, Pal78a, Pop77h, Sjö75, Suz79, Uen75b, Uen75e, Uen75h, BS77, LLS79]. complex-valued [Bar78e, MS75]. Complexe [Băn78, GR79c, Bar75, Cat75, Fou75, Hen75b, Kan75a, KS78b, Lel75a, Lel78b, LBL78, Sec75]. Complexes [Gro77b, Nor75, Nor78c, And77b, Bar75, Bau77d, Blo77, Del79, Duf75a, Duk78, Fey78a, FS78b, Hen75c, Ive78, Jou79f, Kan75b, Kil78, Lel78a, Lel78b, Mis78, Roc75a, Sha79b, Sib77, Sjö78, Suz78a, Gro77f]. Complexity [FR79b, Sim75a]. complicated [FR79k]. components [Cir78, GdM77]. Comportement [Dar77, Duv77, Ren77, Lép78c, Lum76a]. composed [Hei77]. composés [Fer76a]. composite [Eck77]. Composition [Jam78d, Nor78b, Cim77, Han78c, Mok78a]. Compressible [LD77, Ber76e]. Computable [Mic78d, Mic78e]. Computation [Fen75, Kös76b, Mol77e, NV79e, Ruh77, DGS76a, Mol77d, Mol77l, Saf77]. Computational [FR79b, HBT75, Ode75, BR77a, Bab79, GM76c, Roa75, Ode75]. Computations [CT78b, Hof79e, Hof79f, Mol77f, LT76d, Mol77c, Sto78a]. Computer [CS75, CY78, Lan77]. Computing [Wam77, Wuy79a, vdLT79, GLL79]. Concentrated [Hel79a]. concentration [Bou79a]. concept [Bra76c, Bre79c]. conception [Ste78g]. Concepts [BWK77h, Gal76i, Geo79b, HK79b, KLM75, Zie79e]. Conceptual [MR77e]. Concernant [BE78, Ste75c]. concerning [Byc78, ČT78a, Csá77, Dav75b, Dav78h, Edw75a, HL77, Mol77h, Ste78j, TDHJ76]. Concluding [Fei78b]. Conclusions [Hal77]. Concordance [Dav78c, Gol78b, Sto78a]. concordances [BLR75a, Lod79]. Concrete [Cla79a, Dub79i, Ols78a]. condensing [Dei79]. condition [Bla76d, Gre75c, Hal79h, Led76, MW78, Mat77a, Stu76b, Tam78n, Wei75c, Wil77c, ZKB79]. conditional [Bor78a, Che78b, Dub76b, Led79, Mar79d]. conditional-mean [Mar79d]. conditioned [Kaz76]. conditionnelle [Par78]. Conditions [AEK78u, Bak79a, Dei77c, Jac79g, Joh79q, KRZ77b, Les78, Mis78, Sch76e, AB76, BL75, BLP77, Bre75c, Bro78a, CF79, CK75, Cox78, Dei77b, Dei77g, Gri76a, Hah76, Hal79e, Hal78a, KRZ77i, Lév77, Loo75a, Mil78a, NV79b, Nic77, Par78, Sch75r, Sha79b, Sil76a, Tep76, War76g, WK79, Yam76, dP76a, Jac79g]. cone [Ber79b]. cones [Ber76a, CL75a, Cos75]. Conference [AHW77, BCD77, Bec76b, Bec79a, BK76a, BH76b, BGM77, BS77, BOS75, BGS78, CW76a, CS76d, DG75j, ES76, Fen79, GM77n, GR75b, HP79, Hey79, Hil76, HS79b, HS78, HW79b, KK78a, Kel75c, Lit77a, MS77, MOP75, Nat77a, Nat79, Ode75, Pin75a, Ros75b, SZ77a, SZ77b, SZ77c, Ste76c, SW75d, Tem76c, Wat76, Wat78b, Wuy79b, Zel76b]. Configuration [Mac77b, CT78b, Pet75]. configurations [Ber78b, Bol78]. Conformal [DFF78, RR75f, Dra79, LF79, Sch75f]. Congress [GM79d]. congruence [LS77]. congruences [DO77, Ran77, SD77]. Conical [Har75b]. coniques [Har75b]. Conjecture [Erm78, Lam78f, Rat78f, And77j, Ber78b, Dem76a, Ess75h, Fou79, Gel77, Lam78a, Lam78d, Lam78g, Lan78, Lit75, MP78b, Rat78h, Rat78a, Rat78b, Rat78d, Rat78g, Rat78j, Rat78o, Rat78p, Rat78q, SNWC77g, Sch75m, Ste78a]. Conjectures [Gol79, Rat78c, SW78d, Lub79a, Rat78k]. Conjugate [Fle76, Lan79b]. conjugation [BK77b]. Connected [Moi77b, Glu75, GdM77, Kle75, Mag79b, Moi77g, Wor79]. connectedness [Jon78a]. connecting [Bru78s]. Connection [Mol77g, Dek79, May77l, Sri78]. Connections [Dup78e]. connectors [SB77]. Connelly [Kui79]. Connexe [Cha77]. connexes [GS79a, PS79]. connexions [Mol75]. consequences [Dav78i, Ess75j, Laz75h]. conservation [Gat77]. Conservative [New76b]. Considerations [BF75b, Ste78j, Ne’77]. considérés [Sak77]. Consistency [Gre75a, Tak75b, Cun75, Luc75, Szc77a]. consisting [Won76a]. consists [Len76b]. Constant [AD79, Gud76, SW78m, Sza75]. constante [Mar75]. constants [Leg75]. Constrained [Tak76a, GM76a, Pow78a, Pow79]. constraints [Das75, Kno75]. construct [LW75]. Constructibility [Dev77c, Dev77b, Háj78, SA77, Tru78]. Constructible [SA77, Maa78]. constructing [Sim78c]. Construction [And77i, Gér75b, Gib76, Kis77a, Mor79b, Osi78, ST77a, Wei76l, Wei76f, Far76a, Fie78, Gli78, Han78c, Hau76, Her77c, Mah78, Pop77i, Sil76e, Sri79e, Wei76b, Wei76e, Wei76i, Yam78a]. Constructions [Rou78, Ead77, Frö79a, GW76b, Hof79j, Sch79a]. Constructive [GH76, Jer78, Kel79, DS76b, PP79, Sch76s, SZ77a]. Contact [Bla76b, Bla76c, Jen77c, Miu76a, dP76a, Bla76a, Bla76e, Bla76d, Bla76f, Blo77]. Contact-invariant [dP76a]. contain [Hua79]. containing [CM77, Ros78b, Ros78c]. containment [BC77b]. contemporary [BP76]. contenant [Cou79]. contenue [Mok78b]. context [MJ75]. context-free [MJ75]. contiguous [Neu76]. continûment [Nac78]. continu [Cho75b, Cho78b, Mos78b, Tua78]. Continuation [VR75, Ale79, BZ78, GM77h, Het79, Nie77, PP79]. Continue [Bre77g, Bru79b, Del79, Pra79]. continued [Get75h, Get75n, Gro78g, Gil78f]. Continues [BWK77n, Bre77l, Gil78f, Sch76i, Bad78, Buc78, Gui79b, Lum76a, Sch76d, Sch76g, Sch76h]. Continuité [Anc76, Jac79g, Aup79d, Aup79d]. Continuity [Bru78b, Hyl79, Jac79g, Aki78i, Bak76, Bak78a, Bak79a, Bom77, Cha78b, CC76, JS77, KLŠ76, Kun76, Lau76, Sha76a]. Continuous [Bin75d, Blo79f, Geo79c, Moh79, OW75, Pre76c, Wag76, AEK78q, AM79a, Bal76, BK77i, Ber79a, Bin75i, Bru78a, Can75b, DGS76i, Gus76a, Hal76b, HP77, Huš76, Kan79, LY79, Mis79b, MM79b, iN77, Ral79b, Sch76i, Sch78e, Sch75o, Sek76, Haz77e]. continuum [Guz77, Rog76a]. Contopoulos [SV79]. Contrôle [Bis78b, Bis78a, Geo78b, Jen78, Mos78b]. contraction [Bre79d]. contractions [Anc76, Ist78]. contractive [GM77e]. contributions [May77a]. Control [BOS75, Con79g, Cop78i, Kra78a, Åst78, Ban79c, Bro78a, Chr76, Chr78, Dav78a, DH79, Dun78, Geo78b, Gla75b, Gus75, Hal78a, Klo78, Kno75, Köh75, O’M78, Roc78, vH79b]. controllability [Fat75]. controls [FJ75e]. conuclear [Tho78]. convection [GH77]. convective [Sat76b]. Conventions [Bau77a]. Convergence [Att78, Bin75d, Bin75e, Bre77f, Bre77i, BG78d, Chi76c, DC78, DSS76, Gil78b, Jaj75, Mey77a, Sha77b, SW78e, Tay78a, Tay78j, Wah79, Won76d, Ale78, BDR78, BD79, Bin75g, Bin75h, Bou76a, Cai79, Cle75, Edg78, Fey79a, For76a, FK79c, GS76, GB79, Gra76c, Ham75c, Kis77a, Kue76a, Man79, Mil78a, MG79, Neu76, PT76, Sai78, iS78, Sbo79, SS76e, Ste77a, Stu76b, Szu78, Tay78e, TDHJ76, Won76b, dP79, Nit77]. Convergent [Mah76b, DJ77]. Convergentes [Bre77h, Gér78]. converse [dA79]. Convex [CV77c, CV77d, CV77e, Ber79b, Bor78a, Bre75b, Bri76, CV77j, Cha75a, Chi76a, Cho76, Col75, DK79, Die75e, Die75f, Dud79, Egg78, Gri76b, HK78i, Lap79, Mor76, Pet75, Rob77a, Sat76a, Sch75q, Tho76b, TH75, Won79f]. convexe [FP78l, MD78, Tem76a]. Convexes [Lau75c, Att78, Deh76, Kis77b, Lap79, Lel75a, Rap76, Sib76b, VVT77]. Convexite [AN78, Erm76, Ste75b, Nov75]. Convexity [AEK78u, Die75b, Suf77]. Convolution [Hey78, MT76b, Ber76b, Ber79a, Far75a, Haz77e, Hir79b, Mey76g, Sha77b]. Conze [Bor78c]. Coordinate [May77f]. Coordinate-free [May77f]. coordinates [Hab76]. Copenhagen [Bar77b, Løn79b]. coproduct [Dou78]. core [Rei76b]. coreflections [Huš76]. Cornalba [Le 77]. corona [Sch77a]. corps [And79b, ET79, Fer77, Gér75c, Gér75d, Gér79b, Mul75a, Mul79, Szp76, Tit78]. correct [Wei76m]. correcteurs [BT77]. Correction [Mey76a, Zag77a, Böh76b, FU78]. Corrections [MR77f]. corrector [Cla79g]. correspondance [Sch77g, Vie77]. correspondence [Gre77a, Ław79, Mic75, War76f]. Correspondences [Cho78a, Hej76b]. corresponding [KLŠ76]. couche [Bar77a, BT77]. Couette [Bow77]. couleurs [Fou79]. Countable [Dei77d, Eps79b, Tak78b]. Countably [HK78c, HK78d, HK78e]. counter [AEK78p, ČT78a]. counter-example [ČT78a]. countercurrent [LD77]. counterexample [Ess75d]. Counterexamples [BW76]. counterfeit [CT78c]. Counting [Rob76d, Rob77c]. couple [Gui79b]. coupling [Bar78f]. Courants [CH78d, Kan75b, Lel77b]. Courbes [Del75, Des79a, Lig77b, GP78b, MS76c, Ser79b]. cours [Ser75a]. Course [Bar77b]. courses [Ser75a]. Covariance [KV77, Bec78a]. cover [Bre75b, Gup78]. Covering [Cha75a, Oxl79, dG75b, dG75h]. coverings [Per78c, Top76]. Coxeter [HL77, Mat75, Mat77h, War77a]. CP [Emc78]. CP-flows [Emc78]. Cramér [Haz77g, Hol76]. Cramér-Rao [Hol76]. creep [Zie75]. crete [Cup77]. cribles [Des79b]. cristalline [Ill76]. Critère [Bou78c, MS79b]. Criteria [Cop78a, Hil77a, KLŠ76, KRZ77f]. criterion [GW79h, Mor79c, Vaa75]. critical [BP76, Bom76, Raz77]. critique [Dem75]. croisements [CH78e]. Croissance [Ave76, Hec78, Jen78, Kaj75, Lel78b]. croissantes [dSLM75d]. Cross [Far76b, Utr79, PE75]. Cross-validation [Utr79]. Crossed [NT79c, Fis79, MMS78, NT79d, NT79g]. crossings [vdE79]. crosswind [Wac78]. Cruces [AHW77]. crumpled [Bur75a]. crushed [Rau75b]. Crystalline [Ill76]. crystallographic [MS76a]. CS [Han75b]. cube [Cha75b]. cubes [Bur75a, JL77]. cubic [Hua79, Ish76e, Wor79]. Cudnovskiı [Wal77a]. cumulation [Fri78d]. cumulative [Got76]. Cunningham [BDE^{+}77e]. current [Blo79g, MS79a]. currents [CH78d, NV79e]. Curtis [Tan78]. Curvature [Dup78f, CdC77, FJ75r]. Curve [GR79b, GR79a]. curved [BB76]. Curves [Jen77a, Cho77c, DeM76b, Gol79, Jen77e, Jen77d, KK79, LL78, Maz77b, McK79, Pop77d, Pop77i, Rob76a, SW75a, SDB75]. Cusp [Lan76b, DO77, Zee76f]. CW [Bau77d]. CW-complexes [Bau77d]. cycle [GD77b, HR76]. Cycles [Bol78, Bar75, Gre78b, Kre75a, Las76a, Le 77]. Cyclic [Hua79, Rob76e, Ham75a, Her78a, Lev79b, Mic75, Rei78b, Sni79]. Cyclotomic [ACH77c]. cyclotomie [Wei75b]. cyclotomy [MSW76, Wei75b]. cylinder [BLS75, Gua78, LD77]. Cylindres [Hec77]. cylindrical [Chi76c]. cylindriques [Bad78, Che78c].

d [Kaj75, BDPR76]. D. [Bre75a, For79a, Gri77, Lem77]. d’A. [Cho75b]. d’Accélération [Bre77i]. Dade [Voi77s]. Dahlberg [Sjö76b]. d’Aitken [Cor79b]. Dakota [MMP78]. Dal [Fel77c]. d’Alembert [Wil75d, Wil75e, Wil75a, Wil75i]. d’Alexander [Web78d]. Daley [For79a]. d’Algèbre [Mal77, Mal78a, Mal79c]. d’algèbres [Bar78c, Spr77f]. d’Ambrose [dSLM75f]. d’Analyse [BR78a, CDV75, CV77a, CV79]. d’anneaux [Avr79, Cou79, Que78, BWK77b]. dans [Anc76, Aub79, Bad76, Bai76, Bai77b, BWK77g, Buc78, Car76b, Cha79b, CM75, CH78e, Dem77c, Dem77b, Dix75, Erm76, Fit78, GR79c, Gil78g, Gui76b, Hei78, Kaj75, KS78b, Kor76, Kui79, Lap79, Lel78a, Lel78b, Lem77, Les78, Lév77, Lio77b, Loh75, Lum78, Mat75, Mat77b, Mat77h, Mat77c, Mok78b, Nic77, Nie75b, Nov75, Pis76, Pro78, Raï75, Ren79b, Sec75, Sib76b, Sim78b, Sjö76a, Smy78a, Ste75c, Tak79, Wag78, Web78d, Yam78a, Yam78b, de 78]. d’Applications [BR76d, Dlo78b, Nac78]. d’Approximants [Gil78g]. d’aprè [Del75]. d’après [Ali75, Alo79, Bon75, Bon78b, Bor75a, Bou79b, Bre75a, Car78c, Car79e, Cho75a, Del77e, Dem75, Des79a, Des79b, Fer77, Gér79b, Gri77, Hir78a, Ker75f, Kui79, Lan78, Lat79, Le 75, Le 77, Lem77, Lod79, Mey77a, Poi77, Raï78, Ram79b, Ray78b, Ros77, Rou78, Ser79a, Ser75p, Ser77, Ser79b, Sim77, Sjö76a, Sjö76b, Spr75, Ste78e, Tei75, Tit78, dV75a]. Darboux [Bru78c, Bru78d]. d’Ariane [Del77h]. Darmstadt [MWW76]. d’arrêt [Kôn76]. Darstellung [Bau76]. Darstellungen [Voi77i, Voi77j, Voi77n, Voi77c, Voi77p]. d’Artin [Bou78c, Mar75]. d’Asada [Leg75]. Data [Sco75, Bar78e, Bow77, Hal78a, Par79c, Wah79]. d’attaque [Dar77]. d’auteurs [Ano79]. d’autres [Alo79]. Day [Nat77a]. d’Azema [ÉS79]. DC [Sol78b]. decay [Cos75]. December [AT78, AHW77, BGM77, GM77n, GLL79]. Décidabilité [Jac77a]. decidability [AOL79]. Decision [Slo76, FR79d, FR79e]. decisions [BR77a]. deck [Bon78a]. décomposabilité [Art75]. Decomposable [EL77b, Man78b]. Decompose [Bra79d]. Decomposition [AF78, BM77b, Hog75, MM79b, Reb79, Rob79b, Rod77, Sta77c, Bau77f, Bra79o, EL77c, Gli78, Iag75a, Jac75a, Jac78b, Kah78, PE75, Sar78, Sek76, YM76]. Decompositions [Arm75, CMN79, EL77f, Far76c, HWH78, Jac79i, Mém78, Arn77, Bri76, DGS76c, Dou78, EL77a, HR76, McC79, Smi76c, Sna78, Thi77]. décroissance [Rap76]. Deddens [Sta78a]. Dedekind [KL79a]. d’edge [BEGS75]. d’edge-of-the-wedge [BEGS75]. Dedicated [DM75c]. Deep [Wei76g]. Défaut [Bea78g]. defect [Böh76b, FU78, Mic75, Sto77a]. Deficiency [KRZ77d, RB75, EG75, KRZ77c, KRZ77h]. definably [Glo78]. définie [Jen78]. Defining [FM78, FR79k]. definite [DR75d, FJK79, Gór78b, Sle75]. Definition [Bea78a, SNWC77a, SW78f, Joh75a, Zie79d, Maz76]. Definitions [Gra76d, Jac79m, Rat78e, Cen76, Din75, Poh75, Ste79, Wei76s]. deflection [Pau78, RR79a]. deformable [Bre76b]. Deformation [Her79b, Hol78b, Lak79b, Nam79f, Pom75, YF79]. Deformations [Her78b, LL78, Tra76, Tra78a, Tra78b, Lie78, Pal78a, vM79]. Degeneracies [Fis76e, SNWC77i]. Degenerate [Bar79b]. degenerating [Roc77]. dégénère [Fit78]. dégénérée [Rub79]. dégénérés [BLC78]. Degree [Eps79c, GM77d, Bar79h, BC78b, Con75a, Con77, EL76, Fre78, GM77f, GM77e, GM77p, HP76b, Het79, Ish76b, Max78, Moi77g, dB76]. Degrees [Eps79d, Eps79g, Eps79h, Eps79k, Eps79l, Háj78, SA77, Tru78, vW78]. Dehn [Joh79c, Joh79d]. Delay [Had79, Alt79a, AS79, Ban79c, Fit76, Opp78]. delay-differential [Opp78]. delays [Hal79a]. d’Éléments [GR77]. Deligne [Mor77c, Ser75p, Ser77, Sri79e]. Delorme [Gui77a]. delta [Kis77b]. delta-convexes [Kis77b]. delta-plurisousharmoniques [Kis77b]. delta-sousharmoniques [Kis77b]. demi [How77, LR78a]. demi-droite [LR78a]. demi-groupes [How77]. Démonstration [ÉS79, Cor79a, Sch78a, Tho79c, Cor79a, Dem76a]. Démonstrations [Mey77b]. Denmark [Bar77b, DM79b]. Dense [Mac75a, Bro75c]. d’ensembles [Gap76, Mic76a]. Densité [Jou79b, Nac78, Par78, Tua78]. densities [BK77j, GvW76]. Density [Par79c, Bla75, Far76i, Haa79, Kue78, MG79, Ros79]. d’entiers [Bes75b]. d’entrelacement [Mul75a, Mul79, Lio77a]. départ [DB77]. dépendant [Mey79]. dependence [Bra76c, Wol75]. dependent [EW78, GR79j, Phi76, Wil75h]. Déplacements [GKR77d, Bal79]. Déploiement [Poé76a]. Déploiements [Mar76d]. Déployés [Gér75b]. d’Epstein [Fit78]. Depth [Ive78, Rat78f, Hag78, Wei76h]. d’équations [Con76b, Gat77, KS78b, Kré75b, LR78a, Tar78, Yam78a, dM75a]. d’Équivalence [Mol77s, Mol77z, Mol77-27, Mol77-28]. derivates [Bru78g]. derivations [Dal76, Joh79p, Sak78, Yan77n, Bon75, Del77j]. derivative [GR75a, MG78, Roc75d, RY77]. Derivatives [Bru78e, Bru78b, Bru78h, Bru78l, CN76, HW79a, MG79, Sch77f, tM76]. derived [You75]. Derivées [Tho79a, Ver77, Lai78, Bon78c, Krá76b, Tar78, Bon78b, Kré75b, Las78]. d’erreur [Bar77a, Bai77b]. descended [Rat78g]. description [Bin75g, Ded78, Law76, Ver79b]. descriptive [Del77j, vW78]. design [Loo75f, PR77b]. Designs [MSW76, Bre79a, EWW77, Ead77, GW76b, Han78a, Mor79b, MS78e, Pre76a, Rob76c, Rob77b, Str77]. desingularisation [Tog76]. d’Espace [Jac79e]. d’espaces [Ali75, Fd79, Hen75c, HT75b, Lac78, Sch77h, Sub75, Mat78b]. dessus [MM75a]. Desuspension [Car79c]. d’été [And77a, BKK76, DCHRH78, FCG75, HJLN77]. Detecting [BR78b]. Detection [FP78c, FP78d, Gua78]. determinantal [Jam78e, KL79b]. determinants [Cla79b, Bre79f]. Determinateness [Kle77i, Kle77h]. Determination [New77, Nør78d, Gri77]. determined [CSW76]. determining [Bru78n, Tat75]. deterministic [Sza75]. d’Euler [Gro77d]. d’Euler-Poincaré [Gro77d]. Deux [Del77i, Lau75c, And79b, Hen75b]. development [Bro79b, Cod76, Fro77]. developments [Jac78a, Keo77, Kra78b, Pop78, Roa75, Sak78, Sch79e]. deviating [Red79]. deviation [Ros79]. deviations [DV75c]. d’évolution [BCH78, Lum78, LP79b, Paq79a]. d’existence [Gui76a]. d’Experiences [Tik79]. DF [AEK78i]. DFN [Bol75]. DG [KT75c]. Dgln [Jur78e]. d’Heisenberg [Gér75d, Roy75]. d’hélices [dSLM75g]. d’holomorphie [Dlo77]. d’homologie [Gra76b, Lat79]. d’hypersurfaces [Dlo77]. Diagonalisation [MM75a]. Diagrams [Jam78f, Col76b]. Dichotomies [Cop78b, Cop78e, Cop78c, Cop78d, Cop78f, Cop78h]. dichotomy [Cop78a]. Diers [Joh79r]. diffeomorphismes [Ban75, Her75b, Del77e, Ros77]. Diffeomorphisms [Bow75b, AdS75a, Bow75a, Bow75c, CN77, Fra78b, Mañ75b, Man75c, Pal78b, Rob75b, Sch79i]. Difference [BTW75b, Red79, Bra79a, BTW75f, BW75, CS75, KY79a, Mil78a, Per78a]. differences [Bau77g, Rob76b, tM76]. différenciels [Dem77c]. differencing [Wac78]. Different [Col75]. differentiability [Die75b]. Differentiable [Con79b, Con79e, Lem79, Aro76, Bur76, Hal78a, Het79]. Différentiables [BR76d, BR76c, Bur76, Car76b, Las76b, Mar76d, Nac78]. Differential [AT78, Bib79c, Blo79f, BGS78, Cha78a, Cha75c, Cod79, Cor79e, Dei77h, Dup78g, Elw78, Eve75a, ES76, Fis76f, GM77d, Gar78, Gol75, Gro78f, Har75a, KRZ77d, Kuo78, MWW76, PW79, Pin75a, Ros78d, Soc77, AW76, Adl79a, AB75, Alt79a, Ang79, AS79, Axe79, BM79a, BZ78, Beb79, Ben76c, Bih79, Bit76, Bre79d, But76, CV77i, Cha79a, Cod75, Col75, CS78, Cor79f, Dei77d, Dei77e, Düm79, Džu76, Eas75, Egu79, Ess75a, Est76, Eva76, EG75, Far76d, Far76h, FJ75d, Fri78a, GM77c, GM77k, Gam79, GW76a, Gla75b, GS77, Gra76a, Hae78b, Hal76b, Har76a, HB79, IK79, Jur78k, KK75, Kat75, KRZ77c, KRZ77h, Kaw75, Kos75a, Kre75a, KS79b, Kur75, Lam78h, Lut79, Mah76e, Mah76h, Mah76f, Mah76g, Mal79b, McC78a, Meh78a, Meh78b, Neu79, Nix78, Noh79b]. differential [Nus79, Olv79, Opp78, Ort78, Pel78b, Pet79, Ple75, Rab78a, Rau75a, RB75, Ros78i, Sch75c, Ste76d, Tak76a, Tak76b, Van79, Wam79, Wei75a, Wir78, Yam76, You75, vdB79, BR77c, BRP78, Gro76, JMR75, Rog75a, Sch78g]. Differential-Gleichungen [Gro76]. Differentialgleichungen [HK76, Jur78k, Kna79a]. Differentialgleichungssysteme [Fri78a]. Differentialoperatoren [Don76, Hee76]. differentials [Far79, GH75, Kru79]. Differentiation [Bru78f, Cha76b, dG75b, dG75c, dG76b, BW75, HS75a, Joh75a, dG75f, dG75i]. différentiel [Dab79]. différentielle [BR76b]. Différentielles [Cau76a, Jac79h, KS78b, Kre75a, KvQ75, LR78a, Mal79a, Roo78b, Ver79a, Yam78a, GR79c]. différentiels [CDI78, KV77, Kré76a, Las75, Mal79b, Raï78, Rog75a]. Différents [Dem78]. Differenzierbare [GMK75c]. diffusing [AG76]. Diffusion [BL75, iS78, AW75, CS76c, Don75, Had76, Roy76b]. Diffusions [Cau76b, Mal75b, Par78]. digits [Gal76e]. digraphs [HWH78, Rob77c]. digression [GW79c]. Dijon [JMR75]. Dikii [Ste78h]. dilation [Sha77b]. dilations [ACF79, Emc78, Mas78]. d’image [Ste75c]. Dimension [Edw75b, Gil75a, Gil75b, Lac76, Pas79b, Pas79c, Ser75d, Ser75e, Ser75n, Ser75o, And79a, Ari79, Aup79b, Bar75, Bar78b, Bon75, Dyc76, FGR75b, FGR75h, Gal76b, HT79, Hen75b, Kré77, Kré78a, Kré78b, Las78, Man76a, Maz76, Maz77a, McC78a, Nov77b, Nov78, Oso79, Pas79h, Pol79, Pop77a, Sch76b, Uen75b, Jan79b]. Dimensional [Fen79, Aki78b, ACH77g, Ber76a, Buh78h, Car79a, CW76b, Cha78b, Cie75a, Dus79, EH76d, Fat75, Gal78, Gla75b, Gla77, Gra78a, Gus75, Hay78, Haz79a, Jac75b, Law75a, Lev78, Loo75c, Mét78a, MS78b, Moe79, Ørn77, Sib76a, Top76]. Dimensions [CO77, AB75, Din75, FGR75f, Iva79c, Jam78i, Kah76a, Suf77, Uen75a, Uen75g, Utr79]. dimer [Rog79a]. d’incomplétude [Alo79]. d’induction [Rum79]. d’information [Nan75]. d’interface [Lév77]. d’interpolation [Pra78, Bea78a, Bea78b, Bea78c, Bea78e, Bea78f, Bea78g, Bea78h]. d’intersection [Bes75a]. d’inversion [Ber76g]. diophantiennes [Nic75b]. Diophantine [Che76c, Coh79a, Erd75b]. diphasic [Cha76d]. Dirac [Hal77, KSWW75, Mic78a, Ros78c, Ros78e, Ros78g, Ros78h]. Direct [BP79a, FR79c, Geo77, Hir76a, Ioo76, JS76, Rei77b, Smi76c, AEK78t, Arn77, GP77, SV75a, Tak78b]. directe [Ste75c]. Directes [BWK77f]. direction [Fit78]. directions [Blo79g]. Directs [BWK77b]. Dirichlet [Hee76, All76, Anc76, Anc78, Bor76, Con76b, Dou77b, Ede76, Eva76, Gas75, GP77, Gos78, Han79b, Jer76b, Kor76, Le 76, Nag76, Raz77, Rot76b, SNWC77d, SNWC77f, Sil76a, Yor76, dF75]. Dirichlet-Problem [Hee76]. Dirichlet-type [Eva76]. dirigé [Ill77b]. discontinuities [Jer76b]. discontinuity [Oli75]. Discontinuous [Zie79c, HB79, OW75, Tak76a]. discovery [Law76]. Discrepancy [Sto75a, Stu76b]. Discrete [DR75b, Fei78c, Geo79d, Gér75b, Hal76a, Lyc76, Sri79f, Cho78a, FJ75t, Hir79a, Lad77, Sch75b, Spe75b, Var77c, Vas78, vM79, VR75]. Discrètes [Gér75b, And77i, Bor75a]. discretisations [Kre78c]. Discretization [Fra77a]. discretizations [Ste76d]. discrets [Tit77]. discriminant [Hir77]. discriminants [Hof79m, Poi77, Sun79]. Discriminator [Hal79b]. discs [Rud77b]. disintegrations [Dub76b]. disjoint [Hei77]. disjointness [Aus79]. disjunction [BLR75e, BLR75f]. disk [BK77d, BK77e, CM77, Hay77, Rei76a]. disks [BK77d, Dav75a]. dislocation [Blo79b, Blo79h]. Dislocations [Blo79f]. d’isométries [Bop79]. dispersion [Man77]. dispersive [BBB76]. displacement [AV79, Itô75]. disque [Gau75]. dissection [Geo78a]. dissipatifs [Rot76a]. Dissipative [Grm79, Mla78, Dei77c, Hal76a, Haz77b]. dissipativity [BPFS76]. distallity [Dok78]. Distance [GL78, Lau75c, CV77f, DE77, Roc77]. distances [Sto75a, Wol75]. distinct [EN79]. distinguished [AT75f]. distortion [SS76c]. Distributifs [BWK77e]. Distribution [Kuo76, LT76a, Sto77g, Bab76, Bar79a, BP79a, Cor79h, Dam77, Dra79, Iag75a, Kuč75, Kue78, LT76b, LT76e, MR79, Phi76, Ros78f, Ros78h, Sch75d]. Distributionen [Haz77d, Haz77b, Haz77f, Haz77e]. Distributions [Ava77, Blo79f, Cho77b, Cor79c, GK78d, GR76, Har75b, LT76a, RS75, Ros78d, Rub75, BZ76, Bon75, Bop79, GS76, GK78a, GK78c, Haz77e, Ish76i, Jur78a, Kré76a, Kru76, Nic75a, Pir78, Rey78d, Ros78b, Ros78c, Ros78e, Ros78g, Roy76a, Sjö75, Wil75b]. distributive [Eps79b, Eps79e, Eps79m]. disturbances [Zee76e]. d’itèrès [Bru76c]. divergence [Axe79]. Divers [Bre77i]. Diverses [Bou78b]. diviseurs [Cam78b]. divisibilité [BR76c]. Divisibility [HLS79]. divisible [BF79a, Fal79, Fef75a, Kru76]. Division [HW75j, Bof75, Che76f, Gre76, HW75d, HW75e]. divisors [Ber76d, Nam79b]. Dixmier [Bor78c, Ren77]. do [Sch78g]. doctoral [Why78]. Doehlert [Sta78b]. Dolbeault [Kan75a, Kan75b]. domain [Ber76f, Eas78, Gri76b, Hah79, Mol77d, Ole78]. domaine [Fey79b, KS78b, Mok78b]. domaines [Sib76b]. Domains [Gin79, Wil75a, Zie79d, BL75, Bra79e, Bra79i, Bra79n, Cho76, Far75f, Hay76, Pfl75b, Pfl78, Stu76a, Wil75b, Wil75e, Wil75g, Wil75h, Wil75i, ZKB79, Pfl75c]. dominance [NT79e, Nar78]. Dominant [FGR75b]. Domination [Coc78, CL75b, Mis78]. d’ondes [Gat77]. donné [GMP78, Kis77a, Zan77]. dont [Pou79]. doodles [FT79]. d’Opérateurs [dlH79c, BLC78, CDI78, Far75a, Hel78a, Kar79, Sjö78]. d’optimisation [Nic75b]. d’ordre [Ava77, Dam78, Le 77, Smy78b, Mol77u]. d’Ore [WK78]. d’Orlicz [DC75b, Lac78]. Douady [Fel77c, Le 75]. Double [Lat79, Per78c, Sto78a, Zee76f]. double-cusp [Zee76f]. Doubly [Gra76e, Gra76g, Gra76i]. Dreieckig [Jur78e]. dress [ACH77a]. droite [LR78a]. Dual [Gri75, Hla79, Hof79l, Lau75c, OL77, Bol75, CV77h, ER75, Hol79, Mey77b, Moh79, MM79b, Raw79, Rod79]. dual-continuous [Moh79, MM79b]. Dual-mixed [OL77]. dualisant [Fou75]. Dualitaetssatz [RR75a]. Dualité [Bea78b, Del77c, FdlP79, Mey75b, FdlP76, de 78, Băn78]. Duality [CM76, FGR75c, Mic78f, NT79c, Roc78, BM77a, BDE^{+}77g, Bin77, Frö76, Her77c, NT79a, Del77c]. dualizing [Sha79b]. Duals [Lew77]. Dualsants [Gro77b]. Dubins [Cho78b]. Dublin [AD78b, AD78a]. Dubovitskii [Das75]. Dubreil [Mal77, Mal78a, Mal79c]. due [Fel77c, Kre75c, Tol75c]. Duflo [Bor78c, Raï78]. d’un [Ari79, Bar75, Ben77b, Bru76c, Cat75, Cho78b, Dab79, ÉS79, ET79, Fey79b, Gue75, Gui79b, Hen75c, MM75a, Maz77a, Mey79, Mok78b, Poé76b, Pra78, RR78, RR79a, Rap76, Ser76, Zan77, Jac79p]. Dundee [Eve75a, ES76, Wat76, Wat78b]. d’une [BLC78, Bru79b, CB75, Dar77, Duv77, Hen75b, Jac78b, Jen78, JY78, LR78a, Las75, Lat79, Mal75a, Maz76, Meb78, Mok76, Pra79, Reb78, Ren79b, Roc79, Sim78b, Vai78, Yor78a]. durch [Ess76, vG76]. Durchmesser [Küh79]. Durham [FMS79]. Dwyer [Bon75]. Dyer [Lan78, Sna75]. Dynamical [MMP78, OP79, Sma77, Ara78, DGS76j, Gra79a, Grm79, Has78b, HM78b, Shu75, Man75d]. dynamics [Adl79b, Aus79, BS76a, Fix75, Ger79a, Gui77c, Lic77, Roa75, Tul77a, Tul77b, Wal75i]. dynamiques [Sig75]. Dyson [Sto75b].

E. [Bou79b, Car78c, HW75r]. East [Kel75c]. easy [Fin75]. ebener [Fri76b]. Échelle [Lac78]. Ecole [DCHRH78, HJLN77, BGM77, BKK76, FCG75, And77a]. écoulements [Dar77]. eddy [NV79e]. Edwards [Lat79]. effective [Mas75a, Nam79b]. efficiency [MBY79]. efficiency-robustness [MBY79]. Efficient [Bet78]. effilement [Smy78a]. Ehrenfeucht [FR79d, FR79e, KK76]. Eichler [Zag77a, Ral79b]. Eigen [HH77a]. eigenforms [Rib77]. eigenfunction [Wil75f]. eigentlichen [Jur78d]. eigentuple [Blu79]. eigentuple-eigenvector [Blu79]. Eigenvalue [Roc75c, Sch75c, BG78c, KY79b, Kel79, PP79, Pel78a, Sle75, Trè75]. eigenvalues [HH77e, HH77f, MS78a, MS79c, Raţ79, Roc75b, Ruh77, Sat79c]. eigenvector [Blu79]. eigenvectors [Ruh77]. Eilenberg [Gug78]. Eindeutigen [Wec75]. Eindeutigkeitsproblem [Fri76b]. einem [Fri76b, Jan79f, Jan79g]. Einfache [Voi77i]. Einführung [Mét77c]. Einheitengruppen [Kra75f]. einige [Vek76, Col76c]. einiger [Kup76]. Einleitung [Gro79e, Jan79c, Jur78f, Kra75c, KPP^{+}78h]. Eins [Jan79e]. Eintritt [Dic75]. Einzigkeitssatz [Jur78c]. Eisenstein [Ish76a, Lan76c, Lan76g]. elastic [Blo79c, Chi76c, OW75]. elastica [FJ75r]. elasticity [BKM78, Blo79d, Boh78, Cap79, Miy77, OL76, YF79]. Elastizitätstheorie [Kup76]. elastoplastic [Mor76]. ELC [Nov75]. electrical [PR77b]. Electromagnetic [Bar78f]. Element [GM77n, GR79e, BR77a, DN77, Fra77a, FS79b, Gal75, GM77p, Geo77, Geo78a, GR79g, GP77, Hla79, Kik79, KO79, Miy77, NV79e, Nit77, OW75, OL76, OL77, PP79, RT77, SB77, Sco79, YS79]. élémentaire [Bru76d, ÉS79]. Elementary [AT75c, Dub76a, FSW75, NT79d, Sch76p, Sch75i, Fin76, KMS75, Lam78g, Mey75a, Osi75a, Wra75, Zee76a]. elementary-recursive [Mey75a]. Elements [BK77f, Sat79d, Tay78j, AD79, BB76, CL75e, Fix75, Jan75, Luc75, PT76, Tay78h, Spr77f]. elimination [Gen76, Kra75g]. Eliminationstheorie [KPP^{+}78a]. elle [Mar79a]. Elliptic [Cor79d, Cor79e, Mal75b, Mas75b, Moi77c, Ama79, Bre77o, Cho77c, DN77, Dui75, FR76, FJ75b, GM77c, Gol79, Gri76b, Ham75g, Hes79, Mas75c, Mil76, OL77, Rab78a, RT77, Sei75, Stu76a, Tat75, Trè75, Web78a, You75, dF75]. elliptical [Ril79]. elliptique [Fit78]. elliptiques [BLC78, Cha76e, Del75, Dem77c, Hue77, Rot78]. elliptischer [Don76]. elliptisches [BG76b]. Ellitpic [SDB75]. embedded [Bet78, Ecc78, Mor79c, Bin75c]. embedding [Ale75, Bru76a, Can75a, CW76b]. embeddings [Eva75, Gau79, Mac75a]. Emphasis [GK78d]. empirical [CB76, GS76, Kue78, Neu76, Phi76, Rév76, GR76]. endlichen [Voi77n]. endlicher [Voi77s, Voi77h, Voi77f]. endomorphism [Hau77]. Endomorphismen [Kra75a]. Endomorphismen-Schema [Kra75a]. Endomorphismenringe [Voi77j]. endomorphismes [Ari79, Ser75p]. Endomorphisms [GH78a, PPW78]. ends [Gel76c]. Energie [Küh79]. energy [CL75c, Cos75, Wil75b]. Engel [War76g]. engendrés [How77]. Engineering [GLL79, Cot79, HM78b]. enlarging [KT76]. Enoncé [CH78c]. Enriched [Eis77a]. Enriques [Van78b]. Ensemble [CL76a, Gil78g]. Ensembles [Bes75b, GB79, Mok79, Del77j, Gru77, JM76, Maz77a, Nov77b, Nov78]. entière [Rhi75]. entières [Gra78b, Kis77a, Ser76, Yam78b]. entiers [Gér75g, Nic75b]. Entire [Sha77a, BY77, GR75a, RY77]. Entre [Lau75c, Mey75b, Blo77, Gol78a, Lac78, RR78, Yor79]. entropie [Rau75d, DG76a]. Entropy [Bow78, CP77b, DGS76b, Wal75b, DGS76a, DGS76g, DGS76m, Den76, DGS76s, DGS76u, Led79, MN79, Mis79b, Wal75j]. enumeration [Che78a, Ker75a]. énumerative [Le 78]. Envelopes [Fla76]. enveloping [Bor78c, Jos79]. enveloppante [CB75, Dix75, Ren79b]. enveloppantes [Bar78c]. enveloppe [Anc78]. Enveloppes [Dlo77]. Epilogue [FJ75f]. Equadiff [Fáb79]. Equality [Fri75, LeP78a]. Equation [Wil75a, Aiz79, Ati75, BPFS76, Bar78a, BTW75e, BTW75g, Bre79d, Col76e, Con79g, Cop78e, Don75, Eas76, Ell75, GW76a, Gre75b, Gri76b, Hac78a, Hal76b, Hal77, Ham75f, Han79a, LdMP77, Lon79a, Lov75, Meh78a, Meh78b, Noh79a, Noh79b, Olv79, Øvr76b, Raţ79, Rau78, Sim78a, Sko76, Van79, Wah76, Wid78, Wil75d, Wil75e, Wil75i, BCH78, Car78c, Mar75, MR77m, Rab77, Sim78b, dM75a, dV75a, Bre78]. Equational [Hal79c, Smi76f]. Equations [AT78, Bib79c, BGS78, Cha78a, Cha75c, Cop78g, Dei77h, Eve75a, ES76, GM77d, GR79e, Gol75, Lan76g, LS79a, MWW76, PW79, Pin75a, Ree76a, Ros78d, Sch76n, Sch76p, Sch76o, Tem76c, AW76, Adl79a, AB75, Alt79a, Alt79b, Ang79, AL79, AS79, Axe77, Axe79, BM79a, Bak78b, BZ78, Bar79b, Beb79, BP79a, BDNY77, Bih79, Bit76, Böh76b, BBB76, Bor79c, Bra79a, BTW75c, Bre77o, Bro75h, Bru76e, But76, Car76a, CV77i, CG79b, Cla79g, Clé79, Coh79a, Col75, CS76c, CS78, CN76, Cox79, Cus79, Dei77d, DN77, DDP78, Düm79, Dwy78b, Džu76, Eas75, Eav79, Egu79, EW78, FR76, Fit76, Foi75, FT76, Fre78, Fri78a, Fro75, FS79b, GM77c, GM77k, Gaj79, Gam79, Gat77, Geo77, GH76, GR79i, GR79j, Gla75b, Gra76a, GV79b, Gro79h, Gro78f, Had76, Had79]. equations [Hal79a, Hal79i, Ham75g, HB79, Her75d, Hir76a, IK79, Ioo76, Jak79, Jor79, Jou79d, Jur78k, KY79a, Kat75, Kre78c, KS79b, Lak79a, Lam78h, Lan76h, Lev79a, LY79, Lut79, MW79, MM79a, Mah76e, Mah76h, Mah76f, Mah76g, Mar77, Med75, MG78, Neu79, Nix78, Nus79, Ole78, Opp78, Ort78, Pel78b, Pet79, Rab78a, Raţ77, Rau75a, Rei77b, Rhe75, Ros78i, Run76, Sar78, Sch76k, Sch76l, Sch75c, Tak76a, Tak76b, Tem76b, Trè75, Wam79, Web78a, Wed78, Wir78, Yam76, You75, dF75, vdB79, Lai78, Bon78b, Dam78, Jou79b, Krá76b, LS76, Lum78, LP79b, Mal79a, Paq79a, Sec75, Sib77, Tho79a, Vig77a, Cau76c, GR79c, Jac79h, Jou79c, Jou79e, Jou79d, Las78, Lév77, Ver79a]. equi [Kös76b]. equi-spaced [Kös76b]. equicardinal [Bou78a]. equicontinuity [Sch75j]. équicontinus [Deh76]. equidistant [EH78]. Equidistribution [Edr77]. Equilateral [Law75a, Kel75b]. Equilibrium [Bow75b, Geo79e, Wal75c, FJ75r, Fra77a, Kas75, Kas78]. équirépartition [Bes75b, Cou75, Ger75i, Nie75b, Gui76b, MN75a, Rau75d]. équisingularité [HT75a]. Equivalence [Hit78, Kuo75a, Mol77-28, CW78, CMW78, FSW75, GH78b, JS77, Laz75e, Sch75a, Ste79, Mol77z, Mol77-27]. Equivalences [Joh79h, Joh79a, Joh79i, Joh79k, Laz75g, Mun79]. Equivalent [Din75, Hof77, tD79g, tD79h]. Equivariant [HPS77d, tD79d, tD79e, Dov79, Ewi79, Hod75, KS78a, McL79]. équivariante [Poé76c, Poé76d]. équivariantes [Ron75]. Erdélyi [Sne75]. Erdos [CC76]. erections [Rob75a]. Ergebnisse [Jur78s]. Ergodic [Bow75b, Bow75g, Bow75c, Bro78d, DGS76c, DGS76d, Lin79b, Wal75d, Akc79, Aus79, DGS76m, FW78, Gal76j, Gre78a, GK76b, HO79, Kam76, LS79b, MP79a, RN75, DJ79]. ergodicité [Mic76a]. Ergodicity [DGS76e, Aar79, Sch79h]. Ergodique [CK76, Con75b, Cou75, Per77, Sig76]. Ergodiques [CK76]. Errata [Dup78h, Hen75a, Jan79d, Lad76a]. Erratum [Del77d, GR79d, KT75d]. Error [Dah76, Mos77a, SS77a, BC77a, Ste78i, Bai77b]. Ershov [Che76c]. erster [Mei78]. erzeugende [Haz77e]. erzeugenden [Haz77f]. erzeugender [Haz77d]. Espace [Bar75, Ari79, BF75b, Bar78b, Gol78a, GP78b, Jou79e, Kor76, Kré76a, Lac78, Lac79, Lel75a, LBL78, MM75a, Maz76, Maz77a, Rab77, Rab78b, Rap76, Rot76c, Sch76f, Ser76, Suz78b, Szp76, Tak79]. Espaces [BF75c, Bea78a, Bea78c, Bea78e, Bea78f, Bea78h, Fey78b, Fey78a, Jac79q, Mok78a, Sch76f, Sch76g, Sch76h, Sch76i, Tur75, Anc76, Cam78b, Cha79b, CO77, DC75a, DC75b, Deh76, DMY78, Erm78, Far75c, Fd76, Geo78b, Gui76b, Hen79b, Kaj75, Kan75a, Kan75b, Kos76c, LL76, Lap79, Le 75, Lel78a, Lel78b, Lem77, Lod79, Loh75, Pis76, Pou79, PS79, Raï75, Sch76d, Sjö76a, Smy78a, Ste75d, VVT77, de 78]. especially [Col75, Poé76b]. Ésperance [Cau76h]. Essential [Joh79e, Iag75a, PS77]. essentially [Dav76, Sal78]. essentielles [Krá76b, Meb75]. est [Bré76a, Mar79a, Mul75a, Rhi75]. est-elle [Mar79a]. Estimate [Car79b, Ber76a, Mol78]. estimated [CB76, Dur76a, Neu76]. Estimates [Bai77b, And79a, BF76a, DN77, For76a, Ham75c, Mal75b, Mos77a, MG79, Ood78, Ros79, Var78]. Estimation [BBDB^{+}78, GR79b, Geo78b, Gra76f, Gra76g, GM79a, GR79a, Gra76i, Kue78, Par79c, Ste78i, Geo78b]. Estimations [Bai77b, Bar77a]. estimators [MBY79, Rey78c]. etabli [RS78]. etale [Gro77d, DB77, Del77b, Del77b]. État [Ver77]. états [Sig75, Sou77]. etendu [Mok78b]. étude [Ars79, BDR78, CO79a, Reb78, BF75c, Bea78e, Bre77i, GS79b]. Euclidean [Jen77g]. Euler [Bar78a, Mor77a, Tem76b, Zee76c]. Eulerian [Bou78a]. Evaluation [AL77, Bin75i, DV75c]. evaluations [Fer76b, Fer76a]. Evanston [BM78a, BM78b, Ste76c, Zel76b]. even [ACH77g, KY79b, Law75a]. even-dimensional [Law75a]. event [Fin75]. éventuellement [Ver76]. everywhere [Sha77b]. evolution [BPFS76, Dwy78b, Fit76, FS79b, Gaj79, Hir76a, KM78a, Kat75, Kre78c, Lig77a, Kôn76]. evolutions [Pin75b, Too78]. ex [Fou79]. ex-conjecture [Fou79]. Exact [Cza75, Fat75, AB76, Hir76a, Pra79, TY78, Zel76a]. Example [Roc75a, ČT78a, Con76a, FR79e]. Examples [Bar79e, BJM78b, Bra79e, Bra79f, Bra79g, Far76d, Jam78g, KT75e, Sil76c, TW78b, Uen75h, Wei76h, AEK78p, Ass79, BKM78, Bre77e, CW77c, Goo79a, Hub78b, KMS75, Rat78l, Sat79f, TDHJ76, Fei78j]. excellents [Mar79a]. exceptional [Car79b, Hir75b, Jac79a]. exceptionnel [Tak79]. exceptionnelles [Sig75]. excessive [CL75e]. excited [AW78]. exemple [Aub79]. Exemples [Jac79l, dSLM75a, Dem78, Hel78a]. exhaustion [GW79b]. exhaustive [Wei76s]. Existence [AS79, BK76b, Hec76, Jer76a, Lak79a, Mah76c, MR77g, Pre76d, Ree76b, She75, BZ76, Bla76f, BK76c, Chr76, EWW77, Fef75a, GM77h, GvW76, Hes78, HT78, Jer76b, Kis76b, Kos75a, Mos78b, Nor78a, SNWC77a, Sha79b, SW78f, Tem76b, Zan77]. Existentially [Che76f, Che76g, Che76h, HW75e, HW75f, HW75g, Bof75, FSW75, HW75d, Sch75a]. Existenz [Eif76, Fri76b]. Existenz- [Fri76b]. Existenzsatz [Hei75]. exit [Gus76a]. Exkurs [Voi77k]. Exotic [Heg79, Cor75]. expanding [Col76a]. Expansions [Gro78h, Bro78a, Dre77, Eck77, Gal76b, Wil75f, Win75, vR79]. Expansive [DGS76f, Has78a]. Expectation [Fuj76, Cau76h]. Expectations [AHK76b, AHK76a, BP78, DV75c]. expected [Par79a]. Experience [Fro75]. Experiences [Mül78a]. experimentally [Pin79]. Experiments [Lon76, CY78]. explanations [Wei76s]. Explicit [Köh75, Mor77b, Fef75b]. explicites [JY78]. explicitly [Sch79j]. exponent [Man76a, SS78b]. Exponential [Cop78h, GK78d, Gam79, Gla76b, Cop78a, GK78c, GW79j, Kue78, Sch76m]. exponentielle [Roc79]. exponentiellement [Tur75]. Exponentielles [Jac79i, Bai77a, Mém78, MS79b]. exponents [CMN79, War76i]. expose [Str79d, Ber77b, Ste75c]. Exposés [Ano76, Ano77, Ano79, BDV^{+}78, Bou75, And77a]. Exposition [Rat78c, Mos75, Ros75a]. Expressions [KRZ77d, Eva76, EG75, FJ75d, KRZ77c]. Ext [Mek77]. Extendability [Bie77, MM75b, MS76b]. extended [Hal77]. Extending [How78, Laz75b, Gau79]. Extension [Cho78b, Hir76b, Jen78, Ben77a, Bin77, Che79, Cor79a, DS76a, FK79e]. extensional [Fef75c]. Extensions [Bar79f, BWK77c, Cur75a, Dou77a, FGR75i, Geo79i, LT76a, Lub78, PS77, Rog76b, Smi76d, BK77j, Bre79a, CS76a, Dub76a, Far75g, Fie78, Flu75, GS79a, Har79a, HW75a, Ish76h, KN78, Kre76b, Las77, Run76, Sch75r, WK78, Cho75b]. extérieur [Kor76]. Exterior [Wil75a, AF78, Est76, Far76h, Kur75, TW78a, Wil75b, Wil75g, Wil75h, Wil75i]. external [BH76c]. Extremal [Sch75j, AV79, And79a, Duk78, Erd78, ES77, Mul78b]. Extremaleigenschaften [GMK75d]. Extremales [Jac79p, Yor78b]. Extremalisation [Gui79b]. Extremals [FJ75l, FJ75d]. Extreme [Bri76, HH77a, Urb75a, Bru78g, Sch75n]. Extremely [SS77c]. Extremum [Tam78c].

F [Ram79b]. F. [Car79e, Des79a, Lod79, Mey76c, Mic76b, Mol77-28]. Faber [Sch75k]. Facteurs [Vig77a, Com76]. factor [Ply79, SV75e, SV75f, Wei76i, Tak78c]. factor-cells [Wei76i]. Factorisation [Qui79]. Factorisations [Vie78a, Vie78d, HRW78, Rob79c]. Factorization [DW79a, DW79c, Gro77a, Jos76, Ber76f, Har79b, Kas79, MS78c]. Factors [Tak77b, CW75, HR76, Jam78d, Stø78b, YS79, Sai78]. facts [Bra79c, Bra79h, Dav78h, Jur78i, McK75]. faible [Led76, Mey77a, Mok76, Sch76e]. faibles [Kré75b]. fail [Her76]. faisceau [Bar78b, Ste76b]. Faisceaux [BWK77g, Lem79, Ray78b, Ste76a, Car79e, Thi77]. faithful [Rob76e]. Faithfully [Mag79a]. Faktoren [Haz77a]. Fall [Mét77b, Mét77c]. Faltungs [Haz79b]. Faltungs- [Haz79b]. Faltungshalbgruppen [Haz77a, Haz77e, Dri79]. Faltungspotenzen [Ger79d]. Familien [FK79b]. Families [CL75d, Nam79b, Nam79c, Nam79d, Nam79e, Bib79a, Bre75b, Erd78, HPS77m, PS78, Pop77d, Ros78e, Sch75f, Sch75h, Sch75n, She75, Won76a]. Famille [Rit76, Reb78]. Familles [For76b, Fey79a, HT75a, Vie78d]. family [CL75f, Don75, EW76, Kan79, Rod79, Str79a]. famous [Dev77d]. Fano [IŠ79]. fast [Hac78a, Pow78a, STW78]. Favard [FJ75p, FJ75s]. features [Kas78]. February [GR75b, Hey79, NR78b]. feeble [Kel75b]. Feed [vH79b]. Feed-back [vH79b]. feedback [Abr76b, DH79]. Feinstruktur [Jur78g]. Feller [Far75c, Rot78]. fermée [Mar79a]. fermées [Kre75a, KvQ75]. fermés [GS79a, Lel77b]. feuilletage [Gue75, Sec75]. Feuilletages [Hec77, Fed78, Hec78, Jou79f, Rou78, Suz78a]. feuilletées [SW78b]. feuillets [Hen75b]. Feynman [AHK76c, AHK76d, AHK76e, AHK76j, Gaw77]. FGC [Bra79c, Bra79h]. fiber [Gli78, Mun79, Tat75]. fibered [Bir79]. Fibonacci [HLS79]. Fibonacci-type [HLS79]. fibration [Mag76]. Fibrations [BLR75b, BLR75c, Bau77c, Bau77g, Bau77h, BHP78a, Got78a, Heg79, HPS77d, Kau78, LR78b]. Fibre [BHP78b, Erm76, Pop77a, Pou79, Uen75a]. fibres [BE78, Ele78, Sch77h, Ste75b, Cam78b, Dem78, Erm76, Lem77, Sko78, Le 77, Pou79]. field [AHK76c, Che79, DFF78, Dou78, EGS75, FI77, FP76, Gre77b, Hal76b, KL77, Len76b, Man76c, Mul79, Sha76b, Smi79, Ste78g, Sto75b, Str75, TW78a, Tul77a, Yan77e, Yan77g, Yan77h, Yan77j]. Fields [FP78f, Ish76d, Pre76h, Sch76p, ACH77c, Ave78, Bar78f, Car78a, Che75b, Che76c, Che76b, FP78k, FP78j, Fri76a, Gar77a, Gli78, Gol79, Goo79c, Gut75, GdM77, Hir77, Hua79, Ish76a, Ish76c, Ish76e, Ish76f, Koc75, Man78b, Mas79, May77l, Mey76b, Mic78b, Mül78a, Pas79c, Pre76i, Raw79, Saw75, Sch76r, Sta77d, Sta78d, Sun79, Zag77b, dlR75, vO75d]. Fifth [Hey79, Lit77a]. fil [Del77h]. filaments [Gui77c]. Filtered [Hof79g, NV79a, NV79d]. filtering [Gra78a, Mét78a, Moo78]. filters [Hir75c]. Filtrage [Par78, Bré76a, Yor77]. Filtration [Jac79c, JY78, Yor78a]. Filtrierungen [Jan79e, Voi77l]. fin [MD78]. final [Her79a, Mol77h]. find [Olv79]. finding [Hir76a, MJ75, Pin79]. Fine [DG76c, Maa78, Pop77d, Fd76, Jur78g]. finest [Fro79b]. Finetti [Geo79i]. Fini [Gil78g, Ave76, Bop79, Fer77, Le 77, Sab76]. finie [Aup79b, Bar75, Hen75b, Maz76, Maz77a]. Finis [BWK77e, GR77, Gér75d, Ser77]. Finite [Aki78b, BF76a, Con75a, Coo77, Eps79e, FP78f, Fix75, Gab75a, GM77n, GR79e, Hir75c, HW75h, Kik79, Loo75c, Lub79a, Per78a, RR79b, Sch76p, Spr77c, Sri79h, YS79, AM78, AD79, Ass79, BR77a, BB76, But75, CM78b, DGS76q, DN77, EL76, Eps79m, FP76, FP78e, FP78k, FP78j, FGR75b, FGR75h, Fra77a, Fri76a, FS79b, FW78, Gal75, GM77p, Geo77, Geo78a, Gil76, GR79g, GH78a, GP77, Gre75a, Gui77b, HM79b, Han75a, Hin77, Hla79, Hub78b, Jac75b, Jan78, Kan78a, KO79, Kuk76, Kup75, Lin79a, Lou75, McD78, Mic77, Mis78, Miy77, Mül75b, NV79e, Nit77, OW75, OL76, OL77, Ørn77, Par79a, RT77, Rei78c, Rog75b, Rog79b, SB77, SNWC77d, SNWC77f, Sch76r, Sco79, Sha76b, Sri78, Top76, Tra76, Tra78a, tM76, tD79c, tDP79]. Finite-dimensional [Loo75c]. finite-element [OL76]. Finite-time [BF76a]. finite-type [Hin77]. Finitely [Bie79, Bra79d, Cor79f, RS76, Smi79, Mah76f, Tur78]. Finiteness [AOL79, Gre75c, KLM75, Tep76]. Finitistic [AM78]. finitude [Bre75a, Del77g, Jac77a, Tei75]. Finland [LLS79, LS79a]. Finnish [CCJS79]. First [Bec76b, BTW75c, Flu75, Mah76d, MR77h, SW78g, Tho79b, Tol75a, Tol75e, Bru78d, Bru76e, Cos75, Fre78, Gus76a, Hag78, Mac78, Mah76c, MG78, SW78e, Sto77b, Sto77c, Why78, MR77f]. first-kind [Bru76e]. First-order [Flu75]. First-Passage [SW78g, SW78e]. fitting [BC77b, Lam78h]. Five [Kap78, Rau75a]. fixe [BG78a]. Fixed [BS78, Ber77a, BJM78c, Dei79, PW79, Van78a, Bro75g, CK75, Dol78, Dub76a, Ede75, For79b, LT76e, Sha78a, Sha78c, vdLT79]. Fixing [Chi78a]. Flag [Liu78a]. Flat [Hir78b, Bla76f, Dup78a, Dup79, Hau79, Hir75a, KT75a, KT75i, Mol77v]. flatness [Dav75b, Fis76e]. flattening [Ben75a]. flattenings [Can75b]. Flexibles [Kui79]. flips [Joh79c]. flot [Ben75b, dSLM75b]. Flots [Mey75c, Ben77b, Cha76e, dSLM75a, Mey77c]. Flour [And77a, BKK76, DCHRH78, FCG75, HJLN77, BKK76, DCHRH78, FCG75, HJLN77]. Flow [Dei77e, AG76, ADB75, Ber76e, Bow77, Gui77c, HH77h, LD77, McL75, Zie75]. Flows [Fei78k, Gla76i, Bow75g, Bré75f, Cha76d, Ell78, Emc78, Fel79, Gla76h, Gla76g, Gla76j, Per78b, Saf77, Wir76]. ^{≪}Divergentielle^{≫} [Dam78]. fluctuation [Doh75]. fluid [AG76, Ben76a, BS76a, Fix75, Lei79, Lum77, Roa75]. fluid-dynamics [BS76a]. fluids [BF76a, Ber76e, Tem76b]. Fock [Ono77]. Fok [Pes76b]. Fokker [Rau78]. fold [Tho76a]. folds [IŠ79]. Folgen [Ess76]. Folgerungen [Voi77m]. Foliated [Hir75a, KT75f, KT75g, KT75b, KT75e]. foliation [Cor75, Dej78b, Hir75b, SW78a]. Foliations [Eps75, KT75h, Pla75, Sch78g, CN77, Cam78a, Edw75a, ER77, Eps77, Fed78, Gar78, Hae78a, Her77a, HPS77l, Hol78a, Lam75, Mil75b, Pal78b, Sch75u, Sch75v, Ste75a]. following [Bou78a, LY79, Mor77c]. Foncteur [Bou78d, Bou78e]. Fonction [Fit78, Ben75b, Fd79, Jen78, dSLM75b]. fonctionnelle [Mar75]. fonctionnelles [Att78, Kré75b, Vig77a]. fonctionnels [Fey78b]. Fonctions [BWK77n, CL76a, Cup77, Éli79a, Gil78e, Gil78j, Gra78b, GD77a, Kis77b, Lel77a, Sch76i, Ste75b, Tak75a, Yam78b, Ali75, And75, Ave76, Bad78, Bai77a, Bin75a, BR76c, Boc78, Bou79a, Bru75, BG78d, Car75, Che75a, Coe77, Fer75, FdlP76, Gui79b, Han76b, Hou77, Kis77a, Kôn76, LL76, Lac79, Lel78a, Lel78b, Lum76b, Mar75, MD78, Mat77d, Mok76, Mok78a, Mok79, Poé76a, Raï75, Ris76, Sch76d, Sch76g, Sch76h, Ser76, Sjö76b, Smy78b, Tou76, Wal76, Wal77b, Nor75, Nor78c]. fonctorielles [Kré78a]. Fondamentales [BWK77h]. fondamentaux [BF75e]. Forced [HPS77e, Meh78a]. Forcing [HW75j, HW75i, Kec78b, Ada75, Buk77, Hir75c, HW75h, HW75l, Mos75]. Ford [Gol75]. Foreword [Pfl75e]. form [AHK76g, Axe79, Buh77, Buh78b, BLR75e, CH78d, Har76a, Jam78e]. Formal [Lau79a, Laz75a, Laz75c, Laz75d, Mah76e, Sch79a, Yan77g, Adl79a, Gia78, Haz79a, Lau79c, Laz75b, Laz75f, Mah76b, Ste78h, Yan77s, Jur78g, Jur78h]. Formale [Jur78g, Jur78h]. formalen [Jur78l, Jur78p]. formalism [Bow75d, Das75, Rue77]. Formalisme [Rue77]. formally [EG75]. formation [Sat76b, Zee76d]. Forme [CH78d, Anc78, Le 76, Vie77]. formelles [Gér78, WK79]. formels [Sak77]. Formes [Alo79, Bor76, Dab79, KvQ75, All76, Bou79b, CO77, Kre75a, Let79, LV75, Meb75, Pes76a, Roo78b, Rot76b, Sub75, Vig77b, Yor76]. Forms [Bae78e, Far75d, ACH77e, Alo79, And77k, Bae78a, Bae78d, Bla76g, Bre75c, Buh78a, Coh77, DO77, Dup78g, Est76, Far76d, Far76h, FJ75d, Gel76a, Gel77, Hae78b, Hej76d, Hir77, Iit79, Kat77, Kre75a, Kud77, LM79, Lan76b, Mad78, Mas75e, Mas75d, Mas75h, Mas75g, May77l, Mor77b, Pom76, Raw78, SS77b, Sim75b, Sta77d, SD77, TW78a, Zag77b, tD79j]. Formula [Hej76a, Hej76b, Hej76d, Hej76e, Hej76f, Jam78i, Mac78, McL79, Mor79a, Pop77a, Uen75a, War76c, Wil76a, Zag77a]. Formulae [TW79, Cla79h, Oli75]. Formulaire [Del75]. Formulas [Gas75, Gro78i, Bot78, FR79k, Hir79a, Mor77b, Pie78]. formulation [Cha76d, Tul77a, Tul77b]. Formule [Cha77, Gro77d, Gro77c, Loe75, Rot76b, Sub75, Del77a, Del77f, Nic75a]. Formules [Gro77e, Jac79i, JY78, Roo78b, Bes75a]. Fort [Nør78d]. forte [Kos76c]. fortement [Gos78, Mok79, Vai78]. fortes [Cai78]. forth [Cun75, Kue75a]. Fortsetzbarkeit [Wec75]. forward [Fro75]. Foundation [Gol75, Fri78d, GR79f]. Foundations [Lam78b, Mol77m, Bec79b, Ner75, Pét77]. Four [Blu75, Dev77d, Ber78b, HWH78, Hei77, Kel75b]. Four-point [Blu75, Kel75b]. Fourier [AT75d, Bar79a, Ber76c, Bon75, BTW75d, Cha75c, Dre77, ET79, Gui75, HH77b, HH77c, HH77d, JL77, MP79b, MS75, Nic75a, Pir78, Sjö75, Zag77b, dBS76]. Fourth [CW76a, CUW75, ES76, Nov77a, Bre77o]. fractal [Man76a]. Fractals [Man77]. Fractional [Erd75a, Kuč75, Str75, Ask75, Bak79b, BW75, GR75a, HS75a, Joh75a, Kis76a, Lau75a, Mil75a, Ros75a, Ros75b]. fractionnaire [Kis76a]. fractionnaires [Hir76b]. Fractions [Bre77l, Gil78f, Ker75f, Gro78g]. Fragments [Gra79b]. framed [BR78b, Ecc78]. Frames [Cre75, DS76c]. framework [Hal79f]. framings [BS78]. France [BR78a, CV79, Ser75a, Ser75a]. Frank [May77a]. Frankel [Nør78d]. Frattinialgebra [Voi77k]. Fréchet [Bol75, Dub79g]. Fredholm [FJ75c, GM77f, GM77e, GM77j, GM77l, Het79]. Fredholmschen [Häm76]. Free [BLS75, AHK76i, AH79, Arn77, BF76b, Ber77a, Cum76, Eis77a, Fef75c, Gla76b, Har78a, HMRS78, Hof79h, Kaw79b, Lad77, Lio75a, Mac77a, May77f, MJ75, Mek77, Moe79, Rei75, Sch78b, Sch78f, Sij79]. freeness [Hil77a]. Fregean [Sus75]. Freidlin [Ood78]. Freiheitsgrad [Jur78m]. French [Ali75, All76, Alo79, Anc76, Anc78, And77a, And77i, And79b, And75, Ano79, Ari79, Ars79, Art75, Att78, Aub79, Aup79a, Aup79b, Aup79c, Aup79d, Aup79e, Ava77, Ave76, Avr79, AM79b, AY79, BKK76, Bad76, Bad78, Bai77a, Bai76, Bai77b, BF75a, BF75b, BF75d, BF75e, BF75f, BDR78, Bal79, Bam79, Băn78, Ban75, BG78a, Bar77a, Bar75, Bar78b, Bar78c, BE78, Bas75, Bea78a, Bea78b, Bea78c, Bea78e, Bea78d, Bea78f, Bea78g, Bea78h, BDPR76, BBL77, BP79b, BLP77, Ben75b, Ben77b, Ber76b, Ber76c, Ber77b, BM77b, Ber76g, Ber79e, Bes75a, Bes75b, BWK77a, BWK77b, BWK77c, BWK77d, BWK77e, BWK77f, BWK77h, BWK77i, BWK77k, BWK77j, BWK77l, BWK77m, BWK77n, BWK77o, BWK77g, Bin75a, Bis78a, Bis78b, Blo77, BR76b, BR76c, Boc78, Boh75, BCH78, BLC78, Bon75, Bon76]. French [Bon78b, Bon78c, Bop79, Bor75a, Bor76, Bou79a, BT77, Bou79b, Bou76b, Bou78b, Bou78c, Bou78d, Bou78f, Bou78e, Bou78g, Bre75a, Bre78, Bré76a, Bre77g, Bre77h, Bre77i, Bre77l, Bre77j, Bre77m, Bre77n, Bre79f, Bré76c, Bro78b, BEGS75, Bru78r, Bru75, Bru79b, Bru76d, Bru76c, BG78d, Buc78, Bur76, BR76d, CO79b, CO79a, CW77b, Cai78, CG79a, Cai79, Cam78b, CL76a, Car75, Car76b, Car77b, Car78c, Car79e, Cat75, Cau76a, Cau76c, Cau76e, Cau76f, Cau76g, Cau76h, CCP75, Cha77, Cha79b, Cha76e, Che75a, Che78c, CC79, Cho75a, Cho75b, CM75, Cho77b, Cho78b, Cle75, CR79, Clo79, Coe77, CO77, CH78a, CH78b, CH78d, CH78c, CH78e, Com76, Con79f, Con76b, CB75, Con75b, Cor79a, Cor79b, Cou78, Cou79, Cou75, Cow79, Cré76, CL76b, CDI78, Cup77, Dab79, DC75a]. French [DC75b, DCHRH78, DC78, Dam78, Dar77, Deh76, Del75, Del77a, DB77, Del77b, Del77e, Del77c, Del77h, Del77f, Del77g, DM75b, Del77i, Del77j, DMY78, DM78b, Del79, Dem78, Dem75, Dem76a, Dem77a, Dem77c, Dem77b, DT77, Des79a, Des79b, Dix75, Dix79, Dlo77, Dlo78a, Dlo78b, DDC79, Duf75a, Duf79a, Duf79b, Duv77, Duv78, Eis77b, EM77, EW77, El 78, Éli79a, Éli79b, ÉS79, Erm76, Erm78, Eym75, ER75, ET79, FM79, Far75a, Far75b, Far75c, Fed78, FCG75, Fer75, Fer76a, Fer77, FdlP76, Fd76, Fey78b, Fey78a, Fey79a, FdlP79, Fey79b, Fd79, Fit78, FS78b, Foa77c, For76b, For78, Fou75, FP78l, Fou79, Gab75b, GS79a, Gan75, Gap76, GMP78, Gat77, Gau75, GM77o, Geo78b, Gér78, GS79b, Gér75b, Gér75c, Gér75d, Gér75a, Gér75e, Gér75f]. French [Gér75g, Gér75h, Gér79b, Ger75i, GB79, GR77, Gil78b, Gil78c, Gil78a, Gil78d, Gil78f, Gil78e, Gil78g, Gil78h, Gil78i, Gil78j, Gol78a, Gos78, GP79, Gra76b, Gra77, Gra78b, Gri77, GD77b, Gro77b, Gro77d, Gro77c, GD77a, Gro77e, Gro77f, Gru77, GP78b, Gue75, Gui76a, Gui77a, Gui79a, Gui79b, Gui76b, GKR77a, GKR77b, GKR77d, GKR77f, GKR77g, GKR77c, Hae75, HL75, Han76a, Han76b, HL78, Har75b, Hau78b, Hec76, Hec77, Hec78, Hei78, Hel78a, Hen75a, Hen75c, Hen75b, Hen79b, HT75b, HT75a, Her75a, Her75b, Her77b, Her78b, Hes78, Hey78, Hir76b, Hir78a, HT78, Hir79b, HJLN77, Hor77, Hou77, How77, Hue77, Ill76, Ill77a, Imb75, Jac77a, JM76, Jac78b, Jac79c, Jac79g, Jac79d, Jac79f, Jac79b, Jac79e, Jac79h, Jac79i, Jac79k, Jac79j, Jac79l]. French [Jac79m, Jac79n, Jac79o, Jac79p, Jac79q, Jac79r, Jen78, JY78, Joh79q, Jol78, Jou77b, Jou77a, Jou77c, Jou79b, Jou79c, Jou79e, Jou79d, Jou79f, Jou79g, Jou79a, Kah76a, Kaj75, Kan75a, Kan75b, Kar79, KV77, KS78b, Ker75f, Kil78, Kis77a, Kis77b, Kis76a, Kis76b, Kla79, Kôn76, Kor76, Kos75b, Kos76c, Krá76b, Kre75a, Kré75b, Kré76a, Kré77, Kré78a, Kré78b, Kui79, KvQ75, LL76, Lac78, Lac79, LR78a, Lal77b, Lal77a, Lan78, Lap79, Las75, Las78, Las76a, Las77, Las76b, Lat79, Lau75c, dSLM75a, dSLM75b, dSLM75c, dSLM75d, dSLM75f, dSLM75g, dSLM75e, Le 78, Le 76, Le 79, Le 75, Le 77, Led76, Leg75, Lel75a, Lel77b, Lel78a, Lel78b, Lem77, Lem79, Len78, Lép78c, Lép78b, Lép78d, Ler75, Ler78, LS76, Les78, Let76, Let79, LBL78, Lév77, Lig77b, Lio77a, Lio77b]. French [Lod76, Lod79, Loe75, Loh75, Lum76c, Lum76a, Lum76b, Lum78, LP79b, LV75, Mal75a, Mal79b, Mal79a, MM75a, Mal77, Mal78a, Mal78b, Mal79c, Mar79a, Mar75, Mar76d, MD78, Mat75, Mat77b, Mat77h, Mat77c, Mat77d, Mat77e, Mat77f, Mat77g, Mat78b, Maz76, Maz77a, MS76c, MvM77, Meb75, Meb78, MN75a, Mém78, MS79b, MR77m, Mét78b, Mey75b, Mey75c, Mey76a, Mey76d, Mey76e, Mey76f, Mey76i, Mey76g, Mey76c, Mey77a, Mey77b, Mey77c, Mey78, Mey79, Mic76a, Mok76, Mok78b, Mok78a, Mok78c, Mok79, Mol75, Mol77q, Mol77r, Mol77t, Mol77s, Mol77u, Mol77v, Mol77z, Mol77-27, Mol77w, Mol77x, Mol77y, Mol77-29, Mol77-28, Mos78b, MT77c, Mul75a, Mul79, Nac78, Nan75, Nev77, Nic75a, Nic75b, Nic77, Nie75a, Nie75b, Nov75, Nov77b, Nov78, Pag79, Paq79a, Paq79b, Par78, Pau78]. French [Per77, Pes76a, Pes76b, Pha75a, Pir78, Pis76, Poé76a, Poé76b, Poé76e, Poi77, Pou79, Pra78, Pra79, PS79, Pro78, Pue78, Que77, Que78, Rab77, Rab78b, RR78, RR79a, Raï75, Raï78, RS75, Ram78b, Ram78a, Ram78c, Ram79b, Ram79a, Rap76, Rau75d, Ray78b, Reb78, Reb79, RS78, Ren79a, Ren77, Ren79b, Rhi75, Ris76, Rit76, Roc79, Roc75a, Rod75, Rod77, Rog75a, Roo78a, Roo78b, Roo78c, Ros77, Rot76a, Rot76b, Rot76c, Rot78, Rou78, Roy76b, Roy76a, Roy75, Roy78, Rub75, Rub79, Ruc78b, Rue77, Rum79, Sab76, Sak77, Sam76, Sch76a, Sch78a, Sch79g, Sch76d, Sch76e, Sch76f, Sch76g, Sch76h, Sch76i, Sch76j, Sch77g, Sch77h, SW78b, Sec75, Ser79a, Ser75a, Ser75d, Ser75e, Ser75h, Ser75i, Ser75j, Ser75k, Ser75l, Ser75m, Ser75n, Ser75o, Ser75p, Ser77, Ser79b]. French [Ser76, Sib76b, Sib77, Sig75, Sig76, Sim78b, Sjö76a, Sjö76b, Sjö78, Sko78, Smy78b, Smy78a, Sou77, Sou78, Spe75a, Spr75, Spr77f, Ste75b, Ste75c, Ste75d, Ste76b, Ste76a, Ste78e, Str78, Str79d, Str79c, Sub75, Suz78a, Suz78b, Szp76, Tak75a, Tak79, Tar78, Tei75, Tem76a, Thi77, Tho79a, Tho79c, Tik79, Tit77, Tit78, Tou76, Tua78, Tur75, Vai78, Ver76, Ver77, Ver79a, VR75, Ver75, Vie77, Vie78b, Vie78c, Vie78d, Vie78e, Vig77a, Vig77b, VVT77, Voi79, Wag78, Wal76, Wal77b, Wal77a, Web77, Web78d, Wei75b, WK78, WK79, Yam78a, Yam78b, Yam79, Yoe79, Yor76, YM76, Yor77, Yor78a, Yor78b, Yor79, Zan77, de 78, dM75a, dV75a, dlH79c, dlH79a, dlH79b]. Frequency [Lam78h]. Fresnel [AHK76f, AHK76g]. friendly [Liu79b]. Frobenius [Bou76b, Fai79a, Hou77, Kup75, LT76a, Ram79b, Ser75p]. Front [Ser75f, Ser75g]. Frontière [Bro78b, Lio77b, Bai76, Bon78c, Dem77b, Pue78]. frontières [Le 79, Par78]. Frost [APF76]. Fubini [KS76]. Fuchs [Hae75, Kla79, MT77a, Suz79]. Fuchsian [Mar76b, Trè75]. Fuglede [Wei78b]. Fuks [Sch78g, CT78b]. Fuller [Azu79]. Fully [Bea79a, War79b]. Function [Aki78c, AT75a, BK77a, BK77h, BK77i, BK77g, Beh79e, Bin75f, FJ75l, GW79d, Ham75b, KZ76, Sar77, BY77, Bin75h, Bom76, Bru78g, Cla79c, CT78b, DG76c, EGS75, ES77, Far76i, GH76, Hay76, Hej76c, HW79a, KL77, Kue78, MS78c, MG79, Pir77, Roc77, RY77, Sai79a, SNWC77b, SNWC77a, SNWC77g, Sch78e, Sha76a, Sto75b, Tam78o, MWW76]. Functional [BH76a, Bin75g, Gar76b, GN76, HB79, KRZ77e, Lan76g, Adl79a, Ang79, Böh76b, BHP78b, Bre79d, Bro78a, Cop78b, GM77k, GW76a, Hid78, JS77, KS79b, Lan76h, Meh78b, Noh79b, Nus79, Oli76, Ort78, Wai75, Nas79b, PW79]. functionals [Die75e, Fri75, Gri75, Han75a, Kan78b, Muj78, Roc76, Sch75h, Sch75o]. Functions [Acc75d, AT75a, AT75e, BCD77, BØS76, BK75a, BMS76, Bol75, Bru78f, Fei78k, KV79a, Mas75b, Nam79e, Pfl78, SZ77a, Sch75l, Sch75s, Sen76a, Sen76b, SZ77b, SZ77c, Tam78c, Tam78d, AHK76f, And77k, AT75b, AT75f, BK77e, Ber76a, Ber76f, Bra76a, Bru78a, Bru78c, Bru78d, CW79, Can75b, CV77d, CM77, Cho76, CIS76, Cod76, Cop78c, Cur75a, Cza75, Dam77, Din75, Dre77, DR75d, DR75i, Dur76b, Džu76, Ede76, Eic77, Erd75a, Fal79, Far79, Fek79, Fer78a, Fer76b, FR79g, Fis76g, FJ75c, FJ75t, FJK79, For79a, Fri78c, GR75a, Gam76, Gar77b, GM79a, Gei77, Gla76a, Gla76f, Goo76a, Gór78b, GW79b, GW79j, GW79k, Gro77a, Gro78f, Gro78h, Gro78k, Gro78r, Hay77, Hej76f, Hen79a, Hig75, Kaz76, Keo77, Kom75, Koo77, Kor79, Kud77, Mah76a, Mah76d]. functions [Mah76f, Mah76g, Mas75c, Mat77a, McC77, Mei76, MS75, MT77b, Mor79a, Mor77d, iN77, Pfl76, Phi76, Pin79, SNWC77c, SNWC77d, SNWC77e, SNWC77f, SNWC77h, SNWC77l, SNWC77n, Sch76i, Sch77d, Sch75h, Sch77f, Sch76s, Sel77, Sen76c, Sha77a, Sha77c, Sjö75, Ste78b, Str75, Suf76, SDB75, Tak77a, Tam78f, Tam78h, Tam78n, Tho76b, Tho77, Tou76, Utr79, Wei78b, Wil75b, Zag77b, Zie79a, dBS76, tM76]. functor [DS78d, FK79e, Fro79b, PS78]. Functors [Mic78g, BH76c, Brü76b, Cig76, CN75a, CN75b, Ekl75, Gre79, Gru75c, Har79b, Hof76a, Lau79f, May77d, Mic78e, Mic78c, Mic78f, vO75f]. Fundamental [HO79, Uen75i, AEK78i, Bow78, FJ79, Ros75a, Sta77a, BWK77h]. Funktionale [Dri79]. Funktionalen [FM76b, Ess76]. Funktionalinterpretation [DV75b]. Funktionen [Dri79, JLR76, RR75d, RR75g, Bla76i, Bau76, Bin76, FM76b, Mic76c, RR75c]. Funktionenraum [Cro76]. Furstenberg [Gla76c, Tho79c]. Further [Bra79h, Tho77, Geo79g, Sch79e]. fuzzy [Got76].

G [Bar79d, Har79a]. G. [Ben77b, Con75a, Pra78, Spr75, Wal77a]. Gabriel [HT79]. Gabrielov [Mac78]. gaga [Ram78b, Ram78a, Ram78c]. gain [Pre76k]. galbés [Tur75]. Galerkin [Dou77b, DDP78, Jer76b, MG78]. Galois [Buh78c, Buh78g, Gro78j, NT79b, Rib77]. game [FR79e]. games [FR79d, KK76, Sre78]. gamma [Vig77a]. ganzzahligen [vG76]. Gaps [Mol77i]. garden [Hig75]. Garland [Bor75a]. garnir [Jam78h]. Garsia [Cho75a, Cho75b]. Gastrulation [Zee76d]. Gauge [May77l, Gar77a, Hal77, May78b, Ne’78, Wu78]. gauge-symmetry [Gar77a]. Gauss [Bal78, Bor78b, Küh79]. Gaussian [Byc78, Byc79, Car78b, Car79d, CBR76, CIS76, Gua78, Hah79, Hel78b, Hit78, HC78, Kuo75a, Kuo75b, Kuo75c]. gaussien [Tua78]. gaussienne [Nov77b, Nov78]. gaussiennes [Fer75, Kôn76]. gaussiens [Mey76i, Ruc78b, Web77, Fer76a]. GB [Dud79, Rat78g]. GB-conjecture [Rat78g]. Gebiete [Pfl75c]. geblockte [Jur78e]. Gehring [RR75b]. Gelchmässiger [Som76]. Gelfand [Sch78g, Ste78h, CT78b, Gui77a, Hae75, Mac78]. Gelfond [Bro79b]. Genera [Arn77, Tay79]. General [BR75, Bow75d, DS78d, Joh79f, Klo78, Tay78b, TY78, Won79a, CL76a, BZ76, Bin75b, Bin75h, Bre77c, Coh79b, EL77c, Fen75, FJ75g, Gal76d, Gri76b, Hac78a, HvW79, Jen77b, KRZ77h, Laz75e, Mac75a, MBY79, McC78c, Per78c, Rei77b, Szc77b, Van78c, Wid78, dA79, Des79a, Nov77a]. Générale [Jac79o, EM77, EW77]. Générales [Mol77w, Mol77x, Mol77y, Aup79e, Mey76e, Sig75]. Généralisation [Kan75b, Sve77, HL78]. Généralisations [Mol77q, Ger75i]. généralise [Lum76c]. Generalised [MG78, MS79a, Seb79]. généralisées [GKR77c]. généralisés [Paq79a]. Généralités [dSLM75b, Mat77d]. Generalities [Dub79c, FGR75d, vO75b]. generalization [Ded77, Ess75e, Kre75c, Oxl79, Tam78e, Lub79a]. Generalizations [Beh79f, BJM78a, Bru78e, Hal79i, Lau75a, May79, Sne75]. Generalized [Blo79c, Bru78h, DH77, Kom75, Sje78, Tam78f, Ber79d, Bom77, CCMS75, Cur75b, Džu76, Erd75a, Eva77, FI77, GM77h, Gla76d, Gla76e, Hah79, Hof76a, KT75e, KY79b, KV79a, KK76, Lau78, Nas79a, PR77b, Ray78a, Rog76a, Sch76s, Slo76, Str75, Tho77, Fri78a]. generalizing [Tam78i]. Generated [Bra79d, Bro75d, Cor79f, Die75g, Frö76, KM77, Mah79, Sha79c, Smi79, Sni79]. générateur [Fey79b, Mok78b]. générateurs [Rot78]. Generating [Gro78k, Tan78, Haz77e, Wei78b]. Generation [Bro75c, Mey76b, CS75, Gui77b, War79a, Wil77b]. generations [Lub79a]. Generators [Gre76, Hum79, DGS76o, DGS76t, Gla76b]. généraux [Yor78a]. Generic [Alp79, HW75k, Jac77b, Moi77d, FSW75, FT76, Hir75c, KL79b]. generosity [Sax77]. genetics [AW75, Gra79a, Pel78a, iS78, Saw75, Wei78a]. Genre [GP78b, Des79a]. Gentzen [Pra75]. Gentzen-type [Pra75]. Genuine [CT78c]. Genus [Ish76d, Rei78a, Ish76c, Ish76e, Ish76f, KK79]. Geodätischen [GMK75d]. Geometric [Bla76d, Blo79f, CK75, Dol78, EH76c, Gaw77, GR75b, Joh79g, Kij77, Mil78b, Woo78b, tD79f, Ale75, Bru78s, Czy78, MR77g, Sim77, Sim78a, Śni77, Śni78, Suf77, BM78a, BM78b, GR75b]. Geometrical [BR77c, BRP78, Lic77]. geometrically [AB76]. Geometrie [GKM75, Del77b, Ill77b, Ram78c, Sou78, BR76b, Fou75, Gro77e, Gru77, Ste76a, Tei75, Băn78, Bea78c, Le 78]. geometries [Sri78]. géométrique [And77i, Clo79, Des79a, Sou77, Vie77, BF75c]. Géométriques [Bea78e]. Geometry [Bla76c, Bre75d, Del77b, Die75c, EGS75, Fis76d, JMR75, Kel75c, Løn79b, Ols78b, Woy75, Ati75, Bla77, Ded77, Gar78, GS77, Gre78a, GKM75, Mal75b, Mar79b, Rob75a, San79, Sch76r, Sei75, Ste76a, Ste78g, PdC77]. Geophysikalischer [JLR76]. geordnete [Fel76]. geordneten [Hei76a]. Georgia [MS77]. Germ [Man76c, EL76]. Germ-field [Man76c]. German [AE77, Bau76, BG76b, Bin76, Bla76i, Bra76b, Bro76a, Buc75, Col76c, Cré79, Cro76, DG76a, Dic75, DV75b, Don76, Dri79, Eif76, Ess76, Fel76, Fel75, FM76b, FK79a, FK79c, FK79d, Fri78a, Fri76b, Ger79c, Ger79d, GMK75a, GMK75b, GMK75c, GMK75d, GMK75e, GKM75, GMK75g, GMK75f, GMK75h, Gro76, Gro79d, Gro79a, Gro79e, Gro79f, Gro79g, Gro79c, Häm76, Har75a, Haz77a, Haz77b, Haz77c, Haz77d, Haz77g, Haz77f, Haz77e, Haz79b, Hee76, Hei75, Hei76a, HW75o, HJSW75, HK76, Jan79a, Jan79b, Jan79c, Jan79e, Jan79f, Jan79g, Jan79h, Jan79i, Jeg75, JLR76, Jur78b, Jur78e, Jur78f, Jur78d, Jur78c, Jur78g, Jur78h, Jur78i, Jur78j, Jur78k, Jur78l, Jur78n, Jur78p, Jur78o, Jur78m, Jur78r, Jur78q, Jur78s, Kah76b, Kna79a, Kra75a, Kra75b, Kra75c, Kra75e, Kra75f, Küh79, Kup76, KPP^{+}78a, KPP^{+}78b]. German [KPP^{+}78d, KPP^{+}78h, KPP^{+}78i, KPP^{+}78g, KPP^{+}78e, KPP^{+}78f, Maa75, Man78c, MH75, Mei76, Mei78, Mét77c, Mét77d, Mét77a, Mét77e, Mét77f, Mét77b, Mic76c, Nic78, Oss75, Pfe75, Pfl75b, Pfl75c, Pfl75a, Pfl75d, Pfl75e, RR75a, RR75c, RR75e, RR75f, RR75g, RR75b, Sca75, SS76b, Sch77c, Sch78d, Sch76c, Som76, Str76, Vek76, Vog78, Voi77i, Voi77k, Voi77j, Voi77l, Voi77m, Voi77n, Voi77p, Voi77o, Voi77a, Voi77r, Voi77b, Voi77c, Voi77s, Voi77h, Voi77f, Voi77d, Voi77t, Voi77e, Voi77g, Wec75, vG76]. Germany [AT78, BMS76, DJ79, Hey79, KK78a, MWW76, MS78d, SS76a, Bec79a]. germes [Dem75, Tou76]. germs [dP76b]. Gestalt [GMK75b]. Gevrey [Wag78]. Gewicht [Jan79f, Jan79g]. Gewichten [Jan79h]. Gewissen [Bro76a]. Gibbs [AHK76a, Bow75e, Evs78, Geo79i, Pre76d]. Gibbsian [Ave78]. Gilkey [dV75a]. Girsanov [Kaz78]. given [Bos78, DG77, Hua79, Hub78a, Moi77g, Rei76a, Rob76b, Sch79j, Sim78a]. Glättungssatz [FK79a]. Gleichheit [Oss75]. Gleichung [Hee76]. Gleichungen [Gro76]. Gleichverteilungssatz [Ger79c]. Gleitflugbahnen [Dic75]. Glivenko [Kri76]. Global [Ang79, BF76b, BK76c, Cha79a, Gel76a, Gel76c, Lau79b, Lau79c, Ros79, Ste78i, Ber76f, Bor79c, Len76b, McC78a, Sto75b, GM79d, Hec76]. globale [Gru77, Hec76]. globales [Jou79f]. globaux [BEGS75, Che75a]. Godbillon [Gar78, Her77a]. Gödel [Tak78d]. Golay [AS77]. gons [Law75a]. Gordon [Ree78a, Run76]. Gorenstein [FGR75c, FGR75e, Her79b, Jan78, RR78]. Gottlieb [Lem77]. graceful [Che78a, KTR79]. Graded [Kos77, NV79d, NV79b, NV79c, Ne’77, Orz76, Pas79a]. gradient [Fle76]. Graeco [Pre76a]. grammars [MJ75]. grand [Dar77]. graph [AEK78c, AEK78h, Bou78a, EH79a, Ent78, GHM77, MS79a, Tut75, Wei76c, Wei76e, Wei76n, Wei76u, Wil78a]. Graphic [EH79b, EH79a]. Graphs [AL78, Bos78, God78, Wei76l, Bab78, BC78b, Ber78a, BG78b, CM78a, CW78, Che78a, Coc78, Erd78, FS78a, Fra78a, GM76b, GM76c, Gra75, Gre78b, Gro78a, HRW78, Hog75, HS75c, Imr77, Lit75, Lit76, Lit77b, MHG77, Qui79, Sax77, Wei76b, Wei76m, Wor79]. Grassmann [GH78a, Koo77]. Grassmannians [Aki78d, Jen77a, Jen77d]. grassmanniennes [Bru78r, Las77]. Grauertsche [KPP^{+}78f]. Gravity [Ne’78]. greater [Iva79c, Wei76f, Wei76p]. Green [FdlP76, Fd79, Mic75, SNWC77b, SNWC77a, SNWC77g, Tam78o]. Grenoble [DCV78]. Grenzwertsätze [Cré79]. grid [SM79a]. Griffiths [Le 77, Mor77c, Pop77j]. Großen [GKM75]. Grossissement [JY78, Yor78a, DM78b]. Grothendieck [Ill77b, Dre75, Fie78, Kan75a, Kan75b, Lub79b, MR77i, Sch76a]. Grötzsch [Caz79a]. ground [AHK76b]. Group [ACH77f, AHW77, EH76d, Gel76g, Har76a, McG78, Pas79e, Pas79d, Poé76b, Pop77e, Sat79h, SV75d, Yan77h, Yan77i, Yan77t, Ass79, Bae78b, Bot78, Bow78, CN77, Cor79i, DeM76b, Duf75b, Ekl77, Fef75a, Fer78b, Gel76a, Gla76c, Gro78j, GL75, HO79, HM79b, HvW79, Haz79a, Hil78, Hum79, Jam78t, Jan75, Jen77g, Joh79l, Jon79, Kar77, Ker75a, Lev79b, Lin79b, Mac79, Moi77f, Mul79, Ne’78, Olv79, Par76, Pas79a, Pas79g, Ply79, Rog75b, Rou79, Sat79d, Sim78c, Sim78e, Sni79, Sri78, Sta79b, Sto78a, Tay77a, Wal78c, Wat78a, Yan77a, Yan77c, Yan77d, Yan77f, Yan77k, Yan77l, Yan77o, Yan77n, Yan77p, Yan77q, Zel76a, tD79b, tD79g, tDP79, vO75a, BCN77]. group-actions [Bot78]. group-theoretic [Sim78e]. Groupe [Cha77, Foa77a, GKR77d, Hey78, Poé76b, Ban75, Ber76c, Bop79, Cat75, Éli79a, ET79, Gér75g, Gue75, Her75b, LL76, Ler75, Lod76, Mul79, RS75, dlH79b]. Groupes [BWK77d, BKW77, BWK77e, BWK77f, BWK77m, ETFS75, ETFS79, Gér75b, Gér75c, Gér75d, Gér75a, Gro77f, GKR77a, GKR77f, GKR77g, GKR77e, Hey78, Lio77a, Pag79, Roy78, Tit78, Voi79, Ars79, Ave76, Bam79, BBL77, Ber76b, BWK77o, Bor75a, Bou79a, Bré76c, CO79a, Car76b, Car77b, Cow79, Cré76, CL76b, Deh76, Del79, Duf75a, Duf79b, Eym75, Far75c, Fey78a, GS79a, Gui79a, How77, Loe75, LP79b, Mat77b, Mat77f, Mey76g, Mol75, Nan75, Nic77, Paq79a, Paq79b, PS79, Raï75, Ren79a, Rot76c, Rot78, Roy75, Sab76, Sak77, Ser77, Spr77f, Tak75a, Tit77, Ver75]. Groupoids [HK79a]. Groups [Acc75d, BLR75d, FP78f, Gér75b, Goo76d, Hey79, Hum76c, Jam78n, Ker75e, Laz75a, MHG77, RR79b, Sri79h, Tay77a, Var77a, War76b, War76h, Zel76b, tD79l, Bar79a, BW77a, Ber77a, Bie79, Bro76b, Bro78d, Bro78e, But75, Byc79, CM78b, Car79c, CM79, Cha79a, Cha76a, Che76e, Chi76b, Cho78a, Con75a, Con77, Con79a, CM76, Coo77, Cré76, Cur75c, DR75i, Dup78b, Dup78d, Dwy78a, Ell79, Far76g, FP76, FP78e, FP78g, FP78k, Fra78a, Fre78, Fri76a, Gag77, Gel76f, Gia78, God78, Goo76f, GP78a, Gus76b, Har78c, HP76b, HL77, Hil77a, Hil75, Hil77b, Hir79a, Hir78b, Hum76f, Jac77a, JR77, Jan78, Joh79g, Kap78, KZ77, KOS76a, Kov78, Kru79, Lau77, Laz75b, Laz75c, Laz75f, Lev78, Lie77, Lin79a, LS79b, LR79, MS76a, Mag79a, MP79b, Mar76b, Mek77, Mic75, Mic77]. groups [MT76c, Muk79, New77, Obe78, Orz76, Pas79e, Pas79h, Pop77k, RS76, Rei75, Rei78b, Rei78c, Rei78a, Ric77, Rob75a, Rog77b, Rog79b, Sch77e, Sch78f, Sha76b, She75, SS78b, Smi77, Śni77, Spr77c, Sri79i, Sta77a, SV75a, Var77c, Voi77r, Wam77, Wan75, War76a, War76d, War76i, War76l, War77b, Wei76d, Wil77c, Yan77g, Yan77s]. Growth [Ham75c, AW78, Dam77, Gla76b, Hay77, Kue79, Pfl78]. Grundbegriffe [Gro79a]. Grundlegende [Mét77d]. Grundtatsachen [Jur78i]. Gruppen [Bin76, Kra75d, Cré79, Kra75b, Voi77i, Voi77n, Voi77r, Voi77s, Voi77h, Voi77f]. Guide [Dev77c, Ric77]. guided [Wil78a].

h [Lod76, Bol75]. H. [Bor75a, Gér79b, Ker75f, Mal78b, Wil76a]. Haag [Roy76b]. Haar [Ber79a, Far76g, Kah76b, Zie79e]. Hadamard [CMW78, SNWC77g, Seb79]. Hahn [Rob77a]. Halbgruppen [Haz77c, Bur75b, Haz77b, Haz77c, Haz79b]. half [Bla75, KV79a, Raw78, Sil76d, Sim75b]. half-density [Bla75]. half-forms [Raw78, Sim75b]. Hall [LW75, Mor77d, War76c]. halo [dG75g]. Hamilton [Cau76c]. Hamiltonian [Adl79a, Fei78d, Fei78k]. hamiltoniens [Cha76e]. Hamiltons [Ger79a]. Hammerstein [Bro75h]. hand [HB79]. Hans [Nir75]. HAO [dSLM75g]. Hardy [BK77a, BK77c, BK77f, LL76, dG75d]. Harish [Car75, Sri79f]. Harish-Chandra [Car75, Sri79f]. Harmonic [AT75a, Bre77d, CDV75, CV77a, CV79, Fug79, Gla76f, Ham75d, Ham75e, SNWC77e, SNWC77h, SNWC77i, Var77a, AHK76b, AHK76a, BK77e, GH79, Gla76a, GW79j, Hay76, Ike78, Kor79, Rus77, Sch77d]. Harmonics [Koo77]. Harmonique [CDV75, CV77a, CV79, ETFS75, ETFS79, Mat77c, Cle75, Gui79b, Kor76, Mat75, Mat77b, Mat77h, Tak79]. harmoniques [Ava77, Ave76, Éli79a, LL76, Sjö76b]. Harris [Sza75]. Hartogs [Gam76].

Haus [Har79a]. Hausdorff [Kah76b, CV77f, Gal76b, Pol79, Sch76b]. Haven [Ros75b]. having [AGR75, BC78b, Dek79, Ørn77, Wei76m]. heat [Ati75, BTW75e, Col76e, Ham75f]. heating [vH79b]. heavy [AH76]. Heawood [Fou79]. Hecke [Li79]. Heidelberg [GR79b]. height [Dub76a, SW78l]. Heins [Ess75e]. Heisenberg [Pow78b]. Heisuke [Tei75]. Held [AHW77, BCD77, Bar77b, BØS76, BK76a, BH76b, BR77c, BMS76, BCN77, BOS75, BGS78, CW76a, Cha78a, CUW75, CS76d, DKT78, ES76, GM77n, GN76, GDMW76, Hil76, JMR75, Kel75c, KZ76, Lit77a, Man75d, MWW76, Nat77a, Par75a, Pin75a, Ros75b, SS76a, SZ77a, Ste76c, SW75d, Tem76c, Wat78b, Zel76b, BGM77, BS77, DM79b, Eve75a, GR79b, GR75b, Mil78b, MS77, PdC77, Sch78g, Wuy79b]. Hélices [dSLM75c, dSLM75d]. Hellsten [Ess75f]. Helmholtz [Wid78]. Helsinki [LS79a]. Henselsche [KPP^{+}78b]. Hérault [Fel77c]. héréditaires [GP79]. Hereditarily [KR79]. hereditary [MM79a]. Herman [Del77e, Ros77]. Hermite [DDC79, Kra79, Möl77-30]. Hermitian [Kea78]. herrühren [Voi77o]. Hessian [GW79e]. heteroclinic [Dev78]. hexagon [Bad77]. Heyting [Gra79c, MR77b]. Heyting-valued [Gra79c]. Hierarchies [Glo78, KM77]. Hierarchy [LSZ77, MSZ76, HW75k]. high [And79a]. Higher [Dal76, Dus79, Gra76k, Jak79, Jen77c, KRZ77f, Mol77f, Tol75e, GH76, Lev78, Mol77g, SV79, SW75c, Suf77, Tol75c, Mol77u, MS78d]. highest [Jan79f, Jan79g, Wal75a]. highly [Lio76]. Hilbert [AHK76f, Ama76b, BH78, Bar79b, Blu79, Boc78, Bon75, Cha75b, Coh77, Del76, DR75c, GK76a, Haa79, Hah79, HH77e, HH77f, Hir77, Ist78, Kaj75, Mas76, McK75, Mét78a, Nor78b, Rab77, Rab78b, Roq75, Suz79, Wei78b, dF75]. Hilbertian [BGJN76, Sat79a]. Hilbertsche [Mét77a]. Hilfe [Mic76c, Sch76c]. Hill [Raţ79]. Hironaka [Tei75]. histoire [Imb75]. Historic [Gro78l]. historical [GM77a, SNWC77k]. Historische [Jur78j]. history [AHK76c, Dav78g, Gra79b, Imb75, Ros75a]. höchsten [Jan79f, Jan79g, Jan79h]. HOD [Rog76b, Rog77c, Ště78d]. Hodge [LBL78, MM75a, Pro78]. hodograph [BS76a, HS75a]. holder [Aki78i, BF76a]. Holley [Bru76d, Han76a]. Holomorphe [Erm76, Bon75, Ste75b, VR75]. holomorphes [And75, CW77b, Coe77, Dem78, Ele78, Kré78b, LP79b, Paq79b, Pou79, Sch77h, Suz78b]. Holomorphic [Jen77e, Jen77d, Rud77b, BK77e, Ber76f, CW77c, Din75, Far79, Gér77, Hol78a, Nam79c, Ste78c, Suf77]. holomorphie [Kré77, Kré78a]. Holomorphiegebiete [Pfl75b, Pfl75c]. holomorphy [Pfl75c, Pfl78, Pfl75b]. holonômte [Hec78]. holonomy [Pla75]. hom [Che79]. hom-sets [Che79]. homeomorphisms [Alp79, Bru78o, Cha75b, DGS76f, Han75b, Has78a, Sch78c]. homogène [LS76, MM75a, Rot76c]. Homogénéisation [Duv78]. Homogeneity [Bra79b, Stø78b, Eps79f]. Homogenen [Häm76]. Homogeneous [Jen77c, Bra79a, CS76b, For76a, Geo79g, Geo79h, Goo76e, Hay78, Kur78, Sch77b, Shn78]. homogènes [Far75c, Gui76b, Hen79b, PS79]. homogenization [Bab79]. Homological [FGR75f, FGR75i, Fri76a, Bro78e, EH76e, Sje78]. Homologie [Băn78, Avr79]. homologique [And79b]. Homologiques [Ser75d, Ser75e, Pau78]. Homology [CLM76, FP76, FP78f, FP78g, Hau76, Shu75, Bru78q, Coh76b, Coh76a, Dou77a, FP78e, GL75, KS78a, Kan78a, Lan79a, LS77, Lub79a, Mar79c, May76a, May76b, Seg77, Sha76b, tD79d]. homomorphism [Dup78c, Dup78d, Dup78j, HO79, Kna79b, tD79j]. Homomorphismes [Avr79]. homomorphisms [Bru78s, CM78a, Jam78o]. Homotopic [Cha75b]. Homotopie [Lod79]. homotopies [Eav79, HMRS78]. Homotopy [BM78a, BM78b, Bau77e, BF78, CS76a, EH76c, Joh79h, Joh79i, LY79, Liu79a, tD79g, tD79h, AH79, Bau77f, CMPY79, Cur75c, Dwy78a, EH76e, EH76h, EH76j, Gon78, Hil77b, Joh79a, Joh79k, KS79a, Lad76b, Liu78a, Liu78b, Mah79, Mor77c, Mun79, Sta78c, tD79e, tDP79]. homotopy-associative [Gon78]. Homotopy-equivalent [tD79h]. homotopy-obstruction [Sta78c]. honor [Mil78b]. hook [Jam78i]. Hooks [Jam78j]. Hopf [CY78, Gei77, Han78c, Yan77b, Yan77d, Yan77f, Yan77j, Yan77p, Yan77t]. horizontal [Cor75]. Horn [Kea75]. Horrocks [Lam78c]. Host [Cli75]. Host-parasite [Cli75]. Houzel [Ill77b]. Hsiang [Bar79c]. Hubbard [Le 75]. Hughes [PR77c]. hull [Cop78e]. hulls [Fit75]. Hunt [Ber79a, CL75b]. Hurewicz [DS78l]. Hybrid [Bre77o, OL76, OL77]. hydraulique [Bai76]. hydrodynamic [vH79a]. Hyper [BWK77f]. Hyper-Archimediens [BWK77f]. Hyperbolic [Sil76d, ACH77e, BTW75c, Cos75, GW79b, Her75d, HPS77j, HPS77l, KK75, Kol77, Noh79a, Ril79, Shi75, vdB79]. hyperbolicity [GW79h, HPS77h, HPS77m]. hyperbolique [Vai78]. hyperboliques [GM77o, Sub75]. hyperboloïde [Far75b]. hypercomplex [Ede76]. hyperconvexes [Erm78]. hyperelliptic [LL78]. hyperfinite [Ply79]. Hyperfonctions [Bon78b, CCP75, Cho77b]. Hyperfunctions [Pha75b]. hypergeometric [Lov75]. hypergraphs [Ber78a]. hypergroupes [Spe75a]. hyperharmoniques [Smy78b]. Hypermetric [Kel75a]. hyperstereology [Lau75b]. hypersurface [Meb78]. hypoelliptic [MS78a, MS79c]. Hypoellipticité [BCH78]. Hypoellipticity [Roc75b, Roc75c, Roc75d]. hypoelliptique [CDI78]. hypoelliptique-analytiques [CDI78]. hypotheses [Ish76i]. hypothesis [Guz77]. hysteresis [BN77].

I. [RR79b]. ice [Rau75b]. Icosahedral [Buh78c, Buh77, Buh78b, Buh78e]. ideal [Ari79, BF76a, Eva77, Fuj79, Hoc79c, Möl77-30, Pas79f, Pas79i, Rog75b]. ideal-theoretic [Möl77-30]. Idealbasen [KPP^{+}78f]. Ideals [Beh79h, Che75b, Mic78g, Pas79d, CF79, Mat77a, Mic78d, Mic78c, Mic78k, Pas79g, SV75c]. Idealtheorie [KPP^{+}78i]. Idéaux [Dix75, Ser75h, Ser75i]. Idempotent [War76d]. idempotenten [Haz77a]. idempotents [How77]. Identification [Wei76l, Ban79c, Wei76u]. identité [Cor79a]. Identités [Dem77a]. Identities [DW79a, DW79b, DW79d, Jac78a, Mor79b, Oma75, Sch77e]. identity [Cor79a, Obe77, Pea76]. ihrer [Dri79]. II [AD78b, BM78b, Bec79a, BRP78, Erd75b, ETFS79, Nor75, Que78, SD77, Wal77b, Acc75b, AR75, Cho77a, Gla76e, Gra76k, Ker75e, Kle77b, Kré75b, MS79c, Mic75, Ply79, Que78, Ram78b, Roo78b, Roy76b, Sch77f, Sło78, SV75e]. III [Nor78c, PdC77, SW75d, Acc75c, Bon75, CH78a, Com76, Erd77, HRW78, Kle77c, Kur75, Sai78, Sma77, Stø78b, SV75f, Tak77b]. Illinois [Nat79]. im [Cro76, Don76, GKM75]. Images [Lum76c]. imaginaires [Bai77a]. Imaginary [LT76b, MT77b]. imbeddable [Hei76b]. imbedded [Gua78]. Imbeddings [Gro78a, Bou78a]. imcompressible [ADB75]. ImJ [FP78h, FP78j]. immersed [Kos79]. Immersing [GH78b]. Immersions [Aki78e, GHM78, Aki78d, CdC77, Era78]. immigration [Kaw76]. immunological [GW76a]. Implementation [Mor78c]. Implications [And77j]. Implicit [Mos76a, Tak76b, But76, Tak76a]. imply [Hah76]. Impredicativity [Fef75a]. Imprimitive [Wei76i]. imprimitivity [Fel77a, Fel77d]. improve [Man79]. Improved [dP79]. impulse [Car76a]. impulsionnel [Mos78b]. IMU [GN76]. in-core [Rei76b]. inaccessible [McA78]. incidence [Cum76, Leh76]. inclusion [Pom76]. incompleteness [Alo79]. incompressible [BF76a, Cha76d, Gui77c, Hem75, Tem76b]. incorporation [Cox78]. Increasing [Get75d, Hof79a, Tru78]. increment [Kue79]. increments [CC76, Gus76a, Zab75]. Indecomposable [Ham75a, Lou75, AF78, Naz75]. indécomposables [Gab75b]. indéfinies [Vig77b]. indefinite [Fle76, vR79]. Indépendance [Nie75a]. indépendantes [CG79a]. independants [Sid79]. independence [BR75, Hel79b, Iag75b, Mah76g, Sol78b, Sri78]. Independent [Geo79f, Bou78a, Gal76e, Gie76, GMZ79, Gus76a, Kue79, Ros78e, Zab75]. Indescribability [Dev75]. Index [Ano75, KRZ77d, Sha78d, GM77g, GM77j, Gup78, Hen79a, Het79, Hua79, Jan75, KRZ77c, Kni78, Ols78a, Sha77a, Sha78b, Sin77, Wat78a]. Indexed [JPR^{+}78, Woo78a]. indexées [Cai79]. Indiana [Cha78a]. Indices [Lio77a, BS78, EG75, KRZ77h, McA76, RB75]. Induced [Fel77c, Fel77d, Wil78b, dlR75]. inducing [Fel77a]. Induction [tD79i, ACH77a, Poh75, Ptá77]. Inductive [AEK78j, Cen76, Mos78a, AEK78t, Mar79a, Poh75]. induisant [Zan77]. Induzierte [Voi77n, Voi77p, Voi77o]. induzierten [Voi77l]. induzierter [Voi77j]. inégalité [Lép78d]. Inégalités [Mey76a, Cho75a, Han76a, Mét78b, Mey76d, Mey76e]. Inequalities [Ask75, Ale75, BZ78, BL75, BS76a, Caz79a, Che78b, Dei77e, Eve75b, Geo79a, Gro78b, Ive78, Lio75a, Lio76, Mos76a, Mos77a, Pow78b]. inequality [Ess75f, Ess75a, Eve76, Hol76, Pel78b, PR77b, SS76d, Tam78i, Tam78l, Tam78m]. inéquation [Mos78b]. Inequations [Bai77b, Bai76, Bai77b]. inference [MR77a]. inférieure [Anc78]. inférieurement [Bad78]. infini [Ava77]. infinie [Ari79, Bon75, Kré77, Kré78a, Kré78b, Las78, Nov77b, Nov78]. infinis [Sak77]. Infinitary [Kle77h, Kle77i, Kue75b, Bie77, Glo75, Glo78, Gre75c, Kue75a]. Infinite [FP78f, FP78h, Gal76h, Haz79a, HW75l, SW75c, Aar79, BY77, Car79a, Cha78b, Cun75, Cza75, Din75, EW76, EH76d, Fat75, FW78, Gra78a, Hal79a, Has78a, Hay78, Lig77a, May77j, Mét78a, MS78b, Naz75, Sza75, Wal78a, ZKB79, Las75]. infinite-dimensional [Car79a, Gra78a]. infinite-quantifier [Cun75]. Infinitely [BF79a, Fal79, Čud77, Kru76, Maa75]. infinitesimal [Fey79b, Mok78b]. infinitesimalen [Dri79]. infinitesimaux [Rot78, Voi79]. infinitessimal [Mañ75a]. influence [Eas78, SB77]. Information [Ano75, DGS76g, Bak76, BBDB^{+}78, Cir78, Par79a, Pre76k]. Inheritness [Mus76]. Inhomogeneous [BK76f, BP79a]. Inhomogènes [Lio77b]. inhomogenity [Fra77b]. Initial [BTW75b, BTW75f, Her79a, KP77, Dah76, Kra79]. initiales [BLP77]. Initiation [Eym75, dlH79a]. injectifs [BBL77, Ren79a]. Injective [Ban79a, Fai79b, vdK76, BP78, Kai76]. injectives [Jat76]. injectivity [Rei77a]. inner [Bro78e, Ker75b, NS76, Oma75, Ště78d, vR79]. Innerproducts [ACH77f]. inseparable [Yan77m]. insights [Ben76a]. Instantons [DM79a]. Institute [Cro75b, DM75c, Lit77a, MOP75, NR78b]. Instituto [PdC77]. Integer [FR79e, Rob76b, FJ75j, Hua79]. integers [Che75b, Raz77, Sto79]. intégrabilité [MS79b]. Integrability [NT79e, Bor79a]. Integrable [LR78b, McK79, Eas76, Sło78, CW77b, Duf79b]. Intégrables [Bin75a, Sec75]. Integral [BDE^{+}77g, Cha75c, Col76d, Eve75b, Hal78b, Hal78c, LS79a, Mét77b, Rei78b, Rei78c, Rei78a, RR79b, Roc76, Rog79b, Tho78, AHK76f, AHK76g, AHK76d, AL79, Bak78b, Bar79b, BDE^{+}77f, BP79a, Bro75h, Bru76e, But75, CV77b, Cla79g, Cus79, Don75, Gro79h, Har78b, Lak79a, Lev79a, Lov75, MW79, MS75, Rei79, Rob79b, Sch77f, Sto78a, Wal78d, Mét77a, RR79b]. Intégrale [Jac79n, Bis78b, Cai78, FP78l]. intégralement [Que77]. Intégrales [Cau76a, Jac79k, Jac79j, Kre75a, Lio77a, Mul75a, Mul79, Bes75a, CM75, Len78, Roo78b, Suz78a, Yor79]. Integralgleichung [Mei78]. Integralgleichungen [Häm76]. Integrals [AHK76j, Gui75, dG75c, AHK76c, AHK76e, Ded77, Erd75a, Fei78c, Gaw77, Goo76e, Kom75, Mul79, Sha77b, Var78, dG76b, vdAS79, Kre75a]. integrand [CV77e]. integrands [Roc76]. Integrated [Que78]. Integrating [Hal76b]. Integration [AE77, GT77, Mét77c, Ask75, Har76b, Mas76, Mét78a, Wed78, Con79f, Mét77f]. intégrés [Que78]. integro [BM79a, Gra76a]. integro-differential [BM79a]. integrodifferential [CG79b, Han79a, Jor79, Lon79a]. Intensional [vK75]. Intensionale [DV75b]. intensity [YS79]. Interacting [DKT78, Geo79i, Gri79a, Lig77a]. Interaction [DKT78, Geo79c, Geo79d, Kry78, Wah76]. interactions [Cir78, GPR78, Vas78, Yor79]. interchange [Sax77]. interesting [Rog75b]. intérieur [Bar77a]. Interior [DN77]. Interlaced [KTR79]. Interleaving [Mul79]. intermédiaires [Lac78]. intermediären [SS76b]. Intermittent [Man76a]. Internal [Che79, Joh75b, Osi75b, Ros78a]. International [Bec76b, Bec79a, BK75a, BMS76, Cha75c, CS76d, DM75c, DG75j, GLL79, HP79, HS78, MWW76, MOP75, Ode75, Ros75b, SS76a, SZ77b, SZ77c]. interpolating [Ama76b, FJ75o]. Interpolation [Bré75e, Sch76c, Sch77d, BC77a, FJ75j, GP76, Gar77b, Möl77-30, SS77a, Sco79, Wer79, dBS76, dB76]. interpolators [Mar79d]. Interpretation [MR77j, Sou77, Bla76d, Tai75, Tur77]. interrelation [New76a]. Intersecting [Bre75b]. Intersection [Fox79, Jur78a, Las76a, Eva77, Gui75, Las76a]. Intersectional [Pet75]. Intersections [Que78, Bro79a, Fer79, FM78, RT78, Yan77l]. Intertwining [Bla75, ACF79]. interval [dG75e]. Intervall [Nic78]. Intervall-Mathematik [Nic78]. intervals [FJ75o, Wal78a]. Intrinsically [DGS76m]. Introducting [FT79]. Introduction [AEK78k, AHK76h, ACH77d, BF75c, Bib79b, BK76d, Bow75f, Bra79a, Bra79j, Bre77j, Buh78d, Cau76d, CCP75, Che76i, Con79c, Dav78d, Dei77f, DGS76h, Dub79d, DS78e, DS78j, EH76f, EL77d, Far76e, Fei78e, Fel77e, FR79f, FP78i, FJ75h, Foa77b, GM77i, Gal76c, Gér75e, GWdPL76, GW79f, Gri79d, GKR77b, Hal79d, HPS77f, HW75m, HH77g, Hum76a, IP76b, Jac75c, Joh79j, KRZ77g, Ker75d, Lan76d, Lau79d, Law75b, LE75, MR77k, Mat77e, May77g, Mic78h, Miu76b, Moi77e, Mol77r, Pet76c, Pre76e, Rat78i, Ree76c, Rey78a, Rob78, Roc75e, Sai79b, Sch79b, Sch79f, Sch76q, Sha78d, SW78h, Spr77d, Sto77d, Tam78g, Tol75b, Vie78e, Voi77q, Wil75c, Wu78, Eps79g, Eps79h, Iag75a, Jac75e, Rau75a, Seg78, Tay78b, Bre77j, Gér75e, Mat77e, Mol77r, Vie78e]. Introductory [Wal75d]. Intuitionistic [LEV75, Gra79c]. intuitionistically [Fri78c]. Intuitionistichen [Oss75]. Invariance [Joh79k, Ros76a, DFF78, Dei77e, Ne’77, Phi76, Phi79, SS76e]. Invariant [Ama79, DGS76i, EL77e, Far76f, Fly79, HPS77g, Kow75, Mis79b, Pre76f, Spr77e, Var77c, Var77b, AHK76i, Bar79a, BS78, BJM78c, Bih79, Bro79a, BS76b, Cha79a, CS78, Duc77, Eas78, Egu79, FK79e, Gar78, Gol78b, Heg78, Her77a, Hir75a, HPS77j, Ioo76, Jon79, Kos75a, Lin79a, NR78a, Obe78, Rub75, Sch79j, Sta78d, Yan77n, dP76a]. Invarianten [Jur78d, Jur78h]. invarianter [Haz77c]. invariantes [Gui76a, Ker75f, Poé76a, RS75, Rub75]. Invariants [Bae78a, Bro75d, AF78, Bro75b, DC75a, Dav78f, DS78i, Elw78, Far75c, Far76b, Gui77b, Jur78h, Lut79, Pau78, Poé76e, Pro78, Raï78, Rot76c, Spr77b, Wal75e]. Inverse [BG78c, Grü75b, Miu76a, Che76a, Chu76, Col76d, Düm79, EH76e, ET79, FM76a, Gag77, JS76, McF75, Mic76b, Nas79a, New76a]. inverses [Lel78b]. inversion [Ber76g]. inversive [Bea79b]. invertibility [AB76]. inviscid [Gui77c]. Invited [Nov77a]. involutions [Con79b]. Involutivity [Mal79b]. involve [Zag77b]. involving [Pet79]. inwardness [CK75]. irrational [AR79]. irrationality [Gal76g]. irreducibility [Car77a, Gro78o, Roq75]. Irreducible [Ber78b, Hum76b, Blo79a, Jam78k, Sch76k, Sch76l]. irreductibles [Del79, Duf75a, Voi79]. irregular [Geo78a]. Irregularité [Bru75]. irrégulier [Dab79]. irreguliers [Mal79a]. Irrfahrten [Cré79]. ISILC [DM75c, MOP75]. Isogenies [Yan77k, BK75b, Yan77m]. Isojacobic [Wac78]. Isolated [CS78, Per79a, Kni78, SW78a]. isolé [Dem75]. isolée [Mal75a]. isometries [Kad78]. isometry [GW79j]. isomorphes [Ele78]. Isomorphic [HRW78, Rob79c, Qui79]. Isomorphism [Wal75e]. Isomorphisms [Mar76b, BK77d]. isospectral [vM79]. isotopies [Wag77]. Isotopy [Lar79, Sch79i, Smi76a]. isotrope [LS76, Stu76b]. isotropic [Pow78b]. Itô [Sło78, Wal78d]. Iterated [Bau77b, Bau77c, CLM76, FU78, Hei79, Kue76b, Kue76a, Lit79, Poh75]. iterates [Muk79]. Iterative [Axe77, IK79, You75, Gaj79, GP77, Hac78a, Man79, Wir78]. itéré [Ava77, Pis76]. Iteres [BF75d]. ITPFI [Osi78]. IUTAM [GN76]. IUTAM/IMU [GN76]. IV [BK75a, CW76a, Fáb79, FCG75, Nov77a, CH78b, Kle77d, Par76, Saf77]. Iwaniec [Des79b]. IX [Mey75d].

J [For79a, Ill77b]. J. [Alo79, Bor78c, Car78c, Del75, Hir78a, Kos75b, Lan78, Le 75, Lem77, Mey78, Mok78c, Ram79b, Ste78e]. J.-P [Ill77b]. jackknife [Rey78b]. Jacobi [Jou79c, Koo77, MvM77, Sch79j, Tak75a]. Jacobian [Ear79]. Jacobson [Fai79a, Hau77]. Jacod [Mey78]. Jacquet [Li79]. jadis [Wei75b]. Janeiro [PdC77, Sch78g]. Jantzen [Bor78c]. January [CUW75, Gol75, Hey79, NR78b, Sch78g, Ste76c]. Janvier [Nor75]. Japan [AK78, MP76]. Jean [Lai78]. Jensen [Gam76, Ste78e]. jet [Mat76]. Jets [Mol77-29, HT75a, HPS77a]. Jeulin [ÉS79]. jeux [Gri77]. Joensuu [LLS79]. John [RR75b]. join [Bou78a]. Joins [Yan77l]. Joint [GN76]. Jordan [Loo75a, Loo75c, Loo75d, Loo75e]. Jørgen [May77a]. Joseph [Bor78c]. Jost [Sto75b]. Jouanolou [Ill77b]. Journées [CK76, Lel77a, BR78a, DCV78]. Joyal [Osi75b]. Juillet [CDV75, CV77a]. Juin [DCV78, Nor75, Nor78c, Rau75e]. July [BCD77, Bec76b, BH76b, BK75a, BR77c, BRP78, BGS78, CCJS79, Eve75a, FMS79, KK78a, Mil78b, Ols78b, PdC77, PW79, SZ77b, SZ77c, Wat76, Wat78b]. jump [Eps79j]. jumps [Kom76, Ver79b]. June [BH78, Bec79a, BØS76, BK76a, BR78a, CV79, CCJS79, DJ79, GM79d, HL79b, HS79b, JMR75, Kel75c, Lai78, MMP78, Ols78b, Ros75b, SS76a, Tem76c, Wat78b].

K. [RR79b]. Kähler [GW79b, Lie78, Lie79]. kählériennes [Bou79b]. Kakutani [Fla76, dSLM75f]. Kakutani- [Fla76]. Kalkül [Maa75]. Kan [CS76a, FK79e, Har79a]. Karlsruhe [BMS76]. Kármán [Duv78]. Karp [Kue75b]. Karpacz [CUW75]. Kashiwara [Bon78b]. Kawai [Bon78b]. KdV [Raţ79]. Kelley [Fla75, MM75b]. Kent [BCD77, CS76d]. Kentucky [BS77]. Kepler [Dev78, Ver79b]. Kerne [Jan79a]. Kernel [CL75g, GM79a, Clé79, Del76, MG79, Ros79, vO75f]. Kernels [CL75d, CL75f, Wei76j, Yan77m, Ber79a, BF79a]. Kervaire [Bro79a]. Khinchine [Man76b]. Kiefer [CC76]. Kiel [DM75c, MOP75]. kind [Bak78b, Bru76e, Cla79g, Gér77]. Kirby [Ben75a]. Kirillov [Bre75d, Loe75]. Kjellberg [Ess75f]. Klasse [Don76]. Klassen [FM76b]. Klee [Sta78b]. Kleene [Mol77b, Mol77g, Mol77j]. Klein [Iva79a, Ree78a]. Kleinian [Har78c, Kru79]. KMS [Ara78]. Knight [Get75i, Mey76c]. Knot [Tro78, DS76b, Gor78a, Kea78, Lev78, dB76, Hau78a]. knoten [Vog78]. knots [Bir79, Hau78b, KW78, Lic79, Lit79, Sto78a, Tay79]. knotted [Bad77]. known [Pet78a]. Kober [Sne75]. Kochen [Che76c]. Kodaira [Pop77a, Uen75a, Uen75g]. Koeffizienten [vG76]. Koinduzierte [Voi77p]. Kolchin [War76g]. Kolmogorov [Dur76a, FJ75i]. Kombinatorische [Bur75b, Fel75]. Kommutative [Kra75d, Kra75e]. kompakter [FK79b]. komplexer [FK79b, Haz77b]. Konstruktion [FK79b]. Konstruktionen [Fel75]. Konstruktive [Buc75]. Kontrollprobleme [Har75a]. Kontrolltheorie [HJSW75, Wec75]. Konvergenz [DG76a, Jur78l]. Konvergenzbeweis [FK79c]. Konvergenzsatz [Ger79d]. Konvexe [Hei75]. Korteweg [Adl79a, Run76, Wah76]. Kostant [Sch79g, Sim77]. Kreisbereichen [AE77]. Krickeberg [Sek76]. Kripke [Osi75b]. Kristiansand [BØS76]. Kriterium [Voi77m, Voi77a]. Krümmung [GMK75b]. Kuhn [FJ75g]. Künig’s [Tru76]. Kunneth [McL79, Hod75, Mor79a]. Kuros [Tur77]. Kurt [DM75c]. Kurtosis [Man76a]. Kutta [Fri78a, Bak78b, Bet78, Fri78a].

L [Ill77b, Mil78b, Ray78a, Why78, Bre77k]. L. [Cho78b, Fel77c, MT77a]. labelled [RF79]. L’Absolue [Jac79g]. lacunary [Kue78]. Lagrange [Cau76c, Gag77]. Lagrangian [Ler78]. lagrangienne [Ler78]. Laminar [HH77h]. Laminated [SW78b]. laminations [HPS77b, HPS77l]. Landau [FJ75i]. langages [Sak77]. Langen [Maa75]. Langlands [Gér79b, Li79]. language [Fef75b, KMS75, Osi75b, vK75]. languages [Cun75, Glo75, Glo78, Gre75a, Gre75c]. Lansing [Kel75c]. Laplace [Cor79g, SNWC77j]. Laplacians [vM79]. Laplacien [MD78, Ava77, Bor75a]. L’Approximation [CL76a, Lio77b]. Large [DV75c, HH77f, PPW78, BW76, BF76b, Boo75, Dev75, Gie76, GKM75, KM78a, Rhe75, Tay78e, Tay78d, Tay78c, dR79]. largest [Fef75a]. Lashof [Sna75]. later [Ono77]. Latin [PdC77, Hei77, Pre76a]. Lattice [Pre76g, SW78g, CL75a, SW78i, Szu78]. Lattices [Huš76, Pet76b, Pet76d, Eps79b, Eps79e, Eps79m, HK78d, HK78f, HK78g, Jac75a, Sco75, SA77]. Laurent [Mah76b]. law [Hah79, Hei79, Joh79q, Kue76b, Kue76a]. Laws [Tay78e, Tay78d, Tay78c, BW76, DG77, Gie76, Haz79a, SS76e, Urb75b]. Lawvere [Car79e]. Layer [BGM77, FR76, Nør78d, Ver79b]. layers [Dud79, dR79]. leaf [Aki78d]. least [Gen76, GM76a, GV79a, LP79a, Ric75, Spe75b]. leaves [Eps75, Hol78a, Lam75, Sch75u, Sch75v]. Lebesgue [Her77b]. Lebesguesche [Jan79a]. Lebowitz [Car78c]. Lecture [Lan77, Rau75a, Rau75b, Rau75c, Bow77, Cho77a, Man77, Mar77, Raţ77, Saf77, Sma77, Wil77a]. Lectures [Boo75, DL76, LMW75, Wal75d, Wra75]. Lefschetz [Bat78, Bro75g, Gro77c, Gro77e]. Left [Pap79a, Sle75, Yan77n, Bea79a, BS78, BJM78c, Haa79, Pap79b, Sha75]. Left-definite [Sle75]. left-right [Sha75]. Legitimacy [Bör79b]. legitimate [Bon78a]. Lehmann [Sto75b]. Lelong [Ste75c]. Lemma [BLR75e, Rob78, Voi77s, BLR75f, Cor79i, Lau78, Tru76]. Lemmas [Mas75c, Lan76f, Mah76h]. lemme [Bru76d, Kan75b]. length [FS78a, KN79, Par79a, Sch75j, SW78l]. lengths [Bos78, Mol77c]. Lenstra [Ker75f]. lentement [Gat77]. L’équation [Rap76]. Leray [PP79]. Letter [Qui76a]. leurs [Car77b, Voi79]. level [Pop77d, Rat78p, vdLT79]. Levenberg [Mor78c]. Levi [Pfl75c, Pfl75a]. Levi-Problem [Pfl75c, Pfl75a]. Levy [Lap79, Cie75a, CC76, Hid78, Man76b]. Lewy [Nir75]. lexicographic [HSM77]. liaison [LV75]. Liapunov [Kno76, Mar79b]. Libération [Bas75]. Libre [Lio77b, Bai76, Dem77b, Pue78]. Libres [Vie78a, Fer77, Vie78b, Vie78c, Vie78d]. Lie [ETFS75, ETFS79, AdS75a, Bar79a, Bar78c, Bar78d, Bor78c, Bro78a, Bro78e, CN77, Car76b, Cha79a, Cha77, CB75, Cow79, Cré76, Cur75c, Del79, Fed78, Gla76c, Goo76c, Goo76d, Goo76f, Goo79c, Gui79a, Gui77b, GKR77a, GKR77e, Hae78a, Her75a, Hir79a, Jos77, Jos79, KZ77, Kos77, Lio77a, Loe75, Mol75, Ne’77, Poé76b, PS79, Ren77, Ren79b, Roy78, Sch79d, Sch79f, Sch77e, Spr77f, Tak75a, Tit77, Var77b, Ver75, Vie78a, Vie78b, Vie78d, Wal78c, Yan77o, tD79b]. Lie-algebras [Bor78c]. Liealgebra [Voi77k]. Lieb [Car78c]. liée [Ava77]. Liées [BWK77l]. Liénard [LdMP77]. liés [GB79, Kôn76]. life [Wil78b]. Lift [Har79a, Cor75]. Lift-extensions [Har79a]. lifted [Lub79a]. Lifting [Buh78f, Got78a, Gre79, KN78, Lau79e, Coh77, GP78a, Laz75b, vW76]. Liftings [MRN78, Eif76]. light [BS76a, Dui78]. like [Can75a, DS78b, She75]. Likelihood [Bal78, Ish76i]. Limit [And79a, EN76, Fei78f, Gra76h, AH76, Ber76f, EH76e, Ell75, Eva76, Fer78a, For76a, Fuj76, Gar76b, Gin76, Hec76, Hei79, Hel78b, Jai76, Jur75, KRZ77b, KRZ77f, KRZ77i, Ład75, Man76b, Mar76a, MW78, Pac76, Ryc78, SV75a, dA79]. limit-circle [KRZ77i]. limit-point [Eva76, KRZ77b, KRZ77f, KRZ77i]. limite [Bar77a, CR79, Hei78, Mar79a, Mok76, BT77]. Limites [LL76, HT75b, Hec76]. Limiting [Sai79c, Wil75g]. limits [AEK78j, AEK78n, AEK78t, Muk79, Suf76, Vuo79]. linaire [Sch76f]. Lindenbaum [BW77a]. Lindenbaum-algebra [BW77a]. Line [Liu78b, Bom76, Gre77b, Gut75, MM79a, Ste78f]. Linéaire [BR78a, Art75, Duv78, Gue75, Mos78b, Pue78, Sim78b]. Linéaires [MT77c, BP79b, Bis78b, Bré76c, Con76b, Dam78, GM77o, Gos78, Hes78, Jac77a, Let79, Sch77h, Ser77, Sko78, Tar78, Voi79]. Linear [Eps79i, Gra76i, KRZ77h, Kel75c, Liu79b, Mah76h, Mah76f, Mah76g, Mas75d, Ree76a, Ros78e, Sch75d, Tay78j, AEK78q, AEK78s, AL79, AW78, Axe77, BPFS76, BR76a, Bra79a, Bru76e, Byc78, CG79b, CS76c, Con75a, CN76, Cot75, Dah76, Dav78h, DH79, EN76, Fat75, FJ75k, Fly79, GM77e, Gen76, Geo77, Gla75b, Gór78b, Gra76a, Hac78b, HvW79, HP77, Hes79, HPS77h, Kan78b, Kat75, Liu78a, Liu78b, LP79a, Man76d, Mas75e, Mas75h, Mas75g, Nag76, Nam79b, Nam79d, OL76, Ral79a, Ree78a, Rei76b, Rei77b, Sch75e, Sch75h, Sch75o, Sij79, Sri78, Tay78a, Tay78d, Tay78h, You75, dP79]. Lineare [GMK75e]. linearen [Ess76, FM76b, Gro76]. linearer [Kna79a]. linearities [HS75a]. linearization [Per78b]. linearized [Ell75]. Linearly [Bra79k, dlR75]. link [Gol78b, dS79]. linking [Gol78b]. L’Involutivité [Mal79b]. Liouville [Eas76, Kal78]. Lipschitz [Dei77g, Hah76, HPS77a, Mat77a]. lipschitzien [Yam78a]. Lisses [Rod77]. lists [Mol77n]. Littlewood [Mey76a, Jam78l, Mor77d, dG75d]. Livne [Moi77f]. Loads [Lap79]. Local [AB76, Bib79c, Bou78f, Bou78e, Bou78g, Bra79i, Buh78g, Gel76d, Gel76e, HPS77i, Lac77, NV79b, Ser75a, Tem76b, Bae78c, Ber76g, Bra79c, But75, Che76c, Chi76b, Cir78, Ekl75, ET79, HPS77j, Hoc79b, Jan75, Koc75, Pet78a, Pra79, RR78, RR79a, Rin75, Roo79, Ros78j, Sta78d, Vas78, Yan77s, Bou78g]. locale [Bar78c, Ber76g, Bou76b, Bru75, DT77, FS78b, Meb78, Ron75, Ser75a]. Localement [Hey78, Ber76c, Bou79a, CL76b, Deh76, Hel78a, Kos76c, Lap79]. locales [Lép78c, Reb79, Sid79, Str78]. Localisation [Cos79, Ser75h, Ser75i, Len78]. localisées [Vai78]. localizable [Joh79r]. Localization [Bre79b, Hil77b, Ran79a, War76e, Bea76, Bea79b, Hil78, vO75b, vO75g]. Locally [AEK78l, Far76g, Rei75, Wil77c, AEK78a, CV77e, Chi76a, CSW76, CM76, Duf75b, DR75i, Egg78, HK78i, KR79, Lap79, MP79b, Sat76a, Tho76b, DKT78]. locally-convex [Lap79]. locating [CMPY79]. locaux [And79b, Avr79, Ill77a, Le 75, Pau78]. Loewy [CF79]. logarithm [AR79, Hei79, Kue76b, Kue76a]. logarithme [Pis76]. Logarithmic [Gro78b]. logarithmically [Ber79b]. Logic [Cro75b, DM75c, FMS79, KM78b, Kue75b, MR77h, MOP75, Ner75, Par75a, Cre75, Cur75b, Ekl77, Flu75, FS79a, GT75, Imb75, LEV75, MR77f, MR77j, Ras75, Tol75d, Par75a]. Logical [FR79b, Osi75a, Vol75b, Vol75a]. logicians [Cro75a]. logics [Kue75a]. Logik [Oss75]. Logique [Car79e, Imb75]. logiques [Gri77]. Logmodular [BK77j]. loi [Pis76]. loin [And75]. Lois [Bal79]. lokaler [KPP^{+}78c, KPP^{+}78d, KPP^{+}78e]. lomorphic [Bol75]. Long [BH78, Bra79l, Sai79e, Zel76a, Maa75]. Long-Range [Sai79e]. look [Fuc77b]. Loop [CLM76, FP78f, Har75a, CMN79, FP78h, May77j, Pet76e, Roo79]. loops [Mac77a]. Lorentz [Bam79, SS76d]. Lorenz [Lan77, Par79b, Rue76, Wil77a]. l’Orthogonalité [BWK77l]. Losik [Mac78]. loss [Roc75d]. Lösung [Fri78a, Kup76, Mic76c]. Lösungen [HK76, Kna79a, Jur78l, Jur78p]. Lösungsdarstellungen [Hee76]. Lösungsfunktionen [Jur78n]. Lotka [GV79b]. Louis [Foa77a, Nas79b]. Louis-Pasteur [Foa77a]. Low [Fen79, Ecc78]. Low-Dimensional [Fen79]. Lower [Dud79, FR79h, HL79a, Bom76, FR79g, FR79i, GvW76, Mah76f, Mas75e, War76k]. Löwner [Tam78f, Tam78h]. Löwner-functions [Tam78f, Tam78h]. Lubkin [Cor79b]. lubrication [Cap77]. Luminy [CDV75, CV77a, CV79, Rau75e]. Lusztig [Ser77, Spr75, Sri79e]. Lyapunov [Cop78c]. Lyon [BGM77].

M [Cho75a, Mok78c, Poi77, Wil76a]. M.

[Ali75, Bon78b, Bor78c, Car79e, Del77e, Le 77, Raï78, Ros77]. Macaulay [Hoc79a, Len76a]. Macdonald [Dem77a]. machines [Pos76]. Mackey [Voi77g]. Macmahon [And77j]. Macrocausality [Iag75b]. macroscopic [Grm79]. macroscopique [Duv77]. Macroscopy [Abr76a]. Made [Gér75b, Fin75]. magnetic [Gre77b]. magneto [vH79a]. magneto-hydrodynamic [vH79a]. Mahloness [Mol77k]. Mahowald [San79]. Mai [Pha75b]. main [Bra79m, Lan76e, Mah76a, Sto77b, Sto77c, Sto77f]. majoration [Bou79a]. Mal [Tik79]. Mal’cev [Smi76e, War76f, War76l]. Malgrange [Ram79b]. manifold [Joh79g, Lar79, Sta77a]. Manifolds [Aki78f, Aki78h, Bla76c, BLR75d, Dav78e, Far76h, Fen79, GW79d, GW79g, Ham75e, HPS77g, IP76d, Joh79h, SNWC77e, Aki78b, ACH77g, Bar78f, BC78a, BR78b, Bla76a, Bla76b, Bru78q, BLR75b, Cha75b, Dup78j, Ecc78, EH76d, Far76f, FR79a, GH78a, GH78b, GHM78, GW79b, Gut75, GdM77, Ham75f, Hau79, Hay78, Hem75, Hir75a, HPS77c, HPS77j, Hir78b, Joh79c, Joh79d, Joh79e, Joh79l, Joh79m, Koo77, Kos77, Lan79b, LF79, Lie78, Lie79, Liu78a, Mar79c, Moi77a, Nam79b, Nam79c, Nam79f, NR78a, Ray78a, Uen75b, Uen75h, dS79]. manipulation [Per78a]. Mannheim [BH76b]. Mannigfaltigkeiten [GMK75c, GMK75g]. Many [Ras75, Mah76f, Rau75c, Sch76o]. Many-valued [Ras75]. map [Bin75i, EL76, GW79j, Raţ79, Sco77, dP76b]. map-germs [dP76b]. Mapping [RR75f, AEK78d, AV79, BLS75, Cat79, Hum79, Joh79l, Mic76b]. Mappings [WdPL76, AEK78q, Ahl77, Aki78f, Aro76, Ber79c, Car79b, Dav79, GM77f, GM77e, GM77j, GM77l, Ghi79b, Ike79, Iva79b, KY79b, Ław79, Leh79, LF79, Mat76, MR79, Per79a, Por79, Rei76a, Sbo79, SS76c, Sch75f, Sch75p, Sch75t, Sre76, Ste78c, VGK79, Vuo79, Won79g]. Maps [Ada78, Bau77e, Con79e, Con79d, Ham75e, Bal76, Bir79, BHP78b, CY78, Dei79, DS78b, Ham75d, Han75b, Het79, HMRS78, Kan79, Max78, Nam79c, NR78a, Pas79h, Pop77j, Rud77b, Sch78c, Sto77h, Suf77]. Marcel [Bre78]. March [BM78a, BM78b, DKT78, ES76, GR76, Nat79]. marche [Éli79b, LL76]. Marches [ER75, GKR77b, GKR77d, GKR77f, GKR77g, GKR77c, GKR77e, Hen79b, PS79, Roy75, Roy78, Cré76, CL76b]. marching [Fro75]. Marginalia [DJ75]. Margulis [Tit77]. Marie [Del77b, Ill77b]. marker [Fed78]. Markov [DKT78, BD78, DGS76b, DV75c, Dyn76, For76a, Fuk76, Geo78b, Get75e, Gus76a, Hit78, Hud78, KT76, Kry78, Kun76, Let76, Lin79a, Mal78c, Man76c, Par75b, iS78, Sil76b, Sta78d, Vas78]. Markovian [Ruc78a, Wan75]. markovien [Zan77]. markoviennes [Ruc78b]. Marquardt [Mor78c]. marqué [Bré76a]. mars [dlH79c]. Marseille [CDV75, CV77a, CV79, GN76, Rau75e]. Martin [Bro78b, Dev76, Föl75, Gri77]. Martinelli [Roo78a]. martingale [Edg78, GMP78, Gri76a, Jac78b, Kaz76, Pra79, Sło78, Ste77a, Mét77e]. Martingales [Jac79b, Jac79h, Jac79j, Jac79l, Jac79n, Jac79p, Jac79q, Jac79r, Sid79, BM77b, Bré76a, Cai78, Cai79, Cha76c, Cho75b, CM75, JM76, JY78, Lép78c, Lép78b, Lép78d, Mét78a, Mét78b, Pel78b, Reb78, Reb79, Sek76, Str78, Woy75, Yoe79, Yor76, Yor78a, Yor78b, Fei78c]. Maryland [KZ76].

Masani [Sza78]. Maskit [Iva79a]. Maslov [Lio77a, Wei75c]. Mass [Kah76b, Kah76b]. Maße [Haz77b]. Maßen [Jan79a, Haz77c]. Matching [Eck77, Mau79]. Matemática [PdC77]. Material [Blo79d]. Matériaux [Lio77b]. Mathematical [AHK76j, BR77a, BR77c, BRP78, Blo79e, Cro75b, FW79, GR79f, GM79d, HS79b, Nas79b, NR78b, Sat79e, Tay78f, Ber76e, Bri79, GK76a, Law76, Oli76, Ptá77, Rau75b, Sne75, Cop78i, GM77n]. Mathematician [Dev77c]. Mathematics [CW76a, HS78, HW79b, Lit77a, PdC77, SW75d, Dev77a, Fef75b]. Mathematik [SZ77b, SZ77c, Nic78]. Mathématiques [Tik79]. Mather [Ser79a]. matière [Car78c]. Matlis [Ber79e]. matrice [Pha75a, Roc79, Vai78]. matrices [BG78c, Cot75, EW76, Far76c, Hal78b, LW75, Lod76, Mul75c, RS76, Seb79, Sha79c, Sri78, Tro78, Vai78]. matricial [PS77]. Matrix [Bar77b, Jur78n, Dek79, GL78, GK76b, Iag75b, Pea76, Wei78b, Wid78, Wil76b]. Matrix- [Jur78n]. matroids [Oxl79]. Matsusaka [Ray78b]. Mattei [Ram79b]. Matter [Ser75b, Ser75c, Ser75f, Ser75g]. Maunder [AL77]. Maximal [CW75, Hof79a, Ari79, Bra79e, Bra79f, Bra79k, DGS76m, Den76, Far76b, Fis76a, LeP78a, Leh76, MN79, McG78, Mic78c, dG75d, dG75f, Gér75h]. Maximaler [Dic75]. Maximally [Bra79n]. maximaux [Gér75h]. maximum [BC78b, BP79b, Hir79b, Lum76c, Lum78, Pir77, Rit76, Sib77]. Maxwell [EW78]. May [AL78, BMS76, BS77, Gol75, SZ77a, SW75d]. Mazur [Roc77, Ser79b]. mean [GW79k, Mar79d, Per79a]. means [BJM78c, Džu76, Ede76, Glo78, Kaw79b, SS76e]. Measurability [Mas76]. Measurable [CV77c, CV77g, BK77h, Bre79c, CV77b, CV77j, DB78, Roc76]. Measure [DGS76j, Low76, Sto77h, Wal75f, Wal75g, Aar79, Alp79, Bal78, Ber79a, Bor78a, BD78, Bre79b, Chr78, Coh78, Dev77e, Far76g, Get75o, GvW76, Kra78a, Mah77, Mas75f, Pla75, Sat76a, Sat79a, Sch79j, Sto76, BK76a, KK78a]. Measure-preserving [Sto77h, Wal75f, Wal75g]. Measure-theoretic [DGS76j]. measure-valued [BD78]. measured [Pin79]. measurement [Hol76]. Measures [AD78b, AD78a, DGS76k, DGS76m, Den76, DGS76l, Geo79i, Hey79, Kow75, Kuo75c, MN79, MT76a, MT76b, Bak78a, BF79a, Bom77, Bow75e, Bri76, But78, Byc78, Byc79, Car78b, Cha76b, Cha78b, CBR76, Chr76, CS79, DGS76i, Die75d, Duf75b, Eis76, Elw78, Far76f, Fly79, GS76, Gam76, Ghi79a, Gra76c, Gua78, Gue78, Hac78b, Haz77e, Hel78b, Hir75a, Hof79a, KLŠ76, KS76, Kno76, Kuo75a, Kuo75b, Kuo78, Lap79, Lin79a, LS79b, Mis79b, Muj78, Muk79, Mus76, MRN78, NS76, PPW78, Ros79, TDHJ76, Tur78, Vas78, Wal78d]. mécanique [Ler78]. Mechanics [GN76, Ode75, Ben76a, Geo76, HBT75, Kas75, Ler78, Lum77, Mos77b, Rog76a, Śni78, Sze75, Urb75b]. Media [BK76f, APF76, BBB76, Cha76d]. medium [BK76c]. Meeting [GR76, Løn79b]. Mehler [Gas75]. Mehrdimensionale [Sch76c]. mehrfach [Kle77j]. Meilleur [Gil78g]. Meilleures [Bea78d]. Melbourne [Lit77a]. mémoire [And77a]. Memorial [MSZ76, SW75b]. Memoriam [Kue75b]. memory [Gal78]. Menge [Bro76a]. Mengen [Jan79a]. Méromorphe [CH78d, Jur78k]. Meromorphic [Nam79e, CW79, CH78d, Dam77, Fis76g, Gro77a, HW79a, Lut79, MT77b, Ste78c, Jur78k]. mesh [McC78b]. Meßbarkeit [Jan79a]. Messreihen [JLR76]. mesurabilité [Gol78a]. mesurables [BG78d]. Mesure [Cau76g, Her77b, Kré77, Lel78a, Nov77b, Nov78]. Mesures [Jac79k, Kah76a, dSLM75f, Le 76, Nan75, Bad78, Ben75b, Ber76c, Buc78, Che78c, For76b, Gol78a, Gui76a, Lap79, dSLM75d, Sjö76a, Yor78b, Yor79]. Metabolic [ACH77e]. metacyclic [Gro78a]. metalinear [Bla77]. metamathematics [KMS75]. Metaplectic [Gel76f, Gel76g, Gel76a]. Method [Miu76a, Ale78, BR77a, BM79a, Böh76b, Bre75d, Bré75f, BS76a, Bro79b, CMPY79, Chu76, DS76a, Deu75, DDP78, EW78, Fer78a, FM76a, Fra77a, Fri78a, GM77b, GR79g, GR79h, GP77, Hab76, Hac78a, Her75c, Het76, Hir76a, Imb75, KO79, Kru76, Kur78, Lam78e, Luk79, McC78b, Miy77, MS78e, OL77, Opp78, Ort78, PP79, Pop77j, RT77, SB77, Sch75t, Sim78c, Urb75a, Wer79]. Methode [Sch76c, BG78a, Bré76a, Dix79, Imb75, Tar78]. Methoden [Har75a]. Méthodes [Bea78e, Cho75b, GR77, Kré78b, Lem79, Kré78a]. Methods [BR77c, BRP78, BTW75b, Ekl77, GM77n, GN76, GLL79, Lio75a, MR77h, MWW76, Nas79b, Ode75, Pin75a, Rey78f, Sat79h, Ale79, Ama76b, Axe77, Bak78b, BP79a, Bet78, Bor79c, Bre77o, But76, Cla79g, Col76a, Cot79, Čud77, Dah76, Dou77b, Dra79, Dyc76, Fle76, Geo77, GK76a, GH76, GP77, GV79b, HBT75, Hof76b, HM77, IK79, Jel78, Jer76b, Kaw79b, Kel79, Kra79, LP79a, Lyc76, MG78, Moe79, Möl77-30, NV79e, Pop77j, Pow79, Ral79a, Rei77b, Sch75i, Ver79b, Wal78c, Wam79, Wid78, Wir78, vdB79]. Metric [Aki78g, Aki78h, Ale75, Gal76d, Gal76e, Kel75c, Sei75, Wol75, Bla76f, Byc78, Byc79, DGS76k, GW79a, Kle75, Pow79, Sto75a, Val75]. metrics [Ghi79b]. Métriques [Bea78d, Bea78h, Con75b]. metrische [GMK75f]. Metrizable [AEK78m]. Mexico [AHW77]. Michel [And77a]. Michigan [AL78, Kel75c]. Micro [KK75]. Micro-hyperbolic [KK75]. Microanalyticité [Pha75a]. microdifferential [Mal79b]. microdifférentielles [KS78b]. microdifférentiels [Mal79b]. microfunctions [Kaw75]. Microlocal [Iag75a]. microspectre [Cho77b]. mildly [Gla75a]. milieux [Lév77]. Mills [Car78a, Raw79]. Milnor [Qui76a]. Milyutin [Das75]. Mimicry [Day75]. mince [Dar77]. Miniconference [BCN77]. Minimal [Emc78, HP76b, Bar78f, Böh76a, CdC77, Eck75, EN79, GPR78, Hir75b, Hof79m, MP78a, Mic78d, Ros76b, Wil77c, Wol78]. minimaler [Küh79]. minimality [MS78b]. Minimalpunkte [Bro76a]. minimaux [Mic76a]. minimax [Rab78a, Str79b]. Minimierungsaufgaben [Bro76a]. Minimization [FJ75j, FJ75k, FJ75b, FJ75a, FJ75m, FJ75q]. minimizing [Duc77]. Minimum [FJ75l, Hal79e, FJ75r, GM77p]. Minkowski [Cot75]. Minorations [Poi77]. Miscellanea [Gro78m]. Miscellaneous [Bru78i, Gal76f, GK78f, Lan76f, Zie79f, Gra76j]. Mischungen [Haz77d]. Missouri [Nas79b]. mittels [Hee76]. Mittlerer [RR75c, RR75d]. Mixed [CM78a, GR79g, GP77, Miy77, MS78e, OL76, OL77, RT77, Shi75, War77b]. Mixing [DGS76n, Ist78, Lin79a]. Mixtures [Gua78]. mobile [Fed78]. modal [Cre75, GT75]. Model [Che76j, HW75n, LMW75, MR77h, SW75b, Win75, BP76, BBB76, Bow77, Che76d, Dav78h, EH76g, GW76a, Hal79f, HW75h, HW75l, Koc75, Par79b, Pow78b, Sac75, SV79, Ště78d, Str79a, Sza75, Wei78a]. model-companions [HW75n]. Model-completeness [Win75]. Model-completions [HW75n]. Model-Theoretical [MR77h]. Modèle [Rod75]. Modèles [Mol77v]. modeling [Gla77, HS75a]. modelling [Par79c, Zee76b, ZKB79]. Models [GK78d, Bal78, BS75, Bel76, CBR76, Cre75, Cun75, Dub76a, Fit75, FH79, Geo79c, Geo79d, Geo79f, Geo79g, Geo79h, Gra79c, Gua78, HW75f, HW75k, HW75p, HW75q, HW75r, KP77, Lau75a, MR77b, MM75b, MZ78, Mol77v, Mos76b, Pre76c, Pre76g, Rog76b, SW78i]. Modern [Blo79f, FJ75l, Nov77a]. modification [Ne’78, Sch78a]. Modifications [Fis76g, Wei76t]. Modified [BPFS76, Deu75, Gug78, Sch79j]. modulaires [CO77, Lig77b, MS76c, Rum79, Ser79b]. Modular [Hum76c, SS77b, Zag77b, And77k, Buh77, Buh78a, Coh77, Hej76b, Hir77, Hum76b, Kat77, KL77, Maz77b, Moi77f, Sta77d, SDB75, SD77, Tay77a, BK75a, SZ77b, SZ77c]. modulating [Gat77]. module [BDE^{+}77f, Gar76a, Jam78s, Jos79, Par76, Rog75b, Wal78b, Wie79, FW79]. MODULEF [Per79b]. Modules [AGR75, Bra79d, DW79a, Lev79b, NV79d, Tou76, War79b, AF78, Arn77, Bas75, BDE^{+}77g, Beh79e, BF79b, Bra79k, Bra79o, Che76g, Cou78, Del76, Dem75, DW79c, Fer77, FGR75e, Fox79, Gat77, Ham75a, Haz79a, Hoc79a, HK79b, Hum76e, HRW77, Jam78q, Kea78, Kos75a, Lad77, Le 75, Mic77, Moh79, MM79b, NV79a, Nic77, PR77a, Sta79a, Sus79, Tro78, War79a, Web78d, tD79f, Jan79f, Jan79g]. Moduli [LL78, Lau79a, Pal78a, Pop77f, Pop77h, Pop77g, Bom77, Car79a, JS77, Lau79c, Oor77, Pop77a, Pop77b, Pop77c, Pop77d, Pop77i, Pop77j, Tra78a]. Moduln [Jan79e, Kra75f, Voi77o, Gro79g, Jan79f, Jan79g, Jan79h, Voi77l]. Modulo [Rau75e, GD77a, Rhi75, Wei78b]. Moduls [Kra75a, Voi77d]. modulus [CC76, Pir77, Sha76a]. Mokobodzki [Mey77a]. Mokobozki [Ben77b]. Moment [Fei78g, Wed78]. Moments [Far75e, Gro78n, Kue78]. monadic [Mey75a]. monads [Lad76b]. Monash [Cro75b]. Monge [Sib77]. monodromie [Bon76, Kla79]. monodromy [Bir79]. monogenic [Bra76a]. monoidal [Gre79, Par76]. monoïdes [Lal77b, Lal77a, Sak77, Vie78c, Vie78d, Vie78a]. Monomial [Voi77r]. Monomiale [Voi77r]. monomiales [Lod76]. monopole [Gre77b]. monopoles [Bar78f]. monothetic [Bro78d]. monotone [Bré75e, Cen76, Ish76i, Sai78]. monotoner [Nic78]. Monotones [Gil78j, Bré76c]. Monotonic [Too78, Gal78, Gri76b, Kow75]. monotonically [Hof79a]. Monotonicity [Bru78j, Obe77]. Monts [Lai78]. Moore [CCMS75, Dok78, Gug78, MS79a, Why78]. Morgan [Joh79q, Mor77c]. Morlet [BLR75f]. Morphisme [Hou77, Hen75c]. Morphismes [Sko78]. morphisms [Fug79, Lau79e, MR77g]. Morse [Che75a, Cur75a, Fla75, GdM77, MM75b, Mos76b]. Moscow [DKT78]. Moser [Sch78b]. Mostowski [MSZ76]. Motion [Cau76e, Cau76f, Blo79b, Blo79h, Cau75, Cie75a, Goo78, Sil76c, Wil76a]. motions [Jen77g]. Motivation [Wei76k]. mots [Sab76]. Moussu [Ram79b]. Mouvement [Cau76e, Cau76f, Cau76g]. movability [DS78f]. moving [Geo79e, Ole78]. moyen [RS78]. Moyennabilité [Far75a, dlH79b]. moyennables [Ber76b, Eym75, Far75c]. moyenne [MD78]. MU [Boa78]. Muckenhoupts [RR75e]. Mullin [Oxl79]. Multi [Sib76a, Gus75, HM77]. Multi-dimensional [Sib76a, Gus75]. multi-time [HM77]. Multiaxial [Dav78e, Dav78i]. multicolorings [Ber78a]. multiderivative [Jel78]. Multidimensional [Fei78h, For76a, KY79a, KW78]. multidimensionnels [BDR78]. multifunction [CV77j]. Multifunctions [CV77c, CV77g, CV77i, DB78]. multigraph [Gup78]. multiparameter [Man76c, Sle75]. multipartite [HRW78]. Multiple [Hes79, BN77, Bre79a, Ded77, Fri79, Max78, Sat79c, Str77, Vai78]. multiple-valued [Max78]. Multiples [For79a, CG79a, CH78e, Hes78, Wag78]. multiplicatif [GMP78]. multiplicatifs [Yoe79]. Multiplication [Abr78, Coh79a, Har78b, Mas75c, Mas75h, Mas75g, McC77]. Multiplications [Kom78, Ros78f]. multiplicative [Gri75, Jac78b, Sel77]. Multiplicatives [Jac79i, AM79b, Mém78]. Multiplicativity [Dup78i]. Multiplicité [Bes75a, dSLM75g]. Multiplicités [Ser75j, Ser75a, Ser75k]. multiplicities [Ser75a]. multiplicity [Ang79, GM77g, Hit78]. Multiplier [Gro79f]. Multiplierräumen [Gro79c]. multipliers [BTW75d, FW78]. Multiplizität [Jan79e]. multisécantes [Le 78]. multistep [Jel78]. multitype [Kaw76]. multivalued [CV77i]. multivariable [Tua78]. Multivariate [Chi78b, Far76l, GK78c, HP77, KLŠ76, Rév76]. Mumford [Dem76a, Pop77i]. Mycielski [Hal79g].

N [DO77, Ste75c]. N. [Bor78c]. naguère [Wei75b]. Nancy [ETFS75, ETFS79]. nappe [Far75b]. narrow [Rud77a, Sun79]. Nash [Boc78, Ris76, Sch78b]. National [BCN77, NR78b]. natural [CV77h, Han75a, vK75]. nature [BF75b]. Navier [Tem76c, Foi75, FT76, Fro75, GR79e, GR79i, GR79j, Ioo76, Mar77, Ole78, Raţ77]. Navier-Stokes [Mar77, Ole78, Raţ77]. nd [RT77]. Near [MS78e]. Near-self-complementary [MS78e]. nearest [Ros79]. nearness [Ben76b, Hel79a]. nebentypus [Rib77]. necessarily [Obe78]. Necessary [Hal78a, Ros78f, Sha79b, Bro78a, MW78, Stu76b]. negative [vW76]. neighbor [Ros79]. neighborhood [Arm75, Dui78]. neighborhoods [BLS75]. Neroslavskii [Goo79a]. Nerve [Car76a, Rin76, AW75]. Nessel [Sjö76a]. nest [Ded78]. nets [BBL77]. networks [CCMS75, Kur78, Shn78]. neuere [Vek76]. Neumann [Com76, Ars79, BL75, Cho78a, DT77, Joh79p, Mac75b, NT79c, Ols78a, dlH79b]. neuronal [Kry78]. neutre [Sam76]. Nevanlinna [Mal78b, Sko76]. Newton [Bra76b, Deu75, Het76, LY79, McC78b, Var78]. Newton-method [Het76]. Newton-Verfahrens [Bra76b]. Nice [Cha75c, FR79a, HH78b, Pha75b, ST77a]. Nicht [Eif76, Haz77a]. Nicht-Standard-Beweis [Eif76]. nichtklassischer [Kup76]. nichtlinearen [Bra76b, Col76c]. Nichtlinearer [Jeg75, Mic76c]. nichtlineares [BG76b]. Nielsen [Gro77e]. Nigel [May77a]. Nikodým [Cha79c, Chi76a, Die75d, Edg78, Egg78, Gol78a, Hac78b, Pet78d]. Nikodým-property [Edg78, Egg78]. Nilmanifold [AT75a, AT75d]. nilpotence [Pas79i]. Nilpotent [Goo76c, Goo76d, Hil75, War76g, War76h, War76i, Goo76f, Goo79c, HMRS78, Kov78, Mos79, Pas79e, Pas79f, War76d]. nilpotentes [Bar78c]. Nilpotents [GKR77f, Lio77a, GS79a, PS79, Spr77f]. Nirenberg [RR75b]. Nishida [BG78a]. Nisio [dSLM75e]. niveau [Lig77b]. No [HM76b, HM78a, HM79c, Ørn77]. nodes [Kös76b]. Noetherian [Bea79a, Goo79a, Gor76, Hoc79b, Irv79, Kra78b, NV79b, Pap79a, Pap79b, Rob76e, Sta79c, Sta79b, War79b]. noethérien [RR79a]. noetherienne [Mar79a]. noetheriennes [Les78]. noetheriens [Pau78]. Noeuds [Hau78b, Gra77]. noise [AW78, Bal78, Gua78, Hid78]. noisy [Wah79]. Nombre [Her77b, Dar77, dSLM75g]. Nombres [Sig76, Dem75, MN75a, Nic75b, Wal77a]. noms [Ano79]. Non [Bea78d, BR78a, Bla76f, Bra76a, Bre77n, CDV75, CV77a, CV79, Cur75a, Fef75c, KT75i, Kra76a, Lam75, Mac75b, MMS78, Par76, Pre76a, Ree76a, vO75c, AHK76g, Ali75, BPFS76, BR76a, BP79b, Ber76b, Bla77, Bou79a, Bré76c, Can75b, CS76c, Con79f, Con76b, Cor79a, Dah76, Dem78, Duv78, Eff78, Far75c, FJK79, GM77o, Gér75b, Gér75h, Gos78, Gra76c, HM79a, Hei76b, Hes78, Kaw79a, KN78, Krá76b, KvQ75, Le 77, Lem79, Lub79a, Mag79b, Mas75e, MJ75, Mos78b, MT77c, Nag76, Nie78, Osi78, Pet78c, Pol79, Pue78, RS78, Ree78a, Sai78, Sch78a, Shn78, Sij79, Sim78b, Tak79, Tak78c, Tar78, TW78b, Yam78a, Yam79, CDV75, CV77a, CV79, Hör75]. Non- [Bra76a, Tak78c]. non-abelian [KN78]. Non-additive [Par76]. non-atomic [Gra76c]. Non-bounded [Lam75]. Non-Commutative [CV77a, vO75c, Mac75b, Eff78, CDV75, CV77a, CV79]. non-compact [FJK79]. non-complete [Kaw79a, Lub79a]. non-connected [Mag79b]. non-continuous [Can75b]. Non-existence [Bla76f]. Non-extensional [Fef75c]. non-imbeddable [Hei76b]. non-linéaires [Con76b]. non-linear [BR76a, CS76c, Dah76, Nag76, Ree78a]. non-lipschitzien [Yam78a]. non-normal [Cor79a, Pol79]. Non-orthogonal [Pre76a]. non-real [Bla77]. Non-self-adjoint [MMS78]. non-simplicity [HM79a]. Non-singular [Cur75a]. Non-standard [Kra76a]. non-stationary [Nie78, TW78b]. non-terminals [MJ75]. Non-trivial [KT75i, Osi78]. non-uniqueness [Pet78c, Shn78, Hör75]. non-vanishing [Mas75e]. Nonabelian [Con77]. nonbicharacteristic [Cos75]. noncentral [Far76i]. Noncoercifs [MR77m]. noncoercive [Fuč79]. Noncommutative [Hol76, Ale78, Gor76, CS76d]. Noncompactness [HPS77k]. noncomplete [GvW76]. noncomutative [Bea76]. Nonconforming [BB76]. nonconvex [Col75]. nondifferentiable [Hal78a]. Nondiscrete [Gér75b, Ptá77]. nonfocussing [BK76c]. Nonlinear [Ama76a, AW75, Bib79c, Dav78a, FJ75m, FJ75n, Fit76, Fuč79, GM77d, GM77j, GM76a, Had76, HC78, MM79a, Ode75, OW78, Ros78d, AW76, AB76, Axe79, BKM78, Bar79b, Beb79, Ben76a, Bit76, Boh78, BBB76, BN77, Bre79d, Bro75h, Bru76e, Cla79g, Dou77b, DDP78, FJ75d, Fro75, Gaj79, Gal75, Grü75b, HBT75, Het76, Hir76a, HM78b, Jer76b, Jor79, Kel79, Kre78c, LY79, Lio75b, Lon79a, Loo75f, McC78b, Meh78a, Meh78b, Miy77, Neu79, Noh79a, Noh79b, Nus79, OW75, Ort78, PP79, Pel78a, Pet79, Rhe75, Ros78i, Sch75c, Sch75g, Web78a, YF79, You75, dF75, Cha78a, GDMW76]. Nonlinearities [GM77m, GM77k, Jer76a]. nonlinearly [GM76a, Pow78a]. nonlocally [Rob77a]. nonmonotone [Jer76a]. nonnegative [MS78c]. nonoscillation [Meh78b]. Nonparametric [GR79a, Fri79]. nonprojective [Sta79a]. Nonselfadjoint [KO79]. Nonstandard [Roq75, Che75b, Čud77]. Nontrivial [Cus79]. nonuniqueness [Gri79e]. nonwandering [HPS77d]. nonzero [GM77j]. nor [Har78b]. Norm [FJ75l, JW77, Oma75, Blu75, Bra76c]. norm-dependence [Bra76c]. Normal [Kam76, Mol77l, Bra79t, CW79, Chi76b, CdC77, Cor79a, Dav76, Dav78f, HPS77h, How78, Mol77j, Mol77n, Pol79, Rat78j, Roc76, Sal78, Wei76m, Wil77c, vdE79, Fis76a]. normale [Fit78]. normales [Cor79a, Mar75]. Normalform [KPP^{+}78f]. Normalfunktionen [Buc75]. Normalisierung [Jur78p]. normalization [Lei75]. normalized [Sch75h]. normalizer [HO79, Klu77]. Normalizers [Yan77p]. Normally [HPS77l, HPS77j]. normaux [CH78e, Sig76]. norme [Paq79b]. normed [Day75, Day76, DW79b, Por79, Tay78a, Tay78d, Top76]. normes [Far75a, Sch76f]. norming [Beh79d]. Normproblem [BG76b]. Norms [CIS76, Die75b, Duc77, Joh79p, OW78]. Norstad [Ess75f]. North [MMP78]. Northwestern [Ste76c]. Norway [BØS76, Ols78b]. notation [Bau77a]. Notations [Jac79m, Smi76g, Won79a]. Note [McC77, Ood78, Bof75, Bor78a, CF79, Dok78, EWW77, EH78, Evs78, Gag77, HL77, JL77, LR79, Mic76b, Mul75c, Osi75b, SNWC77k, Tru78, War77a, Zab75]. Notes [Fuj79, KM78c, Mol77-28, Uen75f, Chu77b, Gro79h]. notion [Pra75]. Notions [BWK77h, Fis76b]. Notizen [Jur78j]. Nouveau [DM75b]. nouveaux [Che78c, Nie75b]. nouvelle [Tar78]. November [BCN77, BOS75]. Novikov [Mor78a]. Noyau [BLC78, Sib76b, Éli79b, Loh75, Zan77]. Noyaux [Far75b, For78, Fey78a, Hir79b, Roo78a, Roo78c]. NSERC [HS79b]. NSF [Miu76a]. nucléaires [Kré78b]. Nuclear [Dub79g, Won79b, Won79d, Bol75]. null [SNWC77m, Sto78a]. nullité [Ste76b]. Nullstellensatz [HW75o, Irv79]. Number [HR76, Ish76d, Nat79, Tol75e, ACH77c, Bab78, BC78b, Ber76a, Che75b, Erd77, EN79, Hin79, Ish76a, Ish76e, Ish76f, La 76, Mas79, McA76, Mek77, Mol78, Nat77b, Naz75, Sch75a, Sch76r, Sha79a, Sto79, Sun79, Tol75a, Nat77a, Nat79]. Numbered [BG78b]. Numbers [DL76, Gal76h, LT76a, Bar78f, Bat78, BW76, CP77b, Chu79, Dup79, Fin76, Gie76, Han75a, Kam76, Mah76c, Obe77, Pét77, Pol79, Tak78d, Tay78e, Tay78d, Tay78c]. numerators [Edr77].

Numerical [Ale79, Blu79, Cot79, GR79h, KY79b, Kaw79b, LT76c, MW79, Nas79b, Wat76, Axe79, BDNY77, Boh78, Col78, For79b, GM79c, Gus75, IK79, Ist78, JL77, Koß78, Kra79, Lam78h, Mül78a, Nie77, Ort78, PP79, Tou76, YF79, ZKB79, AT78, BGS78, Wat76, Wat78b]. Numérique [Bre77f, Gil78h]. Numériques [BWK77n, Gil78i, LS76, Lum76c, Tou76]. Numerischen [Häm76, Mic76c, AE77, Bra76b].

O [KPP^{+}78g]. Oberwolfach [AT78, Bec79a, BK76a, BOS75, BGS78, DJ79, GR76, Hey79, KK78a, MS78d, SZ77a, Bec76b]. objects [Mol77j, Rob76e]. obliques [Bon78c]. observable [Dav78a]. observables [Jaj75]. observation [Ada75, Bré76a]. Observations [Kre75c]. Obstacle [Bai77b]. obstacles [Eck75, Rau75c]. Obstruction [Bau77d, Bau77e, DM78a, Lau79c, Par79a, Ran79b, Sta78c]. Obstructions [Ewi79, Bau77g, Cla79h, Dov79, KS77, Smi76d]. Occasion [DM75c]. occurring [Pel78a]. ocean [dR79]. October [Zel76b]. Octobre [Nor78c]. Odd [ACH77f, Gre78b, dB76, CM79, Con79d, FP78a, FP78c, Ish76b, Sha76b, Sto79]. Odd-degree [dB76]. odds [Gel76c]. ODE [Ste78j, Ste78i]. ODE-solvers [Ste78j, Ste78i]. Odlyzko [Poi77]. old [Est76]. omitting [Sim75a]. once [Wei75b]. One [BK75a, CW76b, Gal78, Hem75, Mal78c, SZ77c, Buh77, Buh78a, Coh77, CW75, Dav75b, FJ75d, FJ75a, FJK79, Geo78a, Gla75b, Kol77, Lam75, Moe79, Roc75d, Sch75u, SZ77b, Sta77d, Sun79, Utr79, Wei76f, Wei76p, Wei76m]. One-dimensional [CW76b, Gal78, Gla75b, Moe79]. one-factors [CW75]. one-way [Geo78a]. only [Hir75b, Law75a]. onset [GH77]. Open [Rey78e, AEK78d, DS78k, EH76i, Gab75a, Gro77a, Kre75a, SW78d, Tak76b, Har75a, Pop77a]. Open-Loop-Differential-Spiele [Har75a]. operates [Yan77c]. opérateur [Ava77, Bru76c, Dem77c, Fit78, Mey76f, Mok78c, Zan77]. Opérateurs [BF75e, KV77, BF75d, Bon75, Bré76c, Cha76e, HL78, MvM77, Rit76, Rot76a, Rot78, Tur75, Hir78a, Kaj75, Las75, Raï78]. Operations [Hof79h, BDPR76, AL77, Dav79, FI77, Fef75c, FP78g, Gra75, Hof79b, Mic78b, Sna75, Suz78b]. Operator [Far75f, Mic78g, MS77, Rus77, Ben76c, BP78, Cuy79, Eps79j, Fei78a, Gli78, GP77, Gór78b, Gri75, Hac78b, HL79a, Har78b, Her78a, MS78c, Mic78d, Mic78c, Mic78k, PS77, Raţ79, SNWC77j, Shi78, Sin77, Str75, dG75d, dG75f, Haz79b]. Operator-halbgruppen [Haz79b]. Operatoren [Haz77c]. Operators [BH78, Cha75c, Cla79e, Cor79e, Far75h, GDMW76, HH77a, Sai79e, Abr78, Adl79a, Ama76a, Ama76b, BDE^{+}77i, Bla75, BTW75f, Bré75e, Bro78d, CP77b, Cha79a, Cim77, Cla79a, Cla79d, Cla79f, Cod75, Cod79, Col76a, Col76d, Cor79f, Cow78, Dav76, Dui75, EL77b, Far75d, FJ75k, FJ75b, FJ75n, Fit76, Fly79, HH78a, Hal78c, HP77, Her78a, JW77, KSWW75, Kal78, KS77, KK75, KRZ77h, Kaw75, Kos75a, Kur75, Lin79a, Lub78, McC78a, McD78, MS75, MS78a, MS79c, Mic78a, Nas79a, Nor78b, Øvr76b, Pet79, Ple75, RB75, Sal78, iS78, Shi75, Sne75, TY78, Wan75, Wei75a, Wei78b, Wil75i]. optical [BK76c]. Optimal [BOS75, Dun78, GP76, Gla75b, Gra78a, MG79, Bro78a, Das75, El 78, FJ75e, Hal78a, Kno75, Köh75, Mos78b, O’M78, Roc78, Sar76b]. Optimale [Dic75, FM76b, MH75]. optimierung [Bur75b]. Optimierungsprobleme [Hei75]. Optimisation [Lem79, BH76a, FP78l]. Optimization [Tam78j, Tam78k, Col75, Cza75, Eck75, GM76a, Loo75f, Pow78a, Pow79, BOS75]. optionnelle [Del77i]. Orbit [CN77, Dav78h, Hil78]. Orbites [Ren79b, Dix79]. orbits [Dev78, Ver79b, Wan79, vdAS79]. Order [ACH77f, CL75d, HK78f, HK78g, Jen77c, MR77h, Mol77u, Tol75e, Won76d, BY77, Bre75b, BTW75c, Bre77o, Con77, Cos75, DP77, Eff78, EN76, Eva76, Flu75, GK78b, GH76, Gra76g, Gre77a, Gri76b, Hah76, HW75q, HK78d, Jak79, KRZ77f, KRZ77i, Lev79b, LEV75, MR77f, MZ78, Meh78a, Meh78b, Mey75a, MG78, New77, OL77, RT77, Rob76c, Rob77b, Ryc78, SV79, Sch75c, SS78b, TW78b, Tol75a, Tol75c, Tol75d, Trè75, Won76b]. Order-Bounded [Won76d, Won76b]. Order-quasibarrelled [HK78f]. Order-quasiultrabarrelled [HK78g, HK78d]. order-theoretic [Gre77a]. Ordered [HK78b, Ama76a, Che78b, Ell79, Lou75, Naz75, Won76c, Won79f]. ordering [Wil77b]. orderings [Eps79i]. Orders [Won76a, BW77a, CM78b, Jac75a, Kuk76, OP79, Rei75]. ordinals [Fri78b, Kec78a, Ric75, Tak75b]. Ordinalzahlen [Buc75, Pfe75]. Ordinary [Bib79c, Dei77h, Hum76c, Hum76d, KRZ77d, BM79a, But76, Cod75, Cod79, Cop78h, Dei77d, GM77k, Gam79, Kra79, Lam78h, Ple75, Wam79, ES76]. Ordonnés [BWK77b]. ordre [BCH78]. Ore [Les78]. Orientation [May77i, NR78a]. orientations [San79]. oriented [HWH78, May77h]. origin [Tou76]. Original [Boa79]. origine [Tou76]. Orlicz [DK79, Fer78a, Lac78]. orrits [Wol78]. Orsay [Tem76c]. Orthogonal [Fei78i, Fra78a, Rob76c, Rob77b, vR79, Bae78b, Bri76, EWW77, Ead77, GW76b, Mag79b, Pre76a, RS75, Sha76b]. Orthogonality [BWK77l, Gro78n, Kuo75a]. orthogonaux [dSLM75c]. Orthomorphismes [BWK77i]. orthonormal [Gro78r]. oscillating [AS79, Gud76, Lio76, Nør78d]. Oscillation [Meh78b, Meh78a]. Oscillations [Nat77b, HM78b]. oscillator [AHK76b, AHK76a, AHK76e]. oscillatory [Bor79c, Var78]. Oszillation [RR75d, RR75c]. Otaniemi [LS79a]. Other [Ano75, Pop77j, SW78i, Tol75c, Boa78, Gro78r, HM77, SNWC77m, Suf77, Pop77i]. Ottawa [DG75j]. our [Ste78g]. Outils [Ser75l, Ser75m]. Outline [Sha78e, Sim77]. ouvers [Szp76]. ouverts [Kre75a, Rap76]. overview [Cod76].

P [Gér79b, Hir78a, Ill77b, Tit78]. P. [CC76, Le 77, Ser75p, Uen75f, dV75a]. Paare [Jur78r]. packing [HR78]. Packings [Top76]. Padé [Bul79, CC79, Cor79a, Cuy79, DDC79, Edr77, Gil78b, Gil78c, Gil78a, Gil78d, Gil78g, Gil78h, GM79c, Pin79, Wuy79a, Wuy79b]. Padé-approximants [Cuy79]. pairing [Bla75, FR79g]. Pairings [May77j, Fie78, Kea78]. Pairs [Kan79, Loo75e, Cod79, KS77, Loo75a, Loo75b, Loo75c, Loo75d, Ple75].

Palet [Mey76a]. Paley [Egu79, Ess75h]. Palm [Ben75b, dSLM75d, dSLM75f]. paper [Ess75c, Far76n]. Papers [CS76d, Cro75b, FW79, GR76, Kue75b, dBR77, Nov77a]. Parabolic [Sto77g, AW76, Ama79, Beb79, DDP78, FS79b, Gla75b, Gus75, Ham75g, Koß78, Ril79, Sto77c]. paraboliques [CO79b, CO79a, Spr75, Yam78b]. parallel [CdC77, Itô75, Ne’78]. parallel-transport [Ne’78]. parameter [AW78, Gra76a, Han79a, Ish76i, Sha77b]. parameter-excited [AW78]. parameterized [CS78]. parameters [CB76, Dur76a, Neu76, Tam78k]. paramétre [Her78b, Dem77c, Mét78b, Mey79]. Parametrization [Ław79]. parasite [Cli75]. parfaits [FS78b, Gro77f]. Paris [HM76a, Mal77, Mal78a, Mal79c, Tem76c, Alo79, HM76b, HM78a, HM79c]. Paris-Sud [Tem76c]. Park [GR75b]. Part [Nov77a, BP79a, Sch75l, dP79, vO75a, Acc75a, Acc75b, Acc75c, Law75b, Sta78b]. Partial [Cha78a, Cha75c, ES76, Got78b, MWW76, Ros78d, AB75, Axe79, BZ78, Bit76, CCMS75, Düm79, Džu76, Fef75c, GM77c, Gra76a, Kat75, Neu79, Rab78a, Rau75a, Ros78i, Wei75a, Wir78, You75, Gol75]. Partially [Bec78b, Naz75, Che78b, Dav78a, PR77c]. partially-observable [Dav78a]. participants [And77a]. Particle [Geo79i, Gri79a, Mal78c, Ne’77, Ros78h, Tul77b]. particles [Cie75b, Geo79e, Lig77a, Sza75]. Particulier [Poé76b, Ste76b]. partiellement [CDI78]. partiellen [Gro76]. Partielles [Lai78, Tho79a, Bon75, Bon78b, Krá76b, Las78, Tar78]. partitionable [DB78]. Partitions [DGS76o, Ste78f, Buc78, Hel79b, Jam78m, Ros76a]. partout [BG78d]. parts [Bak79b]. Passage [SW78g, SW78e]. Pasternack [Gar78]. Pasteur [Foa77a]. Patch [Bra79s]. Path [AHK76j, LY79, AHK76d, AHK76e, Gaw77, Ril79]. paths [AM79a, BK76c, Bos78, FS78a, Fer78a, Sha75]. Pattern [Bre76b, Sat76b, HM77]. patterns [Ber76e]. Paul [Mal77, Mal78a, Mal79c]. Peaceman [Ale78]. Peano [HW75a, KP77]. Pedagogic [Chu77b, KMS75]. Pelczynski [BCD77]. penalty [Cza75, Kaw79b, Har75a]. penalty-functions [Cza75]. Penalty-Methoden [Har75a]. pencil [Tat75]. Percolation [SW78g, SW78c, SW78i, SW78k]. Perfect [FJ75o, Gre78b, Tem76b]. perfectiness [KL79b]. perfectness [Mus76]. perforée [Duv77]. period [Cat79, Con79d, Pop77j]. Periodic [Alt79a, Bat78, Con79e, GM77k, Nus79, Ang79, BY77, Bib79d, Cop78g, Cus79, Dui75, DR75i, Eas75, Koß78, Kös76b, Nor78a, Sch79i, Wan79, vdAS79, Mei76]. Periodicity [BJM78a]. Périodique [GR77, Duv78]. périodiquement [Duv77]. périodiques [Bes75b]. Periodische [Mei76]. periodischer [FM76b]. permanence [Fey79b]. Permutation [Ker75e, tD79k, EH78, HP76b, HS75c, Wei76d]. permutations [HR76, HL77]. Permutationsgruppen [Kle77j]. Permutative [FP78k]. Permutrization [EK77]. Perron [Sch79j]. perspective [GM77a]. Perturbation [Cla79b, Gal75, Ral79a, Stu76a, BL75, BN77, Cap79, Dem77b, Kat76, KRZ77a, Mil78a, Moe79, Nas79a, Sta78c, vH79a, vdB79]. Perturbations [BLP77, BGM77, GM77o, Hes78, Hue77, NT79f, DH79, GM77f, GM77e, GM77j, GM77l, Gra79a, KS77, LR78b, Lum77, O’M78, dG79]. perturbed [Het79, Jor79, Noh79b, vH79b]. Peterson [Boa79, Lan79a]. petit [Dem77c]. petits [Des79b]. Petresco [War76c]. Pettis [Mas76]. Pfaf [Jou79d]. Pfaff [Gér78, GS79b, GR79c, Jou79b, Jou79e, Jou79d, Jou79g, Kla79, LV75, Sec75]. Pfaffian [Gér77]. Pfister [Bae78c]. Phänomen [Jur78b]. Phase [Föl75, Cla79b, MS75, Sjö75]. phases [Mol78]. phenomena [Sat76b]. phenomènes [Lév77]. phenomenology [Bre75c]. phenomenonological [Saf77]. Physical [Sat79f, Iag75b, Sco76, Ste78g]. physical-region [Iag75b]. Physics [AK78, BR77c, BRP78, Pha75b, Ent78, Grm79, HvW79, Joz79, Ne’77, Sij79, Sne75, Tak78a]. Physique [MT77c, BT77]. PI [Gla76g, Jac75d, Jac75e]. PI-Algebras [Jac75e, Jac75d]. PI-flows [Gla76g]. Picard [Bou78f, Bou78e, Bou78g, Mag76]. pictures [Lan77]. Piecewise [Bru76a, Hof79a, Kow75]. Pierce [Sze76]. Pierre [Ste75c]. pivot [Eav79]. PL [Bru78q, Mun79]. places [Gel76d, Gel76e]. Plain [Smi76f]. planar [BK77g, Lit75, Lit76, Lit77b]. Plancherel [Cha77]. Planck [Rau78]. plane [ADB75, KV79a, Lei79, Sil76d]. Planes [Moi77b, Her77a, PR77c, Rah75]. Planetenatmosphären [Dic75]. Plans [BR76d, Hau78a, dlH79c]. Plans-sur-Bex [BR76d, Hau78a, dlH79c]. plaque [Duv77, HPS77m]. plaques [Duv78]. plasma [BT77, Sij79]. Plasmas [MT77c]. plastic [Chi76c, Zie75]. plasticity [Str79a, Str79b, Zie75]. plate [ADB75, Mar79a, SNWC77g, Wah79]. Platek [Mol77m]. Plates [Mol77v]. Platitude [Hen75b]. Platonov [Tit78]. plats [Cou79]. playful [Bec78b]. ple [Hit78]. plethysms [Hof79b]. plongement [Gau75]. plumbing [NR78a]. plurisousharmoniques [Kis77b, Lel78a, Ste75b]. plus [Hau76]. Plusieurs [Nor75, Nor78c, Ram79a, Wal76, Wal77b]. Po-Hsiang [Bar79c]. poids [Lac78, Lap79, Ser76]. Poincaré [And77b, Cam75, Frö79a, Lie79, Ne’78, Roc75a, Roo79, Gro77d]. Point [Pre76c, AdS75b, AdS77, BG78a, BS78, Ber77a, Blu75, Bro75g, CK75, CS76b, Dab79, Dal76, Dem75, Dol78, Ede75, EGS75, Eva76, Fer76b, GM77m, KLŠ76, KRZ77b, KRZ77f, KRZ77i, Kel75b, Kik79, LP79a, Mat78a, McC78b, Mil78a, Per78a, Sha78a, Sha78c, Sto75b, Urb75a, Vai78, Wal75g]. Pointed [DS78f, DS78c]. Points [MS76c, PW79, Ser79b, Bat78, BJM78c, Bjö77, Bra76a, Chi78b, Dei79, DB77, Far79, FJ75j, For79b, Fra77b, Fuj79, Gér75g, HW75o, Ike78, Kos79, KK79, Mah76d, Mah76f, Mal79a, Mas75d, Maz77b, Pir77, Sch76r, Sch75n, Sto75a, Van78a, vdLT79]. Pointwise [Akc79, GK76b]. Poisson [GK78e, Gra76e, Gra76g, Gra76i, Hac78a, Kom75, Loh75, STW78]. Pol [Con79g, Hal79i]. Polaires [BWK77j, Nov77b, Nov78]. Poland [CUW75, MSZ76, Wer78]. Polar [BWK77j]. polarisations [Sim75b]. Polarité [Smy78a]. Polarity [Smy78a]. polarizations [Bla77]. polarized [Pop77i, Pop77f]. Pole [GW79d]. poles [Pin79]. Pólya [FJ75p, Sch75m]. Polyédriques [Kui79]. Polygons [Grü75a, Var78]. polynôme [Mal75a]. Polynômes [Duf79a, Bas75, Fer77, GP79, Les78, Bon76]. polynômiale [Kaj75]. polynômiales [Nac78]. Polynome [vG76]. Polynomial [Pas79g, Pas79h, Bru76a, Gil75a, Jac78a, Lyc76, Pfl78, Rob79a, SW78a]. Polynomials [Gro78e, Aro76, Bak79b, Bor79a, CMPY79, Far76a, Far76j, GL78, Gro78f, Gro78r, KRZ77c, Koo77, Mag79b, Sch77c, Sch75k, SS76d, Suf76]. polytopes [God78]. ponctuel [Bré76a]. ponctuels [CM75, Jol78, Nev77, Reb78]. Pontifica [Sch78g]. Pontrjagin [Mac78]. Pontryagin [Bin77, SW78a]. population [AW75, Gan75, Gra79a, Par79b, Pel78a, iS78, Wei78a]. poreux [Lév77]. porous [APF76, Cha76d]. posed [Shi75]. posedness [CG79b, Mil79]. Posés [Tik79]. positif [Ber76c, Bop79, Bru76c, Gol78a, Nic75a]. positifs [Lel77b, Sko78]. position [BR75, Bin75b, McC78c]. Positive [Cen76, CL75e, DR75d, FJK79, KS79b, Clé79, Dui75, Fal79, GHM78, Gór78b, GW79j, Hay76, Jac78b, Ros78h, Sch75l]. Positive-definite [DR75d]. positives [Éli79a, Sjö76b]. Possess [GW79d]. Possible [Tay78g]. Postnikov [Bau77f]. postulates [Blu75]. Potential [AL76, CL75d, CL75f, Fuk76, AHK76d, BF79a, CW76b, CL75g, DG76c, FR76, Jac79a]. Potential-theoretic [AL76]. Potentiale [RR75g]. Potentials [Sai79e, KSWW75, Ros78h, Wah76]. Potentiel [HM76a, HM76b, HM78a, HM79c, Bis78a, Éli79b, Le 79]. potentiels [Anc78, Bre78, For78, de 78]. Potenzreihenringe [KPP^{+}78a]. Pour [Bea78e, Cha77, DC78, Jac79g, Lau75c, Anc78, Bai77b, BG78a, Bar77a, Boc78, Boh75, BCH78, BT77, Bou78d, Cai78, CL76a, Cle75, Con76b, CL76b, Del77h, DT77, Duf79a, Fit78, FS78b, GS79a, Gat77, GM77o, Gér79b, GB79, Gol78a, HL78, Hir79b, JM76, KS78b, Le 78, Le 75, Lév77, Loe75, Mat78b, MS79b, Mét78b, Mul75a, Mul79, Ren77, Roo78b, Ver76]. power [Bra79l, Con77, Gil75b, Hof79j, Kis76a, Mah76e, New77, Nie77, Tam78i, Tam78l, Tam78m]. Powers [KRZ77d, AF78, BG76a, EG75, KRZ77b, Pac76, Ros78h]. practical [Ben76a, Est76, Wei76n]. Prädiktoren [Man78c]. Prague [Fáb79, Nov77a]. precedent [Str79d, Ber77b]. Prediction [Nie78, Wer75, Bre77g, Hac78b, Mey76c, YM76, dSLM75c, Yor77]. predictor [Cla79g]. predictor-corrector [Cla79g]. Preface [G.75, SNWC77k]. preimages [Eis76]. Préliminaires [BWK77k, Sch76j]. Preliminaries [And77c, Bar79g, BDE^{+}77h, Beh79l, BJM78d, Bru78k, Dub79e, DS78g, FGR75g, GK78g, Get75f, GW79i, HK78h, Ish76g, Mic78i, SW78j, Wal75h, ACH77b, Blo79e, FR79h, KRZ77e, Sat79e, Tay78f, BWK77k, Sch76j]. preliminary [Dav78h, Ste78h]. Premier [Jac79p]. Premières [Bou79b, Suz78a]. Premiers [Ser75h, Ser75i, Dix75, MN75a]. Preparations [Sto77e]. Preparatory [Sch79c]. prequantization [Kos77]. prescribed [BC78b, Far79]. Presence [Poé76b, Poé76b]. Presentations [RR79b, Rou79, Lal77a, Rog79b]. presented [Bie79, CS76d, RS76]. preserving [Alp79, AdS75a, BHP78b, CN77, Fel79, Fro79b, Han75b, Pla75, Sto77h, Wal75f, Wal75g]. Presque [Mol77s, Bes75b, BG78d, Fey79a, Hec78, Mol75]. presque-connexions [Mol75]. Presque-structures [Mol77s]. preuve [Bru76d]. prevalence [Lor79]. previsible [Get75a, Jac78b, El 78, MS79b]. primaires [Thi77]. Prime [Sun79, Tep76, vO75c, BY77, Con77, Fai79a, FP78b, FP78d, Ish76b, Lev79b, New77]. prime-power [Con77, New77]. Primes [vO75d, FP78a, FP78c]. primitif [Ren79b]. primitifs [Dix75]. Primitive [CV77i, BF79b, Bru78m, Buh78g, CF79, Han75a, HP76b, Jos77, Jos79, SV75c, Wei76q]. Principal [Bau77g, Hoc79c, Sri79g, Bau77b, Bau77c, Bau77h, Car77a, Cla79c, Dem77c, Dup78e, Eva77, HL78, Ren77]. principale [Clo79, Meb75, Sch79g]. principaux [CH78c]. Principe [Anc78, Lum78, Sib77, BP79b, Cha79b, Lum76c, Rit76]. Principes [Hir79b]. principle [Ber79d, Che76c, CL75b, DGS76u, Ger79a, Hal77, KK78b, Led79, May77k, Rab78a, Ros76a, Sai79c, SS76e, Wil75g]. principles [Eck77, Mau79, Phi76, Phi79]. Prinzip [Küh79, Wec75]. priori [BK76e, Don76]. priority [AH76]. pris [Anc78]. Prison [Zee76e]. pro [EH76a, EH76g, Mar76c]. pro- [EH76a]. Pro-categories [Mar76c]. pro-spaces [EH76g]. probabiliste [Pra78]. probabilistes [Mey77b]. probabilistic [Nag76, Pin75a]. Probabilité [Jac79d]. Probabilités [Cau76h, DMW77, DMW78, DMW79, Hey78, Mey75d, Mey76h, And77a, BKK76, Bad76, DCHRH78, FCG75, HJLN77]. Probabilities [Gus76a, Bor78a, Bor78b, Che78b, Dub76b, Goo78]. Probability [BKK76, Bec76b, Bec79a, CUW75, DCHRH78, Fei78k, FCG75, GK78d, HJLN77, HJ77, MP76, Bak78a, Ben78, BF79a, Fel79, Gal76i, Gli78, Haz77e, Lau75a, LS79b, Mac75b, Muk79, Sat76a, Sat79a, Urb75a, Urb75b, Bec76b, Bec79a, Cau76h, CUW75, Hey79, Wer78]. probability-preserving [Fel79]. Probability-Winter [CUW75]. Problême [KS78b]. Problem [Dev78, Fel76, Fis76h, Hee76, Hei76a, Hör75, KRZ77d, Mol77-28, AW76, Bab79, BK76g, Bon78a, BN77, Bru78l, Cap77, Chr76, Chr78, Cot75, Das75, Dev77e, Dou77b, Dra79, Eck75, Ede76, Ess75a, ES77, FI77, FJ75i, FJ75r, FJ75s, GR79g, GR79f, GR79h, GP77, Gri76b, Gus75, Han79b, HvW79, Hit78, Jer76a, KY79b, KRZ77c, Kaz78, Kra78a, Kru79, Lau75b, Leh79, Lub79b, Man76b, Mat78a, MT77a, MS79a, McK75, Mil78a, Miy77, Moe79, Nag76, Nir75, Nun77, Oxl79, Pel78a, PR77b, Pfl75c, Pfl76, Roa75, Rog79a, Rue76, Sal78, SV79, Sch78b, Sha79a, Sij79, Sta78b, Str79b, Sus79, Wei76u, Wil76b, dF75, dG75g, Pfl75c, Pfl75a]. Probleme [Kup76, Gil78g, Jac79p, Jac79r, Lum76a, Lum76b, Mol77s, Mol77-28, AY79, BG78a, Bar77a, Boc78, Bon78c, BT77, Con76b, Dem77b, DT77, Eis77b, Kor76, Le 75, Lel78a, Pue78, Rab78b, Sab76, Ste78e, YM76, Wag78]. Problémes [MT77c, DC78, Jac79b, Jac79h, Jac79l, Jac79n, Le 79, Lio77b, MR77m, Nic75b, Tik79, Bai76, Bis78a, BEGS75, Che75a, GM77o, Gos78, Hec76, Hes78, Hue77, Lal77a, Dem77c, Gri77, Lel78b]. Problems [BK76f, BTW75b, EW78, Erd75b, Erd77, FW79, Far76j, Fis76i, GN76, Hof79i, Øvr76a, Tam78c, Web78c, Åst78, Adl79b, Ama76a, Bab76, BR77a, BM79a, BR76a, Bar79h, Ben76a, BL75, BF76b, Blu79, Boh78, BG78c, Bör75b, BGJN76, BTW75f, Bri79, Caz79b, Che76a, Che76c, Chi76c, Chu77a, Col78, Col76d, Cor79d, Cot79, Cox78, Csá77, Dah76, Dav75b, Dej78b, Dev77d, Duk78, Düm79, DS78k, Eas75, Erd78, FR76, FJ75b, FJ75m, FJ75q, Fuč79, GM77a, GM77c, GM77m, Gen76, Geo77, Geo78a, GK76a, GS77, Gri76a, Gro77a, Grü75b, Hal78a, Han78b, Hec76, Hes79, Hla79, HS75c, HM77, Jer76b, Juh77, Kaw79b, Kel79, Kik79, Kno75, Koß78, Kra79, Kur75, Lio75a, Lio75b, LP79a, Loo75f, LS79b, Lum77, Man76d, McC78b, Mil76, Mil79, Mos76a, Nar78, Neu79, New76b, OL77, OW78]. problems [PP79, Per78a, Pop77a, Pop77j, Ral79a, Rat78m, Rau75a, RT77, Red79, Rog77a, Run76, Sbo79, Sch79h, Sch75c, Sch75e, Sch75g, Sco76, Sei75, Shi78, Shi75, Sle75, Slo76, SW78d, Str79a, Stu76a, Suz79, Tak76b, Tur78, Wal78a, vH79a, vH79b]. Procédé [Bea78b, Bea78g]. Procédés [Bre77m, Cor79b, GB79]. procedure [Cap79, FR79e, Mar79b]. procedures [Brü79a, Čud77, FR79d, Gal75, Pra75, Sar78, ZKB79]. Proceedings [AL78, AT78, AK78, AHW77, AD78b, AD78a, BH78, BCD77, BM78a, BM78b, Bec76b, Bec79a, BK76a, BR78a, BH76b, BK75a, BR77c, BRP78, BMS76, BGM77, BCN77, BS77, BOS75, BGS78, CV79, CW76a, CCJS79, Cha78a, CUW75, Cop78i, DCV78, DJ79, DM75c, DG75j, DKT78, DM79b, Eve75a, ES76, Fáb79, FW79, Fen79, FMS79, GM77n, GR79b, GR75b, GM79d, HL79b, Hau78a, HP79, Hey79, Hil76, HPS78, HS79b, HS78, HW79b, JMR75, KK78a, KM78b, Kel75c, KZ76, Lai78, LLS79, Lit77a, LS79a, Mal78a, Mal79c, Man75d, MMP78, MWW76, Mil78b, MS77, MOP75, MS78d, Nat77a, Nat79, NR78b, Nov77a, Ols78b, PdC77, PW79, Pin75a, Ros75b, SS76a, SZ77a, Sch78g, SZ77b, SZ77c, Ste76c, SW75d, Tem76c, Wat76, Wat78b, Wer78, Wuy79b, Zel76b, MP76]. process [CC76, CB76, DV75c, Fel77a, Fuj76, GK78e, Neu76, Per78b, Rév76, SW78l]. Processes [DKT78, GR76, Get75e, Get75h, Get75g, Gra76e, Sil76b, AM79a, BL75, BLP79, BD78, CW77c, Car79d, Cha78c, DV75c, Dre77, Dyn76, Fei78d, Fei78j, For76a, Fuk76, Get75c, Get75m, Get75n, Gri76a, Gus76a, Hal76a, Hit78, HC78, Hud78, KLŠ76, Kaw76, Kin76, Köh75, Kom76, KT76, Kry78, Kue79, Kun76, Mal78c, Man76c, Man79, Mey76b, MS78b, Mla78, Ood78, Pre76c, Sch78e, Sil76f, SW78e, SW78f, SW78k, Sta78d, TW78b, Zab75]. Processus [Ben75b, DC75a, DCV78, Gan75, Jac79o, Nev77, BDR78, Bré76a, CW77b, CM75, EM77, Fer76a, Jol78, Mey76i, Mey77a, Reb78, Ruc78b, Tua78, Web77]. produced [Kud77, MJ75]. Product [Cla79h, Hof79d, Kad78, Ada75, Car78b, HPS77i, Mic78j, Oma75, Sto79, Tak78d, TH75, Wan75, vR79]. productive [BH79]. Products [DS76c, GM76b, Mic78g, MT76b, NT79c, Ros78g, Won79d, Ber76d, Cig76, FR79c, Fis79, Gue78, Hae78b, Hof79k, HSM77, Jak79, Ker75b, Ker75c, Lew77, MN79, MMS78, McD78, Mic78k, Mor77a, NT79d, NT79g, Tay77b, War76k, Won79e, Won79f]. Produit [Pes76a, LP79b]. Produits [BWK77b, Fd79, Let76, Lal77b]. Produkte [Voi77j, Voi77t]. Produkten [Voi77b]. Professor [Sza78]. profile [BH76a, CV77j, Nør78d]. Program [Gol75]. programme [Loo75f]. programming [Law76, Man76d, Rei76b]. Progrès [Des79b]. progress [Nar78]. progressions [Tho79c]. Projecteurs [Aub79]. projectif [GP78b, Jou79e, LBL78, Szp76]. projectifs [Bas75, BBL77, Fer77, Jou77c]. Projection [DM75b, Jac78b, GM77b, LY79]. projections [BM77a, BDE^{+}77a, Beh79g, CSW76, Moi77d]. Projective [AEK78n, Hum76e, Moi77b, Pet76b, Pet76f, Sac75, Buh78f, Dav79, Guz77, Jen77e, Kec78a, Mic77, Mic78j, Mor77c, Nam79b, Oso79, PR77c, RT78, Sus79]. projectivity [Rei77a]. Prolegomènes [Bad76]. Prolongement [Bon78c, CW77b, Hen75c]. Prolongements [Ber79e, Dlo78b, Dlo77, Dlo78a]. Proof [BZ76, DM75c, Ess75i, Loo76, Ben75a, Doh75, Fri78d, KS76, KM77, Luc75, Mas75a, Sha78e, Ste77a, FK79c]. proofs [Lam78g, Tak75b]. propagating [Lei79]. Propagation [Pir78, dM75a, APF76, AW75, HL78, HBT75]. Propagators [Mas78]. Proper [EH76h, FF78, Hei76b, Jen77g]. Properties [Gro78p, Sch75n, Wei76o, Wei76p, Wei76q, Alp79, Beb79, Blu75, Bor79c, Cun75, DNR79, Dev75, Dub79a, DS78a, DS78c, DS78d, Fer78a, FR79k, Fox79, GK78b, Gra76d, Gra76g, Gre75a, Gre77a, Gro78d, Hal78b, HLS79, Joh79g, Mah76c, Mor77a, MG79, NT79d, Pel77, Peł76, Raw78, Ree76b, RB75, Ros78j, Sch75p, Ste78c, Suf77, Wam79, Wei76a, Zie79b, Zie79f, dG75b, dG75f, BWK77l]. property [Beh79a, Ber76f, Bie77, CV77j, Cha79c, Cic77, Edg78, Egg78, FT76, FGR75c, Gag77, Hud78, Iag75b, Kel75b, Man76c, Rob77a, Sai78]. Propos [Bur76, DM78b, Ram79a]. propositional [GT75]. Propre [Bou78d, Pue78, Ste75c]. propres [BLC78, Lac79, LR78a, Ser75p]. Propriété [Vai78, Foa77c, Gol78a, Pir78, Sjö76b, dlH79b, Roy76a]. Propriétes [Wal76, BF75f, Bea78d, Bea78f, Bea78h, BWK77l, Fey79b, Sig75, WK78, Bré76c, DMY78, Éli79b, Gol78a, Hir76b, Jou79f, Lel78a, LV75, Sec75, Aup79e, Aup79f, Wal77b, dV75a]. prototypes [Bre76b]. provability [Poh75]. provably [Fri78c]. provenant [BT77]. Proximal [Gla76h, Gla76i, Gla76j]. Prozesse [Mét77d]. Pruning [Hag78]. Pseudo [Cor79e, Don76, Fai79a, HPS77m, Kaw75, Adl79a, Can75b, Cho76, Cor79f, CDI78, Dem77c, Fey78a, Gla75a, KK75, Las75, Lel75a, Sib76b, Wag77, Nov75]. pseudo-convex [Cho76]. pseudo-convexes [Lel75a, Sib76b]. Pseudo-convexité [Nov75]. pseudo-différenciels [Dem77c]. Pseudo-Differential [Cor79e, Kaw75, Adl79a, Cor79f, KK75]. Pseudo-Differentialoperatoren [Don76]. pseudo-différentiels [CDI78, Las75]. pseudo-flattenings [Can75b]. Pseudo-Frobenius [Fai79a]. pseudo-isotopies [Wag77]. pseudo-noyaux [Fey78a]. pseudoanalytischer [Bau76]. pseudoconvex [Pfl75c]. Pseudodifferential [Øvr76b]. pseudodifférentiels [Hel78a, Sjö78]. Pseudogroupes [Mol77t]. pseudokonvexe [Pfl75c]. Pseudokonvexität [Pfl75d]. pseudometrics [Rob79b]. pseudoparabolic [GH76]. Pseudopotentials [CT76]. publié [Ste75c]. puissance [Kis76a]. puissances [Hir76b]. pulse [AW75]. Pura [PdC77]. pure [Bar78b, Ish76f, JR77, Rei77a, Wal75g]. purely [Yan77m]. pures [Cou78]. Purkinje [BBDB^{+}78]. Pushchino [DKT78]. q.c [Sch75f, Sch75r, Sch75t]. quadrant [Tho79b]. Quadratic [Bae78d, Bae78e, AHK76g, Bae78a, FJ75d, Gol79, Hir77, Lam78d, LT76b, Ran79a, Sun79, Zag77b, tD79j]. quadratique [Art75]. quadratiques [Bis78b, Vig77b]. quadrature [Gau76]. quadraturformeln [Str76]. quadric [Far79]. quadrilaterals [HWH78]. Qualitatifs [Tho79a]. qualitative [Rau75a]. quand [Bré76a, Rhi75]. Quantifier [Kra75g, BW77a, Cun75, Gre75a]. quantifiers [KK76, Slo76]. quantile [Par79c]. quantique [Kré78b, Ler78]. quantiques [Sou77]. quantisation [Sim77, Sim78a]. quantitative [Nie75b]. quantization [Czy78, Gaw77, Kij77, Śni77, Śni78, Wei75c, Woo78b]. quantized [May77l]. Quantum [Ros78h, AHK76c, EGS75, Gli78, Hol76, Kas75, Ler78, Mos77b, Śni78, Sto75b, Urb75b]. Quasi [Bib79d, BP78, Fey78b, GW79j, Kat75, Kup75, Mañ75b, Rei77a, AHK76i, Ali75, Bai76, BBL77, Dam78, Dra79, FdlP79, GP77, GH75, Kos75a, Lio75a, Mos76a, Mos78b, Ste76b, Ste76a, Str77, RR75f, Ste76a]. quasi- [BBL77]. quasi-analytiques [Ali75]. Quasi-Anosov [Mañ75b]. Quasi-coherent [Ste76a]. quasi-conformal [Dra79, RR75f]. quasi-direct [GP77]. Quasi-expectations [BP78]. quasi-free [AHK76i]. Quasi-Frobenius-Algebras [Kup75]. quasi-invariant [Kos75a]. Quasi-isometry [GW79j]. quasi-linéaires [Dam78]. Quasi-linear [Kat75]. quasi-multiple [Str77]. Quasi-periodic [Bib79d]. quasi-pure [Rei77a]. Quasi-pure-injectivity [Rei77a]. quasi-resolvantes [FdlP79]. quasi-standard [GH75]. Quasi-topologies [Fey78b]. quasi-variationnelle [Mos78b]. quasi-variationnelles [Bai76]. Quasibarrelled [AEK78o, HK78c, HK78f]. Quasibounded [GM77l, GM77k]. Quasiconformal [Por79, Rei76a, Ahl77, Ber79c, Car79b, Gau79, Ghi79b, Leh79, Per79a, Sbo79, SS76c, Sch75p, VGK79]. quasiconformality [Caz79b, Ghi79a]. Quasiharmonic [SNWC77l, SNWC77e, SNWC77i]. quasikonforme [Küh79, RR75f]. quasilinear [Gre75b]. Quasimartingale [Mét77f, Mét77e]. Quasimartingales [Let79]. Quasiregular [MR79, Sre76, Vuo79]. quasisymmetric [Iva79b]. quasitopoi [Dub79i]. quasiultrabarrelled [HK78d, HK78e, HK78g, HK78j]. quaternion [Bra76a]. quatre [Fou79]. Queensland [SW75d]. Quelconque [CL76a, Maz77a]. Quelques [Ars79, Art75, Bea78g, BP79b, Bré76c, BEGS75, Che75a, Che78c, Cle75, Éli79b, Gér75f, Ger75i, Gil78h, Hei78, Hel78a, Jac79l, Jac79m, Jou79f, Kis76b, LV75, Mét78b, Tak79, Ver77, Yor78b, Yor79, Hec76, Jou77a, MvM77, LS76]. quenching [AW76]. question [Edw75a, FK79e, Hal79g, Sta78a]. Questions [Gal76g, Mey75c, Bör75b, EH76i, Eps79f, Moi78, Rat78n]. queueing [AH76]. qui [Fit78]. Quillen [Bre75a, Eis77b, Fer77, Fie78, Hau76, Lam78e, Qui76a]. Quine [Pét77]. Quinn [May77a]. Quivers [Bre75c]. Quotient [Dub79f, Rob77d, Fai79b, Len76a, Pop77b, Pop77i, Rob76e, War76b, Wei76f]. Quotients [Pop77k, Pop77l, BW75, CB75, LL76, Pop77e].

R [Del77e, Don76, Ros77, Sib77, Tou76]. R. [Gér79b, HW75r, Lat79, Moi77f, Ram79b, Ray78a, Ste78e, Why78]. rôle [Buc78, Fd76]. Rachford [Ale78]. racine [Loh75]. racines [Mat77f]. Rademacher [Kwa76, Peł75]. radial [Lau75a]. radiating [BK76g]. Radiation [BK76f, ZKB79]. radiative [Sai79a]. radical [Fai79a, Hau77]. radicals [War76d]. Radon [Cha79c, Chi76a, Die75d, Edg78, Egg78, Gol78a, Gue78, Hac78b, Pet78d]. Raikoff [Haz77g]. Rajkundalia [Seb79]. Raman [Chu76]. Ramanujan [Ran77]. Ramifiées [Gér75b]. ramifiés [Gér75h, Hen75b]. Random [Buk77, Far76k, Fer78a, Kan78b, MP79b, MT76b, Pin75b, Pre76h, Pre76i, Tay78h, Tay78j, Ave78, BW76, BP79a, Cré76, DJ77, Gie76, GMZ79, Gra76c, Gra76f, Gra76i, Gra79a, GH75, Grü75b, HvW79, Jur75, Kue76b, Kue76a, Kur78, Lei79, LR79, Mar76a, MBY79, Mla78, MT76c, Nie78, PT76, Phi76, Phi79, Ros76a, Ryc78, Saw75, Tay78e, Wan75, Wil77b, Sza75]. Randomly [Tay78i]. Randwert [BG76b]. Randwert-Normproblem [BG76b]. Randwertaufgaben [Col76c]. rang [BDR78, Jac77a, Tak79]. Range [Sai79e, CSW76]. rank [Dek79, FJK79, Goo76a, Hub78a, Koo77, Lac76, Leh76, Mah76f, Ree78b, Wei76f, Wei76p]. Rankin [And77k]. ranking [Wil77b]. Rao [Hol76]. Rappels [Gil78i]. Rapport [Del77f, Gra77, Anc78, Ren77]. rapports [Bre79f]. rarefaction [Reb79]. rate [For76a]. Rates [Kue79, Cos75, Wah79]. Ratio [Kle75, Ish76i]. Ratio-sequences [Kle75]. Rational [Cla79g, LT76a, Maz77b, Sta78c, Yan77q, Bru76a, Cor79c, Mor77c, Pin79, Spe75b, Wam79, Wer79]. Rationale [Bla76i]. rationalité [Hou77]. rationality [Gal76g]. rationnelle [Gra76b, Tua78]. Rationnelles [Jou79a, Dab79, Ker75f]. rationnels [MS76c, Ser79b]. ratios [Bal78]. Raum [Don76]. Räume [FK79b, GMK75f]. Räumen [SS76b]. Ray [Get75e, Get75j, May77a, Get75k, Nix78, FdlP79, Get75i]. Rayleigh [Hal79i]. Raymond [Mil78b]. rayon [Coe77, Kis77a]. reacting [AG76]. reactions [Gud76]. Real [BDE^{+}77b, Bro75d, Bru78f, Gal76h, Var77a, AHK76f, AW78, Bab76, Bla77, FJ75j, Hir77, IP76c, Jen77a, Luc75, Sch75l, Sch75q, Ste78f, Sun79, Tog76, Too78, Var77c, Var77b]. Real-Generated [Bro75d]. Reality [HW79a]. realizability [Tai75]. realizations [Cla79a, EH79a]. Realizing [Ben78, Bru78q, Ewi79]. reals [KL79a]. rearrangement [DC75a]. recently [Rat78m]. récents [Des79b, Mey77c, Ris76]. Rechnen [Voi77b]. Recipe [MP78a]. recognition [Bre76b, HM77, May77k]. recollant [And75]. Recollement [Roo78c]. Reconstruction [Bru78m]. Recovering [Ban79b]. rectangles [dG75a]. Rectificatif [Ste75c]. Recurrence [Cré76, Gro78q, Mag79b, PR77b, Cré76]. Recurrent [MT76c]. recurrentes [Bro78b]. Récurrents [GKR77a]. Recursion [MN75b, Mol77o, Mol77n, Fen75, Hin77, HW75i, Kol78, Low76, Maa78, Mol77g, Mol77j, Mol77m, SW75c]. Recursive [Bul79, Fri78c, Han75a, Mey75a]. Recursively [Bör75b, BS75]. Recursivity [Lou78]. rédigées [Mol77-28]. reduced [Pac76]. Reducibility [Gar77a, GPR78, Cop78d]. reducible [Rob76a]. reducing [PS77]. Réductifs [Gér75b, Gér75a, Car77b]. Reduction [Van79, Bof77, Sje78, Tak78b]. reductions [Uen75b]. Reductive [Gér75b, Var77a, Obe78, Pop77k, Var77c, Var77b]. reduit [Bar75]. reduite [Mok78c]. Reduzibilität [Jur78l]. Reeb [Sch78a]. réel [And75, HL78]. réelle [Ava77, Mét77f]. reeller [Mét77b]. réelles [Blo77]. Réels [Bea78c, BWK77f, Duf79b]. reexamined [Bör75b]. references [Bre79e]. refinement [McC78b]. Refining [BV77]. reflecting [Ric75, Yam76]. Reflections [Bon78a, Huš76]. reflexion [Pir78]. reflexive [Ørn77, Sun75]. refractive [BK76c]. Regel [Sca75]. regeneratifs [JM76]. Region [DKT78, Hac78a, Iag75b]. regions [BB76, Jel78, Pfl75c, Wid78, vdLT79]. regression [Fri79, GM79a, MBY79]. Regressive [Ste78b]. Regular [Goo76a, HW75p, HW75q, vdE79, Ben76c, Dek79, Fis76h, HL79a, Hay77, Jam78m, SW75a, Sen76a, Wei76b]. Regularité [Fer75, Sim78b, Cor79b, DT77, Mey76i]. Regularity [LF79, BF76a, Fer78a, Ike78, Kom75, dP76a]. Regularization [Sze75]. Regularly [Sen76b, Sen76c]. regulière [Mos78b]. reguliers [Ren79a, Sch76h]. Reichweite [Dic75]. Reiner [RR79b]. Réiteration [Bea78b]. relate [Mar77]. Related [BWK77l, Dei77i, GK78d, Gol75, MR77h, BDE^{+}77c, Bör75b, Dej78b, Eck75, Egg78, Erd78, FJ79, Fra78a, Gal76b, HvW79, Hof79e, Hof79f, Hof79m, Joh79q, Kue76a, Mor79b, Par79b, Rat78n, Rog77b, Sch79h, Sil76c, Sus79, Vaa75, Zie79e]. Relating [Bra79o]. Relation [Bar78a, Che76a, Hei79, Ral79b, SNWC77m, Chr78, Čud77, PR77b, Sto77a, Wai75]. Relations [And77d, Gol78a, Gro78r, Kud77, Nov77a, Pes76b, RR78, Roo79, AF78, Can75a, Far79, Gil76, Gre76, Gro78q, Jam78h, Ład75, Pom76, Rah75, Sha79c, ST77b, Suz79, tM76, Gil78e, RR78]. relationships [Cha78c, Cli75]. Relative [And77e, Bau77h, Lau77, NT79g, AHK76g, Dre75]. relativement [Bop79]. relatives [Hei78]. Relativiste [Cau76e, Cau76f]. Relativistic [Cau75, AHK76c, Cau76e, Cau76f]. relativity [Szc77b]. Relativizing [Eps79j]. relaxation [Deu75]. Relèvement [Clo79, Spr75]. Relèvements [Gap76]. reliability [BK75c]. Reliable [Cir78, Wer79]. Remark [For79a, BG76a, Ber79d, Gus76b, Kaz76, Kaz78]. Remarks [Beh79m, Beh79n, BK75b, CS76c, Fre78, Gór78b, Iva79b, Lio75b, Oso79, Sel77, Sha77c, Bin75h, Fei78b, Gar77b, Gin76, Hel79b, Kai76, Lau76, Mos76b, Muk79, Sch79c, Sch75b, Seb79, Sin77, Ste78h, Wei76u]. Remarque [KV77, Bin75a, Hor77, Mey79, Str78, Str79d, Yor76]. Remarques [Mal79a, Del77i, Hei78, Kis76b, Mok78c]. Reminiscences [Cro75a]. remplissage [Foa77c]. Rencontre [Pha75b]. Renewal [SW78k, For79a, Rog79a]. Rengel [Caz79a]. Rennes [CK76]. renorming [Die75a]. Renouvellement [HT78, CL76b]. Rényi [CC76]. Reparametrization [Fel79]. Répartition [Rhi75, Rau75e]. repère [Fed78]. replacement [Sta75]. report [Bor78c]. represencation [Sze76]. Représentabilité [Bou78c, Bou78e]. representable [Mag79a]. Representation [FGR75h, Gel76g, Geo79i, Jam78n, Muj78, RR79b, Sch75o, tD79l, BWK77m, BWK77n, BWK77g, Bon75, Foa77a, Jac79n, Lac79, BDE^{+}77f, Bir79, Buh78e, CV77b, Cos79, DR75e, Evs78, Gab75a, Gam79, Jam78a, Kar77, KN79, Ker75b, Kup75, Man76b, Mét78a, Mor77d, Mül75b, Ols78a, Ral79b, Rei79, Rin75, Ruc78a, Sai79d, SNWC77n, Sat79d, All76, Cai78, CM75, JM76, Nan75, Per77, Rot76b, Yor78b]. Representations [Buh78c, DR75f, DR75g, Fel77c, Gal76h, Hof79d, Hum76c, Ker75e, RR79b, Sai79e, Sri79h, SV75d, Wol78, AM79b, Bam79, Bea78d, Duf75a, Gér75g, Gui77a, Ruc78b, Spr77f, Ars79, Bin75h, Blo79a, Buh78g, Buh78f, Buh78h, But75, Don75, DR75c, DR75h, EK77, Fel77d, FJ79, Gal76e, Gal76j, Gel76f, Har75b, Hof79k, Hum76b, Hum76d, Jac77b, Jam78k, KL79a, Ker75b, KZ77, Ład75, Lou75, Mor77a, Mos77b, Naz75, Obe78, Ply79, Qui76b, Rei78b, Rei78c, Rei78a, Rib77, Ril79, Rog79b, Sen78, Sri79c, Sri79f, Sri79g, SV75b, SV75e, SV75f, SD77, Tho78, Wal75a, War76l, Wil78b, Yan77a, Yan77q, tD79g, tD79k, CO79a, Car76b, Car77b, Cow79, Del79, Gér75f, Gér79b, Gui79a, Har75b, Rod75, Rub79, Spr75, Tak75a, Voi79, DG75j, Duf79b, Gab75b, Rod77, Ser77]. Représentations [Ver75]. representatives [EN79]. Representing [Ecc78]. reproducing [Del76]. Research [Cro75b, FMS79, Hil76, Miu76a, NR78b, Blo79g]. residu [Meb75]. residual [Pas79i, War76c, CH78d]. residue [Pas79f]. Résiduels [CH78d, And79b]. residues [SW78a]. résidus [CH78e]. Resistance [Pow78b]. Résoluble [Cha77, Sab76]. Résolubles [GKR77g, Gér75a, Hel78a, Ren77, Ver75, Voi79]. resolues [Fey79a]. résolution [Tar78, Rab77]. resolutions [Eis77a]. resolutive [Ike78]. résolvante [BLC78, Rit76]. resolvantes [Bro78b, FdlP79, For76b, HT78, Kis76b, Fey78a, Lum76c]. Resolvent [CL75d, Clé79, CL75f]. Resolvents [Get75l, Get75j, Gro79h]. resonance [Abr76a, Gos78, Hes78, vdAS79]. resonances [SV79]. resp [Gol78a]. respect [AHK76b, AHK76a, Ber79a, Bre79b, Fri78b, Ghi79b, Mét78a, Wei76j]. restart [vdLT79]. restricted [LEV75, Rob76b]. Restrictions [Sax77]. result [For79a, Kat77, MT77a]. Résultant [Jou79a]. résultat [ÉS79]. Résultats [Mey77c, Nie75b, Tik79, Art75, BDR78, Che78c, CH78c, LS76, Tak79, Ver77, Yor78b, Ris76]. Results [Eas75, Ale78, Ang79, Bak79b, BDE^{+}77c, Byc78, Chr78, Cla79f, Dav75b, DR75a, EWW77, EV77, Erd75b, Erd77, Eva76, GK78f, GK78g, Gel76b, Geo79g, GM76c, Gra76j, Háj78, Hen79a, Hof79m, Kar77, Kuo75d, Lak79a, Lam78a, LT76c, Lev78, Mah76d, Mal79d, Mar76a, Pet78a, Rat78e, Sri79i, Ste77b, Tho77, Van78c, Wei76h, vW76]. Résumé [Jac79o]. Retarded [Hal79a]. Reticle [BWK77a]. Reticules [BWK77a, BWK77b, BWK77m, Mok78a, Fey78a, BKW77]. Retour [Bou78f]. retracts [Arm75]. revêtements [Hen75b]. reversing [NR78a]. Review [Nam79f, Sri79i, Pet78a]. revised [McC78b]. revisited [Li79]. Reynolds [Dar77]. Rham [MM75a, Dup78i, LBL78, SS78a]. Richardson [Jam78l]. Richtungen [Jur78o]. Riemann [Acc75d, Bom76, Ghi79b, Hej76e, Ike79, Nam79d, Nam79e, Roc77, Suz79, Ver76]. Riemannian [Ati75, Bla76c, GKM75, LF79, Mal75b, SNWC77e]. Riemannien [MM75a]. riemanniens [Kos76c]. Riemannigfaltigkeiten [GMK75f]. Riemannsche [GKM75, GMK75g]. Riemannschen [Vek76]. Riesz [Ben77a, Bre78, FJ75c, RR75g, SNWC77n]. Riesz-Potentiale [RR75g]. Right [BJM78a, Get75e, Get75m, Get75n, HB79, Sha75]. rigid [Glu75, She75]. rigid-like [She75]. Rigidité [Kos76c]. Rigidity [Pal78b, Liu78a, Liu78b, Liu79a]. Ring [CS76d, FGR75i, May77a, Rat78c, Bra79s, Coh79b, Fai79a, Gil75b, GH78a, Goo79b, Gre76, Hir77, Hof79h, Mah78, May76a, May77k, May77c, May77b, May78a, NV79c, Par76, Pas79a, RR78, Tou76, War76i, tD79b, tD79c, tD79f, HL79b]. Ringe [Kra75d, KPP^{+}78c, KPP^{+}78b, KPP^{+}78d, KPP^{+}78e]. Ringen [KPP^{+}78i]. Ringkanal [Fri76b]. Rings [Bae78e, Bra79d, Hal79c, Hal79h, HW75j, Hof79c, Hof79b, NV79d, Pas79d, Pet76b, Pet76g, Wie79, And79a, Bae78c, Bae78d, Bae78f, Bea79a, BWK77a, Bof75, BF79b, Bra79c, Bra79f, Bra79g, Bra79h, Bra79k, Bra79l, Bra79q, CF79, CM78b, Che76f, Chi76b, Dre75, Eis77a, Fai79a, Fai79b, Fis76h, Ful79, Gil75a, Gil75b, Goo76a, Goo79a, Gor76, Hal79b, HL79a, HT79, Hau77, HW75e, Hoc79b, Hof79g, Kra78b, Kuk76, La 76, Lad77, Len76a, Lit77b, Mat77a, McC78a, Mül75b, iN77, NV79a, NV79b, Pap79a, Pap79b, Pas79e, Pas79g, Que78, Rob79a, Rob76e, Roo79, SW75a, Sta79c, Sta79b, Tay77a, Tep76, War79b, Yan77s, tD79a, Hof79d]. Ringschemata [Kra75e]. Rio [PdC77, Sch78g]. Ritt [Ram79a]. RKFHB4 [Opp78]. RNP [MRN78]. Robert [Kui79]. robin [BB75]. Robinson [Gre77a, Mac75a, SW75b, Sch77g, Vie77]. Robust [Rey78f, SM79b]. robustness [MBY79]. Roch [Ver76]. Rochlin [Hel79b]. Rockefeller [Nat77a]. rod [AB76]. Rodrigues [Gro78i]. Roggenkamp [RR79b]. role [Ben76b, Sim78d, Ste78g]. rolled [Gui77c]. Romania [CCJS79]. Romanian [CCJS79]. Rome [GM77n]. Ronde [Foa77a]. room [BW77b]. roots [Pea76]. Rosenberg [Gar76a]. Rosser [Maa75, Maa75, Sta75]. rotating [LD77]. Rotation [Her77b, Duc77, Sch75q, Tam78d]. rotation-invariant [Duc77]. rotational [Fra77a]. Roth [Hir78a]. Roughness [Cop78j]. round [BB75]. Route [SW78l]. routes [SW78f]. Row [Gen76]. Royal [Lit77a]. RUCA [BK75a]. Rudiments [Sch76r, Smi76g, War76j]. Ruelle [Roy76b]. rule [Jam78l, Jam78u]. rulers [BG78b]. rules [KM77, MR77a]. Runge [Fri78a, Bak78b, Bet78, Fri78a]. Runge-Kutta [Bak78b].

S [Kos75b, Kuk76]. S. [Bou79b, Kos75b]. sûre [Fey79a]. Saetze [RR75b]. Saint [And77a, BKK76, DCHRH78, FCG75, HJLN77, Lai78, BKK76, DCHRH78, FCG75, HJLN77]. Saint-Flour [And77a, BKK76, DCHRH78, FCG75, HJLN77, BKK76, DCHRH78, FCG75, HJLN77]. Saint-Jean-de-Monts [Lai78]. Saito [Gér79b]. Saks [CS79]. same [Con77, Kan79]. Sample [Sil76e, AM79a, CIS76]. sampling [Rey78d]. sans [Hec78, Jou79b, Nic77]. Santa [Mil78b]. Sard [DS76a, Sch76c]. Sasakian [Bla76e, Bla76g]. satisfaction [Kra76a]. Satisfied [Lan76g]. satisfying [Mat77a, Wil77c]. Sato [Bon78b]. saturated [BS75]. saturateness [Pac76]. saturèe [Dic77]. Satz [Voi77f, Voi77s]. Sätze [Haz77g, Voi77h]. saut [GMP78]. Sauts [Yoe79, Reb79]. SBIBDS [Seb79]. scalaire [Sam76]. Scalaires [Bre77n, Buc78]. scalar [Lon79a]. scale [dR79]. scales [KM78c, Mos78a]. Scattering [GS77, Kur75, Miu76a, Ree76d, Wil75a, AHK76d, Che76a, Chu76, Col76d, FM76a, Mal78c, New76a, Ros78h, Wil75g, Wil75f, Wil75h, Rau75c]. Schalentheorie [Vek76]. Schauder [FJ75c, Nov75, PP79]. Schema [Kra75a, Voi77c, Voi77d, Bou78d, Bou78g, Sca75, Bou78f]. schémas [Jou77a]. scheme [Bou78g, Coh79b, Lie78, Mil78a, Nør78d, Yan77c, Yan77i]. Schemes [Yan77t, BC77a, Che79, KL79b, Oor77, Pop77k, Pop77l, Yan77a, Yan77c, Yan77h, Yan77o, Yan77q]. Schensted [Gre77a, Vie77]. Schiffer [Sch75e, Tam78n]. Schiffer-condition [Tam78n]. Schlicht [Pfl76]. Schließungssatz [Jur78c]. Schmidt [Bon75, Mar79b, Wei78b]. Schnelle [DG76a]. Schnittelimination [Fel75]. Schoenberg [Sch75m]. School [BKK76, BK75a, CUW75, DCHRH78, DKT78, FCG75, GDMW76, HJLN77, PdC77]. School-Seminar [DKT78]. Schottky [Iva79c]. Schröder [Rog77a]. Schrödinger [BTW75g, Car78c, KSWW75, Sai79e, Sim78a, dM75a]. Schuartz [dlH79b]. Schubert [Las76a, Sta77b]. Schur [Jan75, Las77, McC77, Tho77]. Schütte [DM75c, Kre75c, Tol75c]. Schwartz [Won79c, Won79d]. Schweitzer [Sch78a]. science [Cot79]. Sciences [GLL79, Hil76]. Scotland [Eve75a, ES76]. se [And75]. search [Hag78, Wei76s]. Seattle [FW79, Hil76]. secant [Hol78b]. Second [AK78, Bec79a, Fen79, Jac79r, KRZ77i, Tol75d, Trè75, Bak78b, BCH78, Cla79g, Eva76, Gra76g, Gri76b, Hah76, HW75q, LEV75, Meh78b, Mey75a, MG78, OL77, Sch75c, Sto77f, TW78b]. Second-order [Trè75, Eva76, Hah76, LEV75, OL77]. Secondary [FI77, Sen76c]. Section [DM75b, HPS77a]. Sections [BWK77g, Lau79f, Sch77h, Cam78b, Far76b]. seeds [Tou76]. Segal [BDPR76, Pes76b]. segments [KP77]. Seidenberg [Yam79]. Seifert [LR78b, NR78a, Tro78]. seine [HJSW75]. séjour [Kôn76]. Sektoren [Jur78b, Jur78q]. Sektorielle [Jur78m]. Selberg [Zag77a, Hej76b, Hej76a, Hej76c, Sub75]. Selected [Che76j, Die75c, GR76, Sch75s]. selection [Adl79b]. selections [CV77b, Roc76]. Self [Far75g, Far75h, Raw79, Kal78, Kan79, Lan79b, MMS78, MS78e, Rod79]. Self-Adjoint [Far75h, Far75g, Kal78]. self-conjugate [Lan79b]. Self-dual [Raw79, Rod79]. self-maps [Kan79]. selfdecomposable [Jur78a]. sell [Joh78]. selon [Gui77a]. semantical [Vol75b]. semantics [Osi75b, vK75]. Semi [Ber76b, Deh76, Far75c, Ham75g, Hey78, Jac77a, KT75j, LP79b, Mey76g, Paq79a, Paq79b, Rot76c, AH79, Bad78, Bae78c, Bol78, Bor78c, Bré76a, Bré76c, CB75, Del79, Duc77, Duf75a, Duf79b, Fey78a, GP79, Gui77b, HSM77, Hug78, Jac78b, JY78, Jur78a, Kol78, KT76, Lel75a, Lin79a, Loe75, Nan75, Pel78b, Rot78, Sch76f, Sko78, SA77, Wan75, Yan77n, Yoe79, Yor76, Yor78a]. semi-biplanes [Hug78]. semi-continues [Bad78]. semi-derivations [Yan77n]. Semi-elliptic [Ham75g]. semi-free [AH79]. Semi-Groupes [Hey78, Ber76b, Deh76, Far75c, LP79b, Mey76g, Paq79a, Paq79b, Rot76c, Bré76c, Fey78a, Nan75, Rot78]. Semi-groups [Jac77a, Lin79a, Wan75]. semi-héréditaires [GP79]. semi-lattices [SA77]. semi-Markov [KT76]. semi-martingale [Jac78b]. semi-martingales [Bré76a, JY78, Pel78b, Yoe79, Yor76, Yor78a]. semi-normes [Sch76f]. semi-norms [Duc77]. semi-positifs [Sko78]. semi-simple [Bor78c, CB75, Gui77b]. semi-simples [Del79, Duf75a, Duf79b, Loe75]. Semi-simplicial [KT75j]. semi-spector [Kol78]. semi-stability [HSM77]. semi-topological [Bol78]. semi-vectorielles [Lel75a]. Semicellular [Bur75a]. Semicritical [BF79b]. semiflows [Raţ77]. Semifree [GL75]. semigroup [Bre79d, Fuj76, Haz77e, Kat76]. Semigroups [BJM78a, Duf75b, DR75g, Fei78k, MT76b, Pet76b, Pet76h, AL76, BJM78e, Dav78a, DR75b, EWL78, Get75j, Iit79, McF75, MT76a, MT76c, Muk79]. Semilattices [PR77a]. Semilinear [Med75, Ama79, GM77c]. Semilocal [Bae78e]. Semimartingales [Jac79f, Jac79j, Mém78, MS79b]. Semimartingals [Mét77b]. Séminaire [BR76d, ETFS75, ETFS79, Nor75, Nor78c, dlH79c, Ano76, Ano77, Ano79, BDV^{+}78, Bou75, Del77b, DMW77, DMW78, DMW79, HM76a, HM76b, HM78a, HM79c, Ill77b, Lel75b, Lel76, Lel77a, LS78, Mal77, Mal78a, Mal79c, Mey75d, Mey76h, Ste75c]. Seminar [AK78, BØS76, BR77b, CCJS79, Cha78a, Del77b, DKT78, GM79d, KM78b]. Seminars [KZ76]. Seminormal [Cla79d, Cla79e, Cla79a, Cla79f]. semipolar [Zab75]. Semiprime [Fis79]. semisimple [Bar79a, Hir79a, Jos77, Jos79, KZ77, Sch77e]. Semistandard [Jam78o]. Semitopological [DR75g, Gre79]. sens [CL76a]. séparabilité [Del77i, Sch76e]. separable [AHK76f, KR79, Ørn77, Tak78b]. separative [DR75b]. séparée [Dlo78a]. séparés [Sch76h]. September [Cop78i, GN76, GDMW76, HP79, Wer78]. Septembre [HH78b, And77a]. sequence [AEK78i, AEK78r, Bru78s, DS76b, EH79a, Mag76, Mor78a, TY78, Won79g, dB76]. Sequences [Jam78p, AS77, Aus75, AR75, Bab76, BC78a, Bre79e, BS76b, DJ77, EH79b, Gar76a, Get75d, Gie76, Gra76g, Gra76i, HK79a, HLS79, Kle75, Nie78, Rob76b, Rog79a, Tho77, Tho79b, Tru78, Wed78, Zel76a]. sequentially [Bal76, HK78a]. série [Clo79, RR78, Rhi75, Rub79, Sch79g, VR75]. Series [Gal76h, Gér75b, Lan76g, WK79, Gér75b, Bra79l, Bre79e, Buc77, Buh78e, Car77a, Cho77c, FJ75t, FJ79, Frö79a, Gal76b, Gal76j, Gam76, Gil75b, Gro78h, Hir79a, HH77b, HH77c, Jam78r, Kue78, Kwa76, Lan76c, Mah76b, Mah76e, MP79b, Nie77, RR78, Raz77, Roo79, Sch75b, Sri79f, Sri79g, Var77c, War76k, War76m, And77i, Bor76, CO79b, Vig77b]. Serre [Eis77b, Erm78, Her77c, Ill77b, Lam78a, Lam78d, Lam78f, Lam78g]. ses [Led76, Lel78b]. sesquilinéaires [Pes76a]. Session [FW79, Nas79b, Fis76h]. Set [CS76b, Fer79, Glo75, LSZ77, MS78d, Beb79, Che78b, Chu77a, Dev77f, Eas78, FT76, Fra77b, GM77e, Gra79c, Gua78, Juh77, KM78a, MM75b, Mos76b, Osi78, Osi75a, Pol79, Rud77a, Soc76, Soc77, vW78, HW75o, MSZ76]. set-theoretic [Juh77]. Sets [FJ75q, Won76d, AEK78i, Ama79, Ama76b, BK77g, Bec79b, Bih79, Bou78a, BF78, Bri76, Bru78n, Bur75a, Car79b, Che79, Cho76, CS78, CW75, Dei77j, DS76b, Dud79, Fro79b, Fuj79, Gal76b, GH79, Glo78, Got76, Han75b, Hel79b, Hil78, Hir75b, HPS77d, Jac79a, Lou75, Mac77a, Mag79b, MP78a, Mos78a, Naz75, PPW78, Pet75, Rhe75, Sch77a, Sri78, Ste78b, Ste78f, Wil77b, Won76b, Zab75, tD79c]. setting [Das75, IP76c, Mac75a]. seventeenth [McK75]. Several [SZ77a, Cie75a, FJ75d, FJ75a, Möl77-30, Rei77c, Sch77c, Suz79]. several-dimensional [Cie75a]. SGA [Ill77b, Del77h]. Shake [Lic79]. Shape [DS78i, DS78j, Has78b, DH77, DS78a, DS78d, DS78h, DS78l, FK79e, Mar76c, Seg78]. shape-invariant [FK79e]. shapes [Dob75]. shared [RY77]. Sheaf [FMS79, FH79, HK79b, KL79a, Cos79, Gra79b, Kaw75, Tur77]. sheaf-theoretic [Tur77]. shear [Saf77]. Sheaves [FMS79, FS79a, Joz79, vO75e, Ban79a, Ban79b, BM79b, DM79a, Fis76c, Hof77, HK79b, Pap79b, Ste76a, Tra76, Tra78a, Tra78b, vO75g]. sheets [Gui77c]. shell [AB76]. Shidlovski [Mah76a, Mah76h, Mah76g]. shift [Cla79b, DGS76l, HH78a, Har78b]. Shifts [DGS76p, AM78, DGS76b]. Shih [Get75n]. Shilov [KV79a]. Shintani [Gér79b]. ship [BP76]. shock [OW75, Ros78i]. Shoda [Voi77h]. Short [BK76f, Ben75a, FR79k, TY78]. shortest [MJ75]. side [HB79]. sided [Hem75]. Siegel [And77k, Mah76f, Mah76g]. sign [BDNY77]. signaling [Rin76]. signals [Gra78a, Gua78]. Signature [Kau78, ACH77h, Lan79b]. Signatures [Lit79]. significant [MG78]. Signorini [Cap79]. Silver [Ada75, DJ75, Ste78e]. Simon [Car78c]. Simons [Kos75b]. Simple [Goo79a, Gru75c, HW75r, Joh79m, Sch79d, Sta79c, AY79, Bor78c, CB75, Duk78, Gui77b, HP76a, Jan78, Joh79l, Kaz76, Lor79, Luc75, Meb75, Rob76e]. simplement [PS79]. simples [Cow79, Del79, Duf75a, Duf79b, Her78b, Loe75, Tit78]. Simplicial [Bec79b, Dup78j, Dup78i, Geo79a, KT75j]. simpliciaux [Han76b]. simplicity [HM79a]. simplification [Bru78r]. simply [Glu75, Moi77g]. simply-connected [Moi77g]. Simulation [Abr76b]. sine [Run76]. sine-Gordon [Run76]. Singer [Sha78c, Sha78d]. single [CL75g, FR76]. Singular [BN77, Bro75h, DH79, Far79, Goo76e, Joh79o, Joh79n, Kat76, Moe79, O’M78, Sre78, dG79, AHK76g, BG76a, BL75, Ber79a, Bro78a, CN77, Col78, Cur75a, GW76a, Joh79e, KSWW75, Kal78, KV79b, Lub79a, Lum77, Mil78a, Tat75, Tep76, vH79a, BGM77]. singular-splitting [Tep76]. singulären [Jan79a]. singulärer [Bla76i]. Singularitäten [Gro76]. singularité [Mal75a]. Singularités [BR76d, Jou79a, Krá76b, Poé76b, Gér78, HL78, Meb75, Ram79b, dM75a, GS79b, BR76d]. Singularities [Oor77, Poé76b, Ree78a, BM77a, Cam78a, Cor79c, Dek79, Gér77, Gut75, Her79b, Mil76, SW78a, Tra76, Tra78a, Van79, YF79, vdE79]. Singularity [Cha78b, Per79a]. singularly [Jor79, Noh79b, vH79b]. singulier [Dab79, Dar77]. singulière [Dem77b]. singulières [GM77o, Hue77, KvQ75, LR78a, Ver76]. singuliers [Mal79a, Mat78b, Ste78e]. Situation [Voi77s, Voi77h, Voi77f, BK77d, BK77c, BK77h, BK77i]. sixteen [Rob76c]. Sixth [HW79b]. size [Mek77]. sketch [Gro78l]. skew [BW77b, EW76, Jam78j, MN79]. skew-hooks [Jam78j]. Skoda [Mal78b]. Skolem [Win75]. Skolemerweiterungen [Oss75]. Skorokhod [AY79, Ros76b]. slender [Gui77c]. slice [Lic79]. slow [Dam77]. slowly [Gat77]. Smale [Fra78b, GdM77, New76b]. Small [Gra79a, HH77e, Mic77, BK77j, Čud77, Dev75, Gra76a, Hir77, Hoc79a, Kov78, Mah78, Mas79]. Smirnov [Dur76a]. Smith [Jon78b]. Smooth [ACH77g, IP76d, Mad78, Sch78c, WdPL76, Die75f, Hol79, IP76c, Mor77c, SS77c, dP76b]. smoother [Mal79d]. smoothers [Mar79d, SM79b]. smoothing [Las79, Utr79, Wah79, GR79b]. Smoothness [Gri76b, Sun75, BK76c, HPS77e, Raţ79]. Sobolev [Boh75, Duc77, Eva75, Fd76, Gro78b, Lac78, Mat78b]. Society [Cro75b, FW79, HS79b, Nas79b, NR78b]. socle [RR78]. software [Cod76]. Soglio [Fel77c]. solid [HBT75, BWK77o]. Solides [BWK77o]. solids [OW75]. Solitary [BBB76, Lon76]. Solitary-wave [BBB76]. Solitons [Miu76a, Wah76]. soluble [Bie79, Tay77a, Wam77, Wil77c]. Solution [Axe77, Beb79, Eis77b, Geo77, Man76d, Rei77b, AL79, Axe79, AY79, BM79a, BP79a, Ben78, BK76g, Blu79, Bru76e, Don75, Ede76, EW78, FJ75p, Fre78, Gaj79, GR79h, Gri76b, IK79, Ioo76, Jou79b, KY79b, Koß78, Kra79, Lam78h, Mos78b, Ort78, Per78a, Ral79a, Rhe75, Sim78b, Wir78, You75]. Solutions [BK76f, Cox79, Dei77j, Gér78, Jac79n, Jac79p, Kré75b, Ler75, Ros78i, Wil75d, Wil75e, Alt79a, Ang79, AS79, BZ78, Bib79a, Bib79d, BK76b, BBB76, Bon78c, BN77, Cap79, Chi76c, CS76c, Cus79, Dei77a, DT77, DN77, Eas76, FT76, Fro75, GM77b, GM77k, Gam79, Gér77, Gre75b, Gro78f, Hec76, Her75d, Hes78, Hes79, Hir76a, Jer76a, Jer76b, Jos76, Kor76, Krá76b, KS79b, Lev79a, Lio76, Mah76e, May78b, Mil76, Nus79, Rau78, Ree76b, Sko76, Tak76a, Tem76b, Wil75i, Yam76, Yam78a, dM75a]. solvability [Cha79a, Pet79]. solvable [Cha79a]. solved [Rat78m]. solvers [STW78, Ste78j, Ste78i]. solving [Alt79b, Eav79, Fri78a, GT77, Gen76, GR79g, GP77, Hac78a, LY79, Loo75f, Wam79]. Solvmanifolds [Bre77d, Bre77b]. Some [AEK78p, Ale78, Ass79, AH76, Azu79, Bea76, Bel78, Blo79g, Bou76a, Bre79a, Bre77e, Byc78, CW77c, Can75b, Caz79b, CM77, Chr78, Com77, Csá77, Dav78g, Dav78h, Dej78b, Duk78, DS78k, Ead77, Ear79, EH76i, Erd78, Est76, FR79i, FR79j, Fin76, FM76a, For76a, Frö79a, Gal76i, Geo79i, Geo79g, GW76b, Get75o, Gia78, Gin76, GM76c, Gor76, Gor78a, Gra76j, Gre77a, Gro79h, Háj78, Hel79b, Hen79a, Her76, Hof79m, HS75c, HK78i, Ish76h, Jac79a, Jac78a, Jor79, Kai76, Kar77, Keo77, KLM75, Kra78b, KMS75, Kue78, Kuo75d, Lan76h, Lau79g, Lau76, Leh76, Lev78, Lum77, Mah79, Mal79d, Mar76a, Mar79c, McD78, Muk79, Oli76, Pie78, Pre76j, Rah75, Rat78k, Rat78l, Rat78n, Rat78m, Raw78, Ree78b, Sch75b, Sch75p, Seb79, Sen76c, Shi78, Sim78e, Sin77, Smi77]. Some [Sta77b, Ste77b, Ste78h, Sto75b, SV75e, SM79b, Tak76b, Tru76, Van78c, Wam79, Wei76r, Wei76s, Wei76t, Wei76u, Won76b, dG75h, dG75i, vW76, AL77, AB75, Alt79a, Ama76a, Ave78, Bar79a, Bar79h, Beb79, Beh79k, Bin75g, Bin75h, BBB76, Bot78, Bre75c, Che78b, Chi76b, Col78, Cus79, Czy78, Dav78i, DO77, Düm79, Dup78b, Ede75, Fek79, GM77c, Gaj79, Ghi79b, Gro77a, Hal79i, Hau77, Hec76, Hes79, Hla79, HLS79, Ist78, Joh79e, JS77, Kra78a, Laz75b, Lio75b, LS79b, Mol77c, Mos77a, Rob76a, Rob79a, RB75, Sbo79, Sch79e, Sch75g, Sch75n, Shn78, Sne75]. somewhat [Ahl77]. somites [Zee76d]. Sommation [Bre77m]. Sommerfeld [ZKB79]. Sommes [BWK77f, Bru76c, Del77a]. son [ET79]. Sonderforschungsbereich [SZ77b, SZ77c]. sont [Fer77, Pou79]. SOR [EW78]. SOR-method [EW78]. Sorites [Ser75l, Ser75m]. Sound [CN75b, Nix78]. Sources [BK75c]. Sous [BWK77b, BWK77o, Cou78, DC75b, Jac79q, Maz77a, Ste75d, Ben75b, Blo77, CO79a, Cha79b, GS79a, Jen78, dSLM75b, Nac78, Tit77, Zan77]. sous-algèbres [Nac78]. Sous-Directs [BWK77b]. Sous-ensembles [Maz77a]. Sous-Espaces [Jac79q, DC75b, Ste75d]. Sous-groupes [BWK77o, CO79a, GS79a, Tit77]. sous-markovien [Zan77]. Sous-modules [Cou78]. sous-suites [Cha79b]. sous-variété [Jen78]. sous-variétés [Blo77]. sousharmoniques [Kis77b]. Souslin [Fel76]. sousmartingales [AM79b]. Southern [Nat79]. Space [AD78b, AD78a, BH78, HJ77, Acc79, AHK76f, Ama76b, AV79, Ban79b, Bar79b, BW76, Ben78, Bin75i, Bla76g, Bor78b, Bor79a, Bre79c, Car79b, Car79d, Cha75a, DJ77, Dav78h, DGS76l, Die75e, DR75c, EGS75, Eps77, Fer78b, FJ79, FJK79, GK76a, Get75k, Gie76, Gla76f, Goo79b, Hah79, HH77e, HH77f, Ike78, Iva79c, Jai76, Jen77e, Kno75, KT76, Kue76b, Kue76a, Kue79, Kuo75d, Kuo76, LM79, Lak79a, Mad78, Man78a, May77j, Ørn77, Peł76, Rob77a, Ros76a, Sat76a, Sat79a, Sch78e, Sco79, Sil76e, Ste79, Sto75a, Stø78b, TW78b, VGK79, dF75]. space-valued [BW76, Gie76]. spaced [Kös76b]. Spaces [AEK78r, AEK78q, AEK78u, BCD77, Bec76b, Bec79a, BDE^{+}77b, BTW75b, CLM76, Coh76a, Dei77h, DGS76d, Die75c, Dub79g, EL77f, FP78f, FJ75l, HK78b, Jen77c, Kel75c, Kuo75c, May77a, Mic78g, Pet76b, Sto77g, Tay78j, Uen75f, Wer78, Won79d, Ada78, AEK78f, AEK78c, AEK78d, AEK78g, AEK78e, AEK78i, AEK78j, AEK78l, AEK78a, AEK78m, AEK78n, AEK78o, Akc79, Aki78c, Ama76a, Aro76, Bae78c, Bec76a, BDE^{+}77d, BDE^{+}77c, Beh79k, Ben76b, Bin75e, Bin75c, Bin75g, Bin75h, Blu79, Bol75, BHP78b, Bou76a, BF78, Bra79t, BTW75a, Bro75e, Bro75f, BS76b, BM79b, Byc78, Car79a, CP77b, CV77e, Cha78b, CG79b, Chi76a, CS76b, CBR76, Cig76, ČT78a, CT78b, CMN79, CS79, DE77, Dav79, Day75, DGS76k, DK79, Die75f, Die75g, Dub79f, Duc77, Dup78k, Dwy78b, Ede75, EH76g, Egg78, Egu79, Elw78, Fat75, FP78h, FP78j]. spaces [Fis76a, Fla76, Frö76, Fro77, Fuc77a, Gar76b, Geo78b, Get75p, Gin79, Gon78, Goo76c, Goo76e, Gór78b, GvW76, Gua78, Ham75b, Har78a, Hei79, Hel79a, Hel78b, Hil77a, HMRS78, Hof79l, Hof77, HK79b, Hub78a, Hub78b, HK78a, HK78c, HK78d, HK78e, HK78g, HK78i, HK78j, Huš79, Ist78, JW77, Kan78a, KR79, Kel75a, KN79, Kle75, KS79a, Kor79, Kra78a, Kuč75, Kuo75b, Kuo78, Lac77, Lap79, Law75a, LeP78a, Liu79a, Liu79b, Mac77b, MW78, Mas76, Mat76, May76a, May76b, May77k, May77b, May78a, McC79, Nas79a, Nor78b, Oma75, PT76, Pal78a, Pet76f, Pet76j, Pom76, Pop77a, Pop77c, Pop77d, Pop77e, Pop77f, Pop77h, Por79, Raj77, Roo79, Rud77b, Sar77, Sch77b, Sha77c, Sil76a, SS77c, Smy78a, Sna78, Ste77a, Sto77c, Sun75, Tak78d, Tay78a, Tay78d, Tay78h, TW79, Tho76b, Tho78]. spaces [Too78, Top76, TDHJ76, Uen75c, Val75, Wer75, Won76a, Won76c, Won79b, Won79c, Won79f, Won79g, Woy75, Zie79e, dA79, vR79, BØS76, Sch76i, AEK78i]. Spalten [Jur78l]. Spalten-Konvergenz [Jur78l]. span [Law75a]. Sparse [Rei76b, Gen76, Rhe75, Bar77b]. spatial [CN76, Hyl79, Rin76]. Spatially [Geo79h]. Specht [Jam78s, Jam78q, Jam78r]. Special [Cha78a, DR75h, FW79, Fei78k, Ger79a, Huš79, Jen77f, LT76d, Nas79b, Pet76i, Sch75q, Ben75b, Cod76, GM77m, Hig75, Joh79e, Laz75g, Mah76a, dG75i]. species [Tai75]. Specific [Pre76k]. specification [Lin79b]. specifications [Pre76b, Pre76f, Pre76i]. spector [Kec78c, Kol78]. Spectra [May77a, vO75c, Boa78, Boa79, BF78, Cos79, GM76b, GM76c, Kom78, Mah78, May77f, May77c, PS77, dG79]. Spectral [BDE^{+}77i, Cod75, EL77f, Eve76, Sai79d, Sai79e, Wei75a, Ben76c, Bru78s, DNR79, Eas75, EL77a, EL77c, Eve75b, Ist78, KSWW75, KO79, Man76a, Mor78a, Ple75, Sch78e, Tho79b, Wal75e, Eve75a]. Spectrale [Rod77, dSLM75g, Tua78]. Spectrales [Aup79f, LS76, Sjö76a]. Spectre [BWK77g, Car78c, Cha76e, LBL78, Aup79d, Aup79e, MvM77, Ren79b]. Spectrum [Gel76g, NT79h, Pap79b, Ama76b, AdS75b, AdS77, Bra79s, Dui75, GHM77, Jos77, Jos79, KV79b, Ral79b, RB75, Sib76a, Wal75g]. spezielle [Mei78, Jur78n]. spezieller [KPP^{+}78g]. sphere [Bla76h, Dun78, Hau76, Lat79]. spheres [AH79, BLS75, BLR75c, Dav75b, Dav78b, Dav78c, GL75, Kos79, Lar79, Mar79c, Max78, Sch79i, dlR75, tDP79, Boh75, Gra76b, Kui79]. Spherical [FJ75t, Tak77a, Dre77, FJK79, Heg79, Koo77, Lau75a, Mad78]. sphériques [Bam79, Far75b, Mat77d]. Spiegelung [Sch78d]. Spiele [Har75a]. spin [Gaw77, Pet78c]. spines [FR79a]. spinor [Mic78a]. Spitzer [Sza75]. Spline [JLR76, Sch76c, CSW76, Cox78, FJ75c, Lyc76, Mei76, SS77a, Sch76s, Utr79, dBS76, dB76, tM76, BMS76, Mic76c]. Spline-Funktionen [JLR76, Mic76c]. Spline-Interpolation [Sch76c]. Splineapproximation [SS76b]. splinefunktionen [Str76, Mei76, Som76]. Splines [Duc77, JS76, DSS76, DS76b, FJ75o, Kös76b, Wah79, dBS76, Häm76]. split [Aus75, AR75, Ear79]. Splitting [La 76, Tep76]. Splittings [Boa78, AH79, dP79]. Sponsored [FW79, Gol75, HS79b]. spread [Ber76f]. spreading [BS76b]. Square [SW78g, Eas76, Kaz76, Pea76, SW78i, SM79a, Sło78]. square-integrable [Eas76]. squares [BW77b, Gen76, GM76a, GV79a, Hei77, LP79a, Spe75b]. St [Nas79b]. stabilen [Voi77o]. Stabilität [Bra76b]. Stabilitätsbereichen [Sch78d]. Stabilitätsverhalten [Man78c]. Stabilité [Poé76c, Poé76d, Ron75, Car78c]. Stability [Alp76, Bib79e, Bjö77, Böh76a, Bri79, Cop78f, Cop78k, ER77, Gra75, Gue78, Hil76, Joh77, Kas75, McA75, Sat79g, WdPL76, AW78, Bra76c, CN77, Cam78a, Fri76a, Gut75, HPS77i, Hol78a, HSM77, Jel78, Jos76, Loo76, Mañ75a, McA76, Pal78b, Rau78, Rin76, Rob75b, Wed78, Wei76h, Wei76j, vdK76]. stabilization [Wei76g, Wei76t]. Stable [AM79a, FJ75r, Las79, Sil76f, Sna78, Urb75b, War79a, Hir75b, Jur78a, Kah78, Kna79b, KS79a, MW78, MHG77, MP78b, Pap79a, Pop77d, SS76e, Was76, tD79h]. Stables [Jac79q, Bal79, BE78, Mar76d, Mok78a, Sig75]. stacks [Pop77j]. Stampacchia [Pra78]. Standard [Eif76, Mol77v, Bel76, Fei78j, GH75, Jam78s, Kra76a, RS78, Sch78a]. standardization [Cur75b]. Stanton [Oxl79]. Starlike [Suf76, Sch75q]. Starlikeness [Suf77]. stars [Hog75]. State [AHW77, BH78, BCD77, CS76d, Kel75c, MMP78, AHK76b, AHK76a, Goo79b, Kno75, KT76, Stø78b, Wil75g, Wil75f]. Statement [Sha78f, Con76a]. States [Bow75b, AHK76i, Ara78, Geo79e, Kas78, Mal78c, Pre76d, Sta78d, Wal75c]. stationaires [Ben75b]. Stationary [Bru78n, FT76, GR79i, Gra78a, Gra76g, Gra76i, MS78b, Nie78, Sch78e, Ste78b, TW78b, Vas78, Wei76e, Wei76m]. stationnaire [Tua78]. stationnaires [Jol78, Ruc78b, Web77]. Statistical [Moo78, Rey78f, Foi75, Geo76, Gri76a, Kas75, Par79c, Sta78b]. Statistics [Kom76, EN76, GK78b, GK76a]. Statistique [DCV78, Reb78]. Statistiques [DC78, BDR78, Ruc78b, Sig75]. status [MS79a]. Steady [Wil75g, Wil75f, Ioo76]. Steady-state [Wil75g, Wil75f]. Steenrod [EH76c, EH76j]. Stefan [Jer76a]. steifer [Fri78a]. stein [Pou79, Suz78b, Dem78, Erm76]. Steiner [Bjö77]. Steinhaus [AEK78b]. stems [BR78b]. step [Mey76b]. stepping [DDP78]. stetige [Mét77c, Haz77e]. Stetigen [Bro76a, Bin76]. Stetigkeitsmoduli [Ess76]. Steuerprozesse [MH75]. Steuerungen [Dic75]. Stieltjes [Ber79b, Gil78e]. stiff [Dah76, Fri78a, Ste76d]. Stimulated [Chu76]. Stochastic [Åst78, Cha78c, Cie75b, Gra76e, GH75, Her75d, Itô75, KK78a, Mét78a, Tay78j, Dav78a, Dun78, Fei78d, Gra76g, Gra76i, Gri76a, HC78, Kom76, Lig77a, MS78b, Ood78, Pel78b, TW78b, Wed78, Yam76, Fei78c]. stochastically [Gal76e]. Stochastique [Jac79b, Bis78b, Mey79]. Stochastiques [Cau76a, DCV78, Jac79h, Jac79k, Jac79j, CM75, Gan75, Len78, Yam78a, Yor79]. Stochastische [Mét77b, Mét77a]. Stochastischen [Mét77f, Mét77c]. Stokes [Tem76c, Foi75, FT76, Fro75, GR79e, GR79g, GR79f, GR79h, GR79i, GR79j, Ioo76, Mar77, Ole78, Raţ77]. Stokes’sche [Jur78o]. Stokes’sches [Jur78b]. Stone [Beh79i, Ben76c, Beh79a, Beh79b, Beh79f, Bra79p]. stopped [Wil76a]. Stopping [DG77, Get75a, Ros76b]. storage [Cir78]. Störungstheorie [Jeg75]. strain [AD79]. strained [Hab76]. strange [HP76a]. Strasbourg [DMW78, DMW79, ETFS75, ETFS79, Foa77a, DMW77, Mey75d, Mey76h]. strata [Dav78f]. strategy [Alt79b, Deu75, McC78b]. Stratifiable [Tou76]. stratifiables [Tou76]. stratification [Hol78b]. Stratifications [Wir76, Gib76]. Stratified [McC78c]. stratify [Mat76]. stratum [Han75b]. strenge [KPP^{+}78e]. stress [HBT75, YS79]. strict [AEK78t]. Stricte [Mic76a]. strictement [Sib76b]. Strictly [Her78a, Cho76]. Strings [AEK78s, Bar78f, MJ75]. strip [Raz77]. Strong [BC77b, Gie76, Lad76b, Lei75, BW76, Beh79a, Hud78, Kue76a, Rat78o]. Strongly [Gla76j, JR77, Web78a, KSWW75, LD77, Wei76b]. Structural [Cam78a, Gut75, Rob75b, Gal75, Ger79a, Pal78b, Hil76]. Structure [Bae78f, BDE^{+}77b, Beh79i, Beh79k, Dub79g, Eps79d, Goo76d, Goo76f, MMP78, McF75, Nam79g, Adl79a, Bar79f, BDE^{+}77c, Beh79j, BJM78e, Brü76b, BLR75g, CM79, Dav78i, Duv78, EH76g, FS78b, Fra78b, GM77m, GS79a, Gar77a, GPR78, GP78a, HP76a, HPS77i, Hol78b, Jos79, Jur78g, Kij77, Lac77, Laz75h, Lel77b, Maa78, Obe77, Pet78c, Pop77d, Rog77b, Ros78f, Sib76a, Sri78, Sto75b, SV75b, Szc77b, TW79, Wal78b, War76a, War77b, Wei76m, Wil77a, tDP79]. structured [Fri79]. structurelle [Poé76c, Poé76d]. Structures [Dic77, Lau79a, McG78, Mol77u, Mol77v, Mol77-28, SW78b, Yan77r, Yan77s, Aki78g, Bie77, Bla76e, Bla76f, Bro75a, Che76h, Fen75, FSW75, Fis77, Gre79, HW75g, Huš76, Leh76, Mol77t, Mol77s, Mol77p, Oli76, Pom75, Ril79, Sch75a, Shi75, TH75, Pes76b]. Struktur [Haz77f, Jur78p]. Strukturen [Fel76]. Struktursatz [Kra75b]. student [Why78]. studies [BDNY77]. Study [Džu76, Bré75f, Cur75b, Foi75, Kan78b, PP79, Sto75b, Tak76a]. studying [Kur78]. stunted [Dav79]. sturdy [Liu79a]. Sturm [Eas76, Kal78]. Sub [GW79k, Gui77b]. sub-algebras [Gui77b]. Sub-mean-value [GW79k]. Subadditive [Kin76, PR77b]. subalgebras [Yan77b, Yan77d, Yan77f, Yan77p]. subcell [Wei76m]. subclasses [Fek79, FF78]. Subfree [Fuc77a]. subgraphs [Bab78, Chi78a, Qui79]. Subgroup [Imr77, Fef75a, Sim78c]. Subgroups [BWK77o, Con77, Gla76b, How78, JR77, LS77, Pas79b, Pas79c, Pas79h, Wil77c]. subharmonic [GW79k, Hay76]. subidealizer [HT79]. Subidealizers [Tep79]. subject [Bur76]. subloci [Acc79]. Submanifolds [Jen77c, Joh79b, Joh79k, Joh79o, Joh79n]. submartingales [Szu78]. Submeasures [Chr76]. Submersions [Ele78, Aki78e]. subnormal [Lub78]. Subordination [Dur77, Man78b, Haz79b]. subschemes [Yan77d, Yan77f, Yan77i, Yan77l, Yan77p]. Subsemigroups [DR75i]. subsets [Bro75b, Die75e, Mus76, Sol78a]. Subshifts [DGS76q, DGS76p]. subsonic [Bré75f]. Subspaces [BJM78f, Dub79h, Zie79g, AT75f, BJM78c, Bro75c, Cod79, Dub79b, EL77e, Ørn77]. subsquares [Hei77]. substitution [Bro78d, Mic76a]. successive [Sch77f]. successor [Mey75a]. Sud [Tem76c]. Sufficient [Mil78a, Sch75r]. Suffisance [HT75a]. Suisse [dlH79c]. suite [Mok76]. Suites [Bre77h, Bre77n, Gil78e, Gil78i, Gil78j, BG78d, CG79a, Cha79b, GB79, Nie75a]. sujet [Bur76]. Sullivan [Mor77c]. Sum [Mac77a, Arn77, Rob76b, Wol75]. Sum-free [Mac77a]. summability [Luk79]. summable [Tho76b]. summand [Tak78b]. summary [Dav75b, Gel76b]. summation [Dre77, GK76b]. Summer [BKK76, BK75a, Cro75b, DCHRH78, DM75c, FCG75, GDMW76, HJLN77, Løn79b, MOP75, NR78b]. summing [AEK78g, Won79g]. Sums [Moi77b, Tay78j, AEK78t, DS76c, GMZ79, MS78e, PT76, Phi79, Rei77c, Sch76m, Sto75a, Tay78a, Tay78e, Tay78i]. sup [Paq79b]. Super [Bea78h]. Super-Propriétés [Bea78h]. Superalgebras [Sch79f, Sch79d]. superconductivity [Pet78c]. supercuspidales [Gér75f]. Supérieur [Mol77u]. supermartingales [CL75e, Get75d]. Supersingular [LT76e]. supersonic [HH77h]. supersymmetry [Ste77b]. supplement [CV77h]. Support [Die75e, Ros78j, Iag75a, Pra79, Sat76a, Sat79a]. supports [Kru79, Lub79a]. sure [BD79, PT76, Phi76, Phi79]. surface [Des79a, Jen77f, Kaw79b, Lio75a, Mor79c, NV79e, Pie78]. Surfaces [Acc75d, Cat79, Jen77g, Moi77b, Nam79e, Böh76a, Bou78a, Eck75, Fre75, Ghi79b, Glu75, Hej76e, Hem75, Ike79, Joh79e, Joh79i, Joh79o, Kaw79a, Kil78, Moi77a, Moi77c, Moi77d, Moi77g, Moi78, Nam79d, Per78c, Roc77, Van78b, Van78c]. Surgery [And77f, And77g, TW79, And77d, And77e, Cla79h, Dov79, HM79b, Pet78b, Ran79b]. surhamoniques [Han76b]. surharmoniques [Lum76b]. surjectifs [Sko78]. surjective [Ber76f]. surmartingale [Bru79b]. surmartingales [Cho75a, Str79c]. surmédianes [Mok79]. Survey [Sri79h, Ati77, Coc78, Cor79h, Far76e, Gal76j, Gro78b, HL79a, Har78c, Hin77, KW78, Mor77d, Mul78b, Pet78b, Ple75, Sch79e, Slo76, Won76c]. Suslin [Eis77b, Fer77]. suspended [Pal78b]. suspension [Dwy78a, Lat79]. Suspensions [Par75b]. Sussex [Fen79]. Swinnerton [Lan78]. Swinnerton-Dyer [Lan78]. Swirling [McL75]. Switzerland [Hau78a]. symbol [Cow78]. symbole [HL78]. symbolic [Per78a]. symbols [Len76b]. Symétrie [Poé76b, Vai78]. Symétrique [Foa77a, Lac79, Tak79]. symétriques [Aup79c, DC75b, Kos76c, Loh75, Mey76g]. Symmetric [Iit79, Jam78n, Sil76b, vO75f, vO75g, AF78, Car79c, DE77, Egu79, EG75, FJ79, FJK79, Fuk76, Gei77, Gla76f, HWH78, Jam78t, KRZ77c, Kor79, Par76, Pea76, Urb75b, vO75a]. Symmetrization [Ker75b]. Symmetry [Poé76b, Gar77a, Olv79]. symplectic [Ded77, Gar77a, GPR78, KS79a, Lan79b, Sha76b, Szc77b, Tul77a, Tul77b]. symplectique [Ler75]. symplectiques [Ban75]. Symposion [DM75c]. Symposium [BR77c, BMS76, DM79b, Eve75a, FMS79, GN76, GLL79, LS79a, Man75d, MP76, MWW76, Mil78b, Nov77a, Ols78b, Par75a, Sch78g]. syntactic [Joh79r]. syntaxiques [Sak77]. Synthèses [Boh75]. system [Ara78, AH76, CMPY79, Cir78, Got76, BG76b]. Systeme [Buc75, Dab79, MM75a]. Systèmes [Con75b, Jou77c, Jou79g, Mat77c, Bis78b, Con76b, Gér78, GS79b, Kla79, Leg75, Mal79b, Mat75, Mat77h, Sig75, GR79c, Mat77f, Mat77g]. Systems [DKT78, EN79, Geo79i, Gri79a, Man75d, MMP78, Pet76b, Tol75e, Zie79c, Ama79, AG76, AW78, Axe77, BV77, Ban79c, Beh79d, Cie75b, CS78, Cos75, Dav78a, Dei77d, DGS76m, DGS76j, Dun78, Fat75, Fle76, Fri78a, GT77, Gen76, GM77p, Geo77, Geo76, Geo79e, Gér77, Gra79a, Gri79b, Gri79c, Grm79, Grü75b, Hac78b, Has78b, HM78b, KS78b, Klo78, Lig77a, Lor79, Mal79b, McK79, Mor76, Mul78b, Nam79b, Nam79d, Neu79, New76b, OP79, Poh75, Rei77b, Rog77b, Sar78, Shu75, Sma77, Sta75, Ste76d, Sza75, Tol75c, Wan79, Wed78, Wir78, You75, dP79]. Szekeres [Con75a]. Szemerédi [Tho79c].

T [Bou79b]. T. [Bon78b, Bou79b, Gér79b, Ray78b]. tôr [Ste76b]. Table [BK75d, BK75e, BK75f, BK75g, BK75h, BK75i, Foa77a, Ano79, Bul79, GM79b, Jam78c]. Tableau [Tol75e]. tableaux [Foa77c, Jam78f]. Tables [Gil78d, BK75c, Cor79a, Gia78, Tan78]. tabloids [Jam78f]. Tail [Bor78b]. Taketa [Voi77f]. talk [Sza78]. tame [Dav75a, Dwy78a]. Taming [Can75a]. tangent [Bla76h, Goo76c]. tangents [HT75b]. taper [Smy78a]. Taskent [MP76]. Tate [Del75]. tau [Ort78]. Tauberian [Vaa75]. taut [Rat78p]. taut-level [Rat78p]. Taylor [Wil76a]. Taylorlemma [Voi77e]. Tchebycheff [CL76a]. teaching [Sim78d]. Technical [Bar77b]. technique [FR79k, Gau79, Koß78]. Techniques [Bar77b, Blo79f, Far76l, GR79b, Mol77w, Mol77x, Mol77y, Ale75, BP76, Cap77, Cun75, Est76, Far76k, GT77, Utr79, Wuy79a]. technological [Loo75f]. Technology [Lit77a]. Teichmüller [Acc79, Ław79]. Teilräume [Gro79g]. temperate [Kuč75]. temperature [BH76a]. tempérées [Bon75]. Temps [Jac79e, EM77, EW77, Kôn76, Pra79, Str78]. Tensor [Car78b, Cig76, Hof79j, Mic78g, Mic78k, War76k, Won79d, Won79e, Won79f, Bec78a, Ber76d, CdC77, Gue78, Jak79, Ker75b, Lew77, Mic78j]. tensoriel [LP79b]. tensoriels [Fd79]. Tensorprodukte [Jan79i]. tensors [SV75f]. tenue [Foa77a]. term [Mor78a]. Termen [Maa75]. termes [Ill77a]. terminal [MJ75]. terminals [MJ75]. terms [Maa75, MG78, SW75c]. tesselations [Gal78]. Test [Ish76i, Jol78]. testing [BP76]. tests [Dur76a]. Texas [Ode75, Why78]. th [DP77, MZ78]. thêta [Vig77b]. Their [DKT78, FP78f, JPR^{+}78, Miu76a, Pas79d, AEK78q, AG75, Bir79, Cat79, Cha76c, CT76, GM76b, Gro78f, Kea78, Mor77d, Mos77b, MRN78, Rob75a, Run76, Sen78, Sne75, SV75b, Tak76a]. them [Kan78b]. theme [Ram78b, Ram78a, Ram78c]. Theorem [Cam75, Ess75i, Ess75c, Geo79i, Maa75, Sha78d, AEK78c, AEK78d, AEK78b, AEK78h, ACH77h, Alo79, Azu79, Beh79b, Beh79f, Ben77a, Ben75a, Ber76g, Bir79, Bra79m, Bro75g, Bru76a, Buc77, CV77b, Chi76a, Chu77b, Cim77, Col76e, Con76a, CL75b, Dav78i, DJ75, Die75d, Dok78, Ess75b, Ess75e, Ess75j, Eva77, FJ75g, FJ75i, Fuj76, GM77h, Get75n, Gin76, GW79e, GW79k, GK76b, Gug78, Haa79, Hac78b, Hei79, Hel78b, Het79, HPS77a, HH77c, HH77d, Hod75, Ist78, Jai76, Jam78b, Joh78, Jur75, KS76, Kwa76, Ład75, Lam78c, LR78b, Lan76e, Laz75h, Li79, Lie79, Lit76, Loo76, Mac75a, Man78a, MW78, Mas75a, May79, Mic76b, Moi77f, Mol77z, Mol77-27, Mos79, Pet78a, Pet78d, Roq75, Ryc78, Sch78b, SS76c, Sha78a, Sha78b, Sha78c, Sha78f, Sta77c, Ste77a, Sto77b, Sto77c]. theorem [Sto77f, Sto75b, Tog76, TY78, Tur77, Vaa75, Vol75a, War76g, Wei78b, dA79, dBS76, Beh79i]. Théorème [DM75b, Mat78b, Mol77z, Mol77-27, Alo79, BG78a, Ben77b, Ber76g, Bre75a, Cho78b, CR79, Del77j, FS78b, Fou79, GS79a, Gui76a, HL78, Hei78, JM76, Las76b, dSLM75f, dSLM75e, Lel78b, Mey78, Pra78, Ram79a, RS78, Sch76a, Ste75c, Tho79c, Ver76, Rum79]. Théorèmes [CL76b, Tei75, Yor78a, Cle75, Yam79, Del77g]. theorems [Akc79, And79a, AH76, BZ78, Bom77, Bro75g, CV77b, Che76k, Die75a, DS78l, EN76, Egu79, Ell75, Ess75g, Fei78f, FP78c, FP78d, FJ75p, For76a, GM77h, Gar78, Gar76b, Gra76h, Her76, HH77b, Hoc79c, Iag75a, Imr77, Iva79a, JS76, Joh75b, Jos76, Kap78, KZ77, Kre75c, Kri76, Kue76a, Mah76a, Mah76g, MR77g, Mar76a, McC77, McF75, Meh78b, Mét78a, PS78, Pop77b, Rév76, Sch77a, Seb79, Sje78, Sta75, TH75, Uen75i, Van78a, dG75h]. Theoretic [Che76j, MWW76, Sat79h, AL76, Cur75c, DGS76j, GH76, Gre77a, Hal79f, Juh77, KLM75, Kur78, Möl77-30, Rud77a, Sim78e, Tur77]. Theoretical [MR77h, Pha75b, Ent78, Fer79, Hof76b, HK79b, Mal79d, Osi75a]. Theoretische [SZ77b, SZ77c]. Theorie [Mét77c, Mét77f, Voi77n, Cou75, Gat77, Gil78b, Gil78a, HM76a, Jac79o, Mey75c, Mol77-29, Mol77-28, Poé76e, Ser75n, Col76c, Bad78, Ben77b, Bis78a, Buc78, Cho75b, Con79f, Con75b, Del77j, Duv78, EM77, Eym75, FM79, Fed78, Gra77, Hey78, Jou77a, Kar79, Kré78b, Le 79, Lép78b, MN75a, Mey76c, Mey77c, Nie75b, Pro78, Reb78, Sig76, Spe75a, Tho79a, YM76, FK79d, CK76, EW77, HM76b, HM78a, HM79c, Kré77, Roy76b, Ser75o, Bre75a]. Theories [Cho77a, EH76c, FR79b, MR77l, Tol75e, And77d, AB76, Che76e, Fef75c, Fen75, FR79c, FR79g, Fit75, Hil77b, JW78, Kea75, Lac76, May77h, May78b, Mol77b, Mol77e, Mol77l, Obe77, Sna78, Ste78g, Yor77, Kré76a]. théoriques [LS76]. Theory [AEK78u, Aki78h, AHK76j, AHW77, BK77a, BM78a, BM78b, Bau77e, BK76a, Bib79c, Blo79f, Bow75b, BGM77, BCN77, Cop78f, Cop78i, Cor79e, CS76d, DGS76d, DJ79, DM75c, DS78j, Eps79d, Eve75a, FW79, FGR75i, FMS79, GW79d, Hal79c, HL79b, Hau78a, Hil76, HS75b, HS78, Jam78n, KK78a, KZ76, LSZ77, LMW75, MR77h, MSZ76, MP76, MS77, MS78d, Nat77a, Nat79, OL76, Pom75, Pop77g, Rat78c, RR79b, SW75b, SNWC77e, Sat79h, SS76a, SZ77a, Sch79f, Sil76b, Spr77e, Ste76c, Tol75e, Uen75f, Ver79c, Wal75d, Wer78, Wil75a, Yan77t, tD79l, Alt79b, Ati77, AT75c, Aus79, Bak76, BK77e, BK77f, Bar79h, BW77a, Bau77d, BDE^{+}77i, Bel78, Ber76e, Bin77, Bof77, Bor79c, BF78, Bow75g, Bow75c, Bra79a, Bre77c, Bri79, Bro78a, Bro78c, CK75, Che76d, Chu77a, Cod75, Col76d, Cor79h]. theory [CL75f, CL75g, CN76, Cuy79, DM78a, DFF78, Dej78b, DH77, Dev77e, Dev77f, DS76b, Doh75, Dra79, Dus79, Dyc76, DS78a, DS78l, Eas75, Eav79, Ede75, Edw75b, EH76h, EH76j, Ekl77, Ent78, EGS75, Eps79c, Eps79k, Eps79l, Erd77, EN79, ES77, Eva77, Eve75b, Eve76, FI77, Fed78, Fei78a, Fei78c, Fei78g, Fei78h, Fei78i, FR76, FP78j, FJ75c, Fla75, FK79e, Fro77, Fuc77b, Fuk76, GM77g, Gal76d, Gal76e, Gal76j, Gal76i, Gat77, Gel76a, Gel76c, Gel76d, Gel76e, Get75o, Gil75a, Gil75b, Gli78, Glo75, Gor78a, Gra79b, Gra76k, Gra79c, GV79c, Gre78a, Gui75, Hac78b, Han78b, HL79a, Has78b, Hau79, Hin79, HPS77h, HPS77j, HW75h, HW75i, HW75l, HH77f, Hol76, Iag75b, IŠ79, Jac79a, Jac78a, Jam78a, Jen77b, Jen77f, Joh79f, Jon78b, Joz79, KSWW75]. theory [Kam76, KM78a, Kar77, Kat76, KRZ77a, KRZ77h, KM78c, Keo77, Ker75a, Ker75b, Koc75, Kos77, Kra78b, Kuo76, Kur75, Lau79c, LS77, LS79b, Luc75, Maa78, Mac79, Mac75b, Mal78c, Mar76c, MM75b, May77e, May77i, May77j, May78a, McL79, MN75b, Mey75a, Mis79a, Moi78, Mol77m, Mor78c, Mor77b, Mor77c, Mor77d, Mos76b, NT79b, Nam79f, Nar78, Nas79a, NV79c, Nat77b, Ne’78, Nie78, Obe78, Oli76, Ols78a, Par76, Ple75, Pop77a, Pop77i, Pop77h, Pop78, Ran79a, Rau75a, Rau75b, Ree76d, Rei78a, Rei79, Rob76a, Rog77c, Ros75a, Ruc78a, Rus77, RN75, Sac75, Saf77, Sak78, Sat79d, Sch76s, SS76d, Sco77, Sco79, Seg77, Seg78, Sen76c, Sha75, Sha79a, Shi78, Shu75, SS78a, Sin77, Smi77, SW78k, Sna75, Soc76, Soc77, Sri79f, Sta79b, Sta78c, Ste79, Ste76d, Ste78j, Sto76]. theory [Sto75b, Str79b, Szc77a, Tai75, Tak78b, TW78a, Tho76a, Tho79b, Tol75a, Urb75a, Wag76, Wal78c, Was76, Wed78, Wei75a, Wei76d, Wer75, Wil75g, Wil75f, Wil75h, Wil78a, Wu78, tD79e, tD79i, vW78, vW76, AL78, FK79d, Mol77-28]. thermodynamic [Bow75d, Rue77]. Thermodynamics [Blo79h, Pre76j]. Thermodynamique [Sou78, Rue77]. Theta [Acc75d, AT75a, AT75f, Eic77, AT75b, AT75e, Far79, Kud77]. theta-functions [Kud77]. thick [Mus76]. thin [Wah79]. Third [GLL79, MP76, SW75d, Meh78a, Tay77b]. Thirring [Car78c]. Thom [KS79a]. Thomas [Bon75]. Thomsonsches [Küh79]. Thorin [Ben77a]. Three [Rév76, Rog77a, Han78a, KK79, Mil78a, dP79]. three-designs [Han78a]. three-part [dP79]. Thurston [RS78, Rou78, Ser79a]. tidal [BZ76]. Tierney [Car79e]. tightly [Gui77c]. Time [DGS76r, Kre78c, Wil75h, Alt79a, BF76a, CW77a, Cie75a, DDP78, EW78, GR79j, HM77, Raţ79, SW78m]. Time-dependent [Wil75h, GR79j]. Time-discretisations [Kre78c]. time-stepping [DDP78]. times [DV75c, DG77, Get75a, Poh75, Ros76b]. tiny [Rau75c]. Tits [Mat75, Mat77c, Mat77g]. Todd [War77a]. Toeplitz [Cow78, HH77a, TY78]. tools [Gal76i, IP76a, Osi75a, Pin79]. Topics [Che76j, Die75c, Gol75, HS75b, NV79c, RR79b, Rei79, RN75, Sch75s, Bru78i, Dei77i, Egg78, Gal76f, Sus79, NR78b]. Topoi [LMW75, MR77h, Dus79, MR77c, MR77g, MR77i, Mau75, Osi75a, Vol75b, Wra75, Osi75b]. Topological [AEK78t, AEK78u, BJM78a, Bro75d, Bro75e, Brü76b, CV77j, Chu77a, DGS76s, DGS76t, Dup78k, Hal79c, Har79b, Hof76a, Hof79l, HK78b, KS77, LM79, Moi78, MT76b, Nov77a, Ste78c, Wal75i, Wal75j, WdPL76, AEK78e, AEK78j, AEK78l, AEK78a, AEK78m, AEK78n, Adl79b, Aus79, Bar79h, Bol78, Bör79b, Bou76a, BH76c, DGS76u, Egu79, EL76, HM79a, Han78b, Hel79b, Her76, Hof76b, Led79, Loo76, May77e, Mis79b, MT76c, Nel76, Par75b, Smi77, Ste79]. Topologie [Bea78c, Fd76, LV75, Ser76]. Topologies [Lel75a, AEK78s, Bin75i, Bra79s, Fey78b, Kan79, Pap79b, Won76a]. topologique [Gol78a, Maz77a]. Topologiques [Bea78f, BF75f, Con75b, Lel78b]. topologischer [GMK75b]. Topologization [Fla77]. Topology [BH76b, Coh78, DM79b, EH76c, GR75b, HP79, HPS78, HS79b, JMR75, Mil78b, Moi77g, Nov77a, PdC77, Sch78g, Sto76, Won76d, BDE^{+}77h, Ben76b, Ber79c, Bra79o, Com77, EH76a, EV77, Eps77, Get75o, HS79a, Juh77, Rud77a, Trn77, Won76b, Fen79]. topos [KLM75]. toposes [Hyl79, JW78]. Torch [Bra79g, Bra79h]. tordues [WK79]. tore [Bru75, Her75a, Her75b]. Torelli [Pet78a]. Tores [Gér75h].

Tori [Man75c, Gér75h, Kni78, Sri79d]. Tornehave [May77a]. Torsion [ACH77h, CM78b, Har78a, Web78d, And77h, Arn77, Chi76c, Lad77, Lie77, Mun79, Nic77, Sch78f, Wag77]. torsion-complete [Lie77]. torsion-free [Sch78f]. torsors [Dus79]. torus [Has78a, Ioo76, Lit79, McG78]. Total [GV79a, OP79, Ran79b]. Totalement [BWK77b, Gil78j]. Totally [Tho76b, Ell79, Lac76]. Toulouse [Lel77a]. Toupine [Mol77-28]. tour [Wil78a]. tournament [BB75]. Trace [Hej76a, Adl79a, Hej76b, Hej76d, Hej76e, Hej76f, Her77c, LT76e, Zag77a, Lac78]. traces [Del77a, Del77f, Mat78b, Nic75a]. traffic [AH76]. trails [Bos78]. trains [Gat77, Rin76]. Traitement [Tik79]. trajectoires [Fer75, Kôn76]. tranformations [DGS76i]. transcendants [Wal77a]. Transcendence [Mas75b, Mas75f]. transcendency [Mah76g]. Transcendental [Chu79, DL76, Lac76, Mah76c]. Transfer [Che76k, Che76c, Dol78, Fes78]. transfers [Got78b]. transfert [Lem77, WK78]. transfinis [BF75d]. transfinite [Küh79, Poh75]. transform [AB75, AT75d, HH77b, HH77d, New76a]. Transformation [Ber76c, Bre77n, tD79l, AT75c, Bon75, ET79, HO79, Smi77, Wah76]. Transformationen [Jur78m]. Transformations [Bru78o, Miu76a, Bre79e, Che76a, CS76b, FM76a, FW78, Hof79a, Kow75, Lam76, Mis79b, Mos77b, New76a, Rog76a, Run76, Wal75f, Wal75g, Miu76a]. transformées [Nic75a]. Transforms [Far76m, Bar79a, Kaz76, Sco76]. transient [GH77]. Transition [Goo78, CY78, Föl75]. transitive [PR77c]. transitiver [Kle77j]. transitivity [DGS76n, Eas78]. translation [EWL78, HH78a, Rah75]. Translationen [Jan79a]. translations [Rah75]. transmission [Pir78]. transmutations [Her75c]. Transport [MR77m, BLP79, Ne’78]. transversality [Dej78a, Lak79b]. Transverse [Dev78]. travail [DM78b, Mok78c]. Travaux [Kos75b, Mal78b, Sch79g, Tit77, Wal77a]. traveling [CS76c]. treating [Pop77j]. Treatment [AT78, BGS78, Col78, Gus75, Mor76]. Trebels [Sjö76a]. tree [Eps79j, Fri79]. tree-structured [Fri79]. trees [Dob75, Eps79j, GL78, KTR79, McA76, RF79, Rog77b, ST77a]. très [Ray78b]. triangle [Rog77a]. triangles [BC77a]. Triangular [Bou78a, Ład75]. triangulées [Gro77f]. tribus [DM78b, YM76]. Tribute [MSZ76, SW75b]. trigonometric [FJ75s]. trigonométrigues [Del77a]. trisecant [Lau78]. trivalent [CCMS75]. Trivial [FGR75i, KT75i, Osi78, Rou79]. trivialen [Haz77a]. Tromsø [Ols78b]. troncature [Rot76b]. true [Her76]. truncated [GK78a, Jur75, Rob79a]. truth [Pra75]. Trzebieszowice [Wer78]. Tschebyscheff [Cro76, Bra76b]. Tschebyscheff-Approximation [Cro76, Bra76b]. Tucker [FJ75g]. Tukey [Mal79d]. Tulane [Gol75]. tuple [Kos79]. Turbulence [Sch76b, BPFS76, Bar78a, Cho77a, Gla77, LS76, Man76a, Man77, Mar77, Rue76, Sma77, BR77b, Tem76c]. turbulent [Nør78d, Saf77]. turn [Lam76]. turning [Kik79]. Tverberg [TH75]. Tverberg-type [TH75]. twist [Dav78f]. Twisted [Bar78d, Hum76f, Tay77b]. Twistor [Joz79]. twists [Joh79d]. Two [BKM78, Brü79a, Buh78h, Bul79, Cla79f, Ess75j, GM77m, Gar77b, Gla77, HP76a, Hub78b, Jon78b, Juh77, Lam78g, Mos76b, Sch77e, TDHJ76, AB75, Boh78, Bre75b, Coh77, Duk78, Ess75g, Gla76b, Gut75, GdM77, Iva79c, Kom78, Koo77, LP79a, McC78b, Mol77h, New76b, Per78a, Qui79, Ree78b, Sch75h, Ste79, Tho76a, Utr79]. two-cell [Kom78]. two-complexes [Duk78]. Two-dimensional [Buh78h, Gla77]. two-fold [Tho76a]. two-manifolds [Gut75]. Two-point [GM77m, LP79a, McC78b, Per78a]. two-variable [Koo77]. Type [Mat77c, AM78, And77k, Ave76, Ber76c, Bon75, Bop79, Bro75h, Bru76e, Com76, Cou78, Dam78, Dei77c, Dei77g, DGS76q, Des79a, EH79a, Eva76, Fef75c, Fer77, FJ75g, FJ75p, FS79b, Gab75a, Gas75, Goo76e, Gro78i, Gue78, Han75a, Hin77, HLS79, Ish76a, Kik79, Kla79, Kup75, Luc75, MW78, Mar79d, Mat77g, Mol77j, Mül75b, Nic75a, Ood78, Per78c, Pra75, Ree78a, Rin75, Ros78a, Sab76, Sai78, Stø78b, SV75e, Tak79, Tak77b, Tat75, TH75, Van78c, Wie79, SV75f]. type-free [Fef75c]. Types [Mol77f, Adl79a, Bin75g, Bof77, Hil77b, KV79b, Mol77g, Mol77h, PR77c, SW75c, Sco75, Sim75a]. typical [Bru78a]. typically [Sch75q].

Übergangsmatrizen [Jur78o]. UCLA [KM78b]. Ulm [Lie77, Sta77c]. Ultrabarrelled [HK78j, HK78e]. ultrabornological [HK78j]. Ultrafilters [Hin79, Com77]. ultrapowers [HW75r]. umbilic [Zee76f]. unbounded [BF76a, GH79, Sak78, vdLT79]. unbranched [Gér75h]. uncertainty [Ber79d]. unconditional [Lac77, Peł75]. uncorrelation [Tay78e]. Uncountable [Bel76, Osi78]. undecidability [Eps79k, Eps79l]. undirected [Coc78]. Unendlich [Maa75]. Unfoldings [dP76b, Was76]. Ungeichungsnebenbedingungen [Cro76]. uni [Poé76a]. unicité [Kis76a]. Unified [GK78d, AL79, KS76]. Uniform [Aki78i, Bro75d, Bro75f, GH79, Won76d, Bab76, BK76g, EV77, Fro77, GS76, JL77, Peł76, Sch75d, Sib76a, TY78, TDHJ76, Won76b]. uniforme [Aup79d, MS79b]. uniformement [Bam79]. Uniformes [Bea78d, Web77]. uniformities [Pel77]. uniformity [Blo79d, CV77f, HPS77k]. Uniformly [Die75f, Mil78a]. unilateral [BGJN76, Hla79]. unions [Oso79]. Unipotent [War76l]. unipotente [Kra75b]. Unique [BZ78, Jac75a, Kas79, Sch79h, Bos78]. Uniqueness [Gri79e, BZ76, Buc77, Dou78, Eis76, FJ75q, Pet78c, Rau78, Shn78, Yam76, Kis76a, Hör75]. unit [BK77d, BK77e, Gro78n, Hay77, Tay77a]. unitaire [dlH79b]. unitaires [Car76b, Cow79, Del79, Gui79a, Ver75]. Unitarity [Oli75]. unitarizability [Wal75a]. unite [Gau75, Buc78]. Units [KL77]. Univalent [Sch75s, Tam78c, Dur76b, Keo77, Pir77, Sch75h]. Universal [Bin75h, BHP78a, Gla76k, Haz79a, HW75g, Pop77d]. universality [Pac76]. universe [Maa78]. universel [Roo78c]. universelle [Bou79a, Gol78a, Maz76]. universes [Bec78b, Jer78, Mau75]. Universidade [Sch78g]. Université [Cha75c, DMW77, DMW78, DMW79, Foa77a, Mey75d, Mey76h]. University [AHW77, BH78, BCD77, Bar77b, BK75a, BR77c, BRP78, BCN77, BS77, CW76a, Cha78a, CS76d, Cro75b, FW79, Gol75, HL79b, HPS78, HS79b, Kel75c, KZ76, Man75d, MMP78, MS77, Nat77a, NR78b, Pin75a, Ros75b, SZ77b, SZ77c, Ste76c, SW75d, Tem76c]. unknotted [Mor79c]. unlabeled [Rob77c]. unpublished [Ran77]. unramified [Ish76h]. unreliable [Cir78]. Unsolvability [Eps79d, Eps79h]. unsolvable [Bör75b]. unsolved [Nar78]. unstable [FJ75r]. unusual [BG78b]. upper [FR79j, Poh75, Rat78q, War76m]. Urelements [MS76b]. Ursell [BK76g]. USA [AK78]. use [Est76, KMS75, Obe78, Sne75, TW78a]. used [LW75]. Useful [Ano75, Trn77]. users [Far76j]. using [BF76a, Ead77, Far76d, Gli78, Möl77-30, Nør78d, Wei76h, dP79]. USSR [MP76]. Utah [GR75b].

V [BKK76, Lit77a, Rog76b, SZ77b, Wal77a, Kle77e, LSZ77, Man77]. V. [Dem75, Ono77, Tit78]. Valeur [Meb75, Pue78]. Valeurs [And75, LR78a, Ser75p, BLC78, Lum76a]. valid [MR77a]. validation [Utr79]. validity [Man78a, vdB79]. Valiron [Fuc77b]. Valleys [GM79b]. valuated [Coo77, Hil77a, Ric77]. Valuation [Bra79q, Bra79f, Bra79n, Lad77]. Valuations [Bra79r, Lie77]. Value [BTW75b, Sto77g, Ama76a, BK77e, Boh78, BTW75f, Col78, Dah76, Dra79, FR76, Fuč79, GM77c, GM77m, GW79k, Gri76b, Gus75, Hes79, Hla79, Koß78, Kra79, Leh79, Lio75b, LP79a, MR79, McC78b, Mil76, Neu79, Per78a, Red79, Stu76a, Wal78a]. value-distribution [Dra79]. valued [Bar78e, BW76, Bel78, Bom77, Bor79a, BD78, Car79d, Cha76c, CS79, DJ77, DB78, Fuc77a, Gie76, Goo78, Gór78b, Gra79c, Hej76f, Ish76i, Kue76b, Kue76a, Kue79, Kwa76, MM79a, MR77b, Mar76a, Max78, MS75, Mét78a, MS78c, NS76, Phi79, Ras75, Ros76a, Sch78e, Szu78, TW78b]. Values [BWK77e, Hay77, HH77a, Raz77, RY77, AR79, CV77j, Dam77, GMZ79, Mah76d, Mah76f, Mah76g, MT77b, Rei76a, Sko76, Tho76b, VGK79, Vuo79]. Values-Finis [BWK77e]. Vancouver [HPS78]. vanish [Ree78b]. Vanishing [Mas75h, Mas75g, Gar78, Mas75e, Mos79]. vanishing-theorems [Gar78]. Variable [BK75a, Pow79, SZ77b, SZ77c, Alt79a, Bra76a, Coh77, Far76k, FJ75d, Koo77, Qui75]. Variables [Nor75, Nor78c, SZ77a, BW76, CG79a, Coh77, DJ77, FJ75d, FJ75a, Gie76, GMZ79, Gra76f, Gra76i, GH75, Jur75, Kue76b, Kue76a, Las75, Mar76a, Möl77-30, Phi76, Phi79, Ram79a, Rei77c, Ros76a, Sch76o, Sch77h, Suz79, Tay78e, Wal76, Wal77b, Sch77c]. variance [Ros76b]. variant [Cha76a]. Variante [Pfe75]. Variation [Tam78o, Bru79b, Sch75e, Sen76a, Str79c]. Variational [Cap77, Run76, BL75, BS76a, DGS76u, Geo79a, Led79, Lio75a, Lio76, Mos76a, Mos77a, Sch75i, Sch75t, Ste78h]. variationnelle [Mos78b]. variationnelles [Bai76]. Variationnels [MR77m]. Variations [Cam78b, GDMW76, Ram78b, Ram78a, Bal76, BF76b, FP78l, Ram78c, Tem76a, Ded77]. Variationsungleichungen [Jeg75]. Varietäten [Jan79b]. variété [Ali75, Hen75b, Jen78, Le 77, Sib77, dV75a]. variétés [Blo77, Bou79b, Her78b, MvM77, Ver76, HH78b]. Varieties [Kov78, Pop77g, Smi76e, Uen75f, Ear79, Hal79b, Laz75d, Lub79a, Mor77c, Pop77b, Pop77i, Pop77f, Pop77h, Pop78, RT78, Sch76r, Tog76, Uen75b, Uen75c, Uen75e, Yan77k, Yan77n]. variety [Hol79, Jos77, Pop77i]. Various [Eps79m, FS78a, Kue75b, LMW75, Pos76, GP77]. Varying [Sen76b, Sen76c]. Vaught [Ste78a]. Vecteurs [Car76b, Her75a]. Vector [AEK78u, Bin75i, Cha76c, HK78b, NS76, Pet76b, Pet76j, Wal78d, Wer78, Won79g, AEK78e, AEK78j, AEK78m, AEK78n, AdS75b, AdS77, Bel78, But78, CS79, Dek79, DB78, Die75d, FI77, Fuc77a, Goo78, Goo79c, GdM77, Hal76b, Hej76f, Hil77a, HK78d, HK78f, HK78g, Ish76i, Kno76, Kwa76, Lap79, LeP78a, MM79a, Mic78b, Muj78, MRN78, OW78, Won76c, dlR75, tD79h, AD78b, AD78a, Bom77]. Vector-valued [Cha76c, Bel78, Goo78, Hej76f, MM79a, Bom77]. vectoriel [Ari79, Lel75a, Maz77a]. vectorielle [Sam76]. vectorielles [Buc78, Lel75a, Sch76d, Yor79]. Vectoriels [BF75c, Cam78b, Lap79, Le 77, Lel78a, Lel78b]. vectors [KN79, Nie78]. Vektorräumen [Hei76a]. Vektorwertige [Mét77f]. velocity [Gud76, Nør78d, Sza75]. verallgemeinerte [Küh79, Voi77e, Fri78a]. verallgemeinerten [Vek76]. Verbindungsmatrizen [Jur78q]. Verbindungssysteme [Jur78r]. Verbindungssystems [Jur78m]. Verfahren [Fri78a]. Verfahrens [Bra76b]. Vergleichssätze [GMK75h]. verifiant [Rit76]. Verma [Kos75a]. versal [Pal78a]. Verschiebungsprinzip [Jan79i]. verschränkte [Voi77j, Voi77t]. verschränkten [Voi77b]. versel [Poé76a]. verseller [FK79b]. versels [Mar76d]. version [Pra78, Sza75]. vertex [Ham75a]. Vertical [Bin75b]. Vervollständigungen [Gro79b]. Vey [Gar78, Her77a]. VI [HJLN77, HW79b, SZ77c, HJLN77, Kle77f, Rob79c]. via [Ask75, Bru78o, BW75]. Victoria [Pin75a]. vidage [Foa77c]. vidage-remplissage [Foa77c]. video [Abr76b]. view [Eav79, Rud77a]. viewpoint [Boh78]. VII [DCHRH78, ACF79, DCHRH78, Kle77g, Lan77, Wil77a]. VIII [Bow77, Uen75d, dBS76]. Villard [And77a]. Visco [Zie75]. visco-plastic [Zie75]. Visco-plasticity [Zie75]. Viscoélasticité [GR77]. viscoelasticity [BGJN76, Bri79]. viscosity [BDNY77]. Viscous [AG76, ADB75]. Vitesse [DC78]. Vlasov [Rau78]. Vogan [Duf79a]. voisinage [Dar77]. vol [Ano76, Ano77, Ano79, BDV^{+}78, Bou75]. Volterra [Aiz79, Bak78b, Bar79b, Bru76e, CG79b, Cla79g, Clé79, Cus79, GV79b, Gro79h, Jor79, Lak79a, Lon79a, LS79a, MW79, Mil79, Noh79a, Noh79b]. Volume [Hej76a, Hen75a]. vorgeschriebenem [Man78c]. vortex [Gui77c]. vorticity [BZ76]. Vorwort [Pfl75e]. Vries [Adl79a, Run76, Wah76].

W [Bon75, Ker75f, RR79b]. W* [Sai78]. W. [Bor78c, Car78c, Rou78]. w.r.t [TDHJ76]. Wadge [vW78]. Wahrscheinlichkeitsmaßen [Haz77e]. waists [AGR75]. wakes [HH77h]. Waldhausen [Lod79]. walk [HvW79]. Walks [MT76b, Bos78, Che78b, Cré76, LR79, MT76c]. wall [Nør78d]. Walsh [Bor79a, Col76e]. Warfield [HRW77]. Warwick [Man75d]. Washington [FW79]. water [Nor78a]. Waterloo [HL79b, HS79b]. Wave [BK76f, Ree76a, Wil75i, BBB76, CS76c, Gat77, Gre75b, Hal79i, HBT75, Jak79, Med75, Rin76, Wil75b]. waves [BZ76, BBB76, Gud76, Lei79, Lon76, Nor78a, OW75, Ree78a, Ros78i, Wah76]. way [Geo78a, Tho76a]. Wazewski [Con76a]. Weak [BK77k, CB76, Lin79a, Mey75a, Neu76, SS76e, Ben78, Jer76a, Jer76b, Rat78f, Rau78]. weak* [Hal76b]. Weakly [DJ77, Die75g, Hay78, Bal76, DR75i, Phi76, Rod79]. Wecken [Gro77e]. wedge [BEGS75]. Weierstrass [Cox79, Iit79, KK79, Ste75d, KPP^{+}78f]. Weierstraß-Grauertsche [KPP^{+}78f]. weighing [EW76, Mul75c]. weight [Buh77, Buh78a, Hej76d, Jan79f, Jan79g, SS77b, Sta77d]. Weighted [EWL78, HH78a, Tay78j, PT76, RF79, Tay78a, Tay78e, Tay78i]. weights [GHM78, Lap79, Wal75a]. Weil [Li79, Dup78c, Dup78d, Dup78j, Gel76g, KT75j, Ral79b]. Weinstein [Bre78]. Weitere [Jur78s]. Well [CG79b, Mil79, BW77a, Lub79a, Pol79, Shi75, Gue75]. well-orders [BW77a]. Well-posedness [CG79b]. Wellen [Fri76b]. Wellordered [FF78]. Wentzell [Ood78]. Werteverteilung [Kna79a]. wertigen [Bin76]. wet [APF76]. Weyl [Blo79a, Mat77b, Mat77f, Mil75a, Ply79, Rob77d, Spr77f, Vaa75]. Where [Mas79, Hay76]. Which [GW79d, McG78, BY77, Ber79a, Bin75i, Ear79, Her76, Mar79c, MJ75, Oma75, Rah75, Yan77c]. White [Hid78]. Whitehead [And77h, DS78l, Hae78b, Lod76, Mun79, Nun77, Tit78]. Whitney [BM77a, FJ75p]. Whittaker [Rod75]. whose [Bra79d, Ree78b, Tay77a, Zag77b]. Wick [Pes76a, Sło78]. width [FJ75p]. Wiener [Cau76g, Egu79, Gua78, Hud78, Kuo75d]. Wightman [Roy76a]. wild [Dav75a]. Wilder [Jon78a, Mil78b, Ray78a]. Wiles [Lan78]. Wiman [Fuc77b]. winded [Gui77c]. Winter [CUW75, CUW75]. Wirtinger [Eve76]. Wishart [Far76i, Far76n]. Without [AEK78u, GM77m, Lin79a, Mas75h, Obe77, vdLT79]. Witt [ACH77f, Bae78f, Cha76a, Ewi79]. WKB [Woo78b]. work [Bor78c, Con75a, Ran77, Ray78a]. Workshop [GR79b, Miu76a]. Wreath [Hof79d, Hof79k, Ker75c]. writing [FR79k]. written [Mol77-28, Uen75f]. WRNP [MRN78]. Wynn [Cor79a].

X [Mey76h]. XI [DMW77]. XII [DMW78]. XIII [DMW79].

Yang [Car78a, Raw79]. Yau [Bou79b]. years [Ono77, Wil78a]. yields [PE75]. Yor [DM78b]. York [Nat77a]. Young [Foa77c, Jam78u, dBR77].

Zahorski [Bru78p]. Zalesskii [Goo79a]. Zariski [Bra79s]. Zelinsky [Gar76a]. Zerlegungen [Voi77d]. Zerlegungssatz [Jan79a, Voi77g]. Zero [Ber76d, Cho76, HW75o, Mis79b, Uen75b, Szp76]. Zero-divisors [Ber76d]. zero-sets [Cho76]. zeroes [Ber76a]. zeros [Bom76, CMPY79, Edr77, Far79, Gro78p, HW79a, Pin79, Sch77f, Las76b]. zeta [And77k, Bom76, Fit78, Hej76c, Zag77b]. zeta-functions [And77k, Zag77b]. ZF [MM75b, Rog76b]. ZFC [Rog76b]. zonal [Far76a, Far76j]. Zufallsentwicklungen [Haz77d]. Zufallsvariable [Cré79]. zulässigen [Jan79a]. zur [Fri78a, JLR76, Mic76c, AE77, Bau76, Bra76b, Bro76a, Cro76, Häm76, Jeg75, Kna79a, KPP^{+}78i]. Zusammenhang [Gro79c, Nic78]. Zusammenhänge [GMK75e]. Zustandsbeschränkungen [MH75]. Zweidimensionalen [Som76]. zweidimensionaler [KPP^{+}78d]. Zwischen [Gro79c, GMK75b, Nic78].

References

Aaronson:1979:CEW

[Aar79] Jon Aaronson. On the categories of ergodicity when the measure is infinite. Lecture Notes in Mathematics, 729:1–7, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063276/.

Al-Bassam:1975:TTD

[AB75] Mohammed Ali Al-Bassam. H − R transform in two dimensions and some of its applications to partial differential equations. Lecture Notes in Mathematics, 457:91–105, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067099/.

Antman:1976:LIC

[AB76] Stuart S. Antman and Russell C. Browne. Local invertibility conditions for geometrically exact nonlinear rod and shell theories. Lecture Notes in Mathematics, 503:159–165, 1976. CODEN LN-MAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088751/.

Abraham:1976:MR

[Abr76a] Ralph Abraham. Macroscopy of resonance. Lecture Notes in Mathematics, 525:1–9, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077840/.

Abraham:1976:SCV

[Abr76b] Ralph Abraham. Simulation of cascades by video feedback. Lecture Notes in Mathematics, 525:10–14, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077841/.

Abrahamse:1978:MO

[Abr78] M. B. Abrahamse. Multiplication operators. Lecture Notes in Mathematics, 693:17–36, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064658/.

Accola:1975:P

[Acc75a] Prof. Robert D. M. Accola. Part I. Lecture Notes in Mathematics, 483:1–31, 1975. CODEN LNMAA2. ISBN 3-540-07398-1 (print), 3-540-37602-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081274/.

Accola:1975:PIa

[Acc75b] Prof. Robert D. M. Accola. Part II. Lecture Notes in Mathematics, 483:32–65, 1975. CODEN LNMAA2. ISBN 3-540-07398-1 (print), 3-540-37602-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081275/.

Accola:1975:PIb

[Acc75c] Prof. Robert D. M. Accola. Part III. Lecture Notes in Mathematics, 483:66–99, 1975. CODEN LNMAA2. ISBN 3-540-07398-1 (print), 3-540-37602-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081276/.

Accola:1975:RST

[Acc75d] Prof. Robert D. M. Accola. Riemann Surfaces, Theta Functions, and Abelian Automorphisms Groups, volume 483 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07398-1 (print), 3-540-37602-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 105 pp. LCCN QA3

.L28 no. 483. URL http://link.springer.com/book/10.1007/
BFb0081273;   http://www.springerlink.com/content/978-
3-540-37602-6.

Accola:1979:CST

[Acc79] Robert D. M. Accola. On certain subloci of Teichmüller space. Lecture Notes in Mathematics, 743:1–9, 1979. CODEN LN-MAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079480/.

Arsene:1979:IDV

[ACF79] Gr. Arsene, Zoia Ceauşescu, and C. Foiaş. On intertwining dilations. VII. Lecture Notes in Mathematics, 747:24–45, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063957/.

Alexander:1977:DI

[ACH77a] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. W_{∗}(Q, C_{n}) and dress induction. Lecture Notes in Mathematics, 625:93–135, 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069516/.

Alexander:1977:AP

[ACH77b] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. Algebraic preliminaries. Lecture Notes in Mathematics, 625:6–33, 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069513/.

Alexander:1977:CNF

[ACH77c] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. Cyclotomic number fields. Lecture Notes in Mathematics, 625:64– 93, 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069515/.

Alexander:1977:I

[ACH77d] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. Introduction. Lecture Notes in Mathematics, 625:1–6, 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069512/.

Alexander:1977:MHF

[ACH77e] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. Metabolic and hyperbolic forms. Lecture Notes in Mathematics, 625:158–198, 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069518/.

Alexander:1977:OOG

[ACH77f] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. Odd Order Group Actions and Witt Classification of Innerproducts, volume 625 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 202 pp. LCCN QA3 .L28 no. 625. URL http:

//link.springer.com/book/10.1007/BFb0069511;   http://
www.springerlink.com/content/978-3-540-36974-5.

Alexander:1977:SAE

[ACH77g] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. Smooth actions of c_{P^{n}} on even dimensional manifolds. Lecture Notes in Mathematics, 625:33–64, 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069514/.

Alexander:1977:TST

[ACH77h] John P. Alexander, Pierre E. Conner, and Gary C. Hamrick. Torsion signature theorem. Lecture Notes in Mathematics, 625:135– 158, 1977. CODEN LNMAA2. ISBN 3-540-08528-9 (print), 3-540-36974-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069517/.

Aron:1978:VSMb

[AD78a] Richard M. Aron and Seán Dineen, editors. Vector Space Measures and Applications I: Proceedings, Dublin 1977, volume 644 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 644. URL http://link.springer.com/book/10.1007/
BFb0066835;   http://www.springerlink.com/content/978-
3-540-35906-7.

Aron:1978:VSMa

[AD78b] Richard M. Aron and Seán Dineen, editors. Vector Space Measures and Applications II: Proceedings, Dublin 1977, volume 645 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 645. URL http://link.springer.com/book/10.1007/
BFb0069656;   http://www.springerlink.com/content/978-
3-540-35903-6.

Argyris:1979:CSF

[AD79] John H. Argyris and Padraic C. Dunne. Constant strain finite elements. Lecture Notes in Mathematics, 704:75–110, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063616/.

Adamowicz:1975:OPS

[Ada75] Zofia Adamowicz. An observation on the product of Silver’s forcing. Lecture Notes in Mathematics, 499:1–9, 1975. CODEN LN-MAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079415/.

Adams:1978:MBC

[Ada78] J. F. Adams. Maps between classifying spaces. Lecture Notes in Mathematics, 658:1–8, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068707/.

Argyris:1975:PAP

[ADB75] J. H. Argyris, P. C. Dunne, and B. Bichat. A plate analogy for plane imcompressible viscous flow. Lecture Notes in Mathematics, 461:33–46, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074147/.

Adler:1979:TFF

[Adl79a] M. Adler. On a trace functional for formal pseudo-differential operators and the Hamiltonian structure of Korteweg–de Vries types equations. Lecture Notes in Mathematics, 755:1–16, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069801/.

Adler:1979:SPT

[Adl79b] Roy L. Adler. A selection of problems in topological dynamics. Lecture Notes in Mathematics, 729:8–12, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063277/.

Arraut:1975:LAP

[AdS75a] J. L. Arraut and N. M. dos Santos. Lie algebra preserving diffeomorphisms. Lecture Notes in Mathematics, 468:19–20, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082583.

Arraut:1975:PSA

[AdS75b] J. L. Arraut and N. M. dos Santos. The point spectrum of the adjoint to an automorphism of a vector bundle. Lecture Notes in Mathematics, 468:20–21, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082584.

Arraut:1977:PSA

[AdS77] J. L. Arraut and N. Moreira dos Santos. The point spectrum of the adjoint of an automorphism of a vector bundle. Lecture Notes in Mathematics, 597:56–67, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085347/.

Albrecht:1977:NIK

[AE77] Prof. Dr. J. Albrecht and Prof. Dr. H. Engels. Zur numerischen Integration über Kreisbereichen. (German) []. Lecture Notes in Mathematics, 571:1–5, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086560/.

Adasch:1978:LTSb

[AEK78a] Norbert Adasch, Bruno Ernst, and Dieter Keim. σ-locally topological spaces. Lecture Notes in Mathematics, 639:92–97, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068532/.

Adasch:1978:BST

[AEK78b] Norbert Adasch, Bruno Ernst, and Dieter Keim. The Banach– Steinhaus theorem. Lecture Notes in Mathematics, 639:38–40, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068522/.

Adasch:1978:BSC

[AEK78c] Norbert Adasch, Bruno Ernst, and Dieter Keim. Barrelled spaces and the closed graph theorem. Lecture Notes in Mathematics, 639: 41–46, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068523/.

Adasch:1978:BSO

[AEK78d] Norbert Adasch, Bruno Ernst, and Dieter Keim. Barrelled spaces and the open mapping theorem. Lecture Notes in Mathematics, 639:47–50, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068524/.

Adasch:1978:BTV

[AEK78e] Norbert Adasch, Bruno Ernst, and Dieter Keim. Barrelled topological vector spaces. Lecture Notes in Mathematics, 639:31– 37, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068521/.

Adasch:1978:BS

[AEK78f] Norbert Adasch, Bruno Ernst, and Dieter Keim. Bornological spaces. Lecture Notes in Mathematics, 639:60–65, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068526/.

Adasch:1978:BSS

[AEK78g] Norbert Adasch, Bruno Ernst, and Dieter Keim. Boundedly summing spaces. Lecture Notes in Mathematics, 639:74–78, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068529/.

Adasch:1978:CCG

[AEK78h] Norbert Adasch, Bruno Ernst, and Dieter Keim. Completeness and the closed graph theorem. Lecture Notes in Mathematics, 639:51–59, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068525/.

Adasch:1978:DSS

[AEK78i] Norbert Adasch, Bruno Ernst, and Dieter Keim. (DF)-spaces and spaces with a fundamental sequence of compact sets. Lecture Notes in Mathematics, 639:98–108, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068533/.

Adasch:1978:ILT

[AEK78j] Norbert Adasch, Bruno Ernst, and Dieter Keim. Inductive limits of topological vector spaces. Lecture Notes in Mathematics, 639:19–23, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068519/.

Adasch:1978:I

[AEK78k] Norbert Adasch, Bruno Ernst, and Dieter Keim. Introduction. Lecture Notes in Mathematics, 639:1–4, 1978. CODEN LN-MAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068515/.

Adasch:1978:LTSa

[AEK78l] Norbert Adasch, Bruno Ernst, and Dieter Keim. Locally topological spaces. Lecture Notes in Mathematics, 639:79–83, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068530/.

Adasch:1978:MTV

[AEK78m] Norbert Adasch, Bruno Ernst, and Dieter Keim. Metrizable topological vector spaces. Lecture Notes in Mathematics, 639:10– 15, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068517/.

Adasch:1978:PLT

[AEK78n] Norbert Adasch, Bruno Ernst, and Dieter Keim. Projective limits of topological vector spaces. Lecture Notes in Mathematics, 639:16–18, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068518/.

Adasch:1978:QS

[AEK78o] Norbert Adasch, Bruno Ernst, and Dieter Keim. Quasibarrelled spaces. Lecture Notes in Mathematics, 639:70–73, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068528/.

Adasch:1978:SEC

[AEK78p] Norbert Adasch, Bruno Ernst, and Dieter Keim. Some examples and counter examples. Lecture Notes in Mathematics, 639: 109–117, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068534/.

Adasch:1978:SCL

[AEK78q] Norbert Adasch, Bruno Ernst, and Dieter Keim. Spaces of continuous linear mappings and their completion. Lecture Notes in Mathematics, 639:66–69, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068527/.

Adasch:1978:SAS

[AEK78r] Norbert Adasch, Bruno Ernst, and Dieter Keim. Spaces with an absorbing sequence. Lecture Notes in Mathematics, 639:84– 91, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068531/.

Adasch:1978:SLT

[AEK78s] Norbert Adasch, Bruno Ernst, and Dieter Keim. Strings and linear topologies. Lecture Notes in Mathematics, 639:5–9, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068516/.

Adasch:1978:TDS

[AEK78t] Norbert Adasch, Bruno Ernst, and Dieter Keim. Topological direct sums, strict inductive limits. Lecture Notes in Mathematics, 639:24–30, 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068520/.

Adasch:1978:TVS

[AEK78u] Norbert Adasch, Bruno Ernst, and Dieter Keim. Topological Vector Spaces: The Theory Without Convexity Conditions, volume 639 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08662-5 (print), 3-540-35918-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 125 pp. LCCN QA3 .L28 no. 639. URL http:

//link.springer.com/book/10.1007/BFb0068514;   http://
www.springerlink.com/content/978-3-540-35918-0.

Almkvist:1978:DES

[AF78] Gert Almkvist and Robert Fossum. Decomposition of exterior and symmetric powers of indecomposable Z/pZ-modules in characteristic p and relations to invariants. Lecture Notes in Mathematics, 641:1–111, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064842/.

Ames:1975:BAT

[AG75] W. F. Ames and M. Ginsberg. Bilateral algorithms and their applications. Lecture Notes in Mathematics, 461:1–31, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074146/.

Amiel:1976:VFF

[AG76] R. Amiel and G. Geymonat. Viscous fluid flow in chemically reacting and diffusing systems. Lecture Notes in Mathematics, 503: 148–158, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088750/.

Auslander:1975:MHW

[AGR75] M. Auslander, E. L. Green, and I. Reiten. Modules having waists. Lecture Notes in Mathematics, 488:20–28, 1975. CODEN LN-MAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081214/.

Azlarov:1976:SLT

[AH76] T. A. Azlarov and Ya. M. Husainov. Some limit theorems for a queueing system with absolute priority in heavy traffic. Lecture Notes in Mathematics, 550:1–13, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077480/.

Anderson:1979:BSS

[AH79] Douglas R. Anderson and Ian Hambleton. Balanced splittings of semi-free actions on homotopy spheres. Lecture Notes in Mathematics, 741:196–205, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062140/.

Albeverio:1976:ERGb

[AHK76a] Sergio A. Albeverio and Raphael J. Høegh-Krohn. Expectations with respect to the Gibbs state of the harmonic oscillator. Lecture Notes in Mathematics, 523:86–89, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079834/.

Albeverio:1976:ERGa

[AHK76b] Sergio A. Albeverio and Raphael J. Høegh-Krohn. Expectations with respect to the ground state of the harmonic oscillator. Lecture Notes in Mathematics, 523:80–85, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079833/.

Albeverio:1976:FHI

[AHK76c] Sergio A. Albeverio and Raphael J. Høegh-Krohn. The Feynman history integrals for the relativistic quantum boson field. Lecture Notes in Mathematics, 523:105–114, 1976. CODEN LN-MAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079836/.

Albeverio:1976:FPIa

[AHK76d] Sergio A. Albeverio and Raphael J. Høegh-Krohn. The Feynman path integral in potential scattering. Lecture Notes in Mathematics, 523:26–45, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079830/.

Albeverio:1976:FPIb

[AHK76e] Sergio A. Albeverio and Raphael J. Høegh-Krohn. Feynman path integrals for the anharmonic oscillator. Lecture Notes in Mathematics, 523:65–79, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079832/.

Albeverio:1976:FIF

[AHK76f] Sergio A. Albeverio and Raphael J. Høegh-Krohn. The Fresnel integral of functions on a separable real Hilbert space. Lecture Notes in Mathematics, 523:14–25, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079829/.

Albeverio:1976:FIR

[AHK76g] Sergio A. Albeverio and Raphael J. Høegh-Krohn. The Fresnel integral relative to a non singular quadratic form. Lecture Notes in Mathematics, 523:46–64, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079831/.

Albeverio:1976:I

[AHK76h] Sergio A. Albeverio and Raphael J. Høegh-Krohn. Introduction. Lecture Notes in Mathematics, 523:3–13, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079828/.

Albeverio:1976:IQF

[AHK76i] Sergio A. Albeverio and Raphael J. Høegh-Krohn. The invariant quasi-free states. Lecture Notes in Mathematics, 523:90–104, 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079835/.

Albeverio:1976:MTF

[AHK76j] Sergio A. Albeverio and Raphael J. Høegh-Krohn. Mathematical Theory of Feynman Path Integrals, volume 523 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 137 pp. LCCN QA3

.L28 no. 523. URL http://link.springer.com/book/10.1007/
BFb0079827;   http://www.springerlink.com/content/978-
3-540-38250-8.

Ahlfors:1977:SNA

[Ahl77] Lars V. Ahlfors. A somewhat new approach to quasiconformal mappings in R^{n}. Lecture Notes in Mathematics, 599:1–6, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096817/.

Arnold:1977:AGT

[AHW77] David M. Arnold, Roger H. Hunter, and Elbert A. Walker, editors. Abelian Group Theory: Proceedings of the 2nd New Mexico State University Conference, Held at Las Cruces, New Mexico, December 9–12, 1976, volume 616 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 616. URL http:

//link.springer.com/book/10.1007/BFb0068186;   http://
www.springerlink.com/content/978-3-540-37069-7.

Aizicovici:1979:AVE

[Aiz79] Sergiu Aizicovici. On an abstract Volterra equation. Lecture Notes in Mathematics, 737:1–8, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064490/.

Araki:1978:AAP

[AK78] Huzihiro Araki and Richard V. Kadison, editors. C∗-Algebras and Applications to Physics: Proceedings, Second Japan–USA Seminar, Los Angeles, April 18–22, 1977, volume 650 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 650. URL http://link.springer.com/book/10.1007/
BFb0067386;   http://www.springerlink.com/content/978-
3-540-35850-3.

Akcoglu:1979:PET

[Akc79] Mustafa A. Akcoglu. Pointwise ergodic theorems in L_{p} spaces. Lecture Notes in Mathematics, 729:13–15, 1979. CODEN LN-MAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063278/.

Akin:1978:A

[Aki78a] Ethan Akin. Atlases. Lecture Notes in Mathematics, 662:1– 13, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082265/.

Akin:1978:FDM

[Aki78b] Ethan Akin. Finite dimensional manifolds. Lecture Notes in Mathematics, 662:267–300, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082272/.

Akin:1978:FS

[Aki78c] Ethan Akin. Function spaces. Lecture Notes in Mathematics, 662: 14–43, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082266/.

Akin:1978:GLI

[Aki78d] Ethan Akin. Grassmannians and leaf immersions. Lecture Notes in Mathematics, 662:227–266, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082271/.

Akin:1978:IS

[Aki78e] Ethan Akin. Immersions and submersions. Lecture Notes in Mathematics, 662:174–226, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082270/.

Akin:1978:MM

[Aki78f] Ethan Akin. Manifolds of mappings. Lecture Notes in Mathematics, 662:107–153, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082268/.

Akin:1978:MS

[Aki78g] Ethan Akin. Metric structures. Lecture Notes in Mathematics, 662:44–106, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082267/.

Akin:1978:MTB

[Aki78h] Ethan Akin. The Metric Theory of Banach Manifolds, volume 662 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xix + 306 pp. LCCN QA3 .L28 no. 662. URL http:

//link.springer.com/book/10.1007/BFb0082264;   http://
www.springerlink.com/content/978-3-540-35761-2.

Akin:1978:UHC

[Aki78i] Ethan Akin. Uniform and holder continuity. Lecture Notes in Mathematics, 662:154–173, 1978. CODEN LNMAA2. ISBN 3-540-08915-2 (print), 3-540-35761-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082269/.

Arsove:1976:PTS

[AL76] Maynard Arsove and Heinz Leutwiler. Potential-theoretic semigroups. Lecture Notes in Mathematics, 518:6–13, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080399/.

Adem:1977:ESM

[AL77] José Adem and Kee Yuen Lam. Evaluation of some Maunder cohomology operations. Lecture Notes in Mathematics, 597:1–31, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085345/.

Alavi:1978:TAG

[AL78] Yousef Alavi and Don R. Lick, editors. Theory and Applications of Graphs: Proceedings, Michigan May 11–15, 1976, volume 642 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 642. URL http://link.springer.com/book/10.1007/
BFb0070359;   http://www.springerlink.com/content/978-
3-540-35912-8.

Anselone:1979:UAA

[AL79] P. M. Anselone and J. W. Lee. A unified approach to the approximate solution of linear integral equations. Lecture Notes in Mathematics, 701:41–69, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062073/.

Alexander:1975:MET

[Ale75] Ralph Alexander. Metric embedding techniques applied to geometric inequalities. Lecture Notes in Mathematics, 490:57– 65, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081131/.

Alefeld:1978:SCR

[Ale78] G. Alefeld. Some convergence results for the Peaceman–Rachford method in the noncommutative case. Lecture Notes in Mathematics, 631:1–8, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067459/.

Alexander:1979:NCM

[Ale79] J. C. Alexander. Numerical continuation methods and bifurcation. Lecture Notes in Mathematics, 730:1–15, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064307/.

Alinhac:1975:CEF

[Ali75] Serge Alinhac. Caractérisation d’espaces de fonctions analytiques et non quasi-analytiques sur une variété à bord [d’après M. Baouendi]. (French) []. Lecture Notes in Mathematics, 431: 115–123, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066367/.

Allain:1976:RFD

[All76] G. Allain. La représentation des formes de Dirichlet. (French) []. Lecture Notes in Mathematics, 518:1–5, 1976. CODEN LN-MAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080398/.

Aloon:1979:FCD

[Alo79]  Kenneth Mc Aloon.  Formes combinatoires du théorème

d’incomplétude [d’après J. Paris et d’autres]. (French) [Combinatorial forms of the incompleteness theorem [according to J. Paris and others]]. Lecture Notes in Mathematics, 710:263–276, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069984/.

Alperin:1976:S

[Alp76] Roger Alperin. Stability for H_{2}(Su_{n}). Lecture Notes in Mathematics, 551:283–289, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080009/.

Alpern:1979:GPM

[Alp79] Steve Alpern. Generic properties of measure preserving homeomorphisms. Lecture Notes in Mathematics, 729:16–27, 1979. CO-DEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063279/.

Alt:1979:PSS

[Alt79a] Wolfgang Alt. Periodic solutions of some autonomous differential equations with variable time delay. Lecture Notes in Mathematics, 730:16–31, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064308/.

Altman:1979:STS

[Alt79b] Mieczysław Altman. A strategy theory of solving equations. Lecture Notes in Mathematics, 701:1–40, 1979. CODEN LN-MAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062072/.

Adler:1978:FCS

[AM78] Roy Adler and Brian Marcus. Finitistic coding for shifts of finite type. Lecture Notes in Mathematics, 668:1–11, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101775/.

Araujo:1979:SPC

[AM79a] A. Araujo and M. B. Marcus. Stable processes with continuous sample paths. Lecture Notes in Mathematics, 709:9–32, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071945/.

Azema:1979:RMS

[AM79b] J. Azéma and P.-A. Meyer. Représentations multiplicatives de sousmartingales. (French) []. Lecture Notes in Mathematics, 721: 240–249, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070866/.

Amann:1976:NOO

[Ama76a] Herbert Amann. Nonlinear operators in ordered Banach spaces and some applications to nonlinear boundary value problems. Lecture Notes in Mathematics, 543:1–55, 1976. CODEN LN-MAA2. ISBN 3-540-07867-3 (print), 3-540-38075-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079941/.

Amar:1976:HSM

[Ama76b] E. M. Amar. Hilbert space methods and interpolating sets in the spectrum of an algebra of operators. Lecture Notes in Mathematics, 512:1–7, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080018/.

Amann:1979:ISS

[Ama79] H. Amann. Invariant sets for semilinear parabolic and elliptic systems. Lecture Notes in Mathematics, 703:1–4, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067249/.

Andreotti:1978:CA

[AN78] Aldo Andreotti and Mauro Nacinovich. Convexite analytique. Lecture Notes in Mathematics, 670:341–351, 1978. CODEN LN-MAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064408/.

Ancona:1976:CCE

[Anc76] A. Ancona. Continuité des contractions dans les espaces de Dirichlet. (French) []. Lecture Notes in Mathematics, 563:1–26, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087569/.

Ancona:1978:PEI

[Anc78] A. Ancona. Principe de l’enveloppe inférieure pour des potentiels pris par rapport à une forme de Dirichlet. (French) []. Lecture Notes in Mathematics, 681:1–6, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065864/.

Andronikof:1975:VAB

[And75] Emmanuel Andronikof. Valeurs au bord de fonctions holomorphes se recollant loin du réel. (French) []. Lecture Notes in Mathematics, 482:180–189, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080246/.

Anderson:1977:EEC

[And77a] G. A. Anderson. École d’été de calcul des probabilités de Saint-Flour, 21 août–7 septembre 1977: exposés des participants à la mémoire de Michel Villard. (French) [], volume 591 of Annales scientifiques de l’Université de Clermont; no 65: Mathématiques; 15. fasc. [Université de Clermont], Clermont-Ferrand, France, 1977. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). vi + 157 pp. LCCN QA3 .L28 no. 591; QA1 .C63a; QA273.18. URL http://

www.springerlink.com/content/978-3-540-37356-8.

Anderson:1977:PC

[And77b] Gerald A. Anderson. Poincaré complexes. Lecture Notes in Mathematics, 591:39–53, 1977. CODEN LNMAA2. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096748/.

Anderson:1977:P

[And77c] Gerald A. Anderson. Preliminaries. Lecture Notes in Mathematics, 591:1–27, 1977. CODEN LNMAA2. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096746/.

Anderson:1977:RBS

[And77d] Gerald A. Anderson. Relations between surgery theories. Lecture Notes in Mathematics, 591:101–121, 1977. CODEN LN-MAA2. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096751/.

Anderson:1977:RS

[And77e] Gerald A. Anderson. Relative surgery. Lecture Notes in Mathematics, 591:82–100, 1977. CODEN LNMAA2. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096750/.

Anderson:1977:SCa

[And77f] Gerald A. Anderson. Surgery with coefficients. Lecture Notes in Mathematics, 591:54–81, 1977. CODEN LNMAA2. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096749/.

Anderson:1977:SCb

[And77g] Gerald A. Anderson. Surgery with Coefficients, volume 591 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 157 pp. LCCN QA3 .L28 no. 591. URL http:

//link.springer.com/book/10.1007/BFb0096745;   http://
www.springerlink.com/content/978-3-540-37356-8.

Anderson:1977:WT

[And77h] Gerald A. Anderson. Whitehead torsion. Lecture Notes in Mathematics, 591:28–38, 1977. CODEN LNMAA2. ISBN 3-540-08250-6 (print), 3-540-37356-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096747/.

Andrade:1977:CGC

[And77i] Jorge Soto Andrade. Construction géométrique de certaines séries discrètes. (French) []. Lecture Notes in Mathematics, 597:32– 55, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085346/.

Andrews:1977:IMC

[And77j] George E. Andrews. Implications of the Macmahon conjecture. Lecture Notes in Mathematics, 579:287–296, 1977. CODEN LN-MAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090024/.

Andrianov:1977:ZFR

[And77k] A. N. Andrianov. On zeta-functions of Rankin type associated with Siegel modular forms. Lecture Notes in Mathematics, 627: 325–338, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065307/.

Anderson:1979:LTE

[And79a] Glen D. Anderson. Limit theorems and estimates for extremal rings of high dimension. Lecture Notes in Mathematics, 743:10– 34, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079481/.

Andre:1979:AHA

[And79b] Michel André. Algèbre homologique des anneaux locaux à corps résiduels de caractéristique deux. (French) []. Lecture Notes in Mathematics, 740:237–242, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071065/.

Angelstorf:1979:GBM

[Ang79] N. Angelstorf. Global branching and multiplicity results for periodic solutions of functional differential equations. Lecture Notes in Mathematics, 730:32–45, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064309/.

Anonymous:1975:IOU

[Ano75] Anonymous. An Index and Other Useful Information, volume ?? of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CO-DEN LNMAA2. ISBN 1-4899-4583-0 (paperback), 1-4899-4581-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 80 pp. LCCN QA3 .L28 no. ??; QA1-939. URL http://link.

springer.com/book/10.1007/978-1-4899-4581-5.

Anonymous:1976:SBV

[Ano76] Anonymous, editor. Séminaire Bourbaki vol. 1974/75 Exposés 453–470, volume 514 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 514. URL http:

//link.springer.com/book/10.1007/BFb0080053;   http://
www.springerlink.com/content/978-3-540-38218-8.

Anonymous:1977:SBV

[Ano77] Anonymous, editor. Séminaire Bourbaki vol. 1975/76 Exposés 471–488, volume 567 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 567. URL http:

//link.springer.com/book/10.1007/BFb0096057;   http://
www.springerlink.com/content/978-3-540-37514-2.

Anonymous:1979:SBV

[Ano79] Anonymous, editor. Séminaire Bourbaki vol. 1977/78 Exposés 507–524: Avec table par noms d’auteurs de 1967/68 à 1977/ 78. (French) [], volume 710 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 710. URL http:

//link.springer.com/book/10.1007/BFb0069969;   http://
www.springerlink.com/content/978-3-540-35340-9.

Acuna-Ortega:1979:FD

[AOL79] O. Acuña-Ortega and F. E. J. Linton. Finiteness and decidability: I. Lecture Notes in Mathematics, 753:80–100, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061813/.

Aguirre-Puente:1976:FPW

[APF76] J. Aguirre-Puente and M. Frémond. Frost propagation in wet porous media. Lecture Notes in Mathematics, 503:137–147, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088749/.

Auslander:1975:ASSb

[AR75] Maurice Auslander and Idun Reiten. Almost split sequences II. Lecture Notes in Mathematics, 488:9–19, 1975. CODEN LN-MAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081213/.

Alladi:1979:CIV

[AR79] K. Alladi and M. L. Robinson. On certain irrational values of the logarithm. Lecture Notes in Mathematics, 751:1–9, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062698/.

Araki:1978:KSD

[Ara78] Huzihiro Araki. On KMS states of a C∗ dynamical system. Lecture Notes in Mathematics, 650:66–84, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067390/.

Aribaud:1979:IMA

[Ari79] François Aribaud. Un ideal maximal de l’anneau des endomorphismes d’un espace vectoriel de dimension infinie. (French) []. Lecture Notes in Mathematics, 740:184–189, 1979. CODEN LN-MAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071059/.

Armentrout:1975:DAN

[Arm75] Steve Armentrout. Decompositions and absolute neighborhood retracts. Lecture Notes in Mathematics, 438:1–5, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066102/.

Arnold:1977:GDS

[Arn77] David M. Arnold. Genera and direct sum decompositions of torsion free modules. Lecture Notes in Mathematics, 616:197– 218, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068197/.

Aron:1976:CPC

[Aro76] Richard M. Aron. Compact polynomials and compact differentiable mappings between Banach spaces. Lecture Notes in Mathematics, 524:213–222, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078007/.

Arsac:1979:QAA

[Ars79] G. Arsac. Quelques applications des C∗-algèbres et des algèbres de von Neumann à l’étude des representations des groupes. (French) []. Lecture Notes in Mathematics, 725:1–16, 1979. CODEN LN-MAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062612/.

Artzner:1975:QRD

[Art75] Ph. Artzner. Quelques résultats de décomposabilité en algèbre linéaire et en algèbre quadratique aléatoires. (French) []. Lecture Notes in Mathematics, 465:285–293, 1975. CODEN LN-MAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0102998/.

Andres:1977:GS

[AS77] T. H. Andres and R. G. Stanton. Golay sequences. Lecture Notes in Mathematics, 622:44–54, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069179/.

Arino:1979:EOS

[AS79] O. Arino and P. Séguier. Existence of oscillating solutions for certain differential equations with delay. Lecture Notes in Mathematics, 730:46–64, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064310/.

Askey:1975:IFI

[Ask75] Richard Askey. Inequalities via fractional integration. Lecture Notes in Mathematics, 457:106–115, 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067100/.

Assadi:1979:SEF

[Ass79] Amir H. Assadi. Some examples of finite group actions. Lecture Notes in Mathematics, 741:206–221, 1979. CODEN LN-MAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062141/.

Aastrom:1978:SCP

[Åst78] K. J. Åström. Stochastic control problems. Lecture Notes in Mathematics, 680:1–67, 1978. CODEN LNMAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065311/.

Auslander:1975:AHA

[AT75a] Prof. Louis Auslander and Prof. Richard Tolimieri. Abelian Harmonic Analysis, Theta Functions and Function Algebras on a Nilmanifold, volume 436 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07134-2 (print), 3-540-37405-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 99 pp. LCCN QA3 .L28 no. 436. URL http:

//link.springer.com/book/10.1007/BFb0069850;   http://
www.springerlink.com/content/978-3-540-37405-3.

Auslander:1975:CTF

[AT75b] Prof. Louis Auslander and Prof. Richard Tolimieri. Cohomology and theta functions. Lecture Notes in Mathematics, 436: 40–43, 1975. CODEN LNMAA2. ISBN 3-540-07134-2 (print), 3-540-37405-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069854/.

Auslander:1975:ETT

[AT75c] Prof. Louis Auslander and Prof. Richard Tolimieri. Elementary transformation theory. Lecture Notes in Mathematics, 436:30– 39, 1975. CODEN LNMAA2. ISBN 3-540-07134-2 (print), 3-540-37405-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069853/.

Auslander:1975:FTN

[AT75d] Prof. Louis Auslander and Prof. Richard Tolimieri. The Fourier transform and the nilmanifold Γ \ N. Lecture Notes in Mathematics, 436:1–14, 1975. CODEN LNMAA2. ISBN 3-540-07134-2 (print), 3-540-37405-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069851/.

Auslander:1975:FTF

[AT75e] Prof. Louis Auslander and Prof. Richard Tolimieri. Functions on Γ \ N and theta functions. Lecture Notes in Mathematics, 436:

15–29, 1975. CODEN LNMAA2. ISBN 3-540-07134-2 (print), 3-540-37405-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069852/.

Auslander:1975:TFD

[AT75f] Prof. Louis Auslander and Prof. Richard Tolimieri. Theta functions and distinguished subspaces. Lecture Notes in Mathematics, 436:44–67, 1975. CODEN LNMAA2. ISBN 3-540-07134-2 (print), 3-540-37405-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069855/.

Ansorge:1978:NTD

[AT78] Rainer Ansorge and Willi Törnig, editors. Numerical Treatment of Differential Equations in Applications: Proceedings, Oberwolfach, Germany, December 1977, volume 679 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 679. URL http:

//link.springer.com/book/10.1007/BFb0067861;   http://
www.springerlink.com/content/978-3-540-35715-5.

Atiyah:1975:HER

[Ati75] M. F. Atiyah. The heat equation in Riemannian geometry. Lecture Notes in Mathematics, 431:1–11, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066361/.

Atiyah:1977:ST

[Ati77] M. F. Atiyah. A survey of K-theory. Lecture Notes in Mathematics, 575:1–9, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095697/.

Attouch:1978:CFC

[Att78] Hédy Attouch. Convergence de fonctionnelles convexes. (French) []. Lecture Notes in Mathematics, 665:1–40, 1978. CODEN LN-MAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061795/.

Aubert:1979:PEF

[Aub79] P.-L. Aubert. Projecteurs dans U(G): un exemple. (French) []. Lecture Notes in Mathematics, 725:17–18, 1979. CODEN LN-MAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062613.

Aupetit:1979:CABa

[Aup79a] Prof. Dr. Bernard Aupetit. Caractérisation des algèbres de Banach commutatives. (French) []. Lecture Notes in Mathematics, 735:42–62, 1979. CODEN LNMAA2. ISBN 3-540-09531-4 (print), 3-540-35042-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064206/.

Aupetit:1979:CABb

[Aup79b] Prof. Dr. Bernard Aupetit. Caractérisation des algèbres de Banach de dimension finie. (French) []. Lecture Notes in Mathematics, 735:63–106, 1979. CODEN LNMAA2. ISBN 3-540-09531-4 (print), 3-540-35042-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064207/.

Aupetit:1979:CABc

[Aup79c] Prof. Dr. Bernard Aupetit. Caractérisation des algèbres de Banach symétriques. (French) []. Lecture Notes in Mathematics, 735: 107–136, 1979. CODEN LNMAA2. ISBN 3-540-09531-4 (print), 3-540-35042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064208/.

Aupetit:1979:CUC

[Aup79d] Prof. Dr. Bernard Aupetit. Continuité et uniforme continuité du spectre. (French) []. Lecture Notes in Mathematics, 735:137– 156, 1979. CODEN LNMAA2. ISBN 3-540-09531-4 (print), 3-540-35042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064209/.

Aupetit:1979:PGD

[Aup79e] Prof. Dr. Bernard Aupetit. Propriétés générales du spectre. (French) []. Lecture Notes in Mathematics, 735:1–41, 1979. CO-DEN LNMAA2. ISBN 3-540-09531-4 (print), 3-540-35042-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0064205/.

Aupetit:1979:PSA

[Aup79f] Prof. Dr. Bernard Aupetit. Propriétés Spectrales des Algèbres de Banach, volume 735 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09531-4 (print), 3-540-35042-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 192 pp. LCCN QA3 .L28 no. 735. URL http:

//link.springer.com/book/10.1007/BFb0064204;   http://
www.springerlink.com/content/978-3-540-35042-2.

Auslander:1975:ASSa

[Aus75] Maurice Auslander. Almost split sequences I. Lecture Notes in Mathematics, 488:1–8, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081212/.

Auslander:1979:DTD

[Aus79] Joseph Auslander. On disjointness in topological dynamics and ergodic theory. Lecture Notes in Mathematics, 729:28–33, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063280/.

Anderson:1979:EDM

[AV79] G. D. Anderson and M. K. Vamanamurthy. An extremal displacement mapping in n-space. Lecture Notes in Mathematics, 747:1–9, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063955/.

Avanissian:1977:DHO

[Ava77] Vazgain Avanissian. Distributions harmoniques d’ordre infini et l’analyticité (réelle) liée à l’opérateur laplacien itéré. (French) []. Lecture Notes in Mathematics, 581:1–20, 1977. CODEN LN-MAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087182/.

Avez:1976:CGT

[Ave76] André Avez. Croissance des groupes de type fini et fonctions harmoniques. (French) []. Lecture Notes in Mathematics, 532: 35–49, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080171/.

Averintsev:1978:SCG

[Ave78] M. B. Averintsev. On some classes of Gibbsian random fields. Lecture Notes in Mathematics, 653:91–98, 1978. CODEN LN-MAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070086/.

Avramov:1979:HAL

[Avr79] Luchezar L. Avramov. Homomorphismes d’anneaux locaux et homologie. (French) []. Lecture Notes in Mathematics, 740:243– 271, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071066/.

Aronson:1975:NDP

[AW75] D. G. Aronson and H. F. Weinberger. Nonlinear diffusion in population genetics, combustion, and nerve pulse propagation. Lecture Notes in Mathematics, 446:5–49, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070595/.

Acker:1976:QPN

[AW76] Andrew Acker and Wolfgang Walter. The quenching problem for nonlinear parabolic differential equations. Lecture Notes in Mathematics, 564:1–12, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087321/.

Arnold:1978:SGR

[AW78] L. Arnold and V. Wihstutz. On the stability and growth of real noise parameter-excited linear systems. Lecture Notes in Mathematics, 695:211–227, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062668/.

Axelsson:1977:SLS

[Axe77] Owe Axelsson. Solution of linear systems of equations: Iterative methods. Lecture Notes in Mathematics, 572:1–51, 1977. CODEN LNMAA2. ISBN 3-540-08130-5 (print), 3-540-37430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0116614/.

Axelsson:1979:NSN

[Axe79] O. Axelsson. On the numerical solution of nonlinear partial differential equations on divergence form. Lecture Notes in Mathematics, 703:5–15, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067250/.

Azema:1979:SSA

[AY79] Jacques Azéma and Marc Yor. Une solution simple au problème de Skorokhod. (French) []. Lecture Notes in Mathematics, 721: 90–115, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070852/.

Azumaya:1979:SAF

[Azu79] Gorô Azumaya. Some aspects of Fuller’s theorem. Lecture Notes in Mathematics, 700:34–45, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063458/.

Baba:1976:CPU

[Bab76] Yoshikazu Baba. On certain problems of uniform distribution of real sequences. Lecture Notes in Mathematics, 550:14–19, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077481/.

Babai:1978:CNS

[Bab78] László Babai. Chromatic number and subgraphs of Cayley graphs. Lecture Notes in Mathematics, 642:10–22, 1978. CODEN LN-MAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070361/.

Babuska:1979:CAH

[Bab79] I. Babuška. The computational aspects of the homogenization problem. Lecture Notes in Mathematics, 704:309–316, 1979. CO-DEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063628/.

Badrikian:1976:PAC

[Bad76] A. Badrikian. Prolegomènes au calcul des probabilités dans les Banach. (French) []. Lecture Notes in Mathematics, 539:1–166, 1976. CODEN LNMAA2. ISBN 3-540-07858-4 (print), 3-540-38119-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079696/.

Baddeley:1977:KH

[Bad77] Adrian Baddeley. The knotted hexagon. Lecture Notes in Mathematics, 622:55–60, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069180/.

Badrikian:1978:FSC

[Bad78] Albert Badrikian. Les fonctions semi-continues inférieurement et la théorie des mesures cylindriques. (French) []. Lecture Notes in Mathematics, 644:1–43, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066836/.

Baeza:1978:IQF

[Bae78a] Ricardo Baeza. Invariants of quadratic forms. Lecture Notes in Mathematics, 655:22–58, 1978. CODEN LNMAA2. ISBN 3-540-08845-8 (print), 3-540-35816-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070343/.

Baeza:1978:OG

[Bae78b] Ricardo Baeza. The orthogonal group. Lecture Notes in Mathematics, 655:59–88, 1978. CODEN LNMAA2. ISBN 3-540-08845-8 (print), 3-540-35816-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070344/.

Baeza:1978:PSS

[Bae78c] Ricardo Baeza. Pfister spaces over semi local rings. Lecture Notes in Mathematics, 655:89–111, 1978. CODEN LNMAA2. ISBN 3-540-08845-8 (print), 3-540-35816-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070345/.

Baeza:1978:QFR

[Bae78d] Ricardo Baeza. Quadratic forms over rings. Lecture Notes in Mathematics, 655:1–21, 1978. CODEN LNMAA2. ISBN 3-540-08845-8 (print), 3-540-35816-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070342/.

Baeza:1978:QFS

[Bae78e] Ricardo Baeza. Quadratic Forms over Semilocal Rings, volume 655 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08845-8 (print), 3-540-35816-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 199 pp. LCCN QA3 .L28 no. 655. URL http:

//link.springer.com/book/10.1007/BFb0070341;   http://
www.springerlink.com/content/978-3-540-35816-9.

Baeza:1978:SWR

[Bae78f] Ricardo Baeza. Structure of Witt rings. Lecture Notes in Mathematics, 655:112–176, 1978. CODEN LNMAA2. ISBN 3-540-08845-8 (print), 3-540-35816-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070346/.

Baiocchi:1976:IQV

[Bai76] Claudio Baiocchi. Inéquations quasi-variationnelles dans les problémes à frontière libre en hydraulique. (French) []. Lecture Notes in Mathematics, 503:1–7, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088743/.

Baillette:1977:AFP

[Bai77a] Aimée Baillette. Approximation de fonctions par des exponentielles imaginaires. (French) []. Lecture Notes in Mathematics, 578: 158–171, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091467/.

Baiocchi:1977:EEP

[Bai77b] C. Baiocchi. Estimations d’Erreur dans L^{∞} pour les inéquations à obstacle. (French) [Error estimates in L^{∞} for obstacle inequations]. Lecture Notes in Mathematics, 606:27–34, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064453/.

Baker:1976:ACA

[Bak76] Charles R. Baker. Absolute continuity and applications to information theory. Lecture Notes in Mathematics, 526:1–11, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082337/.

Baker:1978:ACC

[Bak78a] Charles R. Baker. Absolute continuity for a class of probability measures. Lecture Notes in Mathematics, 644:44–50, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066837/.

Baker:1978:RKM

[Bak78b] Christopher T. H. Baker. Runge-Kutta methods for Volterra integral equations of the second kind. Lecture Notes in Mathematics, 630:1–13, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067691/.

Baker:1979:CAC

[Bak79a] Charles R. Baker. Conditions for absolute continuity. Lecture Notes in Mathematics, 709:33–44, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071946/.

Baker:1979:RRF

[Bak79b] R. C. Baker. Recent results on fractional parts of polynomials. Lecture Notes in Mathematics, 751:10–17, 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062699/.

Ball:1976:CVS

[Bal76] J. M. Ball. On the calculus of variations and sequentially weakly continuous maps. Lecture Notes in Mathematics, 564:13–25, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087322/.

Balakrishnan:1978:LRG

[Bal78] A. V. Balakrishnan. Likelihood ratios with Gauss measure noise models. Lecture Notes in Mathematics, 695:89–100, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062657/.

Baldi:1979:LSD

[Bal79] P. Baldi. Lois stables sur les déplacements de R^{d}. (French) []. Lecture Notes in Mathematics, 706:1–9, 1979. CODEN LN-MAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063109/.

Bamazi:1979:RSU

[Bam79] Lucien Bamazi. Représentations sphériques uniformement bornées des groupes de Lorentz. (French) []. Lecture Notes in Mathematics, 739:1–14, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062486/.

Banyaga:1975:GDS

[Ban75] A. Banyaga. Sur le groupe des diffeomorphismes symplectiques. (French) []. Lecture Notes in Mathematics, 484:50–56, 1975. CO-DEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082145/.

Banica:1978:HDG

[Băn78] Constantin Bănică. Homologie et Dualité en Géométrie analytique complexe. (French) []. Lecture Notes in Mathematics, 670: 1–21, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064391/.

Banaschewski:1979:IBS

[Ban79a] B. Banaschewski. Injective Banach sheaves. Lecture Notes in Mathematics, 753:101–112, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061814/.

Banaschewski:1979:RSB

[Ban79b] B. Banaschewski. Recovering a space from its Banach sheaves. Lecture Notes in Mathematics, 719:1–12, 1979. CODEN LN-MAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065252/.

Banks:1979:ADS

[Ban79c] H. T. Banks. Approximation of delay systems with applications to control and identification. Lecture Notes in Mathematics, 730: 65–76, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064311/.

Barlet:1975:EAR

[Bar75] Daniel Barlet. Espace analytique reduit des cycles analytiques complexes compacts d’un espace analytique complexe de dimension finie. (French) []. Lecture Notes in Mathematics, 482:1– 158, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080243/.

Baranger:1977:EEI

[Bar77a] J. Baranger. Estimations d’erreur à l’intérieur pour un problème de couche limite. (French) []. Lecture Notes in Mathematics, 594: 1–9, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086077/.

Barker:1977:SMT

[Bar77b] V. A. Barker, editor. Sparse Matrix Techniques: Copenhagen 1976 Advanced Course Held at the Technical University of Denmark Copenhagen, August 9–12, 1976, volume 572 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LN-MAA2. ISBN 3-540-08130-5 (print), 3-540-37430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 572. URL http://link.springer.com/book/10.1007/
BFb0116611;   http://www.springerlink.com/content/978-
3-540-37430-5.

Bardos:1978:EEB

[Bar78a] Claude Bardos. Euler equation and Burger equation — relation with turbulence. Lecture Notes in Mathematics, 648:1–46, 1978. CODEN LNMAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066405/.

Barlet:1978:FEA

[Bar78b] Daniel Barlet. Le faisceau ω′_{X} sur un espace analytique X de dimension pure. (French) []. Lecture Notes in Mathematics, 670: 187–204, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064400/.

Barou:1978:CLA

[Bar78c] Geneviève Barou. Cohomologie locale des algèbres enveloppantes d’algèbres de Lie nilpotentes. (French) []. Lecture Notes in Mathematics, 641:252–280, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064852/.

Barratt:1978:TLA

[Bar78d] M. G. Barratt. Twisted Lie algebras. Lecture Notes in Mathematics, 658:9–15, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068708/.

Barrodale:1978:BAC

[Bar78e] Ian Barrodale. Best approximation of complex-valued data. Lecture Notes in Mathematics, 630:14–22, 1978. CODEN LN-MAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067692/.

Barut:1978:EFM

[Bar78f] A. O. Barut. Electromagnetic fields on manifolds: Betti numbers, monopoles and strings, minimal coupling. Lecture Notes in Mathematics, 676:179–187, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063671/.

Barbasch:1979:FTS

[Bar79a] Dan Barbasch. Fourier transforms of some invariant distribution on semisimple Lie groups and Lie algebras. Lecture Notes in Mathematics, 728:1–7, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063334/.

Barbu:1979:DNV

[Bar79b] V. Barbu. Degenerate nonlinear Volterra integral equations in Hilbert space. Lecture Notes in Mathematics, 737:9–23, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064491/.

Barr:1979:ACA

[Bar79c] Michael Barr. *-Autonomous Categories: With an Appendix by Po-Hsiang Chu, volume 752 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09563-2 (print), 3-540-34850-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 140 pp. LCCN QA3

.L28 no. 752. URL http://link.springer.com/book/10.1007/
BFb0064579;   http://www.springerlink.com/content/978-
3-540-34850-4.

Barr:1979:CG

[Bar79d] Michael Barr. The category g. Lecture Notes in Mathematics, 752:33–42, 1979. CODEN LNMAA2. ISBN 3-540-09563-2 (print), 3-540-34850-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064582/.

Barr:1979:E

[Bar79e] Michael Barr. Examples. Lecture Notes in Mathematics, 752: 43–100, 1979. CODEN LNMAA2. ISBN 3-540-09563-2 (print), 3-540-34850-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064583/.

Barr:1979:ES

[Bar79f] Michael Barr. Extensions of structure. Lecture Notes in Mathematics, 752:19–32, 1979. CODEN LNMAA2. ISBN 3-540-09563-2 (print), 3-540-34850-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064581/.

Barr:1979:P

[Bar79g] Michael Barr. Preliminaries. Lecture Notes in Mathematics, 752: 1–18, 1979. CODEN LNMAA2. ISBN 3-540-09563-2 (print), 3-540-34850-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064580/.

Barrow:1979:TDA

[Bar79h] David L. Barrow. The topological degree applied to some problems in approximation theory. Lecture Notes in Mathematics, 701: 70–79, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062074/.

Bass:1975:LMP

[Bas75] Hyman Bass. Libération des modules projectifs sur certains anneaux de polynômes. (French) []. Lecture Notes in Mathematics, 431:228–254, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066373/.

Batterson:1978:PPL

[Bat78] Steve Batterson. Periodic points and Lefschetz numbers. Lecture Notes in Mathematics, 668:12–20, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101776/.

Bauer:1976:DPF

[Bau76] Karl Wilhelm Bauer. Zur Darstellung pseudoanalytischer Funktionen. (German) []. Lecture Notes in Mathematics, 561:101– 111, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087629/.

Baues:1977:CN

[Bau77a] Hans J. Baues. Conventions and notation. Lecture Notes in Mathematics, 628:1–34, 1977. CODEN LNMAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065145/.

Baues:1977:IPC

[Bau77b] Hans J. Baues. Iterated principal cofibrations. Lecture Notes in Mathematics, 628:171–241, 1977. CODEN LNMAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065148/.

Baues:1977:IPF

[Bau77c] Hans J. Baues. Iterated principal fibrations. Lecture Notes in Mathematics, 628:324–367, 1977. CODEN LNMAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065151/.

Baues:1977:OTC

[Bau77d] Hans J. Baues. Obstruction theory for CW-complexes. Lecture Notes in Mathematics, 628:242–292, 1977. CODEN LN-MAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065149/.

Baues:1977:OTH

[Bau77e] Hans J. Baues. Obstruction Theory: on Homotopy Classification of Maps, volume 628 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 387 pp. LCCN QA3 .L28 no. 628. URL http:

//link.springer.com/book/10.1007/BFb0065144;   http://
www.springerlink.com/content/978-3-540-35979-1.

Baues:1977:PHD

[Bau77f] Hans J. Baues. Postnikov homotopy decomposition. Lecture Notes in Mathematics, 628:293–323, 1977. CODEN LNMAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065150/.

Baues:1977:PFC

[Bau77g] Hans J. Baues. Principal fibrations and cofibrations, obstructions and differences. Lecture Notes in Mathematics, 628:35–120, 1977. CODEN LNMAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065146/.

Baues:1977:RPC

[Bau77h] Hans J. Baues. Relative principal cofibrations and fibrations. Lecture Notes in Mathematics, 628:121–170, 1977. CODEN LN-MAA2. ISBN 3-540-08534-3 (print), 3-540-35979-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065147/.

Brace:1975:ARR

[BB75] Alan Brace and J. Brett. An alternative to the round robin tournament. Lecture Notes in Mathematics, 452:62–78, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069544/.

Barnhill:1976:NFE

[BB76] Robert E. Barnhill and James H. Brown. Nonconforming finite elements for curved regions. Lecture Notes in Mathematics, 506: 1–14, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080110/.

Bona:1976:SWS

[BBB76] J. L. Bona, D. K. Bose, and T. B. Benjamin. Solitary-wave solutions for some model equations for waves in nonlinear dispersive media. Lecture Notes in Mathematics, 503:207–218, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088757/.

Borisyuk:1978:EIC

[BBDB^{+}78] G. N. Borisyuk, R. M. Borisyuk, V. L. Dunin-Barkovsky, V. H. Kovalenko, and E. I. Kovalenko. Estimation of information capacity of Purkinje cells. Lecture Notes in Mathematics, 653:72–90, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070085/.

Benabdallah:1977:GQN

[BBL77] K. Benabdallah, R. Bradley, and A. Laroche. Sur les groupes quasi-p-nets injectifs et projectifs. (French) []. Lecture Notes in Mathematics, 616:228–239, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068199/.

Bohmer:1977:BIS

[BC77a] K. Böhmer and Gh. Coman. Blending interpolation schemes on triangles with error bounds. Lecture Notes in Mathematics, 571: 14–37, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086562/.

Bryce:1977:SCF

[BC77b] R. A. Bryce and John Cossey. Strong containment of fitting classes. Lecture Notes in Mathematics, 573:6–16, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087808/.

Bendersky:1978:CSM

[BC78a] M. Bendersky and E. B. Curtis. Cobordism of sequences of manifolds. Lecture Notes in Mathematics, 658:16–26, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068709/.

Benedict:1978:GHP

[BC78b] James M. Benedict and Phyllis Zweig Chinn. On graphs having prescribed clique number, chromatic number, and maximum degree. Lecture Notes in Mathematics, 642:132–140, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070370/.

Baker:1977:BSA

[BCD77] J. Baker, C. Cleaver, and J. Diestel, editors. Banach Spaces of Analytic Functions: Proceedings of the Pelczynski Conference Held at Kent State University, July 12–16, 1976, volume 604 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 604. URL http://link.springer.com/book/10.1007/
BFb0069197;   http://www.springerlink.com/content/978-
3-540-37262-2.

Bolley:1978:HPE

[BCH78] Pierre Bolley, Jacques Camus, and Bernard Helffer. Hypoellipticité pour une équation d’évolution abstraite du second ordre. (French) []. Lecture Notes in Mathematics, 660:16–32, 1978. CO-DEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082291/.

Bryce:1977:GTP

[BCN77] R. A. Bryce, J. Cossey, and M. F. Newman, editors. Group Theory: Proceedings of a Miniconference Held at the Australian National University, Canberra, November 4–6, 1975, volume 573 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 573. URL http://link.springer.com/book/10.1007/
BFb0087806;   http://www.springerlink.com/content/978-
3-540-37428-2.

Bose:1978:CMV

[BD78] A. Bose and D. Dawson. A class of measure-valued Markov processes. Lecture Notes in Mathematics, 695:115–125, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062659/.

Bellow:1979:CAS

[BD79] A. Bellow and A. Dvoretzky. A characterization of almost sure convergence. Lecture Notes in Mathematics, 709:45–65, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071947/.

Behrends:1977:P

[BDE^{+}77a] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. L^{p}- projections. Lecture Notes in Mathematics, 613:5–15, 1977. CO-DEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0068177/.

Behrends:1977:SRB

[BDE^{+}77b] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. L^{p}- Structure in Real Banach Spaces, volume 613 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 93 (est.) pp. LCCN QA3

.L28 no. 613. URL http://link.springer.com/book/10.1007/
BFb0068175;   http://www.springerlink.com/content/978-
3-540-37086-4.

Behrends:1977:SBS

[BDE^{+}77c] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. The L^{p}-structure of the Bochner spaces and related results. Lecture Notes in Mathematics, 613:83–93, 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068183/.

Behrends:1977:CS

[BDE^{+}77d] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. The classical L^{p}-spaces. Lecture Notes in Mathematics, 613:47– 59, 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068180/.

Behrends:1977:CA

[BDE^{+}77e] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. The Cunningham p-algebra. Lecture Notes in Mathematics, 613:16– 22, 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068178/.

Behrends:1977:IMR

[BDE^{+}77f] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. The integral module representation. Lecture Notes in Mathematics, 613:23–46, 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068179/.

Behrends:1977:IMD

[BDE^{+}77g] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. Integral modules and duality. Lecture Notes in Mathematics, 613:

60–72, 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068181/.

Behrends:1977:PT

[BDE^{+}77h] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. Preliminaries topology. Lecture Notes in Mathematics, 613:1–4, 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068176/.

Behrends:1977:STO

[BDE^{+}77i] Ehrhard Behrends, Rainer Danckwerts, Richard Evans, Silke Göbel, Peter Greim, Konrad Meyfarth, and Winfried Müller. Spectral theory for L^{p}-operators. Lecture Notes in Mathematics, 613:73–82, 1977. CODEN LNMAA2. ISBN 3-540-08441-X (print), 3-540-37086-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068182/.

Berezin:1977:ANS

[BDNY77] Yu. A. Berezin, G. I. Dudnikova, V. A. Novikov, and N. N. Yanenko. Analytical and numerical studies of equations with sign changing viscosity coefficient. Lecture Notes in Mathematics, 594: 30–38, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086079/.

Bellaiche:1976:ODS

[BDPR76] C. Bellaïche, A. Decarreau, D. Pesenti, and A. Robert. Opérations Γ et dΓ de Segal. (French) []. Lecture Notes in Mathematics, 518: 14–36, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080400/.

Balacheff:1978:ARC

[BDR78] Serge Balacheff, Ghislain Dupont, and Jean-Pierre Raoult. Application de résultats de convergence de processus multidimensionnels à l’étude des statistiques de rang. (French) []. Lecture Notes in Mathematics, 636:189–202, 1978. CODEN LNMAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063265/.

Borho:1978:SBV

[BDV^{+}78] Walter Borho, J. J. Duistermaat, A. N. Varchenko, William James Harvey, Michel Raynaud, Jacques Stern, Jean-Michel Bony, Pierre Cartier, Robert MacPherson, Mustapha Raïs, Robert Roussarie, A. Van De Ven, Francis Hirsch, Zvonimir Janko, Serge Lang, Paul Malliavin, and Jacques Tits, editors. Séminaire Bourbaki vol. 1976/77 Exposés 489–506, volume 677 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 677. URL http:

//link.springer.com/book/10.1007/BFb0070748;   http://
www.springerlink.com/content/978-3-540-35719-3.

Barth:1978:CCF

[BE78] W. Barth and G. Elencwajg. Concernant la cohomologie des fibres algébriques stables sur P_{n}(C). (French) []. Lecture Notes in Mathematics, 683:1–24, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063170/.

Beachy:1976:SAN

[Bea76] John A. Beachy. Some aspects of noncomutative localization. Lecture Notes in Mathematics, 545:2–31, 1976. CODEN LN-MAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080304/.

Beauzamy:1978:DEI

[Bea78a] Bernard Beauzamy. Definition des espaces d’Interpolation. (French) []. Lecture Notes in Mathematics, 666:1–23, 1978. CO-DEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0068828/.

Beauzamy:1978:DRD

[Bea78b] Bernard Beauzamy. Dualité et réiteration du procédé d’Interpolation. (French) []. Lecture Notes in Mathematics, 666:59–69, 1978. CO-DEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068831/.

Beauzamy:1978:EIR

[Bea78c] Bernard Beauzamy. Espaces d’Interpolation Réels: Topologie et Géométrie. (French) [], volume 666 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 104 pp. LCCN QA3

.L28 no. 666. URL http://link.springer.com/book/10.1007/
BFb0068827;   http://www.springerlink.com/content/978-
3-540-35754-4.

Beauzamy:1978:MRP

[Bea78d] Bernard Beauzamy. Meilleures représentations et propriétés métriques non uniformes. (French) []. Lecture Notes in Mathematics, 666:82–87, 1978. CODEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068833/.

Beauzamy:1978:MGP

[Bea78e] Bernard Beauzamy. Méthodes géométriques pour l’Étude des espaces d’Interpolation. (French) []. Lecture Notes in Mathematics, 666:38–58, 1978. CODEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068830/.

Beauzamy:1978:PTE

[Bea78f] Bernard Beauzamy. Propriétés topologiques des espaces d’Interpolation. (French) []. Lecture Notes in Mathematics, 666:24–37, 1978. CO-DEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068829/.

Beauzamy:1978:QDD

[Bea78g] Bernard Beauzamy. Quelques ‘défauts’ du procédé d’Interpolation. (French) []. Lecture Notes in Mathematics, 666:88–99, 1978. CO-DEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0068834/.

Beauzamy:1978:SPP

[Bea78h] Bernard Beauzamy. Super-propriétés et propriétés métriques des espaces d’Interpolation. (French) []. Lecture Notes in Mathematics, 666:70–81, 1978. CODEN LNMAA2. ISBN 3-540-08923-3 (print), 3-540-35754-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068832/.

Beachy:1979:FLB

[Bea79a] John Beachy. Fully left bounded left Noetherian rings. Lecture Notes in Mathematics, 700:217, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063468.

Beachy:1979:IL

[Bea79b] John A. Beachy. On inversive localization. Lecture Notes in Mathematics, 700:46–56, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063459/.

Bebernes:1979:SSP

[Beb79] J. W. Bebernes. Solution set properties for some nonlinear parabolic differential equations. Lecture Notes in Mathematics, 703:25–30, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067252/.

Beck:1976:CBS

[Bec76a] Anatole Beck. Cancellation in Banach spaces. Lecture Notes in Mathematics, 526:13–20, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082338/.

Beck:1976:PBS

[Bec76b] Anatole Beck, editor. Probability in Banach Spaces: Proceedings of the First International Conference on Probability in Banach Spaces, 20–26 July 1975, Oberwolfach, volume 526 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 0-387-07793-6, 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 526. URL http:

//link.springer.com/book/10.1007/BFb0082336;   http://
www.springerlink.com/content/978-3-540-38256-0.

Beck:1978:CT

[Bec78a] Anatole Beck. On the covariance tensor. Lecture Notes in Mathematics, 644:51–56, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066838/.

Becker:1978:PPU

[Bec78b] Howard Becker. Partially playful universes. Lecture Notes in Mathematics, 689:55–90, 1978. CODEN LNMAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069295/.

Beck:1979:PBS

[Bec79a] Anatole Beck, editor. Probability in Banach Spaces II: Proceedings of the Second International Conference on Probability in Banach Spaces, 18–24 June 1978, Oberwolfach, Germany, volume 709 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 0-387-09242-0, 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 709. URL http:

//link.springer.com/book/10.1007/BFb0071943;   http://
www.springerlink.com/content/978-3-540-35341-6.

Beck:1979:SSF

[Bec79b] Jon M. Beck. Simplicial sets and the foundations of analysis. Lecture Notes in Mathematics, 753:113–124, 1979. CODEN LN-MAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061815/.

Bros:1975:QAG

[BEGS75] J. Bros, H. Epstein, V. Glaser, and R. Stora. Quelques aspects globaux des problémes d’edge-of-the-wedge. (French) []. Lecture Notes in Mathematics, 449:185–218, 1975. CODEN LN-MAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062923/.

Behrends:1979:BSP

[Beh79a] Ehrhard Behrends. The Banach–Stone property and the strong Banach–Stone property. Lecture Notes in Mathematics, 736:141– 151, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063162/.

Behrends:1979:BST

[Beh79b] Ehrhard Behrends. The Banach–Stone theorem. Lecture Notes in Mathematics, 736:135–140, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063161/.

Behrends:1979:C

[Beh79c] Ehrhard Behrends. The centralizer. Lecture Notes in Mathematics, 736:53–74, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063157/.

Behrends:1979:CNS

[Beh79d] Ehrhard Behrends. Centralizer-norming systems. Lecture Notes in Mathematics, 736:152–166, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063163/.

Behrends:1979:FM

[Beh79e] Ehrhard Behrends. Function modules. Lecture Notes in Mathematics, 736:75–107, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063158/.

Behrends:1979:GBS

[Beh79f] Ehrhard Behrends. Generalizations of the Banach–Stone theorem. Lecture Notes in Mathematics, 736:178–200, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063165/.

Behrends:1979:PP

[Beh79g] Ehrhard Behrends. L-projections and M-projections. Lecture Notes in Mathematics, 736:7–32, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063155/.

Behrends:1979:I

[Beh79h] Ehrhard Behrends. M-ideals. Lecture Notes in Mathematics, 736: 33–52, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063156/.

Behrends:1979:SBS

[Beh79i] Ehrhard Behrends. M-Structure and the Banach–Stone Theorem, volume 736 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 217 pp. LCCN QA3 .L28 no. 736. URL http:

//link.springer.com/book/10.1007/BFb0063153;   http://
www.springerlink.com/content/978-3-540-35038-5.

Behrends:1979:S

[Beh79j] Ehrhard Behrends. M-structure of C_{o}(M, X). Lecture Notes in Mathematics, 736:167–177, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063164/.

Behrends:1979:SSC

[Beh79k] Ehrhard Behrends. M-structure of some classes of Banach spaces. Lecture Notes in Mathematics, 736:108–121, 1979. CODEN LN-MAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063159/.

Behrends:1979:P

[Beh79l] Ehrhard Behrends. Preliminaries. Lecture Notes in Mathematics, 736:1–6, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063154/.

Behrends:1979:Ra

[Beh79m] Ehrhard Behrends. Remarks. Lecture Notes in Mathematics, 736: 122–134, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063160/.

Behrends:1979:Rb

[Beh79n] Ehrhard Behrends. Remarks. Lecture Notes in Mathematics, 736: 201–207, 1979. CODEN LNMAA2. ISBN 3-540-09533-0 (print), 3-540-35038-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063166/.

Bell:1976:USM

[Bel76] J. L. Bell. Uncountable standard models of ZFC + V ≠ L. Lecture Notes in Mathematics, 537:29–36, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096891/.

Bellow:1978:SAT

[Bel78] Alexandra Bellow. Some aspects of the theory of vector-valued amarts. Lecture Notes in Mathematics, 644:57–67, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066839/.

Benson:1975:SPK

[Ben75a] Fred Benson. A short proof of a Kirby flattening theorem. Lecture Notes in Mathematics, 438:6–8, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066103/.

Benveniste:1975:PSM

[Ben75b] A. Benveniste. Processus stationaires et mesures de Palm du flot special sous une fonction. (French) []. Lecture Notes in Mathematics, 465:97–153, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102989/.

Benjamin:1976:APA

[Ben76a] T. Brooke Benjamin. The alliance of practical and analytical insights into the nonlinear problems of fluid mechanics. Lecture Notes in Mathematics, 503:8–29, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088744/.

Bentley:1976:RNS

[Ben76b] H. L. Bentley. The role of nearness spaces in topology. Lecture Notes in Mathematics, 540:1–22, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080851/.

Benzinger:1976:SAS

[Ben76c] Harold E. Benzinger. The spectral analysis of a stone regular differential operator. Lecture Notes in Mathematics, 564:34–40, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087324/.

Bennett:1977:ERT

[Ben77a] G. Bennett. An extension of the Riesz–Thorin theorem. Lecture Notes in Mathematics, 604:1–11, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069198/.

Benveniste:1977:ATG

[Ben77b] A. Benveniste. Application d’un théorème de G. Mokobozki à la théorie des flots. (French) []. Lecture Notes in Mathematics, 581:21–26, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087183/.

Benes:1978:RWS

[Ben78] V. E. Beneš. Realizing a weak solution on a probability space. Lecture Notes in Mathematics, 695:101–113, 1978. CODEN LN-MAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062658/.

Berenstein:1976:ENZ

[Ber76a] Carlos A. Berenstein. An estimate for the number of zeroes of analytic functions in n-dimensional cones. Lecture Notes in Mathematics, 505:1–16, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081096/.

Berg:1976:SGC

[Ber76b] Christian Berg. Semi-groupes de convolution sur les groupes non moyennables. (French) []. Lecture Notes in Mathematics, 518:45– 53, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080402/.

Berg:1976:TFM

[Ber76c] Christian Berg. Transformation de Fourier de mesures de type positif sur un groupe abelien localement compact. (French) []. Lecture Notes in Mathematics, 518:37–44, 1976. CODEN LN-MAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080401/.

Bergman:1976:ZDT

[Ber76d] George M. Bergman. Zero-divisors in tensor products. Lecture Notes in Mathematics, 545:32–82, 1976. CODEN LNMAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080305/.

Bergman:1976:MTF

[Ber76e] Stefan Bergman. On the mathematical theory of flow patterns of compressible fluids. Lecture Notes in Mathematics, 561:1–9, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087620/.

Berner:1976:GFP

[Ber76f] Paul Berner. A global factorization property for holomorphic functions of a domain spread over a surjective limit. Lecture Notes in Mathematics, 524:130–155, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078003/.

Berquier:1976:TIL

[Ber76g] Françoise Berquier. Un théorème d’inversion locale. (French) [A local inversion theorem]. Lecture Notes in Mathematics, 540:24– 42, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080852/.

Berger:1977:FPF

[Ber77a] T. R. Berger. Fixed point free automorphism groups. Lecture Notes in Mathematics, 573:1–5, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087807/.

Bernard:1977:CEP

[Ber77b] Alain Bernard. Complément à l’exposé précédent. (French) []. Lecture Notes in Mathematics, 581:324–326, 1977. CODEN LN-MAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087198/.

Berge:1978:MGH

[Ber78a] Claude Berge. The multicolorings of graphs and hypergraphs. Lecture Notes in Mathematics, 642:23–36, 1978. CODEN LN-MAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070362/.

Bernhart:1978:ICF

[Ber78b] Frank R. Bernhart. Irreducible configurations and the four color conjecture. Lecture Notes in Mathematics, 642:37–52, 1978. CO-DEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070363/.

Berg:1979:HCK

[Ber79a] Christian Berg. Hunt convolution kernels which are continuous singular with respect to Haar measure. Lecture Notes in Mathematics, 706:10–21, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063110/.

Berg:1979:SCL

[Ber79b] Christian Berg. The Stieltjes cone is logarithmically convex. Lecture Notes in Mathematics, 747:46–54, 1979. CODEN LN-MAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063958/.

Bers:1979:AQM

[Ber79c] Lipman Bers. An application of quasiconformal mappings to topology. Lecture Notes in Mathematics, 743:57–65, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079483/.

Berthier:1979:RGU

[Ber79d] Anne-Marie Berthier. A remark on a generalized uncertainty principle. Lecture Notes in Mathematics, 755:17–34, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069802/.

Bertin:1979:PAM

[Ber79e] Jean-Étienne Bertin. Prolongements de Q-anneaux de Matlis. (French) []. Lecture Notes in Mathematics, 740:190–197, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071060/.

Bescheron:1975:MIF

[Bes75a] Dominique Bescheron. Multiplicité d’intersection et formules intégrales. (French) []. Lecture Notes in Mathematics, 482:168– 179, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080245/.

Besineau:1975:EEC

[Bes75b] J. Besineau. Ensembles d’entiers à caractères presque périodiques et équirépartition. (French) []. Lecture Notes in Mathematics, 475: 1–12, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074254/.

Bettis:1978:EER

[Bet78] D. G. Bettis. Efficient embedded Runge–Kutta methods. Lecture Notes in Mathematics, 631:9–18, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067460/.

Bair:1975:AFA

[BF75a] Dr. Jacques Bair and Dr. René Fourneau. Applications. (French) [Applications]. Lecture Notes in Mathematics, 489:75–137, 1975. CODEN LNMAA2. ISBN 3-540-07413-9 (print), 3-540-37938-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080459/.

Bair:1975:CNE

[BF75b] Dr. Jacques Bair and Dr. René Fourneau. Considerations sur la nature de l’espace. (French) []. Lecture Notes in Mathematics, 489:138–157, 1975. CODEN LNMAA2. ISBN 3-540-07413-9 (print), 3-540-37938-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080460/.

Bair:1975:EGE

[BF75c] Dr. Jacques Bair and Dr. René Fourneau. Étude Géométrique des Espaces Vectoriels: Une Introduction, volume 489 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07413-9 (print), 3-540-37938-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 185 pp. LCCN QA3

.L28 no. 489. URL http://link.springer.com/book/10.1007/
BFb0080456;   http://www.springerlink.com/content/978-
3-540-37938-6.

Bair:1975:ITO

[BF75d] Dr. Jacques Bair and Dr. René Fourneau. Iteres transfinis des opérateurs algébriques. (French) []. Lecture Notes in Mathematics, 489:62–74, 1975. CODEN LNMAA2. ISBN 3-540-07413-9 (print), 3-540-37938-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080458/.

Bair:1975:OAF

[BF75e] Dr. Jacques Bair and Dr. René Fourneau. Opérateurs algébriques fondamentaux. (French) []. Lecture Notes in Mathematics, 489: 1–61, 1975. CODEN LNMAA2. ISBN 3-540-07413-9 (print), 3-540-37938-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080457/.

Bair:1975:PTF

[BF75f] Dr. Jacques Bair and Dr. René Fourneau. Propriétés topologiques. (French) []. Lecture Notes in Mathematics, 489:158–170, 1975. CODEN LNMAA2. ISBN 3-540-07413-9 (print), 3-540-37938-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080461/.

Bardos:1976:FTR

[BF76a] C. Bardos and U. Frisch. Finite-time regularity for bounded and unbounded ideal incompressible fluids using holder estimates. Lecture Notes in Mathematics, 565:1–13, 1976. CODEN LN-MAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091443/.

Berger:1976:GFB

[BF76b] M. S. Berger and L. E. Fraenkel. Global free boundary problems and the calculus of variations in the large. Lecture Notes in Mathematics, 503:186–192, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088754/.

Bousfield:1978:HTS

[BF78] A. K. Bousfield and E. M. Friedlander. Homotopy theory of Γspaces, spectra, and bisimplicial sets. Lecture Notes in Mathematics, 658:80–130, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068711/.

Berg:1979:IDP

[BF79a] Christian Berg and Gunnar Forst. Infinitely divisible probability measures and potential kernels. Lecture Notes in Mathematics, 706:22–35, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063111/.

Boyle:1979:SMP

[BF79b] A. K. Boyle and E. H. Feller. Semicritical modules and k-primitive rings. Lecture Notes in Mathematics, 700:57–74, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063460/.

Balcar:1976:RPS

[BG76a] B. Balcar and W. Guzicki. A remark on powers of singular cardinals. Lecture Notes in Mathematics, 537:25–28, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096890/.

Begehr:1976:RNN

[BG76b] Heinrich Begehr and Robert P. Gilbert. Über das Randwert-Normproblem für ein nichtlineares elliptisches System. (German) []. Lecture Notes in Mathematics, 561:112–122, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087630/.

Baouendi:1978:TNP

[BG78a] M. S. Baouendi and C. Goulaouic. Le théorème de Nishida pour le problème de Cauchy abstrait par une méthode de point fixe. (French) []. Lecture Notes in Mathematics, 660:1–8, 1978. CO-DEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082289/.

Bloom:1978:NCG

[BG78b] Gary S. Bloom and Solomon W. Golomb. Numbered complete graphs, unusual rulers, and assorted applications. Lecture Notes in Mathematics, 642:53–65, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070364/.

Boley:1978:IEP

[BG78c] D. Boley and G. H. Golub. Inverse eigenvalue problems for band matrices. Lecture Notes in Mathematics, 630:23–31, 1978. CO-DEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0067693/.

Bucchioni:1978:CPP

[BG78d] D. Bucchioni and A. Goldman. Convergence presque partout des suites de fonctions mesurables et applications. (French) []. Lecture Notes in Mathematics, 645:1–7, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069657/.

Bouc:1976:HUP

[BGJN76] R. Bouc, G. Geymonat, M. Jean, and B. Nayroles. Hilbertian unilateral problems in viscoelasticity. Lecture Notes in Mathematics, 503:219–234, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088758/.

Brauner:1977:SPB

[BGM77] Claude-Michel Brauner, Bernard Gay, and Jean Mathieu, editors. Singular Perturbations and Boundary Layer Theory: Proceedings of the Conference held at the École Centrale de Lyon, December 8–10, 1976, volume 594 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 594. URL http:

//link.springer.com/book/10.1007/BFb0086076;   http://
www.springerlink.com/content/978-3-540-37340-7.

Bulirsch:1978:NTD

[BGS78] R. Bulirsch, R. D. Grigorieff, and J. Schröder, editors. Numerical Treatment of Differential Equations: Proceedings of a Conference, Held at Oberwolfach, July 4–10, 1976, volume 631 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 631. URL http://link.springer.com/book/10.1007/
BFb0067458;   http://www.springerlink.com/content/978-
3-540-35970-8.

Ball:1976:FAA

[BH76a] D. J. Ball and J. R. Hewit. Functional analysis applied to the optimisation of a temperature profile. Lecture Notes in Mathematics, 503:177–185, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088753/.

Binz:1976:CTP

[BH76b] Ernst Binz and Horst Herrlich, editors. Categorical Topology: Proceedings of the Conference Held at Mannheim, 21–25, July, 1975, volume 540 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 540. URL http:

//link.springer.com/book/10.1007/BFb0080850;   http://
www.springerlink.com/content/978-3-540-38118-1.

Brummer:1976:ECT

[BH76c] G. C. L. Brümmer and R.-E. Hoffmann. An external characterization of topological functors. Lecture Notes in Mathematics, 540:136–151, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080856/.

Bachar:1978:HSO

[BH78] John M. Bachar, Jr. and Donald W. Hadwin, editors. Hilbert Space Operators: Proceedings, California State University Long Beach Long Beach, California, 20–24 June, 1977, volume 693 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 693. URL http://link.springer.com/book/10.1007/
BFb0064656;   http://www.springerlink.com/content/978-
3-540-35557-1.

Bentley:1979:CP

[BH79] H. L. Bentley and H. Herrlich. Completeness is productive. Lecture Notes in Mathematics, 719:13–17, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065253/.

Booth:1978:CUF

[BHP78a] Peter I. Booth, Philip R. Heath, and Renzo A. Piccinini. Characterizing universal fibrations. Lecture Notes in Mathematics, 673: 168–184, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064695/.

Booth:1978:FPM

[BHP78b] Peter I. Booth, Philip R. Heath, and Renzo A. Piccinini. Fibre preserving maps and functional spaces. Lecture Notes in Mathematics, 673:158–167, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064694/.

Bibikov:1979:AFS

[Bib79a] Yuri N. Bibikov. Analytic families of solutions. Lecture Notes in Mathematics, 702:3–47, 1979. CODEN LNMAA2. ISBN 3-540-09114-9 (print), 3-540-35527-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064651/.

Bibikov:1979:I

[Bib79b] Yuri N. Bibikov. Introduction. Lecture Notes in Mathematics, 702:1–2, 1979. CODEN LNMAA2. ISBN 3-540-09114-9 (print), 3-540-35527-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064650.

Bibikov:1979:LTN

[Bib79c] Yuri N. Bibikov. Local Theory of Nonlinear Analytic Ordinary Differential Equations, volume 702 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09114-9 (print), 3-540-35527-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 135 (est.) pp. LCCN QA3

.L28 no. 702. URL http://link.springer.com/book/10.1007/
BFb0064649;   http://www.springerlink.com/content/978-
3-540-35527-4.

Bibikov:1979:QPS

[Bib79d] Yuri N. Bibikov. Quasi-periodic solutions. Lecture Notes in Mathematics, 702:91–135, 1979. CODEN LNMAA2. ISBN 3-540-09114-9 (print), 3-540-35527-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064653/.

Bibikov:1979:S

[Bib79e] Yuri N. Bibikov. Stability. Lecture Notes in Mathematics, 702: 48–90, 1979. CODEN LNMAA2. ISBN 3-540-09114-9 (print), 3-540-35527-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064652/.

Bielinski:1977:ESI

[Bie77] Konrad Bielinski. Extendability of structures as infinitary property. Lecture Notes in Mathematics, 619:75–93, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067643/.

Bieri:1979:FPS

[Bie79] Robert Bieri. Finitely presented soluble groups. Lecture Notes in Mathematics, 740:1–8, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071049/.

Bihari:1979:AIS

[Bih79] I. Bihari. Asymptotic invariant sets of autonomous differential equations. Lecture Notes in Mathematics, 703:31–34, 1979. CO-DEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0067253/.

Binder:1975:RCF

[Bin75a] Christa Binder. Une remarque sur la caractérisation des fonctions R − μ-intégrables. (French) []. Lecture Notes in Mathematics, 475:13–17, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074255/.

Bing:1975:VGP

[Bin75b] R. H. Bing. Vertical general position. Lecture Notes in Mathematics, 438:16–41, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066105/.

Binz:1975:ES

[Bin75c] Prof. Dr. Ernst Binz. c-Embedded spaces. Lecture Notes in Mathematics, 469:30–63, 1975. CODEN LNMAA2. ISBN 3-540-07179-2 (print), 3-540-37519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088830/.

Binz:1975:CC

[Bin75d] Prof. Dr. Ernst Binz. Continuous Convergence on C(X), volume 469 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07179-2 (print), 3-540-37519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 125 (est.) pp. LCCN QA3 .L28 no. 469. URL http:

//link.springer.com/book/10.1007/BFb0088826;   http://
www.springerlink.com/content/978-3-540-37519-7.

Binz:1975:CS

[Bin75e] Prof. Dr. Ernst Binz. Convergence spaces. Lecture Notes in Mathematics, 469:1–6, 1975. CODEN LNMAA2. ISBN 3-540-07179-2 (print), 3-540-37519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088827/.

Binz:1975:FA

[Bin75f] Prof. Dr. Ernst Binz. Function algebras. Lecture Notes in Mathematics, 469:7–16, 1975. CODEN LNMAA2. ISBN 3-540-07179-2 (print), 3-540-37519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088828/.

Binz:1975:FAD

[Bin75g] Prof. Dr. Ernst Binz. Functional analytic description of some types of convergence spaces. Lecture Notes in Mathematics, 469: 106–125, 1975. CODEN LNMAA2. ISBN 3-540-07179-2 (print), 3-540-37519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088832/.

Binz:1975:URC

[Bin75h] Prof. Dr. Ernst Binz. Universal representations of convergence algebras and some general remarks of function spaces. Lecture Notes in Mathematics, 469:64–105, 1975. CODEN LNMAA2. ISBN 3-540-07179-2 (print), 3-540-37519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088831/.

Binz:1975:VST

[Bin75i] Prof. Dr. Ernst Binz. Vector space topologies on C(X) for which the evaluation map is continuous. Lecture Notes in Mathematics, 469:16–29, 1975. CODEN LNMAA2. ISBN 3-540-07179-2 (print), 3-540-37519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088829/.

Binz:1976:CGW

[Bin76] E. Binz. Charaktergruppen von Gruppen von S^{1}-wertigen stetigen Funktionen. (German) []. Lecture Notes in Mathematics, 540: 43–92, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080853/.

Binz:1977:EPD

[Bin77] E. Binz. On an extension of Pontryagin’s duality theory. Lecture Notes in Mathematics, 609:1–20, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068668/.

Birman:1979:RTF

[Bir79] Joan S. Birman. A representation theorem for fibered knots and their monodromy maps. Lecture Notes in Mathematics, 722:1– 8, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063183/.

Bismut:1978:ATD

[Bis78a] Jean-Michel Bismut. Applications de la théorie du potentiel à des problémes de contrôle. (French) []. Lecture Notes in Mathematics, 681:7–17, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065865/.

Bismut:1978:CSL

[Bis78b] Jean-Michel Bismut. Contrôle des systèmes linéaires quadratiques: Applications de l’intégrale stochastique. (French) []. Lecture Notes in Mathematics, 649:180–264, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064606/.

Bitsadze:1976:CNP

[Bit76] A. V. Bitsadze. On a class of nonlinear partial differential equations. Lecture Notes in Mathematics, 561:10–16, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087621/.

Berglund:1978:CRT

[BJM78a] John F. Berglund, Hugo D. Junghenn, and Paul Milnes. Compact Right Topological Semigroups and Generalizations of Almost Periodicity, volume 663 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08919-5 (print), 3-540-35759-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 243 pp. LCCN QA3 .L28 no. 663. URL http:

//link.springer.com/book/10.1007/BFb0061381;   http://
www.springerlink.com/content/978-3-540-35759-9.

Berglund:1978:E

[BJM78b] John F. Berglund, Hugo D. Junghenn, and Paul Milnes. Examples. Lecture Notes in Mathematics, 663:166–179, 1978. CODEN LNMAA2. ISBN 3-540-08919-5 (print), 3-540-35759-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061386/.

Berglund:1978:FPL

[BJM78c] John F. Berglund, Hugo D. Junghenn, and Paul Milnes. Fixed points and left invariant means on subspaces of C(S). Lecture Notes in Mathematics, 663:150–165, 1978. CODEN LN-MAA2. ISBN 3-540-08919-5 (print), 3-540-35759-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061385/.

Berglund:1978:P

[BJM78d] John F. Berglund, Hugo D. Junghenn, and Paul Milnes. Preliminaries. Lecture Notes in Mathematics, 663:1–27, 1978. CODEN LNMAA2. ISBN 3-540-08919-5 (print), 3-540-35759-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061382/.

Berglund:1978:SCS

[BJM78e] John F. Berglund, Hugo D. Junghenn, and Paul Milnes. The structure of compact semigroups. Lecture Notes in Mathematics, 663:28–90, 1978. CODEN LNMAA2. ISBN 3-540-08919-5 (print), 3-540-35759-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061383/.

Berglund:1978:SC

[BJM78f] John F. Berglund, Hugo D. Junghenn, and Paul Milnes. Subspaces of C(S) and compactifications of S. Lecture Notes in Mathematics, 663:91–149, 1978. CODEN LNMAA2. ISBN 3-540-08919-5 (print), 3-540-35759-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061384/.

Bjornestaal:1977:SSP

[Bjö77] B. O. Björnestål. Stability of Steiner points. Lecture Notes in Mathematics, 571:6–13, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086561/.

Birch:1975:MFO

[BK75a] Dr. Bryan J. Birch and Prof. Willem Kuyk, editors. Modular Functions of One Variable IV: Proceedings of the International Summer School, University of Antwerp, RUCA, July 17–August 3, 1972, volume 476 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 476. URL http:

//link.springer.com/book/10.1007/BFb0097580;   http://
www.springerlink.com/content/978-3-540-37588-3.

Birch:1975:RI

[BK75b] Dr. Bryan J. Birch and Prof. Willem Kuyk. Remarks on isogenies. Lecture Notes in Mathematics, 476:78–80, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097585/.

Birch:1975:SRT

[BK75c] Dr. Bryan J. Birch and Prof. Willem Kuyk. Sources and reliability of the tables. Lecture Notes in Mathematics, 476:74–77, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097584/.

Birch:1975:Ta

[BK75d] Dr. Bryan J. Birch and Prof. Willem Kuyk. Table 1. Lecture Notes in Mathematics, 476:81–113, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097586/.

Birch:1975:Tb

[BK75e] Dr. Bryan J. Birch and Prof. Willem Kuyk. Table 2. Lecture Notes in Mathematics, 476:114–115, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097587.

Birch:1975:Tc

[BK75f] Dr. Bryan J. Birch and Prof. Willem Kuyk. Table 3. Lecture Notes in Mathematics, 476:116–122, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097588/.

Birch:1975:Td

[BK75g] Dr. Bryan J. Birch and Prof. Willem Kuyk. Table 4. Lecture Notes in Mathematics, 476:123–134, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097589/.

Birch:1975:Te

[BK75h] Dr. Bryan J. Birch and Prof. Willem Kuyk. Table 5. Lecture Notes in Mathematics, 476:135–141, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097590/.

Birch:1975:Tf

[BK75i] Dr. Bryan J. Birch and Prof. Willem Kuyk. Table 6. Lecture Notes in Mathematics, 476:142–144, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097591/.

Bellow:1976:MTP

[BK76a] Alexandra Bellow and Dietrich Kölzow, editors. Measure Theory: Proceedings of the Conference Held at Oberwolfach, 15– 21 June, 1975, volume 541 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 541. URL http:

//link.springer.com/book/10.1007/BFb0081033;   http://
www.springerlink.com/content/978-3-540-38107-5.

Bloom:1976:ES

[BK76b] Clifford O. Bloom and Nicholas D. Kazarinoff. Existence of solutions. Lecture Notes in Mathematics, 522:98–102, 1976. CODEN LNMAA2. ISBN 3-540-07698-0 (print), 3-540-38229-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079582/.

Bloom:1976:GES

[BK76c] Clifford O. Bloom and Nicholas D. Kazarinoff. Global existence, smoothness, and nonfocussing of optical paths in a refractive medium. Lecture Notes in Mathematics, 522:48–83, 1976. CO-DEN LNMAA2. ISBN 3-540-07698-0 (print), 3-540-38229-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079580/.

Bloom:1976:I

[BK76d] Clifford O. Bloom and Nicholas D. Kazarinoff. Introduction. Lecture Notes in Mathematics, 522:1–21, 1976. CODEN LN-MAA2. ISBN 3-540-07698-0 (print), 3-540-38229-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079578/.

Bloom:1976:PB

[BK76e] Clifford O. Bloom and Nicholas D. Kazarinoff. A priori bounds. Lecture Notes in Mathematics, 522:22–47, 1976. CODEN LN-MAA2. ISBN 3-540-07698-0 (print), 3-540-38229-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079579/.

Bloom:1976:SWR

[BK76f] Clifford O. Bloom and Nicholas D. Kazarinoff. Short Wave Radiation Problems in Inhomogeneous Media: Asymptotic Solutions, volume 522 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07698-0 (print), 3-540-38229-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 104 pp. LCCN QA3 .L28 no. 522. URL http:

//link.springer.com/book/10.1007/BFb0079577;   http://
www.springerlink.com/content/978-3-540-38229-4.

Bloom:1976:UAS

[BK76g] Clifford O. Bloom and Nicholas D. Kazarinoff. A uniform approximation to the solution of Ursell’s radiating body problem. Lecture Notes in Mathematics, 522:84–97, 1976. CODEN LNMAA2. ISBN 3-540-07698-0 (print), 3-540-38229-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079581/.

Barbey:1977:AAF

[BK77a] Klaus Barbey and Heinz König. Abstract Analytic Function Theory and Hardy Algebras, volume 593 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 260 pp. LCCN QA3

.L28 no. 593. URL http://link.springer.com/book/10.1007/
BFb0086694;   http://www.springerlink.com/content/978-
3-540-37350-6.

Barbey:1977:AC

[BK77b] Klaus Barbey and Heinz König. The abstract conjugation. Lecture Notes in Mathematics, 593:108–148, 1977. CODEN LN-MAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086700/.

Barbey:1977:AHA

[BK77c] Klaus Barbey and Heinz König. The abstract Hardy algebra situation. Lecture Notes in Mathematics, 593:59–80, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086698/.

Barbey:1977:ADI

[BK77d] Klaus Barbey and Heinz König. Analytic disks and isomorphisms with the unit disk situation. Lecture Notes in Mathematics, 593: 149–171, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086701/.

Barbey:1977:BVT

[BK77e] Klaus Barbey and Heinz König. Boundary value theory for harmonic and holomorphic functions in the unit disk. Lecture Notes in Mathematics, 593:1–21, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086695/.

Barbey:1977:EAH

[BK77f] Klaus Barbey and Heinz König. Elements of abstract Hardy algebra theory. Lecture Notes in Mathematics, 593:81–107, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086699/.

Barbey:1977:FACb

[BK77g] Klaus Barbey and Heinz König. Function algebras on compact planar sets. Lecture Notes in Mathematics, 593:198–235, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086704/.

Barbey:1977:FAB

[BK77h] Klaus Barbey and Heinz König. Function algebras: The boundedmeasurable situation. Lecture Notes in Mathematics, 593:22– 43, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086696/.

Barbey:1977:FACa

[BK77i] Klaus Barbey and Heinz König. Function algebras: The compactcontinuous situation. Lecture Notes in Mathematics, 593:44– 58, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086697/.

Barbey:1977:LDS

[BK77j] Klaus Barbey and Heinz König. Logmodular densities and small extensions. Lecture Notes in Mathematics, 593:181–197, 1977. CODEN LNMAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086703/.

Barbey:1977:WC

[BK77k] Klaus Barbey and Heinz König. Weak compactness of M. Lecture Notes in Mathematics, 593:172–180, 1977. CODEN LN-MAA2. ISBN 3-540-08252-2 (print), 3-540-37350-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086702/.

Badrikian:1976:EEP

[BKK76] Albert Badrikian, J. F. C. (John Frank Charles) Kingman, and James Kuelbs, editors. École d’été de probabilités de Saint-Flour V — 1976. (French) [Saint-Flour Summer School on Probability V — 1976], volume 539 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07858-4 (print), 3-540-38119-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 314 pp. LCCN QA3 .L28 no. 539; QA3 .L28 no. 539

QA273.43. URL http://link.springer.com/book/10.1007/
BFb0079695;   http://www.springerlink.com/content/978-
3-540-38119-8.

Ball:1978:TEN

[BKM78] J. M. Ball, R. J. Knops, and J. E. Marsden. Two examples in nonlinear elasticity. Lecture Notes in Mathematics, 665:41–48, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061796/.

Bigard:1977:GAR

[BKW77] Alain Bigard, Klaus Keimel, and Samuel Wolfenstein. Groupes et Anneaux Réticulés, volume 608 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 334 pp. LCCN QA3

.L28 no. 608. URL http://link.springer.com/book/10.1007/
BFb0067004;   http://www.springerlink.com/content/978-
3-540-37111-3.

Bensoussan:1975:DPB

[BL75] A. Bensoussan and J.-L. Lions. Diffusion processes in bounded domains and singular perturbation problems for variational inequalities with Neumann boundary conditions. Lecture Notes in Mathematics, 451:8–25, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068576/.

Blattner:1975:IOH

[Bla75] Robert J. Blattner. Intertwining operators and the half-density pairing. Lecture Notes in Mathematics, 466:1–12, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082192/.

Blair:1976:ACM

[Bla76a] David Ervin Blair. Almost contact manifolds. Lecture Notes in Mathematics, 509:17–35, 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079309/.

Blair:1976:CM

[Bla76b] David Ervin Blair. Contact manifolds. Lecture Notes in Mathematics, 509:1–16, 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079308/.

Blair:1976:CMR

[Bla76c] David Ervin Blair. Contact Manifolds in Riemannian Geometry, volume 509 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 146 pp. LCCN QA3 .L28 no. 509. URL http:

//link.springer.com/book/10.1007/BFb0079307;   http://
www.springerlink.com/content/978-3-540-38154-9.

Blair:1976:GIC

[Bla76d] David Ervin Blair. Geometric interpretation of the contact condition. Lecture Notes in Mathematics, 509:36–46, 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079310/.

Blair:1976:CSS

[Bla76e] David Ervin Blair. K-contact and Sasakian structures. Lecture Notes in Mathematics, 509:47–91, 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079311/.

Blair:1976:NEF

[Bla76f] David Ervin Blair. Non-existence of flat contact metric structures. Lecture Notes in Mathematics, 509:115–126, 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079313/.

Blair:1976:SSF

[Bla76g] David Ervin Blair. Sasakian space forms. Lecture Notes in Mathematics, 509:92–114, 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079312/.

Blair:1976:TSB

[Bla76h] David Ervin Blair. The tangent sphere bundle. Lecture Notes in Mathematics, 509:127–138, 1976. CODEN LNMAA2. ISBN 3-540-07626-3 (print), 3-540-38154-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079314/.

Blatt:1976:RAS

[Bla76i] Hans-Peter Blatt. Rationale approximierbarkeit singulärer funktionen über [0, ∞]. (German) []. Lecture Notes in Mathematics, 556:1–15, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087393/.

Blattner:1977:MGN

[Bla77] Robert J. Blattner. The metalinear geometry of non-real polarizations. Lecture Notes in Mathematics, 570:11–45, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087780/.

Bolley:1978:NRV

[BLC78] Pierre Bolley, Pham The Lai, and Jacques Camus. Noyau, résolvante et valeurs propres d’une classe d’opérateurs elliptiques et dégénérés. (French) []. Lecture Notes in Mathematics, 660:33– 46, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082292/.

Bloom:1977:CES

[Blo77] Thomas Bloom. Sur le contact entre sous-variétés réelles et sousvariétés complexes de C^{n}. (French) []. Lecture Notes in Mathematics, 578:28–43, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091461/.

Block:1979:IRW

[Blo79a] Richard E. Block. The irreducible representations of the Weyl algebra A_{1}. Lecture Notes in Mathematics, 740:69–79, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071053/.

Bloom:1979:ABD

[Blo79b] Frederick Bloom. Anelastic behavior and dislocation motion. Lecture Notes in Mathematics, 733:101–155, 1979. CODEN LN-MAA2. ISBN 3-540-09528-4 (print), 3-540-35045-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063367/.

Bloom:1979:GEB

[Blo79c] Frederick Bloom. Generalized elastic bodies. Lecture Notes in Mathematics, 733:63–100, 1979. CODEN LNMAA2. ISBN 3-540-09528-4 (print), 3-540-35045-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063366/.

Bloom:1979:MUE

[Blo79d] Frederick Bloom. Material uniformity in elasticity. Lecture Notes in Mathematics, 733:18–62, 1979. CODEN LNMAA2. ISBN 3-540-09528-4 (print), 3-540-35045-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063365/.

Bloom:1979:MP

[Blo79e] Frederick Bloom. Mathematical preliminaries. Lecture Notes in Mathematics, 733:1–17, 1979. CODEN LNMAA2. ISBN 3-540-09528-4 (print), 3-540-35045-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063364/.

Bloom:1979:MDG

[Blo79f] Frederick Bloom. Modern Differential Geometric Techniques in the Theory of Continuous Distributions of Dislocations, volume 733 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09528-4 (print), 3-540-35045-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 202 (est.) pp. LCCN QA3 .L28 no. 733. URL http:

//link.springer.com/book/10.1007/BFb0063363;   http://
www.springerlink.com/content/978-3-540-35045-3.

Bloom:1979:SRD

[Blo79g] Frederick Bloom. Some recent directions in current research. Lecture Notes in Mathematics, 733:199–202, 1979. CODEN LN-MAA2. ISBN 3-540-09528-4 (print), 3-540-35045-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063369/.

Bloom:1979:TDM

[Blo79h] Frederick Bloom. Thermodynamics and dislocation motion. Lecture Notes in Mathematics, 733:156–198, 1979. CODEN LN-MAA2. ISBN 3-540-09528-4 (print), 3-540-35045-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063368/.

Bensoussan:1977:PAC

[BLP77] A. Bensoussan, J.-L. Lions, and G. Papanicolaou. Perturbations et ‘augmentation’ des conditions initiales. (French) []. Lecture Notes in Mathematics, 594:10–29, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086078/.

Bensoussan:1979:ABT

[BLP79] A. Bensoussan, J.-L. Lions, and G. C. Papanicolaou. Asymptotics for branching transport processes. Lecture Notes in Mathematics, 704:317–329, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063629/.

Burghelea:1975:AC

[BLR75a] Prof. Dan Burghelea, Prof. Richard Lashof, and Prof. Melvin Rothenberg. Automorphisms and concordances. Lecture Notes in Mathematics, 473:28–44, 1975. CODEN LNMAA2. ISBN 3-540-07182-2 (print), 3-540-37523-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079984/.

Burghelea:1975:FM

[BLR75b] Prof. Dan Burghelea, Prof. Richard Lashof, and Prof. Melvin Rothenberg. Fibrations over manifolds. Lecture Notes in Mathematics, 473:55–101, 1975. CODEN LNMAA2. ISBN 3-540-07182-2 (print), 3-540-37523-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079986/.

Burghelea:1975:FS

[BLR75c] Prof. Dan Burghelea, Prof. Richard Lashof, and Prof. Melvin Rothenberg. Fibrations over spheres. Lecture Notes in Mathematics, 473:45–54, 1975. CODEN LNMAA2. ISBN 3-540-07182-2 (print), 3-540-37523-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079985/.

Burghelea:1975:GAM

[BLR75d] Prof. Dan Burghelea, Prof. Richard Lashof, and Prof. Melvin Rothenberg. Groups of Automorphisms of Manifolds, volume 473 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07182-2 (print), 3-540-37523-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 156 pp. LCCN QA3 .L28 no. 473. URL http:

//link.springer.com/book/10.1007/BFb0079981;   http://
www.springerlink.com/content/978-3-540-37523-4.

Burghelea:1975:LDF

[BLR75e] Prof. Dan Burghelea, Prof. Richard Lashof, and Prof. Melvin Rothenberg. Lemma of disjunction — 2^{nd} form. Lecture Notes in Mathematics, 473:20–27, 1975. CODEN LNMAA2. ISBN 3-540-07182-2 (print), 3-540-37523-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079983/.

Burghelea:1975:MLD

[BLR75f] Prof. Dan Burghelea, Prof. Richard Lashof, and Prof. Melvin Rothenberg. Morlet’s lemma of disjunction. Lecture Notes in Mathematics, 473:1–19, 1975. CODEN LNMAA2. ISBN 3-540-07182-2 (print), 3-540-37523-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079982/.

Burghelea:1975:S

[BLR75g] Prof. Dan Burghelea, Prof. Richard Lashof, and Prof. Melvin Rothenberg. The structure of A(M ×S^{1}). Lecture Notes in Mathematics, 473:102–109, 1975. CODEN LNMAA2. ISBN 3-540-07182-2 (print), 3-540-37523-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079987/.

Bryant:1975:FSM

[BLS75] J. L. Bryant, R. C. Lacher, and B. J. Smith. Free spheres with mapping cylinder neighborhoods. Lecture Notes in Mathematics, 438:58–65, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066107/.

Blumenthal:1975:FPP

[Blu75] Leonard M. Blumenthal. Four-point properties and norm postulates. Lecture Notes in Mathematics, 490:1–13, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081126/.

Blum:1979:NSE

[Blu79] E. K. Blum. Numerical solution of eigentuple-eigenvector problems in Hilbert spaces. Lecture Notes in Mathematics, 701:80–96, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062075/.

Banchoff:1977:WDS

[BM77a] Thomas Banchoff and Clint McCrory. Whitney duality and singularities of projections. Lecture Notes in Mathematics, 597:68– 81, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085348/.

Bernard:1977:DAM

[BM77b] A. Bernard and B. Maisonneuve. Decomposition atomique de martingales de la classe H^{1}. (French) []. Lecture Notes in Mathematics, 581:303–323, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087197/.

Barratt:1978:GAHa

[BM78a] M. G. Barratt and M. E. Mahowald, editors. Geometric Applications of Homotopy Theory I: Proceedings, Evanston, March 21–26, 1977, volume 657 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 657. URL http:

//link.springer.com/book/10.1007/BFb0069224;   http://
www.springerlink.com/content/978-3-540-35809-1.

Barratt:1978:GAHb

[BM78b] M. G. Barratt and M. E. Mahowald, editors. Geometric Applications of Homotopy Theory II: Proceedings, Evanston, March 21–26, 1977, volume 658 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 658. URL http:

//link.springer.com/book/10.1007/BFb0068706;   http://
www.springerlink.com/content/978-3-540-35808-4.

Bainov:1979:AAM

[BM79a] D. D. Baĭnov and S. D. Milusheva. Application of the averaging method for the solution of boundary problems for ordinary differential and integro-differential equations. Lecture Notes in Mathematics, 703:16–24, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067251/.

Burden:1979:BSC

[BM79b] C. W. Burden and C. J. Mulvey. Banach spaces in categories of sheaves. Lecture Notes in Mathematics, 753:169–196, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061818/.

Bohmer:1976:SFP

[BMS76] Prof. Dr. Klaus Böhmer, Prof. Dr. Günter Meinardus, and Prof. Dr. Walter Schempp, editors. Spline Functions: Proceedings of an International Symposium Held at Karlsruhe, Germany, May 20–23, 1975, volume 501 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 501. URL http:

//link.springer.com/book/10.1007/BFb0079739;   http://
www.springerlink.com/content/978-3-540-38073-3.

Brauner:1977:SPM

[BN77] C. M. Brauner and B. Nicolaenko. Singular perturbation, multiple solutions and hysteresis in a nonlinear problem. Lecture Notes in Mathematics, 594:50–76, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086081/.

Boardman:1978:SMO

[Boa78] J. M. Boardman. Splittings of MU and other spectra. Lecture Notes in Mathematics, 658:27–79, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068710/.

Boardman:1979:OBP

[Boa79] J. M. Boardman. Original Brown–Peterson spectra. Lecture Notes in Mathematics, 741:355–372, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062149/.

Bochnak:1978:PHP

[Boc78] Jacek Bochnak. Sur le 17ème problème de Hilbert pour les fonctions de Nash. (French) []. Lecture Notes in Mathematics, 694: 1–14, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063242/.

Boffa:1975:NEC

[Bof75] Maurice Boffa. A note on existentially complete division rings. Lecture Notes in Mathematics, 498:56–59, 1975. CODEN LN-MAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080974/.

Boffa:1977:RTT

[Bof77] M. Boffa. A reduction of the theory of types. Lecture Notes in Mathematics, 619:95–100, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067644/.

Bohnke:1975:SSP

[Boh75] Georges Bohnke. Synthèses des sphères pour l’algèbre de Sobolev. (French) []. Lecture Notes in Mathematics, 497:1–15, 1975. CO-DEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078010/.

Bohme:1976:SMS

[Böh76a] Reinhold Böhme. Stability of minimal surfaces. Lecture Notes in Mathematics, 561:123–137, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087631/.

Bohmer:1976:DCM

[Böh76b] K. Böhmer. A defect correction method for functional equations. Lecture Notes in Mathematics, 556:16–29, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087394/.

Bohl:1978:TBV

[Boh78] Erich Bohl. On two boundary value problems in nonlinear elasticity from a numerical viewpoint. Lecture Notes in Mathematics, 679:1–14, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067862/.

Boland:1975:HLF

[Bol75] Philip J. Boland. H∂ lomorphic functions on DFN (dual of Fréchet nuclear) spaces. Lecture Notes in Mathematics, 474:109–113, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077402/.

Bollobas:1978:CST

[Bol78] Béla Bollobás. Cycles and semi-topological configurations. Lecture Notes in Mathematics, 642:66–74, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070365/.

Bombieri:1976:LBZ

[Bom76] Enrico Bombieri. A lower bound for the zeros of Riemann’s zeta function on the critical line. Lecture Notes in Mathematics, 514:176–182, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080066/.

Boman:1977:CTG

[Bom77] Jan Boman. Comparison theorems for generalized moduli of continuity. Vector-valued measures. Lecture Notes in Mathematics, 571:38–52, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086563/.

Bonnin:1975:RHD

[Bon75] Olivier Bonnin. Représentation holomorphe des distributions tempérées transformation de Fourier–Borel opérateurs de dérivations partielles de type Hilbert–Schmidt en dimension infinie (d’après Thomas A. W. Dwyer, III). (French) []. Lecture Notes in Mathematics, 474:123–141, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077404/.

Bony:1976:PBM

[Bon76] Jean-Michel Bony. Polynômes de Bernstein et monodromie. (French) []. Lecture Notes in Mathematics, 514:97–110, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080060/.

Bondy:1978:RLD

[Bon78a] J. A. Bondy. Reflections on the legitimate deck problem. Lecture Notes in Mathematics, 686:1–12, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062511/.

Bony:1978:HED

[Bon78b] Jean-Michel Bony. Hyperfonctions et équations aux dérivées partielles [d’après M. Sato, T. Kawai et M. Kashiwara]. (French)

[]. Lecture Notes in Mathematics, 677:73–87, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070755/.

Bony:1978:PFS

[Bon78c] Jean-Michel Bony. Prolongement à la frontière des solutions du problème des derivées obliques. (French) []. Lecture Notes in Mathematics, 660:47–58, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082293/.

Boos:1975:LLC

[Boo75] William Boos. Lectures on large cardinal axioms. Lecture Notes in Mathematics, 499:25–88, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079417/.

Bopp:1979:DTP

[Bop79] Nicole Bopp. Les distributions de type positif relativement à un groupe fini d’isométries. (French) []. Lecture Notes in Mathematics, 739:15–50, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062487/.

Borel:1975:CCG

[Bor75a] Armand Borel. Cohomologie de certains groupes discrètes et laplacien p-adique [d’après H. Garland]. (French) []. Lecture Notes in Mathematics, 431:12–35, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066362/.

Borger:1975:RUA

[Bör75b] Egon Börger. Recursively unsolvable algorithmic problems and related questions reexamined. Lecture Notes in Mathematics, 499: 10–24, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079416/.

Borel:1976:FAS

[Bor76] Armand Borel. Formes automorphes et séries de Dirichlet. (French) []. Lecture Notes in Mathematics, 514:183–222, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080067/.

Borell:1978:NCP

[Bor78a] Christer Borell. A note on conditional probabilities of a convex measure. Lecture Notes in Mathematics, 644:68–72, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066840/.

Borell:1978:TPG

[Bor78b] Christer Borell. Tail probabilities in Gauss space. Lecture Notes in Mathematics, 644:73–82, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066841/.

Borho:1978:RAE

[Bor78c] Walter Borho. Recent advances in enveloping algebras of semisimple Lie-algebras [a report on work of N. Conze, J. Dixmier, M. Duflo, J. C. Jantzen, A. Joseph, W. Borho]. Lecture Notes in Mathematics, 677:1–18, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070749/.

Borell:1979:IBS

[Bor79a] Christer Borell. On the integrability of Banach space valued Walsh polynomials. Lecture Notes in Mathematics, 721:1–3, 1979. CO-DEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070847/.

Borger:1979:LCT

[Bör79b] Reinhard Börger. Legitimacy of certain topological completions. Lecture Notes in Mathematics, 719:18–23, 1979. CODEN LN-MAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065254/.

Boruvka:1979:AMT

[Bor79c] O. Borůvka. Algebraic methods in the theory of global properties of the oscillatory equations Y′′ = Q(t)Y. Lecture Notes in Mathematics, 703:35–45, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067254/.

Bulirsch:1975:OOC

[BOS75] Prof. Roland Bulirsch, Prof. Werner Oettli, and Professor Josef Stoer, editors. Optimization and Optimal Control: Proceedings of a Conference Held at Oberwolfach, November 17–23, 1974, volume 477 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 477. URL http://link.springer.com/book/10.1007/
BFb0079162;   http://www.springerlink.com/content/978-
3-540-37591-3.

Bekken:1976:SAF

[BØS76] Otto B. Bekken, Bernt K. Øksendal, and Arne Stray, editors. Spaces of Analytic Functions: Seminar Held at Kristiansand, Norway, June 9–14, 1975, volume 512 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 512. URL http:

//link.springer.com/book/10.1007/BFb0080017;   http://
www.springerlink.com/content/978-3-540-38201-0.

Bosak:1978:GUW

[Bos78] Juraj Bosák. Graphs with unique walks, trails or paths of given lengths. Lecture Notes in Mathematics, 642:75–85, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070366/.

Bott:1978:SFC

[Bot78] Raoul Bott. On some formulas for the characteristic classes of group-actions. Lecture Notes in Mathematics, 652:25–61, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063501/.

Bourbaki:1975:SBV

[Bou75] Nicolas Bourbaki, editor. Séminaire Bourbaki vol. 1973/74 Exposés 436–452, volume 431 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 0-387-07023-0, 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28

no. 431. URL http://link.springer.com/book/10.1007/
BFb0066360;   http://www.springerlink.com/content/978-
3-540-37310-0.

Bourdaud:1976:SCC

[Bou76a] Gérard Bourdaud. Some cartesian closed topological categories of convergence spaces. Lecture Notes in Mathematics, 540:93– 108, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080854/.

Boutot:1976:FCL

[Bou76b] Jean-François Boutot. Frobenius et cohomologie locale. (French) []. Lecture Notes in Mathematics, 514:1–19, 1976. CODEN LN-MAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080054/.

Bouchet:1978:TIS

[Bou78a] André Bouchet. Triangular imbeddings into surfaces of a join of equicardinal independent sets following an Eulerian graph. Lecture Notes in Mathematics, 642:86–115, 1978. CODEN LN-MAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070367/.

Boutot:1978:ADF

[Bou78b] Jean-François Boutot. Applications diverses. (French) []. Lecture Notes in Mathematics, 632:64–97, 1978. CODEN LNMAA2. ISBN 3-540-08650-1 (print), 3-540-35936-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068427/.

Boutot:1978:CAC

[Bou78c] Jean-François Boutot. Compléments au critère de représentabilité d’Artin. (French) []. Lecture Notes in Mathematics, 632:1– 22, 1978. CODEN LNMAA2. ISBN 3-540-08650-1 (print), 3-540-35936-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068425/.

Boutot:1978:FPS

[Bou78d] Jean-François Boutot. Le foncteur \underline{P ic}_{X̃/k}^{#} pour un R-schéma propre X. (French) []. Lecture Notes in Mathematics, 632:98– 132, 1978. CODEN LNMAA2. ISBN 3-540-08650-1 (print), 3-540-35936-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068428/.

Boutot:1978:RDF

[Bou78e] Jean-François Boutot. Représentabilité du foncteur de Picard local. (French) []. Lecture Notes in Mathematics, 632:23–63, 1978. CODEN LNMAA2. ISBN 3-540-08650-1 (print), 3-540-35936-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068426/.

Boutot:1978:RAS

[Bou78f] Jean-François Boutot. Retour Au Schéma de Picard local. (French) []. Lecture Notes in Mathematics, 632:133–159, 1978. CODEN LNMAA2. ISBN 3-540-08650-1 (print), 3-540-35936-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068429/.

Boutot:1978:SPL

[Bou78g] Jean-François Boutot. Schéma de Picard Local. (French) [Local Picard scheme], volume 632 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08650-1 (print), 3-540-35936-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 165 pp. LCCN QA3

.L28 no. 632. URL http://link.springer.com/book/10.1007/
BFb0068424;   http://www.springerlink.com/content/978-
3-540-35936-4.

Bougerol:1979:MUF

[Bou79a] Philippe Bougerol. Une majoration universelle des fonctions de concentration sur les groupes localement compacts non compacts. (French) []. Lecture Notes in Mathematics, 706:36–40, 1979. CO-DEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063112/.

Bourguignon:1979:PFC

[Bou79b] Jean-Pierre Bourguignon. Premières formes de Chern des variétés kählériennes compactes [d’après E. Calabi, T. Aubin et S. T. Yau]. (French) []. Lecture Notes in Mathematics, 710:1–21, 1979. CO-DEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069970/.

Bowen:1975:AD

[Bow75a] Prof. Rufus Bowen. Axiom a diffeomorphisms. Lecture Notes in Mathematics, 470:68–89, 1975. CODEN LNMAA2. ISBN 3-540-07187-3 (print), 3-540-37534-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081283/.

Bowen:1975:ESE

[Bow75b] Prof. Rufus Bowen. Equilibrium States and the Ergodic Theory of Anosov Diffeomorphisms, volume 470 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07187-3 (print), 3-540-37534-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 108 pp. LCCN QA3

.L28 no. 470. URL http://link.springer.com/book/10.1007/
BFb0081279;   http://www.springerlink.com/content/978-
3-540-37534-0.

Bowen:1975:ETAb

[Bow75c] Prof. Rufus Bowen. Ergodic theory of axiom a diffeomorphisms. Lecture Notes in Mathematics, 470:90–107, 1975. CODEN LN-MAA2. ISBN 3-540-07187-3 (print), 3-540-37534-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081284/.

Bowen:1975:GTF

[Bow75d] Prof. Rufus Bowen. General thermodynamic formalism. Lecture Notes in Mathematics, 470:45–67, 1975. CODEN LNMAA2. ISBN 3-540-07187-3 (print), 3-540-37534-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081282/.

Bowen:1975:GM

[Bow75e] Prof. Rufus Bowen. Gibbs measures. Lecture Notes in Mathematics, 470:4–44, 1975. CODEN LNMAA2. ISBN 3-540-07187-3 (print), 3-540-37534-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081281/.

Bowen:1975:I

[Bow75f] Prof. Rufus Bowen. Introduction. Lecture Notes in Mathematics, 470:1–3, 1975. CODEN LNMAA2. ISBN 3-540-07187-3 (print), 3-540-37534-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081280/.

Bowen:1975:ETAa

[Bow75g] R. Bowen. Ergodic theory of axiom a flows. Lecture Notes in Mathematics, 468:35–36, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082591.

Bowen:1977:LVM

[Bow77] Rufus Bowen. Lecture VIII: a model for Couette flow data. Lecture Notes in Mathematics, 615:117–134, 1977. CODEN LN-MAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068365/.

Bowen:1978:EFG

[Bow78] Rufus Bowen. Entropy and the fundamental group. Lecture Notes in Mathematics, 668:21–29, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101777/.

Bishop:1976:CAC

[BP76] R. E. D. Bishop and W. G. Price. A critical appraisal of certain contemporary ship model testing techniques. Lecture Notes in Mathematics, 503:196–206, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088756/.

Bunce:1978:QEI

[BP78] John W. Bunce and William L. Paschke. Quasi-expectations and injective operator algebras. Lecture Notes in Mathematics, 650: 123–125, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067392/.

Bellomo:1979:DSM

[BP79a] N. Bellomo and G. Pistone. Direct solution methods for a class of integral equations with random distribution of the inhomogeneous part. Lecture Notes in Mathematics, 737:24–34, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064492/.

Benilan:1979:QAN

[BP79b] Philippe Bénilan and Colette Picard. Quelques aspects non linéaires du principe du maximum. (French) []. Lecture Notes in Mathematics, 713:1–37, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071317/.

Bardos:1976:MDN

[BPFS76] C. Bardos, P. Penel, U. Frisch, and P. L. Sulem. Modified dissipativity for a non linear evolution equation arising in turbulence. Lecture Notes in Mathematics, 565:14–23, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091444/.

Berkowitz:1975:GPA

[BR75] H. W. Berkowitz and Prabir Roy. General position and algebraic independence. Lecture Notes in Mathematics, 438:9–15, 1975.

CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066104/.

Barnsley:1976:BBN

[BR76a] M. F. Barnsley and P. D. Robinson. Bivariational bounds on ⟨g, ϕ⟩ for non-linear problems F ϕ = 0. Lecture Notes in Mathematics, 564:26–33, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087323/.

Bochnak:1976:ADG

[BR76b] Jacek Bochnak and Jean-Jacques Risler. Analyse différentielle et géométrie analytique. (French) []. Lecture Notes in Mathematics, 535:63–69, 1976. CODEN LNMAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080496/.

Bochnak:1976:DFD

[BR76c] Jacek Bochnak and Jean-Jacques Risler. Sur la divisibilité des fonctions différentiables. (French) []. Lecture Notes in Mathematics, 535:45–62, 1976. CODEN LNMAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080495/.

Burlet:1976:SAD

[BR76d] Oscar Burlet and Felice Ronga, editors. Singularités d’Applications Différentiables: Séminaire sur les Singularités d’Applications Différentiables Plans-sur-Bex 1975. (French) [], volume 535 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 535. URL http://link.springer.com/book/10.1007/
BFb0080493;   http://www.springerlink.com/content/978-
3-540-38182-2.

Babuska:1977:MPC

[BR77a] I. Babuška and W. C. Rheinboldt. Mathematical problems of computational decisions in the finite element method. Lecture Notes in Mathematics, 606:1–26, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064452/.

Bernard:1977:TSB

[BR77b] Peter Bernard and Tudor Raţiu, editors. Turbulence Seminar: Berkeley 1976/77, volume 615 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 615. URL http:

//link.springer.com/book/10.1007/BFb0068356;   http://
www.springerlink.com/content/978-3-540-37074-1.

Bleuler:1977:DGM

[BR77c] Konrad Bleuler and Axel Reetz, editors. Differential Geometrical Methods in Mathematical Physics: Proceedings of the Symposium Held at the University of Bonn, July 1–4, 1975, volume 570 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 570. URL http://link.springer.com/book/10.1007/
BFb0087778;   http://www.springerlink.com/content/978-
3-540-37498-5.

Benilan:1978:JAN

[BR78a] Philipp Bénilan and Jacques Robert, editors. Journées d’Analyse Non Linéaire: Proceedings, Besançon, France, June 1977, volume 665 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 665. URL http://link.springer.com/book/10.1007/
BFb0061794;   http://www.springerlink.com/content/978-
3-540-35756-8.

Bier:1978:DFM

[BR78b] Thomas Bier and Nigel Ray. Detecting framed manifolds in the 8 and 16 stems. Lecture Notes in Mathematics, 657:32–39, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069226/.

Brackx:1976:NMP

[Bra76a] F. F. Brackx. Non-(k)-monogenic points of functions of a quaternion variable. Lecture Notes in Mathematics, 561:138–149, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087632/.

Braess:1976:NSN

[Bra76b] Prof. Dr. Dietrich Braess. Zur numerischen Stabilität des Newton-Verfahrens bei der nichtlinearen Tschebyscheff-Approximation. (German) []. Lecture Notes in Mathematics, 556:54–62, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087396/.

Brauchli:1976:NDC

[Bra76c] H. Brauchli. On the norm-dependence of the concept of stability. Lecture Notes in Mathematics, 503:235–238, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088759/.

Braaksma:1979:IAT

[Bra79a] Boele Braaksma. Introduction to the asymptotic theory of linear homogeneous difference equations. Lecture Notes in Mathematics, 712:1–19, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062811/.

Brakes:1979:H

[Bra79b] W. R. Brakes. Homogeneity of S^{2} × T^{2}. Lecture Notes in Mathematics, 722:9–15, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063184/.

Brandal:1979:BFA

[Bra79c] Willy Brandal. Basic facts about FGC rings and the local case. Lecture Notes in Mathematics, 723:29–36, 1979. CODEN LN-MAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069026/.

Brandal:1979:CRW

[Bra79d] Willy Brandal. Commutative Rings whose Finitely Generated Modules Decompose, volume 723 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 109 (est.) pp. LCCN QA3

.L28 no. 723. URL http://link.springer.com/book/10.1007/
BFb0069021;   http://www.springerlink.com/content/978-
3-540-35181-8.

Brandal:1979:EAM

[Bra79e] Willy Brandal. Examples of almost maximal Bézout domains. Lecture Notes in Mathematics, 723:98–107, 1979. CODEN LN-MAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069036/.

Brandal:1979:EMV

[Bra79f] Willy Brandal. Examples of maximal valuation rings. Lecture Notes in Mathematics, 723:97, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069035.

Brandal:1979:ETR

[Bra79g] Willy Brandal. Examples of torch rings. Lecture Notes in Mathematics, 723:108–109, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069037.

Brandal:1979:FFA

[Bra79h] Willy Brandal. Further facts about FGC rings and torch rings. Lecture Notes in Mathematics, 723:37–43, 1979. CODEN LN-MAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069027/.

Brandal:1979:LD

[Bra79i] Willy Brandal. h-local domains. Lecture Notes in Mathematics, 723:14–22, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069024/.

Brandal:1979:I

[Bra79j] Willy Brandal. Introduction. Lecture Notes in Mathematics, 723: 3–8, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069022/.

Brandal:1979:LCM

[Bra79k] Willy Brandal. Linearly compact modules and almost maximal rings. Lecture Notes in Mathematics, 723:9–13, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069023/.

Brandal:1979:LPS

[Bra79l] Willy Brandal. Long power series rings. Lecture Notes in Mathematics, 723:82–87, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069033/.

Brandal:1979:MT

[Bra79m] Willy Brandal. The main theorem. Lecture Notes in Mathematics, 723:64–71, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069031/.

Brandal:1979:MCV

[Bra79n] Willy Brandal. Maximally complete valuation domains. Lecture Notes in Mathematics, 723:88–96, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069034/.

Brandal:1979:RTD

[Bra79o] Willy Brandal. Relating topology to the decomposition of modules. Lecture Notes in Mathematics, 723:58–63, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069030/.

Brandal:1979:SCC

[Bra79p] Willy Brandal. The Stone–Cech compactification of N. Lecture Notes in Mathematics, 723:49–57, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069029/.

Brandal:1979:VRB

[Bra79q] Willy Brandal. Valuation rings and Bézout rings. Lecture Notes in Mathematics, 723:23–28, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069025/.

Brandal:1979:V

[Bra79r] Willy Brandal. Valuations. Lecture Notes in Mathematics, 723: 72–81, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069032/.

Brandal:1979:ZPT

[Bra79s] Willy Brandal. The Zariski and patch topologies of the spectrum of a ring. Lecture Notes in Mathematics, 723:44–48, 1979. CODEN LNMAA2. ISBN 3-540-09507-1 (print), 3-540-35181-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069028/.

Brandenburg:1979:NS

[Bra79t] Harald Brandenburg. On E-normal spaces. Lecture Notes in Mathematics, 719:24–34, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065255/.

Breen:1975:TFE

[Bre75a] Lawrence Breen. Un théorème de finitude en K-théorie [d’après D. Quillen]. (French) []. Lecture Notes in Mathematics, 431:36– 57, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066363/.

Breen:1975:IFC

[Bre75b] Marilyn Breen. Intersecting families of convex cover order two. Lecture Notes in Mathematics, 490:218–226, 1975. CODEN LN-MAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081145/.

Brenner:1975:QCC

[Bre75c] Sheila Brenner. Quivers with commutativity conditions and some phenomenology of forms. Lecture Notes in Mathematics, 488:29– 53, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081215/.

Brezin:1975:GMK

[Bre75d] Jonathan Brezin. Geometry and the method of Kirillov. Lecture Notes in Mathematics, 466:13–25, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082193/.

Brezis:1975:ICM

[Bré75e] Haïm Brézis. Interpolation classes for monotone operators. Lecture Notes in Mathematics, 446:65–74, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070597/.

Brezis:1975:NMS

[Bré75f] Haïm Brézis. A new method in the study of subsonic flows. Lecture Notes in Mathematics, 446:50–64, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070596/.

Bremaud:1976:MSM

[Bré76a] P. Brémaud. La méthode des semi-martingales en filtrage quand l’observation est un processus ponctuel marqué. (French) [].

Lecture Notes in Mathematics, 511:1–18, 1976. CODEN LN-MAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101092/.

Bremermann:1976:PRD

[Bre76b] Hans Bremermann. Pattern recognition by deformable prototypes. Lecture Notes in Mathematics, 525:15–57, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077842/.

Brezis:1976:QPO

[Bré76c] H. Brézis. Quelques propriétés des opérateurs monotones et des semi-groupes non linéaires. (French) []. Lecture Notes in Mathematics, 543:56–82, 1976. CODEN LNMAA2. ISBN 3-540-07867-3 (print), 3-540-38075-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079942/.

Brezin:1977:A

[Bre77a] Jonathan Paul Brezin. Afterword. Lecture Notes in Mathematics, 602:171–174, 1977. CODEN LNMAA2. ISBN 3-540-08354-5 (print), 3-540-37281-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069797/.

Brezin:1977:CS

[Bre77b] Jonathan Paul Brezin. Compact solvmanifolds. Lecture Notes in Mathematics, 602:149–170, 1977. CODEN LNMAA2. ISBN 3-540-08354-5 (print), 3-540-37281-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069796/.

Brezin:1977:GT

[Bre77c] Jonathan Paul Brezin. The general theory. Lecture Notes in Mathematics, 602:110–148, 1977. CODEN LNMAA2. ISBN 3-540-08354-5 (print), 3-540-37281-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069795/.

Brezin:1977:HAC

[Bre77d] Jonathan Paul Brezin. Harmonic Analysis on Compact Solvmanifolds, volume 602 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08354-5 (print), 3-540-37281-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 174 (est.) pp. LCCN QA3 .L28 no. 602. URL http:

//link.springer.com/book/10.1007/BFb0069793;   http://
www.springerlink.com/content/978-3-540-37281-3.

Brezin:1977:SBE

[Bre77e] Jonathan Paul Brezin. Some basic examples. Lecture Notes in Mathematics, 602:1–109, 1977. CODEN LNMAA2. ISBN 3-540-08354-5 (print), 3-540-37281-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069794/.

Brezinski:1977:ACA

[Bre77f] Claude Brezinski. Accélération de la Convergence en Analyse Numérique, volume 584 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 313 pp. LCCN QA3

.L28 no. 584. URL http://link.springer.com/book/10.1007/
BFb0089363;   http://www.springerlink.com/content/978-
3-540-37373-5.

Brezinski:1977:APC

[Bre77g] Claude Brezinski. Algorithmes de prédiction continue. (French) []. Lecture Notes in Mathematics, 584:206–246, 1977. CODEN LNMAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089370/.

Brezinski:1977:CSC

[Bre77h] Claude Brezinski. Comparaison de suites convergentes. (French) []. Lecture Notes in Mathematics, 584:2–20, 1977. CODEN LN-MAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089365/.

Brezinski:1977:EDA

[Bre77i] Claude Brezinski. Étude de divers algorithmes d’Accélération de la convergence. (French) []. Lecture Notes in Mathematics, 584: 96–135, 1977. CODEN LNMAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089368/.

Brezinski:1977:IFI

[Bre77j] Claude Brezinski. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 584:1, 1977. CODEN LNMAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089364.

Brezinski:1977:A

[Bre77k] Claude Brezinski. L’ϵ-Algorithme. Lecture Notes in Mathematics, 584:37–95, 1977. CODEN LNMAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089367/.

Brezinski:1977:FCF

[Bre77l] Claude Brezinski. Les fractions continues. (French) []. Lecture Notes in Mathematics, 584:247–295, 1977. CODEN LN-MAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089371/.

Brezinski:1977:PSF

[Bre77m] Claude Brezinski. Les procédés de sommation. (French) []. Lecture Notes in Mathematics, 584:21–36, 1977. CODEN LNMAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089366/.

Brezinski:1977:TSN

[Bre77n] Claude Brezinski. Transformation de suites non scalaires. (French) []. Lecture Notes in Mathematics, 584:136–205, 1977. CODEN LNMAA2. ISBN 3-540-08241-7 (print), 3-540-37373-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089369/.

Brezzi:1977:HMF

[Bre77o] Franco Brezzi. Hybrid methods for fourth order elliptic equations. Lecture Notes in Mathematics, 606:35–46, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064454/.

Brelot:1978:EWP

[Bre78] Marcel Brelot. Équation de Weinstein et potentiels de Marcel Riesz. (French) []. Lecture Notes in Mathematics, 681:18–38, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065866/.

Breach:1979:SDM

[Bre79a] D. R. Breach. Some 2 − (2n + 1, n, n − 1) designs with multiple extensions. Lecture Notes in Mathematics, 748:32–40, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0102682/.

Breitsprecher:1979:LRM

[Bre79b] Siegfried Breitsprecher. Localization with respect to a measure. Lecture Notes in Mathematics, 753:125–156, 1979. CODEN LN-MAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061816/.

Breitsprecher:1979:CMS

[Bre79c] Siegfried Breitsprecher. On the concept of a measurable space I. Lecture Notes in Mathematics, 753:157–168, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061817/.

Brewer:1979:NCS

[Bre79d] Dennis W. Brewer. A nonlinear contraction semigroup for a functional differential equation. Lecture Notes in Mathematics, 737: 35–44, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064493/.

Brezinski:1979:RRS

[Bre79e] Claude Brezinski. Recent references on sequences and series transformations. Lecture Notes in Mathematics, 765:366–374, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085590/.

Brezinski:1979:CCR

[Bre79f] Claude Brezinski. Sur le calcul de certains rapports de déterminants. (French) []. Lecture Notes in Mathematics, 765: 184–210, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085581/.

Briem:1976:EOB

[Bri76] Eggert Briem. Extreme orthogonal boundary measures for A(K) and decompositions for compact convex sets. Lecture Notes in Mathematics, 512:8–16, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080019/.

Brilla:1979:SPM

[Bri79] J. Brilla. Stability problems in mathematical theory of viscoelasticity. Lecture Notes in Mathematics, 703:46–53, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067255/.

Broughan:1975:AS

[Bro75a] Kevin A. Broughan. Algebraic structures. Lecture Notes in Mathematics, 491:104–145, 1975. CODEN LNMAA2. ISBN 3-540-07418-X (print), 3-540-37947-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080890/.

Broughan:1975:CSI

[Bro75b] Kevin A. Broughan. Classes of subsets as invariants. Lecture Notes in Mathematics, 491:1–34, 1975. CODEN LNMAA2. ISBN 3-540-07418-X (print), 3-540-37947-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080887/.

Broughan:1975:GDS

[Bro75c] Kevin A. Broughan. Generation by dense subspaces. Lecture Notes in Mathematics, 491:146–185, 1975. CODEN LNMAA2. ISBN 3-540-07418-X (print), 3-540-37947-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080891/.

Broughan:1975:IRG

[Bro75d] Kevin A. Broughan. Invariants for Real-Generated Uniform Topological and Algebraic Categories, volume 491 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07418-X (print), 3-540-37947-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 185 (est.) pp. LCCN QA3

.L28 no. 491. URL http://link.springer.com/book/10.1007/
BFb0080886;   http://www.springerlink.com/content/978-
3-540-37947-8.

Broughan:1975:TS

[Bro75e] Kevin A. Broughan. Topological spaces. Lecture Notes in Mathematics, 491:63–103, 1975. CODEN LNMAA2. ISBN 3-540-07418-X (print), 3-540-37947-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080889/.

Broughan:1975:US

[Bro75f] Kevin A. Broughan. Uniform spaces. Lecture Notes in Mathematics, 491:35–62, 1975. CODEN LNMAA2. ISBN 3-540-07418-X (print), 3-540-37947-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080888/.

Browder:1975:LFP

[Bro75g] Felix E. Browder. The Lefschetz fixed point theorem and asymptotic fixed point theorems. Lecture Notes in Mathematics, 446: 96–122, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070599/.

Browder:1975:SNI

[Bro75h] Felix E. Browder. Singular nonlinear integral equations of Hammerstein type. Lecture Notes in Mathematics, 446:75–95, 1975.

CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070598/.

Brosowski:1976:SAM

[Bro76a] Prof. Dr. Bruno Brosowski. Zur Stetigen Abhängigkeit der Menge der Minimalpunkte Bei Gewissen Minimierungsaufgaben. (German) []. Lecture Notes in Mathematics, 556:63–72, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087397/.

Brown:1976:CG

[Bro76b] K. S. Brown. Cohomology of groups. Lecture Notes in Mathematics, 551:249–259, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080005/.

Brown:1977:C

[Bro77] Lawrence G. Brown. Characterizing Ext(X). Lecture Notes in Mathematics, 575:10–18, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095698/.

Brockett:1978:LTF

[Bro78a] Roger W. Brockett. Lie theory, functional expansions, and necessary conditions in singular optimal control. Lecture Notes in Mathematics, 680:68–76, 1978. CODEN LNMAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065312/.

Bronner:1978:FMR

[Bro78b] François Bronner. Frontière de Martin des resolvantes recurrentes. (French) []. Lecture Notes in Mathematics, 681:39–53, 1978. CO-DEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0065867/.

Browder:1978:AEC

[Bro78c] William Browder. Algebraic k-theory with coefficients Z/p. Lecture Notes in Mathematics, 657:40–84, 1978. CODEN LN-MAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069227/.

Brown:1978:EGS

[Bro78d] James R. Brown. Ergodic groups of substitution operators associated with algebraically monothetic groups. Lecture Notes in Mathematics, 693:65–70, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064660/.

Brown:1978:HCI

[Bro78e] Robert F. Brown. A homological characterization of inner automorphisms of compact Lie groups. Lecture Notes in Mathematics, 664:38–42, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061688/.

Browder:1979:CIK

[Bro79a] William Browder. Complete intersections and the Kervaire invariant. Lecture Notes in Mathematics, 763:88–108, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088079/.

Brownawell:1979:DGM

[Bro79b] W. Dale Brownawell. On the development of Gelfond’s method. Lecture Notes in Mathematics, 751:18–44, 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062700/.

Bleuler:1978:DGM

[BRP78] Konrad Bleuler, Axel Reetz, and Herbert Rainer Petry, editors. Differential Geometrical Methods in Mathematical Physics II: Proceedings, University of Bonn, July 13–16, 1977, volume 676 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 676. URL http://link.springer.com/book/10.1007/
BFb0063664;   http://www.springerlink.com/content/978-
3-540-35721-6.

Bruneau:1975:ILF

[Bru75] Michel Bruneau. Irregularité locale des fonctions et approximation sur le tore. (French) []. Lecture Notes in Mathematics, 475: 18–25, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074256/.

Brudnyi:1976:PPA

[Bru76a] Ju. A. Brudnyi. Piecewise polynomial approximation, embedding theorem and rational approximation. Lecture Notes in Mathematics, 556:73–98, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087398/.

Brummer:1976:TFS

[Brü76b] G. C. L. Brümmer. Topological functors and structure functors. Lecture Notes in Mathematics, 540:109–135, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080855/.

Brunel:1976:SIO

[Bru76c] Antoine Brunel. Sur les sommes d’itèrès d’un opérateur positif. (French) []. Lecture Notes in Mathematics, 532:19–34, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080170/.

Brunel:1976:PED

[Bru76d] Antoine Brunel. Une preuve élémentaire du lemme de Holley. (French) []. Lecture Notes in Mathematics, 532:15–18, 1976. CO-DEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080169/.

Brunner:1976:ASL

[Bru76e] Hermann Brunner. The approximate solution of linear and nonlinear first-kind integral equations of Volterra type. Lecture Notes in Mathematics, 506:15–27, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080111/.

Bruckner:1978:BTC

[Bru78a] Andrew M. Bruckner. Behavior of typical continuous functions. Lecture Notes in Mathematics, 659:209–223, 1978. CODEN LN-MAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069835/.

Bruckner:1978:CAC

[Bru78b] Andrew M. Bruckner. Continuity and approximate continuity of derivatives. Lecture Notes in Mathematics, 659:45–51, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069825/.

Bruckner:1978:DF

[Bru78c] Andrew M. Bruckner. Darboux functions. Lecture Notes in Mathematics, 659:3–7, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069823/.

Bruckner:1978:DFF

[Bru78d] Andrew M. Bruckner. Darboux functions in the first class of Baire. Lecture Notes in Mathematics, 659:8–44, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069824/.

Bruckner:1978:DG

[Bru78e] Andrew M. Bruckner. Derivatives a.e. and generalizations. Lecture Notes in Mathematics, 659:111–122, 1978. CODEN LN-MAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069830/.

Bruckner:1978:DRF

[Bru78f] Andrew M. Bruckner. Differentiation of Real Functions, volume 659 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 247 pp. LCCN QA3 .L28 no. 659. URL http:

//link.springer.com/book/10.1007/BFb0069821;   http://
www.springerlink.com/content/978-3-540-35776-6.

Bruckner:1978:EDF

[Bru78g] Andrew M. Bruckner. The extreme derivates of a function. Lecture Notes in Mathematics, 659:52–70, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069826/.

Bruckner:1978:GD

[Bru78h] Andrew M. Bruckner. Generalized derivatives. Lecture Notes in Mathematics, 659:147–172, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069832/.

Bruckner:1978:MT

[Bru78i] Andrew M. Bruckner. Miscellaneous topics. Lecture Notes in Mathematics, 659:224–233, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069836/.

Bruckner:1978:M

[Bru78j] Andrew M. Bruckner. Monotonicity. Lecture Notes in Mathematics, 659:173–198, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069833/.

Bruckner:1978:P

[Bru78k] Andrew M. Bruckner. Preliminaries. Lecture Notes in Mathematics, 659:1–2, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069822.

Bruckner:1978:PCD

[Bru78l] Andrew M. Bruckner. The problem of characterizing derivatives. Lecture Notes in Mathematics, 659:98–110, 1978. CODEN LN-MAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069829/.

Bruckner:1978:RP

[Bru78m] Andrew M. Bruckner. Reconstruction of the primitive. Lecture Notes in Mathematics, 659:71–84, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069827/.

Bruckner:1978:SDS

[Bru78n] Andrew M. Bruckner. Stationary and determining sets. Lecture Notes in Mathematics, 659:199–208, 1978. CODEN LN-MAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069834/.

Bruckner:1978:TH

[Bru78o] Andrew M. Bruckner. Transformations via homeomorphisms. Lecture Notes in Mathematics, 659:123–146, 1978. CODEN LN-MAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069831/.

Bruckner:1978:ZC

[Bru78p] Andrew M. Bruckner. The Zahorski classes. Lecture Notes in Mathematics, 659:85–97, 1978. CODEN LNMAA2. ISBN 3-540-08910-1 (print), 3-540-35776-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069828/.

Brumfiel:1978:RHC

[Bru78q] Gregory W. Brumfiel. Realizing homology classes by PL manifolds. Lecture Notes in Mathematics, 664:43–60, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061689/.

Brun:1978:SPG

[Bru78r] Jérôme Brun. Sur la simplification par les grassmanniennes. (French) []. Lecture Notes in Mathematics, 670:205–208, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064401/.

Bruner:1978:AGC

[Bru78s] R. Bruner. Algebraic and geometric connecting homomorphisms in the Adams spectral sequence. Lecture Notes in Mathematics, 658:131–133, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068712/.

Brummer:1979:TPB

[Brü79a] G. C. L. Brümmer. Two procedures in bitopology. Lecture Notes in Mathematics, 719:35–43, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065256/.

Bruneau:1979:VSC

[Bru79b] Michel Bruneau. Sur la p-variation d’une surmartingale continue. (French) []. Lecture Notes in Mathematics, 721:227–232, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070863/.

Barwise:1975:RSM

[BS75] Jon Barwise and John Schlipf. On recursively saturated models of arithmetic. Lecture Notes in Mathematics, 498:42–55, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080973/.

Brezis:1976:HMF

[BS76a] Haïm Brézis and Guido Stampacchia. The hodograph method in fluid-dynamics in the light of variational inequalities. Lecture Notes in Mathematics, 503:239–257, 1976. CODEN LN-MAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0088760/.

Brunel:1976:SIU

[BS76b] Antoine Brunel and Louis Sucheston. On sequences invariant under spreading in Banach spaces. Lecture Notes in Mathematics, 526:21–31, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082339/.

Buckholtz:1977:CAP

[BS77] James D. Buckholtz and Teddy J. Suffridge, editors. Complex Analysis: Proceedings of the Conference held at the University of Kentucky, May 18–22, 1976, volume 599 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 599. URL http:

//link.springer.com/book/10.1007/BFb0096815;   http://
www.springerlink.com/content/978-3-540-37303-2.

Becker:1978:FPI

[BS78] J. C. Becker and R. E. Schultz. Fixed point indices and left invariant framings. Lecture Notes in Mathematics, 657:1–31, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069225/.

Bourgeat:1977:CCP

[BT77] A. Bourgeat and R. Tapiero. Calcul de correcteurs pour un problème de couche limité provenant de la physique du plasma. (French) []. Lecture Notes in Mathematics, 594:39–49, 1977. CO-DEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086080/.

Brenner:1975:BS

[BTW75a] Dr. Philip Brenner, Prof. Vidar Thomée, and Prof. Lars B. Wahlbin. Besov spaces. Lecture Notes in Mathematics, 434:30– 50, 1975. CODEN LNMAA2. ISBN 3-540-07130-X (print), 3-540-37400-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068127/.

Brenner:1975:BSA

[BTW75b] Dr. Philip Brenner, Prof. Vidar Thomée, and Prof. Lars B. Wahlbin. Besov Spaces and Applications to Difference Methods for Initial Value Problems, volume 434 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07130-X (print), 3-540-37400-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 151 (est.) pp. LCCN QA3

.L28 no. 434. URL http://link.springer.com/book/10.1007/
BFb0068125;   http://www.springerlink.com/content/978-
3-540-37400-8.

Brenner:1975:FOH

[BTW75c] Dr. Philip Brenner, Prof. Vidar Thomée, and Prof. Lars B. Wahlbin. First order hyperbolic equations. Lecture Notes in Mathematics, 434:91–131, 1975. CODEN LNMAA2. ISBN 3-540-07130-X (print), 3-540-37400-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068130/.

Brenner:1975:FM

[BTW75d] Dr. Philip Brenner, Prof. Vidar Thomée, and Prof. Lars B. Wahlbin. Fourier multipliers on L_{p}. Lecture Notes in Mathematics, 434:5–29, 1975. CODEN LNMAA2. ISBN 3-540-07130-X (print), 3-540-37400-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068126/.

Brenner:1975:HE

[BTW75e] Dr. Philip Brenner, Prof. Vidar Thomée, and Prof. Lars B. Wahlbin. The heat equation. Lecture Notes in Mathematics, 434: 68–90, 1975. CODEN LNMAA2. ISBN 3-540-07130-X (print), 3-540-37400-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068129/.

Brenner:1975:IVP

[BTW75f] Dr. Philip Brenner, Prof. Vidar Thomée, and Prof. Lars B. Wahlbin. Initial value problems and difference operators. Lecture Notes in Mathematics, 434:51–67, 1975. CODEN LNMAA2. ISBN 3-540-07130-X (print), 3-540-37400-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068128/.

Brenner:1975:SE

[BTW75g] Dr. Philip Brenner, Prof. Vidar Thomée, and Prof. Lars B. Wahlbin. The Schrödinger equation. Lecture Notes in Mathematics, 434:132–151, 1975. CODEN LNMAA2. ISBN 3-540-07130-X (print), 3-540-37400-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068131/.

Buchholz:1975:NKS

[Buc75] W. Buchholz. Normalfunktionen und Konstruktive Systeme von Ordinalzahlen. (German) []. Lecture Notes in Mathematics, 500: 4–25, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0079543/;   http://link.springer.com/chapter/
10.1007/BFb0079544/.

Buckholtz:1977:UTA

[Buc77] J. D. Buckholtz. A uniqueness theorem with application to the Abel series. Lecture Notes in Mathematics, 599:11–16, 1977. CO-DEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096819/.

Buchwalter:1978:RPC

[Buc78] Henri Buchwalter. Le rôle des partitions continues de l’unité dans la théorie des mesures scalaires ou vectorielles. (French) []. Lecture Notes in Mathematics, 644:83–95, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066842/.

Buhler:1977:IMF

[Buh77] Joe Buhler. An icosahedral modular form of weight one. Lecture Notes in Mathematics, 601:289–294, 1977. CODEN LN-MAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063952/.

Buhler:1978:CMF

[Buh78a] Joe P. Buhler. Classical modular forms of weight one. Lecture Notes in Mathematics, 654:70–79, 1978. CODEN LNMAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070424/.

Buhler:1978:IF

[Buh78b] Joe P. Buhler. An icosahedral form. Lecture Notes in Mathematics, 654:80–94, 1978. CODEN LNMAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070425/.

Buhler:1978:IGR

[Buh78c] Joe P. Buhler. Icosahedral Galois Representations, volume 654 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CO-DEN LNMAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 143 pp. LCCN QA3 .L28 no. 654. URL http:

//link.springer.com/book/10.1007/BFb0070418;   http://
www.springerlink.com/content/978-3-540-35818-3.

Buhler:1978:I

[Buh78d] Joe P. Buhler. Introduction. Lecture Notes in Mathematics, 654: 1–5, 1978. CODEN LNMAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070419/.

Buhler:1978:SIR

[Buh78e] Joe P. Buhler. The L-series of an icosahedral representation. Lecture Notes in Mathematics, 654:49–69, 1978. CODEN LN-MAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070423/.

Buhler:1978:LPR

[Buh78f] Joe P. Buhler. Lifting projective representations. Lecture Notes in Mathematics, 654:6–17, 1978. CODEN LNMAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070420/.

Buhler:1978:LPG

[Buh78g] Joe P. Buhler. Local primitive Galois representations. Lecture Notes in Mathematics, 654:18–35, 1978. CODEN LNMAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070421/.

Buhler:1978:TDR

[Buh78h] Joe P. Buhler. Two-dimensional representations over Q. Lecture Notes in Mathematics, 654:36–48, 1978. CODEN LNMAA2. ISBN 3-540-08844-X (print), 3-540-35818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070422/.

Bukovsky:1976:CC

[Buk76] Lev Bukovský. Changing cofinality of X_{2}. Lecture Notes in Mathematics, 537:37–49, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096892/.

Bukowsky:1977:RF

[Buk77] Lev Bukowský. Random forcing. Lecture Notes in Mathematics, 619:101–117, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067645/.

Bultheel:1979:RAP

[Bul79] A. Bultheel. Recursive algorithms for the Padé table: Two approaches. Lecture Notes in Mathematics, 765:211–230, 1979. CO-DEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085582/.

Burgess:1975:SSC

[Bur75a] C. E. Burgess. Semicellular sets in crumpled cubes. Lecture Notes in Mathematics, 438:42–57, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066106/.

Burkard:1975:KOH

[Bur75b] Rainer E. Burkard. Kombinatorische optimierung in halbgruppen. Lecture Notes in Mathematics, 477:1–17, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079163/.

Burlet:1976:PAS

[Bur76] O. Burlet. Propos au sujet des applications différentiables. (French) [On the subject of differentiable applications]. Lecture Notes in Mathematics, 535:187–204, 1976. CODEN LN-MAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080501/.

Butler:1975:CLI

[But75] M. C. R. Butler. On the classification of local integral representations of finite abelian p-groups. Lecture Notes in Mathematics, 488:54–71, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081216/.

Butcher:1976:CIM

[But76] J. C. Butcher. A class of implicit methods for ordinary differential equations. Lecture Notes in Mathematics, 506:28–37, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080112/.

Butkovic:1978:CVM

[But78] Davor Butković. On the completion of vector measures. Lecture Notes in Mathematics, 645:8–15, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069658/.

Balcar:1977:RSB

[BV77] Bohuslav Balcar and Peter Vojtás. Refining systems on Boolean algebras. Lecture Notes in Mathematics, 619:45–58, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067641/.

Butzer:1975:AFD

[BW75] Paul L. Butzer and Ursula Westphal. An access to fractional differentiation via fractional difference quotients. Lecture Notes in Mathematics, 457:116–145, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067101/.

Beck:1976:CSL

[BW76] Anatole Beck and Peter Warren. Counterexamples to strong laws of large numbers for Banach space-valued random variables. Lecture Notes in Mathematics, 532:1–14, 1976. CODEN LN-MAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080168/.

Baudisch:1977:LAT

[BW77a] Andreas Baudisch and Martin Weese. The Lindenbaum-algebra of the theory of well-orders and Abelian groups with the quantifier Q_{α}. Lecture Notes in Mathematics, 619:59–73, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067642/.

Beaman:1977:SRS

[BW77b] I. R. Beaman and W. D. Wallis. On skew room squares. Lecture Notes in Mathematics, 622:61–63, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069181/.

Bigard:1977:ARF

[BWK77a] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Anneaux reticules. (French) [Reticle rings]. Lecture Notes in Mathematics, 608:149–171, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067013/.

Bigard:1977:ARP

[BWK77b] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Anneaux reticules produits sous-directs d’Anneaux totalement ordonnés. (French) []. Lecture Notes in Mathematics, 608:172–199, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067014/.

Bigard:1977:EAF

[BWK77c] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Extensions archimediennes. (French) [Archimedean extensions]. Lecture Notes in Mathematics, 608:90–108, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067010/.

Bigard:1977:GAG

[BWK77d] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Groupes archimediens et groupes complets. (French) []. Lecture Notes in Mathematics, 608:224–240, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067016/.

Bigard:1977:GCD

[BWK77e] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Groupes complètement distributifs et groupes values-finis. (French) []. Lecture Notes in Mathematics, 608:109–127, 1977. CODEN LN-MAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067011/.

Bigard:1977:GHA

[BWK77f] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Groupes hyper-archimediens et sommes directes de groupes réels. (French) []. Lecture Notes in Mathematics, 608:281–295, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067019/.

Bigard:1977:SRP

[BWK77g] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Le spectre et la représentation par des sections dans des faisceaux. (French)

[]. Lecture Notes in Mathematics, 608:200–223, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067015/.

Bigard:1977:NFF

[BWK77h] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Notions fondamentales. (French) [Fundamental concepts]. Lecture Notes in Mathematics, 608:14–35, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067006/.

Bigard:1977:OAA

[BWK77i] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Orthomorphismes et f-anneaux archimediens. (French) []. Lecture Notes in Mathematics, 608:241–257, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067017/.

Bigard:1977:PFPb

[BWK77j] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Polaires. (French) [Polar]. Lecture Notes in Mathematics, 608:52–68, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067008/.

Bigard:1977:PFPa

[BWK77k] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Préliminaires. (French) [Preliminaries]. Lecture Notes in Mathematics, 608:1– 13, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067005/.

Bigard:1977:PLO

[BWK77l] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Propriétés liées à l’Orthogonalité. (French) [Properties related to orthogonality]. Lecture Notes in Mathematics, 608:128–148, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067012/.

Bigard:1977:RGR

[BWK77m] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Représentation des groupes réticules. (French) []. Lecture Notes in Mathematics, 608:69–89, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067009/.

Bigard:1977:RPF

[BWK77n] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Représentation par des fonctions numériques continues. (French) []. Lecture Notes in Mathematics, 608:258–280, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067018/.

Bigard:1977:SGS

[BWK77o] Alain Bigard, Samuel Wolfenstein, and Klaus Keimel. Sousgroupes solides. (French) [Solid subgroups]. Lecture Notes in Mathematics, 608:36–51, 1977. CODEN LNMAA2. ISBN 3-540-08436-3 (print), 3-540-37111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067007/.

Baker:1977:IOP

[BY77] I. N. Baker and Chung-Chun Yang. An infinite order periodic entire function which is prime. Lecture Notes in Mathematics, 599: 7–10, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096818/.

Byczkowski:1978:SRC

[Byc78] T. Byczkowski. Some results concerning Gaussian measures on metric linear spaces. Lecture Notes in Mathematics, 656:1– 16, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068805/.

Byczkowski:1979:GMM

[Byc79] T. Byczkowski. Gaussian measures on metric abelian groups. Lecture Notes in Mathematics, 706:41–53, 1979. CODEN LN-MAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063113/.

Beyer:1976:PEU

[BZ76] K. Beyer and E. Zeidler. Proof of existence and uniqueness of tidal waves with general vorticity distributions. Lecture Notes in Mathematics, 503:193–195, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088755/.

Baouendi:1978:UCT

[BZ78] M. S. Baouendi and E. C. Zachmanoglou. Unique continuation theorems for solutions of partial differential equations and inequalities. Lecture Notes in Mathematics, 660:9–15, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082290/.

Cairoli:1978:RIP

[Cai78] R. Cairoli. Une représentation intégrale pour les martingales fortes. (French) []. Lecture Notes in Mathematics, 649:162– 169, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064604/.

Cairoli:1979:CMI

[Cai79] R. Cairoli. Sur la convergence des martingales indexées par N×N. (French) []. Lecture Notes in Mathematics, 721:162–173, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070859/.

Camacho:1975:PBT

[Cam75] C. Camacho. A Poincaré–Bendixson theorem for R^{2}-actions. Lecture Notes in Mathematics, 468:12–13, 1975. CODEN LN-MAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082580.

Camacho:1978:SSF

[Cam78a] César Camacho. Structural stability of foliations with singularities. Lecture Notes in Mathematics, 652:128–137, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063508/.

Campana:1978:VFV

[Cam78b] Frédéric Campana. Variations de fibrés vectoriels applications aux espaces de sections et aux diviseurs. (French) []. Lecture Notes in Mathematics, 670:312–336, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064406/.

Cannon:1975:TCL

[Can75a] J. W. Cannon. Taming cell-like embedding relations. Lecture Notes in Mathematics, 438:66–118, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066108/.

Cantrell:1975:SNC

[Can75b] J. C. Cantrell. Some non-continuous functions and pseudoflattenings. Lecture Notes in Mathematics, 438:119–121, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066109/.

Capriz:1977:VTA

[Cap77] G. Capriz. Variational techniques for the analysis of a lubrication problem. Lecture Notes in Mathematics, 606:47–55, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064455/.

Capriz:1979:BSS

[Cap79] G. Capriz. On the branching of solutions and Signorini’s perturbation procedure in elasticity. Lecture Notes in Mathematics, 703:54–63, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067256/.

Carmona:1975:FHC

[Car75] Jacques Carmona. Sur les fonctions c_{w} de Harish-Chandra. (French) []. Lecture Notes in Mathematics, 466:26–30, 1975. CO-DEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082194/.

Carpenter:1976:NIE

[Car76a] Gail A. Carpenter. Nerve impulse equations. Lecture Notes in Mathematics, 525:58–76, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077843/.

Cartier:1976:VDR

[Car76b] Pierre Cartier. Vecteurs différentiables dans les représentations unitaires des groupes de Lie. (French) []. Lecture Notes in Mathematics, 514:20–34, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080055/.

Carmona:1977:IPS

[Car77a] Jacques Carmona. On irreducibility of the principal series. Lecture Notes in Mathematics, 587:1–31, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087913/.

Cartier:1977:RGR

[Car77b] Pierre Cartier. Les représentations des groupes réductifs p-adiques et leurs caractères. (French) []. Lecture Notes in Mathematics, 567: 1–22, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096058/.

Carmeli:1978:CCY

[Car78a] M. Carmeli. Classification of classical Yang–Mills fields. Lecture Notes in Mathematics, 676:105–149, 1978. CODEN LN-MAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063667/.

Carmona:1978:TPG

[Car78b] René Carmona. Tensor product of Gaussian measures. Lecture Notes in Mathematics, 644:96–124, 1978. CODEN LN-MAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066843/.

Cartier:1978:SES

[Car78c] Pierre Cartier. Spectre de l’équation de Schrödinger, application à la stabilité de la matière [d’après J. Lebowitz, E. Lieb, B. Simon et W. Thirring]. (French) []. Lecture Notes in Mathematics, 677:88–104, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070756/.

Caraman:1979:ACM

[Car79a] Petru Caraman. About capacities and moduli in infinitedimensional spaces. Lecture Notes in Mathematics, 747:73–91, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063960/.

Caraman:1979:EES

[Car79b] Petru Caraman. Estimate of exceptional sets for quasiconformal mappings in space. Lecture Notes in Mathematics, 743:66– 107, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079484/.

Carlsson:1979:DSG

[Car79c] Gunnar Carlsson. Desuspension in the symmetric L-groups. Lecture Notes in Mathematics, 763:175–197, 1979. CODEN LN-MAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088086/.

Carmona:1979:BSV

[Car79d] René Carmona. Banach space valued Gaussian processes. Lecture Notes in Mathematics, 709:67–73, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071948/.

Cartier:1979:LCF

[Car79e] Pierre Cartier. Logique, catégories et faisceaux [d’après F. Lawvere et M. Tierney]. (French) []. Lecture Notes in Mathematics, 710:123–146, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069976/.

Cathelineau:1975:CDC

[Cat75] J. L. Cathelineau. d′′-cohomologie du classifiant d’un groupe complexe. (French) []. Lecture Notes in Mathematics, 484:201– 223, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082153/.

Catanese:1979:STP

[Cat79] Fabrizio Catanese. Surfaces with K^{2} = p_{g} = 1 and their period mapping. Lecture Notes in Mathematics, 732:1–29, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066634/.

Caubet:1975:RBM

[Cau75] Jean-Pierre Caubet. Relativistic Brownian motion. Lecture Notes in Mathematics, 451:113–142, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068583/.

Caubet:1976:DIS

[Cau76a] Jean-Pierre Caubet. Différentielles et intégrales stochastiques. (French) []. Lecture Notes in Mathematics, 559:67–95, 1976. CO-DEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0089083/.

Caubet:1976:D

[Cau76b] Jean-Pierre Caubet. Diffusions. Lecture Notes in Mathematics, 559:96–131, 1976. CODEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089084/.

Caubet:1976:ELH

[Cau76c] Jean-Pierre Caubet. Équations de Lagrange et de Hamilton. (French) []. Lecture Notes in Mathematics, 559:180–193, 1976. CODEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089086/.

Caubet:1976:I

[Cau76d] Jean-Pierre Caubet. Introduction. Lecture Notes in Mathematics, 559:1–18, 1976. CODEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089080/.

Caubet:1976:MBRa

[Cau76e] Jean-Pierre Caubet. Le mouvement brownien relativiste. (French) [Relativistic Brownian motion]. Lecture Notes in Mathematics, 559:132–179, 1976. CODEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089085/.

Caubet:1976:MBRb

[Cau76f] Jean-Pierre Caubet. Le Mouvement Brownien Relativiste. (French) [Relativistic Brownian Motion], volume 559 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 212 pp. LCCN QA3

.L28 no. 559. URL http://link.springer.com/book/10.1007/
BFb0089079;   http://www.springerlink.com/content/978-
3-540-37538-8.

Caubet:1976:MWM

[Cau76g] Jean-Pierre Caubet. Mesure de Wiener et mouvement brownien. (French) []. Lecture Notes in Mathematics, 559:53–66, 1976.

CODEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089082/.

Caubet:1976:PEF

[Cau76h] Jean-Pierre Caubet. Probabilités et Ésperance. (French) [Probability and expectation]. Lecture Notes in Mathematics, 559:19– 52, 1976. CODEN LNMAA2. ISBN 3-540-08052-X (print), 3-540-37538-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089081/.

Cazacu:1979:GRI

[Caz79a] Cabiria Andreian Cazacu. On the Grötzsch and Rengel inequalities. Lecture Notes in Mathematics, 747:10–23, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063956/.

Cazacu:1979:SPQ

[Caz79b] Cabiria Andreian Cazacu. Some problems in quasiconformality. Lecture Notes in Mathematics, 743:35–56, 1979. CODEN LN-MAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079482/.

Conze-Berline:1975:CQA

[CB75] Nicole Conze-Berline. Sur certains quotients de l’algèbre enveloppante d’une algèbre de Lie semi-simple. (French) []. Lecture Notes in Mathematics, 466:31–37, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082195/.

Csorgo:1976:WAE

[CB76] M. Csörgő and M. D. Burke. Weak approximations of the empirical process when parameters are estimated. Lecture Notes in Mathematics, 566:1–16, 1976. CODEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096875/.

Christensen:1976:AMG

[CBR76] Mark J. Christensen and A. T. Bharucha-Reid. Algebraic models for Gaussian measures on Banach spaces. Lecture Notes in Mathematics, 526:53–65, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082341/.

Csorgo:1976:ERI

[CC76] Miklós Csörgő and Arthur H. C. Chan. On the Erdős–Rényi increments and the P. Lévy modulus of continuity of a Kiefer process. Lecture Notes in Mathematics, 566:17–32, 1976. CODEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096876/.

Chisholm:1979:PCA

[CC79] J. S. R. Chisholm and A. K. Common. Padé–Chebychev approximants. (French) []. Lecture Notes in Mathematics, 765:1–19, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085572/.

Cazacu:1979:RFS

[CCJS79] Cabiria Andreian Cazacu, Aurel Cornea, Martin Jurchescu, and Ion Suciu, editors. Romanian–Finnish Seminar on Complex Analysis: Proceedings, Bucharest, Romania, June 27–July 2, 1976, volume 743 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 743. URL http:

//link.springer.com/book/10.1007/BFb0079479;   http://
www.springerlink.com/content/978-3-540-34861-0.

Cerf:1975:PCT

[CCMS75] V. G. Cerf, D. D. Cowan, R. C. Mullin, and R. G. Stanton. A partial census of trivalent generalized Moore networks. Lecture Notes in Mathematics, 452:1–27, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069540/.

Cerezo:1975:IHF

[CCP75] A. Cerezo, J. Chazarain, and A. Piriou. Introduction aux hyperfonctions. (French) []. Lecture Notes in Mathematics, 449:1– 53, 1975. CODEN LNMAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062914/.

Colares:1977:MIP

[CdC77] A. G. Colares and M. P. do Carmo. On minimal immersions with parallel normal curvature tensor. Lecture Notes in Mathematics, 597:104–113, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085350/.

Croc:1978:COP

[CDI78] Elisabeth Croc, Yves Dermenjian, and Viorel Iftimie. Une classe d’opérateurs pseudo-différentiels partiellement hypoelliptiqueanalytiques. (French) []. Lecture Notes in Mathematics, 660: 59–65, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082294/.

Carmona:1975:NCH

[CDV75] Prof. Jacques Carmona, Prof. Jaques Dixmier, and Prof. Michèle Vergne, editors. Non-Commutative Harmonic Analysis: Actes du Colloque d’Analyse Harmonique Non Commutative, Marseille– Luminy, 1 au 5 Juillet 1974, volume 466 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 466. URL http:

//link.springer.com/book/10.1007/BFb0082191;   http://
www.springerlink.com/content/978-3-540-37524-1.

Cenzer:1976:IDP

[Cen76] Douglas Cenzer. Inductive definitions: Positive and monotone. Lecture Notes in Mathematics, 537:51–63, 1976. CODEN LN-MAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096893/.

Camillo:1979:NLR

[CF79] V. P. Camillo and K. R. Fuller. A note on Loewy rings and chain conditions on primitive ideals. Lecture Notes in Mathematics, 700:75–86, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063461/.

Cairoli:1979:ACS

[CG79a] R. Cairoli and J.-P. Gabriel. Arrêt de certaines suites multiples de variables aléatoires indépendantes. (French) []. Lecture Notes in Mathematics, 721:174–198, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070860/.

Chen:1979:WPA

[CG79b] Goong Chen and Ronald Grimmer. Well-posedness and approximations of linear Volterra integrodifferential equations in Banach spaces. Lecture Notes in Mathematics, 737:83–87, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064497/.

Coleff:1978:CIFa

[CH78a] Nicolas R. Coleff and Miguel E. Herrera. Chapitre III. (French) [Chapter III]. Lecture Notes in Mathematics, 633:116–157, 1978. CODEN LNMAA2. ISBN 3-540-08651-X (print), 3-540-35935-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068352/.

Coleff:1978:CIFb

[CH78b] Nicolas R. Coleff and Miguel E. Herrera. Chapitre IV. (French) [Chapter IV]. Lecture Notes in Mathematics, 633:158–204, 1978. CODEN LNMAA2. ISBN 3-540-08651-X (print), 3-540-35935-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068353/.

Coleff:1978:ERP

[CH78c] Nicolas R. Coleff and Miguel E. Herrera. Enoncé des résultats principaux. (French) []. Lecture Notes in Mathematics, 633:1–53, 1978. CODEN LNMAA2. ISBN 3-540-08651-X (print), 3-540-35935-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068350/.

Coleff:1978:CRA

[CH78d] Nicolas R. Coleff and Miguel E. Herrera. Les Courants Résiduels Associés à une Forme Méromorphe. (French) [Residual currents associated with a meromorphic form], volume 633 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08651-X (print), 3-540-35935-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 211 pp. LCCN QA3

.L28 no. 633. URL http://link.springer.com/book/10.1007/
BFb0068349;   http://www.springerlink.com/content/978-
3-540-35935-7.

Coleff:1978:RMC

[CH78e] Nicolas R. Coleff and Miguel E. Herrera. Les résidus multiples dans les croisements normaux. (French) []. Lecture Notes in Mathematics, 633:54–115, 1978. CODEN LNMAA2. ISBN 3-540-08651-X (print), 3-540-35935-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068351/.

Chakerian:1975:CSC

[Cha75a] Gulbank D. Chakerian. Covering space with convex bodies. Lecture Notes in Mathematics, 490:188–193, 1975. CODEN LN-MAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081141/.

Chapman:1975:HHH

[Cha75b] T. A. Chapman. Homotopic homeomorphisms of Hilbert cube manifolds. Lecture Notes in Mathematics, 438:122–136, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066110/.

Chazarain:1975:FIO

[Cha75c] Prof. Jacques Chazarain, editor. Fourier Integral Operators and Partial Differential Equations: Colloque International, Université de Nice, 1974, volume 459 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 459. URL http:

//link.springer.com/book/10.1007/BFb0074188;   http://
www.springerlink.com/content/978-3-540-37521-0.

Chase:1976:VWB

[Cha76a] Stephen U. Chase. On a variant of the Witt and Brauer groups. Lecture Notes in Mathematics, 549:148–187, 1976. CODEN LN-MAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077341/.

Chatterji:1976:DM

[Cha76b] S. D. Chatterji. Differentiation of measures. Lecture Notes in Mathematics, 541:173–179, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081050/.

Chatterji:1976:VVM

[Cha76c] S. D. Chatterji. Vector-valued martingales and their applications. Lecture Notes in Mathematics, 526:33–51, 1976. CODEN LN-MAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082340/.

Chavent:1976:NFD

[Cha76d] G. Chavent. A new formulation of diphasic incompressible flows in porous media. Lecture Notes in Mathematics, 503:258–270, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088761/.

Chazarain:1976:SOE

[Cha76e] Jacques Chazarain. Spectre des opérateurs elliptiques et flots hamiltoniens. (French) []. Lecture Notes in Mathematics, 514: 111–123, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080061/.

Charbonnel:1977:FPP

[Cha77] J. Y. Charbonnel. La formule de Plancherel pour un groupe de Lie résoluble connexe. (French) []. Lecture Notes in Mathematics, 587:32–76, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087914/.

Chadam:1978:NPD

[Cha78a] J. M. Chadam, editor. Nonlinear Partial Differential Equations and Applications: Proceedings of a Special Seminar Held at Indiana University, 1976–1977, volume 648 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 648. URL http:

//link.springer.com/book/10.1007/BFb0066404;   http://
www.springerlink.com/content/978-3-540-35868-8.

Chatterji:1978:SAC

[Cha78b] S. D. Chatterji. Singularity and absolute continuity of measures in infinite dimensional spaces. Lecture Notes in Mathematics, 656: 17–23, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068806/.

Chatterji:1978:SPC

[Cha78c] S. D. Chatterji. Stochastic processes and commutation relationships. Lecture Notes in Mathematics, 645:16–26, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069659/.

Chang:1979:GSB

[Cha79a] Weita Chang. Global solvability of bi-invariant differential operators on solvable Lie groups. Lecture Notes in Mathematics, 728: 8–16, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063335/.

Chatterji:1979:PSS

[Cha79b] S. D. Chatterji. Le principe de sous-suites dans les espaces de Banach. (French) []. Lecture Notes in Mathematics, 721:4–21, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070848/.

Chatterji:1979:RNP

[Cha79c] S. D. Chatterji. The Radon–Nikodým property. Lecture Notes in Mathematics, 709:75–86, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071949/.

Chenciner:1975:QPG

[Che75a] A. Chenciner. Quelques problémes globaux sur les fonctions de Morse. (French) []. Lecture Notes in Mathematics, 484:1–7, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082140/.

Cherlin:1975:IIN

[Che75b] Gregory L. Cherlin. Ideals of integers in nonstandard number fields. Lecture Notes in Mathematics, 498:60–90, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080975/.

Chen:1976:RBB

[Che76a] Hsing-Hen Chen. Relation between Bäcklund transformations and inverse scattering problems. Lecture Notes in Mathematics, 515: 241–252, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081171/.

Cherlin:1976:CF

[Che76b] Greg Cherlin. ℵ_{1}-categorical fields. Lecture Notes in Mathematics, 521:207–222, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079574/.

Cherlin:1976:AKE

[Che76c] Greg Cherlin. The Ax–Kochen–Ershov transfer principle: (Diophantine problems over local fields). Lecture Notes in Mathematics, 521:32–66, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079569/.

Cherlin:1976:BMT

[Che76d] Greg Cherlin. Basic model theory. Lecture Notes in Mathematics, 521:4–20, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079567/.

Cherlin:1976:CTA

[Che76e] Greg Cherlin. Complete theories of abelian groups. Lecture Notes in Mathematics, 521:185–206, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079573/.

Cherlin:1976:ECD

[Che76f] Greg Cherlin. Existentially complete division rings. Lecture Notes in Mathematics, 521:108–155, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079571/.

Cherlin:1976:ECM

[Che76g] Greg Cherlin. Existentially complete modules. Lecture Notes in Mathematics, 521:156–184, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079572/.

Cherlin:1976:ECS

[Che76h] Greg Cherlin. Existentially complete structures. Lecture Notes in Mathematics, 521:67–107, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079570/.

Cherlin:1976:I

[Che76i] Greg Cherlin. Introduction. Lecture Notes in Mathematics, 521: 1–3, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079566/.

Cherlin:1976:MTA

[Che76j] Greg Cherlin. Model Theoretic Algebra Selected Topics, volume 521 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 234 pp. LCCN QA3 .L28 no. 521. URL http:

//link.springer.com/book/10.1007/BFb0079565;   http://
www.springerlink.com/content/978-3-540-38228-7.

Cherlin:1976:TTA

[Che76k] Greg Cherlin. Transfer theorems in algebra. Lecture Notes in Mathematics, 521:21–31, 1976. CODEN LNMAA2. ISBN 3-540-07696-4 (print), 3-540-38228-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079568/.

Chen:1978:ECG

[Che78a] C. C. Chen. On the enumeration of certain graceful graphs. Lecture Notes in Mathematics, 686:111–115, 1978. CODEN LN-MAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062523/.

Chetayev:1978:WPO

[Che78b] A. N. Chetayev. On walks over a partially ordered set (some inequalities for conditional probabilities). Lecture Notes in Mathematics, 653:149–172, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070091/.

Chevet:1978:QNR

[Che78c] Simone Chevet. Quelques nouveaux résultats sur les mesures cylindriques. (French) []. Lecture Notes in Mathematics, 644:125– 158, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066844/.

Cherenack:1979:IHS

[Che79] Paul Cherenack. Internal hom-sets in an extension of affine schemes over a field. Lecture Notes in Mathematics, 732:30– 42, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066635/.

Chi:1976:RNT

[Chi76a] G. Y. H. Chi. On the Radon–Nikodým theorem in locally convex spaces. Lecture Notes in Mathematics, 541:199–210, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081053/.

Childs:1976:BGS

[Chi76b] Lindsay N. Childs. On Brauer groups of some normal local rings. Lecture Notes in Mathematics, 549:1–15, 1976. CODEN LN-MAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077329/.

Chipot:1976:CSP

[Chi76c] Michel Chipot. Convergence of solutions in problems of elastic plastic torsion of cylindrical bars. Lecture Notes in Mathematics, 503:271–278, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088762/.

Chidzey:1978:FS

[Chi78a] Keith Chidzey. Fixing subgraphs of K_{m,n}. Lecture Notes in Mathematics, 686:116–125, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062524/.

Chisholm:1978:MAB

[Chi78b] J. S. R. Chisholm. Multivariate approximants with branch points. Lecture Notes in Mathematics, 630:32–40, 1978. CODEN LN-MAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067694/.

Chou:1975:ISA

[Cho75a] Ching-Sung Chou. Les inégalités des surmartingales d’après A. M. Garsia. (French) []. Lecture Notes in Mathematics, 465:206– 212, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102991/.

Chou:1975:MGT

[Cho75b] Ching-Sung Chou. Les méthodes d’A. Garsia en théorie des martingales. Extensions au cas continu. (French) []. Lecture Notes in Mathematics, 465:213–225, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0102992/.

Chollet:1976:BZS

[Cho76] Anne-Marie Chollet. Boundary zero-sets of A^{∞} functions on strictly pseudo-convex domains. Lecture Notes in Mathematics, 512:17–23, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080020/.

Chorin:1977:LIT

[Cho77a] Alexandre Chorin. Lecture II theories of turbulence. Lecture Notes in Mathematics, 615:36–47, 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068359/.

Chou:1977:DHM

[Cho77b] Chin-Cheng Chou. Distributions, hyperfonctions et le microspectre analytique. (French) []. Lecture Notes in Mathematics, 578: 172–182, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091468/.

Chowla:1977:SEC

[Cho77c] S. Chowla. L-series and elliptic curves. Lecture Notes in Mathematics, 626:1–42, 1977. CODEN LNMAA2. ISBN 3-540-08529-7 (print), 3-540-35985-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063063/.

Choda:1978:CBN

[Cho78a] Hisashi Choda. Correspondences between von Neumann algebras and discrete automorphism groups. Lecture Notes in Mathematics, 650:41–54, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067388/.

Chou:1978:EAC

[Cho78b] Ching-Sung Chou. Extension au cas continu d’un théorème de L. E. Dubins. (French) []. Lecture Notes in Mathematics, 649: 132–133, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064600.

Christensen:1976:SPE

[Chr76] Jens Peter Reus Christensen. Submeasures and the problem on the existence of control measures. Lecture Notes in Mathematics, 541:49–51, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081040/.

Christensen:1978:SRR

[Chr78] Jens Peter Reus Christensen. Some results with relation to the control measure problem. Lecture Notes in Mathematics, 645: 27–34, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069660/.

Chu:1976:SRB

[Chu76] Flora Ying Fun Chu. Stimulated Raman and Brillouin scattering and the inverse method. Lecture Notes in Mathematics, 515: 25–39, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081161/.

Chudacek:1977:TPA

[Chu77a] Jaroslav Chudáček. Topological problems in alternative set theory. Lecture Notes in Mathematics, 619:119–133, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067646/.

Chung:1977:PNB

[Chu77b] Kai Lai Chung. Pedagogic notes on the barrier theorem. Lecture Notes in Mathematics, 581:27–33, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087184/.

Chudnovsky:1979:TN

[Chu79] G. V. Chudnovsky. Transcendental numbers. Lecture Notes in Mathematics, 751:45–69, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062701/.

Cichon:1977:BPB

[Cic77] Jacek Cichoń. On the Baire property of Boolean algebras. Lecture Notes in Mathematics, 619:135–141, 1977. CODEN LN-MAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067647/.

Ciesielski:1975:LBM

[Cie75a] Z. Ciesielski. On Lévy’s Brownian motion with several-dimensional time. Lecture Notes in Mathematics, 472:29–56, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081944/.

Ciesielski:1975:SSP

[Cie75b] Z. Ciesielski. Stochastic systems of particles. Lecture Notes in Mathematics, 472:1–28, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081943/.

Cigler:1976:TPF

[Cig76] J. Cigler. Tensor products of functors on categories of Banach spaces. Lecture Notes in Mathematics, 540:164–187, 1976. CO-DEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080858/.

Cima:1977:TCO

[Cim77] Joseph A. Cima. A theorem on composition operators. Lecture Notes in Mathematics, 604:21–24, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069200/.

Cirelson:1978:RSI

[Cir78] B. S. Cirel’son. Reliable storage of information in a system of unreliable components with local interactions. Lecture Notes in Mathematics, 653:15–30, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070081/.

Cirelson:1976:NGS

[CIS76] B. S. Cirel’son, I. A. Ibragimov, and V. N. Sudakov. Norms of Gaussian sample functions. Lecture Notes in Mathematics, 550: 20–41, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077482/.

Caristi:1975:GFP

[CK75] James Caristi and William A. Kirk. Geometric fixed point theory and inwardness conditions. Lecture Notes in Mathematics, 490: 74–83, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081133/.

Conze:1976:TEA

[CK76] Jean-Pierre Conze and Michael S. Keane, editors. Théorie Ergodique: Actes des Journées Ergodiques, Rennes 1973/ 1974, volume 532 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 532. URL http:

//link.springer.com/book/10.1007/BFb0080167;   http://
www.springerlink.com/content/978-3-540-38217-1.

Cornea:1975:LC

[CL75a] Dr. Aurel Cornea and Dr. Gabriela Licea. δ-lattice cones. Lecture Notes in Mathematics, 494:1–35, 1975. CODEN LNMAA2. ISBN 3-540-07531-3 (print), 3-540-38019-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087301/.

Cornea:1975:DPH

[CL75b] Dr. Aurel Cornea and Dr. Gabriela Licea. Domination principle and Hunt’s theorem. Lecture Notes in Mathematics, 494:85– 119, 1975. CODEN LNMAA2. ISBN 3-540-07531-3 (print), 3-540-38019-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087304/.

Cornea:1975:E

[CL75c] Dr. Aurel Cornea and Dr. Gabriela Licea. The energy. Lecture Notes in Mathematics, 494:120–134, 1975. CODEN LN-MAA2. ISBN 3-540-07531-3 (print), 3-540-38019-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087305/.

Cornea:1975:OPR

[CL75d] Dr. Aurel Cornea and Dr. Gabriela Licea. Order and Potential Resolvent Families of Kernels, volume 494 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07531-3 (print), 3-540-38019-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 154 pp. LCCN QA3

.L28 no. 494. URL http://link.springer.com/book/10.1007/
BFb0087300;   http://www.springerlink.com/content/978-
3-540-38019-1.

Cornea:1975:PSE

[CL75e] Dr. Aurel Cornea and Dr. Gabriela Licea. Positive supermartingales as excessive elements. Lecture Notes in Mathematics, 494: 135–151, 1975. CODEN LNMAA2. ISBN 3-540-07531-3 (print), 3-540-38019-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087306/.

Cornea:1975:PTA

[CL75f] Dr. Aurel Cornea and Dr. Gabriela Licea. Potential theory associated with a resolvent family of kernels. Lecture Notes in Mathematics, 494:46–84, 1975. CODEN LNMAA2. ISBN 3-540-07531-3 (print), 3-540-38019-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087303/.

Cornea:1975:PTS

[CL75g] Dr. Aurel Cornea and Dr. Gabriela Licea. The potential theory of a single kernel. Lecture Notes in Mathematics, 494:36–45, 1975. CODEN LNMAA2. ISBN 3-540-07531-3 (print), 3-540-38019-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087302/.

Carasso:1976:AGP

[CL76a] C. Carasso and P. J. Laurent. Un algorithme général pour L’Approximation au sens de Tchebycheff de fonctions bornées sur un ensemble quelconque. (French) []. Lecture Notes in Mathematics, 556:99–121, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087399/.

Crepel:1976:TRP

[CL76b] P. Crépel and J. Lacroix. Théorèmes de renouvellement pour les marches aléatoires sur les groupes localement compacts. (French) []. Lecture Notes in Mathematics, 563:27–42, 1976. CODEN LN-MAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087570/.

Clancey:1979:CRS

[Cla79a] Kevin Clancey. Concrete realizations of seminormal operators. Lecture Notes in Mathematics, 742:37–62, 1979. CODEN LN-MAA2. ISBN 3-540-09547-0 (print), 3-540-34865-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065644/.

Clancey:1979:PDP

[Cla79b] Kevin Clancey. Perturbation determinants and the phase shift. Lecture Notes in Mathematics, 742:79–91, 1979. CODEN LN-MAA2. ISBN 3-540-09547-0 (print), 3-540-34865-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065646/.

Clancey:1979:PF

[Cla79c] Kevin Clancey. The principal function. Lecture Notes in Mathematics, 742:92–118, 1979. CODEN LNMAA2. ISBN 3-540-09547-0 (print), 3-540-34865-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065647/.

Clancey:1979:SOa

[Cla79d] Kevin Clancey. Seminormal operators. Lecture Notes in Mathematics, 742:1–36, 1979. CODEN LNMAA2. ISBN 3-540-09547-0 (print), 3-540-34865-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065643/.

Clancey:1979:SOb

[Cla79e] Kevin Clancey. Seminormal Operators, volume 742 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09547-0 (print), 3-540-34865-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 125 pp. LCCN QA3

.L28 no. 742. URL http://link.springer.com/book/10.1007/
BFb0065642;   http://www.springerlink.com/content/978-
3-540-34865-8.

Clancey:1979:TBR

[Cla79f] Kevin Clancey. Two basic results on seminormal operators. Lecture Notes in Mathematics, 742:63–78, 1979. CODEN LN-MAA2. ISBN 3-540-09547-0 (print), 3-540-34865-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065645/.

Clarysse:1979:RPC

[Cla79g] T. H. Clarysse. Rational predictor-corrector methods for nonlinear Volterra integral equations of the second kind. Lecture Notes in Mathematics, 765:278–294, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085586/.

Clauwens:1979:PFS

[Cla79h] F. J.-B. J. Clauwens. Product formulae for surgery obstructions. Lecture Notes in Mathematics, 763:198–211, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088087/.

Clerc:1975:QTC

[Cle75] Jean-Louis Clerc. Quelques théorèmes de convergence pour l’analyse harmonique de SU(2). (French) []. Lecture Notes in Mathematics, 497:16–25, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078011/.

Clement:1979:AVE

[Clé79] Ph. Clément. On abstract Volterra equations with kernel of positive resolvent. Lecture Notes in Mathematics, 737:45–49, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064494/.

Clifford:1975:HPR

[Cli75] H. T. Clifford. Host-parasite relationships. Lecture Notes in Mathematics, 452:79–82, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069545/.

Cohen:1976:HIL

[CLM76] Frederick R. Cohen, Thomas J. Lada, and J. Peter May. The Homology of Iterated Loop Spaces, volume 533 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07984-X (print), 3-540-37985-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 490 pp. LCCN QA3

.L28 no. 533. URL http://link.springer.com/book/10.1007/
BFb0080464;   http://www.springerlink.com/content/978-
3-540-37985-0.

Clozel:1979:BCG

[Clo79] L. Clozel. “Base change” géométrique: Relèvement de la série principale de GL(n, C/R). (French) []. Lecture Notes in Mathematics, 728:17–41, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063336/.

Chou:1975:RMC

[CM75] Ching-Sung Chou and P.-A. Meyer. Sur la représentation des martingales comme intégrales stochastiques dans les processus ponctuels. (French) []. Lecture Notes in Mathematics, 465:226– 236, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102993/.

Cooper:1976:DCL

[CM76] J. B. Cooper and P. Michor. Duality of compactological and locally compact groups. Lecture Notes in Mathematics, 540:188– 207, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080859/.

Chang:1977:SAB

[CM77] S. Y. Chang and D. E. Marshall. Some algebras of bounded analytic functions containing the disk algebra. Lecture Notes in Mathematics, 604:12–20, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069199/.

Capobianco:1978:MGH

[CM78a] Michael Capobianco and John Molluzzo. Mixed graphs of homomorphisms. Lecture Notes in Mathematics, 642:116–120, 1978.

CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070368/.

Carlsson:1978:TRO

[CM78b] G. Carlsson and R. James Milgram. Torsion with rings for orders and finite groups. Lecture Notes in Mathematics, 657:85– 105, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069228/.

Carlsson:1979:SOG

[CM79] G. Carlsson and R. James Milgram. The structure of odd Lgroups. Lecture Notes in Mathematics, 741:1–72, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062134/.

Cohen:1979:DLS

[CMN79] F. R. Cohen, J. C. Moore, and J. A. Neisendorfer. Decompositions of loop spaces and applications to exponents. Lecture Notes in Mathematics, 763:1–12, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088074/.

Chow:1979:HML

[CMPY79] Shui-Nee Chow, John Mallet-Paret, and James A. Yorke. A homotopy method for locating all zeros of a system of polynomials. Lecture Notes in Mathematics, 730:77–88, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064312/.

Cooper:1978:HE

[CMW78] Joan Cooper, James Milas, and W. D. Wallis. Hadamard equivalence. Lecture Notes in Mathematics, 686:126–135, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062525/.

Crossley:1975:CF

[CN75a] J. N. Crossley and Anil Nerode. Combinatorial functors. Lecture Notes in Mathematics, 453:1–21, 1975. CODEN LNMAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064869/.

Crossley:1975:SF

[CN75b] J. N. Crossley and Anil Nerode. Sound functors. Lecture Notes in Mathematics, 500:26–43, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079545/.

Coroi-Nedelcu:1976:TLE

[CN76] Mariana Coroi-Nedelcu. On the theory of linear equations with spatial derivatives. Lecture Notes in Mathematics, 561:150– 157, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087633/.

Camacho:1977:OPD

[CN77] César Camacho and Alcides Lins Neto. Orbit preserving diffeomorphisms and the stability of Lie group actions and singular foliations. Lecture Notes in Mathematics, 597:82–103, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085349/.

Cohen:1977:DEF

[CO77] H. Cohen and J. Oesterlé. Dimensions des espaces de formes modulaires. (French) []. Lecture Notes in Mathematics, 627:69–78, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065297/.

Cailliez:1979:SGP

[CO79a] J. Cailliez and J. Oberdoerffer. Les sous-groupes paraboliques de SU(p, q) et Sp(n, R) et applications à l’étude des représentations. (French) []. Lecture Notes in Mathematics, 739:51–106, 1979. CO-DEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062488/.

Cailliez:1979:SCA

[CO79b] J. Cailliez and J. Oberdoerffer. Séries complémentaires associées à certains paraboliques de SU(n, n). (French) []. Lecture Notes in Mathematics, 739:107–121, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062489/.

Cockayne:1978:DUG

[Coc78] E. J. Cockayne. Domination of undirected graphs — a survey. Lecture Notes in Mathematics, 642:141–147, 1978. CODEN LN-MAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070371/.

Coddington:1975:STO

[Cod75] Earl A. Coddington. Spectral theory of ordinary differential operators. Lecture Notes in Mathematics, 448:1–24, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067079/.

Cody:1976:OSD

[Cod76] W. J. Cody. An overview of software development for special functions. Lecture Notes in Mathematics, 506:38–48, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080113/.

Coddington:1979:DSA

[Cod79] E. A. Coddington. Differential subspaces associated with pairs of ordinary differential operators. Lecture Notes in Mathematics, 703:64–72, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067257/.

Coeure:1977:RBF

[Coe77] G. Coeuré. Sur le rayon de bornologie des fonctions holomorphes. (French) []. Lecture Notes in Mathematics, 578:183–194, 1977.

CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091469/.

Cohen:1976:HS

[Coh76a] Fred Cohen. The homology of C_{n+1}-spaces, n ≥ 0. Lecture Notes in Mathematics, 533:207–351, 1976. CODEN LNMAA2. ISBN 3-540-07984-X (print), 3-540-37985-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080467/.

Cohen:1976:H

[Coh76b] Fred Cohen. The homology of SF(n + 1). Lecture Notes in Mathematics, 533:352–398, 1976. CODEN LNMAA2. ISBN 3-540-07984-X (print), 3-540-37985-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080468/.

Cohen:1977:LMF

[Coh77] Henri Cohen. A lifting of modular forms in one variable to Hilbert modular forms in two variables. Lecture Notes in Mathematics, 627:175–196, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065301/.

Cohen:1978:TM

[Coh78] Leon W. Cohen. Topology and measure. Lecture Notes in Mathematics, 664:61–75, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061690/.

Cohn:1979:DEC

[Coh79a] Harvey Cohn. Diophantine equations over C(t) and complex multiplication. Lecture Notes in Mathematics, 751:70–81, 1979. CO-DEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062702/.

Cohn:1979:ASG

[Coh79b] P. M. Cohn. The affine scheme of a general ring. Lecture Notes in Mathematics, 753:197–211, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061819/.

Collatz:1975:DAC

[Col75] L. Collatz. Different applications of convex and nonconvex optimization, especially to differential equations. Lecture Notes in Mathematics, 448:112–125, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067082/.

Collatz:1976:AME

[Col76a] L. Collatz. Approximation methods for expanding operators. Lecture Notes in Mathematics, 506:49–59, 1976. CODEN LN-MAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080114/.

Collatz:1976:BD

[Col76b] L. Collatz. Bifurcation diagrams. Lecture Notes in Mathematics, 564:41–53, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087325/.

Collatz:1976:EAN

[Col76c] Lothar Collatz. Einige Anwendungen der nichtlinearen Approximationstheorie auf Randwertaufgaben. (German) []. Lecture Notes in Mathematics, 556:136–143, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087401/.

Colton:1976:IOI

[Col76d] David Colton. Integral operators and inverse problems in scattering theory. Lecture Notes in Mathematics, 561:17–28, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087622/.

Colton:1976:WTH

[Col76e] David Colton. Walsh’s theorem for the heat equation. Lecture Notes in Mathematics, 564:54–60, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087326/.

Collatz:1978:NTS

[Col78] Lothar Collatz. The numerical treatment of some singular boundary value problems. Lecture Notes in Mathematics, 630:41– 50, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067695/.

Combes:1976:FNT

[Com76] François Combes. Les facteurs de von Neumann de type III. (French) []. Lecture Notes in Mathematics, 514:124–137, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080062/.

Comfort:1977:SRA

[Com77] W. W. Comfort. Some recent applications of ultrafilters to topology. Lecture Notes in Mathematics, 609:34–42, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068670/.

Conlon:1975:FLG

[Con75a] S. B. Conlon. Finite linear p-groups of degree p and the work of G. Szekeres. Lecture Notes in Mathematics, 488:72–78, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081217/.

Conze:1975:STM

[Con75b] J.-P. Conze. Systèmes topologiques et métriques en théorie ergodique. (French) []. Lecture Notes in Mathematics, 480:97– 187, 1975. CODEN LNMAA2. ISBN 3-540-07396-5 (print), 3-540-37600-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080191/.

Conley:1976:NSW

[Con76a] C. Conley. A new statement of Wazewski’s theorem and an example. Lecture Notes in Mathematics, 564:61–71, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087327/.

Constantinescu:1976:PDP

[Con76b] Corneliu Constantinescu. Le problème de Dirichlet pour des systèmes d’équations non-linéaires. (French) []. Lecture Notes in Mathematics, 518:54–73, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080403/.

Conlon:1977:NSP

[Con77] S. B. Conlon. Nonabelian subgroups of prime-power order of classical groups of the same prime degree. Lecture Notes in Mathematics, 573:17–50, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087809/.

Conner:1979:BG

[Con79a] P. E. Conner. Bordism groups. Lecture Notes in Mathematics, 738:3–66, 1979. CODEN LNMAA2. ISBN 3-540-09535-7 (print), 3-540-35032-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063219/.

Conner:1979:DI

[Con79b] P. E. Conner. Differentiable involutions. Lecture Notes in Mathematics, 738:67–122, 1979. CODEN LNMAA2. ISBN 3-540-09535-7 (print), 3-540-35032-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063220/.

Conner:1979:I

[Con79c] P. E. Conner. Introduction. Lecture Notes in Mathematics, 738: 1–2, 1979. CODEN LNMAA2. ISBN 3-540-09535-7 (print), 3-540-35032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063218.

Conner:1979:MOP

[Con79d] P. E. Conner. Maps of odd period. Lecture Notes in Mathematics, 738:123–175, 1979. CODEN LNMAA2. ISBN 3-540-09535-7 (print), 3-540-35032-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063221/.

Conner:1979:DPM

[Con79e] Pierre E. Conner. Differentiable Periodic Maps, volume 738 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1979. CODEN LNMAA2. ISBN 3-540-09535-7 (print), 3-540-35032-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 181 pp. LCCN QA3 .L28 no. 738. URL http:

//link.springer.com/book/10.1007/BFb0063217;   http://
www.springerlink.com/content/978-3-540-35032-3.

Connes:1979:TNC

[Con79f] Alain Connes. Sur la théorie non commutative de l’intégration. (French) []. Lecture Notes in Mathematics, 725:19–143, 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062614/.

Conti:1979:CVP

[Con79g] R. Conti. Control and the van der Pol equation. Lecture Notes in Mathematics, 703:73–80, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067258/.

Cooper:1977:FVG

[Coo77] Yonina S. Cooper. Finite valuated groups. Lecture Notes in Mathematics, 616:124–139, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068192/.

Coppel:1978:CED

[Cop78a] W. A. Coppel. Criteria for an exponential dichotomy. Lecture Notes in Mathematics, 629:47–58, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067786/.

Coppel:1978:DFA

[Cop78b] W. A. Coppel. Dichotomies and functional analysis. Lecture Notes in Mathematics, 629:20–27, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067783/.

Coppel:1978:DLF

[Cop78c] W. A. Coppel. Dichotomies and Lyapunov functions. Lecture Notes in Mathematics, 629:59–66, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067787/.

Coppel:1978:DR

[Cop78d] W. A. Coppel. Dichotomies and reducibility. Lecture Notes in Mathematics, 629:38–46, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067785/.

Coppel:1978:DHE

[Cop78e] W. A. Coppel. Dichotomies and the hull of an equation. Lecture Notes in Mathematics, 629:74–86, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067789/.

Coppel:1978:DST

[Cop78f] W. A. Coppel. Dichotomies in Stability Theory, volume 629 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CO-DEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 98 pp. LCCN QA3 .L28 no. 629. URL http:

//link.springer.com/book/10.1007/BFb0067780;   http://
www.springerlink.com/content/978-3-540-35976-0.

Coppel:1978:EAP

[Cop78g] W. A. Coppel. Equations on R and almost periodic equations. Lecture Notes in Mathematics, 629:67–73, 1978. CODEN LN-MAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067788/.

Coppel:1978:EOD

[Cop78h] W. A. Coppel. Exponential and ordinary dichotomies. Lecture Notes in Mathematics, 629:10–19, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067782/.

Coppel:1978:MCT

[Cop78i] W. A. Coppel, editor. Mathematical Control Theory: Proceedings, Canberra, Australia, August 23–September 2, 1977, volume 680 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 680. URL http://link.springer.com/book/10.1007/
BFb0065310;   http://www.springerlink.com/content/978-
3-540-35714-8.

Coppel:1978:R

[Cop78j] W. A. Coppel. Roughness. Lecture Notes in Mathematics, 629: 28–37, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067784/.

Coppel:1978:S

[Cop78k] W. A. Coppel. Stability. Lecture Notes in Mathematics, 629: 1–9, 1978. CODEN LNMAA2. ISBN 3-540-08536-X (print), 3-540-35976-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067781/.

Cordero:1975:HLF

[Cor75] L. A. Cordero. The horizontal lift of a foliation and its exotic classes. Lecture Notes in Mathematics, 484:192–200, 1975. CO-DEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082152/.

Cordellier:1979:DAE

[Cor79a] F. Cordellier. Démonstration algébrique de l’extension de l’identité de Wynn aux tables de Padé non normales. (French) [Algebraic demonstration of the extension of the identity of Wynn to nonnormal Padé tables]. Lecture Notes in Mathematics, 765:36– 60, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085574/.

Cordellier:1979:RPA

[Cor79b] F. Cordellier. Sur la régularité des procédés δ^{2} d’Aitken et W de Lubkin. (French) []. Lecture Notes in Mathematics, 765:20– 35, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085573/.

Cordes:1979:DRS

[Cor79c] H. O. Cordes. Distributions with rational singularities. Lecture Notes in Mathematics, 756:58–82, 1979. CODEN LNMAA2. ISBN 3-540-09704-X (print), 3-540-38463-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069815/.

Cordes:1979:EBP

[Cor79d] H. O. Cordes. Elliptic boundary problems. Lecture Notes in Mathematics, 756:181–252, 1979. CODEN LNMAA2. ISBN 3-540-09704-X (print), 3-540-38463-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069818/.

Cordes:1979:EPD

[Cor79e] H. O. Cordes. Elliptic Pseudo-Differential Operators — an Abstract Theory, volume 756 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09704-X (print), 3-540-38463-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 331 pp. LCCN QA3

.L28 no. 756. URL http://link.springer.com/book/10.1007/
BFb0069813;   http://www.springerlink.com/content/978-
3-540-38463-2.

Cordes:1979:FGP

[Cor79f] H. O. Cordes. Finitely generated pseudo-differential operators. Lecture Notes in Mathematics, 756:83–126, 1979. CODEN LN-MAA2. ISBN 3-540-09704-X (print), 3-540-38463-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069816/.

Cordes:1979:LCA

[Cor79g] H. O. Cordes. On the Laplace comparison algebras for L^{2}(R^{n}). Lecture Notes in Mathematics, 756:127–180, 1979. CODEN LN-MAA2. ISBN 3-540-09704-X (print), 3-540-38463-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069817/.

Cordes:1979:SDT

[Cor79h] H. O. Cordes. A survey of distribution theory. Lecture Notes in Mathematics, 756:1–57, 1979. CODEN LNMAA2. ISBN 3-540-09704-X (print), 3-540-38463-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069814/.

Cornell:1979:ALC

[Cor79i] Gary Cornell. Abhyankar’s lemma and the class group. Lecture Notes in Mathematics, 751:82–88, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062703/.

Costa:1975:DRB

[Cos75] David G. Costa. L^{p} decay rates, p bit (≤ ∞), and energy decay in nonbicharacteristic cones for first order hyperbolic systems. Lecture Notes in Mathematics, 446:123–143, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070600/.

Coste:1979:LSS

[Cos79] Michel Coste. Localisation, spectra and sheaf representation. Lecture Notes in Mathematics, 753:212–238, 1979. CODEN LN-MAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061820/.

Cottle:1975:MML

[Cot75] Richard W. Cottle. On Minkowski matrices and the linear complementarity problem. Lecture Notes in Mathematics, 477:18– 26, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079164/.

Cottle:1979:NMC

[Cot79] Richard W. Cottle. Numerical methods for complementarity problems in engineering and applied science. Lecture Notes in Mathematics, 704:37–52, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063613/.

Couot:1975:TEE

[Cou75] J. Couot. Théorie ergodique de l’équirépartition. (French) []. Lecture Notes in Mathematics, 475:26–88, 1975. CODEN LN-MAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074257/.

Couchot:1978:SMP

[Cou78] François Couchot. Sous-modules pures et modules de type cofini. (French) []. Lecture Notes in Mathematics, 641:198–208, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064848/.

Couchot:1979:CAC

[Cou79] François Couchot. Classes d’anneaux contenant les V -anneaux et les anneaux absolument plats. (French) []. Lecture Notes in Mathematics, 740:170–183, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071058/.

Cowen:1978:CAT

[Cow78] Carl C. Cowen. Commutants of analytic Toeplitz operators with automorphic symbol. Lecture Notes in Mathematics, 693:71– 75, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064661/.

Cowling:1979:CRU

[Cow79] Michael Cowling. Sur les coéfficients des représentations unitaires des groupes de Lie simples. (French) []. Lecture Notes in Mathematics, 739:132–178, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062491/.

Cox:1978:IBC

[Cox78] M. G. Cox. The incorporation of boundary conditions in spline approximation problems. Lecture Notes in Mathematics, 630:51– 63, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067696/.

Cox:1979:SWE

[Cox79] David A. Cox. Solutions of Weierstrass equations. Lecture Notes in Mathematics, 732:43–59, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066636/.

Carey:1977:ACA

[CP77a] Richard W. Carey and Joel D. Pincus. Almost commuting algebras. Lecture Notes in Mathematics, 575:19–43, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095699/.

Carl:1977:ENO

[CP77b] B. Carl and A. Pietsch. Entropy numbers of operators in Banach spaces. Lecture Notes in Mathematics, 609:21–33, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068669/.

Clerc:1979:TCL

[CR79] J.-L. Clerc and B. Roynette. Un théorème central-limite. (French) []. Lecture Notes in Mathematics, 739:122–131, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062490/.

Cresswell:1975:FMM

[Cre75] Max J. Cresswell. Frames and models in modal logic. Lecture Notes in Mathematics, 450:63–86, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062851/.

Crepel:1976:RMA

[Cré76] Pierre Crépel. Recurrence des marches aléatoires sur les groupes de Lie. (French) [Recurrence of random walks on Lie groups]. Lecture Notes in Mathematics, 532:50–69, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080172/.

Crepel:1979:GAZ

[Cré79] Pierre Crépel. Grenzwertsätze für abhängige Zufallsvariable und Irrfahrten auf Gruppen. (German) []. Lecture Notes in Mathematics, 706:54–66, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063114/.

Crossley:1975:RL

[Cro75a] J. N. Crossley. Reminiscences of logicians. Lecture Notes in Mathematics, 450:1–62, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062850/.

Crossley:1975:ALP

[Cro75b] Prof. John Newsome Crossley, editor. Algebra and Logic: Papers from the 1974 Summer Research Institute of the Australian Mathematical Society, Monash University, Australia, volume 450 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 450. URL http://link.springer.com/book/10.1007/
BFb0062849;   http://www.springerlink.com/content/978-
3-540-37480-0.

Cromme:1976:TAU

[Cro76] Dr. L. Cromme. Zur Tschebyscheff-Approximation bei Ungeichungsnebenbedingungen im Funktionenraum. (German) []. Lecture Notes in Mathematics, 556:144–153, 1976. CODEN LN-MAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087402/.

Collens:1975:CAG

[CS75] R. J. Collens and R. G. Stanton. Computer assisted generation of difference blocks. Lecture Notes in Mathematics, 452:83–89, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069546/.

Calder:1976:HKE

[CS76a] Allan Calder and Jerrold Siegel. Homotopy and Kan extensions. Lecture Notes in Mathematics, 540:152–163, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080857/.

Choksi:1976:SPT

[CS76b] J. R. Choksi and R. R. Simha. Set and point transformations on homogeneous spaces. Lecture Notes in Mathematics, 541:1– 4, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081034/.

Conley:1976:RTW

[CS76c] C. Conley and Joel Smoller. Remarks on traveling wave solutions of non-linear diffusion equations. Lecture Notes in Mathematics, 525:77–89, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077844/.

Cozzens:1976:NRT

[CS76d] Jonh Henry Cozzens and Francis Louis Sandomierski, editors. Noncommutative Ring Theory: Papers presented at the International Conference Held at Kent State University April 4–5, 1975, volume 545 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 545. URL http:

//link.springer.com/book/10.1007/BFb0080303;   http://
www.springerlink.com/content/978-3-540-37983-6.

Conley:1978:IIS

[CS78] C. Conley and Joel Smoller. Isolated invariant sets of parameterized systems of differential equations. Lecture Notes in Mathematics, 668:30–47, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101778/.

Cooper:1979:SSV

[CS79] J. B. Cooper and W. Schachermayer. Saks spaces and vector valued measures. Lecture Notes in Mathematics, 719:44–54, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065257/.

Csaszar:1977:SPC

[Csá77] Á. Császár. Some problems concerning C(X). Lecture Notes in Mathematics, 609:43–55, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068671/.

Chui:1976:RCL

[CSW76] C. K. Chui, P. W. Smith, and J. D. Ward. On the range of certain locally determined spline projections. Lecture Notes in Mathematics, 556:122–135, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087400/.

Corones:1976:PTA

[CT76] James P. Corones and Frank J. Testa. Pseudopotentials and their applications. Lecture Notes in Mathematics, 515:184–198, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081168/.

Cobanjan:1978:CEC

[ČT78a] S. A. Čobanjan and V. I. Tarieladze. A counter-example concerning CLT in Banach spaces. Lecture Notes in Mathematics, 656: 25–30, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068807/.

Cohen:1978:CGF

[CT78b] F. R. Cohen and L. R. Taylor. Computations of Gelfand–Fuks cohomology, the cohomology of function spaces, and the cohomology of configuration spaces. Lecture Notes in Mathematics, 657:106–143, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069229/.

Curtis:1978:GC

[CT78c] Morton Curtis and George Terrell. HP^{∞}, genuine and counterfeit. Lecture Notes in Mathematics, 664:76–82, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061691/.

Cuda:1977:RBP

[Čud77] Karel Čuda. The relation between ϵ–δ procedures and the infinitely small in nonstandard methods. Lecture Notes in Mathematics, 619:143–152, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067648/.

Cummings:1976:CFC

[Cum76] Larry J. Cummings. Comma-free codes and incidence algebras. Lecture Notes in Mathematics, 560:1–6, 1976. CODEN LN-MAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097363/.

Cunningham:1975:CMA

[Cun75] Ellen Cunningham. Chain models: Applications of consistency properties and back-and-forth techniques in infinite-quantifier languages. Lecture Notes in Mathematics, 492:125–142, 1975. CO-DEN LNMAA2. ISBN 3-540-07419-8 (print), 3-540-37949-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081122/.

Cuppens:1977:FCF

[Cup77] Roger Cuppens. Fonctions à crete. (French) []. Lecture Notes in Mathematics, 578:195–206, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091470/.

Curley:1975:NSE

[Cur75a] C. A. Curley. Non-singular extensions of Morse functions. Lecture Notes in Mathematics, 484:8–22, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082141/.

Curry:1975:SGS

[Cur75b] Haskell B. Curry. A study of generalized standardization in combinatory logic. Lecture Notes in Mathematics, 500:44–55, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079546/.

Curtis:1975:HTA

[Cur75c] Morton Curtis. A homotopy theoretic approach to Lie groups. Lecture Notes in Mathematics, 438:137–141, 1975. CODEN LN-MAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066111/.

Cushing:1979:NPS

[Cus79] J. M. Cushing. Nontrivial periodic solutions of some Volterra integral equations. Lecture Notes in Mathematics, 737:50–66, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064495/.

Ciesielski:1975:PWS

[CUW75] Prof. Z. Ciesielski, Prof. K. Urbanik, and Prof. W. A. Woyczyński, editors. Probability-Winter School: Proceedings of the Fourth Winter School on Probability Held at Karpacz, Poland, January 1975, volume 472 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 0-387-07190-3, 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 472. URL http:

//link.springer.com/book/10.1007/BFb0081942;   http://
www.springerlink.com/content/978-3-540-37556-2.

Cuyt:1979:APA

[Cuy79] Annie A. M. Cuyt. Abstract Padé-approximants in operator theory. Lecture Notes in Mathematics, 765:61–87, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085575/.

Carmona:1977:NCH

[CV77a] Jacques Carmona and Michèle Vergne, editors. Non-Commutative Harmonic Analysis: Actes du Colloque d’Analyse Harmonique Non-Commutative, Marseille–Luminy, 5 au 9 Juillet, 1976, volume 587 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 587. URL http:/

/link.springer.com/book/10.1007/BFb0087912;   http://
www.springerlink.com/content/978-3-540-37365-0.

Castaing:1977:CTM

[CV77b] Charles Castaing and Michel Valadier. Compactness theorems of measurable selections and integral representation theorem. Lecture Notes in Mathematics, 580:125–161, 1977. CODEN LN-MAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087690/.

Castaing:1977:CAM

[CV77c] Charles Castaing and Michel Valadier. Convex Analysis and Measurable Multifunctions, volume 580 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 278 pp. LCCN QA3

.L28 no. 580. URL http://link.springer.com/book/10.1007/
BFb0087685;   http://www.springerlink.com/content/978-
3-540-37384-1.

Castaing:1977:CF

[CV77d] Charles Castaing and Michel Valadier. Convex functions. Lecture Notes in Mathematics, 580:1–36, 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087686/.

Castaing:1977:CIL

[CV77e] Charles Castaing and Michel Valadier. Convex integrand on locally convex spaces. And its applications. Lecture Notes in Mathematics, 580:195–229, 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087692/.

Castaing:1977:HDH

[CV77f] Charles Castaing and Michel Valadier. Hausdorff distance and Hausdorff uniformity. Lecture Notes in Mathematics, 580:37– 57, 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087687/.

Castaing:1977:MM

[CV77g] Charles Castaing and Michel Valadier. Measurable multifunctions. Lecture Notes in Mathematics, 580:59–90, 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087688/.

Castaing:1977:NSD

[CV77h] Charles Castaing and Michel Valadier. A natural supplement of L^{↑} in the dual of L^{∞}. applications. Lecture Notes in Mathematics, 580:231–272, 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087693/.

Castaing:1977:PMM

[CV77i] Charles Castaing and Michel Valadier. Primitive of multifunctions and multivalued differential equations. Lecture Notes in Mathematics, 580:163–194, 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087691/.

Castaing:1977:TPP

[CV77j] Charles Castaing and Michel Valadier. Topological property of the profile of a measurable multifunction with compact convex values. Lecture Notes in Mathematics, 580:91–123, 1977. CODEN LNMAA2. ISBN 3-540-08144-5 (print), 3-540-37384-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087689/.

Carmona:1979:NCH

[CV79] Jacques Carmona and Michèle Vergne, editors. Non-Commutative Harmonic Analysis: Proceedings, Marseille–Luminy, France, June 26 to 30, 1978 Actes du Colloque d’Analyse Harmonique Non Commutative, volume 728 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 728. URL http:

//link.springer.com/book/10.1007/BFb0063333;   http://
www.springerlink.com/content/978-3-540-35131-3.

Cousins:1975:MSO

[CW75] Elizabeth Cousins and W. D. Wallis. Maximal sets of one-factors. Lecture Notes in Mathematics, 452:90–94, 1975. CODEN LN-MAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069547/.

Casse:1976:CMI

[CW76a] Louis R. A. Casse and Walter D. Wallis, editors. Combinatorial Mathematics IV: Proceedings of the Fourth Australian Conference Held at the University of Adelaide August 27–29, 1975, volume 560 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 560. URL http://link.springer.com/book/10.1007/
BFb0097362;   http://www.springerlink.com/content/978-
3-540-37537-1.

Chacon:1976:ODP

[CW76b] R. V. Chacon and J. B. Walsh. One-dimensional potential embedding. Lecture Notes in Mathematics, 511:19–23, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101093/.

Cairoli:1977:CT

[CW77a] R. Cairoli and J. B. Walsh. On changing time. Lecture Notes in Mathematics, 581:349–355, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087201/.

Cairoli:1977:PPH

[CW77b] R. Cairoli and J. B. Walsh. Prolongement de processus holomorphes cas ‘carré intégrable’. (French) []. Lecture Notes in Mathematics, 581:327–339, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087199/.

Cairoli:1977:SEH

[CW77c] R. Cairoli and J. B. Walsh. Some examples of holomorphic processes. Lecture Notes in Mathematics, 581:340–348, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087200/.

Chao:1978:CEG

[CW78] Chong-Yun Chao and Earl Glen Whitehead, Jr. On chromatic equivalence of graphs. Lecture Notes in Mathematics, 642:121– 131, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070369/.

Campbell:1979:CNM

[CW79] Douglas M. Campbell and Gene Wickes. Characterizations of normal meromorphic functions. Lecture Notes in Mathematics, 747: 55–72, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063959/.

Curry:1978:THB

[CY78] James H. Curry and James A. Yorke. A transition from Hopf bifurcation to chaos: Computer experiments with maps on R^{2}. Lecture Notes in Mathematics, 668:48–66, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101779/.

Czap:1975:EPF

[Cza75] Hans Czap. Exact penalty-functions in infinite optimization. Lecture Notes in Mathematics, 477:27–43, 1975. CODEN LN-MAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079165/.

Czyz:1978:SAG

[Czy78] J. Czyż. On some approach to geometric quantization. Lecture Notes in Mathematics, 676:315–328, 1978. CODEN LN-MAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063678/.

deAcosta:1979:GCC

[dA79] Alejandro de Acosta. On the general converse central limit theorem in Banach spaces. Lecture Notes in Mathematics, 709:1– 7, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071944/.

Dabeche:1979:FCR

[Dab79] A. Dabèche. Formes canoniques rationnelles d’un système différentiel á point singulier irrégulier. (French) []. Lecture Notes in Mathematics, 712:20–32, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062812/.

Dahlquist:1976:EAC

[Dah76] Germund Dahlquist. Error analysis for a class of methods for stiff non-linear initial value problems. Lecture Notes in Mathematics, 506:60–72, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080115/.

Dales:1976:HPD

[Dal76] H. G. Dales. Higher point derivations on commutative Banach algebras. Lecture Notes in Mathematics, 512:24–30, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080021/.

Damodaran:1977:DVM

[Dam77] Meledath Damodaran. On the distribution of values of meromorphic functions of slow growth. Lecture Notes in Mathematics, 599:17–21, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096820/.

Damlamian:1978:CEQ

[Dam78] Alain Damlamian. Sur certaines équations quasi-linéaires de type ≪divergentielle^{≫} d’ordre arbitraire. (French) []. Lecture Notes in Mathematics, 665:49–55, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061797/.

Darrozes:1977:CSE

[Dar77] Professeur J.-S. Darrozes. Comportement singulier des écoulements à grand nombre de Reynolds au voisinage du bord d’attaque d’une allé mince. (French) []. Lecture Notes in Mathematics, 594:77– 107, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086082/.

Das:1975:ADM

[Das75] P. C. Das. Application of Dubovitskii–Milyutin formalism to optimal setting problem with constraints. Lecture Notes in Mathematics, 477:44–58, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079166/.

Daverman:1975:ATD

[Dav75a] Robert J. Daverman. On the absence of tame disks in certain wild cells. Lecture Notes in Mathematics, 438:142–155, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066112/.

Daverman:1975:SRP

[Dav75b] Robert J. Daverman. A summary of results and problems concerning flatness of codimension one spheres in E^{n}. Lecture Notes in Mathematics, 438:156–165, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066113/.

Davie:1976:CEN

[Dav76] A. M. Davie. Classification of essentially normal operators. Lecture Notes in Mathematics, 512:31–55, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080022/.

Davis:1978:NSC

[Dav78a] M. H. A. Davis. Nonlinear semigroups in the control of partiallyobservable stochastic systems. Lecture Notes in Mathematics, 695:

37–49, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062653/.

Davis:1978:AS

[Dav78b] Michael Davis. Actions on spheres. Lecture Notes in Mathematics, 643:66–84, 1978. CODEN LNMAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065349/.

Davis:1978:CCA

[Dav78c] Michael Davis. Concordance classes of actions on spheres. Lecture Notes in Mathematics, 643:85–102, 1978. CODEN LN-MAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065350/.

Davis:1978:I

[Dav78d] Michael Davis. Introduction. Lecture Notes in Mathematics, 643: 1–2, 1978. CODEN LNMAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065344.

Davis:1978:MAM

[Dav78e] Michael Davis. Multiaxial Actions on Manifolds, volume 643 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 141 pp. LCCN QA3

.L28 no. 643. URL http://link.springer.com/book/10.1007/
BFb0065343;   http://www.springerlink.com/content/978-
3-540-35911-1.

Davis:1978:NBS

[Dav78f] Michael Davis. The normal bundles of the strata and twist invariants. Lecture Notes in Mathematics, 643:24–33, 1978. CODEN LNMAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065347/.

Davis:1978:SH

[Dav78g] Michael Davis. Some history. Lecture Notes in Mathematics, 643: 3–17, 1978. CODEN LNMAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065345/.

Davis:1978:SPF

[Dav78h] Michael Davis. Some preliminary facts concerning the linear model and its orbit space. Lecture Notes in Mathematics, 643: 18–23, 1978. CODEN LNMAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065346/.

Davis:1978:STM

[Dav78i] Michael Davis. A structure theorem for multiaxial actions and some of its consequences. Lecture Notes in Mathematics, 643: 34–65, 1978. CODEN LNMAA2. ISBN 3-540-08667-6 (print), 3-540-35911-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065348/.

Davis:1979:BOM

[Dav79] Donald M. Davis. Bp-operations and mappings of stunted complex projective spaces. Lecture Notes in Mathematics, 741:373– 393, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062150/.

Day:1975:MNS

[Day75] Mahlon M. Day. Mimicry in normed spaces. Lecture Notes in Mathematics, 490:91–106, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081135/.

Dayton:1976:CNA

[Day76] Barry H. Dayton. SK_{1} of commutative normed algebras. Lecture Notes in Mathematics, 551:30–43, 1976. CODEN LN-MAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079991/;   http:
//link.springer.com/chapter/10.1007/BFb0079992/.

deBoor:1976:ODS

[dB76] Carl de Boor. Odd-degree spline interpolation at a biinfinite knot sequence. Lecture Notes in Mathematics, 556:30–53, 1976. CO-DEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087395/.

Deligne:1977:CEP

[DB77] P. Deligne and J. F. Boutot. Cohomologie étale: les points de départ. (French) []. Lecture Notes in Mathematics, 569:4–75, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091518/.

Delanghe:1978:MPV

[DB78] R. Delanghe and C. Blondia. On measurable and partitionable vector valued multifunctions. Lecture Notes in Mathematics, 645: 35–47, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069661/.

Robinson:1977:PAY

[dBR77] Gilbert de B. Robinson. The papers of Alfred Young 1873–1940. Lecture Notes in Mathematics, 579:11–28, 1977. CODEN LN-MAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0090010/.

deBoor:1976:CIS

[dBS76] Carl de Boor and I. J. Schoenberg. Cardinal interpolation and spline functions VIII. The Budan–Fourier theorem for splines and applications. Lecture Notes in Mathematics, 501:1–79, 1976. CO-DEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079740/.

Dacunha-Castelle:1975:PEB

[DC75a] D. Dacunha-Castelle. Processus et espaces de Banach invariants par rearrangement. (French) []. Lecture Notes in Mathematics, 465:246–267, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102996/.

Dacunha-Castelle:1975:SES

[DC75b] D. Dacunha-Castelle. Sous-espaces symétriques des espaces d’Orlicz. (French) []. Lecture Notes in Mathematics, 465:268– 284, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102997/.

Dacunha-Castelle:1978:VCP

[DC78] D. Dacunha-Castelle. Vitesse de convergence pour certains problèmes statistiques. (French) []. Lecture Notes in Mathematics, 678:1–172, 1978. CODEN LNMAA2. ISBN 3-540-08938-1 (print), 3-540-35716-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062797/.

Dacunha-Castelle:1978:EEP

[DCHRH78] Didier Dacunha-Castelle, Herbert Heyer, B. Roynette, and Paul Louis Hennequin, editors. Ecole d’été de probabilités de Saint-Flour VII — 1977. (French) [Saint-Flour Summer School on Probability VII — 1977], volume 678 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 0-387-08938-1, 3-540-08938-1 (print), 3-540-35716-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 379 pp. LCCN QA3 .L28 no. 678; QA3 .L28 no. 678 QA273.18. URL http:

//link.springer.com/book/10.1007/BFb0062796;   http://
www.springerlink.com/content/978-3-540-35716-2.

Dacunha-Castelle:1978:JSP

[DCV78] Didier Dacunha-Castelle and Bernard Van Cutsem, editors. Journées de Statistique des Processus Stochastiques: Proceedings, Grenoble, Juin 1977, volume 636 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 636. URL http:

//link.springer.com/book/10.1007/BFb0063258;   http://
www.springerlink.com/content/978-3-540-35925-8.

Dora:1979:APH

[DDC79] J. Della Dora and C. Di-Crescenzo. Approximation de Padé– Hermite. (French) [Padé–Hermite approximation]. Lecture Notes in Mathematics, 765:88–115, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085576/.

Douglas:1978:TSM

[DDP78] Jim Douglas, Jr., Todd Dupont, and Peter Percell. A timestepping method for Galerkin approximations for nonlinear parabolic equations. Lecture Notes in Mathematics, 630:64– 75, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067697/.

deLaPradelle:1978:CPE

[de 78] A. de La Pradelle. Cônes de potentiels dans des espaces de Banach adaptés et dualité. (French) []. Lecture Notes in Mathematics, 681:215–233, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065878/.

Davis:1977:DSS

[DE77] W. Davis and Per Enflo. The distance of symmetric spaces from ℓ_{p}^{(n)}. Lecture Notes in Mathematics, 604:25–28, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069201/.

Dedecker:1977:GSG

[Ded77] Paul Dedecker. On the generalization of symplectic geometry to multiple integrals in the Calculus of Variations. Lecture Notes in Mathematics, 570:395–456, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087794/.

Deddens:1978:ADN

[Ded78] James A. Deddens. Another description of nest algebras. Lecture Notes in Mathematics, 693:77–86, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064662/.

Dehen:1976:SGL

[Deh76] Danièle Dehen. Semi-groupes localement équicontinus sur des espaces localement convexes. (French) []. Lecture Notes in Mathematics, 518:74–82, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080404/.

Deimling:1977:AS

[Dei77a] Klaus Deimling. Approximate solutions. Lecture Notes in Mathematics, 596:95–112, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091644/.

Deimling:1977:CC

[Dei77b] Klaus Deimling. Compactness conditions. Lecture Notes in Mathematics, 596:18–31, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091639/.

Deimling:1977:CDT

[Dei77c] Klaus Deimling. Conditions of dissipative type. Lecture Notes in Mathematics, 596:32–48, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091640/.

Deimling:1977:CSO

[Dei77d] Klaus Deimling. Countable systems of ordinary differential equations. Lecture Notes in Mathematics, 596:78–94, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091643/.

Deimling:1977:FID

[Dei77e] Klaus Deimling. Flow invariance and differential inequalities. Lecture Notes in Mathematics, 596:65–77, 1977. CODEN LN-MAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091642/.

Deimling:1977:I

[Dei77f] Klaus Deimling. Introduction. Lecture Notes in Mathematics, 596:1–3, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091637/.

Deimling:1977:LTC

[Dei77g] Klaus Deimling. Lipschitz type conditions. Lecture Notes in Mathematics, 596:4–17, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091638/.

Deimling:1977:ODE

[Dei77h] Klaus Deimling. Ordinary Differential Equations in Banach Spaces, volume 596 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 137 pp. LCCN QA3 .L28 no. 596. URL http:

//link.springer.com/book/10.1007/BFb0091636;   http://
www.springerlink.com/content/978-3-540-37338-4.

Deimling:1977:RT

[Dei77i] Klaus Deimling. Related topics. Lecture Notes in Mathematics, 596:113–125, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091645/.

Deimling:1977:SCS

[Dei77j] Klaus Deimling. Solutions in closed sets. Lecture Notes in Mathematics, 596:49–64, 1977. CODEN LNMAA2. ISBN 3-540-08260-3 (print), 3-540-37338-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091641/.

Deimling:1979:FPC

[Dei79] Klaus Deimling. Fixed points of condensing maps. Lecture Notes in Mathematics, 737:67–82, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064496/.

Dejter:1978:T

[Dej78a] Italo José Dejter. G-transversality to CP^{n}. Lecture Notes in Mathematics, 652:222–239, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063515/.

Dejter:1978:SPF

[Dej78b] Italo José Dejter. Some problems in foliation theory and related areas. Lecture Notes in Mathematics, 652:240–252, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063516/.

Dekkers:1979:MCH

[Dek79] W. Dekkers. The matrix of a connection having regular singularities on a vector bundle of rank 2 on IP^{1}(c). Lecture Notes in Mathematics, 712:33–43, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062813/.

Deligne:1975:CEF

[Del75] P. Deligne. Courbes elliptiques: Formulaire d’aprè J. Tate. (French) []. Lecture Notes in Mathematics, 476:53–73, 1975. CO-DEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0097583/.

Delanghe:1976:HMR

[Del76] R. Delanghe. On Hilbert modules with reproducing kernel. Lecture Notes in Mathematics, 561:158–170, 1976. CODEN LN-MAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087634/.

Deligne:1977:AFT

[Del77a] Pierre Deligne. Applications de la formule des traces aux sommes trigonométrigues. (French) []. Lecture Notes in Mathematics, 569: 168–232, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091523/.

Deligne:1977:CES

[Del77b] Pierre Deligne, editor. Cohomologie Étale: Séminaire de Géométrie Algébrique du Bois–Marie SGA41/2. (French) [Étale cohomology: Bois–Marie Algebraic Geometry Seminar at rmSGA41/2], volume 569 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 569. URL http:/

/link.springer.com/book/10.1007/BFb0091516;   http://
www.springerlink.com/content/978-3-540-37507-4.

Deligne:1977:DFD

[Del77c] Pierre Deligne. Dualité. (French) [Duality]. Lecture Notes in Mathematics, 569:154–167, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091522/.

Deligne:1977:E

[Del77d] Pierre Deligne. Erratum. Lecture Notes in Mathematics, 569: 312, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091526.

Deligne:1977:DDC

[Del77e] Pierre Deligne. Les difféomorphismes du cercle [d’après M. R. Herman]. (French) []. Lecture Notes in Mathematics, 567:99–121, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096064/.

Deligne:1977:RFT

[Del77f] Pierre Deligne. Rapport sur la formule des traces. (French) []. Lecture Notes in Mathematics, 569:76–109, 1977. CODEN LN-MAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091519/.

Deligne:1977:TFC

[Del77g] Pierre Deligne. Théorèmes de finitude en cohomologie ℓ-adique. (French) []. Lecture Notes in Mathematics, 569:233–261, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091524/.

Deligne:1977:FAP

[Del77h] Pierre Deligne. Un fil d’Ariane pour SGA 4, SGA 41/2 et SGA 5. (French) []. Lecture Notes in Mathematics, 569:1–3, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091517/.

Dellacherie:1977:DRS

[Del77i] C. Dellacherie. Deux remarques sur la séparabilité optionnelle. (French) []. Lecture Notes in Mathematics, 581:47–50, 1977. CO-DEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087186/.

Dellacherie:1977:DTD

[Del77j] C. Dellacherie. Les dérivations en théorie descriptive des ensembles et le théorème de la borne. (French) []. Lecture Notes in Mathematics, 581:34–46, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087185/.

Delorme:1979:CCR

[Del79] P. Delorme. Sur la chomologie continue des représentations unitaires irreductibles des groupes de Lie semi-simples complexes. (French) []. Lecture Notes in Mathematics, 739:179–191, 1979.

CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062492/.

Demazure:1975:CGP

[Dem75] Michel Demazure. Classification des germes à point critique isolé et à nombres de modules 0 ou 1 [d’après V. I. Arnol’d]. (French) []. Lecture Notes in Mathematics, 431:124–142, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066368/.

Demazure:1976:DCM

[Dem76a] Michel Demazure. Démonstration de la conjecture de Mumford. (French) []. Lecture Notes in Mathematics, 514:138–144, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080063/.

DeMeyer:1976:BGA

[DeM76b] F. R. DeMeyer. The Brauer group of affine curves. Lecture Notes in Mathematics, 549:16–24, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077330/.

Demazure:1977:IMF

[Dem77a] Michel Demazure. Identités de Macdonald. (French) []. Lecture Notes in Mathematics, 567:191–201, 1977. CODEN LN-MAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096070/.

Demidov:1977:PSP

[Dem77b] A. S. Demidov. Sur la perturbation ‘singulière’ dans un problème à frontière libre. (French) []. Lecture Notes in Mathematics, 594: 123–130, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086084/.

Demidov:1977:PEP

[Dem77c] A. S. Demidov. Sur les problèmes elliptiques pseudo-différenciels, à petit paramètre dans l’opérateur principal. (French) []. Lecture Notes in Mathematics, 594:108–122, 1977. CODEN LN-MAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086083/.

Demailly:1978:DEF

[Dem78] J.-P. Demailly. Différents exemples de fibrés holomorphes non de Stein. (French) []. Lecture Notes in Mathematics, 694:15–41, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063243/.

Denker:1976:MMEb

[Den76] Manfred Denker. Measures with maximal entropy. Lecture Notes in Mathematics, 532:70–112, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080173/.

Deschamps:1979:CGG

[Des79a] Mireille Deschamps. Courbes de genre géométrique borné sur une surface de type général [d’après F. A. Bogomolov]. (French) []. Lecture Notes in Mathematics, 710:233–247, 1979. CODEN LN-MAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069982/.

Deshouillers:1979:PRP

[Des79b] Jean-Marc Deshouillers. Progrès récents des petits cribles arithmétiques [d’après Chen, iwaniec, …]. (French) []. Lecture Notes in Mathematics, 710:248–262, 1979. CODEN LN-MAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069983/.

Deuflhard:1975:RSM

[Deu75] Peter Deuflhard. A relaxation strategy for the modified Newton method. Lecture Notes in Mathematics, 477:59–73, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079167/.

Devlin:1975:IPS

[Dev75] Keith J. Devlin. Indescribability properties and small large cardinals. Lecture Notes in Mathematics, 499:89–114, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079418/.

Devlin:1976:AMA

[Dev76] Keith J. Devlin. An alternative to Martin’s axiom. Lecture Notes in Mathematics, 537:65–76, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096894/.

Devlin:1977:AM

[Dev77a] Keith J. Devlin. Applications of V = L in mathematics. Lecture Notes in Mathematics, 617:36–64, 1977. CODEN LNMAA2. ISBN 3-540-08520-3 (print), 3-540-37034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070212/.

Devlin:1977:AC

[Dev77b] Keith J. Devlin. The axiom of constructibility. Lecture Notes in Mathematics, 617:27–35, 1977. CODEN LNMAA2. ISBN 3-540-08520-3 (print), 3-540-37034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070211/.

Devlin:1977:ACG

[Dev77c] Keith J. Devlin. The Axiom of Constructibility: a Guide for the Mathematician, volume 617 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08520-3 (print), 3-540-37034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 96 pp. LCCN QA3

.L28 no. 617. URL http://link.springer.com/book/10.1007/
BFb0070208;   http://www.springerlink.com/content/978-
3-540-37034-5.

Devlin:1977:FFP

[Dev77d] Keith J. Devlin. Four famous problems. Lecture Notes in Mathematics, 617:1–10, 1977. CODEN LNMAA2. ISBN 3-540-08520-3 (print), 3-540-37034-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070209/.

Devlin:1977:PMT

[Dev77e] Keith J. Devlin. A problem in measure theory. Lecture Notes in Mathematics, 617:65–80, 1977. CODEN LNMAA2. ISBN 3-540-08520-3 (print), 3-540-37034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070213/.

Devlin:1977:WST

[Dev77f] Keith J. Devlin. What is set theory? Lecture Notes in Mathematics, 617:11–26, 1977. CODEN LNMAA2. ISBN 3-540-08520-3 (print), 3-540-37034-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070210/.

Devaney:1978:THO

[Dev78] Robert L. Devaney. Transverse heteroclinic orbits in the anisotropic Kepler problem. Lecture Notes in Mathematics, 668: 67–87, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101780/.

deFigueiredo:1975:DPN

[dF75] Djairo Guedes de Figueiredo. The Dirichlet problem for nonlinear elliptic equations: a Hilbert space approach. Lecture Notes in Mathematics, 446:144–165, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070601/.

DeAlfaro:1978:CIF

[DFF78] V. De Alfaro, S. Fubini, and G. Furlan. Conformal invariance in field theory. Lecture Notes in Mathematics, 676:255–293, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063675/.

deGuzman:1975:BR

[dG75a] Prof. Miguel de Guzmán. The basis of rectangles B_{∋}. Lecture Notes in Mathematics, 481:109–133, 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081991/.

deGuzman:1975:DCP

[dG75b] Prof. Miguel de Guzmán. Differentiation and covering properties. Lecture Notes in Mathematics, 481:148–176, 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081993/.

deGuzman:1975:DI

[dG75c] Prof. Miguel de Guzmán. Differentiation of Integrals in R^{n}, volume 481 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 189 (est.) pp. LCCN QA3 .L28 no. 481. URL http:

//link.springer.com/book/10.1007/BFb0081986;   http://
www.springerlink.com/content/978-3-540-37604-0.

deGuzman:1975:HLM

[dG75d] Prof. Miguel de Guzmán. The Hardy–Littlewood maximal operator. Lecture Notes in Mathematics, 481:35–64, 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081988/.

deGuzman:1975:IB

[dG75e] Prof. Miguel de Guzmán. The interval basis B_{∈}. Lecture Notes in Mathematics, 481:92–108, 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081990/.

deGuzman:1975:MOD

[dG75f] Prof. Miguel de Guzmán. The maximal operator and the differentiation properties of a basis. Lecture Notes in Mathematics, 481:65–91, 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081989/.

deGuzman:1975:HP

[dG75g] Prof. Miguel de Guzmán. On the halo problem. Lecture Notes in Mathematics, 481:177–189, 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081994/.

deGuzman:1975:SCT

[dG75h] Prof. Miguel de Guzmán. Some covering theorems. Lecture Notes in Mathematics, 481:1–34, 1975. CODEN LNMAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081987/.

deGuzman:1975:SSD

[dG75i] Prof. Miguel de Guzmán. Some special differentiation bases. Lecture Notes in Mathematics, 481:134–147, 1975. CODEN LN-MAA2. ISBN 3-540-07399-X (print), 3-540-37604-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081992/.

Dlab:1975:RAP

[DG75j] Prof. Vlastimil Dlab and Prof. Peter Gabriel, editors. Representations of Algebras: Proceedings of the International Conference Ottawa 1974, volume 488 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 488. URL http:

//link.springer.com/book/10.1007/BFb0081211;   http://
www.springerlink.com/content/978-3-540-37921-8.

Dahmen:1976:SKC

[DG76a] W. Dahmen and E. Görlich. Schnelle Konvergenz: Charakterisierung der besten Approximation und Entropie. (German) []. Lecture Notes in Mathematics, 556:154, 1976. CODEN LN-MAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087403.

deGuzman:1976:DI

[dG76b] Miguel de Guzmán. Differentiation of integrals in R^{n}. Lecture Notes in Mathematics, 541:181–185, 1976. CODEN LN-MAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081051/.

Debiard:1976:FPF

[DG76c] A. Debiard and B. Gaveau. Fine potential and function algebras. Lecture Notes in Mathematics, 512:56–64, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080023/.

Dudley:1977:STG

[DG77] R. M. Dudley and Sam Gutmann. Stopping times with given laws. Lecture Notes in Mathematics, 581:51–58, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087187/.

deGroen:1979:SPS

[dG79] P. P. N. de Groen. Singular perturbations of spectra. Lecture Notes in Mathematics, 711:9–32, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062945/.

Denker:1976:CE

[DGS76a] Manfred Denker, Christian Grillenberger, and Karl Sigmund. The computation of entropy. Lecture Notes in Mathematics, 527:62– 67, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082376/.

Denker:1976:EBM

[DGS76b] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Entropy for Bernoulli and Markov shifts. Lecture Notes in Mathematics, 527:68–72, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082377/.

Denker:1976:ED

[DGS76c] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Ergodic decompositions. Lecture Notes in Mathematics, 527:73– 81, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082378/.

Denker:1976:ETC

[DGS76d] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Ergodic Theory on Compact Spaces, volume 527 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 360 pp. LCCN QA3

.L28 no. 527. URL http://link.springer.com/book/10.1007/
BFb0082364;   http://www.springerlink.com/content/978-
3-540-38263-8.

Denker:1976:E

[DGS76e] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Ergodicity. Lecture Notes in Mathematics, 527:23–29, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082370/.

Denker:1976:EH

[DGS76f] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Expansive homeomorphisms. Lecture Notes in Mathematics, 527: 103–116, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082381/.

Denker:1976:IE

[DGS76g] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Information and entropy. Lecture Notes in Mathematics, 527:56– 61, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082375/.

Denker:1976:I

[DGS76h] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Introduction. Lecture Notes in Mathematics, 527:1–2, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082365.

Denker:1976:IMC

[DGS76i] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Invariant measures for continuous tranformations. Lecture Notes in Mathematics, 527:17–19, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082368/.

Denker:1976:MTD

[DGS76j] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Measure-theoretic dynamical systems. Lecture Notes in Mathematics, 527:3–7, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082366/.

Denker:1976:MCM

[DGS76k] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Measures on compact metric spaces. Lecture Notes in Mathematics, 527:8–16, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082367/.

Denker:1976:MSS

[DGS76l] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Measures on the shift space. Lecture Notes in Mathematics, 527: 41–48, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082373/.

Denker:1976:MMEa

[DGS76m] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Measures with maximal entropy — intrinsically ergodic systems.

Lecture Notes in Mathematics, 527:147–159, 1976. CODEN LN-MAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082384/.

Denker:1976:MT

[DGS76n] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Mixing and transitivity. Lecture Notes in Mathematics, 527:30– 35, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082371/.

Denker:1976:PG

[DGS76o] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Partitions and generators. Lecture Notes in Mathematics, 527:49– 55, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082374/.

Denker:1976:SS

[DGS76p] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Shifts and subshifts. Lecture Notes in Mathematics, 527:36– 40, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082372/.

Denker:1976:SFT

[DGS76q] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Subshifts of finite type. Lecture Notes in Mathematics, 527:117– 130, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082382/.

Denker:1976:TA

[DGS76r] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Time averages. Lecture Notes in Mathematics, 527:20–22, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082369/.

Denker:1976:TE

[DGS76s] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Topological entropy. Lecture Notes in Mathematics, 527:82– 91, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082379/.

Denker:1976:TG

[DGS76t] Manfred Denker, Christian Grillenberger, and Karl Sigmund. Topological generators. Lecture Notes in Mathematics, 527:92– 102, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082380/.

Denker:1976:VPT

[DGS76u] Manfred Denker, Christian Grillenberger, and Karl Sigmund. The variational principle for topological entropy. Lecture Notes in Mathematics, 527:131–146, 1976. CODEN LNMAA2. ISBN 3-540-07797-9 (print), 3-540-38263-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082383/.

Deleanu:1977:GST

[DH77] Aristide Deleanu and Peter Hilton. Generalized shape theory. Lecture Notes in Mathematics, 609:56–65, 1977. CODEN LN-MAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068672/.

Dragan:1979:SPL

[DH79] V. Drăgan and A. Halanay. Singular perturbations and linear feedback control. Lecture Notes in Mathematics, 703:86–92, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067260/.

Dickmanns:1975:OSG

[Dic75] E. D. Dickmanns. Optimale Steuerungen für Gleitflugbahnen Maximaler Reichweite beim Eintritt in Planetenatmosphären. (German) []. Lecture Notes in Mathematics, 477:74–94, 1975.

CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079168/.

Dickman:1977:SS

[Dic77] M. A. Dickman. Structures Σ-saturèe. Lecture Notes in Mathematics, 619:153–168, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067649/.

Diestel:1975:CRT

[Die75a] Prof. Joseph Diestel. The classical renorming theorems. Lecture Notes in Mathematics, 485:94–127, 1975. CODEN LN-MAA2. ISBN 3-540-07402-3 (print), 3-540-37913-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082083/.

Diestel:1975:CDN

[Die75b] Prof. Joseph Diestel. Convexity and differentiability of norms. Lecture Notes in Mathematics, 485:20–53, 1975. CODEN LN-MAA2. ISBN 3-540-07402-3 (print), 3-540-37913-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082081/.

Diestel:1975:GBS

[Die75c] Prof. Joseph Diestel. Geometry of Banach Spaces — Selected Topics, volume 485 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07402-3 (print), 3-540-37913-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 282 pp. LCCN QA3 .L28 no. 485. URL http:

//link.springer.com/book/10.1007/BFb0082079;   http://
www.springerlink.com/content/978-3-540-37913-3.

Diestel:1975:RNT

[Die75d] Prof. Joseph Diestel. The Radon–Nikodým theorem for vector measures. Lecture Notes in Mathematics, 485:199–282, 1975. CO-DEN LNMAA2. ISBN 3-540-07402-3 (print), 3-540-37913-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082085/.

Diestel:1975:SFC

[Die75e] Prof. Joseph Diestel. Support functionals for closed bounded convex subsets of a Banach space. Lecture Notes in Mathematics, 485:1–19, 1975. CODEN LNMAA2. ISBN 3-540-07402-3 (print), 3-540-37913-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082080/.

Diestel:1975:UCU

[Die75f] Prof. Joseph Diestel. Uniformly convex and uniformly smooth Banach spaces. Lecture Notes in Mathematics, 485:54–93, 1975. CODEN LNMAA2. ISBN 3-540-07402-3 (print), 3-540-37913-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082082/.

Diestel:1975:WCG

[Die75g] Prof. Joseph Diestel. Weakly compactly generated Banach spaces. Lecture Notes in Mathematics, 485:128–198, 1975. CODEN LN-MAA2. ISBN 3-540-07402-3 (print), 3-540-37913-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082084/.

Dineen:1975:EDH

[Din75] Seán Dineen. Equivalent definitions of holomorphic functions in infinite dimensions. Lecture Notes in Mathematics, 474:114– 122, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077403/.

Dixmier:1975:IPC

[Dix75] Jacques Dixmier. Idéaux primitifs complètement premiers dans l’algèbre enveloppante de sl(3, C). (French) []. Lecture Notes in Mathematics, 466:38–55, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082196/.

Dixmier:1979:MOF

[Dix79] J. Dixmier. Sur la méthode des orbites. (French) []. Lecture Notes in Mathematics, 728:42–63, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063337/.

Devlin:1975:MTS

[DJ75] Keith I. Devlin and R. B. Jensen. Marginalia to a theorem of Silver. Lecture Notes in Mathematics, 499:115–142, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079419/.

Davis:1977:WCS

[DJ77] W. J. Davis and W. B. Johnson. Weakly convergent sequences of Banach space valued random variables. Lecture Notes in Mathematics, 604:29–31, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069202/.

Denker:1979:ETP

[DJ79] Manfred Denker and Konrad Jacobs, editors. Ergodic Theory: Proceedings, Oberwolfach, Germany, June 11–17, 1978, volume 729 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 729. URL http://link.springer.com/book/10.1007/
BFb0063275;   http://www.springerlink.com/content/978-
3-540-35130-6.

Denker:1979:COS

[DK79] Manfred Denker and Rainer Kombrink. On B-convex Orlicz spaces. Lecture Notes in Mathematics, 709:87–95, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071950/.

Dobrushin:1978:LIS

[DKT78] R. L. Dobrushin, V. I. Kryukov, and A. L. Toom, editors. Locally Interacting Systems and Their Application in Biology: Proceedings of the School-Seminar on Markov Interaction Processes in Biology, Held in Pushchino, Moscow Region, March, 1976, volume 653 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 653. URL http://link.springer.com/book/10.1007/
BFb0070079;   http://www.springerlink.com/content/978-
3-540-37044-4.

Divis:1976:LTN

[DL76] B. Diviš and W. J. Le Veque. Lectures on Transcendental Numbers, volume 546 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 212 (est.) pp. LCCN QA3 .L28 no. 546. URL http:

//link.springer.com/book/10.1007/BFb0081107;   http://
www.springerlink.com/content/978-3-540-37981-2.

delaHarpe:1979:IAC

[dlH79a] P. de la Harpe. Initiation à l’algèbre de Calkin. (French) []. Lecture Notes in Mathematics, 725:180–219, 1979. CODEN LN-MAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062617/.

delaHarpe:1979:MDG

[dlH79b] P. de la Harpe. Moyennabilité du groupe unitaire et propriété P de Schuartz des algèbres de von Neumann. (French) []. Lecture Notes in Mathematics, 725:220–227, 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062618/.

delaHarpe:1979:AOS

[dlH79c] Pierre de la Harpe, editor. Algèbres d’Opérateurs: Séminaire sur les Algèbres d’Opérateurs, les Plans-sur-Bex, Suisse, 13–18 mars 1978. (French) [], volume 725 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 725. URL http:

//link.springer.com/book/10.1007/BFb0062611;   http://
www.springerlink.com/content/978-3-540-35162-7.

Dloussky:1977:EHP

[Dlo77] G. Dloussky. Enveloppes d’holomorphie et prolongements d’hypersurfaces. (French) []. Lecture Notes in Mathematics, 578: 217–235, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091472/.

Dloussky:1978:ASP

[Dlo78a] G. Dloussky. Analycité séparée et prolongements analytiques. (French) []. Lecture Notes in Mathematics, 683:179–202, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063178/.

Dloussky:1978:PAA

[Dlo78b] Georges Dloussky. Prolongements d’applications analytiques. (French) []. Lecture Notes in Mathematics, 694:42–95, 1978. CO-DEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063244/.

delaRocque:1975:LIV

[dlR75] Palis Gilda de la Rocque. Linearly induced vector fields and R^{2}-actions on spheres. Lecture Notes in Mathematics, 468:13– 17, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082581/.

deMonvel:1975:PSS

[dM75a] Louis Boutet de Monvel. Propagation des singularités des solutions d’équations analogues à l’équation de Schrödinger. (French) []. Lecture Notes in Mathematics, 459:1–14, 1975. CODEN LN-MAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074189/.

Dellacherie:1975:NTP

[DM75b] C. Dellacherie and P.-A. Meyer. Un nouveau théorème de projection et de section. (French) []. Lecture Notes in Mathematics, 465:239–245, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102995/.

Diller:1975:IPT

[DM75c] Prof. Justus Diller and Prof. Gert H. Müller, editors. [vDash] ISILC Proof Theory Symposion: Dedicated to Kurt Schütte on the Occasion of His 65th Birthday Proceedings of the International Summer Institute and Logic Colloquium, Kiel 1974, volume 500 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 500. URL http://link.springer.com/book/10.1007/
BFb0079542;   http://www.springerlink.com/content/978-
3-540-38020-7.

Davis:1978:OTT

[DM78a] Donald M. Davis and Mark Mahowald. Obstruction theory and K-theory. Lecture Notes in Mathematics, 658:134–164, 1978. CO-DEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0068713/.

Dellacherie:1978:PDT

[DM78b] C. Dellacherie and P.-A. Meyer. A propos du travail de Yor sur le grossissement des tribus. (French) []. Lecture Notes in Mathematics, 649:70–77, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064596/.

Drinfeld:1979:IS

[DM79a] V. G. Drinfel’d and Ju. I. Manin. Instantons and sheaves on CIP^{3}. Lecture Notes in Mathematics, 732:60–81, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066637/.

Dupont:1979:ATA

[DM79b] Johan Louis Dupont and Ib Henning Madsen, editors. Algebraic Topology Aarhus 1978: Proceedings of a Symposium held at Aarhus, Denmark, August 7–12, 1978, volume 763 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 763. URL http://link.springer.com/book/10.1007/
BFb0088073;   http://www.springerlink.com/content/978-
3-540-38520-2.

Dellacherie:1977:SPX

[DMW77] C. Dellacherie, P.-A. Meyer, and M. Weil. Séminaire de Probabilités XI: Université de Strasbourg, volume 581 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 573 pp. LCCN QA3

.L28 no. 581. URL http://link.springer.com/book/10.1007/
BFb0087181;   http://www.springerlink.com/content/978-
3-540-37383-4.

Dellacherie:1978:SPX

[DMW78] C. Dellacherie, P.-A. Meyer, and M. Weil, editors. Séminaire de Probabilités XII: Université de Strasbourg 1976/77, volume 649 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 649. URL http://link.springer.com/book/10.1007/
BFb0064586;   http://www.springerlink.com/content/978-
3-540-35856-5.

Dellacherie:1979:SPX

[DMW79] C. Dellacherie, P.-A. Meyer, and M. Weil, editors. Séminaire de Probabilités XIII: Université de Strasbourg 1977/78, volume 721 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 721. URL http://link.springer.com/book/10.1007/
BFb0070846;   http://www.springerlink.com/content/978-
3-540-35189-4.

Dellacherie:1978:CPE

[DMY78] C. Dellacherie, P.-A. Meyer, and M. Yor. Sur certaines propriétés des espaces de Banach H^{1} et BMO. (French) []. Lecture Notes in Mathematics, 649:98–113, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064598/.

Descloux:1977:IEF

[DN77] Jean Descloux and Nabil Nassif. Interior L^{∞} estimates for finite element approximations of solutions of elliptic equations. Lecture Notes in Mathematics, 606:56–63, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064456/.

Descloux:1979:PSA

[DNR79] J. Descloux, N. Nassif, and J. Rappaz. On properties of spectral approximations. Lecture Notes in Mathematics, 703:81–85, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067259/.

Doi:1977:SCB

[DO77] Koji Doi and Masami Ohta. On some congruences between cusp forms on Γ_{o} (n). Lecture Notes in Mathematics, 601:91–105, 1977. CODEN LNMAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063946/.

Dobson:1975:CST

[Dob75] Annette J. Dobson. Comparing the shapes of trees. Lecture Notes in Mathematics, 452:95–100, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069548/.

Doherty:1975:APF

[Doh75] Margaret Doherty. An amusing proof in fluctuation theory. Lecture Notes in Mathematics, 452:101–104, 1975. CODEN LN-MAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069549/.

Dokken:1978:NDT

[Dok78] Douglas Dokken. A note on a distallity theorem of C. C. Moore. Lecture Notes in Mathematics, 668:88–94, 1978. CODEN LN-MAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101781/.

Dold:1978:GCF

[Dol78] Albrecht Dold. Geometric Cobordism and the fixed point transfer. Lecture Notes in Mathematics, 673:32–87, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064688/.

Donaldson:1975:FIR

[Don75] James A. Donaldson. A family of integral representations for the solution of the diffusion equation. Lecture Notes in Mathematics, 457:146–150, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067102/.

Donig:1976:PAK

[Don76] J. Donig. A priori Abschätzungen für eine Klasse elliptischer Pseudo-Differentialoperatoren im Raum L^{p} (R^{n} ). (German) []. Lecture Notes in Mathematics, 561:171–191, 1976. CODEN LN-MAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087635/.

Douglas:1977:EAH

[Dou77a] R. G. Douglas. Extensions of C∗-algebras and K-homology. Lecture Notes in Mathematics, 575:44–52, 1977. CODEN LN-MAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095700/.

Douglas:1977:GMN

[Dou77b] Jim Douglas, Jr. H^{1}-Galerkin methods for a nonlinear Dirichlet problem. Lecture Notes in Mathematics, 606:64–86, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064457/.

Douglas:1978:UCD

[Dou78] Roy Douglas. The uniqueness of coproduct decompositions for algebras over a field. Lecture Notes in Mathematics, 673:1–6, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064686/.

Dovermann:1979:AES

[Dov79] Karl Heinz Dovermann. Addition of equivariant surgery obstructions. Lecture Notes in Mathematics, 741:244–271, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062143/.

duPlessis:1976:CIR

[dP76a] Andrew du Plessis. Contact-invariant regularity conditions. Lecture Notes in Mathematics, 535:205–236, 1976. CODEN LN-MAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080502/.

duPlessis:1976:USM

[dP76b] Andrew du Plessis. Unfoldings of smooth map-germs. Lecture Notes in Mathematics, 552:64–123, 1976. CODEN LN-MAA2. ISBN 3-540-07997-1 (print), 3-540-37957-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0095248/.

Delvos:1977:TOB

[DP77] Dr. Franz-Jürgen Delvos and Dipl.-Math. Horst Posdorf. N-th order blending. Lecture Notes in Mathematics, 571:53–64, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086564/.

dePillis:1979:ICL

[dP79] John de Pillis. Improved convergence for linear systems using three-part splittings. Lecture Notes in Mathematics, 701:97– 114, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062076/.

Dunkl:1975:BR

[DR75a] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Basic results. Lecture Notes in Mathematics, 435:1–9, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068936/.

Dunkl:1975:DSS

[DR75b] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Discrete separative semigroups. Lecture Notes in Mathematics, 435:52–79, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068939/.

Dunkl:1975:HSR

[DR75c] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Hilbert space representations. Lecture Notes in Mathematics, 435:138–162, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068943/.

Dunkl:1975:PDC

[DR75d] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Positive-definite and completely positive functions. Lecture Notes in Mathematics, 435:27–51, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068938/.

Dunkl:1975:RAa

[DR75e] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. The representation algebra. Lecture Notes in Mathematics, 435:10–26, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068937/.

Dunkl:1975:RAb

[DR75f] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Representations in Q-algebras. Lecture Notes in Mathematics, 435:99–117, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068941/.

Dunkl:1975:RCS

[DR75g] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Representations of Commutative Semitopological Semigroups, volume 435 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 181 pp. LCCN QA3

.L28 no. 435. URL http://link.springer.com/book/10.1007/
BFb0068935;   http://www.springerlink.com/content/978-
3-540-37402-2.

Dunkl:1975:SCR

[DR75h] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Special cases of Q-representations. Lecture Notes in Mathematics, 435:118– 137, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068942/.

Dunkl:1975:SLC

[DR75i] Dr. Charles F. Dunkl and Dr. Donald E. Ramirez. Subsemigroups of locally compact abelian groups and weakly almost periodic functions. Lecture Notes in Mathematics, 435:80–98, 1975. CODEN LNMAA2. ISBN 3-540-07132-6 (print), 3-540-37402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068940/.

deRuijter:1979:BLL

[dR79] Will de Ruijter. Boundary layers in large scale ocean circulation. Lecture Notes in Mathematics, 711:125–145, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062950/.

Drasin:1979:AQC

[Dra79] David Drasin. An application of quasi-conformal methods to a problem in value-distribution theory. Lecture Notes in Mathematics, 747:92–100, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063961/.

Dress:1975:RGR

[Dre75] Andreas W. M. Dress. On relative Grothendieck rings. Lecture Notes in Mathematics, 488:79–131, 1975. CODEN LN-MAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081218/.

Dreseler:1977:SPF

[Dre77] Dr. Bernd Dreseler. On summation processes of Fourier expansions for spherical functions. Lecture Notes in Mathematics, 571: 65–84, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086565/.

Drisch:1979:FAF

[Dri79] Th. Drisch. Faltungshalbgruppen sind analytische Funktionen ihrer infinitesimalen Funktionale. (German) []. Lecture Notes in Mathematics, 706:67–87, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063115/.

Delvos:1976:ESM

[DS76a] Dr. Franz-Jürgen Delvos and Prof. Dr. Walter Schempp. An extension of Sard’s method. Lecture Notes in Mathematics, 501: 80–91, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079741/.

Devore:1976:CTA

[DS76b] R. Devore and K. Scherer. A constructive theory for approximation by splines with an arbitrary sequence of knot sets. Lecture Notes in Mathematics, 556:167–183, 1976. CODEN LN-MAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087405/.

Dowker:1976:PSC

[DS76c] C. H. Dowker and Dona Strauss. Products and sums in the category of frames. Lecture Notes in Mathematics, 540:208–219, 1976.

CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080860/.

Dydak:1978:APA

[DS78a] Jerzy Dydak and Jack Segal. Algebraic properties associated with shape theory. Lecture Notes in Mathematics, 688:77–87, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067578/.

Dydak:1978:CLM

[DS78b] Jerzy Dydak and Jack Segal. Cell-like maps. Lecture Notes in Mathematics, 688:123–140, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067582/.

Dydak:1978:CPP

[DS78c] Jerzy Dydak and Jack Segal. Characterizations and properties of pointed ANSR’s. Lecture Notes in Mathematics, 688:111– 122, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067581/.

Dydak:1978:GPS

[DS78d] Jerzy Dydak and Jack Segal. General properties of the shape category and the shape functor. Lecture Notes in Mathematics, 688:47–63, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067576/.

Dydak:1978:I

[DS78e] Jerzy Dydak and Jack Segal. Introduction. Lecture Notes in Mathematics, 688:1–5, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067573/.

Dydak:1978:PM

[DS78f] Jerzy Dydak and Jack Segal. Pointed 1-movability. Lecture Notes in Mathematics, 688:88–102, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067579/.

Dydak:1978:P

[DS78g] Jerzy Dydak and Jack Segal. Preliminaries. Lecture Notes in Mathematics, 688:6–19, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067574/.

Dydak:1978:SC

[DS78h] Jerzy Dydak and Jack Segal. The shape category. Lecture Notes in Mathematics, 688:20–46, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067575/.

Dydak:1978:SI

[DS78i] Jerzy Dydak and Jack Segal. Shape invariants. Lecture Notes in Mathematics, 688:64–76, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067577/.

Dydak:1978:STI

[DS78j] Jerzy Dydak and Jack Segal. Shape Theory: An Introduction, volume 688 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 150 pp. LCCN QA3 .L28 no. 688. URL http:

//link.springer.com/book/10.1007/BFb0067572;   http://
www.springerlink.com/content/978-3-540-35684-4.

Dydak:1978:SOP

[DS78k] Jerzy Dydak and Jack Segal. Some open problems. Lecture Notes in Mathematics, 688:141–142, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067583.

Dydak:1978:WHT

[DS78l] Jerzy Dydak and Jack Segal. Whitehead and Hurewicz theorems in shape theory. Lecture Notes in Mathematics, 688:103–110, 1978. CODEN LNMAA2. ISBN 3-540-08955-1 (print), 3-540-35684-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067580/.

deSa:1979:LCM

[dS79] Eugénia César de Sá. A link calculus for 4-manifolds. Lecture Notes in Mathematics, 722:16–30, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063185/.

Lazaro:1975:EFF

[dSLM75a] J. de Sam Lazaro and P.-A. Meyer. Exemples de K-flots. (French) []. Lecture Notes in Mathematics, 465:30–37, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0102984/.

Lazaro:1975:GFS

[dSLM75b] J. de Sam Lazaro and P.-A. Meyer. Généralités. Flot sous une fonction. (French) []. Lecture Notes in Mathematics, 465:2–14, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0102982/.

Lazaro:1975:HAO

[dSLM75c] J. de Sam Lazaro and P.-A. Meyer. Hélices à accroissements orthogonaux et prédiction. (French) []. Lecture Notes in Mathematics, 465:52–72, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102986/.

Lazaro:1975:HCM

[dSLM75d] J. de Sam Lazaro and P.-A. Meyer. Hélices croissantes et mesures de Palm. (French) []. Lecture Notes in Mathematics, 465:38– 51, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102985/.

Lazaro:1975:TNF

[dSLM75e] J. de Sam Lazaro and P.-A. Meyer. Le théorème de Nisio. (French) []. Lecture Notes in Mathematics, 465:89–96, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0102988/.

Lazaro:1975:MPT

[dSLM75f] J. de Sam Lazaro and P.-A. Meyer. Mesures de Palm, théorème d’Ambrose–Kakutani. (French) []. Lecture Notes in Mathematics, 465:15–29, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102983/.

Lazaro:1975:MSN

[dSLM75g] J. de Sam Lazaro and P.-A. Meyer. Multiplicité spectrale, nombre d’hélices HAO. (French) []. Lecture Notes in Mathematics, 465:73–88, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102987/.

Delvos:1976:CAS

[DSS76] Dr. F. J. Delvos, Dr. W. Schäfer, and Prof. Dr. W. Schempp. Convergence of abstract splines. Lecture Notes in Mathematics, 556:155–166, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087404/.

Derridj:1977:RLS

[DT77] M. Derridj and D. Tartakoff. Sur la régularité locale des solutions du problème de Neumann pour ¯∂. (French) []. Lecture Notes in Mathematics, 578:207–216, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091471/.

Dubiel:1976:EEM

[Dub76a] M. Dubiel. Elementary extensions of β-models of A_{2} with fixed height. Lecture Notes in Mathematics, 537:77–81, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096895/.

Dubins:1976:DCP

[Dub76b] Lester E. Dubins. On disintegrations and conditional probabilities. Lecture Notes in Mathematics, 541:53–59, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081041/.

Dubinsky:1979:AP

[Dub79a] Ed Dubinsky. Approximation properties. Lecture Notes in Mathematics, 720:155–183, 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067204/.

Dubinsky:1979:CS

[Dub79b] Ed Dubinsky. Complemented subspaces. Lecture Notes in Mathematics, 720:128–154, 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067203/.

Dubinsky:1979:G

[Dub79c] Ed Dubinsky. Generalities. Lecture Notes in Mathematics, 720: 23–52, 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067200/.

Dubinsky:1979:I

[Dub79d] Ed Dubinsky. Introduction. Lecture Notes in Mathematics, 720: 1–5, 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067198/.

Dubinsky:1979:P

[Dub79e] Ed Dubinsky. Preliminaries. Lecture Notes in Mathematics, 720: 6–22, 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067199/.

Dubinsky:1979:QS

[Dub79f] Ed Dubinsky. Quotient spaces. Lecture Notes in Mathematics, 720:108–127, 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067202/.

Dubinsky:1979:SNF

[Dub79g] Ed Dubinsky. The Structure of Nuclear Fréchet Spaces, volume 720 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 187 pp. LCCN QA3 .L28 no. 720. URL http:

//link.springer.com/book/10.1007/BFb0067197;   http://
www.springerlink.com/content/978-3-540-35191-7.

Dubinsky:1979:S

[Dub79h] Ed Dubinsky. Subspaces. Lecture Notes in Mathematics, 720: 53–107, 1979. CODEN LNMAA2. ISBN 3-540-09504-7 (print), 3-540-35191-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067201/.

Dubuc:1979:CQ

[Dub79i] Eduardo J. Dubuc. Concrete quasitopoi. Lecture Notes in Mathematics, 753:239–254, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061821/.

Duchon:1977:SMR

[Duc77] Jean Duchon. Splines minimizing rotation-invariant semi-norms in Sobolev spaces. Lecture Notes in Mathematics, 571:85–100, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086566/.

Dudley:1979:LLC

[Dud79] R. M. Dudley. Lower layers in R^{2} and convex sets in R^{3} are not GB classes. Lecture Notes in Mathematics, 709:97–102, 1979.

CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071951/.

Duflo:1975:RIG

[Duf75a] Michel Duflo. Représentations irreductibles des groupes semisimples complexes. (French) []. Lecture Notes in Mathematics, 497:26–88, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078012/.

Duflo:1975:SCM

[Duf75b] Michel Duflo. Semigroups of complex measures on a locally compact group. Lecture Notes in Mathematics, 466:56–64, 1975. CO-DEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082197/.

Duflo:1979:PVP

[Duf79a] Michel Duflo. Polynômes de Vogan pour SL(n, C). (French) []. Lecture Notes in Mathematics, 728:64–76, 1979. CODEN LN-MAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063338/.

Duflo:1979:RCI

[Duf79b] Michel Duflo. Représentations de carré intégrable des groupes semi-simples réels. (French) []. Lecture Notes in Mathematics, 710: 22–40, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069971/.

Duistermaat:1975:SPE

[Dui75] J. J. Duistermaat. On the spectrum of positive elliptic operators and periodic bicharacteristics. Lecture Notes in Mathematics, 459:15–22, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074190/.

Duistermaat:1978:LNC

[Dui78] J. J. Duistermaat. The light in the neighborhood of a caustic. Lecture Notes in Mathematics, 677:19–29, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070750/.

Duke:1978:SEP

[Duk78] Richard A. Duke. Some extremal problems for simple twocomplexes. Lecture Notes in Mathematics, 642:148–159, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070372/.

Dummel:1979:SIP

[Düm79] S. Dümmel. On some inverse problems for partial differential equations. Lecture Notes in Mathematics, 703:93–98, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067261/.

Duncan:1978:OCS

[Dun78] T. E. Duncan. Optimal control of stochastic systems in a sphere bundle. Lecture Notes in Mathematics, 695:51–61, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062654/.

Dupont:1978:AFB

[Dup78a] Johan L. Dupont. Applications to flat bundles. Lecture Notes in Mathematics, 640:144–164, 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065373/.

Dupont:1978:CCS

[Dup78b] Johan L. Dupont. Characteristic classes for some classical groups. Lecture Notes in Mathematics, 640:97–113, 1978. CODEN LN-MAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065371/.

Dupont:1978:CWHa

[Dup78c] Johan L. Dupont. The Chern–Weil homomorphism. Lecture Notes in Mathematics, 640:61–70, 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065368/.

Dupont:1978:CWHb

[Dup78d] Johan L. Dupont. The Chern–Weil homomorphism for compact groups. Lecture Notes in Mathematics, 640:114–143, 1978. CO-DEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0065372/.

Dupont:1978:CPB

[Dup78e] Johan L. Dupont. Connections in principal bundles. Lecture Notes in Mathematics, 640:38–60, 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065367/.

Dupont:1978:CCC

[Dup78f] Johan L. Dupont. Curvature and Characteristic Classes, volume 640 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 175 pp. LCCN QA3 .L28 no. 640. URL http:

//link.springer.com/book/10.1007/BFb0065364;   http://
www.springerlink.com/content/978-3-540-35914-2.

Dupont:1978:DFC

[Dup78g] Johan L. Dupont. Differential forms and cohomology. Lecture Notes in Mathematics, 640:1–19, 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065365/.

Dupont:1978:E

[Dup78h] Johan L. Dupont. Errata. Lecture Notes in Mathematics, 640: 165, 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065374.

Dupont:1978:MSR

[Dup78i] Johan L. Dupont. Multiplicativity. The simplicial de Rham complex. Lecture Notes in Mathematics, 640:20–37, 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065366/.

Dupont:1978:SMC

[Dup78j] Johan L. Dupont. Simplicial manifolds. The Chern–Weil homomorphism for BG. Lecture Notes in Mathematics, 640:89–96, 1978. CODEN LNMAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065370/.

Dupont:1978:TBC

[Dup78k] Johan L. Dupont. Topological bundles and classifying spaces. Lecture Notes in Mathematics, 640:71–88, 1978. CODEN LN-MAA2. ISBN 3-540-08663-3 (print), 3-540-35914-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065369/.

Dupont:1979:BCN

[Dup79] Johan L. Dupont. Bounds for characteristic numbers of flat bundles. Lecture Notes in Mathematics, 763:109–119, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088080/.

Durbin:1976:KST

[Dur76a] James Durbin. Kolmogorov–Smirnov tests when parameters are estimated. Lecture Notes in Mathematics, 566:33–44, 1976. CO-DEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096877/.

Duren:1976:ABC

[Dur76b] Peter L. Duren. Asymptotic behavior of coefficients of univalent functions. Lecture Notes in Mathematics, 505:17–23, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081097/.

Duren:1977:S

[Dur77] Peter Duren. Subordination. Lecture Notes in Mathematics, 599: 22–29, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096821/.

Duskin:1979:HDT

[Dus79] J. Duskin. Higher dimensional torsors and the cohomology of topoi: The abelian theory. Lecture Notes in Mathematics, 753: 255–279, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061822/.

Duvaut:1977:CMP

[Duv77] G. Duvaut. Comportement macroscopique d’une plaque perforée périodiquement. (French) []. Lecture Notes in Mathematics, 594: 131–145, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086085/.

Duvaut:1978:HPS

[Duv78] G. Duvaut. Homogénéisation des plaques à structure périodique en théorie non linéaire de von Kármán. (French) []. Lecture Notes in Mathematics, 665:56–69, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061798/.

deVerdiere:1975:PAE

[dV75a] Y. Colin de Verdière. Propriétés asymptotiques de l’équation de la chaleur sur une variété compacte [d’après P. Gilkey]. (French) []. Lecture Notes in Mathematics, 431:58–68, 1975. CODEN LN-MAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066364/.

Diller:1975:IFA

[DV75b] Justus Diller and Helmut Vogel. Intensionale Funktionalinterpretation der Analysis. (German) []. Lecture Notes in Mathematics, 500:56–72, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079547/.

Donsker:1975:LDM

[DV75c] M. D. Donsker and S. R. S. Varadhan. Large deviations for Markov processes and the asymptotic evaluation of certain Markov process expectations for large times. Lecture Notes in Mathematics, 451:82–88, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068580/.

Doran:1979:AIF

[DW79a] Robert S. Doran and Josef Wichmann. Approximate Identities and Factorization in Banach Modules, volume 768 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09725-2 (print), 3-540-38533-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 305 pp. LCCN QA3

.L28 no. 768. URL http://link.springer.com/book/10.1007/
BFb0098610;   http://www.springerlink.com/content/978-
3-540-38533-2.

Doran:1979:AIN

[DW79b] Robert S. Doran and Josef Wichmann. Approximate identities in normed algebras. Lecture Notes in Mathematics, 768:1–88, 1979. CODEN LNMAA2. ISBN 3-540-09725-2 (print), 3-540-38533-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098611/.

Doran:1979:FBM

[DW79c] Robert S. Doran and Josef Wichmann. Factorization in Banach modules. Lecture Notes in Mathematics, 768:89–181, 1979. CO-DEN LNMAA2. ISBN 3-540-09725-2 (print), 3-540-38533-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098612/.

Doran:1979:MAA

[DW79d] Robert S. Doran and Josef Wichmann. More about approximate identities. Lecture Notes in Mathematics, 768:182–236, 1979. CO-DEN LNMAA2. ISBN 3-540-09725-2 (print), 3-540-38533-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0098613/.

Dwyer:1978:THG

[Dwy78a] W. G. Dwyer. The tame homotopy groups of a suspension. Lecture Notes in Mathematics, 658:165–168, 1978. CODEN LN-MAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068714/.

Dwyer:1978:AEE

[Dwy78b] Thomas A. W. Dwyer III. Analytic evolution equations in Banach spaces. Lecture Notes in Mathematics, 645:48–61, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069662/.

Dyckhoff:1976:CMD

[Dyc76] Roy Dyckhoff. Categorical methods in dimension theory. Lecture Notes in Mathematics, 540:220–242, 1976. CODEN LN-MAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080861/.

Dynkin:1976:NAM

[Dyn76] E. B. Dynkin. On a new approach to Markov processes. Lecture Notes in Mathematics, 550:42–62, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077483/.

Dzuraev:1976:SPD

[Džu76] A. Džuraev. Study of partial differential equations by the means of generalized analytical functions. Lecture Notes in Mathematics, 561:29–38, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087623/.

Eades:1977:SNC

[Ead77] Peter Eades. Some new constructions for orthogonal designs using circulants. Lecture Notes in Mathematics, 622:64–75, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069182/.

Earle:1979:SJV

[Ear79] Clifford J. Earle. Some Jacobian varieties which split. Lecture Notes in Mathematics, 747:101–107, 1979. CODEN LN-MAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063962/.

Eastham:1975:RPS

[Eas75] M. S. P. Eastham. Results and problems in the spectral theory of periodic differential equations. Lecture Notes in Mathematics, 448:126–135, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067083/.

Eastham:1976:ASI

[Eas76] M. S. P. Eastham. On the absence of square-integrable solutions of the Sturm–Liouville equation. Lecture Notes in Mathematics, 564:72–77, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087328/.

Easton:1978:CTD

[Eas78] Robert Easton. Chain transitivity and the domain of influence of an invariant set. Lecture Notes in Mathematics, 668:95–102, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101782/.

Eaves:1979:VCP

[Eav79] B. Curtis Eaves. A view of complementary pivot theory (or solving equations with homotopies). Lecture Notes in Mathematics, 730:89–111, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064313/.

Eccles:1978:RFB

[Ecc78] Peter John Eccles. Representing framed bordism classes by manifolds embedded in low codimension. Lecture Notes in Mathematics, 657:150–155, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069231/.

Eckhardt:1975:OPR

[Eck75] Ulrich Eckhardt. On an optimization problem related to minimal surfaces with obstacles. Lecture Notes in Mathematics, 477: 95–101, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079169/.

Eckhaus:1977:MPC

[Eck77] Wiktor Eckhaus. Matching principles and composite expansions. Lecture Notes in Mathematics, 594:146–177, 1977. CODEN LN-MAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086086/.

Edelstein:1975:SAF

[Ede75] Michael Edelstein. On some aspects of fixed point theory in Banach spaces. Lecture Notes in Mathematics, 490:84–90, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081134/.

Edenhofer:1976:SBD

[Ede76] J. Edenhofer. A solution of the biharmonic Dirichlet problem by means of hypercomplex analytic functions. Lecture Notes in Mathematics, 561:192–202, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087636/.

Edgar:1978:RNP

[Edg78] G. A. Edgar. On the Radon–Nikodým-property and martingale convergence. Lecture Notes in Mathematics, 645:62–76, 1978. CO-DEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069663/.

Edrei:1977:EAZ

[Edr77] Albert Edrei. Equidistribution of the arguments of the zeros of Padé numerators. Lecture Notes in Mathematics, 599:30–39, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096822/.

Edwards:1975:QCC

[Edw75a] R. Edwards. A question concerning compact foliations. Lecture Notes in Mathematics, 468:2–4, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082574/.

Edwards:1975:DT

[Edw75b] Robert D. Edwards. Dimension theory, I. Lecture Notes in Mathematics, 438:195–211, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066115/.

Effros:1978:ANC

[Eff78] Edward G. Effros. Aspects of non-commutative order. Lecture Notes in Mathematics, 650:1–40, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067387/.

Everitt:1975:DIP

[EG75] W. N. Everitt and M. Giertz. On the deficiency indices of powers of formally symmetric differential expressions. Lecture Notes in Mathematics, 448:167–181, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067086/.

Egghe:1978:RNP

[Egg78] L. Egghe. On the Radon–Nikodým-property, and related topics in locally convex spaces. Lecture Notes in Mathematics, 645:77– 90, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069664/.

Epstein:1975:GPS

[EGS75] H. Epstein, V. Glaser, and R. Stora. Geometry of the N point P space function of quantum field theory. Lecture Notes in Mathematics, 449:143–162, 1975. CODEN LNMAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062921/.

Eguchi:1979:ATP

[Egu79] Masaaki Eguchi. An application of topological Paley–Wiener theorems to invariant differential equations on symmetric spaces. Lecture Notes in Mathematics, 739:192–206, 1979. CODEN LN-MAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062493/.

Edwards:1976:ATP

[EH76a] David A. Edwards and Harold M. Hastings. The algebraic topology of pro-C. Lecture Notes in Mathematics, 542:172–212, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081088/.

Edwards:1976:B

[EH76b] David A. Edwards and Harold M. Hastings. Background. Lecture Notes in Mathematics, 542:4–55, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081085/.

Edwards:1976:CSH

[EH76c] David A. Edwards and Harold M. Hastings. Čech and Steenrod Homotopy Theories with Applications to Geometric Topology, volume 542 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 296 pp. LCCN QA3 .L28 no. 542. URL http:

//link.springer.com/book/10.1007/BFb0081083;   http://
www.springerlink.com/content/978-3-540-38103-7.

Edwards:1976:GAI

[EH76d] David A. Edwards and Harold M. Hastings. Group actions on infinite dimensional manifolds. Lecture Notes in Mathematics, 542: 233–244, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081090/.

Edwards:1976:HIL

[EH76e] David A. Edwards and Harold M. Hastings. The homotopy inverse limit and its applications to homological algebra. Lecture Notes in Mathematics, 542:129–171, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081087/.

Edwards:1976:I

[EH76f] David A. Edwards and Harold M. Hastings. Introduction. Lecture Notes in Mathematics, 542:1–3, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081084/.

Edwards:1976:MSP

[EH76g] David A. Edwards and Harold M. Hastings. The model structure on pro-spaces. Lecture Notes in Mathematics, 542:56–128, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081086/.

Edwards:1976:PHT

[EH76h] David A. Edwards and Harold M. Hastings. Proper homotopy theory. Lecture Notes in Mathematics, 542:213–232, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081089/.

Edwards:1976:SOQ

[EH76i] David A. Edwards and Harold M. Hastings. Some open questions. Lecture Notes in Mathematics, 542:279–280, 1976. CODEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081092.

Edwards:1976:SHT

[EH76j] David A. Edwards and Harold M. Hastings. Steenrod homotopy theory. Lecture Notes in Mathematics, 542:245–278, 1976. CO-DEN LNMAA2. ISBN 3-540-07863-0 (print), 3-540-38103-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081091/.

Eggleton:1978:NEP

[EH78] R. B. Eggleton and A. Hartman. A note on equidistant permutation arrays. Lecture Notes in Mathematics, 686:136–147, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062526/.

Eggleton:1979:GTR

[EH79a] R. B. Eggleton and D. A. Holton. The graph of type (0, ∞, ∞) realizations of a graphic sequence. Lecture Notes in Mathematics, 748:41–54, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102683/.

Eggleton:1979:GS

[EH79b] R. B. Eggleton and D. A. Holton. Graphic sequences. Lecture Notes in Mathematics, 748:1–10, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0102679/.

Eichler:1977:TF

[Eic77] M. Eichler. Theta functions over Q and over Q(\sqrt{q}). Lecture Notes in Mathematics, 627:197–225, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065302/.

Eifrig:1976:SBE

[Eif76] B. Eifrig. Ein Nicht-Standard-Beweis für die Existenz eines Liftings. (German) []. Lecture Notes in Mathematics, 541:133– 135, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081046/.

Eisele:1976:UPM

[Eis76] K.-Th. Eisele. On the uniqueness of preimages of measures. Lecture Notes in Mathematics, 541:5–12, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081035/.

Eisenbud:1977:EFR

[Eis77a] David Eisenbud. Enriched free resolutions and change of rings. Lecture Notes in Mathematics, 586:1–8, 1977. CODEN LN-MAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087115/.

Eisenbud:1977:SDP

[Eis77b] David Eisenbud. Solution du problème de Serre par Quillen– Suslin. (French) []. Lecture Notes in Mathematics, 586:9–19, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087116/.

Esper:1977:PR

[EK77] Norbert Esper and Adalbert Kerber. Permutrization of representations. Lecture Notes in Mathematics, 579:267–280, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090022/.

Eklof:1975:CLF

[Ekl75] Paul C. Eklof. Categories of local functors. Lecture Notes in Mathematics, 498:91–116, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080976/.

Eklof:1977:MLA

[Ekl77] Paul C. Eklof. Methods of logic in abelian group theory. Lecture Notes in Mathematics, 616:251–269, 1977. CODEN LN-MAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068201/.

ElKaroui:1978:AOP

[El 78] Nicole El Karoui. Arrêt optimal prévisible. (French) []. Lecture Notes in Mathematics, 695:1–11, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062650/.

Eisenbud:1976:TDF

[EL76] David Eisenbud and Harold I. Levine. The topological degree of a finite C^{∞} map germ. Lecture Notes in Mathematics, 525: 90–98, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077845/.

Erdelyi:1977:ASD

[EL77a] Ivan Erdélyi and Ridgley Lange. Asymptotic spectral decompositions. Lecture Notes in Mathematics, 623:50–72, 1977. CODEN LNMAA2. ISBN 3-540-08525-4 (print), 3-540-36991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065480/.

Erdelyi:1977:DO

[EL77b] Ivan Erdélyi and Ridgley Lange. Decomposable operators. Lecture Notes in Mathematics, 623:73–105, 1977. CODEN LN-MAA2. ISBN 3-540-08525-4 (print), 3-540-36991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065481/.

Erdelyi:1977:GSD

[EL77c] Ivan Erdélyi and Ridgley Lange. The general spectral decomposition. Lecture Notes in Mathematics, 623:36–49, 1977. CODEN LNMAA2. ISBN 3-540-08525-4 (print), 3-540-36991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065479/.

Erdelyi:1977:I

[EL77d] Ivan Erdélyi and Ridgley Lange. Introduction. Lecture Notes in Mathematics, 623:1–2, 1977. CODEN LNMAA2. ISBN 3-540-08525-4 (print), 3-540-36991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065477.

Erdelyi:1977:IS

[EL77e] Ivan Erdélyi and Ridgley Lange. Invariant subspaces. Lecture Notes in Mathematics, 623:3–35, 1977. CODEN LNMAA2. ISBN 3-540-08525-4 (print), 3-540-36991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065478/.

Erdelyi:1977:SDB

[EL77f] Ivan Erdélyi and Ridgley Lange. Spectral Decompositions on Banach Spaces, volume 623 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08525-4 (print), 3-540-36991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 122 pp. LCCN QA3

.L28 no. 623. URL http://link.springer.com/book/10.1007/
BFb0065476;   http://www.springerlink.com/content/978-
3-540-36991-2.

Elencwajg:1978:SHF

[Ele78] G. Elencwajg. Submersions holomorphes a fibres isomorphes. Lecture Notes in Mathematics, 670:209–223, 1978. CODEN LN-MAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064402/.

Elie:1979:FHP

[Éli79a] Laure Élie. Fonctions harmoniques positives sur le groupe affine. (French) []. Lecture Notes in Mathematics, 706:96–110, 1979. CO-DEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063117/.

Elie:1979:QPD

[Éli79b] Laure Élie. Quelques propriétés du noyau potentiel associé à une marche aléatoire. (French) []. Lecture Notes in Mathematics, 706:88–95, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063116/.

Ellis:1975:ALT

[Ell75] Richard S. Ellis. Asymptotics and limit theorems for the linearized Boltzmann equation. Lecture Notes in Mathematics, 451: 143–151, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068584/.

Ellis:1978:CF

[Ell78] Robert Ellis. Cohomology of flows. Lecture Notes in Mathematics, 668:103–116, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101783/.

Elliott:1979:TOG

[Ell79] George A. Elliott. On totally ordered groups, and K_{0}. Lecture Notes in Mathematics, 734:1–49, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103152/.

Elworthy:1978:DIM

[Elw78] K. D. Elworthy. Differential invariants of measures on Banach spaces. Lecture Notes in Mathematics, 644:159–187, 1978. CO-DEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0066845/.

ElKaroui:1977:CTT

[EM77] Nicole El Karoui and P.-A. Meyer. Les changements de temps en théorie générale des processus. (French) []. Lecture Notes in Mathematics, 581:65–78, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087189/.

Emch:1978:MDC

[Emc78] Gérard G. Emch. Minimal dilations of CP-flows. Lecture Notes in Mathematics, 650:156–159, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067398/.

Egorov:1976:LTL

[EN76] V. A. Egorov and V. B. Nevzorov. Limit theorems for linear combinations of order statistics. Lecture Notes in Mathematics, 550:63–79, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077484/.

Erdos:1979:SDR

[EN79] Paul Erdős and Melvyn B. Nathanson. Systems of distinct representatives and minimal bases in additive number theory. Lecture Notes in Mathematics, 751:89–107, 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062704/.

Enting:1978:CAG

[Ent78] I. G. Enting. The combinatorics of algebraic graph theory in theoretical physics. Lecture Notes in Mathematics, 686:148–156, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062527/.

Epstein:1975:FAL

[Eps75] David Epstein. Foliations with all leaves compact. Lecture Notes in Mathematics, 468:1–2, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082573.

Epstein:1977:TSF

[Eps77] D. B. A. Epstein. A topology for the space of foliations. Lecture Notes in Mathematics, 597:132–150, 1977. CODEN LN-MAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085352/.

Epstein:1979:D

[Eps79a] Richard L. Epstein. ⇒^{∗} D. Lecture Notes in Mathematics, 759: 33–46, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067138/.

Epstein:1979:CDL

[Eps79b] Richard L. Epstein. Countable distributive lattices ⇒^{∗} D. Lecture Notes in Mathematics, 759:80–94, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067142/.

Epstein:1979:DTA

[Eps79c] Richard L. Epstein. Degree theory and analysis. Lecture Notes in Mathematics, 759:115–137, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067145/.

Epstein:1979:DUS

[Eps79d] Richard L. Epstein. Degrees of Unsolvability: Structure and Theory, volume 759 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 240 pp. LCCN QA3 .L28 no. 759. URL http:

//link.springer.com/book/10.1007/BFb0067135;   http://
www.springerlink.com/content/978-3-540-38480-9.

Epstein:1979:FDL

[Eps79e] Richard L. Epstein. Finite distributive lattices ⇒^{∗} D. Lecture Notes in Mathematics, 759:65–72, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067140/.

Epstein:1979:HQ

[Eps79f] Richard L. Epstein. The homogeneity questions. Lecture Notes in Mathematics, 759:105–114, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067144/.

Epstein:1979:ID

[Eps79g] Richard L. Epstein. An introduction to degrees ≤ 0′. Lecture Notes in Mathematics, 759:138–168, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067146/.

Epstein:1979:IDU

[Eps79h] Richard L. Epstein. An introduction to degrees of unsolvability. Lecture Notes in Mathematics, 759:1–25, 1979. CODEN LN-MAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067136/.

Epstein:1979:LO

[Eps79i] Richard L. Epstein. Linear orderings ⇒^{∗} D. Lecture Notes in Mathematics, 759:73–79, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067141/.

Epstein:1979:RTT

[Eps79j] Richard L. Epstein. Relativizing, a tree of trees, the jump operator. Lecture Notes in Mathematics, 759:95–104, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067143/.

Epstein:1979:UTDa

[Eps79k] Richard L. Epstein. The undecidability of the theory of degrees. Lecture Notes in Mathematics, 759:26–32, 1979. CODEN LN-MAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067137/.

Epstein:1979:UTDb

[Eps79l] Richard L. Epstein. The undecidability of the theory of degrees ≤ 0′. Lecture Notes in Mathematics, 759:169–185, 1979. CODEN LNMAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067147/.

Epstein:1979:VFD

[Eps79m] Richard L. Epstein. Various finite distributive lattices ⇒^{∗} D. Lecture Notes in Mathematics, 759:47–64, 1979. CODEN LN-MAA2. ISBN 3-540-09710-4 (print), 3-540-38480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067139/.

Eymard:1975:MAD

[ER75] Pierre Eymard and Bernard Roynette. Marches aléatoires sur le dual de su(2). (French) []. Lecture Notes in Mathematics, 497: 108–152, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078014/.

Epstein:1977:SCF

[ER77] D. B. A. Epstein and H. Rosenberg. Stability of compact foliations. Lecture Notes in Mathematics, 597:151–160, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085353/.

Erard:1978:I

[Era78] P. J. Erard. On immersions CP^{n} ↪ R^{4n−2α(n)}. Lecture Notes in Mathematics, 673:106–115, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064691/.

Erdelyi:1975:FIG

[Erd75a] Arthur Erdélyi. Fractional integrals of generalized functions. Lecture Notes in Mathematics, 457:151–170, 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067103/.

Erdos:1975:PRD

[Erd75b] P. Erdős. Problems and results on Diophantine approximations (II). Lecture Notes in Mathematics, 475:89–99, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074258/.

Erdos:1977:PRC

[Erd77] Paul Erdős. Problems and results on combinatorial number theory III. Lecture Notes in Mathematics, 626:43–72, 1977. CODEN LNMAA2. ISBN 3-540-08529-7 (print), 3-540-35985-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063064/.

Erdos:1978:SEP

[Erd78] P. Erdős. Some extremal problems on families of graphs and related problems. Lecture Notes in Mathematics, 686:13–21, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062512/.

Ermine:1976:HCF

[Erm76] Jean-Louis Ermine. Holomorphe convexité des fibrés analytiques à base de Stein et à fibre dans C. (French) []. Lecture Notes in Mathematics, 524:96–105, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078001/.

Ermine:1978:CSE

[Erm78] Jean-Louis Ermine. Conjecture de Serre et espaces hyperconvexes. (French) []. Lecture Notes in Mathematics, 670:124–139, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064398/.

Everitt:1976:OPD

[ES76] William N. Everitt and Brian D. Sleeman, editors. Ordinary and Partial Differential Equations: Proceedings of the Fourth Conference Held at Dundee, Scotland March 30–April 2, 1976, volume 564 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 564. URL http://link.springer.com/book/10.1007/

BFb0087320;   http://www.springerlink.com/content/978-
3-540-37517-3.

Essen:1977:EPF

[ES77] Matts Essén and Daniel F. Shea. An extremal problem in function theory. Lecture Notes in Mathematics, 599:40–45, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096823/.

Emery:1979:DER

[ÉS79] Michel Émery and C. Stricker. Démonstration élémentaire d’un résultat d’Azema et Jeulin. (French) []. Lecture Notes in Mathematics, 721:116–117, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070853.

Essen:1975:PDI

[Ess75a] Prof. Matts R. Essén. A cos πλ-problem and a differential inequality. Lecture Notes in Mathematics, 467:40–52, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094974/.

Essen:1975:E

[Ess75b] Prof. Matts R. Essén. The cos πλ-theorem. Lecture Notes in Mathematics, 467:10–17, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094971/.

Essen:1975:TPC

[Ess75c] Prof. Matts R. Essén. The cos πλ Theorem: With a paper by Christer Borell, volume 467 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 112 pp. LCCN QA3 .L28 no. 467. URL http:

//link.springer.com/book/10.1007/BFb0094969;   http://
www.springerlink.com/content/978-3-540-37513-5.

Essen:1975:C

[Ess75d] Prof. Matts R. Essén. A counterexample. Lecture Notes in Mathematics, 467:53–55, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094975/.

Essen:1975:GAH

[Ess75e] Prof. Matts R. Essén. A generalization of the Ahlfors–Heins theorem. Lecture Notes in Mathematics, 467:18–30, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094972/.

Essen:1975:HKN

[Ess75f] Prof. Matts R. Essén. The Hellsten–Kjellberg–Norstad inequality. Lecture Notes in Mathematics, 467:1–9, 1975. CODEN LN-MAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0094970/.

Essen:1975:TTB

[Ess75g] Prof. Matts R. Essén. On two theorems of A. Baernstein. Lecture Notes in Mathematics, 467:73–89, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094978/.

Essen:1975:PC

[Ess75h] Prof. Matts R. Essén. The Paley conjecture. Lecture Notes in Mathematics, 467:31–39, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094973/.

Essen:1975:PT

[Ess75i] Prof. Matts R. Essén. Proof of theorem 5.2. Lecture Notes in Mathematics, 467:56–64, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094976/.

Essen:1975:TCT

[Ess75j] Prof. Matts R. Essén. Two consequences of a theorem of A. Baernstein. Lecture Notes in Mathematics, 467:65–72, 1975. CODEN LNMAA2. ISBN 3-540-07176-8 (print), 3-540-37513-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094977/.

Esser:1976:ADS

[Ess76] Henning Esser. Abschätzungen durch Stetigkeitsmoduli bei Folgen von linearen Funktionalen. (German) []. Lecture Notes in Mathematics, 556:184–190, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087406/.

Estabrook:1976:SON

[Est76] Frank B. Estabrook. Some old and new techniques for the practical use of exterior differential forms. Lecture Notes in Mathematics, 515:136–161, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081166/.

Eymard:1979:TFS

[ET79] Pierre Eymard and Marianne Terp. La transformation de Fourier et son inverse sur le groupe des ax+b d’un corps local. (French) []. Lecture Notes in Mathematics, 739:207–248, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062494/.

Eymard:1975:AHG

[ETFS75] Pierre Eymard, Reiji Takahashi, Jacques Faraut, and Gérard Schiffmann, editors. Analyse Harmonique sur les Groupes de Lie: Séminaire Nancy–Strasbourg 1973–75, volume 497 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 497. URL http://link.springer.com/book/10.1007/
BFb0078009;   http://www.springerlink.com/content/978-
3-540-38047-4.

Eymard:1979:AHG

[ETFS79] Pierre Eymard, Reiji Takahashi, Jacques Faraut, and Gérard Schiffmann, editors. Analyse Harmonique sur les Groupes de Lie II: Séminaire Nancy–Strasbourg 1976–78, volume 739 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 739. URL http://link.springer.com/book/10.1007/
BFb0062485;   http://www.springerlink.com/content/978-
3-540-35023-1.

Efremovic:1977:NRU

[EV77] V. A. Efremovič and A. G. Vaĭnšteĭn. New results in uniform topology. Lecture Notes in Mathematics, 609:66–80, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068673/.

Evans:1975:SE

[Eva75] W. D. Evans. Sobolev embeddings. Lecture Notes in Mathematics, 448:136–147, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067084/.

Evans:1976:LPD

[Eva76] W. D. Evans. On limit-point and Dirichlet-type results for secondorder differential expressions. Lecture Notes in Mathematics, 564: 78–92, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087329/.

Evans:1977:GPI

[Eva77] E. Graham Evans, Jr. A generalized principal ideal theorem with applications to intersection theory. Lecture Notes in Mathematics, 586:20–24, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087117/.

Everitt:1975:STD

[Eve75a] Prof. William N. Everitt, editor. Spectral Theory and Differential Equations: Proceedings of the Symposium held at Dundee, Scotland, 1–19 July, 1974, volume 448 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 448. URL http:

//link.springer.com/book/10.1007/BFb0067078;   http://
www.springerlink.com/content/978-3-540-37444-2.

Everitt:1975:IIS

[Eve75b] W. N. Everitt. Integral inequalities and spectral theory. Lecture Notes in Mathematics, 448:148–166, 1975. CODEN LN-MAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067085/.

Everitt:1976:STW

[Eve76] W. N. Everitt. Spectral theory of the Wirtinger inequality. Lecture Notes in Mathematics, 564:93–105, 1976. CODEN LN-MAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087330/.

Evstigneev:1978:NGR

[Evs78] I. V. Evstigneev. A note on Gibbs representation. Lecture Notes in Mathematics, 653:194–202, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070093/.

Eades:1976:IFS

[EW76] Peter Eades and Jennifer Seberry Wallis. An infinite family of skew weighing matrices. Lecture Notes in Mathematics, 560:27– 40, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097365/.

ElKaroui:1977:TGC

[EW77] Nicole El Karoui and Gérard Weidenfeld. Théorie générale et changement de temps. (French) []. Lecture Notes in Mathematics, 581:79–108, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087190/.

Elben:1978:PAS

[EW78] W. Elben and W. Wolff. Problems in applying the SOR-method to the solution of the Maxwell’s time dependent equations. Lecture Notes in Mathematics, 679:25–40, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067864/.

Ewing:1979:ORE

[Ewi79] John Ewing. Obstructions to realizing equivariant Witt classes. Lecture Notes in Mathematics, 741:272–282, 1979. CODEN LN-MAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062144/.

Embry-Wardrop:1978:WTS

[EWL78] Mary Embry-Wardrop and Alan Lambert. Weighted translation semigroups on L^{2}[0, ∞]. Lecture Notes in Mathematics, 693:87– 91, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064663/.

Eades:1977:NAE

[EWW77] Peter Eades, Jennifer Seberry Wallis, and Nicholas Wormald. A note on asymptotic existence results for orthogonal designs. Lecture Notes in Mathematics, 622:76–90, 1977. CODEN LN-MAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069183/.

Eymard:1975:ITG

[Eym75] Pierre Eymard. Initiation à la théorie des groupes moyennables. (French) []. Lecture Notes in Mathematics, 497:89–107, 1975. CO-DEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0078013/.

Fabera:1979:EIP

[Fáb79] Jiří Fábera, editor. Equadiff IV: Proceedings, Prague, August 22–26, 1977, volume 703 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 703. URL http:

//link.springer.com/book/10.1007/BFb0067248;   http://
www.springerlink.com/content/978-3-540-35519-9.

Faith:1979:BPR

[Fai79a] Carl Faith. Bounded prime rings, pseudo-Frobenius rings, the Jacobson radical of a ring. Lecture Notes in Mathematics, 700: 218–223, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063469/.

Faith:1979:IQR

[Fai79b] Carl Faith. Injective quotient rings of commutative rings. Lecture Notes in Mathematics, 700:151–203, 1979. CODEN LN-MAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063466/.

Falkowski:1979:IDP

[Fal79] B.-J. Falkowski. Infinitely divisible positive functions on SO(3)R^{3}. Lecture Notes in Mathematics, 706:111–115, 1979. CODEN LN-MAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063118/.

Faraut:1975:MNO

[Far75a] Jacques Faraut. Moyennabilité et normes d’opérateurs de convolution. (French) []. Lecture Notes in Mathematics, 497:153– 163, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078015/.

Faraut:1975:NSH

[Far75b] Jacques Faraut. Noyaux sphériques sur un hyperboloïde à une nappe. (French) []. Lecture Notes in Mathematics, 497:172– 210, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078017/.

Faraut:1975:SGF

[Far75c] Jacques Faraut. Semi-groupes de Feller invariants sur les espaces homogènes non moyennables. (French) []. Lecture Notes in Mathematics, 497:164–171, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078016/.

Faris:1975:FO

[Far75d] William G. Faris. Forms and operators. Lecture Notes in Mathematics, 433:1–47, 1975. CODEN LNMAA2. ISBN 3-540-07030-3 (print), 3-540-37317-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068568/.

Faris:1975:M

[Far75e] William G. Faris. Moments. Lecture Notes in Mathematics, 433: 99–108, 1975. CODEN LNMAA2. ISBN 3-540-07030-3 (print), 3-540-37317-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068571/.

Faris:1975:OD

[Far75f] William G. Faris. Operator domains. Lecture Notes in Mathematics, 433:48–82, 1975. CODEN LNMAA2. ISBN 3-540-07030-3 (print), 3-540-37317-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068569/.

Faris:1975:SAE

[Far75g] William G. Faris. Self-adjoint extensions. Lecture Notes in Mathematics, 433:83–98, 1975. CODEN LNMAA2. ISBN 3-540-07030-3 (print), 3-540-37317-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068570/.

Faris:1975:SAO

[Far75h] William G. Faris. Self-Adjoint Operators, volume 433 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07030-3 (print), 3-540-37317-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 115 pp. LCCN QA3

.L28 no. 433. URL http://link.springer.com/book/10.1007/
BFb0068567;   http://www.springerlink.com/content/978-
3-540-37317-9.

Farrell:1976:CZP

[Far76a] Roger H. Farrell. The construction of zonal polynomials. Lecture Notes in Mathematics, 520:231–304, 1976. CODEN LN-MAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079675/.

Farrell:1976:CSM

[Far76b] Roger H. Farrell. Cross sections and maximal invariants. Lecture Notes in Mathematics, 520:158–184, 1976. CODEN LN-MAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079673/.

Farrell:1976:DM

[Far76c] Roger H. Farrell. Decompositions of matrices. Lecture Notes in Mathematics, 520:116–137, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079671/.

Farrell:1976:EUD

[Far76d] Roger H. Farrell. Examples using differential forms. Lecture Notes in Mathematics, 520:138–157, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079672/.

Farrell:1976:IBS

[Far76e] Roger H. Farrell. Introduction and brief survey. Lecture Notes in Mathematics, 520:1–12, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079664/.

Farrell:1976:IMM

[Far76f] Roger H. Farrell. Invariant measures on manifolds. Lecture Notes in Mathematics, 520:100–115, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079670/.

Farrell:1976:LCG

[Far76g] Roger H. Farrell. Locally compact groups and Haar measure. Lecture Notes in Mathematics, 520:32–55, 1976. CODEN LN-MAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079666/.

Farrell:1976:MED

[Far76h] Roger H. Farrell. Manifolds and exterior differential forms. Lecture Notes in Mathematics, 520:73–99, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079669/.

Farrell:1976:NWD

[Far76i] Roger H. Farrell. The noncentral Wishart density function. Lecture Notes in Mathematics, 520:64–72, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079668/.

Farrell:1976:PUZ

[Far76j] Roger H. Farrell. Problems for users of zonal polynomials. Lecture Notes in Mathematics, 520:305–327, 1976. CODEN LN-MAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079676/.

Farrell:1976:RVT

[Far76k] Roger H. Farrell. Random variable techniques. Lecture Notes in Mathematics, 520:185–230, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079674/.

Farrell:1976:TMC

[Far76l] Roger H. Farrell. Techniques of Multivariate Calculation, volume 520 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 337 pp. LCCN QA3 .L28 no. 520. URL http:

//link.springer.com/book/10.1007/BFb0079663;   http://
www.springerlink.com/content/978-3-540-38227-0.

Farrell:1976:T

[Far76m] Roger H. Farrell. Transforms. Lecture Notes in Mathematics, 520: 13–31, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079665/.

Farrell:1976:WP

[Far76n] Roger H. Farrell. Wishart’s paper. Lecture Notes in Mathematics, 520:56–63, 1976. CODEN LNMAA2. ISBN 3-540-07695-6 (print), 3-540-38227-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079667/.

Farkas:1979:SPT

[Far79] H. M. Farkas. Singular points of theta functions, quadric relations and holomorphic differentials with prescribed zeros. Lecture Notes in Mathematics, 747:108–122, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063963/.

Fattorini:1975:ECL

[Fat75] H. O. Fattorini. Exact controllability of linear systems in infinite dimensional spaces. Lecture Notes in Mathematics, 446:166– 183, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070602/.

Fernique:1975:EEP

[FCG75] X. M. (Xavier M.) Fernique, Jean Pierre Conze, and J. M. (Joseph Mark) Gani, editors. École d’été de probabilités de Saint-Flour IV — 1974. (French) [Saint-Flour Summer School on Probability IV — 1974], volume 480 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 0-387-07396-5, 3-540-07396-5 (print), 3-540-37600-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 480; QA3 .L28 no. 480 QA273.A1. URL http:

//link.springer.com/book/10.1007/BFb0080189;   http://
www.springerlink.com/content/978-3-540-37600-2.

Feyel:1976:RES

[Fd76] D. Feyel and A. de La Pradelle. Le rôle des espaces de Sobolev en topologie fine. (French) []. Lecture Notes in Mathematics, 563: 43–61, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087571/.

Feyel:1979:PTE

[Fd79] D. Feyel and A. de La Pradelle. Produits tensoriels d’espaces de Banach adaptés. Applications à la fonction de Green. (French) []. Lecture Notes in Mathematics, 713:56–66, 1979. CODEN LN-MAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071320/.

Feyel:1976:CDA

[FdlP76] D. Feyel and A. de la Pradelle. Cônes en dualité. Applications aux fonctions de Green. (French) []. Lecture Notes in Mathematics, 563:62–99, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087572/.

Feyel:1979:DQR

[FdlP79] D. Feyel and Arnaud de la Pradelle. Dualité des quasi-resolvantes de Ray. (French) []. Lecture Notes in Mathematics, 713:67– 88, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071321/.

Fedida:1978:TFA

[Fed78] Edmond Fedida. Sur la théorie des feuilletages associée au repère mobile: cas des feuilletages de Lie. (French) [On the theory of foliations associated with the mobile marker: the case of Lie foliations]. Lecture Notes in Mathematics, 652:183–195, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063512/.

Feferman:1975:IEL

[Fef75a] Solomon Feferman. Impredicativity of the existence of the largest divisible subgroup of an Abelian p-group. Lecture Notes in Mathematics, 498:117–130, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080977/.

Feferman:1975:LAE

[Fef75b] Solomon Feferman. A language and axioms for explicit mathematics. Lecture Notes in Mathematics, 450:87–139, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062852/.

Feferman:1975:NET

[Fef75c] Solomon Feferman. Non-extensional type-free theories of partial operations and classifications, I. Lecture Notes in Mathematics, 500:73–118, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079548/.

Feinsilver:1978:BOT

[Fei78a] Philip J. Feinsilver. Basic operator theory. Lecture Notes in Mathematics, 696:3–7, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064534/.

Feinsilver:1978:CR

[Fei78b] Philip J. Feinsilver. Concluding remarks. Lecture Notes in Mathematics, 696:107–108, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064542.

Feinsilver:1978:DTM

[Fei78c] Philip J. Feinsilver. Discrete theory. Martingales. Stochastic integrals. Lecture Notes in Mathematics, 696:63–76, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064539/.

Feinsilver:1978:HPS

[Fei78d] Philip J. Feinsilver. Hamiltonian processes and stochastic processes. Lecture Notes in Mathematics, 696:98–106, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064541/.

Feinsilver:1978:I

[Fei78e] Philip J. Feinsilver. Introduction. Lecture Notes in Mathematics, 696:1–2, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064533.

Feinsilver:1978:LT

[Fei78f] Philip J. Feinsilver. Limit theorems. Lecture Notes in Mathematics, 696:55–62, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064538/.

Feinsilver:1978:MTa

[Fei78g] Philip J. Feinsilver. Moment theory. Lecture Notes in Mathematics, 696:8–21, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064535/.

Feinsilver:1978:MTb

[Fei78h] Philip J. Feinsilver. Multidimensional theory. Lecture Notes in Mathematics, 696:77–97, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064540/.

Feinsilver:1978:OT

[Fei78i] Philip J. Feinsilver. Orthogonal theory. Lecture Notes in Mathematics, 696:22–34, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064536/.

Feinsilver:1978:SPE

[Fei78j] Philip J. Feinsilver. The standard processes. Examples. Lecture Notes in Mathematics, 696:35–54, 1978. CODEN LNMAA2. ISBN 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064537/.

Feinsilver:1978:SFP

[Fei78k] Philip J. (Philip Joel) Feinsilver. Special Functions, Probability Semigroups, and Hamiltonian Flows, volume 696 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 0-387-09100-9, 3-540-09100-9 (print), 3-540-35554-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 112 pp. LCCN QA3

.L28 no. 696. URL http://link.springer.com/book/10.1007/
BFb0064532;   http://www.springerlink.com/content/978-
3-540-35554-0.

Fekete:1979:SSB

[Fek79] Otto Fekete. On some subclasses of Bazilevič functions. Lecture Notes in Mathematics, 743:268–273, 1979. CODEN LN-MAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079499/.

Felscher:1975:KKB

[Fel75] Walter Felscher. Kombinatorische Konstruktionen mit Beweisen und Schnittelimination. (German) []. Lecture Notes in Mathematics, 500:119–151, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079549/.

Felgner:1976:PSG

[Fel76] Ulrich Felgner. Das Problem von Souslin für geordnete algebraische Strukturen. (German) []. Lecture Notes in Mathematics, 537:83–107, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096896/.

Fell:1977:AIP

[Fel77a] J. M. G. Fell. The abstract inducing process and abstract imprimitivity. Lecture Notes in Mathematics, 582:28–98, 1977. CODEN LNMAA2. ISBN 3-540-08147-X (print), 3-540-69977-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086206/.

Fell:1977:BAB

[Fel77b] J. M. G. Fell. Banach *-algebraic bundles. Lecture Notes in Mathematics, 582:99–192, 1977. CODEN LNMAA2. ISBN 3-540-08147-X (print), 3-540-69977-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086207/.

Fell:1977:IRB

[Fel77c] J. M. G. Fell. Induced Representations and Banach^{∗}-Algebraic Bundles: with an Appendix due to A. Douady and L. Dal Soglio– Hérault, volume 582 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08147-X (print), 3-540-69977-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 349 pp. LCCN QA3 .L28 no. 582. URL http:

//link.springer.com/book/10.1007/BFb0086204;   http://
www.springerlink.com/content/978-3-540-69977-4.

Fell:1977:IRI

[Fel77d] J. M. G. Fell. Induced representations and imprimitivity for Banach *-algebraic bundles. Lecture Notes in Mathematics, 582:193– 323, 1977. CODEN LNMAA2. ISBN 3-540-08147-X (print), 3-540-69977-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086208/.

Fell:1977:I

[Fel77e] J. M. G. Fell. Introduction. Lecture Notes in Mathematics, 582: 1–27, 1977. CODEN LNMAA2. ISBN 3-540-08147-X (print), 3-540-69977-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086205/.

Feldman:1979:RPP

[Fel79] J. Feldman. Reparametrization of probability-preserving n-flows. Lecture Notes in Mathematics, 729:34–42, 1979. CODEN LN-MAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063281/.

Fenstad:1975:CTA

[Fen75] Jens Erik Fenstad. Computation theories: an axiomatic approach to recursion on general structures. Lecture Notes in Mathematics, 499:143–168, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079420/.

Fenn:1979:TLD

[Fen79] Roger Fenn, editor. Topology of Low-Dimensional Manifolds: Proceedings of the Second Sussex Conference, 1977, volume 722 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 722. URL http://link.springer.com/book/10.1007/
BFb0063182;   http://www.springerlink.com/content/978-
3-540-35186-3.

Fernique:1975:RTF

[Fer75] X. Fernique. Regularité des trajectoires des fonctions aléatoires gaussiennes. (French) []. Lecture Notes in Mathematics, 480:1– 96, 1975. CODEN LNMAA2. ISBN 3-540-07396-5 (print), 3-540-37600-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080190/.

Fernique:1976:EPG

[Fer76a] X. Fernique. Évaluations de processus Gaussiens composés. (French) []. Lecture Notes in Mathematics, 526:67–83, 1976. CO-DEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082342/.

Fernstrom:1976:BPE

[Fer76b] Claes Fernstrøm. Bounded point evaluations and approximation in L^{P} by analytic functions. Lecture Notes in Mathematics, 512: 65–68, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080024/.

Ferrand:1977:MPT

[Fer77] Daniel Ferrand. Les modules projectifs de type fini sur un anneau de polynômes sur un corps sont libres [d’après Quillen [11] et Suslin [16]]. (French) []. Lecture Notes in Mathematics, 567: 202–221, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096071/.

Fernique:1978:RFO

[Fer78a] X. Fernique. Random functions and Orlicz’s method regularity of paths and limit properties. Lecture Notes in Mathematics, 656:31–36, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068808/.

Ferry:1978:SGA

[Fer78b] S. Ferry. On a space of group actions. Lecture Notes in Mathematics, 664:83–86, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061692/.

Ferrand:1979:STC

[Fer79] Daniel Ferrand. Set theoretical complete intersections in characteristic p > 0. Lecture Notes in Mathematics, 732:82–89, 1979.

CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066638/.

Feshbach:1978:TCC

[Fes78] Mark Feshbach. The transfer and characteristic classes. Lecture Notes in Mathematics, 657:156–162, 1978. CODEN LN-MAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069232/.

Feyel:1978:ECR

[Fey78a] D. Feyel. Espaces complètement réticules de pseudo-noyaux. Applications aux résolvantes et aux semi-groupes complexes. (French) []. Lecture Notes in Mathematics, 681:54–80, 1978. CODEN LN-MAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065868/.

Feyel:1978:EBF

[Fey78b] D. Feyel. Espaces de Banach fonctionnels adaptés quasi-topologies et balayage. (French) []. Lecture Notes in Mathematics, 681:81– 102, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065869/.

Feyel:1979:CCP

[Fey79a] D. Feyel. Complements sur la convergence presque sûre des familles resolues. (French) []. Lecture Notes in Mathematics, 713: 51–55, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071319/.

Feyel:1979:PPD

[Fey79b] D. Feyel. Propriétés de permanence du domaine d’un générateur infinitesimal. (French) []. Lecture Notes in Mathematics, 713:38– 50, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071318/.

Felgner:1978:WSP

[FF78] U. Felgner and T. B. Flannagan. Wellordered subclasses of proper classes. Lecture Notes in Mathematics, 669:1–14, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103097/.

Fossum:1975:C

[FGR75a] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Coherence. Lecture Notes in Mathematics, 456:24–34, 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065407/.

Fossum:1975:DDF

[FGR75b] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Dominant dimension of finite algebras. Lecture Notes in Mathematics, 456:104–112, 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065411/.

Fossum:1975:DGP

[FGR75c] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Duality and the Gorenstein property. Lecture Notes in Mathematics, 456:35–51, 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065408/.

Fossum:1975:G

[FGR75d] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Generalities. Lecture Notes in Mathematics, 456:3–23, 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065406/.

Fossum:1975:GM

[FGR75e] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Gorenstein modules. Lecture Notes in Mathematics, 456: 87–103, 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065410/.

Fossum:1975:HD

[FGR75f] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Homological dimensions in A[leftnormalfactorsemidirectproduct]F. Lecture Notes in Mathematics, 456:52–86, 1975. CODEN LN-MAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065409/.

Fossum:1975:P

[FGR75g] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Preliminaries. Lecture Notes in Mathematics, 456:1–2, 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065405.

Fossum:1975:RDF

[FGR75h] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Representation dimension of finite algebras. Lecture Notes in Mathematics, 456:113–116, 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065412/.

Fossum:1975:TEA

[FGR75i] Dr. Robert M. Fossum, Dr. Phillip A. Griffith, and Dr. Idun Reiten. Trivial Extensions of Abelian Categories: Homological Algebra of Trivial Extensions of Abelian Categories with Applications to Ring Theory, volume 456 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07159-8 (print), 3-540-37487-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 122 pp. LCCN QA3

.L28 no. 456. URL http://link.springer.com/book/10.1007/
BFb0065404;   http://www.springerlink.com/content/978-
3-540-37487-9.

Fourman:1979:SMA

[FH79] M. P. Fourman and J. M. E. Hyland. Sheaf models for analysis. Lecture Notes in Mathematics, 753:280–301, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061823/.

Feder:1977:SOT

[FI77] S. Feder and W. Iberkleid. Secondary operations in K-theory and the generalized vector field problem. Lecture Notes in Mathematics, 597:161–175, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085354/.

Fiedorowicz:1978:QGC

[Fie78] Zbigniew Fiedorowicz. The Quillen–Grothendieck construction and extensions of pairings. Lecture Notes in Mathematics, 657: 163–169, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069233/.

Finucan:1975:CEM

[Fin75] H. M. Finucan. ’Combination of events’ made easy. Lecture Notes in Mathematics, 452:105–115, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069550/.

Finucan:1976:SEA

[Fin76] H. M. Finucan. Some elementary aspects of the Catalan numbers. Lecture Notes in Mathematics, 560:41–45, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097366/.

Fischer:1976:ANM

[Fis76a] Gerd Fischer. Appendix. Normal and maximal complex spaces. Lecture Notes in Mathematics, 538:108–130, 1976. CODEN LN-MAA2. ISBN 3-540-07857-6 (print), 3-540-38121-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080342/.

Fischer:1976:BN

[Fis76b] Gerd Fischer. Basic notions. Lecture Notes in Mathematics, 538: 1–43, 1976. CODEN LNMAA2. ISBN 3-540-07857-6 (print), 3-540-38121-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080339/.

Fischer:1976:CS

[Fis76c] Gerd Fischer. Coherent sheaves. Lecture Notes in Mathematics, 538:44–76, 1976. CODEN LNMAA2. ISBN 3-540-07857-6 (print), 3-540-38121-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080340/.

Fischer:1976:CAG

[Fis76d] Gerd Fischer. Complex Analytic Geometry, volume 538 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07857-6 (print), 3-540-38121-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 201 pp. LCCN QA3

.L28 no. 538. URL http://link.springer.com/book/10.1007/
BFb0080338;   http://www.springerlink.com/content/978-
3-540-38121-1.

Fischer:1976:DF

[Fis76e] Gerd Fischer. Degeneracies and flatness. Lecture Notes in Mathematics, 538:131–161, 1976. CODEN LNMAA2. ISBN 3-540-07857-6 (print), 3-540-38121-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080343/.

Fischer:1976:DC

[Fis76f] Gerd Fischer. Differential calculus. Lecture Notes in Mathematics, 538:77–108, 1976. CODEN LNMAA2. ISBN 3-540-07857-6 (print), 3-540-38121-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080341/.

Fischer:1976:MMF

[Fis76g] Gerd Fischer. Modifications and meromorphic functions. Lecture Notes in Mathematics, 538:162–190, 1976. CODEN LN-MAA2. ISBN 3-540-07857-6 (print), 3-540-38121-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080344/.

Fisher:1976:PSR

[Fis76h] Joe W. Fisher. Problem session on regular rings. Lecture Notes in Mathematics, 545:195–199, 1976. CODEN LNMAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080312/.

Fisher:1976:P

[Fis76i] Joe W. Fisher. Problems. Lecture Notes in Mathematics, 545: 200–212, 1976. CODEN LNMAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080313/.

Fisher:1977:AS

[Fis77] Edward R. Fisher. Abelian structures I. Lecture Notes in Mathematics, 616:270–322, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068202/.

Fisher:1979:SCP

[Fis79] Joe W. Fisher. Semiprime crossed products. Lecture Notes in Mathematics, 734:50–63, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103153/.

Fittler:1975:CMH

[Fit75] Robert Fittler. Closed models and hulls of theories. Lecture Notes in Mathematics, 499:169–189, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079421/.

Fitzgibbon:1976:NEO

[Fit76] W. E. Fitzgibbon. Nonlinear evolution operators and delay equations. Lecture Notes in Mathematics, 564:106–113, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087331/.

Fitouhi:1978:FZE

[Fit78] Ahmed Fitouhi. Fonction zeta d’Epstein pour un opérateur elliptique qui dégénère dans la direction normale. (French) []. Lecture Notes in Mathematics, 660:66–80, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082295/.

Fix:1975:FEF

[Fix75] George Fix. Finite elements and fluid dynamics. Lecture Notes in Mathematics, 461:47–70, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074148/.

Fisher:1975:MOS

[FJ75a] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. L^{1} minimization in one and several variables. Lecture Notes in Mathematics, 479:54–63, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097065/.

Fisher:1975:MPE

[FJ75b] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. L^{∞} minimization problems for elliptic operators. Lecture Notes in Mathematics, 479:47–53, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097064/.

Fisher:1975:ARF

[FJ75c] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Application of the Riesz–Fredholm–Schauder theory to spline functions. Lecture Notes in Mathematics, 479:195–205, 1975. CODEN LN-MAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097076/.

Fisher:1975:AEN

[FJ75d] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Approximation by extremals of nonlinear differential expressions in one variable and quadratic forms in several variables. Lecture Notes in Mathematics, 479:107–133, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097070/.

Fisher:1975:BBO

[FJ75e] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Bang–Bang optimal controls. Lecture Notes in Mathematics, 479:73–84, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097067/.

Fisher:1975:E

[FJ75f] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Epilogue. Lecture Notes in Mathematics, 479:206–207, 1975. CODEN LN-MAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097077.

Fisher:1975:GTK

[FJ75g] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. A general theorem of Kuhn–Tucker type. Lecture Notes in Mathematics, 479:85–89, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097068/.

Fisher:1975:I

[FJ75h] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Introduction. Lecture Notes in Mathematics, 479:1–11, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097060/.

Fisher:1975:LPK

[FJ75i] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. The Landau problem and Kolmogorov’s theorem. Lecture Notes in Mathematics, 479:164–171, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097073/.

Fisher:1975:MII

[FJ75j] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Minimization and interpolation at integer points of the real axis. Lecture Notes in Mathematics, 479:151–163, 1975. CODEN LN-MAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097072/.

Fisher:1975:MLO

[FJ75k] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Minimization with linear operators. Lecture Notes in Mathematics, 479: 23–32, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097062/.

Fisher:1975:MNE

[FJ75l] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Minimum Norm Extremals in Function Spaces: With Applications to Classical and Modern Analysis, volume 479 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 209 pp. LCCN QA3

.L28 no. 479. URL http://link.springer.com/book/10.1007/
BFb0097059;   http://www.springerlink.com/content/978-
3-540-37599-9.

Fisher:1975:NMP

[FJ75m] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Nonlinear minimization problems. Lecture Notes in Mathematics, 479: 12–22, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097061/.

Fisher:1975:NO

[FJ75n] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Nonlinear operators in L^{P}, 1 < p ≤ ∞. Lecture Notes in Mathematics, 479:33–46, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097063/.

Fisher:1975:PIS

[FJ75o] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Perfect interpolating splines on compact intervals. Lecture Notes in Mathematics, 479:172–176, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097074/.

Fisher:1975:PAF

[FJ75p] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. A Pólya algorithm for the Favard solution, N-width characterizations and Whitney type theorems. Lecture Notes in Mathematics, 479:177– 194, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097075/.

Fisher:1975:SUM

[FJ75q] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Sets of uniqueness in L^{∞} minimization problems. Lecture Notes in Mathematics, 479:64–72, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097066/.

Fisher:1975:SUE

[FJ75r] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. Stable and unstable elastica equilibrium and the problem of minimum curvature. Lecture Notes in Mathematics, 479:90–106, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097069/.

Fisher:1975:TAF

[FJ75s] Prof. Stephen D. Fisher and Prof. Joseph W. Jerome. The trigonometric and algebraic Favard problem. Lecture Notes in Mathematics, 479:134–150, 1975. CODEN LNMAA2. ISBN 3-540-07394-9 (print), 3-540-37599-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097071/.

Flensted-Jensen:1975:SFD

[FJ75t] Mogens Flensted-Jensen. Spherical functions and discrete series. Lecture Notes in Mathematics, 466:65–78, 1975. CODEN LN-MAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082198/.

Flensted-Jensen:1979:FSR

[FJ79] Mogens Flensted-Jensen. On a fundamental series of representations related to an affine symmetric space. Lecture Notes in Mathematics, 728:77–96, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063339/.

Flensted-Jensen:1979:PDS

[FJK79] Mogens Flensted-Jensen and Tom H. Koornwinder. Positive definite spherical functions on a non-compact, rank one symmetric space. Lecture Notes in Mathematics, 739:249–282, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062495/.

Forster:1979:GG

[FK79a] Otto Forster and Knut Knorr. Glättungssatz. (German) []. Lecture Notes in Mathematics, 705:50–98, 1979. CODEN LNMAA2. ISBN 3-540-09122-X (print), 3-540-35507-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070097/.

Forster:1979:KVF

[FK79b] Otto Forster and Knut Knorr. Konstruktion verseller Familien kompakter komplexer Räume, volume 705 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09122-X (print), 3-540-35507-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 141 pp. LCCN QA3

.L28 no. 705. URL http://link.springer.com/book/10.1007/
BFb0070095;   http://www.springerlink.com/content/978-
3-540-35507-6.

Forster:1979:KGP

[FK79c] Otto Forster and Knut Knorr. Konvergenzbeweis. (German) [Proof of convergence]. Lecture Notes in Mathematics, 705:99– 134, 1979. CODEN LNMAA2. ISBN 3-540-09122-X (print), 3-540-35507-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070098/.

Forster:1979:TGT

[FK79d] Otto Forster and Knut Knorr. Theorie. (German) [Theory]. Lecture Notes in Mathematics, 705:1–49, 1979. CODEN LN-MAA2. ISBN 3-540-09122-X (print), 3-540-35507-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070096/.

Frei:1979:QCS

[FK79e] Armin Frei and Heinrich Kleisli. A question in categorical shape theory: When is a shape-invariant functor a Kan extension? Lecture Notes in Mathematics, 719:55–62, 1979. CODEN LN-MAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065258/.

Flannagan:1975:ACM

[Fla75] T. B. Flannagan. Axioms of choice in Morse–Kelley class theory. Lecture Notes in Mathematics, 499:190–247, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079422/.

Flachsmeyer:1976:ECK

[Fla76] Jürgen Flachsmeyer. Envelopes in the category of Kakutani-m-spaces. Lecture Notes in Mathematics, 540:243–253, 1976. CO-DEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080862/.

Flachsmeyer:1977:TBA

[Fla77] J. Flachsmeyer. Topologization of Boolean algebras. Lecture Notes in Mathematics, 609:81–97, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068674/.

Fletcher:1976:CGM

[Fle76] R. Fletcher. Conjugate gradient methods for indefinite systems. Lecture Notes in Mathematics, 506:73–89, 1976. CODEN LN-MAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080116/.

Flum:1975:FOL

[Flu75] J. Flum. First-order logic and its extensions. Lecture Notes in Mathematics, 499:248–310, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079423/.

Flytzanis:1979:IML

[Fly79] E. Flytzanis. Invariant measures for linear operators. Lecture Notes in Mathematics, 709:103–110, 1979. CODEN LN-MAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071952/.

Flaschka:1976:SCB

[FM76a] Hermann Flaschka and David W. McLaughlin. Some comments on Bäcklund transformations, canonical transformations, and the inverse scattering method. Lecture Notes in Mathematics, 515: 253–295, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081172/.

Forst:1976:OAL

[FM76b] Wilhelm Forst and Mary Mikhail. Optimale Approximation von linearen Funktionalen auf Klassen periodischer Funktionen. (German) []. Lecture Notes in Mathematics, 556:191–200, 1976. CO-DEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087407/.

Fulton:1978:DAI

[FM78] William Fulton and Robert MacPherson. Defining algebraic intersections. Lecture Notes in Mathematics, 687:1–30, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062926/.

Fack:1979:ATA

[FM79] Th. Fack and O. Maréchal. Application de la K-théorie algébrique aux C∗-algèbres. (French) []. Lecture Notes in Mathematics, 725: 144–169, 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062615/.

Fourman:1979:ASP

[FMS79] Michael Fourman, Christopher Mulvey, and Dana Scott, editors. Applications of Sheaves: Proceedings of the Research Symposium on Applications of Sheaf Theory to Logic, Algebra, and Analysis, Durham, July 9–21, 1977, volume 753 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 753. URL http:

//link.springer.com/book/10.1007/BFb0061811;   http://
www.springerlink.com/content/978-3-540-34849-8.

Foata:1977:CRD

[Foa77a] Dominique Foata, editor. Combinatoire et Représentation du Groupe Symétrique: Actes de la Table Ronde du C.N.R.S. tenue à l’Université Louis-Pasteur de Strasbourg, 26 au 30 avril 1976, volume 579 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 579. URL http:

//link.springer.com/book/10.1007/BFb0090008;   http://
www.springerlink.com/content/978-3-540-37385-8.

Foata:1977:I

[Foa77b] Dominique Foata. Introduction. Lecture Notes in Mathematics, 579:1–10, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090009/.

Foata:1977:PDV

[Foa77c] Dominique Foata. Une propriété du vidage-remplissage des tableaux de Young. (French) []. Lecture Notes in Mathematics, 579:121–135, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090014/.

Foias:1975:SSN

[Foi75] Ciprian Foiaş. On the statistical study of the Navier–Stokes equations. Lecture Notes in Mathematics, 446:184–197, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070603/.

Follmer:1975:PTM

[Föl75] Hans Föllmer. Phase transition and Martin boundary. Lecture Notes in Mathematics, 465:305–317, 1975. CODEN LN-MAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103000/.

Formanov:1976:SER

[For76a] Š. K. Formanov. Some estimates of the rate of convergence in multidimensional limit theorems for homogeneous Markov processes. Lecture Notes in Mathematics, 550:80–98, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077485/.

Forst:1976:FRM

[For76b] Gunnar Forst. Familles resolvantes de mesures. (French) []. Lecture Notes in Mathematics, 518:83–90, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080405/.

Forst:1978:NPF

[For78] Gunnar Forst. Noyaux potentiels sur R_{+}. (French) []. Lecture Notes in Mathematics, 681:103–120, 1978. CODEN LN-MAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065870/.

Forst:1979:MRF

[For79a] Gunnar Forst. Multiples of renewal functions: Remark on a result of D. J. Daley. Lecture Notes in Mathematics, 706:116–119, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063119/.

Forster:1979:NAF

[For79b] W. Forster. On numerical approximation of fixed points in C[0, 1]. Lecture Notes in Mathematics, 730:112–125, 1979. CODEN LN-MAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064314/.

Fouche:1975:CDG

[Fou75] François Fouché. Un complexe dualisant en géométrie analytique. (French) []. Lecture Notes in Mathematics, 482:282–332, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080252/.

Fournier:1979:TDC

[Fou79] Jean-Claude Fournier. Le théorème du coloriage des cartes [exconjecture de Heawood et conjecture des quatre couleurs]. (French) []. Lecture Notes in Mathematics, 710:41–64, 1979. CODEN LN-MAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069972/.

Foxby:1979:IPM

[Fox79] Hans-Bjørn Foxby. Intersection properties of modules. Lecture Notes in Mathematics, 732:90–97, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066639/.

Fiedorowicz:1976:HCG

[FP76] Zbigniew Fiedorowicz and Stewart Priddy. Homology of classical groups over a finite field. Lecture Notes in Mathematics, 551: 269–277, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080007/.

Fiedorowicz:1978:COP

[FP78a] Zbigniew Fiedorowicz and Stewart Priddy. Calculations at odd primes. Lecture Notes in Mathematics, 674:293–312, 1978. CO-DEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062830/.

Fiedorowicz:1978:CP

[FP78b] Zbigniew Fiedorowicz and Stewart Priddy. Calculations at the prime 2. Lecture Notes in Mathematics, 674:250–292, 1978. CO-DEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062829/.

Fiedorowicz:1978:DTO

[FP78c] Zbigniew Fiedorowicz and Stewart Priddy. Detection theorems at odd primes. Lecture Notes in Mathematics, 674:355–385, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062833/.

Fiedorowicz:1978:DTP

[FP78d] Zbigniew Fiedorowicz and Stewart Priddy. Detection theorems at the prime 2. Lecture Notes in Mathematics, 674:343–354, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062832/.

Fiedorowicz:1978:HCF

[FP78e] Zbigniew Fiedorowicz and Stewart Priddy. The homology of certain finite groups. Lecture Notes in Mathematics, 674:313– 342, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062831/.

Fiedorowicz:1978:HCG

[FP78f] Zbigniew Fiedorowicz and Stewart Priddy. Homology of Classical Groups Over Finite Fields and Their Associated Infinite Loop Spaces, volume 674 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 434 pp. LCCN QA3 .L28 no. 674. URL http:

//link.springer.com/book/10.1007/BFb0062824;   http://
www.springerlink.com/content/978-3-540-35735-3.

Fiedorowicz:1978:HOA

[FP78g] Zbigniew Fiedorowicz and Stewart Priddy. Homology operations associated with the classical groups. Lecture Notes in Mathematics, 674:386–406, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062834/.

Fiedorowicz:1978:ILS

[FP78h] Zbigniew Fiedorowicz and Stewart Priddy. Infinite loop spaces associated with ImJ. Lecture Notes in Mathematics, 674:9–48, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062826/.

Fiedorowicz:1978:I

[FP78i] Zbigniew Fiedorowicz and Stewart Priddy. Introduction. Lecture Notes in Mathematics, 674:1–8, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062825/.

Fiedorowicz:1978:TFF

[FP78j] Zbigniew Fiedorowicz and Stewart Priddy. K-theory of finite fields and the ImJ spaces. Lecture Notes in Mathematics, 674: 166–249, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062828/.

Fiedorowicz:1978:PCC

[FP78k] Zbigniew Fiedorowicz and Stewart Priddy. Permutative categories of classical groups over finite fields. Lecture Notes in Mathematics, 674:49–165, 1978. CODEN LNMAA2. ISBN 3-540-08932-2 (print), 3-540-35735-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062827/.

Fougeres:1978:AAC

[FP78l] Alain Fougères and Jean-Claude Peralba. Application au calcul des variations de l’optimisation intégrale convexe. (French) []. Lecture Notes in Mathematics, 665:70–99, 1978. CODEN LN-MAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061799/.

Fichera:1976:SLP

[FR76] Gaetano Fichera and Paolo Emilio Ricci. The single layer potential approach in the theory of boundary value problems for elliptic equations. Lecture Notes in Mathematics, 561:39–50, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087624/.

Fenn:1979:NSM

[FR79a] Roger Fenn and Colin Rourke. Nice spines of 3-manifolds. Lecture Notes in Mathematics, 722:31–36, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063186/.

Ferrante:1979:CCL

[FR79b] Jeanne Ferrante and Charles W. Rackoff. The Computational Complexity of Logical Theories, volume 718 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 243 pp. LCCN QA3

.L28 no. 718. URL http://link.springer.com/book/10.1007/
BFb0062837;   http://www.springerlink.com/content/978-
3-540-35197-9.

Ferrante:1979:DPT

[FR79c] Jeanne Ferrante and Charles W. Rackoff. Direct products of theories. Lecture Notes in Mathematics, 718:128–147, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062842/.

Ferrante:1979:EGD

[FR79d] Jeanne Ferrante and Charles W. Rackoff. Ehrenfeucht games and decision procedures. Lecture Notes in Mathematics, 718:28– 46, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062839/.

Ferrante:1979:IAE

[FR79e] Jeanne Ferrante and Charles W. Rackoff. Integer addition — an example of an Ehrenfeucht game decision procedure. Lecture Notes in Mathematics, 718:47–54, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062840/.

Ferrante:1979:IB

[FR79f] Jeanne Ferrante and Charles W. Rackoff. Introduction and background. Lecture Notes in Mathematics, 718:1–27, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062838/.

Ferrante:1979:LBT

[FR79g] Jeanne Ferrante and Charles W. Rackoff. A lower bound on the theories of pairing functions. Lecture Notes in Mathematics, 718:162–186, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062845/.

Ferrante:1979:LBP

[FR79h] Jeanne Ferrante and Charles W. Rackoff. Lower bound preliminaries. Lecture Notes in Mathematics, 718:148–152, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062843/.

Ferrante:1979:SAL

[FR79i] Jeanne Ferrante and Charles W. Rackoff. Some additional lower bounds. Lecture Notes in Mathematics, 718:187–233, 1979. CO-DEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062846/.

Ferrante:1979:SAU

[FR79j] Jeanne Ferrante and Charles W. Rackoff. Some additional upper bounds. Lecture Notes in Mathematics, 718:55–127, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062841/.

Ferrante:1979:TWS

[FR79k] Jeanne Ferrante and Charles W. Rackoff. A technique for writing short formulas defining complicated properties. Lecture Notes in Mathematics, 718:153–161, 1979. CODEN LNMAA2. ISBN 3-540-09501-2 (print), 3-540-35197-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062844/.

FraeijsdeVeubeke:1977:DRE

[Fra77a] B. M. Fraeijs de Veubeke. Discretization of rotational equilibrium in the finite element method. Lecture Notes in Mathematics, 606:87–112, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064458/.

Frankiewicz:1977:ISP

[Fra77b] R. Frankiewicz. On the inhomogenity of the set of P (m)-points of ω∗. Lecture Notes in Mathematics, 619:169–179, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067650/.

Frame:1978:OGR

[Fra78a] J. Sutherland Frame. Orthogonal groups over GF(2) and related graphs. Lecture Notes in Mathematics, 642:174–185, 1978. CO-DEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070374/.

Franks:1978:SSD

[Fra78b] John Franks. The structure of Smale diffeomorphisms. Lecture Notes in Mathematics, 668:117–126, 1978. CODEN LN-MAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101784/.

Freedman:1975:ACB

[Fre75] Michael H. Freedman. Automorphisms of circle bundles over surfaces. Lecture Notes in Mathematics, 438:212–214, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066116/.

Freedman:1978:RSF

[Fre78] Michael H. Freedman. Remarks on the solution of first degree equations in groups. Lecture Notes in Mathematics, 664:87– 93, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061693/.

Friedman:1975:EBF

[Fri75] Harvey Friedman. Equality between functionals. Lecture Notes in Mathematics, 453:22–37, 1975. CODEN LNMAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064870/.

Friedlander:1976:HSC

[Fri76a] Eric M. Friedlander. Homological stability for classical groups over finite fields. Lecture Notes in Mathematics, 551:290–302, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080010/.

Friedrich:1976:EEA

[Fri76b] Norbert Friedrich. Existenz- und Eindeutigkeitsproblem bei der Abstrahlung ebener Wellen aus einem angeströmten Ringkanal. (German) []. Lecture Notes in Mathematics, 561:203–217, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087637/.

Friedli:1978:VRK

[Fri78a]  A. Friedli.  Verallgemeinerte Runge–Kutta Verfahren zur

Lösung steifer Differentialgleichungssysteme. (German) [Generalized Runge–Kutta method for solving stiff systems of differential equations]. Lecture Notes in Mathematics, 631:35–50, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067462/.

Friedman:1978:CRO

[Fri78b] Harvey Friedman. Categoricity with respect to ordinals. Lecture Notes in Mathematics, 669:17–20, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103099/.

Friedman:1978:CIP

[Fri78c] Harvey Friedman. Classically and intuitionistically provably recursive functions. Lecture Notes in Mathematics, 669:21–27, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103100/.

Friedman:1978:PFA

[Fri78d] Harvey Friedman. A proof of foundation from axioms of cumulation. Lecture Notes in Mathematics, 669:15–16, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103098.

Friedman:1979:TSA

[Fri79] Jerome H. Friedman. A tree-structured approach to nonparametric multiple regression. Lecture Notes in Mathematics, 757: 5–22, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098488/.

Fromm:1975:EFM

[Fro75] J. E. Fromm. Experience with forward marching nonlinear solutions of the Navier–Stokes equations. Lecture Notes in Mathematics, 461:71–73, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074149/.

Frolicher:1976:CGS

[Frö76] Alfred Frölicher. Compactly generated spaces and duality. Lecture Notes in Mathematics, 540:254–264, 1976. CODEN LN-MAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080863/.

Frolik:1977:RDT

[Fro77] Zdeněk Frolík. Recent development of theory of uniform spaces. Lecture Notes in Mathematics, 609:98–108, 1977. CODEN LN-MAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068675/.

Froberg:1979:SCC

[Frö79a] Ralf Fröberg. Some complex constructions with applications to Poincaré series. Lecture Notes in Mathematics, 740:272–284, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071067/.

Frolik:1979:FFP

[Fro79b] Zdeněk Frolík. The finest functor preserving the Baire sets. Lecture Notes in Mathematics, 719:63–73, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065259/.

Faudree:1978:VLP

[FS78a] R. J. Faudree and R. H. Schelp. Various length paths in graphs. Lecture Notes in Mathematics, 642:160–173, 1978. CODEN LN-MAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070373/.

Flexor:1978:TSL

[FS78b] Marguerite Flexor and Lucien Szpiro. Un théorème de structure locale pour les complexes parfaits. (French) []. Lecture Notes in Mathematics, 687:236–244, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062934/.

Fourman:1979:SL

[FS79a] M. P. Fourman and D. S. Scott. Sheaves and logic. Lecture Notes in Mathematics, 753:302–401, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061824/.

Fujita:1979:FEA

[FS79b] Hiroshi Fujita and Takashi Suzuki. On the finite element approximation for evolution equations of parabolic type. Lecture Notes in Mathematics, 704:207–221, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063621/.

Fisher:1975:EEC

[FSW75] E. Fisher, H. Simmons, and W. Wheeler. Elementary equivalence classes of generic structures and existentially complete structures. Lecture Notes in Mathematics, 498:131–169, 1975. CODEN LN-MAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080978/.

Foias:1976:GPS

[FT76] C. Foiaş and R. Temam. A generic property of the set of stationary solutions of Navier Stokes equations. Lecture Notes in Mathematics, 565:24–28, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091445/.

Fenn:1979:ID

[FT79] Roger Fenn and Paul Taylor. Introducting doodles. Lecture Notes in Mathematics, 722:37–43, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063187/.

Frank:1978:CID

[FU78] Reinhard Frank and Christoph W. Ueberhuber. Collocation and iterated defect correction. Lecture Notes in Mathematics, 631: 19–34, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067461/.

Fuchs:1977:SVV

[Fuc77a] László Fuchs. Subfree valued vector spaces. Lecture Notes in Mathematics, 616:158–167, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068194/.

Fuchs:1977:LWV

[Fuc77b] W. H. J. Fuchs. A look at Wiman–Valiron theory. Lecture Notes in Mathematics, 599:46–50, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096824/.

Fucik:1979:NNB

[Fuč79] S. Fučík. Nonlinear noncoercive boundary value problems. Lecture Notes in Mathematics, 703:99–109, 1979. CODEN LN-MAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067262/.

Fuglede:1979:HM

[Fug79] Bent Fuglede. Harmonic morphisms. Lecture Notes in Mathematics, 747:123–131, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063964/.

Fujimagari:1976:ESC

[Fuj76] Tetsuo Fujimagari. Expectation semigroup of a cascade process and a limit theorem. Lecture Notes in Mathematics, 550:99– 118, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077486/.

Fujiie:1979:NCS

[Fuj79] Tatsuo Fuji’i’e. Notes on cluster sets at ideal boundary points. Lecture Notes in Mathematics, 747:132–135, 1979. CODEN LN-MAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063965/.

Fukushima:1976:PTS

[Fuk76] M. Fukushima. Potential theory of symmetric Markov processes and its applications. Lecture Notes in Mathematics, 550:119– 133, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077487/.

Fuller:1979:BR

[Ful79] Kent R. Fuller. Biserial rings. Lecture Notes in Mathematics, 734: 64–90, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103154/.

Furstenberg:1978:FMI

[FW78] Hillel Furstenberg and Benjamin Weiss. The finite multipliers of infinite ergodic transformations. Lecture Notes in Mathematics, 668:127–132, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101785/.

Faith:1979:MTP

[FW79] Carl Faith and Sylvia Wiegand, editors. Module Theory: Papers and Problems from The Special Session Sponsored by The American Mathematical Society at The University of Washington Proceedings, Seattle, August 15–18, 1977, volume 700 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 700. URL http://link.springer.com/book/10.1007/
BFb0063455;   http://www.springerlink.com/content/978-
3-540-35538-0.

G:1975:P

[G.75] J. A. G. Preface. Lecture Notes in Mathematics, 446:3–4, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070593.

Gabriel:1975:FRT

[Gab75a] Peter Gabriel. Finite representation type is open. Lecture Notes in Mathematics, 488:132–155, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081219/.

Gabriel:1975:RIF

[Gab75b] Pierre Gabriel. Représentations indécomposables. (French) []. Lecture Notes in Mathematics, 431:143–169, 1975. CODEN LN-MAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066369/.

Gagen:1977:NGI

[Gag77] T. M. Gagen. A note on groups with the inverse Lagrange property. Lecture Notes in Mathematics, 573:51–52, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087810.

Gajewski:1979:ISS

[Gaj79] H. Gajewski. On the iterative solution of some nonlinear evolution equations. Lecture Notes in Mathematics, 703:110–117, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067263/.

Gallagher:1975:PPN

[Gal75] Richard H. Gallagher. Perturbation procedures in nonlinear finite element structural analysis. Lecture Notes in Mathematics, 461:75–89, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074150/.

Galambos:1976:A

[Gal76a] Prof. János Galambos. The algorithms. Lecture Notes in Mathematics, 502:4–20, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081644/.

Galambos:1976:HDC

[Gal76b] Prof. János Galambos. The Hausdorff dimension of certain sets related to series expansions. Lecture Notes in Mathematics, 502: 110–119, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081650/.

Galambos:1976:I

[Gal76c] Prof. János Galambos. Introduction. Lecture Notes in Mathematics, 502:1–3, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081643/.

Galambos:1976:MTG

[Gal76d] Prof. János Galambos. Metric theory in the general case. Lecture Notes in Mathematics, 502:83–109, 1976. CODEN LN-MAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081649/.

Galambos:1976:MTR

[Gal76e] Prof. János Galambos. Metric theory of representations with stochastically independent digits. Lecture Notes in Mathematics, 502:50–69, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081647/.

Galambos:1976:MT

[Gal76f] Prof. János Galambos. Miscellaneous topics. Lecture Notes in Mathematics, 502:120–134, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081651/.

Galambos:1976:QRI

[Gal76g] Prof. János Galambos. Questions of rationality and irrationality. Lecture Notes in Mathematics, 502:21–31, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081645/.

Galambos:1976:RRN

[Gal76h] Prof. János Galambos. Representations of Real Numbers by Infinite Series, volume 502 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 146 pp. LCCN QA3 .L28 no. 502. URL http:

//link.springer.com/book/10.1007/BFb0081642;   http://
www.springerlink.com/content/978-3-540-38087-0.

Galambos:1976:SCT

[Gal76i] Prof. János Galambos. Some concepts and tools from probability theory. Lecture Notes in Mathematics, 502:32–49, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081646/.

Galambos:1976:SAE

[Gal76j] Prof. János Galambos. A survey on the applications of ergodic theory to series representations. Lecture Notes in Mathematics, 502:70–82, 1976. CODEN LNMAA2. ISBN 3-540-07547-X (print), 3-540-38087-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081648/.

Galperin:1978:ODM

[Gal78] G. A. Gal’perin. One-dimensional monotonic tesselations with memory. Lecture Notes in Mathematics, 653:56–71, 1978. CO-DEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070084/.

Gamelin:1976:HSH

[Gam76] T. W. Gamelin. Hartogs series, Hartogs functions and Jensen measures. Lecture Notes in Mathematics, 512:69–83, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080025/.

Gamkrelidze:1979:ERS

[Gam79] R. V. Gamkrelidze. Exponential representation of solutions of ordinary differential equations. Lecture Notes in Mathematics, 703: 118–129, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067264/.

Gani:1975:PSP

[Gan75] J. Gani. Processus stochastiques de population. (French) []. Lecture Notes in Mathematics, 480:189–293, 1975. CODEN LN-MAA2. ISBN 3-540-07396-5 (print), 3-540-37600-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080192/.

Gapaillard:1976:RAE

[Gap76] Jacques Gapaillard. Relèvements sur une algèbre d’ensembles. (French) []. Lecture Notes in Mathematics, 541:137–153, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081047/.

Garfinkel:1976:MAC

[Gar76a] Gerald S. Garfinkel. A module approach to the Chase–Rosenberg– Zelinsky sequences. Lecture Notes in Mathematics, 549:50–62, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077332/.

Garling:1976:FCL

[Gar76b] D. J. H. Garling. Functional central limit theorems in Banach spaces. Lecture Notes in Mathematics, 526:85–88, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082343/.

Garcia:1977:RSS

[Gar77a] Pedro L. García. Reducibility of the symplectic structure of classical fields with gauge-symmetry. Lecture Notes in Mathematics, 570:365–376, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087792/.

Garnett:1977:TRI

[Gar77b] John B. Garnett. Two remarks on interpolation by bounded analytic functions. Lecture Notes in Mathematics, 604:32–40, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069203/.

Gardner:1978:DGF

[Gar78] Robert B. Gardner. Differential geometry and foliations: The Godbillon–Vey invariant and the Bott–Pasternack vanishingtheorems. Lecture Notes in Mathematics, 652:75–94, 1978. CO-DEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063503/.

Gasper:1975:FDM

[Gas75] George Gasper. Formulas of the Dirichlet–Mehler type. Lecture Notes in Mathematics, 457:207–215, 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067105/.

Gatignol:1977:TAT

[Gat77] Ph. Gatignol. Théorie asymptotique des trains d’ondes lentement modules pour certaines classes d’équations de conservation. (French) [Asymptotic theory of slowly modulating wave trains for certain classes of conservation equations]. Lecture Notes in Mathematics, 594:178–200, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086087/.

Gauthier:1975:PDD

[Gau75] Paul Gauthier. Un plongement du disque unite. (French) []. Lecture Notes in Mathematics, 482:333–336, 1975. CODEN LN-MAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080253/.

Gautschi:1976:ACQ

[Gau76] Walter Gautschi. Advances in Chebyshev quadrature. Lecture Notes in Mathematics, 506:100–121, 1976. CODEN LN-MAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080118/.

Gauld:1979:TEQ

[Gau79] David B. Gauld. A technique for extending quasiconformal embeddings. Lecture Notes in Mathematics, 747:136–143, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063966/.

Gawedzki:1977:GQF

[Gaw77] K. Gawedzki. Geometric quantization and Feynman path integrals for spin. Lecture Notes in Mathematics, 570:67–71, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087782/.

Germain-Bonne:1979:ESP

[GB79] B. Germain-Bonne. Ensembles de suites et de procédés liés pour l’accélération de la convergence. (French) []. Lecture Notes in Mathematics, 765:116–134, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085577/.

Grothendieck:1977:FMM

[GD77a] A. Grothendieck and P. Deligne. Fonctions L modulo ℓ^{n} et modulo p. (French) []. Lecture Notes in Mathematics, 569:110–128, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091520/.

Grothendieck:1977:CCA

[GD77b] A. Grothendieck and P. Deligne. La classe de cohomologie associée à un cycle. (French) []. Lecture Notes in Mathematics, 569:129–153, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091521/.

Gutierrez:1977:CCM

[GdM77] C. Gutiérrez and W. de Melo. The connected components of Morse–Smale vector fields on two manifolds. Lecture Notes in Mathematics, 597:230–251, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085357/.

Gossez:1976:NOC

[GDMW76] Jean Pierre Gossez, Enrique José Lami Dozo, Jean Mawhin, and Lucien Waelbroeck, editors. Nonlinear Operators and the Calculus of Variations: Summer School Held in Bruxelles 8– 19 September 1975, volume 543 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07867-3 (print), 3-540-38075-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 543. URL http:

//link.springer.com/book/10.1007/BFb0079940;   http://
www.springerlink.com/content/978-3-540-38075-7.

Geissinger:1977:HAS

[Gei77] Ladnor Geissinger. Hopf algebras of symmetric functions and class functions. Lecture Notes in Mathematics, 579:168–181, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090017/.

Gelbart:1976:AFM

[Gel76a] Stephen Samuel Gelbart. Automorphic forms on the metaplectic group: Global theory. Lecture Notes in Mathematics, 530: 51–71, 1976. CODEN LNMAA2. ISBN 3-540-07799-5 (print), 3-540-38265-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077833/.

Gelbart:1976:BSR

[Gel76b] Stephen Samuel Gelbart. Background and summary of results. Lecture Notes in Mathematics, 530:1–10, 1976. CODEN LN-MAA2. ISBN 3-540-07799-5 (print), 3-540-38265-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077831/.

Gelbart:1976:GTO

[Gel76c] Stephen Samuel Gelbart. Global theory and odds and ends. Lecture Notes in Mathematics, 530:121–133, 1976. CODEN LN-MAA2. ISBN 3-540-07799-5 (print), 3-540-38265-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077836/.

Gelbart:1976:LTAa

[Gel76d] Stephen Samuel Gelbart. Local theory: the Archimedean places. Lecture Notes in Mathematics, 530:72–95, 1976. CODEN LN-MAA2. ISBN 3-540-07799-5 (print), 3-540-38265-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077834/.

Gelbart:1976:LTAb

[Gel76e] Stephen Samuel Gelbart. Local theory: the p-adic places. Lecture Notes in Mathematics, 530:96–120, 1976. CODEN LN-MAA2. ISBN 3-540-07799-5 (print), 3-540-38265-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077835/.

Gelbart:1976:MGR

[Gel76f] Stephen Samuel Gelbart. Metaplectic groups and representations. Lecture Notes in Mathematics, 530:11–50, 1976. CODEN LN-MAA2. ISBN 3-540-07799-5 (print), 3-540-38265-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077832/.

Gelbart:1976:WRS

[Gel76g] Stephen Samuel Gelbart. Weil’s Representation and the Spectrum of the Metaplectic Group, volume 530 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07799-5 (print), 3-540-38265-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 133 (est.) pp. LCCN QA3

.L28 no. 530. URL http://link.springer.com/book/10.1007/
BFb0077830;   http://www.springerlink.com/content/978-
3-540-38265-2.

Gelbart:1977:AFA

[Gel77] Stephen Gelbart. Automorphic forms and Artin’s conjecture. Lecture Notes in Mathematics, 627:241–276, 1977. CODEN LN-MAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065304/.

Gentleman:1976:RES

[Gen76] W. Morven Gentleman. Row elimination for solving sparse linear systems and least squares problems. Lecture Notes in Mathematics, 506:122–133, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080119/.

Georgii:1976:SBS

[Geo76] Hans-Otto Georgii. On K-systems and Bernoullicity in statistical mechanics. Lecture Notes in Mathematics, 532:216–227, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080182/.

George:1977:SLS

[Geo77] J. Alan George. Solution of linear systems of equations: Direct methods for finite element problems. Lecture Notes in Mathematics, 572:52–101, 1977. CODEN LNMAA2. ISBN 3-540-08130-5 (print), 3-540-37430-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0116615/.

George:1978:AOW

[Geo78a] Alan George. An automatic one-way dissection algorithm for irregular finite element problems. Lecture Notes in Mathematics, 630:76–89, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067698/.

Georgin:1978:ECC

[Geo78b] J.-P. Georgin. Estimation et contrôle des chaines de Markov sur des espaces arbitraires. (French) [Estimation and control of Markov chains on arbitrary spaces]. Lecture Notes in Mathematics, 636:71–113, 1978. CODEN LNMAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063261/.

Georg:1979:ASA

[Geo79a] Kurt Georg. An application of simplicial algorithms to variational inequalities. Lecture Notes in Mathematics, 730:126–135, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064315/.

Georgii:1979:BC

[Geo79b] H. O. Georgii. Basic concepts. Lecture Notes in Mathematics, 760:1–28, 1979. CODEN LNMAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068558/.

Georgii:1979:CMI

[Geo79c] H. O. Georgii. Continuous models with interaction. Lecture Notes in Mathematics, 760:139–169, 1979. CODEN LNMAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068563/.

Georgii:1979:DMI

[Geo79d] H. O. Georgii. Discrete models with interaction. Lecture Notes in Mathematics, 760:110–138, 1979. CODEN LNMAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068562/.

Georgii:1979:ESS

[Geo79e] H. O. Georgii. Equilibrium states for systems of moving particles. Lecture Notes in Mathematics, 760:29–63, 1979. CODEN LNMAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068559/.

Georgii:1979:IM

[Geo79f] H. O. Georgii. Independent models. Lecture Notes in Mathematics, 760:91–109, 1979. CODEN LNMAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068561/.

Georgii:1979:SFR

[Geo79g] H. O. Georgii. Some further results on homogeneous models. Lecture Notes in Mathematics, 760:170–181, 1979. CODEN LN-MAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068564/.

Georgii:1979:SHM

[Geo79h] H. O. Georgii. Spatially homogeneous models. Lecture Notes in Mathematics, 760:64–90, 1979. CODEN LNMAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068560/.

Georgii:1979:CGM

[Geo79i] Hans-Otto Georgii. Canonical Gibbs Measures: Some Extensions of de Finetti’s Representation Theorem for Interacting Particle Systems, volume 760 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09712-0 (print), 3-540-38484-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 190 pp. LCCN QA3 .L28 no. 760. URL http:

//link.springer.com/book/10.1007/BFb0068557;   http://
www.springerlink.com/content/978-3-540-38484-7.

Gerardin:1975:GRG

[Gér75a] Paul Gérardin. Groupes reductifs et groupes résolubles. (French) []. Lecture Notes in Mathematics, 466:79–85, 1975. CODEN LN-MAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082199/.

Gerardin:1975:CSD

[Gér75b] Prof. Paul Gérardin. Construction de Séries Discrètes padiques: Sur les Séries Discrètes non Ramifiées des Groupes Réductifs Déployés p-adiques. (French) [Construction of p-Adic: Discrete Series: On Nondiscrete Series of Branched Reductive Groups Made p-adic], volume 462 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07172-5 (print), 3-540-37506-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 180 pp. LCCN QA3

.L28 no. 462. URL http://link.springer.com/book/10.1007/
BFb0082161;   http://www.springerlink.com/content/978-
3-540-37506-7.

Gerardin:1975:GCC

[Gér75c] Prof. Paul Gérardin. Groupes de Chevalley sur les corps ℘-adiques. (French) []. Lecture Notes in Mathematics, 462:55–95, 1975. CO-DEN LNMAA2. ISBN 3-540-07172-5 (print), 3-540-37506-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082164/.

Gerardin:1975:GHC

[Gér75d] Prof. Paul Gérardin. Groupes d’Heisenberg sur les corps finis. (French) []. Lecture Notes in Mathematics, 462:5–54, 1975. CO-DEN LNMAA2. ISBN 3-540-07172-5 (print), 3-540-37506-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0082163/.

Gerardin:1975:IFI

[Gér75e] Prof. Paul Gérardin. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 462:1–4, 1975. CODEN LN-MAA2. ISBN 3-540-07172-5 (print), 3-540-37506-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082162/.

Gerardin:1975:QRS

[Gér75f] Prof. Paul Gérardin. Quelques représentations supercuspidales. (French) []. Lecture Notes in Mathematics, 462:162–178, 1975. CODEN LNMAA2. ISBN 3-540-07172-5 (print), 3-540-37506-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082167/.

Gerardin:1975:RDG

[Gér75g] Prof. Paul Gérardin. Représentations du groupe des points entiers. (French) []. Lecture Notes in Mathematics, 462:122–161, 1975. CODEN LNMAA2. ISBN 3-540-07172-5 (print), 3-540-37506-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082166/.

Gerardin:1975:TMN

[Gér75h] Prof. Paul Gérardin. Tores maximaux non ramifiés. (French) [Maximal unbranched tori]. Lecture Notes in Mathematics, 462: 96–121, 1975. CODEN LNMAA2. ISBN 3-540-07172-5 (print), 3-540-37506-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082165/.

Gerl:1975:QGE

[Ger75i] P. Gerl. Quelques généralisations de l’équirépartition. (French) []. Lecture Notes in Mathematics, 475:100–103, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074259/.

Gerard:1977:HSC

[Gér77] R. Gérard. On holomorphic solutions of certain kind of Pfaffian systems with singularities. Lecture Notes in Mathematics, 597: 176–190, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085355/.

Gerard:1978:SFC

[Gér78] Raymond Gérard. Solutions formelles et convergentes de certains systèmes de Pfaff analytiques avec singularités. (French) []. Lecture Notes in Mathematics, 670:352–364, 1978. CODEN LN-MAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064409/.

Geradin:1979:SAH

[Ger79a] M. Geradin. Special applications of Hamiltons’s principle to structural dynamics. Lecture Notes in Mathematics, 704:222–238, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063622/.

Gerardin:1979:CDC

[Gér79b] Paul Gérardin. Changement du corps de base pour les représentations de gl(2) [d’après R. P. Langlands, H. Saito et T. Shintani]. (French) []. Lecture Notes in Mathematics, 710:

65–88, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069973/.

Gerl:1979:GG

[Ger79c] Peter Gerl. Ein Gleichverteilungssatz auf F_{2}. (German) []. Lecture Notes in Mathematics, 706:126–130, 1979. CODEN LN-MAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063121/.

Gerl:1979:KFG

[Ger79d] Peter Gerl. Ein Konvergenzsatz für Faltungspotenzen. (German) []. Lecture Notes in Mathematics, 706:120–125, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063120/.

Getoor:1975:CPS

[Get75a] Prof. Ronald K. Getoor. Characterization of previsible stopping times. Lecture Notes in Mathematics, 440:33–42, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070068/.

Getoor:1975:C

[Get75b] Prof. Ronald K. Getoor. Comparison of (X_{t−}) and (X_{t−}^{∗}). Lecture Notes in Mathematics, 440:88–95, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070074/.

Getoor:1975:CP

[Get75c] Prof. Ronald K. Getoor. Comparison of processes. Lecture Notes in Mathematics, 440:66–77, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070072/.

Getoor:1975:ISS

[Get75d] Prof. Ronald K. Getoor. Increasing sequences of supermartingales. Lecture Notes in Mathematics, 440:17–18, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070065.

Getoor:1975:MPR

[Get75e] Prof. Ronald K. Getoor. Markov Processes: Ray Processes and Right Processes, volume 440 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 118 pp. LCCN QA3

.L28 no. 440. URL http://link.springer.com/book/10.1007/
BFb0070061;   http://www.springerlink.com/content/978-
3-540-37422-0.

Getoor:1975:Pa

[Get75f] Prof. Ronald K. Getoor. Preliminaries. Lecture Notes in Mathematics, 440:1–2, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070062.

Getoor:1975:Pb

[Get75g] Prof. Ronald K. Getoor. Processes. Lecture Notes in Mathematics, 440:19–27, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070066/.

Getoor:1975:PC

[Get75h] Prof. Ronald K. Getoor. Processes continued. Lecture Notes in Mathematics, 440:28–32, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070067/.

Getoor:1975:RKC

[Get75i] Prof. Ronald K. Getoor. The Ray Knight compactification. Lecture Notes in Mathematics, 440:57–65, 1975. CODEN LN-MAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070071/.

Getoor:1975:RRS

[Get75j] Prof. Ronald K. Getoor. Ray resolvents and semigroups. Lecture Notes in Mathematics, 440:8–16, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070064/.

Getoor:1975:RS

[Get75k] Prof. Ronald K. Getoor. The ray space. Lecture Notes in Mathematics, 440:102–112, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070076/.

Getoor:1975:R

[Get75l] Prof. Ronald K. Getoor. Resolvents. Lecture Notes in Mathematics, 440:3–7, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070063/.

Getoor:1975:RP

[Get75m] Prof. Ronald K. Getoor. Right processes. Lecture Notes in Mathematics, 440:50–56, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070070/.

Getoor:1975:RPC

[Get75n] Prof. Ronald K. Getoor. Right processes continued: Shih’s theorem. Lecture Notes in Mathematics, 440:78–87, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070073/.

Getoor:1975:STM

[Get75o] Prof. Ronald K. Getoor. Some topology and measure theory. Lecture Notes in Mathematics, 440:43–49, 1975. CODEN LN-MAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070069/.

Getoor:1975:S

[Get75p] Prof. Ronald K. Getoor. U-spaces. Lecture Notes in Mathematics, 440:96–101, 1975. CODEN LNMAA2. ISBN 3-540-07140-7 (print), 3-540-37422-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070075/.

Greenwood:1975:SDQ

[GH75] P. Greenwood and R. Hersh. Stochastic differentials and quasistandard random variables. Lecture Notes in Mathematics, 451: 35–62, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068578/.

Gilbert:1976:CFT

[GH76] R. P. Gilbert and G. C. Hsiao. Constructive function theoretic methods for higher order pseudoparabolic equations. Lecture Notes in Mathematics, 561:51–67, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087625/.

Gordon:1977:ATB

[GH77] Noam Gordon and Frank C. Hoppensteadt. An analysis of transient behavior in the onset of convection. Lecture Notes in Mathematics, 594:231–243, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086089/.

Glover:1978:ECR

[GH78a] Henry Glover and Bill Homer. Endomorphisms of the cohomology ring of finite Grassmann manifolds. Lecture Notes in Mathematics, 657:170–193, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069234/.

Glover:1978:IME

[GH78b] Henry Glover and Bill Homer. Immersing manifolds and 2equivalence. Lecture Notes in Mathematics, 657:194–197, 1978.

CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069235/.

Gauthier:1979:UHA

[GH79] P. M. Gauthier and W. Hengartner. Uniform harmonic approximation on unbounded sets. Lecture Notes in Mathematics, 747: 144–149, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063967/.

Ghisa:1979:AMQ

[Ghi79a] D. Ghişa. Associated measures and the quasiconformality. Lecture Notes in Mathematics, 747:150–152, 1979. CODEN LN-MAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063968/.

Ghisa:1979:BSM

[Ghi79b] Dorin Ghişa. The behaviour of some metrics on Riemann surfaces in respect with quasiconformal mappings. Lecture Notes in Mathematics, 743:108–113, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079485/.

Godsil:1977:SG

[GHM77] C. Godsil, D. A. Holton, and B. McKay. The spectrum of a graph. Lecture Notes in Mathematics, 622:91–117, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069184/.

Glover:1978:IMP

[GHM78] Henry Glover, Bill Homer, and Guido Mislin. Immersions in manifolds of positive weights. Lecture Notes in Mathematics, 673: 88–92, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064689/.

Giambalvo:1978:STF

[Gia78] V. Giambalvo. Some tables for formal groups and BP. Lecture Notes in Mathematics, 658:169–176, 1978. CODEN LN-MAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068715/.

Gibson:1976:CCS

[Gib76] Christopher Gibson. Construction of canonical stratifications. Lecture Notes in Mathematics, 552:8–34, 1976. CODEN LN-MAA2. ISBN 3-540-07997-1 (print), 3-540-37957-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0095246/.

Giesy:1976:SLL

[Gie76] Daniel P. Giesy. Strong laws of large numbers for independent sequences of Banach space-valued random variables. Lecture Notes in Mathematics, 526:89–99, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082344/.

Gilmer:1975:DTC

[Gil75a] Robert Gilmer. Dimension theory of commutative polynomial rings. Lecture Notes in Mathematics, 450:140–154, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062853/.

Gilmer:1975:DTP

[Gil75b] Robert Gilmer. Dimension theory of power series rings over a commutative ring. Lecture Notes in Mathematics, 450:155–162, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062854/.

Gillam:1976:BFR

[Gil76] D. W. H. Gillam. Bounds of finite relations. Lecture Notes in Mathematics, 560:55–60, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097368/.

Gilewicz:1978:APT

[Gil78a] Jacek Gilewicz. Approximants de Padé — théorie algébrique. (French) []. Lecture Notes in Mathematics, 667:161–237, 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061332/.

Gilewicz:1978:APA

[Gil78b] Jacek Gilewicz. Approximants de Padé et ϵ-algorithme théorie de la convergence. (French) []. Lecture Notes in Mathematics, 667: 239–337, 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061333/.

Gilewicz:1978:APF

[Gil78c] Jacek Gilewicz. Approximants de Padé. (French) [Padé Approximants], volume 667 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 511 pp. LCCN QA3 .L28 no. 667. URL http:

//link.springer.com/book/10.1007/BFb0061327;   http://
www.springerlink.com/content/978-3-540-35752-0.

Gilewicz:1978:CAP

[Gil78d] Jacek Gilewicz. Calcul des approximants de Padé, des tables de Padé et des tables C. (French) []. Lecture Notes in Mathematics, 667:339–381, 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061334/.

Gilewicz:1978:FSF

[Gil78e] Jacek Gilewicz. Fonctions de Stieltjes et fonctions de classe. Relations avec les suites. (French) []. Lecture Notes in Mathematics, 667:97–118, 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061330/.

Gilewicz:1978:FCF

[Gil78f] Jacek Gilewicz. Fractions continues. (French) [Continued fractions]. Lecture Notes in Mathematics, 667:119–160, 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061331/.

Gilewicz:1978:PDM

[Gil78g] Jacek Gilewicz. Problème du meilleur approximant de Padé dans un ensemble fini d’Approximants. (French) []. Lecture Notes in Mathematics, 667:383–420, 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061335/.

Gilewicz:1978:QAA

[Gil78h] Jacek Gilewicz. Quelques applications des approximants de Padé en analyse Numérique. (French) []. Lecture Notes in Mathematics, 667:421–472, 1978. CODEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061336/.

Gilewicz:1978:RSN

[Gil78i] Jacek Gilewicz. Rappels sur les suites numériques. (French) []. Lecture Notes in Mathematics, 667:1–31, 1978. CODEN LN-MAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061328/.

Gilewicz:1978:SFT

[Gil78j] Jacek Gilewicz. Suites et fonctions totalement monotones. (French) []. Lecture Notes in Mathematics, 667:33–96, 1978. CO-DEN LNMAA2. ISBN 3-540-08924-1 (print), 3-540-35752-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0061329/.

GineM:1976:SRC

[Gin76] Evarist Giné M. Some remarks on the central limit theorem in C(S). Lecture Notes in Mathematics, 526:101–106, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082345/.

Gine:1979:DAB

[Gin79] Evarist Giné. Domains of attraction in Banach spaces. Lecture Notes in Mathematics, 721:22–40, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070849/.

Gerlein:1976:HSM

[GK76a] O. V. Gerleĭn and A. M. Kagan. Hilbert space methods in classical problems of mathematical statistics. Lecture Notes in Mathematics, 550:134–159, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077488/.

Grillenberger:1976:MSP

[GK76b] Chr. Grillenberger and U. Krengel. On matrix summation and the pointwise ergodic theorem. Lecture Notes in Mathematics, 532: 113–124, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080174/.

Galambos:1978:CBT

[GK78a] Janos Galambos and Samuel Kotz. Characterizations based on truncated distributions. Lecture Notes in Mathematics, 675:21– 35, 1978. CODEN LNMAA2. ISBN 3-540-08933-0 (print), 3-540-35733-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069532/.

Galambos:1978:CPO

[GK78b] Janos Galambos and Samuel Kotz. Characterizations by properties of order statistics. Lecture Notes in Mathematics, 675:36– 63, 1978. CODEN LNMAA2. ISBN 3-540-08933-0 (print), 3-540-35733-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069533/.

Galambos:1978:CME

[GK78c] Janos Galambos and Samuel Kotz. Characterizations of multivariate exponential distributions. Lecture Notes in Mathematics, 675:103–132, 1978. CODEN LNMAA2. ISBN 3-540-08933-0 (print), 3-540-35733-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069535/.

Galambos:1978:CPD

[GK78d] János Galambos and Samuel Kotz. Characterizations of Probability Distributions: A Unified Approach with an Emphasis on Exponential and Related Models, volume 675 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 0-387-08933-0, 3-540-08933-0 (print), 3-540-35733-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 169 pp. LCCN QA3 .L28 no. 675; QA3 .L28 no. 675 QA273.6. URL http:

//link.springer.com/book/10.1007/BFb0069530;   http://
www.springerlink.com/content/978-3-540-35733-9.

Galambos:1978:CPP

[GK78e] Janos Galambos and Samuel Kotz. Characterizations of the Poisson process. Lecture Notes in Mathematics, 675:64–102, 1978. CODEN LNMAA2. ISBN 3-540-08933-0 (print), 3-540-35733-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069534/.

Galambos:1978:MR

[GK78f] Janos Galambos and Samuel Kotz. Miscellaneous results. Lecture Notes in Mathematics, 675:133–142, 1978. CODEN LN-MAA2. ISBN 3-540-08933-0 (print), 3-540-35733-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069536/.

Galambos:1978:PBR

[GK78g] Janos Galambos and Samuel Kotz. Preliminaries and basic results. Lecture Notes in Mathematics, 675:1–20, 1978. CODEN LNMAA2. ISBN 3-540-08933-0 (print), 3-540-35733-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069531/.

Gromoll:1975:RGI

[GKM75] Detlef Gromoll, Wilhelm Klingenberg, and Wolfgang Meyer. Riemannsche Geometrie im Großen. (German) [Riemannian geometry in the large], volume 55 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 287 pp. LCCN QA3 .L28 no. 55. URL http:

//link.springer.com/book/10.1007/BFb0079185;   http://
www.springerlink.com/content/978-3-540-35852-7.

Guivarch:1977:CCC

[GKR77a] Yves Guivarc’h, Michael Keane, and Bernard Roynette. Caractérisation de certaines classes de groupes de Lie récurrents. (French) []. Lecture Notes in Mathematics, 624:176–191, 1977. CODEN LNMAA2. ISBN 3-540-08526-2 (print), 3-540-36987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061344/.

Guivarch:1977:IMA

[GKR77b] Yves Guivarc’h, Michael Keane, and Bernard Roynette. Introduction aux marches aléatoires. (French) []. Lecture Notes in Mathematics, 624:1–56, 1977. CODEN LNMAA2. ISBN 3-540-08526-2 (print), 3-540-36987-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061340/.

Guivarch:1977:MAGd

[GKR77c] Yves Guivarc’h, Michael Keane, and Bernard Roynette. Marches aléatoires généralisées. (French) []. Lecture Notes in Mathematics, 624:192–284, 1977. CODEN LNMAA2. ISBN 3-540-08526-2 (print), 3-540-36987-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061345/.

Guivarch:1977:MAGa

[GKR77d] Yves Guivarc’h, Michael Keane, and Bernard Roynette. Marches aléatoires sur le groupe de déplacements de R^{d}. (French) []. Lecture Notes in Mathematics, 624:57–100, 1977. CODEN LN-MAA2. ISBN 3-540-08526-2 (print), 3-540-36987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061341/.

Guivarch:1977:MAGe

[GKR77e] Yves Guivarc’h, Michael Keane, and Bernard Roynette. Marches Aléatoires sur les Groupes de Lie, volume 624 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2.

ISBN 3-540-08526-2 (print), 3-540-36987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 292 pp. LCCN QA3

.L28 no. 624. URL http://link.springer.com/book/10.1007/
BFb0061339;   http://www.springerlink.com/content/978-
3-540-36987-5.

Guivarch:1977:MAGb

[GKR77f] Yves Guivarc’h, Michael Keane, and Bernard Roynette. Marches aléatoires sur les groupes nilpotents. (French) []. Lecture Notes in Mathematics, 624:101–159, 1977. CODEN LNMAA2. ISBN 3-540-08526-2 (print), 3-540-36987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061342/.

Guivarch:1977:MAGc

[GKR77g] Yves Guivarc’h, Michael Keane, and Bernard Roynette. Marches aléatoires sur les groupes résolubles. (French) []. Lecture Notes in Mathematics, 624:160–175, 1977. CODEN LNMAA2. ISBN 3-540-08526-2 (print), 3-540-36987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061343/.

Gutierrez:1975:SGA

[GL75] M. A. Gutiérrez and R. C. Lacher. Semifree group actions and homology spheres. Lecture Notes in Mathematics, 438:240–244, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066119/.

Graham:1978:DMP

[GL78] R. L. Graham and L. Lovász. Distance matrix polynomials of trees. Lecture Notes in Mathematics, 642:186–190, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070375/.

Glaser:1975:MCP

[Gla75a] Leslie C. Glaser. On mildly cellular pseudo cells. Lecture Notes in Mathematics, 438:215–224, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066117/.

Glashoff:1975:OCO

[Gla75b] Klaus Glashoff. Optimal control of one-dimensional linear parabolic differential equations. Lecture Notes in Mathematics, 477:102–120, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079170/.

Glasner:1976:HF

[Gla76a] Shmuel Glasner. μ-harmonic functions. Lecture Notes in Mathematics, 517:55–70, 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080144/.

Glasner:1976:EGA

[Gla76b] Shmuel Glasner. Exponential growth, amenability and free subgroups on two generators. Lecture Notes in Mathematics, 517: 87–97, 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080146/.

Glasner:1976:FBL

[Gla76c] Shmuel Glasner. The Furstenberg boundary of a Lie group. Lecture Notes in Mathematics, 517:42–54, 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080143/.

Glasner:1976:GBCa

[Gla76d] Shmuel Glasner. The generalized Bohr compactification, I. Lecture Notes in Mathematics, 517:98–111, 1976. CODEN LN-MAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080147/.

Glasner:1976:GBCb

[Gla76e] Shmuel Glasner. The generalized Bohr compactification, II. Lecture Notes in Mathematics, 517:112–126, 1976. CODEN LN-MAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080148/.

Glasner:1976:HFS

[Gla76f] Shmuel Glasner. Harmonic functions on a symmetric space. Lecture Notes in Mathematics, 517:71–86, 1976. CODEN LN-MAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080145/.

Glasner:1976:PFb

[Gla76g] Shmuel Glasner. PI-flows. Lecture Notes in Mathematics, 517: 127–150, 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080149/.

Glasner:1976:PFa

[Gla76h] Shmuel Glasner. Proximal flows. Lecture Notes in Mathematics, 517:17–29, 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080141/.

Glasner:1976:PFc

[Gla76i] Shmuel Glasner. Proximal Flows, volume 517 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 153 pp. LCCN QA3

.L28 no. 517. URL http://link.springer.com/book/10.1007/
BFb0080139;   http://www.springerlink.com/content/978-
3-540-38224-9.

Glasner:1976:SPF

[Gla76j] Shmuel Glasner. Strongly proximal flows. Lecture Notes in Mathematics, 517:30–41, 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080142/.

Glasner:1976:UA

[Gla76k] Shmuel Glasner. The universal approach. Lecture Notes in Mathematics, 517:1–16, 1976. CODEN LNMAA2. ISBN 3-540-07689-1 (print), 3-540-38224-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080140/.

Glaz:1977:TAM

[Gla77] Harland M. Glaz. Two attempts at modeling two-dimensional turbulence. Lecture Notes in Mathematics, 615:135–155, 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068366/.

Glimm:1978:CDQ

[Gli78] J. Glimm. The construction and decomposition of quantum fields using operator theory, probability and fiber bundles. Lecture Notes in Mathematics, 650:55–65, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067389/.

Glowinski:1979:CMA

[GLL79] R. Glowinski, J.-L. Lions, and Iria Laboria, editors. Computing Methods in Applied Sciences and Engineering, 1977, I: Third International Symposium December 5–9, 1977, volume 704 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 704. URL http://link.springer.com/book/10.1007/
BFb0063609;   http://www.springerlink.com/content/978-
3-540-35411-6.

Gloede:1975:STI

[Glo75] K. Gloede. Set theory in infinitary languages. Lecture Notes in Mathematics, 499:311–362, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079424/.

Gloede:1978:HSD

[Glo78] Klaus Gloede. Hierarchies of sets definably by means of infinitary languages. Lecture Notes in Mathematics, 669:29–54, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103101/.

Gluck:1975:AAS

[Glu75] Herman Gluck. Almost all simply connected closed surfaces are rigid. Lecture Notes in Mathematics, 438:225–239, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066118/.

Gill:1976:NLS

[GM76a] Philip E. Gill and Walter Murray. Nonlinear least squares and nonlinearly constrained optimization. Lecture Notes in Mathematics, 506:134–147, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080120/.

Godsil:1976:PGT

[GM76b] C. Godsil and B. Mckay. Products of graphs and their spectra. Lecture Notes in Mathematics, 560:61–72, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097369/.

Godsil:1976:SCR

[GM76c] C. Godsil and B. Mckay. Some computational results on the spectra of graphs. Lecture Notes in Mathematics, 560:73–92, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097370/.

Gaines:1977:APH

[GM77a] Robert E. Gaines and Jean L. Mawhin. Alternative problems: an historical perspective. Lecture Notes in Mathematics, 568: 4–9, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089539/.

Gaines:1977:ASP

[GM77b] Robert E. Gaines and Jean L. Mawhin. Approximation of solutions — the projection method. Lecture Notes in Mathematics, 568:104–133, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089543/.

Gaines:1977:BVP

[GM77c] Robert E. Gaines and Jean L. Mawhin. Boundary-value problems for some semilinear elliptic partial differential equations. Lecture Notes in Mathematics, 568:151–165, 1977. CODEN LN-MAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089545/.

Gaines:1977:CDN

[GM77d] Robert E. Gaines and Jean L. Mawhin. Coincidence Degree, and Nonlinear Differential Equations, volume 568 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 262 pp. LCCN QA3

.L28 no. 568. URL http://link.springer.com/book/10.1007/
BFb0089537;   http://www.springerlink.com/content/978-
3-540-37501-2.

Gaines:1977:CDS

[GM77e] Robert E. Gaines and Jean L. Mawhin. Coincidence degree for k-set contractive perturbations of linear Fredholm mappings. Lecture Notes in Mathematics, 568:209–228, 1977. CODEN LN-MAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089548/.

Gaines:1977:CDP

[GM77f] Robert E. Gaines and Jean L. Mawhin. Coincidence degree for perturbations of Fredholm mappings. Lecture Notes in Mathematics, 568:10–25, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089540/.

Gaines:1977:CIM

[GM77g] Robert E. Gaines and Jean L. Mawhin. Coincidence index, multiplicity and bifurcation theory. Lecture Notes in Mathematics, 568: 189–208, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089547/.

Gaines:1977:GCT

[GM77h] Robert E. Gaines and Jean L. Mawhin. A generalized continuation theorem and existence theorems for Lx = Nx. Lecture Notes in Mathematics, 568:26–35, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089541/.

Gaines:1977:I

[GM77i] Robert E. Gaines and Jean L. Mawhin. Introduction. Lecture Notes in Mathematics, 568:1–3, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089538/.

Gaines:1977:NPF

[GM77j] Robert E. Gaines and Jean L. Mawhin. Nonlinear perturbations of Fredholm mappings of nonzero index. Lecture Notes in Mathematics, 568:229–241, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089549/.

Gaines:1977:PSO

[GM77k] Robert E. Gaines and Jean L. Mawhin. Periodic solutions of ordinary differential equations with quasibounded nonlinearities and of functional differential equations. Lecture Notes in Mathematics, 568:166–188, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089546/.

Gaines:1977:QPF

[GM77l] Robert E. Gaines and Jean L. Mawhin. Quasibounded perturbations of Fredholm mappings. Lecture Notes in Mathematics, 568: 134–150, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089544/.

Gaines:1977:TPB

[GM77m] Robert E. Gaines and Jean L. Mawhin. Two-point boundary value problems: Nonlinearities without special structure. Lecture Notes in Mathematics, 568:36–103, 1977. CODEN LNMAA2. ISBN 3-540-08067-8 (print), 3-540-37501-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089542/.

Galligani:1977:MAF

[GM77n] Ilio Galligani and Enrico Magenes, editors. Mathematical Aspects of Finite Element Methods: Proceedings of the Conference Held in Rome, December 10–12, 1975, volume 606 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LN-MAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 606. URL http://link.springer.com/book/10.1007/
BFb0064451;   http://www.springerlink.com/content/978-
3-540-37158-8.

Genet:1977:PSP

[GM77o] J. Genet and M. Madaune. Perturbations singulières pour une classe de problémes hyperboliques non linéaires. (French) []. Lecture Notes in Mathematics, 594:201–230, 1977. CODEN LN-MAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086088/.

George:1977:AMD

[GM77p] Alan George and David R. McIntyre. On the application of the minimum degree algorithm to finite element systems. Lecture Notes in Mathematics, 606:122–149, 1977. CODEN LN-MAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064460/.

Gasser:1979:KER

[GM79a] Theo Gasser and Hans-Georg Müller. Kernel estimation of regression functions. Lecture Notes in Mathematics, 757:23–68, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098489/.

Gilewicz:1979:VT

[GM79b] Jacek Gilewicz and Alphonse Magnus. Valleys in c-table. Lecture Notes in Mathematics, 765:135–149, 1979. CODEN LN-MAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085578/.

Graves-Morris:1979:NCP

[GM79c] P. R. Graves-Morris. The numerical calculation of Padé approximants. Lecture Notes in Mathematics, 765:231–245, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085583/.

Grmela:1979:GAP

[GM79d] Miroslav Grmela and Jerrold Eldon Marsden, editors. Global Analysis: Proceedings of the Biennial Seminar of the Canadian Mathematical Congress, Calgary, Alberta, June 12–27, 1978, volume 755 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 755. URL http://link.springer.com/book/10.1007/
BFb0069800;   http://www.springerlink.com/content/978-
3-540-38462-5.

Gromoll:1975:AGA

[GMK75a] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Anhang. (German) [Appendix]. Lecture Notes in Mathematics, 55: 272–282, 1975. CODEN LNMAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079193/.

Gromoll:1975:BZK

[GMK75b] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Beziehungen zwischen Krümmung und topologischer Gestalt. (German) []. Lecture Notes in Mathematics, 55:197–271, 1975. CODEN LNMAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079192/.

Gromoll:1975:DMA

[GMK75c] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Differenzierbare Mannigfaltigkeiten und Abbildungen. (German) []. Lecture Notes in Mathematics, 55:1–34, 1975. CODEN LN-MAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079186/.

Gromoll:1975:EGG

[GMK75d] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Extremaleigenschaften von Geodätischen. (German) []. Lecture Notes in Mathematics, 55:121–155, 1975. CODEN LNMAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079189/.

Gromoll:1975:LZG

[GMK75e] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Lineare Zusammenhänge. (German) []. Lecture Notes in Mathematics, 55:35–68, 1975. CODEN LNMAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079187/.

Gromoll:1975:RMR

[GMK75f] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Riemannigfaltigkeiten als metrische Räume. (German) []. Lecture Notes in Mathematics, 55:156–173, 1975. CODEN LN-MAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079190/.

Gromoll:1975:RMG

[GMK75g] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Riemannsche Mannigfaltigkeiten. (German) []. Lecture Notes in Mathematics, 55:69–120, 1975. CODEN LNMAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079188/.

Gromoll:1975:VG

[GMK75h] Detlef Gromoll, Wolfgang Meyer, and Wilhelm Klingenberg. Vergleichssätze. (German) []. Lecture Notes in Mathematics, 55:174– 196, 1975. CODEN LNMAA2. ISBN 3-540-07133-4 (print), 3-540-35852-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079191/.

Garcia:1978:MSM

[GMP78] M. Garcia, P. Maillard, and Y. Peltraut. Une martingale de saut multiplicatif donné. (French) []. Lecture Notes in Mathematics, 649:51–52, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064592.

Gine:1979:SIR

[GMZ79] Evarist Giné, V. Mandrekar, and Joel Zinn. On sums of independent random variables with values in L_{p}(2 ≤ p < ∞). Lecture Notes in Mathematics, 709:111–124, 1979. CODEN LN-MAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071953/.

Germain:1976:AMF

[GN76] Paul Germain and Bernard Nayroles, editors. Applications of Methods of Functional Analysis to Problems in Mechanics: Joint Symposium IUTAM/IMU Held in Marseille, September 1–6, 1975, volume 503 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 503. URL http:

//link.springer.com/book/10.1007/BFb0088742;   http://
www.springerlink.com/content/978-3-540-38165-5.

Godsil:1978:GGP

[God78] C. D. Godsil. Graphs, groups and polytopes. Lecture Notes in Mathematics, 686:157–164, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062528/.

Goldstein:1975:PDE

[Gol75] Prof. Jerome A. Goldstein, editor. Partial Differential Equations and Related Topics: Ford Foundation Sponsored Program at Tulane University, January to May, 1974, volume 446 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 446. URL http://link.springer.com/book/10.1007/
BFb0070592;   http://www.springerlink.com/content/978-
3-540-37440-4.

Goldman:1978:REP

[Gol78a] André Goldman. Relations entre les propriétés de mesurabilité universelle pour un espace topologique T et la propriété de Radon– Nikodým pour le cône positif des mesures de Radon (resp, de Baire) sur T. (French) []. Lecture Notes in Mathematics, 645: 91–96, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069665/.

Goldsmith:1978:LIC

[Gol78b] Deborah L. Goldsmith. A linking invariant of classical link concordance. Lecture Notes in Mathematics, 685:135–170, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062970/.

Goldfeld:1979:CEC

[Gol79] Dorian Goldfeld. Conjectures on elliptic curves over quadratic fields. Lecture Notes in Mathematics, 751:108–118, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062705/.

Goncalves:1978:HAS

[Gon78] Daciberg Lima Gonçalves. mod 2 homotopy-associative H-spaces. Lecture Notes in Mathematics, 657:198–216, 1978. CODEN LN-MAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069236/.

Goodearl:1976:RRR

[Goo76a] K. R. Goodearl. Regular rings and rank functions. Lecture Notes in Mathematics, 545:83–104, 1976. CODEN LNMAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080306/.

Goodman:1976:A

[Goo76b] Roe William Goodman. Applications. Lecture Notes in Mathematics, 562:117–174, 1976. CODEN LNMAA2. ISBN 3-540-08055-4 (print), 3-540-37529-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087598/.

Goodman:1976:NLA

[Goo76c] Roe William Goodman. Nilpotent Lie algebras as tangent spaces. Lecture Notes in Mathematics, 562:33–66, 1976. CODEN LN-MAA2. ISBN 3-540-08055-4 (print), 3-540-37529-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087596/.

Goodman:1976:NLG

[Goo76d] Roe William Goodman. Nilpotent Lie Groups: Structure and Applications to Analysis, volume 562 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08055-4 (print), 3-540-37529-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 210 pp. LCCN QA3

.L28 no. 562. URL http://link.springer.com/book/10.1007/
BFb0087594;   http://www.springerlink.com/content/978-
3-540-37529-6.

Goodman:1976:SIS

[Goo76e] Roe William Goodman. Singular integrals on spaces of homogeneous type. Lecture Notes in Mathematics, 562:67–116, 1976. CODEN LNMAA2. ISBN 3-540-08055-4 (print), 3-540-37529-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087597/.

Goodman:1976:SNL

[Goo76f] Roe William Goodman. Structure of nilpotent Lie algebras and Lie groups. Lecture Notes in Mathematics, 562:1–32, 1976. CODEN LNMAA2. ISBN 3-540-08055-4 (print), 3-540-37529-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087595/.

Goodman:1978:TPV

[Goo78] Victor Goodman. Transition probabilities for vector-valued Brownian motion with boundaries. Lecture Notes in Mathematics, 644: 188–195, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066846/.

Goodearl:1979:SNR

[Goo79a] K. R. Goodearl. Simple Noetherian rings — the Zalesskii– Neroslavskii examples. Lecture Notes in Mathematics, 734:118– 130, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103156/.

Goodearl:1979:SSR

[Goo79b] K. R. Goodearl. The state space of K_{O} of a ring. Lecture Notes in Mathematics, 734:91–117, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103155/.

Goodman:1979:ALA

[Goo79c] Roe Goodman. Approximating Lie algebras of vector fields by nilpotent Lie algebras. Lecture Notes in Mathematics, 739:293– 307, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062497/.

Gordon:1976:SAN

[Gor76] Robert Gordon. Some aspects of noncommutative Noetherian rings. Lecture Notes in Mathematics, 545:105–127, 1976. CODEN LNMAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080307/.

Gordon:1978:SAC

[Gor78a] C. McA. Gordon. Some aspects of classical knot theory. Lecture Notes in Mathematics, 685:1–60, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062968/.

Gorniak:1978:RPD

[Gór78b] Janusz Górniak. Remarks on positive definite operator valued functions in linear spaces. Lecture Notes in Mathematics, 656: 37–44, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068809/.

Gossez:1978:CPD

[Gos78] Jean-Pierre Gossez. Sur certains problémes de Dirichlet fortement non linéaires à la resonance. (French) []. Lecture Notes in Mathematics, 665:100–107, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061800/.

Gottwald:1976:CSF

[Got76] Siegfried Gottwald. A cumulative system of fuzzy sets. Lecture Notes in Mathematics, 537:109–119, 1976. CODEN LN-MAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096897/.

Gottlieb:1978:LAF

[Got78a] Daniel Henry Gottlieb. Lifting actions in fibrations. Lecture Notes in Mathematics, 657:217–254, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069237/.

Gottlieb:1978:PT

[Got78b] Daniel Henry Gottlieb. Partial transfers. Lecture Notes in Mathematics, 657:255–266, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069238/.

Gaffney:1976:OI

[GP76] P. W. Gaffney and M. J. D. Powell. Optimal interpolation. Lecture Notes in Mathematics, 506:90–99, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080117/.

Glowinski:1977:SMF

[GP77] R. Glowinski and O. Pironneau. On solving a mixed finite element approximation of the Dirichlet problem for the biharmonic operator by a ‘quasi-direct’ method and various iterative methods. Lecture Notes in Mathematics, 606:167–193, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064462/.

Greub:1978:LSG

[GP78a] Werner Greub and Herbert-Rainer Petry. On the lifting of structure groups. Lecture Notes in Mathematics, 676:217–246, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063673/.

Gruson:1978:GCE

[GP78b] Laurent Gruson and Christian Peskine. Genre des courbes de l’espace projectif. (French) []. Lecture Notes in Mathematics, 687: 31–59, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062927/.

Goursaud:1979:APS

[GP79] J.-M. Goursaud and J.-L. Pascaud. Anneaux de polynômes semihéréditaires. (French) []. Lecture Notes in Mathematics, 734:131– 157, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103157/.

Garcia:1978:RSS

[GPR78] Pedro L. García and Antonio Pérez-Rendón. Reducibility of the symplectic structure of minimal interactions. Lecture Notes in Mathematics, 676:409–433, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063684/.

Gaer:1975:FDE

[GR75a] Marvin C. Gaer and Lee A. Rubel. The fractional derivative and entire functions. Lecture Notes in Mathematics, 457:171– 206, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067104/.

Glaser:1975:GTP

[GR75b] Prof. Dr. Leslie Curtis Glaser and Prof. Dr. Thomas Benjamin Rushing, editors. Geometric Topology: Proceedings of the Geometric Topology Conference held at Park City, Utah, February 19–22, 1974, volume 438 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 438. URL http:

//link.springer.com/book/10.1007/BFb0066101;   http://
www.springerlink.com/content/978-3-540-37412-1.

Ganssler:1976:EDP

[GR76] Peter Gänssler and Pál Révész, editors. Empirical Distributions and Processes: Selected Papers from a Meeting at Oberwolfach, March 28–April 3, 1976, volume 566 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 0-387-08061-9, 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 566; QA273.6 .E46; QA3 .L28 no. 566 QA273.6. URL http:

//link.springer.com/book/10.1007/BFb0096874;   http://
www.springerlink.com/content/978-3-540-37515-9.

Geymonat:1977:MEF

[GR77] G. Geymonat and M. Raous. Méthodes d’Éléments finis en viscoélasticité périodique. (French) []. Lecture Notes in Mathematics, 606:150–166, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064461/.

Gasser:1979:NCE

[GR79a] Th. Gasser and M. Rosenblatt. Nonparametric curve estimation. Lecture Notes in Mathematics, 757:1–4, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098487/.

Gasser:1979:STC

[GR79b] Th. Gasser and M. Rosenblatt, editors. Smoothing Techniques for Curve Estimation: Proceedings of a Workshop held in Heidelberg, April 2–4, 1979, volume 757 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 757; QA3

.L35 v.757. URL http://link.springer.com/book/10.1007/
BFb0098486;   http://www.springerlink.com/content/978-
3-540-38475-5.

Gerard:1979:EDS

[GR79c] Raymond Gérard and Jean-Pierre Ramis. Équations Différentielles et Systèmes de Pfaff dans le Champ Complexe, volume 712 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 364 pp. LCCN QA3

.L28 no. 712. URL http://link.springer.com/book/10.1007/
BFb0062810;   http://www.springerlink.com/content/978-
3-540-35314-0.

Girault:1979:E

[GR79d] Vivette Girault and Pierre-Arnaud Raviart. Erratum. Lecture Notes in Mathematics, 749:192, 1979. CODEN LNMAA2. ISBN 3-540-09557-8 (print), 3-540-34856-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063453.

Girault:1979:FEA

[GR79e] Vivette Girault and Pierre-Arnaud Raviart. Finite Element Approximation of the Navier–Stokes Equations, volume 749 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09557-8 (print), 3-540-34856-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 192 (est.) pp. LCCN QA3

.L28 no. 749. URL http://link.springer.com/book/10.1007/
BFb0063447;   http://www.springerlink.com/content/978-
3-540-34856-6.

Girault:1979:MFS

[GR79f] Vivette Girault and Pierre-Arnaud Raviart. Mathematical foundation of the Stokes problem. Lecture Notes in Mathematics, 749: 1–57, 1979. CODEN LNMAA2. ISBN 3-540-09557-8 (print), 3-540-34856-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063448/.

Girault:1979:MFE

[GR79g] Vivette Girault and Pierre-Arnaud Raviart. A mixed finite element method for solving the Stokes problem. Lecture Notes in Mathematics, 749:86–103, 1979. CODEN LNMAA2. ISBN 3-540-09557-8 (print), 3-540-34856-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063450/.

Girault:1979:NSS

[GR79h] Vivette Girault and Pierre-Arnaud Raviart. Numerical solution of the Stokes problem a classical method. Lecture Notes in Mathematics, 749:58–85, 1979. CODEN LNMAA2. ISBN 3-540-09557-8 (print), 3-540-34856-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063449/.

Girault:1979:SNS

[GR79i] Vivette Girault and Pierre-Arnaud Raviart. The stationary Navier–Stokes equations. Lecture Notes in Mathematics, 749:104– 147, 1979. CODEN LNMAA2. ISBN 3-540-09557-8 (print), 3-540-34856-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063451/.

Girault:1979:TDN

[GR79j] Vivette Girault and Pierre-Arnaud Raviart. The time-dependent Navier–Stokes equations. Lecture Notes in Mathematics, 749:148– 191, 1979. CODEN LNMAA2. ISBN 3-540-09557-8 (print), 3-540-34856-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063452/.

Grant:1975:SOG

[Gra75] Douglas D. Grant. Stability and operations on graphs. Lecture Notes in Mathematics, 452:116–135, 1975. CODEN LN-MAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069551/.

Grabmuller:1976:LPI

[Gra76a] Hans Grabmüller. On linear partial integro differential equations with a small parameter. Lecture Notes in Mathematics, 564:125– 134, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087333/.

Gramain:1976:SHR

[Gra76b] André Gramain. Sphères d’homologie rationnelle. (French) []. Lecture Notes in Mathematics, 514:35–52, 1976. CODEN LN-MAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080056/.

Grandell:1976:CCN

[Gra76c] Jan Grandell. Characterization and convergence of non-atomic random measures. Lecture Notes in Mathematics, 529:65–68, 1976. CODEN LNMAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077761/.

Grandell:1976:DBP

[Gra76d] Jan Grandell. Definitions and basic properties. Lecture Notes in Mathematics, 529:1–30, 1976. CODEN LNMAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077759/.

Grandell:1976:DSP

[Gra76e] Jan Grandell. Doubly Stochastic Poisson Processes, volume 529 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 234 pp. LCCN QA3

.L28 no. 529. URL http://link.springer.com/book/10.1007/
BFb0077758;   http://www.springerlink.com/content/978-
3-540-38258-4.

Grandell:1976:ERV

[Gra76f] Jan Grandell. Estimation of random variables. Lecture Notes in Mathematics, 529:86–157, 1976. CODEN LNMAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077763/.

Grandell:1976:ESO

[Gra76g] Jan Grandell. Estimation of second order properties of stationary doubly stochastic Poisson sequences. Lecture Notes in Mathematics, 529:190–204, 1976. CODEN LNMAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077765/.

Grandell:1976:LT

[Gra76h] Jan Grandell. Limit theorems. Lecture Notes in Mathematics, 529:68–86, 1976. CODEN LNMAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077762/.

Grandell:1976:LER

[Gra76i] Jan Grandell. Linear estimation of random variables in stationary doubly stochastic Poisson sequences. Lecture Notes in Mathematics, 529:158–189, 1976. CODEN LNMAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077764/.

Grandell:1976:SMR

[Gra76j] Jan Grandell. Some miscellaneous results. Lecture Notes in Mathematics, 529:31–64, 1976. CODEN LNMAA2. ISBN 3-540-07795-2 (print), 3-540-38258-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077760/.

Grayson:1976:HAT

[Gra76k] Daniel Grayson. Higher algebraic K-theory: II. Lecture Notes in Mathematics, 551:217–240, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080003/.

Gramain:1977:RTC

[Gra77] André Gramain. Rapport sur la théorie classique des noeuds. (French) []. Lecture Notes in Mathematics, 567:222–237, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096072/.

Graef:1978:OFI

[Gra78a] F. Graef. Optimal filtering of infinite-dimensional stationary signals. Lecture Notes in Mathematics, 695:63–75, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062655/.

Gramain:1978:FEA

[Gra78b] François Gramain. Fonctions entières arithmétiques. (French) []. Lecture Notes in Mathematics, 694:96–125, 1978. CODEN LN-MAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063245/.

Grasman:1979:SRP

[Gra79a] J. Grasman. Small random perturbations of dynamical systems with applications in population genetics. Lecture Notes in Mathematics, 711:158–175, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062952/.

Gray:1979:FHS

[Gra79b] John W. Gray. Fragments of the history of sheaf theory. Lecture Notes in Mathematics, 753:1–79, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061812/.

Grayson:1979:HVM

[Gra79c] R. J. Grayson. Heyting-valued models for intuitionistic set theory. Lecture Notes in Mathematics, 753:402–414, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061825/.

Green:1975:CPF

[Gre75a] Judy Green. Consistency properties for finite quantifier languages. Lecture Notes in Mathematics, 492:73–123, 1975. CODEN LN-MAA2. ISBN 3-540-07419-8 (print), 3-540-37949-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081121/.

Greenberg:1975:ABS

[Gre75b] J. M. Greenberg. Asymptotic behavior of solutions to the quasilinear wave equation. Lecture Notes in Mathematics, 446:198– 246, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070604/.

Gregory:1975:FCI

[Gre75c] John Gregory. On a finiteness condition for infinitary languages. Lecture Notes in Mathematics, 492:143–206, 1975. CODEN LN-MAA2. ISBN 3-540-07419-8 (print), 3-540-37949-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081123/.

Green:1976:GRD

[Gre76] Sherry M. Green. Generators and relations for K_{2} of a division ring. Lecture Notes in Mathematics, 551:74–76, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079996/.

Greene:1977:SOT

[Gre77a] Curtis Greene. Some order-theoretic properties of the Robinson– Schensted correspondence. Lecture Notes in Mathematics, 579: 114–120, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090013/.

Greub:1977:CLB

[Gre77b] Werner H. Greub. Complex line bundles and the magnetic field of a monopole. Lecture Notes in Mathematics, 570:350–354, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087790/.

Green:1978:AET

[Gre78a] Leon W. Green. Applications of ergodic theory to geometry. Lecture Notes in Mathematics, 668:133–141, 1978. CODEN LN-MAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101786/.

Greenwell:1978:OCP

[Gre78b] Don Greenwell. Odd cycles and perfect graphs. Lecture Notes in Mathematics, 642:191–193, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070376/.

Greve:1979:LCM

[Gre79] Georg Greve. Lifting closed and monoidal structures along semitopological functors. Lecture Notes in Mathematics, 719:74– 83, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065260/.

Griego:1975:DMO

[Gri75] Richard Griego. Dual multiplicative operator functionals. Lecture Notes in Mathematics, 451:152–162, 1975. CODEN LN-MAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068585/.

Grigelionis:1976:MAS

[Gri76a] B. Grigelionis. On the martingale approach to statistical problems for stochastic processes with boundary conditions. Lecture Notes in Mathematics, 550:160–180, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077489/.

Grisvard:1976:SSM

[Gri76b] Pierre Grisvard. Smoothness of the solution of a monotonic boundary value problem for a second order elliptic equation in a general convex domain. Lecture Notes in Mathematics, 564:135–151, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087334/.

Grigorieff:1977:DJB

[Gri77] Serge Grigorieff. Détermination des jeux boréliens et problèmes logiques associés [d’après D. Martin]. (French) []. Lecture Notes in Mathematics, 567:122–135, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096065/.

Griffeath:1979:ACI

[Gri79a] David Griffeath. Additive and Cancellative Interacting Particle Systems, volume 724 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09508-X (print), 3-540-35174-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 108 pp. LCCN QA3 .L28 no. 724. URL http:

//link.springer.com/book/10.1007/BFb0067306;   http://
www.springerlink.com/content/978-3-540-35174-0.

Griffeath:1979:AS

[Gri79b] David Griffeath. Additive systems. Lecture Notes in Mathematics, 724:14–65, 1979. CODEN LNMAA2. ISBN 3-540-09508-X (print), 3-540-35174-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067308/.

Griffeath:1979:CS

[Gri79c] David Griffeath. Cancellative systems. Lecture Notes in Mathematics, 724:66–88, 1979. CODEN LNMAA2. ISBN 3-540-09508-X (print), 3-540-35174-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067309/.

Griffeath:1979:I

[Gri79d] David Griffeath. Introduction. Lecture Notes in Mathematics, 724:1–13, 1979. CODEN LNMAA2. ISBN 3-540-09508-X (print), 3-540-35174-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067307/.

Griffeath:1979:UN

[Gri79e] David Griffeath. Uniqueness and nonuniqueness. Lecture Notes in Mathematics, 724:89–100, 1979. CODEN LNMAA2. ISBN 3-540-09508-X (print), 3-540-35174-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067310/.

Grmela:1979:DDS

[Grm79] Miroslav Grmela. Dissipative dynamical systems of macroscopic physics. Lecture Notes in Mathematics, 755:35–51, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069803/.

Gronau:1976:BSL

[Gro76] Detlef Gronau. Bewegliche Singularitäten von linearen partiellen Differential-Gleichungen. (German) []. Lecture Notes in Mathematics, 561:218–226, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087638/.

Gross:1977:FMF

[Gro77a] Fred Gross. Factorization of meromorphic functions and some open problems. Lecture Notes in Mathematics, 599:51–67, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096825/.

Grothendieck:1977:CDF

[Gro77b] A. Grothendieck. Complexes Dualsants. (French) [Dualsants complexes]. Lecture Notes in Mathematics, 589:1–72, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096803/.

Grothendieck:1977:FLF

[Gro77c] A. Grothendieck. Formule de Lefschetz. (French) []. Lecture Notes in Mathematics, 589:73–137, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096804/.

Grothendieck:1977:FDP

[Gro77d] A. Grothendieck. Formule d’Euler-Poincaré en cohomologie etale. (French) []. Lecture Notes in Mathematics, 589:372–406, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096810/.

Grothendieck:1977:FNW

[Gro77e] A. Grothendieck. Formules de Nielsen–Wecken et de Lefschetz en géométrie algébrique. (French) []. Lecture Notes in Mathematics, 589:407–441, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096811/.

Grothendieck:1977:GCC

[Gro77f] A. Grothendieck. Groupes de classes des catégories abeliennes et triangulées. Complexes parfaits. (French) []. Lecture Notes in Mathematics, 589:351–371, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096809/.

Gross:1978:IMC

[Gro78a] Jonathan L. Gross. Imbeddings of metacyclic Cayley graphs. Lecture Notes in Mathematics, 642:194–203, 1978. CODEN LN-MAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070377/.

Gross:1978:LSI

[Gro78b] Leonard Gross. Logarithmic Sobolev inequalities — a survey. Lecture Notes in Mathematics, 644:196–203, 1978. CODEN LN-MAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066847/.

Grosswald:1978:A

[Gro78c] Emil Grosswald. Applications. Lecture Notes in Mathematics, 698: 131–149, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063149/.

Grosswald:1978:APB

[Gro78d] Emil Grosswald. Asymptotic properties of the BP. Lecture Notes in Mathematics, 698:124–130, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063148/.

Grosswald:1978:BP

[Gro78e] Emil Grosswald. Bessel Polynomials, volume 698 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 182 pp. LCCN QA3

.L28 no. 698. URL http://link.springer.com/book/10.1007/
BFb0063135;   http://www.springerlink.com/content/978-
3-540-35548-9.

Grosswald:1978:BPB

[Gro78f] Emil Grosswald. Bessel polynomials and Bessel functions: Differential equations and their solutions. Lecture Notes in Mathematics, 698:4–17, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063137/.

Grosswald:1978:BCF

[Gro78g] Emil Grosswald. The BP and continued fractions. Lecture Notes in Mathematics, 698:59–63, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063143/.

Grosswald:1978:EFS

[Gro78h] Emil Grosswald. Expansions of functions in series of BP. Lecture Notes in Mathematics, 698:64–74, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063144/.

Grosswald:1978:FRT

[Gro78i] Emil Grosswald. Formulas of Rodrigues’ type. Lecture Notes in Mathematics, 698:51–58, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063142/.

Grosswald:1978:GGB

[Gro78j] Emil Grosswald. The Galois group of B.P. Lecture Notes in Mathematics, 698:116–123, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063147/.

Grosswald:1978:GF

[Gro78k] Emil Grosswald. Generating functions. Lecture Notes in Mathematics, 698:41–50, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063141/.

Grosswald:1978:HS

[Gro78l] Emil Grosswald. Historic sketch. Lecture Notes in Mathematics, 698:1–3, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063136/.

Grosswald:1978:M

[Gro78m] Emil Grosswald. Miscellanea. Lecture Notes in Mathematics, 698: 150–161, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063150/.

Grosswald:1978:MOU

[Gro78n] Emil Grosswald. Moments and orthogonality on the unit circle. Lecture Notes in Mathematics, 698:25–33, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063139/.

Grosswald:1978:AIB

[Gro78o] Emil Grosswald. On the algebraic irreducibility of the BP. Lecture Notes in Mathematics, 698:99–115, 1978. CODEN LN-MAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063146/.

Grosswald:1978:PZB

[Gro78p] Emil Grosswald. Properties of the zeros of BP. Lecture Notes in Mathematics, 698:75–98, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063145/.

Grosswald:1978:RR

[Gro78q] Emil Grosswald. Recurrence relations. Lecture Notes in Mathematics, 698:18–24, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063138/.

Grosswald:1978:RBC

[Gro78r] Emil Grosswald. Relations of the BP to classical orthonormal polynomials and to other functions. Lecture Notes in Mathematics, 698:34–40, 1978. CODEN LNMAA2. ISBN 3-540-09104-1 (print), 3-540-35548-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063140/.

Grosser:1979:BGG

[Gro79a] Michael Grosser. Bezeichnungen, Grundbegriffe. (German) []. Lecture Notes in Mathematics, 717:5–10, 1979. CODEN LN-MAA2. ISBN 3-540-09257-9 (print), 3-540-35277-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064149/.

Grosser:1979:BVB

[Gro79b] Michael Grosser. Bidualräume und Vervollständigungen von Banachmoduln, volume 717 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09257-9 (print), 3-540-35277-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 209 pp. LCCN QA3

.L28 no. 717. URL http://link.springer.com/book/10.1007/
BFb0064147;   http://www.springerlink.com/content/978-
3-540-35277-8.

Grosser:1979:ZZB

[Gro79c] Michael Grosser. Der Zusammenhang Zwischen Den Bidualräumen Und Den Multiplierräumen. (German) []. Lecture Notes in Mathematics, 717:144–198, 1979. CODEN LNMAA2. ISBN 3-540-09257-9 (print), 3-540-35277-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064153/.

Grosser:1979:BG

[Gro79d] Michael Grosser. Die Bidualräume. (German) []. Lecture Notes in Mathematics, 717:11–45, 1979. CODEN LNMAA2. ISBN 3-540-09257-9 (print), 3-540-35277-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064150/.

Grosser:1979:EG

[Gro79e] Michael Grosser. Einleitung. (German) []. Lecture Notes in Mathematics, 717:1–4, 1979. CODEN LNMAA2. ISBN 3-540-09257-9 (print), 3-540-35277-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064148/.

Grosser:1979:MCG

[Gro79f] Michael Grosser. Multiplier und Centralizer. (German) []. Lecture Notes in Mathematics, 717:46–112, 1979. CODEN LN-MAA2. ISBN 3-540-09257-9 (print), 3-540-35277-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064151/.

Grosser:1979:TMG

[Gro79g] Michael Grosser. Teilräume von H(X, X) als Moduln. (German) []. Lecture Notes in Mathematics, 717:113–143, 1979. CODEN LNMAA2. ISBN 3-540-09257-9 (print), 3-540-35277-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064152/.

Grossman:1979:SNR

[Gro79h] Stanley I. Grossman. Some notes on the resolvents of Volterra integral equations. Lecture Notes in Mathematics, 737:88–91, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064498/.

Grunbaum:1975:P

[Grü75a] Branko Grünbaum. Polygons. Lecture Notes in Mathematics, 490: 147–184, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081139/.

Grunbaum:1975:IPN

[Grü75b] F. Alberto Grünbaum. Inverse problems for nonlinear random systems. Lecture Notes in Mathematics, 446:247–263, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070605/.

Gruson:1975:SCF

[Gru75c] Laurent Gruson. Simple coherent functors. Lecture Notes in Mathematics, 488:156–159, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081220/.

Gruman:1977:GGE

[Gru77] Lawrence Gruman. La géométrie globale des ensembles analytiques. (French) []. Lecture Notes in Mathematics, 578:236– 245, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091473/.

Gaenssler:1976:UCM

[GS76] P. Gaenssler and W. Stute. On uniform convergence of measures with applications to uniform convergence of empirical distributions. Lecture Notes in Mathematics, 566:45–56, 1976. CODEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096878/.

Gluck:1977:SPD

[GS77] Herman Gluck and David Singer. Scattering problems in differential geometry. Lecture Notes in Mathematics, 597:191–229, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085356/.

Gallardo:1979:TSP

[GS79a] Léonard Gallardo and René Schott. Un théorème de structure pour les sous-groupes fermés, connexes des groupes extensions compactes de groupes nilpotents. (French) []. Lecture Notes in Mathematics, 739:283–292, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062496/.

Gerard:1979:ECS

[GS79b] R. Gérard and Y. Sibuya. Étude de certains systèmes de Pfaff avec singularitès. (French) []. Lecture Notes in Mathematics, 712: 131–288, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062819/.

Goldblatt:1975:ACP

[GT75] R. I. Goldblatt and S. K. Thomason. Axiomatic classes in propositional modal logic. Lecture Notes in Mathematics, 450:163– 173, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062855/.

Galligani:1977:ITS

[GT77] Ilio Galligani and Donato Trigiante. Integration techniques for solving algebraic systems. Lecture Notes in Mathematics, 606: 113–121, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064459/.

Gualtierotti:1978:MGC

[Gua78] A. F. Gualtierotti. Mixtures of Gaussian cylinder set measures and abstract Wiener spaces as models for detection of signals imbedded in noise. Lecture Notes in Mathematics, 656:45–57, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068810/.

Gudkenheimer:1976:CVW

[Gud76] John Gudkenheimer. Constant velocity waves in oscillating chemical reactions. Lecture Notes in Mathematics, 525:99–103, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077846/.

Guelorget:1975:AWD

[Gue75] S. Guelorget. Algèbre de Well du groupe linéaire application aux classes caractéristiques d’un feuilletage. (French) []. Lecture Notes in Mathematics, 484:179–191, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082151/.

Guerra:1978:STP

[Gue78] P. J. Guerra. Stability of tensor products of Radon measures of type (H). Lecture Notes in Mathematics, 645:97–108, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069666/.

Gugenheim:1978:MEM

[Gug78] V. K. A. M. Gugenheim. On a modified Eilenberg–Moore theorem. Lecture Notes in Mathematics, 658:177–190, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068716/.

Guillemin:1975:CIT

[Gui75] Victor Guillemin. Clean intersection theory and Fourier integrals. Lecture Notes in Mathematics, 459:23–35, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074191/.

Guichardet:1976:TEM

[Gui76a] A. Guichardet. Un théorème d’existence de mesures invariantes. (French) []. Lecture Notes in Mathematics, 532:125–130, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080175/.

Guivarch:1976:EEH

[Gui76b] Yves Guivarc’h. Équirépartition dans les espaces homogènes. (French) []. Lecture Notes in Mathematics, 532:131–142, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080176/.

Guichardet:1977:RSG

[Gui77a] Alain Guichardet. Représentations de G^{X} selon Gelfand et Delorme. (French) []. Lecture Notes in Mathematics, 567:238– 255, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096073/.

Guillemonat:1977:FGI

[Gui77b] Alain Guillemonat. On finite generation of invariants for certain sub-algebras of a semi-simple Lie algebra. Lecture Notes in Mathematics, 587:77–90, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087915/.

Guiraud:1977:DRV

[Gui77c] Jean-Pierre Guiraud. The dynamics of rolled vortex sheets tightly winded around slender vortex filaments in inviscid incompressible flow. Lecture Notes in Mathematics, 594:244–259, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086090/.

Guichardet:1979:CRU

[Gui79a] Alain Guichardet. Sur la 1-cohomologie des représentations unitaires de certains groupes de Lie. (French) []. Lecture Notes in Mathematics, 739:308–314, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062498/.

Guillerme:1979:EHC

[Gui79b] Jean Guillerme. Extremalisation harmonique d’un couple de fonctions continues. (French) []. Lecture Notes in Mathematics, 713:

89–112, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071322/.

Gupta:1978:CIC

[Gup78] Ram P. Gupta. On the chromatic index and the cover index of a multigraph. Lecture Notes in Mathematics, 642:204–215, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070378/.

Gustafson:1975:NTM

[Gus75] Sven-Åke Gustafson. On the numerical treatment of a multidimensional parabolic boundary-value control problem. Lecture Notes in Mathematics, 477:121–126, 1975. CODEN LN-MAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079171/.

Gusak:1976:PFE

[Gus76a] D. V. Gusak. Probabilities of the first exit for continuous processes with independent increments on a Markov chain. Lecture Notes in Mathematics, 550:181–193, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077490/.

Gustafson:1976:RCG

[Gus76b] William H. Gustafson. A remark on class groups. Lecture Notes in Mathematics, 549:92–99, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077336/.

Gutierrez:1975:SSL

[Gut75] Carlos Gutiérrez. Structural stability for line fields with singularities on two-manifolds. Lecture Notes in Mathematics, 468: 17–19, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082582/.

Guzicki:1977:PCC

[Guz77] Wojciech Guzicki. On the projective class of the continuum hypothesis. Lecture Notes in Mathematics, 619:181–185, 1977. CO-DEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0067651/.

Golub:1979:TLS

[GV79a] Gene H. Golub and Charles Van Loan. Total least squares. Lecture Notes in Mathematics, 757:69–76, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098490/.

Grasman:1979:AMV

[GV79b] J. Grasman and E. J. M. Veling. Asymptotic methods for the Volterra–Lotka equations. Lecture Notes in Mathematics, 711: 146–157, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062951/.

Greco:1979:TA

[GV79c] Silvio Greco and Paolo Valabrega. On the theory of adjoints. Lecture Notes in Mathematics, 732:98–123, 1979. CODEN LN-MAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066640/.

Graf:1976:ELD

[GvW76] Siegfried Graf and Heinrich von Weizsäcker. On the existence of lower densities in noncomplete measure spaces. Lecture Notes in Mathematics, 541:155–158, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081048/.

Gatica:1976:SFD

[GW76a] Juan A. Gatica and Paul Waltman. A singular functional differential equation arising in an immunological model. Lecture Notes in Mathematics, 564:114–124, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087332/.

Geramita:1976:SNC

[GW76b] Anthony V. Geramita and Jennifer Seberry Wallis. Some new constructions for orthogonal designs. Lecture Notes in Mathematics, 560:46–54, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097367/.

Greene:1979:BM

[GW79a] R. E. Greene and H. Wu. The Bergman metric. Lecture Notes in Mathematics, 699:141–179, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063422/.

Greene:1979:BEF

[GW79b] R. E. Greene and H. Wu. Bounded exhaustion functions and a new class of hyperbolic Kähler manifolds. Lecture Notes in Mathematics, 699:99–115, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063420/.

Greene:1979:D

[GW79c] R. E. Greene and H. Wu. A digression. Lecture Notes in Mathematics, 699:116–140, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063421/.

Greene:1979:FTM

[GW79d] R. E. Greene and H. Wu. Function Theory on Manifolds Which Possess a Pole, volume 699 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 215 pp. LCCN QA3

.L28 no. 699. URL http://link.springer.com/book/10.1007/
BFb0063413;   http://www.springerlink.com/content/978-
3-540-35536-6.

Greene:1979:HCT

[GW79e] R. E. Greene and H. Wu. The Hessian comparison theorem. Lecture Notes in Mathematics, 699:19–42, 1979. CODEN LN-MAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063416/.

Greene:1979:I

[GW79f] R. E. Greene and H. Wu. Introduction. Lecture Notes in Mathematics, 699:1–4, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063414/.

Greene:1979:MB

[GW79g] R. E. Greene and H. Wu. Manifolds biholomorphic to C^{n}. Lecture Notes in Mathematics, 699:180–205, 1979. CODEN LN-MAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063423/.

Greene:1979:NCH

[GW79h] R. E. Greene and H. Wu. New criterion of hyperbolicity. Lecture Notes in Mathematics, 699:81–98, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063419/.

Greene:1979:P

[GW79i] R. E. Greene and H. Wu. Preliminaries. Lecture Notes in Mathematics, 699:5–18, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063415/.

Greene:1979:QIE

[GW79j] R. E. Greene and H. Wu. Quasi-isometry of the exponential map and the absence of positive harmonic functions. Lecture Notes in Mathematics, 699:56–80, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063418/.

Greene:1979:SMV

[GW79k] R. E. Greene and H. Wu. Sub-mean-value theorem for subharmonic functions. Lecture Notes in Mathematics, 699:43–55, 1979. CODEN LNMAA2. ISBN 3-540-09108-4 (print), 3-540-35536-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063417/.

Gibson:1976:I

[GWdPL76] Christopher G. Gibson, Klaus Wirthmüller, Andrew A. du Plessis, and Eduard J. N. Looijenga. Introduction. Lecture Notes in Mathematics, 552:1–7, 1976. CODEN LNMAA2. ISBN 3-540-07997-1 (print), 3-540-37957-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095245/.

Haagerup:1979:DTL

[Haa79] Uffe Haagerup. A density theorem for left Hilbert algebras. Lecture Notes in Mathematics, 725:170–179, 1979. CODEN LN-MAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062616/.

Habets:1976:MSC

[Hab76] P. Habets. On the method of strained coordinates. Lecture Notes in Mathematics, 564:152–162, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087335/.

Hackbusch:1978:FIM

[Hac78a] W. Hackbusch. A fast iterative method for solving Poisson’s equation in a general region. Lecture Notes in Mathematics, 631:51– 62, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067463/.

Hackenbroch:1978:RNT

[Hac78b] Wolfgang Hackenbroch. On the Radon–Nikodým theorem for operator measures and its applications to prediction and linear systems theory. Lecture Notes in Mathematics, 695:193–206, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062666/.

Hadeler:1976:NDE

[Had76] K. P. Hadeler. Nonlinear diffusion equations in biology. Lecture Notes in Mathematics, 564:163–206, 1976. CODEN LN-MAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087336/.

Hadeler:1979:DEB

[Had79] K. P. Hadeler. Delay equations in biology. Lecture Notes in Mathematics, 730:136–156, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064316/.

Haefliger:1975:CGF

[Hae75] A. Haefliger. Sur la cohomologie de Gelfand–Fuchs. (French) []. Lecture Notes in Mathematics, 484:121–152, 1975. CODEN LN-MAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082148/.

Haefliger:1978:CLA

[Hae78a] André Haefliger. Cohomology of Lie algebras and foliations. Lecture Notes in Mathematics, 652:1–12, 1978. CODEN LN-MAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063499/.

Haefliger:1978:WPD

[Hae78b] André Haefliger. Whitehead products and differential forms. Lecture Notes in Mathematics, 652:13–24, 1978. CODEN LN-MAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063500/.

Haggard:1978:PDF

[Hag78] Gary Haggard. Pruning and depth first search. Lecture Notes in Mathematics, 642:216–224, 1978. CODEN LNMAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070379/.

Hahn:1976:WSO

[Hah76] Marjorie G. Hahn. What second-order Lipschitz conditions imply the CLT? Lecture Notes in Mathematics, 526:107–111, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082346/.

Hahn:1979:GDA

[Hah79] Marjorie G. Hahn. The generalized domain of attraction of a Gaussian law on Hilbert space. Lecture Notes in Mathematics, 709:125–144, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071954/.

Hajek:1978:SRD

[Háj78] Petr Hájek. Some results on degrees of constructibility. Lecture Notes in Mathematics, 669:55–71, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103102/.

Hale:1976:DDP

[Hal76a] Jack K. Hale. Discrete dissipative processes. Lecture Notes in Mathematics, 564:207–224, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087337/.

Hall:1976:IDE

[Hal76b] William S. Hall. Integrating a differential equation with a weak* continuous vector field. Lecture Notes in Mathematics, 564:225– 230, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087338/.

Halpern:1977:CEG

[Hal77] Leopold Halpern. Conclusions from an extended gauge principle of Dirac’s equation. Lecture Notes in Mathematics, 570:355– 364, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087791/.

Halkin:1978:NCO

[Hal78a] Hubert Halkin. Necessary conditions for optimal control problems with differentiable or nondifferentiable data. Lecture Notes in Mathematics, 680:77–118, 1978. CODEN LNMAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065313/.

Hall:1978:IPC

[Hal78b] Marshall Hall, Jr. Integral properties of combinatorial matrices. Lecture Notes in Mathematics, 686:22–33, 1978. CODEN LN-MAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062513/.

Halmos:1978:IO

[Hal78c] P. R. Halmos. Integral operators. Lecture Notes in Mathematics, 693:1–15, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064657/.

Hale:1979:REI

[Hal79a] Jack K. Hale. Retarded equations with infinite delays. Lecture Notes in Mathematics, 730:157–193, 1979. CODEN LN-MAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064317/.

Haley:1979:DVR

[Hal79b] David K. Haley. Discriminator varieties and m-rings. Lecture Notes in Mathematics, 745:85–136, 1979. CODEN LN-MAA2. ISBN 3-540-09548-9 (print), 3-540-34863-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062806/.

Haley:1979:ECR

[Hal79c] David K. Haley. Equational Compactness in Rings: With Applications to the Theory of Topological Rings, volume 745 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09548-9 (print), 3-540-34863-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 167 pp. LCCN QA3

.L28 no. 745. URL http://link.springer.com/book/10.1007/
BFb0062801;   http://www.springerlink.com/content/978-
3-540-34863-4.

Haley:1979:I

[Hal79d] David K. Haley. Introduction. Lecture Notes in Mathematics, 745:1–8, 1979. CODEN LNMAA2. ISBN 3-540-09548-9 (print), 3-540-34863-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062802/.

Haley:1979:MCC

[Hal79e] David K. Haley. Minimum conditions and compactifications. Lecture Notes in Mathematics, 745:40–64, 1979. CODEN LN-MAA2. ISBN 3-540-09548-9 (print), 3-540-34863-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062804/.

Haley:1979:MTF

[Hal79f] David K. Haley. The model theoretic framework. Lecture Notes in Mathematics, 745:9–39, 1979. CODEN LNMAA2. ISBN 3-540-09548-9 (print), 3-540-34863-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062803/.

Haley:1979:MQ

[Hal79g] David K. Haley. The Mycielski question. Lecture Notes in Mathematics, 745:137–158, 1979. CODEN LNMAA2. ISBN 3-540-09548-9 (print), 3-540-34863-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062807/.

Haley:1979:RAC

[Hal79h] David K. Haley. Rings with the ascending chain condition. Lecture Notes in Mathematics, 745:65–84, 1979. CODEN LNMAA2. ISBN 3-540-09548-9 (print), 3-540-34863-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062805/.

Hall:1979:RVP

[Hal79i] W. S. Hall. The Rayleigh and van der Pol wave equations, some generalizations. Lecture Notes in Mathematics, 703:130–138, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067265/.

Hamernik:1975:IMC

[Ham75a] Wolfgang Hamernik. Indecomposable modules with cyclic vertex. Lecture Notes in Mathematics, 488:160–167, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081221/.

Hamilton:1975:FS

[Ham75b] Prof. Richard S. Hamilton. Function spaces. Lecture Notes in Mathematics, 471:9–55, 1975. CODEN LNMAA2. ISBN 3-540-07185-7 (print), 3-540-37530-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087229/.

Hamilton:1975:GEC

[Ham75c] Prof. Richard S. Hamilton. Growth estimates and convergence. Lecture Notes in Mathematics, 471:125–165, 1975. CODEN LN-MAA2. ISBN 3-540-07185-7 (print), 3-540-37530-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087232/.

Hamilton:1975:HM

[Ham75d] Prof. Richard S. Hamilton. Harmonic maps. Lecture Notes in Mathematics, 471:3–8, 1975. CODEN LNMAA2. ISBN 3-540-07185-7 (print), 3-540-37530-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087228/.

Hamilton:1975:HMM

[Ham75e] Prof. Richard S. Hamilton. Harmonic Maps of Manifolds with Boundary, volume 471 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07185-7 (print), 3-540-37530-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 168 pp. LCCN QA3 .L28 no. 471. URL http:

//link.springer.com/book/10.1007/BFb0087227;   http://
www.springerlink.com/content/978-3-540-37530-2.

Hamilton:1975:HEM

[Ham75f] Prof. Richard S. Hamilton. The heat equation for manifolds. Lecture Notes in Mathematics, 471:101–124, 1975. CODEN LN-MAA2. ISBN 3-540-07185-7 (print), 3-540-37530-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087231/.

Hamilton:1975:SEP

[Ham75g] Prof. Richard S. Hamilton. Semi-elliptic and parabolic equations. Lecture Notes in Mathematics, 471:56–100, 1975. CODEN LN-MAA2. ISBN 3-540-07185-7 (print), 3-540-37530-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087230/.

Hammerlin:1976:NBH

[Häm76] Prof. Dr. Günther Hämmerlin. Zur Numerischen Behandlung von Homogenen Fredholmschen Integralgleichungen 2. Art mit Splines. (German) []. Lecture Notes in Mathematics, 501:92–98, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079742/.

Hanatani:1975:CPR

[Han75a] Yoshito Hanatani. Calculability of the primitive recursive functionals of finite type over the natural numbers. Lecture Notes in Mathematics, 500:152–163, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079550/.

Handel:1975:ASP

[Han75b] Michael Handel. Approximating stratum preserving CE maps between CS sets by stratum preserving homeomorphisms. Lecture Notes in Mathematics, 438:245–250, 1975. CODEN LN-MAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066120/.

Hansel:1976:IHF

[Han76a] G. Hansel. Sur les inégalités de Holley. (French) []. Lecture Notes in Mathematics, 532:143–151, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080177/.

Hansen:1976:CSF

[Han76b] Wolfhard Hansen. Cônes simpliciaux de fonctions surhamoniques. (French) []. Lecture Notes in Mathematics, 518:91–94, 1976. CO-DEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080406/.

Hanani:1978:CTD

[Han78a] Haim Hanani. A class of three-designs. Lecture Notes in Mathematics, 686:34–46, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062514/.

Handel:1978:ATP

[Han78b] David Handel. Algebraic-topological problems in approximation theory. Lecture Notes in Mathematics, 657:267–270, 1978. CO-DEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069239/.

Hanks:1978:HCA

[Han78c] C. H. Hanks. On the Hopf construction associated with a composition. Lecture Notes in Mathematics, 658:191–205, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068717/.

Hannsgen:1979:IEP

[Han79a] Kenneth B. Hannsgen. An integrodifferential equation with parameter. Lecture Notes in Mathematics, 737:92–98, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064499/.

Hansen:1979:DP

[Han79b] W. Hansen. The Dirichlet problem. Lecture Notes in Mathematics, 703:139–144, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067266/.

Hartung:1975:PMK

[Har75a] Joachim Hartung. Penalty-Methoden für Kontrollprobleme und Open-Loop-Differential-Spiele. (German) []. Lecture Notes in Mathematics, 477:127–144, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079172/.

Harzallah:1975:DCR

[Har75b] Khélifa Harzallah. Distributions coniques et représentations associées à SO_{o}(1, q). (French) [Conical distributions and representations associated with SO_{o}(1, q)]. Lecture Notes in Mathematics, 497:211–229, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078018/.

Harris:1976:GCC

[Har76a] Bruno Harris. Group cohomology classes with differential form coefficients. Lecture Notes in Mathematics, 551:278–282, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080008/.

Harris:1976:AI

[Har76b] W. A. Harris, Jr. On asymptotic integration. Lecture Notes in Mathematics, 564:231–238, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087339/.

Harper:1978:TFS

[Har78a] John R. Harper. Torsion free modp H-spaces. Lecture Notes in Mathematics, 657:144–149, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069230/.

Harvey:1978:OSI

[Har78b] Bernard N. Harvey. An operator not a shift, integral, nor multiplication. Lecture Notes in Mathematics, 693:101–102, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064665.

Harvey:1978:KGS

[Har78c] William James Harvey. Kleinian groups [a survey]. Lecture Notes in Mathematics, 677:30–45, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070752/.

Hardie:1979:KLE

[Har79a] K. A. Hardie. Kan lift-extensions in C. G. Haus. Lecture Notes in Mathematics, 719:94–101, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065262/.

Harvey:1979:TFF

[Har79b] J. Martin Harvey. Topological functors from factorization. Lecture Notes in Mathematics, 719:102–111, 1979. CODEN LN-MAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065263/.

Hastings:1978:EHI

[Has78a] Harold M. Hastings. On expansive homeomorphisms of the infinite torus. Lecture Notes in Mathematics, 668:142–149, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101787/.

Hastings:1978:STD

[Has78b] Harold M. Hastings. Shape theory and dynamical systems. Lecture Notes in Mathematics, 668:150–159, 1978. CODEN LN-MAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101788/.

Hausmann:1976:HSB

[Hau76] Jean-Claude Hausmann. Homology sphere bordism and Quillen plus construction. Lecture Notes in Mathematics, 551:170–181, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080000/.

Hausen:1977:JRS

[Hau77] Jutta Hausen. The Jacobson radical of some endomorphism rings. Lecture Notes in Mathematics, 616:332–336, 1977. CODEN LN-MAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068204/.

Hausmann:1978:KTP

[Hau78a] Jean-Claude Hausmann, editor. Knot Theory: Proceedings, Planssur-Bex, Switzerland 1977, volume 685 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 685. URL http:

//link.springer.com/book/10.1007/BFb0062967;   http://
www.springerlink.com/content/978-3-540-35705-6.

Hausmann:1978:NAF

[Hau78b] Jean-Claude Hausmann. Noeuds antisimples. (French) [Antisimple knots]. Lecture Notes in Mathematics, 685:171–202, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062971/.

Hausmann:1979:ATF

[Hau79] Jean-Claude Hausmann. Algebraic K-theory and flat manifolds. Lecture Notes in Mathematics, 763:212–234, 1979. CODEN LN-MAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088088/.

Hayman:1976:DWH

[Hay76] W. K. Hayman. On the domains where a harmonic or subharmonic function is positive. Lecture Notes in Mathematics, 505: 24–55, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081098/.

Hayman:1977:VGF

[Hay77] W. K. Hayman. Values and growth of functions regular in the unit disk. Lecture Notes in Mathematics, 599:68–75, 1977. CO-DEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096826/.

Hayes:1978:WHI

[Hay78] Sandra Hayes. Weakly homogeneous infinite dimensional manifolds. Lecture Notes in Mathematics, 670:303–311, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064405/.

Hazod:1977:AFT

[Haz77a] Wilfried Hazod. Anhang. Faltungshalbgruppen mit nicht trivialen idempotenten Faktoren. (German) []. Lecture Notes in Mathematics, 595:136–146, 1977. CODEN LNMAA2. ISBN 3-540-08259-X (print), 3-540-37339-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091069/.

Hazod:1977:HKM

[Haz77b] Wilfried Hazod. Halbgruppen komplexer Maße und dissipative Distributionen. (German) []. Lecture Notes in Mathematics, 595: 107–135, 1977. CODEN LNMAA2. ISBN 3-540-08259-X (print), 3-540-37339-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091068/.

Hazod:1977:HMH

[Haz77c] Wilfried Hazod. Halbgruppen von Maßen und Halbgruppen invarianter Operatoren. (German) []. Lecture Notes in Mathematics, 595:1–27, 1977. CODEN LNMAA2. ISBN 3-540-08259-X (print), 3-540-37339-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091064/.

Hazod:1977:MED

[Haz77d] Wilfried Hazod. Mischungen erzeugender Distributionen und Zufallsentwicklungen. (German) []. Lecture Notes in Mathematics, 595:66–88, 1977. CODEN LNMAA2. ISBN 3-540-08259-X (print), 3-540-37339-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091066/.

Hazod:1977:SFW

[Haz77e] Wilfried Hazod. Stetige Faltungshalbgruppen von Wahrscheinlichkeitsmaßen und erzeugende Distributionen. (German) [Continuous convolution semigroup of probability measures and generating distributions], volume 595 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 0-387-08259-X, 3-540-08259-X (print), 3-540-37339-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 157 pp. LCCN QA3 .L28 no. 595; QA3 .L28 no. 595 QA273.43. URL http:

//link.springer.com/book/10.1007/BFb0091063;   http://
www.springerlink.com/content/978-3-540-37339-1.

Hazod:1977:SED

[Haz77f] Wilfried Hazod. Struktur der erzeugenden Distributionen. (German) []. Lecture Notes in Mathematics, 595:28–65, 1977. CODEN LNMAA2. ISBN 3-540-08259-X (print), 3-540-37339-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091065/.

Hazod:1977:SCR

[Haz77g] Wilfried Hazod. Über Sätze von Cramér, Raikoff und Bernstein. (German) []. Lecture Notes in Mathematics, 595:89–106, 1977. CODEN LNMAA2. ISBN 3-540-08259-X (print), 3-540-37339-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091067/.

Hazewinkel:1979:IDU

[Haz79a] Michiel Hazewinkel. Infinite dimensional universal formal group laws and formal a-modules. Lecture Notes in Mathematics, 732: 124–143, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066641/.

Hazod:1979:SFO

[Haz79b]  W. Hazod.  Subordination von Faltungs- und Operator-

halbgruppen. (German) []. Lecture Notes in Mathematics, 706: 144–202, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063123/.

Herdman:1979:FDE

[HB79] T. L. Herdman and J. A. Burns. Functional differential equations with discontinuous right hand side. Lecture Notes in Mathematics, 737:99–106, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064500/.

Herrmann:1975:CMS

[HBT75] W. Herrmann, L. D. Bertholf, and S. L. Thompson. Computational methods for stress wave propagation in nonlinear solid mechanics. Lecture Notes in Mathematics, 461:91–127, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074151/.

Huang:1978:GPN

[HC78] Steel T. Huang and Stamatis Cambanis. Gaussian processes: Nonlinear analysis and stochastic calculus. Lecture Notes in Mathematics, 695:165–177, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062664/.

Hecquet:1976:EGS

[Hec76] Gérard Hecquet. Existence globale des solutions de quelques problémes aux limites. (French) [Global existence of solutions of some limit problems]. Lecture Notes in Mathematics, 564:239– 248, 1976. CODEN LNMAA2. ISBN 3-540-08058-9 (print), 3-540-37517-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087340/.

Hector:1977:FCF

[Hec77] Gilbert Hector. Feuilletages en cylindres. (French) []. Lecture Notes in Mathematics, 597:252–270, 1977. CODEN LN-MAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085358/.

Hector:1978:CFP

[Hec78] Gilbert Hector. Croissance des feuilletages presque sans holonômte. (French) []. Lecture Notes in Mathematics, 652:141– 182, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063511/.

Heersink:1976:LMD

[Hee76] Rudolf Heersink. Lösungsdarstellungen mittels Differentialoperatoren für das Dirichlet-Problem der Gleichung Δu + c(x, y)u = 0. (German) []. Lecture Notes in Mathematics, 561:227–238, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087639/.

Hegenbarth:1978:AI

[Heg78] Friedrich Hegenbarth. An application of the ρ-invariant. Lecture Notes in Mathematics, 652:212–221, 1978. CODEN LN-MAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063514/.

Hegenbarth:1979:ECC

[Heg79] Friedrich Hegenbarth. Exotic characteristic classes of spherical fibrations. Lecture Notes in Mathematics, 763:120–131, 1979. CO-DEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0088081/.

Heindl:1975:EKO

[Hei75] Gerhard Heindl. Ein Existenzsatz für Konvexe Optimierungsprobleme. (German) []. Lecture Notes in Mathematics, 477:145– 149, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079173/.

Heindl:1976:PBG

[Hei76a] Gerhard Heindl. Ein Problem der Bestapproximation in geordneten Vektorräumen. (German) []. Lecture Notes in Mathematics, 556:213–221, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087409/.

Heinrich:1976:NIP

[Hei76b] Katherine Heinrich. A non-imbeddable proper colouring. Lecture Notes in Mathematics, 560:93–115, 1976. CODEN LN-MAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097371/.

Heinrich:1977:LSC

[Hei77] Katherine Heinrich. Latin squares composed of four disjoint subsquares. Lecture Notes in Mathematics, 622:118–127, 1977. CO-DEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069185/.

Heinkel:1978:QRR

[Hei78] Bernard Heinkel. Quelques remarques relatives au théorème central-limite dans C(S). (French) []. Lecture Notes in Mathematics, 644:204–211, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066848/.

Heinkel:1979:RBC

[Hei79] B. Heinkel. Relation between central-limit theorem and law of the iterated logarithm in Banach spaces. Lecture Notes in Mathematics, 709:145–150, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071955/.

Hejhal:1976:STFb

[Hej76a] Dennis A. Hejhal. The Selberg Trace Formula for PSL(2, R) Volume I, volume 548 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07988-2 (print), 3-540-37979-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 499 (est.) pp. LCCN QA3 .L28 no. 548. URL http:

//link.springer.com/book/10.1007/BFb0079608;   http://
www.springerlink.com/content/978-3-540-37979-9.

Hejhal:1976:STFa

[Hej76b] Dennis A. Hejhal. The Selberg trace formula for modular correspondences. Lecture Notes in Mathematics, 548:462–499, 1976. CODEN LNMAA2. ISBN 3-540-07988-2 (print), 3-540-37979-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079613/.

Hejhal:1976:SZF

[Hej76c] Dennis A. Hejhal. The Selberg zeta function. Lecture Notes in Mathematics, 548:39–325, 1976. CODEN LNMAA2. ISBN 3-540-07988-2 (print), 3-540-37979-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079610/.

Hejhal:1976:TFA

[Hej76d] Dennis A. Hejhal. The trace formula for automorphic forms of weight m. Lecture Notes in Mathematics, 548:355–461, 1976. CO-DEN LNMAA2. ISBN 3-540-07988-2 (print), 3-540-37979-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079612/.

Hejhal:1976:TFC

[Hej76e] Dennis A. Hejhal. The trace formula for compact Riemann surfaces. Lecture Notes in Mathematics, 548:1–38, 1976. CODEN LNMAA2. ISBN 3-540-07988-2 (print), 3-540-37979-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079609/.

Hejhal:1976:TFV

[Hej76f] Dennis A. Hejhal. The trace formula for vector-valued functions. Lecture Notes in Mathematics, 548:326–354, 1976. CODEN LN-MAA2. ISBN 3-540-07988-2 (print), 3-540-37979-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079611/.

Helffer:1978:QEO

[Hel78a] B. Helffer. Quelques exemples d’opérateurs pseudodifférentiels localement résolubles. (French) []. Lecture Notes in Mathematics, 660:88–116, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082297/.

Helm:1978:GMC

[Hel78b] Werner E. Helm. On Gaussian measures and the central limit theorem in certain f-spaces. Lecture Notes in Mathematics, 656: 59–65, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068811/.

Heldermann:1979:CNS

[Hel79a] N. C. Heldermann. Concentrated nearness spaces. Lecture Notes in Mathematics, 719:122–136, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065265/.

Helmberg:1979:SRI

[Hel79b] Gilbert Helmberg. Some remarks on ϵ-independence of partitions and on topological Rochlin sets. Lecture Notes in Mathematics, 729:58–65, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063283/.

Hempel:1975:OSI

[Hem75] John Hempel. One sided incompressible surfaces in 3-manifolds. Lecture Notes in Mathematics, 438:251–258, 1975. CODEN LN-MAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066121/.

Henaut:1975:EDV

[Hen75a] Alain Henaut. Errata du volume 409. (French) []. Lecture Notes in Mathematics, 482:367, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080257.

Henaut:1975:PRA

[Hen75b] Alain Henaut. Platitude des revêtements analytiques ramifiés à deux feuillets d’une variété analytique complexe de dimension finie. (French) []. Lecture Notes in Mathematics, 482:358– 366, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080256/.

Henaut:1975:PME

[Hen75c] Alain Henaut. Prolongement d’un morphisme d’espaces analytiques complexes. (French) []. Lecture Notes in Mathematics, 482: 274–281, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080251/.

Hennekemper:1979:SRF

[Hen79a] Wilhelm Hennekemper. Some results on functions of bounded index. Lecture Notes in Mathematics, 747:158–160, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063970/.

Hennion:1979:MAE

[Hen79b] H. Hennion. Marches aléatoires sur les espaces homogènes. (French) []. Lecture Notes in Mathematics, 706:203–207, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063124/.

Herman:1975:ALC

[Her75a] M. R. Herman. Sur l’algèbre de Lie des champs de vecteurs R-analytiques du tore. (French) []. Lecture Notes in Mathematics, 484:43–49, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082144/.

Herman:1975:GDA

[Her75b] M. R. Herman. Sur le groupe des diffeomorphismes R-analytiques du tore. (French) []. Lecture Notes in Mathematics, 484:36– 42, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082143/.

Hersh:1975:MT

[Her75c] Reuben Hersh. The method of transmutations. Lecture Notes in Mathematics, 446:264–282, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070606/.

Hersh:1975:SSH

[Her75d] Reuben Hersh. Stochastic solutions of hyperbolic equations. Lecture Notes in Mathematics, 446:283–300, 1975. CODEN LN-MAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070607/.

Herrlich:1976:STT

[Her76] Horst Herrlich. Some topological theorems which fail to be true. Lecture Notes in Mathematics, 540:265–285, 1976. CODEN LN-MAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080864/.

Herman:1977:GVI

[Her77a] Michael-Robert Herman. The Godbillon–Vey invariant of foliations by planes of T^{3}. Lecture Notes in Mathematics, 597:294– 307, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085360/.

Herman:1977:MLN

[Her77b] Michael-Robert Herman. Mesure de Lebesgue et nombre de rotation. (French) []. Lecture Notes in Mathematics, 597:271– 293, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085359/.

Herrera:1977:CTS

[Her77c] M. Herrera. On the construction of the trace in Serre duality. Lecture Notes in Mathematics, 597:308–318, 1977. CODEN LN-MAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085361/.

Herrero:1978:SCO

[Her78a] D. A. Herrero. Strictly cyclic operator algebras and approximation of operators. Lecture Notes in Mathematics, 693:103–107, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064666/.

Hervier:1978:DPV

[Her78b] Y. Hervier. Deformations à un paramétre de variétés simples. (French) []. Lecture Notes in Mathematics, 683:42–61, 1978. CO-DEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063173/.

Herrlich:1979:IFC

[Her79a] Horst Herrlich. Initial and final completions. Lecture Notes in Mathematics, 719:137–149, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065266/.

Herzog:1979:DCG

[Her79b] Jürgen Herzog. Deformation of certain Gorenstein singularities. Lecture Notes in Mathematics, 740:230–236, 1979. CODEN LN-MAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071064/.

Hess:1978:PNL

[Hes78] Peter Hess. Perturbations non linéaires de problémes linéaires à la resonance: existence de multiples solutions. (French) []. Lecture Notes in Mathematics, 665:108–120, 1978. CODEN LN-MAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061801/.

Hess:1979:MSS

[Hes79] P. Hess. Multiple solutions of some asymptotically linear elliptic boundary value problems. Lecture Notes in Mathematics, 703: 145–151, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067267/.

Hettich:1976:NMN

[Het76] R. Hettich. A Newton-method for nonlinear Chebyshev approximation. Lecture Notes in Mathematics, 556:222–236, 1976. CO-DEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087410/.

Hetzer:1979:DCT

[Het79] Georg Hetzer. A degree continuation theorem for a class of compactly perturbed differentiable Fredholm maps of index 0. Lecture Notes in Mathematics, 730:194–203, 1979. CODEN LN-MAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064318/.

Heyer:1978:SGC

[Hey78] H. Heyer. Semi-groupes de convolution sur un groupe localement compact et applications à la théorie des probabilités. (French) []. Lecture Notes in Mathematics, 678:173–236, 1978. CODEN LN-MAA2. ISBN 3-540-08938-1 (print), 3-540-35716-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062798/.

Heyer:1979:PMG

[Hey79] Herbert Heyer, editor. Probability Measures on Groups: Proceedings of the Fifth Conference Oberwolfach, Germany, January 29th–February 4, 1978, volume 706 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 0-387-09124-6, 3-540-09124-6 (print), 3-540-35406-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 706; QA3 .L28 no. 706; QA273.A1. URL http:

//link.springer.com/book/10.1007/BFb0063108;   http://
www.springerlink.com/content/978-3-540-35406-2.

Hirschman:1977:EEV

[HH77a] I. I. Hirschman, Jr. and Daniel E. Hughes. Extreme Eigen Values of Toeplitz Operators, volume 618 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-07147-4 (print), 3-540-37439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 138 (est.) pp. LCCN QA3

.L28 no. 618. URL http://link.springer.com/book/10.1007/
BFb0068398;   http://www.springerlink.com/content/978-
3-540-37439-8.

Hirschman:1977:FSF

[HH77b] I. I. Hirschman, Jr. and Daniel E. Hughes. The Fourier series and Fourier transform theorems. Lecture Notes in Mathematics, 618:105–138, 1977. CODEN LNMAA2. ISBN 3-540-07147-4 (print), 3-540-37439-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068404/.

Hirschman:1977:FST

[HH77c] I. I. Hirschman, Jr. and Daniel E. Hughes. The Fourier series theorem. Lecture Notes in Mathematics, 618:74–96, 1977. CODEN LNMAA2. ISBN 3-540-07147-4 (print), 3-540-37439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068402/.

Hirschman:1977:FTT

[HH77d] I. I. Hirschman, Jr. and Daniel E. Hughes. The Fourier transform theorem. Lecture Notes in Mathematics, 618:31–73, 1977. CODEN LNMAA2. ISBN 3-540-07147-4 (print), 3-540-37439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068401/.

Hirschman:1977:HSB

[HH77e] I. I. Hirschman, Jr. and Daniel E. Hughes. Hilbert space background — small eigenvalues. Lecture Notes in Mathematics, 618: 16–30, 1977. CODEN LNMAA2. ISBN 3-540-07147-4 (print), 3-540-37439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068400/.

Hirschman:1977:HST

[HH77f] I. I. Hirschman, Jr. and Daniel E. Hughes. Hilbert space theory — large eigenvalues. Lecture Notes in Mathematics, 618:97– 104, 1977. CODEN LNMAA2. ISBN 3-540-07147-4 (print), 3-540-37439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068403/.

Hirschman:1977:I

[HH77g] I. I. Hirschman, Jr. and Daniel E. Hughes. Introduction. Lecture Notes in Mathematics, 618:1–15, 1977. CODEN LNMAA2. ISBN 3-540-07147-4 (print), 3-540-37439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068399/.

Holt:1977:LWS

[HH77h] Maurice Holt and Mohammed Y. Hussaini. Laminar wakes in supersonic flow. Lecture Notes in Mathematics, 594:260–274, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086091/.

Hadwin:1978:WTW

[HH78a] Donald W. Hadwin and Thomas B. Hoover. Weighted translation and weighted shift operators. Lecture Notes in Mathematics, 693:93–99, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064664/.

Hervier:1978:VAC

[HH78b] Y. Hervier and A. Hirschowitz, editors. Variétés Analytiques Compactes: Colloque, Nice, 19–23 Septembre 1977, volume 683 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 683. URL http://link.springer.com/book/10.1007/
BFb0063169;   http://www.springerlink.com/content/978-
3-540-35710-0.

Hida:1978:WNL

[Hid78] Takeyuki Hida. White noise and Lévy’s functional analysis. Lecture Notes in Mathematics, 695:155–163, 1978. CODEN LN-MAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062663/.

Higgins:1975:CGS

[Hig75] Theodore Parker Higgins. A child’s garden of special functions. Lecture Notes in Mathematics, 457:216–225, 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067106/.

Hilton:1975:NAN

[Hil75] Peter Hilton. Nilpotent actions on nilpotent groups. Lecture Notes in Mathematics, 450:174–196, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062856/.

Hilton:1976:SST

[Hil76] Peter Hilton, editor. Structural Stability, the Theory of Catastrophes, and Applications in the Sciences: Proceedings of the Conference Held at Battelle Seattle Research Center 1975, volume 525 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 525. URL http://link.springer.com/book/10.1007/
BFb0077839;   http://www.springerlink.com/content/978-
3-540-38254-6.

Hill:1977:CFG

[Hil77a] Paul Hill. Criteria for freeness in groups and valuated vector spaces. Lecture Notes in Mathematics, 616:140–157, 1977. CO-DEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0068193/.

Hilton:1977:LTG

[Hil77b] Peter Hilton. Localization theories for groups and homotopy types. Lecture Notes in Mathematics, 597:319–329, 1977. CODEN LN-MAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085362/.

Hilton:1978:OSG

[Hil78] Peter Hilton. On orbit sets for group actions and localization. Lecture Notes in Mathematics, 673:185–201, 1978. CODEN LN-MAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064696/.

Hinman:1977:SFT

[Hin77] Peter G. Hinman. A survey of finite-type recursion. Lecture Notes in Mathematics, 619:187–209, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067652/.

Hindman:1979:UCN

[Hin79] Neil Hindman. Ultrafilters and combinatorial number theory. Lecture Notes in Mathematics, 751:119–184, 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062706/.

Hirsch:1975:FBF

[Hir75a] M. Hirsch. Foliated bundles, flat manifolds and invariant measures. Lecture Notes in Mathematics, 468:8–9, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082577.

Hirsch:1975:SAF

[Hir75b] M. Hirsch. A stable analytic foliation with only exceptional minimal sets. Lecture Notes in Mathematics, 468:9–10, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082578.

Hirschfeld:1975:FFG

[Hir75c] J. Hirschfeld. Finite forcing and generic filters in arithmetic. Lecture Notes in Mathematics, 498:172–199, 1975. CODEN LN-MAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080980/.

Hirota:1976:DMF

[Hir76a] Ryogo Hirota. Direct method of finding exact solutions of nonlinear evolution equations. Lecture Notes in Mathematics, 515: 40–68, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081162/.

Hirsch:1976:EPP

[Hir76b] Francis Hirsch. Extension des propriétés des puissances fractionnaires. (French) []. Lecture Notes in Mathematics, 563:100– 120, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087573/.

Hirzebruch:1977:RHM

[Hir77] F. Hirzebruch. The ring of Hilbert modular forms for real quadratic fields of small discriminant. Lecture Notes in Mathematics, 627:287–323, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065306/.

Hirsch:1978:OCD

[Hir78a] Francis Hirsch. Opérateurs carré du champ [d’après J. P. Roth]. (French) []. Lecture Notes in Mathematics, 677:167–182, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070761/.

Hirsch:1978:FMC

[Hir78b] Morris W. Hirsch. Flat manifolds and the cohomology of groups. Lecture Notes in Mathematics, 664:94–103, 1978. CODEN LN-MAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061694/.

Hirai:1979:CFD

[Hir79a] Takeshi Hirai. Character formulas for the discrete series for semisimple Lie groups. Lecture Notes in Mathematics, 739:315– 340, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062499/.

Hirsch:1979:PDM

[Hir79b] Francis Hirsch. Principes du maximum pour les noyaux de convolution. (French) []. Lecture Notes in Mathematics, 713:113– 136, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071323/.

Hitsuda:1978:EPG

[Hit78] Masuyuki Hitsuda. Equivalence problem on Gaussian N-ple Markov processes with multiplicity N. Lecture Notes in Mathematics, 695:139–143, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062661/.

Hoffmann-Jorgensen:1977:PBS

[HJ77] J. Hoffmann-Jørgensen. Probability in Banach space. Lecture Notes in Mathematics, 598:1–186, 1977. CODEN LNMAA2. ISBN 3-540-08340-5 (print), 3-540-37307-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097492/.

Hoffmann-Jorgensen:1977:EEP

[HJLN77] J. (Jørgen) Hoffmann-Jørgensen, Thomas M. (Thomas Milton) Liggett, and J. (Jacques) Neveu, editors. Ecole d’été de probabilités de Saint-Flour VI — 1976. (French) [Saint-Flour Summer School on Probability VI — 1976], volume 598 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 0-387-08340-5, 3-540-08340-5 (print), 3-540-37307-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 447 pp. LCCN QA3 .L28 no. 598; QA3 .L28 no. 598 QA273.A1. URL http:

//link.springer.com/book/10.1007/BFb0097491;   http://
www.springerlink.com/content/978-3-540-37307-0.

Hoffmann:1975:ASA

[HJSW75] K.-H. Hoffmann, E. Jörn, E. Schäfer, and H. Weber. Ein Approximationssatz und seine Anwendung in der Kontrolltheorie.

(German) []. Lecture Notes in Mathematics, 477:150–157, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079174/.

Hoffmann:1976:ALD

[HK76] K.-H. Hoffmann and A. Klostermair. Approximationen mit Lösungen von Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 556:237–273, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087411/.

Husain:1978:BSB

[HK78a] T. Husain and S. M. Khaleelulla. σ-barrelled and sequentially barrelled spaces. Lecture Notes in Mathematics, 692:159–191, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063935/.

Husain:1978:BTO

[HK78b] T. Husain and S. M. Khaleelulla. Barrelledness in Topological and Ordered Vector Spaces, volume 692 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 258 pp. LCCN QA3

.L28 no. 692. URL http://link.springer.com/book/10.1007/
BFb0063929;   http://www.springerlink.com/content/978-
3-540-35559-5.

Husain:1978:CBC

[HK78c] T. Husain and S. M. Khaleelulla. Countably barrelled and countably quasibarrelled spaces. Lecture Notes in Mathematics, 692: 130–158, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063934/.

Husain:1978:COQ

[HK78d] T. Husain and S. M. Khaleelulla. Countably order-quasiultrabarrelled vector lattices and spaces. Lecture Notes in Mathematics, 692: 207–227, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063937/.

Husain:1978:CUC

[HK78e] T. Husain and S. M. Khaleelulla. Countably ultrabarrelled and countably quasiultrabarrelled spaces. Lecture Notes in Mathematics, 692:192–206, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063936/.

Husain:1978:OQVa

[HK78f] T. Husain and S. M. Khaleelulla. Order-quasibarrelled vector lattices. Lecture Notes in Mathematics, 692:103–129, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063933/.

Husain:1978:OQVb

[HK78g] T. Husain and S. M. Khaleelulla. Order-quasiultrabarrelled vector lattices and spaces. Lecture Notes in Mathematics, 692:228– 244, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063938/.

Husain:1978:P

[HK78h] T. Husain and S. M. Khaleelulla. Preliminaries. Lecture Notes in Mathematics, 692:1–41, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063930/.

Husain:1978:SCL

[HK78i] T. Husain and S. M. Khaleelulla. Some classes of locally convex spaces. Lecture Notes in Mathematics, 692:42–82, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063931/.

Husain:1978:UUQ

[HK78j] T. Husain and S. M. Khaleelulla. Ultrabarrelled, ultrabornological and quasiultrabarrelled spaces. Lecture Notes in Mathematics, 692:83–102, 1978. CODEN LNMAA2. ISBN 3-540-09096-7 (print), 3-540-35559-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063932/.

Heath:1979:GCS

[HK79a] Philip R. Heath and Klaus Heiner Kamps. Groupoids and classification sequences. Lecture Notes in Mathematics, 719:112– 121, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065264/.

Hofmann:1979:STC

[HK79b] Karl Heinrich Hofmann and Klaus Keimel. Sheaf theoretical concepts in analysis: Bundles and sheaves of Banach spaces, Banach C(X)-modules. Lecture Notes in Mathematics, 753:415–441, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061826/.

Halperin:1975:CCC

[HL75] Stephan Halperin and Daniel Lehmann. Cohomologies et classes caractéristiques des choux de Bruxelles. (French) []. Lecture Notes in Mathematics, 484:79–120, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082147/.

Herzog:1977:NCC

[HL77] Marcel Herzog and Gustav I. Lehrer. A note concerning Coxeter groups and permutations. Lecture Notes in Mathematics, 573: 53–56, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087811/.

Hardin:1978:GDT

[HL78] T. Hardin and A. Laville. Une généralisation du théorème de propagation des singularités pour les opérateurs à symbole principal réel. (French) []. Lecture Notes in Mathematics, 660:81–87, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082296/.

Handelman:1979:LTR

[HL79a] David Handelman and John Lawrence. Lower K-theory, regular rings and operator algebras — a survey. Lecture Notes in Mathematics, 734:158–173, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103158/.

Handelman:1979:RTW

[HL79b] David Handelman and John Lawrence, editors. Ring Theory Waterloo 1978 Proceedings, University of Waterloo, Canada, 12–16 June, 1978, volume 734 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 734. URL http:

//link.springer.com/book/10.1007/BFb0103151;   http://
www.springerlink.com/content/978-3-540-35043-9.

Hlavacek:1979:DFE

[Hla79] I. Hlaváček. Dual finite element analysis for some unilateral boundary value problems. Lecture Notes in Mathematics, 703: 152–159, 1979. CODEN LNMAA2. ISBN 3-540-09116-5 (print), 3-540-35519-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067268/.

Horadam:1979:DPS

[HLS79] A. F. Horadam, R. P. Loh, and A. G. Shannon. Divisibility properties of some Fibonacci-type sequences. Lecture Notes in Mathematics, 748:55–64, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102684/.

Hirsch:1976:STDa

[HM76a] Francis Hirsch and Gabriel Mokobodzki, editors. Séminaire de Théorie du Potentiel Paris 1972–1974, volume 518 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 518. URL http://link.springer.com/book/10.1007/
BFb0080397;   http://www.springerlink.com/content/978-
3-540-38225-6.

Hirsch:1976:STDb

[HM76b] Francis Hirsch and Gabriel Mokobodzki. Séminaire de Théorie du Potentiel Paris, No. 2, volume 563 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 292 pp. LCCN QA3

.L28 no. 563. URL http://link.springer.com/book/10.1007/
BFb0087568;   http://www.springerlink.com/content/978-
3-540-37526-5.

Hoppensteadt:1977:AMT

[HM77] Frank C. Hoppensteadt and Willard L. Miranker. Applications of multi-time methods to pattern recognition and other problems. Lecture Notes in Mathematics, 594:275–287, 1977. CODEN LN-MAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086092/.

Hirsch:1978:STD

[HM78a] Francis Hirsch and Gabriel Mokobodzki. Séminaire de Théorie du Potentiel Paris, No. 3, volume 681 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 294 pp. LCCN QA3

.L28 no. 681. URL http://link.springer.com/book/10.1007/
BFb0065863;   http://www.springerlink.com/content/978-
3-540-35712-4.

Holmes:1978:BDS

[HM78b] Philip J. Holmes and Jerrold E. Marsden. Bifurcations of dynamical systems and nonlinear oscillations in engineering systems. Lecture Notes in Mathematics, 648:163–206, 1978. CODEN LN-MAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066411/.

Hajek:1979:NST

[HM79a] Darrell W. Hajek and Adam Mysior. On non-simplicity of topological categories. Lecture Notes in Mathematics, 719:84–93, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065261/.

Hambleton:1979:SGF

[HM79b] Ian Hambleton and R. James Milgram. The surgery group L_{3}^{h}(Z(G)) for G a finite 2-group. Lecture Notes in Mathematics, 741:73–89, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062135/.

Hirsch:1979:STD

[HM79c] Francis Hirsch and Gabriel Mokobodzki. Séminaire de Théorie du Potentiel Paris, No. 4, volume 713 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 281 pp. LCCN QA3

.L28 no. 713. URL http://link.springer.com/book/10.1007/
BFb0071316;   http://www.springerlink.com/content/978-
3-540-35308-9.

Hilton:1978:FMF

[HMRS78] Peter Hilton, Guido Mislin, Joseph Roitberg, and Richard Steiner. On free maps and free homotopies into nilpotent spaces. Lecture Notes in Mathematics, 673:202–218, 1978. CODEN LN-MAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064697/.

Hamachi:1979:FHN

[HO79] Toshihiro Hamachi and Motosige Osikawa. Fundamental homomorphism of normalizer group of ergodic transformation. Lecture Notes in Mathematics, 729:43–57, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063282/.

Hochster:1979:BSC

[Hoc79a] Melvin Hochster. Big and small Cohen–Macaulay modules. Lecture Notes in Mathematics, 700:119–142, 1979. CODEN LN-MAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063464/.

Hochster:1979:CNL

[Hoc79b] Melvin Hochster. Commutative Noetherian local rings. Lecture Notes in Mathematics, 700:224–225, 1979. CODEN LN-MAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063470.

Hochster:1979:PIT

[Hoc79c] Melvin Hochster. Principal ideal theorems. Lecture Notes in Mathematics, 734:174–206, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103159/.

Hodgkin:1975:EKT

[Hod75] Luke Hodgkin. The equivariant Künneth theorem in K-theorem. Lecture Notes in Mathematics, 496:1–101, 1975. CODEN LN-MAA2. ISBN 3-540-07536-4 (print), 3-540-38026-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082285/.

Hoffmann:1976:TFA

[Hof76a] Rudolf-E. Hoffmann. Topological functors admitting generalized Cauchy-completions. Lecture Notes in Mathematics, 540:286– 344, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080865/.

Hofmann:1976:CTM

[Hof76b] Karl Heinrich Hofmann. Category theoretical methods in topological algebra. Lecture Notes in Mathematics, 540:345–403, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080866/.

Hofmann:1977:BSE

[Hof77] Karl Heinrich Hofmann. Bundles and sheaves are equivalent in the category of Banach spaces. Lecture Notes in Mathematics, 575: 53–69, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095701/.

Hofbauer:1979:MMP

[Hof79a] Franz Hofbauer. Maximal measures for piecewise monotonically increasing transformations on [0, 1]. Lecture Notes in Mathematics, 729:66–77, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063284/.

Hoffman:1979:RAO

[Hof79b] Peter Hoffman. λ-rings, Adams operations and plethysms. Lecture Notes in Mathematics, 746:58–71, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067683/.

Hoffman:1979:R

[Hof79c] Peter Hoffman. τ-rings. Lecture Notes in Mathematics, 746:19– 26, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067680/.

Hoffman:1979:RWP

[Hof79d] Peter Hoffman. τ-Rings and Wreath Product Representations, volume 746 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 148 pp. LCCN QA3 .L28 no. 746. URL http:

//link.springer.com/book/10.1007/BFb0067677;   http://
www.springerlink.com/content/978-3-540-34860-3.

Hoffman:1979:CRa

[Hof79e] Peter Hoffman. Computations related to R(Σ_{k}). Lecture Notes in Mathematics, 746:34–57, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067682/.

Hoffman:1979:CRb

[Hof79f] Peter Hoffman. Computations related to R(Σ_{k}⟨Γ⟩). Lecture Notes in Mathematics, 746:72–99, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067684/.

Hoffman:1979:FR

[Hof79g] Peter Hoffman. Filtered τ-rings. Lecture Notes in Mathematics, 746:100–109, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067685/.

Hoffman:1979:OFR

[Hof79h] Peter Hoffman. Operations and the free τ-ring S. Lecture Notes in Mathematics, 746:27–33, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067681/.

Hoffman:1979:PA

[Hof79i] Peter Hoffman. Problems and analogues. Lecture Notes in Mathematics, 746:110–117, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067686/.

Hoffman:1979:TPC

[Hof79j] Peter Hoffman. Tensor power constructions. Lecture Notes in Mathematics, 746:10–18, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067679/.

Hoffman:1979:WPR

[Hof79k] Peter Hoffman. Wreath products and representations. Lecture Notes in Mathematics, 746:1–9, 1979. CODEN LNMAA2. ISBN 3-540-09551-9 (print), 3-540-34860-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067678/.

Hoffmann:1979:TSA

[Hof79l] Rudolf-E. Hoffmann. Topological spaces admitting a ‘dual’. Lecture Notes in Mathematics, 719:157–166, 1979. CODEN LN-MAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065268/.

Hoffstein:1979:SRR

[Hof79m] Jeffrey Hoffstein. Some results related to minimal discriminants. Lecture Notes in Mathematics, 751:185–194, 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062707/.

Hogarth:1975:DCG

[Hog75] Pauline Cain Hogarth. Decomposition of complete graphs into 6-stars and into 10-stars. Lecture Notes in Mathematics, 452:136– 142, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069552/.

Holevo:1976:NAC

[Hol76] A. S. Holevo. Noncommutative analogues of the Cramér-Rao inequality in the quantum measurement theory. Lecture Notes in Mathematics, 550:194–222, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077491/.

Holmann:1978:SHF

[Hol78a] Harald Holmann. On the stability of holomorphic foliations with all leaves compact. Lecture Notes in Mathematics, 683:217– 248, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063180/.

Holme:1978:DSS

[Hol78b] Audun Holme. Deformation and stratification of secant structure. Lecture Notes in Mathematics, 687:60–91, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062928/.

Holme:1979:DSV

[Hol79] Audun Holme. On the dual of a smooth variety. Lecture Notes in Mathematics, 732:144–156, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066642/.

Hormander:1975:NUC

[Hör75] Lars Hörmander. Non-uniqueness for the Cauchy problem. Lecture Notes in Mathematics, 459:36–72, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074192/.

Horowitz:1977:RBF

[Hor77] Joseph Horowitz. Une remarque sur les bimesures. (French) []. Lecture Notes in Mathematics, 581:59–64, 1977. CODEN LN-MAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087188/.

Houzel:1977:MFR

[Hou77] C. Houzel. Morphisme de Frobenius et rationalité des fonctions L. (French) []. Lecture Notes in Mathematics, 589:442–480, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096812/.

Howie:1977:DGE

[How77] Monsieur J. M. Howie. Sur les demi-groupes engendrés par des idempotents. (French) []. Lecture Notes in Mathematics, 586:110– 115, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087127/.

Howlett:1978:ECN

[How78] Robert B. Howlett. Extending characters from normal subgroups. Lecture Notes in Mathematics, 697:1–7, 1978. CODEN LN-MAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103119/.

Henon:1976:TSA

[HP76a] M. Hénon and Y. Pomeau. Two strange attractors with a simple structure. Lecture Notes in Mathematics, 565:29–68, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091446/.

Herzog:1976:MDP

[HP76b] Marcel Herzog and Cheryl E. Praeger. Minimal degree of primitive permutation groups. Lecture Notes in Mathematics, 560:116– 122, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097372/.

Haussmann:1977:MAC

[HP77] Werner Haußmann and Peter Pottinger. On multivariate approximation by continuous linear operators. Lecture Notes in Mathematics, 571:101–108, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086567/.

Herrlich:1979:CTP

[HP79] Horst Herrlich and Gerhard Preuß, editors. Categorical Topology: Proceedings of the International Conference, Berlin, August 27th to September 2nd, 1978, volume 719 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 719. URL http:

//link.springer.com/book/10.1007/BFb0065251;   http://
www.springerlink.com/content/978-3-540-35193-1.

Hirsch:1977:STL

[HPS77a] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. The C^{r} section theorem and Lipschitz jets. Lecture Notes in Mathematics, 583:25–38, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092045/.

Hirsch:1977:BL

[HPS77b] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Branched laminations. Lecture Notes in Mathematics, 583:110–114, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092051/.

Hirsch:1977:CM

[HPS77c] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Center manifolds. Lecture Notes in Mathematics, 583:64–67, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092048/.

Hirsch:1977:EFN

[HPS77d] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Equivariant fibrations and nonwandering sets. Lecture Notes in Mathematics, 583:136–144, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092054/.

Hirsch:1977:FS

[HPS77e] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Forced smoothness of i: V → M. Lecture Notes in Mathematics, 583: 108–110, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092050/.

Hirsch:1977:I

[HPS77f] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Introduction. Lecture Notes in Mathematics, 583:1–5, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092043/.

Hirsch:1977:IM

[HPS77g] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Invariant Manifolds, volume 583 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 144 (est.) pp. LCCN QA3 .L28 no. 583. URL http:

//link.springer.com/book/10.1007/BFb0092042;   http://
www.springerlink.com/content/978-3-540-37382-7.

Hirsch:1977:LTNa

[HPS77h] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. The linear theory of normal hyperbolicity. Lecture Notes in Mathematics, 583:5–24, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092044/.

Hirsch:1977:LPS

[HPS77i] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Local product structure and local stability. Lecture Notes in Mathematics, 583:132–136, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092053/.

Hirsch:1977:LTNb

[HPS77j] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. The local theory of normally hyperbolic, invariant, compact manifolds. Lecture Notes in Mathematics, 583:39–53, 1977. CODEN LN-MAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0092046/.

Hirsch:1977:NU

[HPS77k] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Noncompactness and uniformity. Lecture Notes in Mathematics, 583: 67–108, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092049/.

Hirsch:1977:NHF

[HPS77l] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Normally hyperbolic foliations and laminations. Lecture Notes in Mathematics, 583:115–132, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092052/.

Hirsch:1977:PHP

[HPS77m] Morris W. Hirsch, Charles C. Pugh, and Michael Shub. Pseudo hyperbolicity and plaque families. Lecture Notes in Mathematics, 583:53–64, 1977. CODEN LNMAA2. ISBN 3-540-08148-8 (print), 3-540-37382-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092047/.

Hoffman:1978:ATP

[HPS78] Peter Hoffman, Renzo A. Piccinini, and Denis Sjerve, editors. Algebraic Topology: Proceedings, University of British Columbia, Vancouver, August 1977, volume 673 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 673. URL http:

//link.springer.com/book/10.1007/BFb0064685;   http://
www.springerlink.com/content/978-3-540-35737-7.

Herzog:1976:NFC

[HR76] Marcel Herzog and K. B. Reid. Number of factors in K-cycle decompositions of permutations. Lecture Notes in Mathematics, 560:123–131, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097373/.

Holton:1978:BP

[HR78] D. A. Holton and J. A. Rickard. Brick packing. Lecture Notes in Mathematics, 686:174–183, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062530/.

Hunter:1977:WM

[HRW77] Roger Hunter, Fred Richman, and Elbert Walker. Warfield modules. Lecture Notes in Mathematics, 616:87–123, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068191/.

Harary:1978:IFI

[HRW78] Frank Harary, Robert W. Robinson, and Nicholas C. Wormald. Isomorphic factorisations III: Complete multipartite graphs. Lecture Notes in Mathematics, 686:47–54, 1978. CODEN LN-MAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062515/.

Hatch:1975:AFD

[HS75a] Douglas E. Hatch and J. Richard Shanebrook. Application of fractional differentiation to the modeling of hodograph linearities. Lecture Notes in Mathematics, 457:289–293, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067113/.

Hodgkin:1975:TT

[HS75b] Dr. Luke H. Hodgkin and Dr. Victor P. Snaith. Topics in K-Theory, volume 496 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07536-4 (print), 3-540-38026-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 294 (est.) pp. LCCN QA3 .L28 no. 496. URL http:

//link.springer.com/book/10.1007/BFb0082284;   http://
www.springerlink.com/content/978-3-540-38026-9.

Holton:1975:SPP

[HS75c] D. A. Holton and K. C. Stacey. Some problems in permutation graphs. Lecture Notes in Mathematics, 452:143–155, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069553/.

Holton:1978:CMP

[HS78] D. A. Holton and Jennifer Seberry, editors. Combinatorial Mathematics: Proceedings of the International Conference on Combinatorial Theory Canberra, August 16–27, 1977, volume 686 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 686. URL http://link.springer.com/book/10.1007/
BFb0062510;   http://www.springerlink.com/content/978-
3-540-35702-5.

Herrlich:1979:AT

[HS79a] Horst Herrlich and George E. Strecker. Algebra ∪ topology. Lecture Notes in Mathematics, 719:150–156, 1979. CODEN LN-MAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065267/.

Hoffman:1979:ATW

[HS79b] Peter Hoffman and Victor Snaith, editors. Algebraic Topology Waterloo 1978: Proceedings of a Conference Sponsored by the Canadian Mathematical Society, NSERC (Canada), and the University of Waterloo, June 1978, volume 741 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 741. URL http:

//link.springer.com/book/10.1007/BFb0062133;   http://
www.springerlink.com/content/978-3-540-35009-5.

Holton:1977:SSL

[HSM77] D. A. Holton, K. C. Stacey, and K. L. McAvaney. The semistability of lexicographic products. Lecture Notes in Mathematics, 622:128–132, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069186/.

Henry:1975:SFJ

[HT75a] Jean-Pierre Henry and Bernard Teissier. Suffisance des familles de jets et équisingularité. (French) []. Lecture Notes in Mathematics, 482:351–357, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080255/.

Henry:1975:LET

[HT75b] Jean-Pierre G. Henry and Lê Dûng Tráng. Limités d’espaces tangents. (French) []. Lecture Notes in Mathematics, 482:251– 265, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080249/.

Hirsch:1978:RER

[HT78] F. Hirsch and J. C. Taylor. Renouvellement et existence de resolvantes. (French) []. Lecture Notes in Mathematics, 681:121– 133, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065871/.

Hansen:1979:GDS

[HT79] Friedhelm Hansen and Mark L. Teply. On the Gabriel dimension and subidealizer rings. Lecture Notes in Mathematics, 700: 95–118, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063463/.

Huard:1979:CCF

[Hua79] James G. Huard. Cyclic cubic fields that contain an integer of given index. Lecture Notes in Mathematics, 751:195–199, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062708/.

Hubbuck:1978:SGR

[Hub78a] J. R. Hubbuck. H-spaces of a given rank. Lecture Notes in Mathematics, 657:271–281, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069240/.

Hubbuck:1978:TEF

[Hub78b] J. R. Hubbuck. Two examples on finite H-spaces. Lecture Notes in Mathematics, 657:282–291, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069241/.

Hudson:1978:SMP

[Hud78] R. L. Hudson. The strong Markov property for canonical Wiener processes. Lecture Notes in Mathematics, 645:109–113, 1978. CO-DEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069667/.

Huet:1977:PSP

[Hue77] Denise Huet. Perturbations singulières de problémes elliptiques. (French) []. Lecture Notes in Mathematics, 594:288–300, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086093/.

Hughes:1978:BSB

[Hug78] Daniel Hughes. Biplanes and semi-biplanes. Lecture Notes in Mathematics, 686:55–58, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062516/.

Humphreys:1976:I

[Hum76a] James E. Humphreys. Introduction. Lecture Notes in Mathematics, 528:1–4, 1976. CODEN LNMAA2. ISBN 3-540-07796-0 (print), 3-540-38261-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079106/.

Humphreys:1976:IMR

[Hum76b] James E. Humphreys. Irreducible modular representations. Lecture Notes in Mathematics, 528:5–20, 1976. CODEN LN-MAA2. ISBN 3-540-07796-0 (print), 3-540-38261-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079107/.

Humphreys:1976:OMR

[Hum76c] James E. Humphreys. Ordinary and Modular Representations of Chevalley Groups, volume 528 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07796-0 (print), 3-540-38261-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 127 pp. LCCN QA3

.L28 no. 528. URL http://link.springer.com/book/10.1007/
BFb0079105;   http://www.springerlink.com/content/978-
3-540-38261-4.

Humphreys:1976:OR

[Hum76d] James E. Humphreys. Ordinary representations. Lecture Notes in Mathematics, 528:59–94, 1976. CODEN LNMAA2. ISBN 3-540-07796-0 (print), 3-540-38261-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079109/.

Humphreys:1976:PM

[Hum76e] James E. Humphreys. Projective modules. Lecture Notes in Mathematics, 528:21–58, 1976. CODEN LNMAA2. ISBN 3-540-07796-0 (print), 3-540-38261-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079108/.

Humphreys:1976:TG

[Hum76f] James E. Humphreys. Twisted groups. Lecture Notes in Mathematics, 528:95–99, 1976. CODEN LNMAA2. ISBN 3-540-07796-0 (print), 3-540-38261-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079110/.

Humphries:1979:GMC

[Hum79] Stephen P. Humphries. Generators for the mapping class group. Lecture Notes in Mathematics, 722:44–47, 1979. CODEN LN-MAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063188/.

Husek:1976:LRC

[Huš76] Miroslav Hušek. Lattices of reflections and coreflections in continuous structures. Lecture Notes in Mathematics, 540:404–424, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080867/.

Husek:1979:SCC

[Huš79] M. Hušek. Special classes of compact spaces. Lecture Notes in Mathematics, 719:167–175, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065269/.

Hantsch:1979:RWG

[HvW79] Lutz Hantsch and Wilhelm von Waldenfels. A random walk on the general linear group related to a problem of atomic physics. Lecture Notes in Mathematics, 706:131–143, 1979. CODEN LN-MAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063122/.

Hirschfeld:1975:ACE

[HW75a] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Applications to complete extensions of Peano’s arithmetic. Lecture Notes in Mathematics, 454:187–189, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064096/.

Hirschfeld:1975:AC

[HW75b] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Approximating chains for. Lecture Notes in Mathematics, 454:76–85, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064087/.

Hirschfeld:1975:A

[HW75c] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Axiomatizations. Lecture Notes in Mathematics, 454:111–116, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064089/.

Hirschfeld:1975:CEC

[HW75d] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Classes of existentially complete division algebras. Lecture Notes in Mathematics, 454:232–252, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064099/.

Hirschfeld:1975:ECD

[HW75e] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Existentially complete division rings. Lecture Notes in Mathematics, 454:198– 222, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064097/.

Hirschfeld:1975:ECM

[HW75f] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Existentially complete models. Lecture Notes in Mathematics, 454:141– 154, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064091/.

Hirschfeld:1975:ECS

[HW75g] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Existentially complete structures and existentially universal structures. Lecture Notes in Mathematics, 454:15–43, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064084/.

Hirschfeld:1975:FFM

[HW75h] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Finite forcing in model theory. Lecture Notes in Mathematics, 454:86–110, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064088/.

Hirschfeld:1975:FRT

[HW75i] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Forcing and recursion theory. Lecture Notes in Mathematics, 454:117–136, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064090/.

Hirschfeld:1975:FAD

[HW75j] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Forcing, Arithmetic, Division Rings, volume 454 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 266 pp. LCCN QA3

.L28 no. 454. URL http://link.springer.com/book/10.1007/
BFb0064082;   http://www.springerlink.com/content/978-
3-540-37485-5.

Hirschfeld:1975:GMA

[HW75k] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Generic models and the analytic hierarchy. Lecture Notes in Mathematics, 454: 179–186, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064095/.

Hirschfeld:1975:IFM

[HW75l] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Infinite forcing in model theory. Lecture Notes in Mathematics, 454:55–75, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064086/.

Hirschfeld:1975:I

[HW75m] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Introduction. Lecture Notes in Mathematics, 454:1–10, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064083/.

Hirschfeld:1975:MCM

[HW75n] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Modelcompletions and model-companions. Lecture Notes in Mathematics, 454:44–54, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064085/.

Hirschfeld:1975:NGS

[HW75o] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Nullstellensatz. (German) [Set of zero points]. Lecture Notes in Mathematics, 454:223–231, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064098/.

Hirschfeld:1975:RM

[HW75p] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Regular models. Lecture Notes in Mathematics, 454:160–167, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064093/.

Hirschfeld:1975:RMS

[HW75q] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Regular models and second order models for arithmetic. Lecture Notes in Mathematics, 454:168–178, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064094/.

Hirschfeld:1975:SMR

[HW75r] Dr. Joram Hirschfeld and Dr. William H. Wheeler. Simple models and R. E. ultrapowers. Lecture Notes in Mathematics, 454: 155–159, 1975. CODEN LNMAA2. ISBN 3-540-07157-1 (print), 3-540-37485-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064092/.

Hellerstein:1979:RZD

[HW79a] Simon Hellerstein and Jack Williamson. Reality of the zeros of derivatives of a meromorphic function. Lecture Notes in Mathematics, 747:153–157, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063969/.

Horadam:1979:CMV

[HW79b] A. F. Horadam and W. D. Wallis, editors. Combinatorial Mathematics VI: Proceedings of the Sixth Australian Conference on Combinatorial Mathematics, Armidale, Australia, August 1978, volume 748 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 748. URL http:

//link.springer.com/book/10.1007/BFb0102678;   http://
www.springerlink.com/content/978-3-540-34857-3.

Harary:1978:DCS

[HWH78] Frank Harary, W. D. Wallis, and Katherine Heinrich. Decompositions of complete symmetric digraphs into the four oriented quadrilaterals. Lecture Notes in Mathematics, 686:165–173, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062529/.

Hyland:1979:CST

[Hyl79] J. M. E. Hyland. Continuity in spatial toposes. Lecture Notes in Mathematics, 753:442–465, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061827/.

Iagolnitzer:1975:AME

[Iag75a] D. Iagolnitzer. Appendix microlocal essential support of a distribution and decomposition theorems — an introduction. Lecture Notes in Mathematics, 449:121–132, 1975. CODEN LN-MAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062919/.

Iagolnitzer:1975:MPR

[Iag75b] D. Iagolnitzer. Macrocausality, physical-region analyticity and independence property in S-matrix theory. Lecture Notes in Mathematics, 449:102–120, 1975. CODEN LNMAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062918/.

Iitaka:1979:SFW

[Iit79] Shigeru Iitaka. Symmetric forms and Weierstrass semigroups. Lecture Notes in Mathematics, 732:157–170, 1979. CODEN LN-MAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066643/.

Ilin:1979:IMN

[IK79] V. P. Il’in and Yu. A. Kuznetsov. Iterative methods in numerical solution of differential equations. Lecture Notes in Mathematics, 704:23–36, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/3-
540-09123-8_3/.

Ikegami:1978:RBP

[Ike78] Teruo Ikegami. On the regularity of boundary points in a resolutive compactification of a harmonic space. Lecture Notes in Mathematics, 681:134–142, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065872/.

Ikegami:1979:BBA

[Ike79] Teruo Ikegami. The boundary behavior of analytic mappings of Riemann surfaces. Lecture Notes in Mathematics, 747:161– 166, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063971/.

Illusie:1976:CCF

[Ill76] Luc Illusie. Cohomologie cristalline. (French) [Crystalline cohomology]. Lecture Notes in Mathematics, 514:53–60, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080057/.

Illusie:1977:CTL

[Ill77a] Luc Illusie. Calculs de termes locaux. (French) []. Lecture Notes in Mathematics, 589:138–203, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096805/.

Illusie:1977:SGA

[Ill77b] Luc Illusie, editor. Séminaire de Géométrie Algébrique du Bois– Marie 1965–66 SGA 5: dirigé par A. Grothendieck avec la collaboration de I. Bucur, C. Houzel, L. Illusie, J.-P. Jouanolou et J.-P. Serre, volume 589 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 589. URL http:

//link.springer.com/book/10.1007/BFb0096802;   http://
www.springerlink.com/content/978-3-540-37359-9.

Imbert:1975:MHL

[Imb75] C. Imbert. Sur la méthode en histoire de la logique. (French) [On the method of logic in history]. Lecture Notes in Mathematics, 499:363–383, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079425/.

Imrich:1977:STG

[Imr77] Wilfried Imrich. Subgroup theorems and graphs. Lecture Notes in Mathematics, 622:1–27, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069177/.

Nagata:1977:RCF

[iN77] Jun iti Nagata. On rings of continuous functions. Lecture Notes in Mathematics, 609:136–153, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068679/.

Iooss:1976:DBS

[Ioo76] Gérard Iooss. Direct bifurcation of a steady solution of the Navier– Stokes equations into an invariant torus. Lecture Notes in Mathematics, 565:69–84, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091447/.

Iberkleid:1976:AT

[IP76a] Wolf Iberkleid and Ted Petrie. The algebraic tools. Lecture Notes in Mathematics, 557:25–99, 1976. CODEN LNMAA2. ISBN 3-540-08002-3 (print), 3-540-37551-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091047/.

Iberkleid:1976:I

[IP76b] Wolf Iberkleid and Ted Petrie. Introduction. Lecture Notes in Mathematics, 557:1–23, 1976. CODEN LNMAA2. ISBN 3-540-08002-3 (print), 3-540-37551-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091046/.

Iberkleid:1976:SSA

[IP76c] Wolf Iberkleid and Ted Petrie. A setting for smooth S^{1} actions with applications to real algebraic actions on P (C^{4n}). Lecture Notes in Mathematics, 557:101–160, 1976. CODEN LN-MAA2. ISBN 3-540-08002-3 (print), 3-540-37551-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091048/.

Iberkleid:1976:SM

[IP76d] Wolf Iberkleid and Ted Petrie. Smooth S^{1} Manifolds, volume 557 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08002-3 (print), 3-540-37551-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 160 (est.) pp. LCCN QA3 .L28 no. 557. URL http:

//link.springer.com/book/10.1007/BFb0091045;   http://
www.springerlink.com/content/978-3-540-37551-7.

Irving:1979:NAN

[Irv79] Ronald S. Irving. Noetherian algebras and the Nullstellensatz. Lecture Notes in Mathematics, 740:80–87, 1979. CODEN LN-MAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071054/.

Sato:1978:DOP

[iS78] Ken iti Sato. Diffusion operators in population genetics and convergence of Markov chains. Lecture Notes in Mathematics, 695: 127–137, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062660/.

Iskovskih:1979:BTF

[IŠ79] V. A. Iskovskih and V. V. Šokurov. Biregular theory of Fano 3folds. Lecture Notes in Mathematics, 732:171–182, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066644/.

Ishida:1976:ANF

[Ish76a] Makoto Ishida. Algebraic number fields of Eisenstein type. Lecture Notes in Mathematics, 555:9–22, 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100831/.

Ishida:1976:COP

[Ish76b] Makoto Ishida. The case of odd prime degree. Lecture Notes in Mathematics, 555:54–75, 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100834/.

Ishida:1976:GF

[Ish76c] Makoto Ishida. The genus fields. Lecture Notes in Mathematics, 555:33–53, 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100833/.

Ishida:1976:GFA

[Ish76d] Makoto Ishida. The Genus Fields of Algebraic Number Fields, volume 555 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 116 pp. LCCN QA3 .L28 no. 555. URL http:

//link.springer.com/book/10.1007/BFb0100829;   http://
www.springerlink.com/content/978-3-540-37553-1.

Ishida:1976:GFC

[Ish76e] Makoto Ishida. The genus fields of cubic number fields. Lecture Notes in Mathematics, 555:76–86, 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100835/.

Ishida:1976:GFP

[Ish76f] Makoto Ishida. The genus fields of pure number fields. Lecture Notes in Mathematics, 555:87–96, 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100836/.

Ishida:1976:P

[Ish76g] Makoto Ishida. Preliminaries. Lecture Notes in Mathematics, 555: 1–8, 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100830/.

Ishida:1976:SUA

[Ish76h] Makoto Ishida. Some unramified abelian extensions. Lecture Notes in Mathematics, 555:23–32, 1976. CODEN LNMAA2. ISBN 3-540-08000-7 (print), 3-540-37553-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100832/.

Ishii:1976:THD

[Ish76i] Gorô Ishii. Test of hypotheses for distributions with monotone likelihood ratio: case of vector valued parameter. Lecture Notes in Mathematics, 550:223–231, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077492/.

Istratescu:1978:SMT

[Ist78] Vasile I. Istrăţescu. On the spectral mixing theorem for some classes of Banach spaces and for the numerical contractions on Hilbert spaces. Lecture Notes in Mathematics, 656:67–75, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068812/.

Ito:1975:SPD

[Itô75] Kiyosi Itô. Stochastic parallel displacement. Lecture Notes in Mathematics, 451:1–7, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068575/.

Ivascu:1979:KMC

[Iva79a] D. Ivaşcu. On Klein–Maskit combination theorems. Lecture Notes in Mathematics, 743:114–124, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079486/.

Ivascu:1979:RCQ

[Iva79b] D. Ivaşcu. Remarks on a class of quasisymmetric mappings. Lecture Notes in Mathematics, 743:125–131, 1979. CODEN LN-MAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079487/.

Ivascu:1979:SSD

[Iva79c] D. Ivaşcu. The Schottky space in dimensions greater than two. Lecture Notes in Mathematics, 747:167–177, 1979. CODEN LN-MAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063972/.

Iversen:1978:DIC

[Ive78] Birger Iversen. Depth inequalities for complexes. Lecture Notes in Mathematics, 687:92–111, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062929/.

Jacobinski:1975:UDL

[Jac75a] H. Jacobinski. Unique decomposition of lattices over orders. Lecture Notes in Mathematics, 488:168–176, 1975. CODEN LN-MAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081222/.

Jacobson:1975:AFD

[Jac75b] Prof. Nathan Jacobson. Applications to finite dimensional algebras. Lecture Notes in Mathematics, 441:67–114, 1975. CODEN LNMAA2. ISBN 3-540-07143-1 (print), 3-540-37427-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070024/.

Jacobson:1975:I

[Jac75c] Prof. Nathan Jacobson. Introduction. Lecture Notes in Mathematics, 441:1–10, 1975. CODEN LNMAA2. ISBN 3-540-07143-1 (print), 3-540-37427-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070022/.

Jacobson:1975:PA

[Jac75d] Prof. Nathan Jacobson. PI-algebras. Lecture Notes in Mathematics, 441:11–66, 1975. CODEN LNMAA2. ISBN 3-540-07143-1 (print), 3-540-37427-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070023/.

Jacobson:1975:PAI

[Jac75e] Prof. Nathan Jacobson. PI-Algebras: An introduction, volume 441 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07143-1 (print), 3-540-37427-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 115 pp. LCCN QA3 .L28 no. 441. URL http:

//link.springer.com/book/10.1007/BFb0070021;   http://
www.springerlink.com/content/978-3-540-37427-5.

Jacob:1977:SGL

[Jac77a] Gérard Jacob. Semi-groups linéaires de rang borné. décidabilité de la finitude. (French) []. Lecture Notes in Mathematics, 586: 116–135, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087128/.

Jacquet:1977:GR

[Jac77b] Hervé Jacquet. Generic representations. Lecture Notes in Mathematics, 587:91–101, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087916/.

Jacobson:1978:SRD

[Jac78a] N. Jacobson. Some recent developments in the theory of algebras with polynomial identities. Lecture Notes in Mathematics, 697:8–46, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103120/.

Jacod:1978:PPD

[Jac78b] Jean Jacod. Projection previsible et decomposition multiplicative d’une semi-martingale positive. (French) []. Lecture Notes in Mathematics, 649:22–34, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064589/.

Jackson:1979:SES

[Jac79a] Howard L. Jackson. Some exceptional sets in potential theory. Lecture Notes in Mathematics, 747:178–188, 1979. CODEN LN-MAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063973/.

Jacod:1979:CSP

[Jac79b] Jean Jacod. Calcul Stochastique et Problèmes de Martingales. (French) [], volume 714 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 539 pp. LCCN QA3 .L28 no. 714. URL http:

//link.springer.com/book/10.1007/BFb0064907;   http://
www.springerlink.com/content/978-3-540-35301-0.

Jacod:1979:CFF

[Jac79c] Jean Jacod. Changements de filtration. (French) []. Lecture Notes in Mathematics, 714:278–310, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064917/.

Jacod:1979:CPF

[Jac79d] Jean Jacod. Changements de probabilité. (French) []. Lecture Notes in Mathematics, 714:211–248, 1979. CODEN LN-MAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064915/.

Jacod:1979:CTC

[Jac79e] Jean Jacod. Changements de temps et changements d’Espace. (French) []. Lecture Notes in Mathematics, 714:311–336, 1979.

CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064918/.

Jacod:1979:CSF

[Jac79f] Jean Jacod. Complements sur les semimartingales. (French) []. Lecture Notes in Mathematics, 714:158–189, 1979. CODEN LN-MAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064913/.

Jacod:1979:CPA

[Jac79g] Jean Jacod. Conditions pour L’Absolue continuité. (French) [Conditions for absolute continuity]. Lecture Notes in Mathematics, 714:249–277, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064916/.

Jacod:1979:EDS

[Jac79h] Jean Jacod. Équations différentielles stochastiques et problèmes de martingales. (French) []. Lecture Notes in Mathematics, 714: 447–494, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064922/.

Jacod:1979:FED

[Jac79i] Jean Jacod. Formules exponentielles et decompositions multiplicatives. (French) []. Lecture Notes in Mathematics, 714:190– 210, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064914/.

Jacod:1979:MSI

[Jac79j] Jean Jacod. Martingales, semimartingales et intégrales stochastiques. (French) []. Lecture Notes in Mathematics, 714:26–65, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064910/.

Jacod:1979:MAI

[Jac79k] Jean Jacod. Mesures aléatoires et intégrales stochastiques. (French) []. Lecture Notes in Mathematics, 714:66–112, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064911/.

Jacod:1979:PMQ

[Jac79l] Jean Jacod. Problèmes de martingales: Quelques exemples. (French) []. Lecture Notes in Mathematics, 714:406–446, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064921/.

Jacod:1979:QND

[Jac79m] Jean Jacod. Quelques notations et definitions. (French) []. Lecture Notes in Mathematics, 714:1–7, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064908/.

Jacod:1979:RIS

[Jac79n] Jean Jacod. Représentation intégrale des solutions de problèmes de martingales. (French) []. Lecture Notes in Mathematics, 714: 495–517, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064923/.

Jacod:1979:RTG

[Jac79o] Jean Jacod. Résumé de la théorie générale des processus. (French) []. Lecture Notes in Mathematics, 714:8–25, 1979. CODEN LN-MAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064909/.

Jacod:1979:SEP

[Jac79p] Jean Jacod. Solutions extremales d’Un premier problème de martingales. (French) []. Lecture Notes in Mathematics, 714:337– 361, 1979. CODEN LNMAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064919/.

Jacod:1979:SES

[Jac79q] Jean Jacod. Sous-espaces stables de martingales. (French) []. Lecture Notes in Mathematics, 714:113–157, 1979. CODEN LN-MAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064912/.

Jacod:1979:SPM

[Jac79r] Jean Jacod. Un second problème de martingales. (French) []. Lecture Notes in Mathematics, 714:362–405, 1979. CODEN LN-MAA2. ISBN 3-540-09253-6 (print), 3-540-35301-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064920/.

Jain:1976:CLT

[Jai76] Naresh C. Jain. Central limit theorem in a Banach space. Lecture Notes in Mathematics, 526:113–130, 1976. CODEN LN-MAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082347/.

Jajte:1975:CO

[Jaj75] R. Jajte. Convergence of observables. Lecture Notes in Mathematics, 472:57–68, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081945/.

Jakobsen:1979:HOT

[Jak79] Hans Plesner Jakobsen. Higher order tensor products of wave equations. Lecture Notes in Mathematics, 728:97–115, 1979. CO-DEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063340/.

James:1978:BRT

[Jam78a] G. D. James. Background from representation theory. Lecture Notes in Mathematics, 682:1–4, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067709/.

James:1978:BT

[Jam78b] G. D. James. The branching theorem. Lecture Notes in Mathematics, 682:34–35, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067717.

James:1978:CT

[Jam78c] G. D. James. The character table of. Lecture Notes in Mathematics, 682:22–26, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067714/.

James:1978:CF

[Jam78d] G. D. James. Composition factors. Lecture Notes in Mathematics, 682:42–43, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067720.

James:1978:DF

[Jam78e] G. D. James. The determinantal form. Lecture Notes in Mathematics, 682:74–76, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067727/.

James:1978:DTT

[Jam78f] G. D. James. Diagrams, tableaux and tabloids. Lecture Notes in Mathematics, 682:8–12, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067711/.

James:1978:E

[Jam78g] G. D. James. Examples. Lecture Notes in Mathematics, 682: 18–21, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067713/.

James:1978:GR

[Jam78h] G. D. James. The garnir relations. Lecture Notes in Mathematics, 682:27–28, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067715.

James:1978:HFD

[Jam78i] G. D. James. The hook formula for dimensions. Lecture Notes in Mathematics, 682:77–78, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067728.

James:1978:HSH

[Jam78j] G. D. James. Hooks and skew-hooks. Lecture Notes in Mathematics, 682:73, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067726.

James:1978:IR

[Jam78k] G. D. James. The irreducible representations of. Lecture Notes in Mathematics, 682:39–41, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067719/.

James:1978:LRR

[Jam78l] G. D. James. The Littlewood–Richardson rule. Lecture Notes in Mathematics, 682:60–64, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067724/.

James:1978:RP

[Jam78m] G. D. James. p-regular partitions. Lecture Notes in Mathematics, 682:36–38, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067718/.

James:1978:RTS

[Jam78n] G. D. James. The Representation Theory of the Symmetric Groups, volume 682 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 156 pp. LCCN QA3 .L28 no. 682. URL http:

//link.springer.com/book/10.1007/BFb0067708;   http://
www.springerlink.com/content/978-3-540-35711-7.

James:1978:SH

[Jam78o] G. D. James. Semistandard homomorphisms. Lecture Notes in Mathematics, 682:44–50, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067721/.

James:1978:S

[Jam78p] G. D. James. Sequences. Lecture Notes in Mathematics, 682: 54–59, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067723/.

James:1978:SM

[Jam78q] G. D. James. Specht modules. Lecture Notes in Mathematics, 682:13–17, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067712/.

James:1978:SS

[Jam78r] G. D. James. A Specht series for M^{μ}. Lecture Notes in Mathematics, 682:65–72, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067725/.

James:1978:SBS

[Jam78s] G. D. James. The standard basis of the Specht module. Lecture Notes in Mathematics, 682:29–33, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067716/.

James:1978:SG

[Jam78t] G. D. James. The symmetric group. Lecture Notes in Mathematics, 682:5–7, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067710/.

James:1978:YR

[Jam78u] G. D. James. Young’s rule. Lecture Notes in Mathematics, 682: 51–53, 1978. CODEN LNMAA2. ISBN 3-540-08948-9 (print), 3-540-35711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067722/.

Janusz:1975:LIE

[Jan75] Gerald J. Janusz. The local index of elements in the Schur group. Lecture Notes in Mathematics, 488:177–183, 1975. CODEN LN-MAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081223/.

Janko:1978:FSG

[Jan78] Zvonimir Janko. On the finite simple groups [according to Aschbacher and Gorenstein]. Lecture Notes in Mathematics, 677: 183–188, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070762/.

Janssen:1979:MMZ

[Jan79a] Arnold Janssen. Über die Meßbarkeit der Mengen der zulässigen und singulären Translationen von Maßen: Der Lebesguesche Zerlegungssatz für Kerne. (German) []. Lecture Notes in Mathematics, 706:208–212, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063125/.

Jantzen:1979:AVB

[Jan79b] Jens Carsten Jantzen. Assoziierte Varietäten und Bernštein– Dimension. (German) []. Lecture Notes in Mathematics, 750:78– 111, 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069525/.

Jantzen:1979:EG

[Jan79c] Jens Carsten Jantzen. Einleitung. (German) []. Lecture Notes in Mathematics, 750:1–10, 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069522/.

Jantzen:1979:E

[Jan79d] Jens Carsten Jantzen. Errata. Lecture Notes in Mathematics, 750:182, 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069528.

Jantzen:1979:FMM

[Jan79e] Jens Carsten Jantzen. Filtrierungen der Moduln — Multiplizität Eins. (German) []. Lecture Notes in Mathematics, 750:144– 181, 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069527/.

Jantzen:1979:MEHa

[Jan79f] Jens Carsten Jantzen. Moduln mit einem höchsten Gewicht. (German) [Modules with a highest weight]. Lecture Notes in Mathematics, 750:11–41, 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069523/.

Jantzen:1979:MEHb

[Jan79g] Jens Carsten Jantzen. Moduln mit einem höchsten Gewicht. (German) [Modules with a highest weight], volume 750 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 195 pp. LCCN QA3

.L28 no. 750. URL http://link.springer.com/book/10.1007/
BFb0069521;   http://www.springerlink.com/content/978-
3-540-34854-2.

Jantzen:1979:MHG

[Jan79h] Jens Carsten Jantzen. Moduln mit höchsten Gewichten über k-Algebren. (German) []. Lecture Notes in Mathematics, 750:112– 143, 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069526/.

Jantzen:1979:TVG

[Jan79i] Jens Carsten Jantzen. Tensorprodukte und Verschiebungsprinzip. (German) []. Lecture Notes in Mathematics, 750:42–77, 1979. CODEN LNMAA2. ISBN 3-540-09558-6 (print), 3-540-34854-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069524/.

Jategaonkar:1976:CIA

[Jat76] Arun Vinayak Jategaonkar. Certain injectives are Artinian. Lecture Notes in Mathematics, 545:128–139, 1976. CODEN LN-MAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080308/.

Jeggle:1975:SNV

[Jeg75] Prof. Dr. Hansgeorg Jeggle. Zur Störungstheorie Nichtlinearer Variationsungleichungen. (German) []. Lecture Notes in Mathematics, 477:158–176, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079175/.

Jeltsch:1978:SRM

[Jel78] Rolf Jeltsch. On the stability regions of multistep multiderivative methods. Lecture Notes in Mathematics, 631:63–80, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067464/.

Jensen:1977:CRG

[Jen77a] Gary R. Jensen. Curves in real Grassmannians. Lecture Notes in Mathematics, 610:58–65, 1977. CODEN LNMAA2. ISBN 3-540-08433-9 (print), 3-540-37120-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068418/.

Jensen:1977:GT

[Jen77b] Gary R. Jensen. The general theory. Lecture Notes in Mathematics, 610:1–43, 1977. CODEN LNMAA2. ISBN 3-540-08433-9 (print), 3-540-37120-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068416/.

Jensen:1977:HOC

[Jen77c] Gary R. Jensen. Higher Order Contact of Submanifolds of Homogeneous Spaces, volume 610 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08433-9 (print), 3-540-37120-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 154 pp. LCCN QA3

.L28 no. 610. URL http://link.springer.com/book/10.1007/
BFb0068415;   http://www.springerlink.com/content/978-
3-540-37120-5.

Jensen:1977:HCCb

[Jen77d] Gary R. Jensen. Holomorphic curves in complex Grassmannians. Lecture Notes in Mathematics, 610:81–114, 1977. CODEN LNMAA2. ISBN 3-540-08433-9 (print), 3-540-37120-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068420/.

Jensen:1977:HCCa

[Jen77e] Gary R. Jensen. Holomorphic curves in complex projective space. Lecture Notes in Mathematics, 610:66–80, 1977. CODEN LN-MAA2. ISBN 3-540-08433-9 (print), 3-540-37120-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068419/.

Jensen:1977:SAS

[Jen77f] Gary R. Jensen. Special affine surface theory. Lecture Notes in Mathematics, 610:115–150, 1977. CODEN LNMAA2. ISBN 3-540-08433-9 (print), 3-540-37120-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068421/.

Jensen:1977:SUE

[Jen77g] Gary R. Jensen. Surfaces in R^{3} under the Euclidean group of proper motions. Lecture Notes in Mathematics, 610:44–58, 1977. CODEN LNMAA2. ISBN 3-540-08433-9 (print), 3-540-37120-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068417/.

Jennane:1978:EFD

[Jen78] B. Jennane. Extension d’une fonction définie sur une sous-variété avec contrôle de la croissance. (French) []. Lecture Notes in Mathematics, 694:126–133, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063246/.

Jerome:1976:EAW

[Jer76a] Joseph W. Jerome. Existence and approximation of weak solutions of the Stefan problem with nonmonotone nonlinearities. Lecture Notes in Mathematics, 506:148–156, 1976. CODEN LN-MAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080121/.

Jerome:1976:GME

[Jer76b] Joseph W. Jerome. Galerkin methods for the existence and approximation of weak solutions of nonlinear Dirichlet problems with discontinuities. Lecture Notes in Mathematics, 556:274–290, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087412/.

Jervell:1978:CU

[Jer78] Herman Ruge Jervell. Constructive universes I. Lecture Notes in Mathematics, 669:73–98, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103103/.

Jetter:1977:NNF

[JL77] Dr. K. Jetter and Prof. Dr. F. Locher. A note on numerical Fourier analysis and uniform approximation on cubes. Lecture Notes in Mathematics, 571:109–118, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086568/.

Jentzsch:1976:ASF

[JLR76] Dipl.-Geophys. Gerhald Jentzsch, Dipl.-Math. Gerald Lange, and Prof. Dr.-Ing. Otto Rosenbach. Anwendung der Spline-Funktionen zur Bearbeitung Geophysikalischer Messreihen. (German) []. Lecture Notes in Mathematics, 501:99–115, 1976. CODEN LN-MAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079743/.

Jacod:1976:TRM

[JM76] J. Jacod and J. Mémin. Un théorème de représentation des martingales pour les ensembles regeneratifs. (French) []. Lecture Notes in Mathematics, 511:24–39, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101094/.

Joubert:1975:DTG

[JMR75] Dr. Gérard P. Joubert, Dr. Robert P. Moussu, and Dr. Robert H. Roussarie, editors. Differential Topology and Geometry: Proceedings of the Colloquium Held at Dijon, 17–22 June, 1974, volume 484 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 484. URL http://link.springer.com/book/10.1007/
BFb0082139;   http://www.springerlink.com/content/978-
3-540-37919-5.

Johnson:1975:ADF

[Joh75a] Peter D. Johnson, Jr. An algebraic definition of fractional differentiation. Lecture Notes in Mathematics, 457:226–231, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067107/.

Johnstone:1975:ICC

[Joh75b] P. T. Johnstone. Internal categories and classification theorems. Lecture Notes in Mathematics, 445:103–113, 1975. CODEN LN-MAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061295/.

Johnson:1977:SBA

[Joh77] B. E. Johnson. Stability of Banach algebras. Lecture Notes in Mathematics, 609:109–114, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068676/.

Johnson:1978:TS

[Joh78] Russell A. Johnson. On a theorem of sell. Lecture Notes in Mathematics, 668:160–172, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101789/.

Johannson:1979:AHE

[Joh79a] Klaus Johannson. Attaching homotopy equivalences. Lecture Notes in Mathematics, 761:174–186, 1979. CODEN LN-MAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085415/.

Johannson:1979:CS

[Joh79b] Klaus Johannson. Characteristic submanifolds. Lecture Notes in Mathematics, 761:83–103, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085410/.

Johannson:1979:DFM

[Joh79c] Klaus Johannson. Dehn flips of 3-manifolds. Lecture Notes in Mathematics, 761:227–249, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085417/.

Johannson:1979:DTM

[Joh79d] Klaus Johannson. Dehn twists of 3-manifolds. Lecture Notes in Mathematics, 761:187–226, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085416/.

Johannson:1979:ESS

[Joh79e] Klaus Johannson. Essential singular surfaces in some special 3manifolds. Lecture Notes in Mathematics, 761:47–82, 1979. CO-DEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085409/.

Johannson:1979:GT

[Joh79f] Klaus Johannson. General theory. Lecture Notes in Mathematics, 761:19–46, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085408/.

Johannson:1979:GPM

[Joh79g] Klaus Johannson. Geometric properties of 3-manifold groups. Lecture Notes in Mathematics, 761:280–296, 1979. CODEN LN-MAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085419/.

Johannson:1979:HEM

[Joh79h] Klaus Johannson. Homotopy Equivalences of 3-Manifolds with Boundaries, volume 761 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 303 pp. LCCN QA3

.L28 no. 761. URL http://link.springer.com/book/10.1007/
BFb0085406;   http://www.springerlink.com/content/978-
3-540-38486-1.

Johannson:1979:HES

[Joh79i] Klaus Johannson. Homotopy equivalences of surfaces and Ibundles. Lecture Notes in Mathematics, 761:250–279, 1979. CO-DEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085418/.

Johannson:1979:I

[Joh79j] Klaus Johannson. Introduction. Lecture Notes in Mathematics, 761:3–18, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085407/.

Johannson:1979:ICS

[Joh79k] Klaus Johannson. Invariance of the characteristic submanifolds under homotopy equivalences. Lecture Notes in Mathematics, 761: 134–158, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085413/.

Johannson:1979:MCG

[Joh79l] Klaus Johannson. On the mapping class group of simple 3manifolds. Lecture Notes in Mathematics, 722:48–66, 1979. CO-DEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063189/.

Johannson:1979:SM

[Joh79m] Klaus Johannson. Simple 3-manifolds. Lecture Notes in Mathematics, 761:159–173, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085414/.

Johannson:1979:SSCb

[Joh79n] Klaus Johannson. Singular submanifolds and characteristic submanifolds. Lecture Notes in Mathematics, 761:120–133, 1979. CO-DEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085412/.

Johannson:1979:SSCa

[Joh79o] Klaus Johannson. Singular surfaces and characteristic submanifolds. Lecture Notes in Mathematics, 761:104–119, 1979. CODEN LNMAA2. ISBN 3-540-09714-7 (print), 3-540-38486-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085411/.

Johnson:1979:CND

[Joh79p] B. Johnson. Characterization and norms of derivations on von Neumann algebras. Lecture Notes in Mathematics, 725:228– 236, 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062619/.

Johnstone:1979:CRM

[Joh79q] P. T. Johnstone. Conditions related to de Morgan’s law. (French) []. Lecture Notes in Mathematics, 753:479–491, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061829/.

Johnstone:1979:SAD

[Joh79r] P. T. Johnstone. A syntactic approach to Diers’ localizable categories. Lecture Notes in Mathematics, 753:466–478, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061828/.

Jolivet:1978:CTD

[Jol78] E. Jolivet. Caractérisation et test du caractère agrégatif des processus ponctuels stationnaires sur R^{2}. (French) []. Lecture Notes in Mathematics, 636:1–25, 1978. CODEN LNMAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063259/.

Jones:1978:WC

[Jon78a] F. Burton Jones. Wilder on connectedness. Lecture Notes in Mathematics, 664:1–6, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061685/.

Jones:1978:TCC

[Jon78b] Lowell Jones. Two characteristic classes and Smith theory. Lecture Notes in Mathematics, 664:104–122, 1978. CODEN LN-MAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061695/.

Jones:1979:IGA

[Jon79] V. F. R. Jones. An invariant for group actions. Lecture Notes in Mathematics, 725:237–253, 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062620/.

Jordan:1979:SNS

[Jor79] G. S. Jordan. Some nonlinear singularly perturbed Volterra integrodifferential equations. Lecture Notes in Mathematics, 737: 107–119, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064501/.

Joseph:1976:FTS

[Jos76] Daniel D. Joseph. Factorization theorems for the stability of bifurcating solutions. Lecture Notes in Mathematics, 565:85–93, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091448/.

Joseph:1977:CVP

[Jos77] A. Joseph. A characteristic variety for the primitive spectrum of a semisimple Lie algebra. Lecture Notes in Mathematics, 587: 102–118, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087917/.

Joseph:1979:MSP

[Jos79] A. Joseph. W -module structure in the primitive spectrum of the enveloping algebra of a semisimple Lie algebra. Lecture Notes in Mathematics, 728:116–135, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063341/.

Jouanolou:1977:CQS

[Jou77a] J. P. Jouanolou. Cohomologie de quelques schémas classiques et théorie cohomologique des classes de Chern. (French) []. Lecture Notes in Mathematics, 589:282–350, 1977. CODEN LN-MAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096808/.

Jouanolou:1977:CAF

[Jou77b] J. P. Jouanolou. Cohomologie L-adique. (French) [L-adic cohomology]. Lecture Notes in Mathematics, 589:251–281, 1977. CODEN LNMAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096807/.

Jouanolou:1977:SPA

[Jou77c] J. P. Jouanolou. Systèmes projectifs J-adiques. (French) []. Lecture Notes in Mathematics, 589:204–250, 1977. CODEN LN-MAA2. ISBN 3-540-08248-4 (print), 3-540-37359-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096806/.

Jouanolou:1979:SRD

[Jou79a] J. P. Jouanolou. Singularités rationnelles du résultant. (French) []. Lecture Notes in Mathematics, 732:183–213, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066645/.

Jouanolou:1979:DEP

[Jou79b] Jean-Pierre Jouanolou. Densité des équations de Pfaff algébriques sans solution algébrique. (French) []. Lecture Notes in Mathematics, 708:157–197, 1979. CODEN LNMAA2. ISBN 3-540-09239-0 (print), 3-540-35344-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063397/.

Jouanolou:1979:EJF

[Jou79c] Jean-Pierre Jouanolou. Équations de Jacobi. (French) []. Lecture Notes in Mathematics, 708:1–79, 1979. CODEN LNMAA2. ISBN 3-540-09239-0 (print), 3-540-35344-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063394/.

Jouanolou:1979:EPAb

[Jou79d] Jean-Pierre Jouanolou. Équations de Pfaff algébriques. (French) [Algebraic Pfaf equations], volume 708 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09239-0 (print), 3-540-35344-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 255 pp. LCCN QA3

.L28 no. 708. URL http://link.springer.com/book/10.1007/
BFb0063393;   http://www.springerlink.com/content/978-
3-540-35344-7.

Jouanolou:1979:EPAa

[Jou79e] Jean-Pierre Jouanolou. Équations de Pfaff algébriques sur un espace projectif. (French) []. Lecture Notes in Mathematics, 708: 80–135, 1979. CODEN LNMAA2. ISBN 3-540-09239-0 (print), 3-540-35344-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063395/.

Jouanolou:1979:QPG

[Jou79f] Jean-Pierre Jouanolou. Quelques propriétés globales des feuilletages analytiques complexes. (French) []. Lecture Notes in Mathematics, 708:198–251, 1979. CODEN LNMAA2. ISBN 3-540-09239-0 (print), 3-540-35344-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063398/.

Jouanolou:1979:SPA

[Jou79g] Jean-Pierre Jouanolou. Systèmes de Pfaff algébriques. (French) []. Lecture Notes in Mathematics, 708:136–156, 1979. CODEN LNMAA2. ISBN 3-540-09239-0 (print), 3-540-35344-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063396/.

Jozsa:1979:SPT

[Joz79] Richard Jozsa. Sheaves in physics — twistor theory. Lecture Notes in Mathematics, 753:492–499, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061830/.

Johnstone:1978:ICT

[JPR^{+}78] Peter T. Johnstone, Robert Paré, R. D. Rosebrugh, D. Schumacher, R. J. Wood, and G. C. Wraith. Indexed Categories and Their Applications, volume 661 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08914-4 (print), 3-540-35762-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 260 pp. LCCN QA3

.L28 no. 661. URL http://link.springer.com/book/10.1007/
BFb0061360;   http://www.springerlink.com/content/978-
3-540-35762-9.

Janakiraman:1977:SPS

[JR77] S. Janakiraman and K. M. Rangaswamy. Strongly pure subgroups of Abelian groups. Lecture Notes in Mathematics, 573:57–65, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087812/.

Johnen:1976:DIT

[JS76] H. Johnen and K. Scherer. Direct and inverse theorems for best approximation by Λ-splines. Lecture Notes in Mathematics, 501: 116–131, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079744/.

Johnen:1977:EFM

[JS77] H. Johnen and K. Scherer. On the equivalence of the K-functional and moduli of continuity and some applications. Lecture Notes in Mathematics, 571:119–140, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086569/.

Juhasz:1977:TST

[Juh77] I. Juhász. Two set-theoretic problems in topology. Lecture Notes in Mathematics, 609:115–123, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068677/.

Jurek:1975:LTT

[Jur75] Z. Jurek. A limit theorem for truncated random variables. Lecture Notes in Mathematics, 472:69–76, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081946/.

Jurek:1978:ICS

[Jur78a] Zbigniew J. Jurek. Intersection of the classes of s-selfdecomposable and s-semi-stable distributions. Lecture Notes in Mathematics, 656:77–93, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068813/.

Jurkat:1978:ASS

[Jur78b] W. B. Jurkat. Asymptotische Sektoren und Stokes’sches Phänomen. (German) []. Lecture Notes in Mathematics, 637: 108–114, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068246/.

Jurkat:1978:ESG

[Jur78c] W. B. Jurkat. Der Einzigkeitssatz und der Schließungssatz. (German) []. Lecture Notes in Mathematics, 637:88–101, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068244/.

Jurkat:1978:EIG

[Jur78d] W. B. Jurkat. Die eigentlichen Invarianten. (German) []. Lecture Notes in Mathematics, 637:124–129, 1978. CODEN LN-MAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068248/.

Jurkat:1978:DGD

[Jur78e] W. B. Jurkat. Dreieckig geblockte Dgln. (German) []. Lecture Notes in Mathematics, 637:130–145, 1978. CODEN LN-MAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068249/.

Jurkat:1978:EG

[Jur78f] W. B. Jurkat. Einleitung. (German) []. Lecture Notes in Mathematics, 637:1–10, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068237/.

Jurkat:1978:FFG

[Jur78g] W. B. Jurkat. Formale Feinstruktur. (German) [Formal fine structure]. Lecture Notes in Mathematics, 637:50–62, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068241/.

Jurkat:1978:FIG

[Jur78h] W. B. Jurkat. Formale Invarianten. (German) [Formal invariants]. Lecture Notes in Mathematics, 637:34–50, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068240/.

Jurkat:1978:GGB

[Jur78i] W. B. Jurkat. Grundtatsachen. (German) [Basic facts]. Lecture Notes in Mathematics, 637:11–16, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068238/.

Jurkat:1978:HNG

[Jur78j] W. B. Jurkat. Historische Notizen. (German) []. Lecture Notes in Mathematics, 637:184–188, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068253/.

Jurkat:1978:MDG

[Jur78k] W. B. Jurkat. Meromorphe Differentialgleichungen. (German) [Meromorphic differential equations], volume 637 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 194 pp. LCCN QA3

.L28 no. 637. URL http://link.springer.com/book/10.1007/
BFb0068236;   http://www.springerlink.com/content/978-
3-540-35923-4.

Jurkat:1978:RSK

[Jur78l] W. B. Jurkat. Reduzibilität und Spalten-Konvergenz bei formalen Lösungen. (German) []. Lecture Notes in Mathematics, 637:145– 159, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068250/.

Jurkat:1978:STF

[Jur78m] W. B. Jurkat. Sektorielle Transformationen und Freiheitsgrad des Verbindungssystems. (German) []. Lecture Notes in Mathematics, 637:102–108, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068245/.

Jurkat:1978:SML

[Jur78n] W. B. Jurkat. Spezielle Matrix- und Lösungsfunktionen. (German) []. Lecture Notes in Mathematics, 637:159–176, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068251/.

Jurkat:1978:SRU

[Jur78o] W. B. Jurkat. Stokes’sche Richtungen und Übergangsmatrizen. (German) []. Lecture Notes in Mathematics, 637:63–75, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068242/.

Jurkat:1978:SNF

[Jur78p] W. B. Jurkat. Struktur und Normalisierung der formalen Lösungen. (German) []. Lecture Notes in Mathematics, 637:17– 33, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068239/.

Jurkat:1978:VAS

[Jur78q] W. B. Jurkat. Verbindungsmatrizen und asymptotische Sektoren. (German) []. Lecture Notes in Mathematics, 637:76–87, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068243/.

Jurkat:1978:VAP

[Jur78r] W. B. Jurkat. Verbindungssysteme und der Äquivalenzsatz für Paare. (German) []. Lecture Notes in Mathematics, 637:115– 123, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068247/.

Jurkat:1978:WEB

[Jur78s] W. B. Jurkat. Weitere Ergebnisse und Bemerkungen. (German) []. Lecture Notes in Mathematics, 637:176–183, 1978. CODEN LNMAA2. ISBN 3-540-08659-5 (print), 3-540-35923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068252/.

Johnson:1977:NAO

[JW77] Jerry Johnson and John Wolfe. Norm attaining operators on C(S) spaces. Lecture Notes in Mathematics, 604:41–43, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069204/.

Johnstone:1978:ATT

[JW78] Peter T. Johnstone and Gavin C. Wraith. Algebraic theories in toposes. Lecture Notes in Mathematics, 661:141–242, 1978. CO-DEN LNMAA2. ISBN 3-540-08914-4 (print), 3-540-35762-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0061363/.

Jeulin:1978:GFS

[JY78] T. Jeulin and M. Yor. Grossissement d’une filtration et semimartingales: Formules explicites. (French) []. Lecture Notes in Mathematics, 649:78–97, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064597/.

Kadison:1978:PIA

[Kad78] Richard V. Kadison. Product isometries and automorphisms of the Car algebra. Lecture Notes in Mathematics, 650:173–185, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067401/.

Kahane:1976:MDF

[Kah76a] J.-P. Kahane. Mesures et dimensions. (French) []. Lecture Notes in Mathematics, 565:94–103, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091449/.

Kahnert:1976:HMH

[Kah76b] D. Kahnert. Haar–Mass und Hausdorff–Mass. (German) []. Lecture Notes in Mathematics, 541:13–23, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081036/.

Kahn:1978:SD

[Kah78] Daniel S. Kahn. On the stable decomposition of Ω^{∞}S^{∞}A. Lecture Notes in Mathematics, 658:206–214, 1978. CODEN LN-MAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068718/.

Kaijser:1976:SRI

[Kai76] Sten Kaijser. Some remarks on injective Banach algebras. Lecture Notes in Mathematics, 512:84–95, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080026/.

Kajiwara:1975:OEH

[Kaj75] Joji Kajiwara. Opérateurs d’ dans les espaces de Hilbert avec croissance polynômiale. (French) []. Lecture Notes in Mathematics, 474:91–108, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077401/.

Kalisch:1978:SSA

[Kal78] Gerhard K. Kalisch. On singular self-adjoint Sturm–Liouville operators. Lecture Notes in Mathematics, 693:109–114, 1978. CO-DEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0064667/.

Kamae:1976:NNE

[Kam76] Teturo Kamae. Normal numbers and ergodic theory. Lecture Notes in Mathematics, 550:253–269, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077494/.

Kantor:1975:CDG

[Kan75a] J.-M. Kantor. Le complexe de Dolbeault–Grothendieck sur les espaces analytiques. (French) []. Lecture Notes in Mathematics, 474:142–154, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077405/.

Kantor:1975:CEA

[Kan75b] Jean-Michel Kantor. Courants sur les espaces analytiques complexes: Généralisation du lemme de Dolbeault–Grothendieck. (French) []. Lecture Notes in Mathematics, 482:159–167, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080244/.

Kane:1978:BHF

[Kan78a] Richard Kane. Bp homology and finite H-spaces. Lecture Notes in Mathematics, 673:93–105, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064690/.

Kanter:1978:RLF

[Kan78b] Marek Kanter. Random linear functionals and why we study them. Lecture Notes in Mathematics, 645:114–123, 1978. CODEN LN-MAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069668/.

Kannan:1979:PTS

[Kan79] V. Kannan. Pairs of topologies with same family of continuous self-maps. Lecture Notes in Mathematics, 719:176–184, 1979. CO-DEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0065270/.

Kaplansky:1978:FTA

[Kap78] Irving Kaplansky. Five theorems on abelian groups. Lecture Notes in Mathematics, 697:47–51, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103121/.

Karpilovsky:1977:SRG

[Kar77] G. Karpilovsky. Some results in group representation theory. Lecture Notes in Mathematics, 573:143–146, 1977. CODEN LN-MAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087820/.

Karoubi:1979:TAC

[Kar79] Max Karoubi. K-théorie algébrique de certaines algèbres d’opérateurs. (French) []. Lecture Notes in Mathematics, 725:254– 290, 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062621/.

Kastler:1975:SEQ

[Kas75] Daniel Kastler. Stability and equilibrium in quantum statistical mechanics. Lecture Notes in Mathematics, 466:86–100, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082200/.

Kastler:1978:AFE

[Kas78] Daniel Kastler. Algebraic features of equilibrium states. Lecture Notes in Mathematics, 650:145–155, 1978. CODEN LN-MAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067397/.

Kasube:1979:UAU

[Kas79] Herbert E. Kasube. Unique and almost unique factorization. Lecture Notes in Mathematics, 751:200–205, 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062709/.

Kato:1975:QLE

[Kat75] Tosio Kato. Quasi-linear equations of evolution, with applications to partial differential equations. Lecture Notes in Mathematics, 448:25–70, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067080/.

Kato:1976:SPS

[Kat76] Tosio Kato. Singular perturbation and semigroup theory. Lecture Notes in Mathematics, 565:104–112, 1976. CODEN LN-MAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091450/.

Katz:1977:RMF

[Kat77] Nicholas M. Katz. A result on modular forms in characteristic p. Lecture Notes in Mathematics, 601:53–61, 1977. CODEN LN-MAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063944/.

Kauffman:1978:SBF

[Kau78] Louis H. Kauffman. Signature of branched fibrations. Lecture Notes in Mathematics, 685:203–217, 1978. CODEN LN-MAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062972/.

Kawai:1975:PDO

[Kaw75] Takahiro Kawai. Pseudo-differential operators acting on the sheaf of microfunctions. Lecture Notes in Mathematics, 449:54–69, 1975. CODEN LNMAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062915/.

Kawazu:1976:MBP

[Kaw76] Kiyoshi Kawazu. On multitype branching processes with immigration. Lecture Notes in Mathematics, 550:270–275, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077495/.

Kawamata:1979:CNC

[Kaw79a] Yujiro Kawamata. On the classification of non-complete algebraic surfaces. Lecture Notes in Mathematics, 732:215–232, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066646/.

Kawarada:1979:NMF

[Kaw79b] H. Kawarada. Numerical methods for free surface problems by means of penalty. Lecture Notes in Mathematics, 704:282– 291, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063626/.

Kazamaki:1976:SRC

[Kaz76] N. Kazamaki. A simple remark on the conditioned square functions for martingale transforms. Lecture Notes in Mathematics, 511:40–43, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101095/.

Kazamaki:1978:RPG

[Kaz78] N. Kazamaki. A remark on a problem of Girsanov. Lecture Notes in Mathematics, 649:47–50, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064591/.

Keane:1975:AHT

[Kea75] Orville Keane. Abstract Horn theories. Lecture Notes in Mathematics, 445:15–50, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061292/.

Kearton:1978:ACK

[Kea78] C. Kearton. Attempting to classify knot modules and their Hermitian pairings. Lecture Notes in Mathematics, 685:227–242, 1978.

CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062974/.

Kechris:1978:APO

[Kec78a] Alexander S. Kechris. Ad and projective ordinals. Lecture Notes in Mathematics, 689:91–132, 1978. CODEN LNMAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069296/.

Kechris:1978:FA

[Kec78b] Alexander S. Kechris. Forcing in analysis. Lecture Notes in Mathematics, 669:277–302, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103105/.

Kechris:1978:SC

[Kec78c] Alexander S. Kechris. On spector classes. Lecture Notes in Mathematics, 689:245–277, 1978. CODEN LNMAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069303/.

Kelly:1975:HS

[Kel75a] John B. Kelly. Hypermetric spaces. Lecture Notes in Mathematics, 490:17–31, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081128/.

Kelly:1975:EFF

[Kel75b] Leroy M. Kelly. On the equilateral feeble four-point property. Lecture Notes in Mathematics, 490:14–16, 1975. CODEN LN-MAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081127/.

Kelly:1975:GML

[Kel75c] Prof. L. M. Kelly, editor. The Geometry of Metric and Linear Spaces: Proceedings of a Conference Held at Michigan State University, East Lansing, June 17–19, 1974, volume 490 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 490. URL http://link.springer.com/book/10.1007/
BFb0081125;   http://www.springerlink.com/content/978-
3-540-37946-1.

Keller:1979:CMB

[Kel79] Herbert B. Keller. Constructive methods for bifurcation and nonlinear eigenvalue problems. Lecture Notes in Mathematics, 704: 241–251, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063623/.

Keogh:1977:SRD

[Keo77] F. R. Keogh. Some recent developments in the theory of univalent functions. Lecture Notes in Mathematics, 599:76–92, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096827/.

Kerber:1975:ACT

[Ker75a] Prof. Dr. Adalbert Kerber. An application to combinatorics: The theory of enumeration under group action. Lecture Notes in Mathematics, 495:103–166, 1975. CODEN LNMAA2. ISBN 3-540-07535-6 (print), 3-540-38024-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085744/.

Kerber:1975:ART

[Ker75b] Prof. Dr. Adalbert Kerber. An application to representation theory: Symmetrization of inner tensor products of representations. Lecture Notes in Mathematics, 495:62–102, 1975. CODEN LN-MAA2. ISBN 3-540-07535-6 (print), 3-540-38024-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085743/.

Kerber:1975:CWP

[Ker75c] Prof. Dr. Adalbert Kerber. Characters of wreath products. Lecture Notes in Mathematics, 495:3–61, 1975. CODEN LNMAA2. ISBN 3-540-07535-6 (print), 3-540-38024-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085742/.

Kerber:1975:I

[Ker75d] Prof. Dr. Adalbert Kerber. Introduction. Lecture Notes in Mathematics, 495:1–2, 1975. CODEN LNMAA2. ISBN 3-540-07535-6 (print), 3-540-38024-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085741.

Kerber:1975:RPG

[Ker75e] Prof. Dr. Adalbert Kerber. Representations of Permutation Groups II, volume 495 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07535-6 (print), 3-540-38024-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 175 pp. LCCN QA3 .L28 no. 495. URL http:

//link.springer.com/book/10.1007/BFb0085740;   http://
www.springerlink.com/content/978-3-540-38024-5.

Kervaire:1975:FRI

[Ker75f] Michel Kervaire. Fractions rationnelles invariantes [d’après H. W. Lenstra]. (French) []. Lecture Notes in Mathematics, 431:170– 189, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066370/.

Kijowski:1977:GSQ

[Kij77] Jerzy Kijowski. Geometric structure of quantization. Lecture Notes in Mathematics, 570:97–108, 1977. CODEN LN-MAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087785/.

Kikuchi:1979:FEA

[Kik79] F. Kikuchi. Finite element approximations to bifurcation problems of turning point type. Lecture Notes in Mathematics, 704: 252–266, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063624/.

Kilambi:1978:CSA

[Kil78] Srinivasacharyulu Kilambi. Sur certaines surfaces analytiques complexes compactes. (French) []. Lecture Notes in Mathematics, 670:337–340, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064407/.

Kingman:1976:SP

[Kin76] J. F. C. Kingman. Subadditive processes. Lecture Notes in Mathematics, 539:167–223, 1976. CODEN LNMAA2. ISBN 3-540-07858-4 (print), 3-540-38119-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079697/.

Kishi:1976:LPF

[Kis76a] M. Kishi. L’unicité de la puissance fractionnaire. (French) [Uniqueness of the fractional power]. Lecture Notes in Mathematics, 563: 130–141, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087575/.

Kishi:1976:QRE

[Kis76b] M. Kishi. Quelques remarques sur l’existence de resolvantes. (French) []. Lecture Notes in Mathematics, 563:121–129, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087574/.

Kiselman:1977:CFE

[Kis77a] Christer O. Kiselman. Construction de fonctions entières à rayon de convergence donné. (French) []. Lecture Notes in Mathematics, 578:246–253, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091474/.

Kiselman:1977:FDC

[Kis77b] Christer O. Kiselman. Fonctions delta-convexes, delta-sousharmoniques et delta-plurisousharmoniques. (French) []. Lecture Notes in Mathematics, 578:93–107, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091464/.

Kashiwara:1975:MHP

[KK75] Masaki Kashiwara and Takahiro Kawai. Micro-hyperbolic pseudodifferential operators. Lecture Notes in Mathematics, 449:70– 82, 1975. CODEN LNMAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062916/.

Krawczyk:1976:EGG

[KK76] A. Krawczyk and M. Krynicki. Ehrenfeucht games for generalized quantifiers. Lecture Notes in Mathematics, 537:145–152, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096899/.

Kallianpur:1978:MTA

[KK78a] G. Kallianpur and D. Kölzowa, editors. Measure Theory Applications to Stochastic Analysis: Proceedings, Oberwolfach Conference, Germany, July 3–9, 1977, volume 695 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 695. URL http://link.springer.com/book/10.1007/
BFb0062649;   http://www.springerlink.com/content/978-
3-540-35556-4.

Kluvanek:1978:BBP

[KK78b] Igor Kluvánek and Greg Knowles. The Bang–Bang principle. Lecture Notes in Mathematics, 680:138–151, 1978. CODEN LN-MAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065315/.

Kuribayashi:1979:WPA

[KK79] Akikazu Kuribayashi and Kaname Komiya. On Weierstrass points and automorphisms of curves of genus three. Lecture Notes in Mathematics, 732:253–299, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066649/.

Kubert:1977:UMF

[KL77] Dan Kubert and Serge Lang. Units in the modular function field. Lecture Notes in Mathematics, 601:247–275, 1977. CODEN LN-MAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063950/.

Kennison:1979:SRD

[KL79a] John F. Kennison and Carl S. Ledbetter. Sheaf representations and the Dedekind reals. Lecture Notes in Mathematics, 753:500– 513, 1979. CODEN LNMAA2. ISBN 3-540-09564-0 (print), 3-540-34849-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061831/.

Kleppe:1979:GPD

[KL79b] H. Kleppe and D. Laksov. The generic perfectiness of determinantal schemes. Lecture Notes in Mathematics, 732:244–252, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066648/.

Klares:1979:MSP

[Kla79] B. Klarès. Sur la monodromie des systèmes de Pfaff du type de Fuchs sur P_{m}(C). (French) []. Lecture Notes in Mathematics, 712:293–324, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062821/.

Klee:1975:RSC

[Kle75] Victor Klee. Ratio-sequences of chains in connected metric spaces. Lecture Notes in Mathematics, 490:134–146, 1975. CODEN LN-MAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081138/.

Kleinberg:1977:C

[Kle77a] Eugene M. Kleinberg. Chapter I. Lecture Notes in Mathematics, 612:9–21, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069387/.

Kleinberg:1977:CIa

[Kle77b] Eugene M. Kleinberg. Chapter II. Lecture Notes in Mathematics, 612:22–48, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069388/.

Kleinberg:1977:CIb

[Kle77c] Eugene M. Kleinberg. Chapter III. Lecture Notes in Mathematics, 612:49–54, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069389/.

Kleinberg:1977:CIc

[Kle77d] Eugene M. Kleinberg. Chapter IV. Lecture Notes in Mathematics, 612:55–78, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069390/.

Kleinberg:1977:CVa

[Kle77e] Eugene M. Kleinberg. Chapter V. Lecture Notes in Mathematics, 612:79–116, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069391/.

Kleinberg:1977:CVb

[Kle77f] Eugene M. Kleinberg. Chapter VI. Lecture Notes in Mathematics, 612:117–128, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069392/.

Kleinberg:1977:CVc

[Kle77g] Eugene M. Kleinberg. Chapter VII. Lecture Notes in Mathematics, 612:129–148, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069393/.

Kleinberg:1977:ICAa

[Kle77h] Eugene M. Kleinberg. Infinitary combinatorics and the axiom of determinateness. Lecture Notes in Mathematics, 612:1–8, 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069386/.

Kleinberg:1977:ICAb

[Kle77i] Eugene M. Kleinberg. Infinitary Combinatorics and the Axiom of Determinateness, volume 612 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08440-1 (print), 3-540-37097-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 150 pp. LCCN QA3

.L28 no. 612. URL http://link.springer.com/book/10.1007/
BFb0069385;   http://www.springerlink.com/content/978-
3-540-37097-0.

Klemm:1977:CMT

[Kle77j] Michael Klemm. Charaktere mehrfach transitiver permutationsgruppen. Lecture Notes in Mathematics, 579:281–286, 1977. CO-DEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0090023/.

Kock:1975:STT

[KLM75] A. Kock, P. Lecouturier, and C. J. Mikkelsen. Some topos theoretic concepts of finiteness. Lecture Notes in Mathematics, 445: 209–283, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061297/.

Kloeden:1978:GCS

[Klo78] P. E. Kloeden. General control systems. Lecture Notes in Mathematics, 680:119–137, 1978. CODEN LNMAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065314/.

Kabanov:1976:CAC

[KLŠ76] Yu. M. Kabanov, R. Š. Lipcer, and A. N. Širyaev. Criteria of absolute continuity of measures corresponding to multivariate point processes. Lecture Notes in Mathematics, 550:232–252, 1976. CO-DEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077493/.

Kluit:1977:N

[Klu77] P. G. Kluit. On the normalizer of γ_{0}(N). Lecture Notes in Mathematics, 601:239–246, 1977. CODEN LNMAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063949/.

Krawczyk:1977:RPG

[KM77] A. Krawczyk and W. Marek. On the rules of proof generated by hierarchies. Lecture Notes in Mathematics, 619:227–239, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067654/.

Kanamori:1978:ELC

[KM78a] A. Kanamori and M. Magidor. The evolution of large cardinal axioms in set theory. Lecture Notes in Mathematics, 669:99– 275, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103104/.

Kechris:1978:CSP

[KM78b] Alexander S. Kechris and Yiannis N. Moschovakis, editors. Cabal Seminar 76–77: Proceedings, Caltech–UCLA Logic Seminar 1976–77, volume 689 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 689. URL http:

//link.springer.com/book/10.1007/BFb0069293;   http://
www.springerlink.com/content/978-3-540-35626-4.

Kechris:1978:NTS

[KM78c] Alexander S. Kechris and Yiannis N. Moschovakis. Notes on the theory of scales. Lecture Notes in Mathematics, 689:1–53, 1978. CODEN LNMAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069294/.

Kreisel:1975:UAL

[KMS75] G. Kreisel, G. E. Mints, and S. G. Simpson. The use of abstract language in elementary metamathematics: Some pedagogic examples. Lecture Notes in Mathematics, 453:38–131, 1975. CODEN LNMAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064871/.

Keynes:1978:LNA

[KN78] H. B. Keynes and D. Newton. Lifting in non-abelian (G, τ)- extensions. Lecture Notes in Mathematics, 668:173–178, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101790/.

Kempf:1979:LVR

[KN79] George Kempf and Linda Ness. The length of vectors in representation spaces. Lecture Notes in Mathematics, 732:233–243, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066647/.

Knab:1979:WLL

[Kna79a] Otto Knab. Zur Werteverteilung der Lösungen linearer Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 747: 189–204, 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063974/.

Knapp:1979:BSH

[Kna79b] K. Knapp. On the bi-stable J-homomorphism. Lecture Notes in Mathematics, 763:13–22, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088075/.

Knill:1978:IIC

[Kni78] R. J. Knill. On the index of isolated closed tori. Lecture Notes in Mathematics, 652:196–211, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063513/.

Knobloch:1975:OCP

[Kno75] H. W. Knobloch. On optimal control problems with state space constraints. Lecture Notes in Mathematics, 477:177–187, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079176/.

Knowles:1976:LVM

[Kno76] Greg Knowles. On Liapunov vector measures. Lecture Notes in Mathematics, 541:61–67, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081042/.

Kolata:1979:NSA

[KO79] William Kolata and John Osborn. Nonselfadjoint spectral approximation and the finite element method. Lecture Notes in Mathematics, 701:115–133, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062077/.

Kochen:1975:MTL

[Koc75] Simon Kochen. The model theory of local fields. Lecture Notes in Mathematics, 499:384–425, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079426/.

Kohler:1975:EAO

[Köh75] Michael Köhler. Explicit approximation of optimal control processes. Lecture Notes in Mathematics, 477:188–201, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079177/.

Kollmer:1977:HAC

[Kol77] Heinrich Kollmer. On hyperbolic attractors of codimension one. Lecture Notes in Mathematics, 597:330–334, 1977. CODEN LN-MAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085363/.

Kolaitis:1978:RSS

[Kol78] Phokion G. Kolaitis. On recursion in E and semi-spector classes. Lecture Notes in Mathematics, 689:209–243, 1978. CODEN LN-MAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069302/.

Komatsu:1975:GPI

[Kom75] Hikosaburo Komatsu. Generalized Poisson integrals and regularity of functions. Lecture Notes in Mathematics, 457:232–248, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067108/.

Komatsu:1976:SSP

[Kom76] Takashi Komatsu. Statistics of stochastic processes with jumps. Lecture Notes in Mathematics, 550:276–289, 1976. CODEN LN-MAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077496/.

Komornicki:1978:MTC

[Kom78] Wojciech Komornicki. Multiplications in two-cell spectra. Lecture Notes in Mathematics, 658:215–223, 1978. CODEN LN-MAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068719/.

Kono:1976:EAT

[Kôn76] Norio Kôno. Évolution asymptotique des temps d’arrêt et des temps de séjour liés aux trajectoires de certaines fonctions aléatoires gaussiennes. (French) []. Lecture Notes in Mathematics, 550:290–296, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077497/.

Koornwinder:1977:HSF

[Koo77] Tom Koornwinder. Harmonics and spherical functions on Grassmann manifolds of rank two and two-variable analogues of Jacobi polynomials. Lecture Notes in Mathematics, 571:141–154, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086570/.

Kori:1976:CSD

[Kor76] Tosiaki Kori. Sur une classe des solutions du problème de Dirichlet extérieur dans un espace harmonique de Brelot. (French) []. Lecture Notes in Mathematics, 563:142–160, 1976. CODEN LN-MAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087576/.

Koranyi:1979:CSS

[Kor79] Adam Korányi. Compactifications of symmetric spaces and harmonic functions. Lecture Notes in Mathematics, 739:341– 366, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062500/.

Kostant:1975:VME

[Kos75a] Bertram Kostant. Verma modules and the existence of quasiinvariant differential operators. Lecture Notes in Mathematics, 466:101–128, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082201/.

Koszul:1975:TCJ

[Kos75b] Jean-Louis Koszul. Travaux de S. S. Chern et J. Simons sur les classes. (French) []. Lecture Notes in Mathematics, 431:69– 88, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066365/.

Knus:1976:BGC

[KOS76a] M. A. Knus, M. Ojanguren, and D. J. Saltman. On Brauer groups in characteristic p. Lecture Notes in Mathematics, 549:25–49, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077331/.

Kosters:1976:CPS

[Kös76b] Dr. Heinz-Walter Kösters. Computation of periodic M-splines with equi-spaced nodes. Lecture Notes in Mathematics, 501:132– 143, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079745/.

Koszul:1976:RFE

[Kos76c] Jean-Louis Koszul. Rigidité forte des espaces riemanniens localement symétriques. (French) []. Lecture Notes in Mathematics, 514:223–237, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080068/.

Kostant:1977:GMG

[Kos77] Bertram Kostant. Graded manifolds, graded Lie theory, and prequantization. Lecture Notes in Mathematics, 570:177–306, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087788/.

Koschorke:1979:TPI

[Kos79] Ulrich Koschorke. On the (n + 1)-tuple points of immersed nspheres. Lecture Notes in Mathematics, 763:132–134, 1979. CO-DEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0088082/.

Kossmann:1978:BVT

[Koß78] H. Koßmann. Boundary-value technique for the numerical solution of periodic parabolic problems. Lecture Notes in Mathematics, 679:41–55, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067865/.

Kovacs:1978:VNG

[Kov78] L. G. Kovács. Varieties of nilpotent groups of small class. Lecture Notes in Mathematics, 697:205–229, 1978. CODEN LN-MAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103131/.

Kowalski:1975:IMP

[Kow75] Z. S. Kowalski. Invariant measures for piecewise monotonic transformations. Lecture Notes in Mathematics, 472:77–94, 1975. CO-DEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081947/.

Kirby:1977:ISM

[KP77] L. A. S. Kirby and J. B. Paris. Initial segments of models of Peano’s axioms. Lecture Notes in Mathematics, 619:211–226, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067653/.

Kurke:1978:AEP

[KPP^{+}78a] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Anhang: Eine Eliminationstheorie für Potenzreihenringe. (German) []. Lecture Notes in Mathematics, 634:59–79, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068286/.

Kurke:1978:AHR

[KPP^{+}78b] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Approximationssätze für Henselsche Ringe. (German) []. Lecture Notes in Mathematics, 634:7–58, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068285/.

Kurke:1978:ALR

[KPP^{+}78c] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Die Approximationseigenschaft lokaler Ringe, volume 634 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 204 pp. LCCN QA3 .L28 no. 634. URL http:

//link.springer.com/book/10.1007/BFb0068283;   http://
www.springerlink.com/content/978-3-540-35928-9.

Kurke:1978:AZL

[KPP^{+}78d] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Die Approximationseigenschaft zweidimensionaler lokaler Ringe. (German) []. Lecture Notes in Mathematics, 634:187–199, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068291/.

Kurke:1978:SAL

[KPP^{+}78e] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Die strenge Approximationseigenschaft lokaler Ringe. (German) []. Lecture Notes in Mathematics, 634: 80–137, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068287/.

Kurke:1978:WGN

[KPP^{+}78f] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Die Weierstraß-Grauertsche Normalform von Idealbasen. (German) []. Lecture Notes in Mathematics, 634:160–178, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068289/.

Kurke:1978:SAC

[KPP^{+}78g] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Ein spezieller Approximationssatz in Charakteristik O. (German) []. Lecture Notes in Mathematics, 634:138–159, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068288/.

Kurke:1978:EG

[KPP^{+}78h] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Einleitung. (German) []. Lecture Notes in Mathematics, 634:1–5, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068284/.

Kurke:1978:IRA

[KPP^{+}78i] Herbert Kurke, Gerhard Pfister, Dorin Popescu, Marco Roczen, and Tadeusz Mostowski. Zur Idealtheorie von Ringen mit Approximationseigenschaft. (German) []. Lecture Notes in Mathematics, 634:179–186, 1978. CODEN LNMAA2. ISBN 3-540-08656-0 (print), 3-540-35928-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068290/.

Kannan:1979:HLC

[KR79] V. Kannan and M. Rajagopalan. Hereditarily locally compact separable spaces. Lecture Notes in Mathematics, 719:185–195, 1979. CODEN LNMAA2. ISBN 3-540-09503-9 (print), 3-540-35193-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065271/.

Kraft:1975:AES

[Kra75a] Prof. Dr. Hanspeter Kraft. Anhang: Das Endomorphismen-Schema eines R-Moduls. (German) []. Lecture Notes in Mathematics, 455:158–162, 1975. CODEN LNMAA2. ISBN 3-540-07158-X (print), 3-540-37486-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063303/.

Kraft:1975:CSU

[Kra75b] Prof. Dr. Hanspeter Kraft. Cohenschemata, Struktursatz für unipotente Gruppen. (German) []. Lecture Notes in Mathematics, 455:13–75, 1975. CODEN LNMAA2. ISBN 3-540-07158-X (print), 3-540-37486-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063300/.

Kraft:1975:EG

[Kra75c] Prof. Dr. Hanspeter Kraft. Einleitung. (German) []. Lecture Notes in Mathematics, 455:1–12, 1975. CODEN LNMAA2. ISBN 3-540-07158-X (print), 3-540-37486-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063299/.

Kraft:1975:KAG

[Kra75d] Prof. Dr. Hanspeter Kraft. Kommutative algebraische Gruppen und Ringe, volume 455 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07158-X (print), 3-540-37486-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 163 pp. LCCN QA3 .L28 no. 455. URL http:

//link.springer.com/book/10.1007/BFb0063298;   http://
www.springerlink.com/content/978-3-540-37486-2.

Kraft:1975:KRG

[Kra75e] Prof. Dr. Hanspeter Kraft. Kommutative Ringschemata. (German) []. Lecture Notes in Mathematics, 455:131–157, 1975. CO-DEN LNMAA2. ISBN 3-540-07158-X (print), 3-540-37486-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063302/.

Kraft:1975:MEG

[Kra75f] Prof. Dr. Hanspeter Kraft. R-Moduln und Einheitengruppen. (German) []. Lecture Notes in Mathematics, 455:76–130, 1975. CODEN LNMAA2. ISBN 3-540-07158-X (print), 3-540-37486-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063301/.

Krauss:1975:QE

[Kra75g] Peter Krauss. Quantifier elimination. Lecture Notes in Mathematics, 499:426–444, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079427/.

Krajewski:1976:NSS

[Kra76a] S. Krajewski. Non-standard satisfaction classes. Lecture Notes in Mathematics, 537:121–144, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096898/.

Kral:1976:SNE

[Krá76b] Joseph Král. Singularités non essentielles des solutions des équations aux derivées partielles. (French) []. Lecture Notes in Mathematics, 518:95–106, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080407/.

Kranz:1978:CMP

[Kra78a] Przemysław Kranz. Control measure problem in some classes of F -spaces. Lecture Notes in Mathematics, 645:124–130, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069669/.

Krause:1978:SRD

[Kra78b] Günter Krause. Some recent developments in the theory of Noetherian rings. Lecture Notes in Mathematics, 641:209–219, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064849/.

Kramarz:1979:HMN

[Kra79] Luis Kramarz. Hermite methods for the numerical solution of ordinary initial value problems. Lecture Notes in Mathematics, 701: 134–148, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062078/.

Krebs:1975:IFD

[Kre75a] Hélène Krebs. Intégrales de formes différentielles d′′-fermées sur les cycles analytiques compacts de certains ouverts de IP_{n}(c). (French) [Integrals of differential forms d′′-closed on certain compact analytical open cycles IP_{n}(c)]. Lecture Notes in Mathematics, 482:216–250, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080248/.

Kree:1975:SFE

[Kré75b] Paul Krée. Solutions faibles d’équations aux dérivées fonctionnelles II. (French) []. Lecture Notes in Mathematics, 474:16– 47, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077397/.

Kreisel:1975:ORG

[Kre75c] G. Kreisel. Observations on a recent generalization of completeness theorems due to Schütte. Lecture Notes in Mathematics, 500:164–181, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079551/.

Kree:1976:TDC

[Kré76a] Paul Krée. Théories des distributions et calculs différentiels sur un espace de Banach. (French) []. Lecture Notes in Mathematics, 524:163–192, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078005/.

Kreimer:1976:AEC

[Kre76b] H. F. Kreimer. Abelian p-extensions and cohomology. Lecture Notes in Mathematics, 549:104–111, 1976. CODEN LN-MAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077338/.

Kree:1977:TMH

[Kré77] Paul Krée. Théorie de la mesure et holomorphie en dimension infinie. (French) []. Lecture Notes in Mathematics, 578:44–70, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091462/.

Kree:1978:MFA

[Kré78a] Paul Krée. Méthodes fonctorielles en analyse de dimension infinie et holomorphie anticommutative. (French) []. Lecture Notes in Mathematics, 694:134–171, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063247/.

Kree:1978:MHM

[Kré78b] Paul Krée. Méthodes holomorphes et méthodes nucléaires en analyse de dimension infinie et en théorie quantique des champs. (French) []. Lecture Notes in Mathematics, 644:212–254, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066849/.

Kreth:1978:TDN

[Kre78c] H. Kreth. Time-discretisations for nonlinear evolution equations. Lecture Notes in Mathematics, 679:57–63, 1978. CODEN LN-MAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067866/.

Krickeberg:1976:AAG

[Kri76] Klaus Krickeberg. An alternative approach to Glivenko–Cantelli theorems. Lecture Notes in Mathematics, 566:57–67, 1976. CO-DEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096879/.

Kruglov:1976:MAI

[Kru76] V. M. Kruglov. The method of accompanying infinitely divisible distributions. Lecture Notes in Mathematics, 550:316–323, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077499/.

Krushkal:1979:PSB

[Kru79] S. L. Krushkal’. To the problem of the supports of Beltrami differentials for Kleinian groups. Lecture Notes in Mathematics, 743:132–139, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079488/.

Kryukov:1978:MIP

[Kry78] V. I. Kryukov. Markov interaction processes and neuronal activity. Lecture Notes in Mathematics, 653:122–139, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070089/.

Kauffman:1977:APT

[KRZ77a] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. Applications of perturbation theory. Lecture Notes in Mathematics, 621:79–88, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064284/.

Kauffman:1977:CCA

[KRZ77b] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. Conditions on the coefficients for all powers to be limit-point. Lecture Notes in Mathematics, 621:89–97, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064285/.

Kauffman:1977:DIPa

[KRZ77c] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. The deficiency index problem for polynomials in symmetric differential expressions. Lecture Notes in Mathematics, 621:55–78, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064283/.

Kauffman:1977:DIPb

[KRZ77d] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. The Deficiency Index Problem for Powers of Ordinary Differential Expressions, volume 621 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 112 pp. LCCN QA3 .L28 no. 621. URL http:

//link.springer.com/book/10.1007/BFb0064277;   http://
www.springerlink.com/content/978-3-540-37024-6.

Kauffman:1977:FAP

[KRZ77e] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. Functional analytic preliminaries. Lecture Notes in Mathematics, 621: 3–5, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064279/.

Kauffman:1977:HOL

[KRZ77f] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. Higher order limit-point criteria. Lecture Notes in Mathematics, 621:41– 54, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064282/.

Kauffman:1977:I

[KRZ77g] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. Introduction. Lecture Notes in Mathematics, 621:1–2, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064278.

Kauffman:1977:LDO

[KRZ77h] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. Linear differential operators and the general classification theory of deficiency indices. Lecture Notes in Mathematics, 621:6–21, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064280/.

Kauffman:1977:SOL

[KRZ77i] Robert M. Kauffman, Thomas T. Read, and Anton Zettl. Second order limit-point, limit-circle conditions. Lecture Notes in Mathematics, 621:22–40, 1977. CODEN LNMAA2. ISBN 3-540-08523-8 (print), 3-540-37024-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064281/.

Kelley:1976:UPF

[KS76] J. L. Kelley and T. P. Srinivasan. A unified proof of Fubini theorem for Baire and Borel measures. Lecture Notes in Mathematics, 541:25–29, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081037/.

Kaminker:1977:TOP

[KS77] Jerome Kaminker and Claude Schochet. Topological obstructions to perturbations of pairs of operators. Lecture Notes in Mathematics, 575:70–77, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095702/.

Kaminker:1978:AEH

[KS78a] Jerome Kaminker and Claude Schochet. Analytic equivariant Khomology. Lecture Notes in Mathematics, 657:292–302, 1978. CO-DEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069242/.

Kashiwara:1978:PCP

[KS78b] M. Kashiwara and P. Schapira. Problême de Cauchy pour les systems d’équations différentielles et microdifférentielles dans le domaine complexe. (French) []. Lecture Notes in Mathematics, 660: 117–127, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082298/.

Kochman:1979:SHS

[KS79a] Stanley O. Kochman and Victor P. Snaith. On the stable homotopy of symplectic classifying and Thom spaces. Lecture Notes in Mathematics, 741:394–448, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062151/.

Kunisch:1979:PSF

[KS79b] K. Kunisch and W. Schappacher. Positive solutions of functional differential equations. Lecture Notes in Mathematics, 730:238– 246, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064321/.

Kalf:1975:STS

[KSWW75] H. Kalf, U.-W. Schmincke, J. Walter, and R. Wüst. On the spectral theory of Schrödinger and Dirac operators with strongly singular potentials. Lecture Notes in Mathematics, 448:182–226, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067087/.

Kamber:1975:CCFa

[KT75a] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Characteristic classes of flat bundles. Lecture Notes in Mathematics, 493: 39–53, 1975. CODEN LNMAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081561/.

Kamber:1975:CCFb

[KT75b] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Characteristic classes of foliated bundles. Lecture Notes in Mathematics, 493:54–92, 1975. CODEN LNMAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081562/.

Kamber:1975:CDA

[KT75c] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Cohomology of g-DG-algebras. Lecture Notes in Mathematics, 493:93– 132, 1975. CODEN LNMAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081563/.

Kamber:1975:E

[KT75d] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Erratum. Lecture Notes in Mathematics, 493:e1, 1975. CODEN LN-MAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081567.

Kamber:1975:EGC

[KT75e] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Examples of generalized characteristic classes for foliated bundles. Lecture Notes in Mathematics, 493:153–185, 1975. CODEN LN-MAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081565/.

Kamber:1975:FB

[KT75f] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Foliated bundles. Lecture Notes in Mathematics, 493:20–38, 1975. CODEN LNMAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081560/.

Kamber:1975:FBC

[KT75g] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Foliated Bundles and Characteristic Classes, volume 493 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 208 pp. LCCN QA3

.L28 no. 493. URL http://link.springer.com/book/10.1007/
BFb0081558;   http://www.springerlink.com/content/978-
3-540-37956-0.

Kamber:1975:F

[KT75h] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Foliations. Lecture Notes in Mathematics, 493:1–19, 1975. CODEN LN-MAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081559/.

Kamber:1975:NTC

[KT75i] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Non-trivial characteristic classes for flat bundles. Lecture Notes in Mathematics, 493:133–152, 1975. CODEN LNMAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081564/.

Kamber:1975:SSW

[KT75j] Prof. Franz W. Kamber and Prof. Philippe Tondeur. Semisimplicial Weil algebras. Lecture Notes in Mathematics, 493:186– 201, 1975. CODEN LNMAA2. ISBN 3-540-07420-1 (print), 3-540-37956-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081566/.

Korolyuk:1976:AES

[KT76] V. S. Korolyuk and A. F. Turbin. Asymptotic enlarging of semi-Markov processes with an arbitrary state space. Lecture Notes in Mathematics, 550:297–315, 1976. CODEN LNMAA2. ISBN 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077498/.

Koh:1979:ITC

[KTR79] K. M. Koh, T. Tan, and D. G. Rogers. Interlaced trees: a class of graceful trees. Lecture Notes in Mathematics, 748:65–78, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0102685/.

Kucera:1975:FST

[Kuč75] Jan Kučera. Fractional spaces of temperate distribution. Lecture Notes in Mathematics, 457:249–255, 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067109/.

Kudla:1977:RBA

[Kud77] Stephen S. Kudla. Relations between automorphic forms produced by theta-functions. Lecture Notes in Mathematics, 627: 277–285, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065305/.

Kueker:1975:BFA

[Kue75a] David W. Kueker. Back-and-forth arguments and infinitary logics. Lecture Notes in Mathematics, 492:17–71, 1975. CODEN LNMAA2. ISBN 3-540-07419-8 (print), 3-540-37949-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081120/.

Kueker:1975:ILM

[Kue75b] Prof. David W. Kueker. Infinitary Logic: In Memoriam Carol Karp: A Collection of Papers by Various Authors, volume 492 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07419-8 (print), 3-540-37949-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 206 pp. LCCN QA3

.L28 no. 492. URL http://link.springer.com/book/10.1007/
BFb0081118;   http://www.springerlink.com/content/978-
3-540-37949-2.

Kuelbs:1976:LILb

[Kue76a] J. Kuelbs. The law of the iterated logarithm and related strong convergence theorems for Banach space valued random variables. Lecture Notes in Mathematics, 539:224–314, 1976. CODEN LN-MAA2. ISBN 3-540-07858-4 (print), 3-540-38119-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079698/.

Kuelbs:1976:LILa

[Kue76b] J. Kuelbs. The law of the iterated logarithm for Banach space valued random variables. Lecture Notes in Mathematics, 526:131– 142, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082348/.

Kuelbs:1978:SEM

[Kue78] J. Kuelbs. Some exponential moments with applications to density estimation, the empirical distribution function, and lacunary series. Lecture Notes in Mathematics, 644:255–269, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066850/.

Kuelbs:1979:RGB

[Kue79] J. Kuelbs. Rates of growth for Banach space valued independent increment processes. Lecture Notes in Mathematics, 709: 151–169, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071956/.

Kuhnau:1979:GTP

[Küh79] Reiner Kühnau. Gauss–Thomsonsches Prinzip minimaler Energie, verallgemeinerte transfinite Durchmesser und quasikonforme Abbildungen. (German) []. Lecture Notes in Mathematics, 743:140– 164, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079489/.

Kuiper:1979:SPF

[Kui79] Nicolaas H. Kuiper. Sphères Polyédriques flexibles dans E^{3}, d’après Robert connelly. (French) []. Lecture Notes in Mathematics, 710:147–168, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069977/.

Kuku:1976:OFR

[Kuk76] A. O. Kuku. SK_{n} of orders and G_{n} of finite rings. Lecture Notes in Mathematics, 551:60–68, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079994/.

Kunita:1976:ACM

[Kun76] Hiroshi Kunita. Absolute continuity of Markov processes. Lecture Notes in Mathematics, 511:44–77, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101096/.

Kuo:1975:EOG

[Kuo75a] Prof. Hui-Hsiung Kuo. Equivalence and orthogonality of Gaussian measures. Lecture Notes in Mathematics, 463:110–152, 1975. CODEN LNMAA2. ISBN 3-540-07173-3 (print), 3-540-37508-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082009/.

Kuo:1975:GMBa

[Kuo75b] Prof. Hui-Hsiung Kuo. Gaussian measures in Banach spaces. Lecture Notes in Mathematics, 463:1–109, 1975. CODEN LN-MAA2. ISBN 3-540-07173-3 (print), 3-540-37508-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082008/.

Kuo:1975:GMBb

[Kuo75c] Prof. Hui-Hsiung Kuo. Gaussian Measures in Banach Spaces, volume 463 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07173-3 (print), 3-540-37508-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 224 pp. LCCN QA3 .L28 no. 463. URL http:

//link.springer.com/book/10.1007/BFb0082007;   http://
www.springerlink.com/content/978-3-540-37508-1.

Kuo:1975:SRA

[Kuo75d] Prof. Hui-Hsiung Kuo. Some results about abstract Wiener space. Lecture Notes in Mathematics, 463:153–217, 1975. CODEN LN-MAA2. ISBN 3-540-07173-3 (print), 3-540-37508-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082010/.

Kuo:1976:DTB

[Kuo76] Hui-Hsiung Kuo. Distribution theory on Banach space. Lecture Notes in Mathematics, 526:143–156, 1976. CODEN LN-MAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082349/.

Kuo:1978:DCM

[Kuo78] Hui-Hsiung Kuo. Differential calculus for measures on Banach spaces. Lecture Notes in Mathematics, 644:270–285, 1978. CO-DEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0066851/.

Kupisch:1975:QFA

[Kup75] H. Kupisch. Quasi-Frobenius-algebras of finite representation type. Lecture Notes in Mathematics, 488:184–200, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081224/.

Kupradze:1976:LEN

[Kup76] V. D. Kupradze. Über die Lösung einiger nichtklassischer Probleme der Elastizitätstheorie. (German) []. Lecture Notes in Mathematics, 561:68–72, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087626/.

Kuroda:1975:STD

[Kur75] S. T. Kuroda. Scattering theory for differential operators, III; exterior problems. Lecture Notes in Mathematics, 448:227–241, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067088/.

Kurdyumov:1978:ATM

[Kur78] G. L. Kurdyumov. An algorithm-theoretic method in studying homogeneous random networks. Lecture Notes in Mathematics, 653:37–55, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070083/.

Kashiwara:1977:RCC

[KV77] Masaki Kashiwara and Michèle Vergne. Remarque sur la covariance de certains opérateurs différentiels. (French) []. Lecture Notes in Mathematics, 587:119–137, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087918/.

Kashiwara:1979:FSB

[KV79a] M. Kashiwara and M. Vergne. Functions on the Shilov boundary of the generalized half plane. Lecture Notes in Mathematics, 728:136–176, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063342/.

Kashiwara:1979:TSS

[KV79b] M. Kashiwara and M. Vergne. K-types and singular spectrum. Lecture Notes in Mathematics, 728:177–200, 1979. CODEN LN-MAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063343/.

Kupka:1975:FDF

[KvQ75] Ivan Kupka and Ngô van Quê. Formes différentielles fermées non singulières. (French) []. Lecture Notes in Mathematics, 484:239– 256, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082156/.

Kervaire:1978:SMK

[KW78] M. Kervaire and C. Weber. A survey of multidimensional knots. Lecture Notes in Mathematics, 685:61–134, 1978. CODEN LN-MAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062969/.

Kwapien:1976:TRS

[Kwa76] S. Kwapień. A theorem on the Rademacher series with vector valued coefficients. Lecture Notes in Mathematics, 526:157–158, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082350.

Kaplan:1979:CBM

[KY79a] James L. Kaplan and James A. Yorke. Chaotic behavior of multidimensional difference equations. Lecture Notes in Mathematics, 730:204–227, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064319/.

Kaplan:1979:NSG

[KY79b] James L. Kaplan and James A. Yorke. Numerical solution of a generalized eigenvalue problem for even mappings. Lecture Notes in Mathematics, 730:228–237, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064320/.

Kirwan:1976:ACF

[KZ76] William E. Kirwan and Lawrence Zalcman, editors. Advances in Complex Function Theory: Proceedings of Seminars Held at Maryland University, 1973/74, volume 505 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 505. URL http:

//link.springer.com/book/10.1007/BFb0081095;   http://
www.springerlink.com/content/978-3-540-38088-7.

Knapp:1977:CTR

[KZ77] A. W. Knapp and Gregg Zuckerman. Classification theorems for representations of semisimple Lie groups. Lecture Notes in Mathematics, 587:138–159, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087919/.

LaFollette:1976:SAA

[La 76] Philip La Follette. Splitting of Azumaya algebras over number rings. Lecture Notes in Mathematics, 549:100–103, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077337/.

Lachlan:1976:DTT

[Lac76] A. H. Lachlan. Dimension and totally transcendental theories of rank 2. Lecture Notes in Mathematics, 537:153–183, 1976. CO-DEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096900/.

Lacey:1977:LUS

[Lac77] H. Elton Lacey. Local unconditional structure in Banach spaces. Lecture Notes in Mathematics, 604:44–56, 1977. CODEN LN-MAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069205/.

Lacroix:1978:EEI

[Lac78] Marie-Thérèse Lacroix. Échelle d’espaces intermédiaires entre un espace de Sobolev–Orlicz et un espace d’Orlicz tracé d’espaces de Sobolev–Orlicz avec poids. (French) []. Lecture Notes in Mathematics, 665:121–139, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061802/.

Lacroix:1979:RFP

[Lac79] Jean Lacroix. Représentation des fonctions propres sur un espace symétrique. (French) []. Lecture Notes in Mathematics, 713: 137–144, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071324/.

Ladogorski:1975:LTT

[Ład75] C. Ładogórski. A limit theorem for triangular arrays of representations of canonical anti-commutation relations. Lecture Notes in Mathematics, 472:95–106, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081948/.

Lada:1976:EA

[Lad76a] Thomas J. Lada. Errata and addenda to [A], [G], and [G′]. Lecture Notes in Mathematics, 533:484–490, 1976. CODEN LN-MAA2. ISBN 3-540-07984-X (print), 3-540-37985-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080470/.

Lada:1976:SHA

[Lad76b] Thomas J. Lada. Strong homotopy algebras over monads. Lecture Notes in Mathematics, 533:399–479, 1976. CODEN LN-MAA2. ISBN 3-540-07984-X (print), 3-540-37985-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080469/.

Lady:1977:CTF

[Lad77] E. L. Lady. On classifying torsion free modules over discrete valuation rings. Lecture Notes in Mathematics, 616:168–172, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068195/.

Lai:1978:EDP

[Lai78] Pham The Lai, editor. Equations aux Dérivées Partielles: Proceedings, Saint-Jean-de-Monts, June 1–4, 1977, volume 660 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 660. URL http://link.springer.com/book/10.1007/
BFb0082288;   http://www.springerlink.com/content/978-
3-540-35764-3.

Lakshmikantham:1979:ECR

[Lak79a] V. Lakshmikantham. Existence and comparison results for Volterra integral equations in a Banach space. Lecture Notes in Mathematics, 737:120–126, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064502/.

Laksov:1979:DT

[Lak79b] Dan Laksov. Deformation and transversality. Lecture Notes in Mathematics, 732:300–316, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066650/.

Lallement:1977:PMP

[Lal77a] Gérard Lallement. Présentations de monoïdes et problémes algorithmiques. (French) []. Lecture Notes in Mathematics, 586: 136–144, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087129/.

Lallement:1977:PAM

[Lal77b] Gérard Lallement. Sur les produits amalgames de monoïdes. (French) []. Lecture Notes in Mathematics, 586:25–33, 1977. CO-DEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087118/.

Lamoureux:1975:NBL

[Lam75] C. Lamoureux. Non-bounded leaves in codimension one foliations. Lecture Notes in Mathematics, 484:257–272, 1975. CODEN LN-MAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082157/.

Lamb:1976:BTT

[Lam76] George L. Lamb, Jr. Bäcklund transformations at the turn of the century. Lecture Notes in Mathematics, 515:69–79, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081163/.

Lam:1978:CRS

[Lam78a] T. Y. Lam. The ‘classical’ results on Serre’s conjecture. Lecture Notes in Mathematics, 635:45–78, 1978. CODEN LNMAA2. ISBN 3-540-08657-9 (print), 3-540-35926-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068342/.

Lam:1978:F

[Lam78b] T. Y. Lam. Foundations. Lecture Notes in Mathematics, 635: 1–44, 1978. CODEN LNMAA2. ISBN 3-540-08657-9 (print), 3-540-35926-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068341/.

Lam:1978:HT

[Lam78c] T. Y. Lam. Horrocks’ theorem. Lecture Notes in Mathematics, 635:97–121, 1978. CODEN LNMAA2. ISBN 3-540-08657-9 (print), 3-540-35926-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068344/.

Lam:1978:QAS

[Lam78d] T. Y. Lam. The quadratic analogue of Serre’s conjecture. Lecture Notes in Mathematics, 635:167–204, 1978. CODEN LN-MAA2. ISBN 3-540-08657-9 (print), 3-540-35926-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068346/.

Lam:1978:QM

[Lam78e] T. Y. Lam. Quillen’s method. Lecture Notes in Mathematics, 635: 123–165, 1978. CODEN LNMAA2. ISBN 3-540-08657-9 (print), 3-540-35926-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068345/.

Lam:1978:SC

[Lam78f] T. Y. Lam. Serre’s Conjecture, volume 635 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08657-9 (print), 3-540-35926-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 204 (est.) pp. LCCN QA3

.L28 no. 635. URL http://link.springer.com/book/10.1007/
BFb0068340;   http://www.springerlink.com/content/978-
3-540-35926-5.

Lam:1978:TEP

[Lam78g] T. Y. Lam. Two elementary proofs of Serre’s conjecture. Lecture Notes in Mathematics, 635:79–95, 1978. CODEN LNMAA2. ISBN 3-540-08657-9 (print), 3-540-35926-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068343/.

Lambert:1978:FFN

[Lam78h] J. D. Lambert. Frequency fitting in the numerical solution of ordinary differential equations. Lecture Notes in Mathematics, 679: 65–72, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067867/.

Langlands:1976:A

[Lan76a] Robert P. Langlands. The assumptions. Lecture Notes in Mathematics, 544:3–36, 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079931/.

Langlands:1976:CF

[Lan76b] Robert P. Langlands. Cusp forms. Lecture Notes in Mathematics, 544:37–60, 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079932/.

Langlands:1976:ES

[Lan76c] Robert P. Langlands. Eisenstein series. Lecture Notes in Mathematics, 544:61–99, 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079933/.

Langlands:1976:I

[Lan76d] Robert P. Langlands. Introduction. Lecture Notes in Mathematics, 544:1–2, 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079930.

Langlands:1976:MT

[Lan76e] Robert P. Langlands. The main theorem. Lecture Notes in Mathematics, 544:161–233, 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079936/.

Langlands:1976:ML

[Lan76f] Robert P. Langlands. Miscellaneous lemmas. Lecture Notes in Mathematics, 544:100–118, 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079934/.

Langlands:1976:FES

[Lan76g] Robert P. Langlands. On the Functional Equations Satisfied by Eisenstein Series, volume 544 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 337 pp. LCCN QA3

.L28 no. 544. URL http://link.springer.com/book/10.1007/
BFb0079929;   http://www.springerlink.com/content/978-
3-540-38070-2.

Langlands:1976:SFE

[Lan76h] Robert P. Langlands. Some functional equations. Lecture Notes in Mathematics, 544:119–160, 1976. CODEN LNMAA2. ISBN 3-540-07872-X (print), 3-540-38070-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079935/.

Lanford:1977:ALV

[Lan77] Oscar E. Lanford, III. Appendix to lecture VII: Computer pictures of the Lorenz attractor. Lecture Notes in Mathematics, 615: 113–116, 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068364/.

Lang:1978:CBS

[Lan78] Serge Lang. Sur la conjecture de Birch–Swinnerton-Dyer [d’après J. Coates et A. Wiles]. (French) []. Lecture Notes in Mathematics, 677:189–200, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070763/.

Landweber:1979:NAC

[Lan79a] Peter S. Landweber. New applications of commutative algebra to Brown–Peterson homology. Lecture Notes in Mathematics, 741: 449–460, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062152/.

Landweber:1979:SSS

[Lan79b] Peter S. Landweber. The signature of symplectic and selfconjugate manifolds. Lecture Notes in Mathematics, 741:461– 472, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062153/.

Lapreste:1979:CPM

[Lap79] Jean-Thierry Lapresté. Charges, poids et mesures de Levy dans les espaces vectoriels localement convexes. (French) [loads, weights, and Levy measures in locally-convex vector spaces]. Lecture Notes in Mathematics, 721:41–71, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070850/.

Larmore:1979:ICS

[Lar79] Lawrence L. Larmore. Isotopy classification of spheres in a manifold. Lecture Notes in Mathematics, 763:135–140, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088083/.

Lascar:1975:OPD

[Las75] B. Lascar. Opérateurs pseudo-différentiels d’une infinité de variables. (French) []. Lecture Notes in Mathematics, 474:83–90, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077400/.

Lascoux:1976:ICS

[Las76a] Alain Lascoux. Intersection de cycles de Schubert. (French) [Intersection of Schubert cycles]. Lecture Notes in Mathematics, 535: 177–186, 1976. CODEN LNMAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080500/.

Lassalle:1976:TZD

[Las76b] Guy Lassalle. Sur le théorème des zéros différentiables. (French) []. Lecture Notes in Mathematics, 535:70–97, 1976. CODEN LN-MAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080497/.

Lascoux:1977:CSE

[Las77] A. Lascoux. Calcul de Schur et extensions grassmanniennes des λ-anneaux. (French) []. Lecture Notes in Mathematics, 579:182– 216, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090018/.

Lascar:1978:EDP

[Las78] Bernard Lascar. Équations aux dérivées partielles en dimension infinie. (French) []. Lecture Notes in Mathematics, 644:286– 313, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066852/.

Lashof:1979:SS

[Las79] Richard Lashof. Stable G-smoothing. Lecture Notes in Mathematics, 741:283–306, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062145/.

Latour:1979:DSS

[Lat79] François Latour. Double suspension d’une sphère d’homologie [d’après R. Edwards]. (French) []. Lecture Notes in Mathematics, 710:169–186, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069978/.

Laurent:1975:AFC

[Lau75a] Andre G. Laurent. Applications of fractional calculus to spherical (radial) probability models and generalizations. Lecture Notes in Mathematics, 457:256–266, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067110/.

Laurent:1975:PH

[Lau75b] Andre G. Laurent. A problem of hyperstereology. Lecture Notes in Mathematics, 457:267–271, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067111/.

Laurent:1975:ADP

[Lau75c] P. J. Laurent. Un algorithme dual pour le calcul de la distance entre deux convexes. (French) []. Lecture Notes in Mathematics, 477:202–228, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079178/.

Laursen:1976:SRA

[Lau76] K. B. Laursen. Some remarks on automatic continuity. Lecture Notes in Mathematics, 512:96–108, 1976. CODEN LN-MAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080027/.

Lausch:1977:RCG

[Lau77] Hans Lausch. Relative cohomology of groups. Lecture Notes in Mathematics, 573:66–72, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087813/.

Laudal:1978:GTL

[Lau78] O. A. Laudal. A generalized trisecant lemma. Lecture Notes in Mathematics, 687:112–149, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062930/.

Laudal:1979:FMA

[Lau79a] Olav Arnfinn Laudal. Formal Moduli of Algebraic Structures, volume 754 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09702-3 (print), 3-540-38532-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 161 pp. LCCN QA3 .L28 no. 754. URL http:

//link.springer.com/book/10.1007/BFb0065055;   http://
www.springerlink.com/content/978-3-540-38532-5.

Laudal:1979:GC

[Lau79b] Olav Arnfinn Laudal. Global cohomology. Lecture Notes in Mathematics, 754:48–77, 1979. CODEN LNMAA2. ISBN 3-540-09702-3 (print), 3-540-38532-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065059/.

Laudal:1979:GOT

[Lau79c] Olav Arnfinn Laudal. Global obstruction theory and formal moduli. Lecture Notes in Mathematics, 754:78–124, 1979. CODEN LNMAA2. ISBN 3-540-09702-3 (print), 3-540-38532-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065060/.

Laudal:1979:I

[Lau79d] Olav Arnfinn Laudal. Introduction. Lecture Notes in Mathematics, 754:1–5, 1979. CODEN LNMAA2. ISBN 3-540-09702-3 (print), 3-540-38532-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065056/.

Laudal:1979:LAM

[Lau79e] Olav Arnfinn Laudal. Lifting algebras and morphisms of algebras. Lecture Notes in Mathematics, 754:17–47, 1979. CODEN LNMAA2. ISBN 3-540-09702-3 (print), 3-540-38532-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065058/.

Laudal:1979:SF

[Lau79f] Olav Arnfinn Laudal. Sections of functors. Lecture Notes in Mathematics, 754:6–16, 1979. CODEN LNMAA2. ISBN 3-540-09702-3 (print), 3-540-38532-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065057/.

Laudal:1979:SA

[Lau79g] Olav Arnfinn Laudal. Some applications. Lecture Notes in Mathematics, 754:125–151, 1979. CODEN LNMAA2. ISBN 3-540-09702-3 (print), 3-540-38532-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065061/.

Lawrence:1975:EGS

[Law75a] Jim Lawrence. k-equilateral (2k + 1)-gons span only evendimensional spaces. Lecture Notes in Mathematics, 490:185– 186, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081140.

Lawvere:1975:IP

[Law75b] F. William Lawvere. Introduction to Part I. Lecture Notes in Mathematics, 445:3–14, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061291/.

Lawson:1976:DDM

[Law76] Charles L. Lawson. On the discovery and description of mathematical programming algorithms. Lecture Notes in Mathematics, 506:157–165, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080122/.

Lawrynowicz:1979:PBC

[Ław79] Julian Ławrynowicz. Parametrization and boundary correspondence for Teichmüller mappings in an annulus. Lecture Notes in Mathematics, 743:165–183, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079490/.

Lazard:1975:CFG

[Laz75a] Prof. Dr. Michel Lazard. Commutative Formal Groups, volume 443 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 236 pp. LCCN QA3 .L28 no. 443. URL http:

//link.springer.com/book/10.1007/BFb0070554;   http://
www.springerlink.com/content/978-3-540-37431-2.

Lazard:1975:ELS

[Laz75b] Prof. Dr. Michel Lazard. Extending and lifting some formal groups. Lecture Notes in Mathematics, 443:199–230, 1975. CO-DEN LNMAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070561/.

Lazard:1975:FGB

[Laz75c] Prof. Dr. Michel Lazard. Formal groups and buds. Lecture Notes in Mathematics, 443:31–57, 1975. CODEN LNMAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070556/.

Lazard:1975:FV

[Laz75d] Prof. Dr. Michel Lazard. Formal varieties. Lecture Notes in Mathematics, 443:8–31, 1975. CODEN LNMAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070555/.

Lazard:1975:GEC

[Laz75e] Prof. Dr. Michel Lazard. The general equivalence of categories. Lecture Notes in Mathematics, 443:57–91, 1975. CODEN LN-MAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070557/.

Lazard:1975:FGC

[Laz75f] Prof. Dr. Michel Lazard. On formal groups in characteristic p. Lecture Notes in Mathematics, 443:163–199, 1975. CODEN LN-MAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070560/.

Lazard:1975:SEC

[Laz75g] Prof. Dr. Michel Lazard. The special equivalences of categories. Lecture Notes in Mathematics, 443:92–118, 1975. CODEN LN-MAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070558/.

Lazard:1975:STC

[Laz75h] Prof. Dr. Michel Lazard. The structure theorem and its consequences. Lecture Notes in Mathematics, 443:118–162, 1975. CO-DEN LNMAA2. ISBN 3-540-07145-8 (print), 3-540-37431-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0070559/.

Levy-Bruhl-Laperriere:1978:SDR

[LBL78] Anne Lévy-Bruhl-Laperrière. Spectre du de Rham Hodge sur l’espace projectif complexe. (French) []. Lecture Notes in Mathematics, 641:163–188, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064846/.

Louvet:1977:CCF

[LD77] Pierre Louvet and Jean Durivault. Compressible countercurrent flow in a strongly rotating cylinder. Lecture Notes in Mathematics, 594:312–333, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086095/.

Lins:1977:LE

[LdMP77] A. Lins, W. de Melo, and C. C. Pugh. On Liénard’s equation. Lecture Notes in Mathematics, 597:335–357, 1977. CODEN LN-MAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085364/.

LePotier:1975:PML

[Le 75] Joseph Le Potier. Le problème des modules locaux pour les espaces C-analytiques compacts [d’après A. Douady et J. Hubbard]. (French) []. Lecture Notes in Mathematics, 431:255–272, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066374/.

LeJan:1976:MAF

[Le 76] Yves Le Jan. Mesures associées à une forme de Dirichlet. (French) []. Lecture Notes in Mathematics, 563:179–192, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087578/.

LePotier:1977:FVC

[Le 77] Joseph Le Potier. Fibrés vectoriels et cycles d’ordre fini sur une variété algébrique non compacte [d’après M. Cornalba et P. Griffiths]. (French) []. Lecture Notes in Mathematics, 567:38–53, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096060/.

LeBarz:1978:GEP

[Le 78] Patrick Le Barz. Géométrie énumerative pour les multisécantes. (French) []. Lecture Notes in Mathematics, 683:116–167, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063176/.

LeJan:1979:PFT

[Le 79] Yves Le Jan. Problèmes frontières en théorie du potentiel. (French) []. Lecture Notes in Mathematics, 713:145–155, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071325/.

Lopez-Escobar:1975:I

[LE75] E. G. K. López-Escobar. Introduction. Lecture Notes in Mathematics, 492:1–16, 1975. CODEN LNMAA2. ISBN 3-540-07419-8 (print), 3-540-37949-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081119/.

Ledrappier:1976:CBF

[Led76] F. Ledrappier. Sur la condition de Bernoulli faible et ses applications. (French) []. Lecture Notes in Mathematics, 532:152– 159, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080178/.

Ledrappier:1979:VPT

[Led79] F. Ledrappier. A variational principle for the topological conditional entropy. Lecture Notes in Mathematics, 729:78–88, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063285/.

Leger:1975:SAC

[Leg75] Christian Leger. Les systèmes d’Asada à coefficients constants. (French) []. Lecture Notes in Mathematics, 482:266–273, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080250/.

Lehrer:1976:SIS

[Leh76] G. I. Lehrer. Some incidence structures of maximal rank. Lecture Notes in Mathematics, 560:132–135, 1976. CODEN LN-MAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097374/.

Lehto:1979:BVP

[Leh79] Olli Lehto. On the boundary value problem for quasiconformal mappings. Lecture Notes in Mathematics, 743:184–196, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079491/.

Leivant:1975:SNA

[Lei75] Daniel Leivant. Strong normalization for arithmetic. Lecture Notes in Mathematics, 500:182–197, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079552/.

Leitman:1979:PWP

[Lei79] M. J. Leitman. On plane waves propagating into a random fluid: Asymptotic behavior. Lecture Notes in Mathematics, 737:127– 136, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064503/.

Lelong:1975:TSV

[Lel75a] Pierre Lelong. Topologies semi-vectorielles et topologies pseudoconvexes sur un espace vectoriel complexe. (French) []. Lecture Notes in Mathematics, 474:1–15, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077396/.

Lelong:1975:SPL

[Lel75b] Prof. Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1973/74, volume 474 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 474. URL http:

//link.springer.com/book/10.1007/BFb0077395;   http://
www.springerlink.com/content/978-3-540-37554-8.

Lelong:1976:SPL

[Lel76] Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1974/75, volume 524 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 524. URL http:

//link.springer.com/book/10.1007/BFb0077993;   http://
www.springerlink.com/content/978-3-540-38251-5.

Lelong:1977:SPL

[Lel77a] Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1975/76: Journées sur les Fonctions Analytiques, Toulouse, 1976, volume 578 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 578. URL http:

//link.springer.com/book/10.1007/BFb0091458;   http://
www.springerlink.com/content/978-3-540-37341-4.

Lelong:1977:SCP

[Lel77b] Pierre Lelong. Sur la structure des courants positifs fermés. (French) []. Lecture Notes in Mathematics, 578:136–156, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091466/.

Lelong:1978:APF

[Lel78a] P. Lelong. Application des propriétés des fonctions plurisousharmoniques à un problème de mesure dans les espaces vectoriels complexes. (French) []. Lecture Notes in Mathematics, 645:131– 135, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069670/.

Lelong:1978:TFI

[Lel78b] Pierre Lelong. Un théorème de fonctions inverses dans les espaces vectoriels topologiques complexes et ses applications à des problèmes de croissance en analyse complexe. (French) []. Lecture Notes in Mathematics, 694:172–195, 1978. CODEN LN-MAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063248/.

Lemaire:1977:TEF

[Lem77] J.-M. Lemaire. Le transfert dans les espaces fibrés [d’après J. Becker et D. Gottlieb]. (French) []. Lecture Notes in Mathematics, 567:23–37, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096059/.

Lemarechal:1979:OND

[Lem79] C. Lemarechal. Optimisation non différentiable: Méthodes de faisceaux. (French) []. Lecture Notes in Mathematics, 704:53–61, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063614/.

Lenagan:1976:AQR

[Len76a] T. H. Lenagan. Artinian quotient rings of Macaulay rings. Lecture Notes in Mathematics, 545:140–150, 1976. CODEN LN-MAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080309/.

Lenstra:1976:GFC

[Len76b] H. W. Lenstra, Jr. K_{2} of a global field consists of symbols. Lecture Notes in Mathematics, 551:69–73, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079995/.

Lenglart:1978:LIS

[Len78] E. Lenglart. Sur la localisation des intégrales stochastiques. (French) []. Lecture Notes in Mathematics, 649:53–56, 1978. CO-DEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0064593/.

LePage:1978:MEA

[LeP78a] Raoul LePage. A maximal equality and its application in vector spaces. Lecture Notes in Mathematics, 645:136–146, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069671/.

Lepingle:1978:CCT

[Lép78b] D. Lépingle. Sur certains commutateurs de la théorie des martingales. (French) []. Lecture Notes in Mathematics, 649:138– 147, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064602/.

Lepingle:1978:CAM

[Lép78c] D. Lépingle. Sur le comportement asymptotique des martingales locales. (French) []. Lecture Notes in Mathematics, 649: 148–161, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064603/.

Lepingle:1978:IMF

[Lép78d] D. Lépingle. Une inégalité de martingales. (French) []. Lecture Notes in Mathematics, 649:134–137, 1978. CODEN LN-MAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064601/.

Leray:1975:SAG

[Ler75] Jean Leray. Solutions asymptotiques et groupe symplectique. (French) []. Lecture Notes in Mathematics, 459:73–97, 1975. CO-DEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074193/.

Leray:1978:ALM

[Ler78] Jean Leray. Analyse lagrangienne et mécanique quantique. (French) [Lagrangian analysis and quantum mechanics]. Lecture Notes in Mathematics, 660:128–129, 1978. CODEN LN-MAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082299.

Lesieur:1978:CNA

[Les78] Léonce Lesieur. Conditions noetheriennes dans l’anneau de polynômes de Ore A[X, σ, δ]. (French) []. Lecture Notes in Mathematics, 641:220–234, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064850/.

Letac:1976:PCM

[Let76] Gérard Letac. Produits de chaines de Markov. (French) []. Lecture Notes in Mathematics, 532:160–188, 1976. CODEN LN-MAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080179/.

Letta:1979:QFL

[Let79] G. Letta. Quasimartingales et formes linéaires associées. (French) []. Lecture Notes in Mathematics, 721:216–226, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070862/.

Lopez-Escobar:1975:ICR

[LEV75] E. G. K. López-Escobar and W. Veldman. Intuitionistic completeness of a restricted second-order logic. Lecture Notes in Mathematics, 500:198–232, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079553/.

Levy:1977:ECI

[Lév77] Th. Lévy. Équations et conditions d’interface pour des phenomènes acoustiques dans des milieux poreux. (French) []. Lecture Notes in Mathematics, 594:301–311, 1977. CODEN LN-MAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086094/.

Levine:1978:SRH

[Lev78] J. Levine. Some results on higher dimensional knot groups. Lecture Notes in Mathematics, 685:243–269, 1978. CODEN LN-MAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062975/.

Levin:1979:ABS

[Lev79a] J. J. Levin. On the asymptotic behavior of solutions of integral equations. Lecture Notes in Mathematics, 737:137–148, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064504/.

Levy:1979:MCG

[Lev79b] Lawrence S. Levy. Modules over the cyclic group of prime order. Lecture Notes in Mathematics, 734:207–222, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103160/.

Lewis:1977:DTP

[Lew77] D. R. Lewis. Duals of tensor products. Lecture Notes in Mathematics, 604:57–66, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069206/.

Lelong-Ferrand:1979:RCM

[LF79] Jacqueline Lelong-Ferrand. Regularity of conformal mappings of Riemannian manifolds. Lecture Notes in Mathematics, 743:197– 203, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079492/.

Li:1979:HWJ

[Li79] Wen-Ch’ing Winnie Li. Hecke–Weil–Jacquet–Langlands theorem revisited. Lecture Notes in Mathematics, 751:206–220, 1979. CO-DEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062710/.

Lichnerowicz:1977:NGD

[Lic77] André Lichnerowicz. New geometrical dynamics. Lecture Notes in Mathematics, 570:377–394, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087793/.

Lickorish:1979:SSK

[Lic79] W. B. Raymond Lickorish. Shake–slice knots. Lecture Notes in Mathematics, 722:67–70, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063190/.

Liebert:1977:UVC

[Lie77] Wolfgang Liebert. Ulm valuations and co-valuations on torsioncomplete p-groups. Lecture Notes in Mathematics, 616:337– 353, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068205/.

Lieberman:1978:CCS

[Lie78] David I. Lieberman. Compactness of the Chow scheme: Applications to automorphisms and deformations of Kähler manifolds. Lecture Notes in Mathematics, 670:140–186, 1978. CODEN LN-MAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064399/.

Lieberman:1979:PBT

[Lie79] David I. Lieberman. A Poincaré–Bendixson theorem for compact Kähler manifolds. Lecture Notes in Mathematics, 712:60–65, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062815/.

Liggett:1977:SEI

[Lig77a] T. M. Liggett. The stochastic evolution of infinite systems of interacting particles. Lecture Notes in Mathematics, 598:187–248, 1977. CODEN LNMAA2. ISBN 3-540-08340-5 (print), 3-540-37307-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097493/.

Ligozat:1977:CMN

[Lig77b] Gérard Ligozat. Courbes modulaires de niveau 11. (French) []. Lecture Notes in Mathematics, 601:149–237, 1977. CODEN LN-MAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063948/.

Lin:1979:WMS

[Lin79a] Michael Lin. Weak mixing for semi-groups of Markov operators without finite invariant measures. Lecture Notes in Mathematics, 729:89–92, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063286/.

Lind:1979:EGA

[Lin79b] D. A. Lind. Ergodic group automorphisms and specification. Lecture Notes in Mathematics, 729:93–104, 1979. CODEN LN-MAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063287/.

Lions:1975:FSP

[Lio75a] J.-L. Lions. On free surface problems: Methods of variational and quasi variational inequalities. Lecture Notes in Mathematics, 461:129–148, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074152/.

Lions:1975:RSN

[Lio75b] J.-L. Lions. Remarks on some new nonlinear boundary value problems. Lecture Notes in Mathematics, 446:301–328, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070608/.

Lions:1976:ABS

[Lio76] J.-L. Lions. Asymptotic behaviour of solutions of variational inequalities with highly oscillating coefficients. Lecture Notes in Mathematics, 503:30–55, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088745/.

Lion:1977:IEG

[Lio77a] Gérard Lion. Intégrales d’Entrelacement sur des groupes de Lie nilpotents et indices de Maslov. (French) []. Lecture Notes in Mathematics, 587:160–176, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087920/.

Lions:1977:APF

[Lio77b] J.-L. Lions. Sur l’Approximation de problèmes à frontière libre dans les matériaux inhomogènes. (French) []. Lecture Notes in Mathematics, 606:194–203, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064463/.

Little:1975:CAC

[Lit75] Charles H. C. Little. A conjecture about circuits in planar graphs. Lecture Notes in Mathematics, 452:171–175, 1975. CODEN LN-MAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069555/.

Little:1976:TPG

[Lit76] Charles H. C. Little. A theorem on planar graphs. Lecture Notes in Mathematics, 560:136–141, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097375/.

Little:1977:CMV

[Lit77a] Charles H. C. Little, editor. Combinatorial Mathematics V: Proceedings of the Fifth Australian Conference, Held at the Royal Melbourne Institute of Technology, August 24–26, 1976, volume 622 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 622. URL http://link.springer.com/book/10.1007/
BFb0069176;   http://www.springerlink.com/content/978-
3-540-37020-8.

Little:1977:RCP

[Lit77b] Charles H. C. Little. On rings of circuits in planar graphs. Lecture Notes in Mathematics, 622:133–140, 1977. CODEN LN-MAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069187/.

Litherland:1979:SIT

[Lit79] R. A. Litherland. Signatures of iterated torus knots. Lecture Notes in Mathematics, 722:71–84, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063191/.

Liulevicius:1978:FMH

[Liu78a] Arunas Liulevicius. Flag manifolds and homotopy rigidity of linear actions. Lecture Notes in Mathematics, 673:254–261, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064701/.

Liulevicius:1978:LBC

[Liu78b] Arunas Liulevicius. Line bundles, cohomology automorphisms, and homotopy rigidity of linear actions. Lecture Notes in Mathematics, 658:224–233, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068720/.

Liulevicius:1979:HRS

[Liu79a] Arunas Liulevicius. Homotopy rigidity of sturdy spaces. Lecture Notes in Mathematics, 763:141–151, 1979. CODEN LN-MAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088084/.

Liulevicius:1979:LAF

[Liu79b] Arunas Liulevicius. Linear actions on friendly spaces. Lecture Notes in Mathematics, 741:307–315, 1979. CODEN LN-MAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062146/.

Lacroix:1976:LQF

[LL76] J. Lacroix and E. Le Page. Limites de quotients de fonctions harmoniques et espaces de Hardy associés à une marche aléatoire sur un groupe abélien. (French) []. Lecture Notes in Mathematics, 563:161–178, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087577/.

Laudal:1978:DCM

[LL78] O. A. Laudal and K. Lønsted. Deformations of curves I. moduli for hyperelliptic curves. Lecture Notes in Mathematics, 687: 150–167, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062931/.

Laine:1979:CAJ

[LLS79] Ilpo Laine, Olli Lehto, and Tuomas Sorvali, editors. Complex Analysis Joensuu 1978: Proceedings of the Colloquium on Complex Analysis, Joensuu, Finland, August 24–27, 1978, volume 747 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09553-5 (print), 3-540-34859-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 747. URL http://link.springer.com/book/10.1007/
BFb0063954;   http://www.springerlink.com/content/978-
3-540-34859-7.

Laitinen:1979:TCS

[LM79] Erkki Laitinen and Ib Madsen. Topological classifications of Sℓ_{2}(F_{p}) space forms. Lecture Notes in Mathematics, 763:235– 261, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088089/.

Lawvere:1975:MTT

[LMW75] Dr. F. William Lawvere, Dr. Christian Maurer, and Dr. Gavin C. Wraith. Model Theory and Topoi: A Collection of Lectures by Various Authors, volume 445 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 354 (est.) pp. LCCN QA3

.L28 no. 445. URL http://link.springer.com/book/10.1007/
BFb0061290;   http://www.springerlink.com/content/978-
3-540-37495-4.

Loday:1976:MMG

[Lod76] Jean-Louis Loday. Les matrices monomiales et le groupe de Whitehead ω h_{2}. (French) []. Lecture Notes in Mathematics, 551:155– 163, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079998/.

Loday:1979:HEC

[Lod79] Jean-Louis Loday. Homotopie des espaces de concordances [d’après F. Waldhausen]. (French) []. Lecture Notes in Mathematics, 710:187–205, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069979/.

Loeb:1975:FKP

[Loe75] Jean Jacques Loeb. Formule de Kirillov pour les groupes de Lie semi-simples compacts. (French) []. Lecture Notes in Mathematics, 497:230–256, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078019/.

Lohoue:1975:RCD

[Loh75] Noël Lohoue. Sur la racine carrée du noyau de Poisson dans les espaces symétriques. (French) []. Lecture Notes in Mathematics, 466:129–134, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082202/.

Lonngren:1976:ESW

[Lon76] Karl E. Lonngren. Experiments on solitary waves. Lecture Notes in Mathematics, 515:12–24, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081160/.

Londen:1979:ANS

[Lon79a] Stig-Olof Londen. On the asymptotics of a nonlinear scalar Volterra integrodifferential equation. Lecture Notes in Mathematics, 737:149–172, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064505/.

Lonsted:1979:AGS

[Løn79b] Knud Lønsted, editor. Algebraic Geometry: Summer Meeting, Copenhagen, August 7–12, 1978, volume 732 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 732. URL http://link.springer.com/book/10.1007/
BFb0066633;   http://www.springerlink.com/content/978-
3-540-35049-1.

Loos:1975:AJP

[Loo75a] Dr. Ottmar Loos. Alternative and Jordan pairs with chain conditions. Lecture Notes in Mathematics, 460:103–140, 1975. CODEN LNMAA2. ISBN 3-540-07166-0 (print), 3-540-37499-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080846/.

Loos:1975:AP

[Loo75b] Dr. Ottmar Loos. Alternative pairs. Lecture Notes in Mathematics, 460:57–102, 1975. CODEN LNMAA2. ISBN 3-540-07166-0 (print), 3-540-37499-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080845/.

Loos:1975:FDJ

[Loo75c] Dr. Ottmar Loos. Finite-dimensional Jordan pairs. Lecture Notes in Mathematics, 460:141–214, 1975. CODEN LNMAA2. ISBN 3-540-07166-0 (print), 3-540-37499-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080847/.

Loos:1975:JPa

[Loo75d] Dr. Ottmar Loos. Jordan pairs. Lecture Notes in Mathematics, 460:1–56, 1975. CODEN LNMAA2. ISBN 3-540-07166-0 (print), 3-540-37499-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080844/.

Loos:1975:JPb

[Loo75e] Dr. Ottmar Loos. Jordan Pairs, volume 460 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07166-0 (print), 3-540-37499-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 218 pp. LCCN QA3

.L28 no. 460. URL http://link.springer.com/book/10.1007/
BFb0080843;   http://www.springerlink.com/content/978-
3-540-37499-2.

Lootsma:1975:DNO

[Loo75f] F. A. Lootsma. The design of a nonlinear optimization programme for solving technological problems. Lecture Notes in Mathematics, 477:229–243, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079179/.

Looijenga:1976:PTS

[Loo76] Eduard Looijenga. Proof of the topological stability theorem. Lecture Notes in Mathematics, 552:124–149, 1976. CODEN LN-MAA2. ISBN 3-540-07997-1 (print), 3-540-37957-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0095249/.

Lorenz:1979:PAS

[Lor79] Edward N. Lorenz. On the prevalence of aperiodicity in simple systems. Lecture Notes in Mathematics, 755:53–75, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069804/.

Loupias:1975:IRF

[Lou75] Michèle Loupias. Indecomposable representations of finite ordered sets. Lecture Notes in Mathematics, 488:201–209, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081225/.

Louveau:1978:RC

[Lou78] A. Louveau. Recursivity and compactness. Lecture Notes in Mathematics, 669:303–337, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103106/.

Love:1975:HIE

[Lov75] Eric Russell Love. A hypergeometric integral equation. Lecture Notes in Mathematics, 457:272–288, 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067112/.

Lowenthal:1976:MCR

[Low76] F. Lowenthal. Measure and categoricity in α-recursion. Lecture Notes in Mathematics, 537:185–201, 1976. CODEN LN-MAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096901/.

Locker:1979:LSM

[LP79a] John Locker and P. M. Prenter. On least squares methods for linear two-point boundary value problems. Lecture Notes in Mathematics, 701:149–168, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062079/.

Lumer:1979:SGH

[LP79b] Gunter Lumer and Luc Paquet. Semi-groupes holomorphes, produit tensoriel de semi-groupes et équations d’évolution. (French) []. Lecture Notes in Mathematics, 713:156–177, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071326/.

Lai:1978:VPC

[LR78a] Pha.m The La.i and Didier Robert. Valeurs propres d’une classe d’équations différentielles singulières sur une demi-droite. (French) []. Lecture Notes in Mathematics, 660:164–171, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082302/.

Langevin:1978:IPF

[LR78b] Rémi Langevin and Harold Rosenberg. Integrable perturbations of fibrations and a theorem of Seifert. Lecture Notes in Mathematics, 652:122–127, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063507/.

Losert:1979:NRW

[LR79] V. Losert and H. Rindler. A note on random walks on compact groups. Lecture Notes in Mathematics, 706:213–219, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063126/.

Lesieur:1976:EST

[LS76] M. Lesieur and P. L. Sulem. Les équations spectrales en turbulence homogène et isotrope. Quelques résultats théoriques et numériques. (French) []. Lecture Notes in Mathematics, 565:113– 120, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091451/.

Lee:1977:ATH

[LS77] Ronnie Lee and R. H. Szczarba. On algebraic K-theory and the homology of congruence subgroups. Lecture Notes in Mathematics, 575:78–87, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095703/.

Lelong:1978:SPL

[LS78] Pierre Lelong and Henri Skoda, editors. Séminaire Pierre Lelong — Henri Skoda (Analyse) Année 1976/77, volume 694 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 694. URL http://link.springer.com/book/10.1007/
BFb0063241;   http://www.springerlink.com/content/978-
3-540-35552-6.

Londen:1979:VEP

[LS79a] Stig-Olof Londen and Olof J. Staffans, editors. Volterra Equations: Proceedings of the Helsinki Symposium on Integral Equations, Otaniemi, Finland, August 11–14, 1978, volume 737 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 737. URL http://link.springer.com/book/10.1007/
BFb0064489;   http://www.springerlink.com/content/978-
3-540-35035-4.

Losert:1979:CPM

[LS79b] Viktor Losert and Klaus Schmidt. A class of probability measures on groups arising from some problems in ergodic theory. Lecture Notes in Mathematics, 706:220–238, 1979. CODEN LN-MAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063127/.

Lachlan:1977:STH

[LSZ77] Alistair Lachlan, Marian Srebrny, and Andrzej Zarach. Set Theory and Hierarchy Theory V, volume 619 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 358 pp. LCCN QA3

.L28 no. 619. URL http://link.springer.com/book/10.1007/
BFb0067639;   http://www.springerlink.com/content/978-
3-540-37032-1.

Lang:1976:FDE

[LT76a] Serge Lang and Hale Freeman Trotter. Frobenius Distributions in GL_{2}-Extensions: Distribution of Frobenius Automorphisms in GL2-Extensions of the Rational Numbers, volume 504 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07550-X (print), 3-540-38094-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 272 (est.) pp. LCCN QA3

.L28 no. 504. URL http://link.springer.com/book/10.1007/
BFb0082087;   http://www.springerlink.com/content/978-
3-540-38094-8.

Lang:1976:IQD

[LT76b] Serge Lang and Hale Freeman Trotter. Imaginary quadratic distribution. Lecture Notes in Mathematics, 504:67–154, 1976. CODEN LNMAA2. ISBN 3-540-07550-X (print), 3-540-38094-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082089/.

Lang:1976:NR

[LT76c] Serge Lang and Hale Freeman Trotter. Numerical results. Lecture Notes in Mathematics, 504:229–272, 1976. CODEN LN-MAA2. ISBN 3-540-07550-X (print), 3-540-38094-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082091/.

Lang:1976:SC

[LT76d] Serge Lang and Hale Freeman Trotter. Special computations. Lecture Notes in Mathematics, 504:155–227, 1976. CODEN LN-MAA2. ISBN 3-540-07550-X (print), 3-540-38094-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082090/.

Lang:1976:SFT

[LT76e] Serge Lang and Hale Freeman Trotter. Supersingular and fixed trace distribution. Lecture Notes in Mathematics, 504:13–66, 1976. CODEN LNMAA2. ISBN 3-540-07550-X (print), 3-540-38094-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082088/.

Lubin:1978:ECS

[Lub78] Arthur Lubin. Extensions of commuting subnormal operators. Lecture Notes in Mathematics, 693:115–120, 1978. CODEN LN-MAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064668/.

Lubkin:1979:FGL

[Lub79a] Saul Lubkin. Finite generations of lifted P -adic homology with compact supports. Generalization of the well conjectures to singular, non-complete algebraic varieties. Lecture Notes in Mathematics, 732:317–373, 1979. CODEN LNMAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066651/.

Lubotzky:1979:PG

[Lub79b] Alexander Lubotzky. On a problem of Grothendieck. Lecture Notes in Mathematics, 732:374–383, 1979. CODEN LN-MAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066652/.

Luckhardt:1975:REC

[Luc75] Horst Luckhardt. The real elements in a consistency proof for simple type theory I. Lecture Notes in Mathematics, 500:233– 256, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079554/.

Luke:1979:SM

[Luk79] Yudell L. Luke. On a summability method. Lecture Notes in Mathematics, 765:295–337, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085587/.

Lumer:1976:PCA

[Lum76a] G. Lumer. Problème de Cauchy avec valeurs au bord continues, comportement asymptotique, et applications. (French) []. Lecture Notes in Mathematics, 563:193–201, 1976. CODEN LN-MAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087579/.

Lumer:1976:PCF

[Lum76b] G. Lumer. Problème de Cauchy et fonctions surharmoniques. (French) []. Lecture Notes in Mathematics, 563:202–218, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087580/.

Lumer:1976:INP

[Lum76c] Gunter Lumer. Images numériques, principe du maximum généralise, et résolvantes. (French) []. Lecture Notes in Mathematics, 518:107–119, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080408/.

Lumley:1977:SAS

[Lum77] John L. Lumley. Some applications of singular perturbations to problems in fluid mechanics. Lecture Notes in Mathematics, 594: 334–350, 1977. CODEN LNMAA2. ISBN 3-540-08258-1 (print), 3-540-37340-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086096/.

Lumer:1978:PDM

[Lum78] G. Lumer. Principe du maximum et équations d’évolution dans L^{2}. (French) []. Lecture Notes in Mathematics, 681:143–156, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065873/.

Lutz:1979:BIM

[Lut79] Donald A. Lutz. Birkhoff invariants and meromorphic differential equations. Lecture Notes in Mathematics, 712:66–76, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062816/.

Lutz:1975:QPF

[LV75] R. Lutz and F. Varela. Quelques propriétés des formes de Pfaff en liaison avec la co-topologie. (French) []. Lecture Notes in Mathematics, 484:234–238, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082155/.

Lakein:1975:MUC

[LW75] Richard B. Lakein and Jennifer Seberry Wallis. On the matrices used to construct Baumert–Hall arrays. Lecture Notes in Mathematics, 452:156–170, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069554/.

Li:1979:PFA

[LY79] Tien-Yien Li and James A. Yorke. Path following approaches for solving nonlinear equations: Homotopy, continuous Newton and projection. Lecture Notes in Mathematics, 730:257–264, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064323/.

Lyche:1976:DPS

[Lyc76] Dr. Tom Lyche. Discrete polynomial spline approximation methods. Lecture Notes in Mathematics, 501:144–176, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079746/.

Maass:1975:CRT

[Maa75] W. Maaß. Church Rosser Theorem für λ-Kalkül mit Unendlich Langen Termen. (German) [Church–Rosser theorem for λ-calculus with infinitely long terms]. Lecture Notes in Mathematics, 500: 257–263, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079555/.

Maass:1978:FST

[Maa78] Wolfgang Maass. Fine structure theory of the constructible universe in α- and β-recursion theory. Lecture Notes in Mathematics, 669:339–359, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103107/.

Macintyre:1975:DET

[Mac75a] Angus Macintyre. Dense embeddings I: a theorem of Robinson in a general setting. Lecture Notes in Mathematics, 498:200–219, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080981/.

Maczynski:1975:NCP

[Mac75b] M. J. Maczyński. Non-commutative probability theory on von Neumann algebras. Lecture Notes in Mathematics, 472:107– 117, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081949/.

Macdonald:1977:SFS

[Mac77a] Sheila Oates Macdonald. Sum-free sets in loops. Lecture Notes in Mathematics, 622:141–147, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069188/.

MacDuff:1977:CS

[Mac77b] Dusa MacDuff. Configuration spaces. Lecture Notes in Mathematics, 575:88–95, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095704/.

MacPherson:1978:CFG

[Mac78] Robert MacPherson. The combinatorial formula of Gabrielov, Gelfand and Losik for the first Pontrjagin class. Lecture Notes in Mathematics, 677:105–124, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070757/.

MacDonald:1979:CBG

[Mac79] Sheila Oates MacDonald. Combinatorics — a branch of group theory? Lecture Notes in Mathematics, 748:11–20, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0102680/.

Madsen:1978:SSS

[Mad78] Ib Madsen. Smooth spherical space forms. Lecture Notes in Mathematics, 657:303–352, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069243/.

Magid:1976:PSF

[Mag76] Andy R. Magid. The Picard sequence of a fibration. Lecture Notes in Mathematics, 549:71–85, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077334/.

Magid:1979:FRA

[Mag79a] Andy R. Magid. Faithfully representable analytic groups. Lecture Notes in Mathematics, 732:384–395, 1979. CODEN LN-MAA2. ISBN 3-540-09527-6 (print), 3-540-35049-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066653/.

Magnus:1979:RCO

[Mag79b] Alphonse Magnus. Recurrence coefficients for orthogonal polynomials on connected and non-connected sets. Lecture Notes in Mathematics, 765:150–171, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085579/.

Mahler:1976:ASM

[Mah76a] Kurt Mahler. Applications of Shidlovski’s main theorems to special functions. Lecture Notes in Mathematics, 546:145–184, 1976. CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081114/.

Mahler:1976:CLS

[Mah76b] Kurt Mahler. Convergent Laurent series and formal Laurent series. Lecture Notes in Mathematics, 546:29–44, 1976. CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081109/.

Mahler:1976:EFP

[Mah76c] Kurt Mahler. Existence and first properties of transcendental numbers. Lecture Notes in Mathematics, 546:1–28, 1976. CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081108/.

Mahler:1976:FRV

[Mah76d] Kurt Mahler. First results on the values of analytic functions at algebraic points. Lecture Notes in Mathematics, 546:45–59, 1976.

CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081110/.

Mahler:1976:FPS

[Mah76e] Kurt Mahler. Formal power series as solutions of algebraic differential equations. Lecture Notes in Mathematics, 546:185–212, 1976. CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081115/.

Mahler:1976:LDEb

[Mah76f] Kurt Mahler. Linear differential equations: a lower bound for the rank of the values of finitely many Siegel E-functions at algebraic points. Lecture Notes in Mathematics, 546:90–114, 1976. CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081112/.

Mahler:1976:LDEc

[Mah76g] Kurt Mahler. Linear differential equations: Shidlovski’s theorems on the transcendency and algebraic independence of values of Siegel E-functions. Lecture Notes in Mathematics, 546:115– 144, 1976. CODEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081113/.

Mahler:1976:LDEa

[Mah76h] Kurt Mahler. Linear differential equations: The lemmas of Shidlovski. Lecture Notes in Mathematics, 546:60–89, 1976. CO-DEN LNMAA2. ISBN 3-540-07986-6 (print), 3-540-37981-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081111/.

Maharam:1977:CBA

[Mah77] D. Maharam. Category, Boolean algebras and measure. Lecture Notes in Mathematics, 609:124–135, 1977. CODEN LN-MAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068678/.

Mahowald:1978:CSR

[Mah78] Mark Mahowald. The construction of small ring spectra. Lecture Notes in Mathematics, 658:234–239, 1978. CODEN LN-MAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068721/.

Mahowald:1979:SHC

[Mah79] Mark Mahowald. Some homotopy classes generated by η_{j}. Lecture Notes in Mathematics, 763:23–37, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088076/.

Malgrange:1975:PBS

[Mal75a] B. Malgrange. Le polynôme de Bernstein d’une singularité isolée. (French) []. Lecture Notes in Mathematics, 459:98–119, 1975. CO-DEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0074194/.

Malliavin:1975:EED

[Mal75b] Paul Malliavin. Elliptic estimates and diffusions in Riemannian geometry and complex analysis. Lecture Notes in Mathematics, 451:26–34, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068577/.

Malliavin:1977:SAP

[Mal77] Marie Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil Paris 1975–1976 (29ème Année). (French) [], volume 586 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LN-MAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 586. URL http://link.springer.com/book/10.1007/
BFb0087114;   http://www.springerlink.com/content/978-
3-540-37369-8.

Malliavin:1978:SAP

[Mal78a] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil Proceedings, Paris 1976–1977 (30ème Année). (French) [], volume 641 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 641. URL http://link.springer.com/book/10.1007/
BFb0064841;   http://www.springerlink.com/content/978-
3-540-35913-5.

Malliavin:1978:THS

[Mal78b] Paul Malliavin. Travaux de H. Skoda sur la classe de Nevanlinna. (French) []. Lecture Notes in Mathematics, 677:201–217, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070764/.

Malyshev:1978:OPS

[Mal78c] V. A. Malyshev. One particle states and scattering theory for Markov processes. Lecture Notes in Mathematics, 653:173– 193, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070092/.

Malgrange:1979:RED

[Mal79a] B. Malgrange. Remarques sur les équations différentielles à points singuliers irreguliers. (French) []. Lecture Notes in Mathematics, 712:77–86, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062817/.

Malgrange:1979:LCS

[Mal79b] Bernard Malgrange. L’Involutivité des caractéristiques des systèmes différentiels et microdifférentiels. (French) [Involutivity of the characteristics of differential and microdifferential systems]. Lecture Notes in Mathematics, 710:277–289, 1979. CODEN LN-MAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069985/.

Malliavin:1979:SAP

[Mal79c] Marie-Paule Malliavin, editor. Séminaire d’Algèbre Paul Dubreil: Proceedings, Paris 1977–78 (31ème Année). (French) [], volume 740 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 740. URL http://link.springer.com/book/10.1007/
BFb0071048;   http://www.springerlink.com/content/978-
3-540-35020-0.

Mallows:1979:STR

[Mal79d] C. L. Mallows. Some theoretical results on Tukey’s 3R smoother. Lecture Notes in Mathematics, 757:77–90, 1979. CODEN LN-MAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098491/.

Mane:1975:AIS

[Mañ75a] R. Mañé. Absolute and infinitessimal stability. Lecture Notes in Mathematics, 468:24–26, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082586/.

Mane:1975:QAD

[Mañ75b] R. Mañé. Quasi-Anosov diffeomorphisms. Lecture Notes in Mathematics, 468:27–29, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082588/.

Manning:1975:CAD

[Man75c] Anthony Manning. Classification of Anosov diffeomorphisms on tori. Lecture Notes in Mathematics, 468:26–27, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082587.

Manning:1975:DSW

[Man75d] Dr. Anthony Manning, editor. Dynamical Systems — Warwick 1974: Proceedings of a Symposium Held at the University of Warwick 1973/74, volume 468 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 468. URL http:

//link.springer.com/book/10.1007/BFb0082572;   http://
www.springerlink.com/content/978-3-540-37525-8.

Mandelbrot:1976:ITF

[Man76a] Benoît Mandelbrot. Intermittent turbulence and fractal dimension: Kurtosis and the spectral exponent 5/3 + B. Lecture Notes in Mathematics, 565:121–145, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091452/.

Mandrekar:1976:CLP

[Man76b] V. Mandrekar. Central limit problem on L_{p} (p ≥ 2) I. Lévy– Khinchine representation. Lecture Notes in Mathematics, 526: 159–166, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082351/.

Mandrekar:1976:GFM

[Man76c] V. Mandrekar. Germ-field Markov property for multiparameter processes. Lecture Notes in Mathematics, 511:78–85, 1976. CO-DEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0101097/.

Mangasarian:1976:SLC

[Man76d] O. L. Mangasarian. Solution of linear complementarity problems by linear programming. Lecture Notes in Mathematics, 506:166– 175, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080123/.

Mandelbrot:1977:LVF

[Man77] Benoît B. Mandelbrot. Lecture V. Fractals and turbulence: Attractors and dispersion. Lecture Notes in Mathematics, 615:83– 93, 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068362/.

Mandrekar:1978:CBS

[Man78a] V. Mandrekar. Characterization of Banach space through validity of Bochner theorem. Lecture Notes in Mathematics, 644:314– 326, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066853/.

Mandrekar:1978:SDF

[Man78b] V. Mandrekar. On subordination of decomposable fields. Lecture Notes in Mathematics, 695:207–210, 1978. CODEN LN-MAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062667/.

Mannshardt:1978:PVS

[Man78c] R. Mannshardt. Prädiktoren mit vorgeschriebenem Stabilitätsverhalten. (German) []. Lecture Notes in Mathematics, 631:81–96, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067465/.

Mann:1979:AIC

[Man79] W. Robert Mann. Averaging to improve convergence of iterative processes. Lecture Notes in Mathematics, 701:169–179, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062080/.

Martinet:1975:BNC

[Mar75] Jacques Martinet. Bases normales et constante de l’équation fonctionnelle des fonctions L d’Artin. (French) []. Lecture Notes in Mathematics, 431:273–294, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066375/.

Marcus:1976:SNR

[Mar76a] Michael B. Marcus. Some new results on central limit theorems for C(S)-valued random variables. Lecture Notes in Mathematics, 526:167–186, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082352/.

Marden:1976:IBF

[Mar76b] Albert Marden. Isomorphisms between Fuchsian groups. Lecture Notes in Mathematics, 505:56–78, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081099/.

Mardesic:1976:PCS

[Mar76c] Sibe Mardešić. Pro-categories and shape theory. Lecture Notes in Mathematics, 540:425–434, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080868/.

Martinet:1976:DVA

[Mar76d] Jean Martinet. Déploiements versels des applications différentiables et classification des applications stables. (French) []. Lecture Notes in Mathematics, 535:1–44, 1976. CODEN LNMAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080494/.

Marsden:1977:LAR

[Mar77] Jerry Marsden. Lecture I. Attempts to relate the Navier-Stokes equations to turbulence. Lecture Notes in Mathematics, 615:1– 22, 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068357/.

Marot:1979:CAE

[Mar79a] Jean Marot. La classe des anneaux excellents est-elle fermée par limite inductive noetherienne plate ?. (French) []. Lecture Notes in Mathematics, 740:198–204, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071061/.

Marsden:1979:GLS

[Mar79b] Jerrold E. Marsden. On the geometry of the Liapunov–Schmidt procedure. Lecture Notes in Mathematics, 755:77–82, 1979. CO-DEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069805/.

Martin:1979:SHS

[Mar79c] Nigel Martin. Some homology 3-spheres which bound acyclic 4manifolds. Lecture Notes in Mathematics, 722:85–92, 1979. CO-DEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063192/.

Martin:1979:ACM

[Mar79d] R. Douglas Martin. Approximate conditional-mean type smoothers and interpolators. Lecture Notes in Mathematics, 757: 117–143, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098493/.

Masser:1975:EPT

[Mas75a] Dr. D. W. Masser. An effective proof of a theorem of Coates. Lecture Notes in Mathematics, 437:44–48, 1975. CODEN LN-MAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069436/.

Masser:1975:EFT

[Mas75b] Dr. D. W. Masser. Elliptic Functions and Transcendence, volume 437 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 112 (est.) pp. LCCN QA3 .L28 no. 437. URL http:

//link.springer.com/book/10.1007/BFb0069432;   http://
www.springerlink.com/content/978-3-540-37410-7.

Masser:1975:LEF

[Mas75c] Dr. D. W. Masser. Lemmas on elliptic functions with complex multiplication. Lecture Notes in Mathematics, 437:63–76, 1975.

CODEN LNMAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069438/.

Masser:1975:LFA

[Mas75d] Dr. D. W. Masser. Linear forms in algebraic points. Lecture Notes in Mathematics, 437:77–112, 1975. CODEN LNMAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069439/.

Masser:1975:LBN

[Mas75e] Dr. D. W. Masser. A lower bound for non-vanishing linear forms. Lecture Notes in Mathematics, 437:49–62, 1975. CODEN LN-MAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069437/.

Masser:1975:TM

[Mas75f] Dr. D. W. Masser. A transcendence measure. Lecture Notes in Mathematics, 437:1–15, 1975. CODEN LNMAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069433/.

Masser:1975:VLFb

[Mas75g] Dr. D. W. Masser. Vanishing of linear forms with complex multiplication. Lecture Notes in Mathematics, 437:36–43, 1975. CO-DEN LNMAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069435/.

Masser:1975:VLFa

[Mas75h] Dr. D. W. Masser. Vanishing of linear forms without complex multiplication. Lecture Notes in Mathematics, 437:16–35, 1975. CODEN LNMAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069434/.

Masani:1976:MPI

[Mas76] P. Masani. Measurability and Pettis integration in Hilbert spaces. Lecture Notes in Mathematics, 541:69–105, 1976. CODEN LN-MAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081043/.

Masani:1978:PD

[Mas78] P. Masani. Propagators and dilations. Lecture Notes in Mathematics, 656:95–117, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068814/.

Masley:1979:WNF

[Mas79] John Myron Masley. Where are number fields with small class number? Lecture Notes in Mathematics, 751:221–242, 1979. CO-DEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062711/.

Matsumoto:1975:AHC

[Mat75] Hideya Matsumoto. Analyse harmonique dans certains systèmes de Coxeter et de Tits. (French) []. Lecture Notes in Mathematics, 497:257–276, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078020/.

Mather:1976:HSM

[Mat76] John N. Mather. How to stratify mappings and jet spaces. Lecture Notes in Mathematics, 535:128–176, 1976. CODEN LN-MAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080499/.

Matheson:1977:CIR

[Mat77a] A. L. Matheson. Closed ideals in rings of analytic functions satisfying a Lipschitz condition. Lecture Notes in Mathematics, 604: 67–72, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069207/.

Matsumoto:1977:AHG

[Mat77b] Hideya Matsumoto. Analyse harmonique dans les groupes de Weyl affines. (French) []. Lecture Notes in Mathematics, 590:

83–132, 1977. CODEN LNMAA2. ISBN 3-540-08249-2 (print), 3-540-37358-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086712/.

Matsumoto:1977:AHSb

[Mat77c] Hideya Matsumoto. Analyse Harmonique dans les Systèmes de Tits Bornologiques de Type Affine. (French) [], volume 590 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LN-MAA2. ISBN 3-540-08249-2 (print), 3-540-37358-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 219 pp. LCCN QA3

.L28 no. 590. URL http://link.springer.com/book/10.1007/
BFb0086707;   http://www.springerlink.com/content/978-
3-540-37358-2.

Matsumoto:1977:GFS

[Mat77d] Hideya Matsumoto. Généralités sur les fonctions sphériques. (French) []. Lecture Notes in Mathematics, 590:8–22, 1977. CO-DEN LNMAA2. ISBN 3-540-08249-2 (print), 3-540-37358-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086709/.

Matsumoto:1977:IFI

[Mat77e] Hideya Matsumoto. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 590:1–7, 1977. CODEN LNMAA2. ISBN 3-540-08249-2 (print), 3-540-37358-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086708/.

Matsumoto:1977:SRG

[Mat77f] Hideya Matsumoto. Systèmes de racines et groupes de Weyl affines. (French) []. Lecture Notes in Mathematics, 590:57– 82, 1977. CODEN LNMAA2. ISBN 3-540-08249-2 (print), 3-540-37358-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086711/.

Matsumoto:1977:STB

[Mat77g] Hideya Matsumoto. Systèmes de Tits bornologiques de type affine. (French) []. Lecture Notes in Mathematics, 590:133–215, 1977. CODEN LNMAA2. ISBN 3-540-08249-2 (print), 3-540-37358-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086713/.

Matsumoto:1977:AHSa

[Mat77h] Hideya Matsumoto. Une analyse harmonique dans les systèmes de Coxeter. (French) []. Lecture Notes in Mathematics, 590:23– 56, 1977. CODEN LNMAA2. ISBN 3-540-08249-2 (print), 3-540-37358-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086710/.

Mathias:1978:PP

[Mat78a] A. R. D. Mathias. O^{#} and the p-point problem. Lecture Notes in Mathematics, 669:375–384, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103109/.

Mattera:1978:TTP

[Mat78b] C. Mattera. Théorème de traces pour une classe d’Espaces de Sobolev singuliers. (French) []. Lecture Notes in Mathematics, 660: 130–156, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082300/.

Maurer:1975:UT

[Mau75] Christian Maurer. Universes in topoi. Lecture Notes in Mathematics, 445:284–296, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061298/.

Mauss:1979:MP

[Mau79] J. Mauss. On matching principles. Lecture Notes in Mathematics, 711:1–8, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062944/.

Maxwell:1978:DMV

[Max78] Charles N. Maxwell. The degree of multiple-valued maps of spheres. Lecture Notes in Mathematics, 664:123–141, 1978. CO-DEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0061696/.

May:1976:HRS

[May76a] J. P. May. The homology of E_{∞} ring spaces. Lecture Notes in Mathematics, 533:69–206, 1976. CODEN LNMAA2. ISBN 3-540-07984-X (print), 3-540-37985-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080466/.

May:1976:HS

[May76b] J. P. May. The homology of E_{∞} spaces. Lecture Notes in Mathematics, 533:1–68, 1976. CODEN LNMAA2. ISBN 3-540-07984-X (print), 3-540-37985-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080465/.

May:1977:RSR

[May77a] J. Peter May. E_{∞} Ring Spaces and E_{∞} Ring Spectra: with contributions by Frank Quinn, Nigel Ray, and Jørgen Tornehave, volume 577 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 256 (est.) pp. LCCN QA3 .L28 no. 577. URL http:

//link.springer.com/book/10.1007/BFb0097608;   http://
www.springerlink.com/content/978-3-540-37409-1.

May:1977:RSB

[May77b] J. Peter May. E_{∞} ring spaces and bipermutative categories. Lecture Notes in Mathematics, 577:139–167, 1977. CODEN LN-MAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097615/.

May:1977:RS

[May77c] J. Peter May. E_{∞} ring spectra. Lecture Notes in Mathematics, 577:63–86, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097613/.

May:1977:F

[May77d] J. Peter May. I functors. Lecture Notes in Mathematics, 577: 9–24, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097610/.

May:1977:ATT

[May77e] J. Peter May. Algebraic and topological K-theory. Lecture Notes in Mathematics, 577:201–243, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097617/.

May:1977:CFS

[May77f] J. Peter May. Coordinate-free spectra. Lecture Notes in Mathematics, 577:25–48, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097611/.

May:1977:I

[May77g] J. Peter May. Introduction. Lecture Notes in Mathematics, 577: 1–8, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097609/.

May:1977:OBT

[May77h] J. Peter May. On kO-oriented bundle theories. Lecture Notes in Mathematics, 577:87–138, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097614/.

May:1977:OT

[May77i] J. Peter May. Orientation theory. Lecture Notes in Mathematics, 577:49–62, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097612/.

May:1977:PIL

[May77j] J. Peter May. Pairings in infinite loop space theory. Lecture Notes in Mathematics, 577:244–256, 1977. CODEN LNMAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097618/.

May:1977:RPR

[May77k] J. Peter May. The recognition principle for E_{∞} ring spaces. Lecture Notes in Mathematics, 577:168–200, 1977. CODEN LN-MAA2. ISBN 3-540-08136-4 (print), 3-540-37409-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097616/.

Mayer:1977:GFQ

[May77l] Meinhard E. Mayer. Gauge fields as quantized connection forms. Lecture Notes in Mathematics, 570:307–349, 1977. CODEN LN-MAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087789/.

May:1978:RSA

[May78a] J. P. May. A_{∞} ring spaces and algebraic K-theory. Lecture Notes in Mathematics, 658:240–315, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068722/.

Mayer:1978:CCS

[May78b] Meinhard E. Mayer. Characteristic classes and solutions of gauge theories. Lecture Notes in Mathematics, 676:81–104, 1978. CO-DEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063666/.

May:1979:AGA

[May79] J. P. May. Applications and generalizations of the approximation theorem. Lecture Notes in Mathematics, 763:38–69, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088077/.

Mazet:1976:DAU

[Maz76] Pierre Mazet. Définition d’une application universelle sur un espace analytique en dimension finie. (French) []. Lecture Notes in Mathematics, 524:67–78, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077998/.

Mazet:1977:SEA

[Maz77a] P. Mazet. Sous-ensembles analytiques de dimension finie d’un espace vectoriel topologique de dimension quelconque. (French) []. Lecture Notes in Mathematics, 578:254–264, 1977. CODEN LN-MAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091475/.

Mazur:1977:RPM

[Maz77b] B. Mazur. Rational points on modular curves. Lecture Notes in Mathematics, 601:107–148, 1977. CODEN LNMAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063947/.

Maronna:1979:BER

[MBY79] Ricardo Maronna, Oscar Bustos, and Víctor Yohai. Bias- and efficiency-robustness of general M-estimators for regression with random carriers. Lecture Notes in Mathematics, 757:91–116, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098492/.

McAvaney:1975:SC

[McA75] K. L. McAvaney. Stability and cacti. Lecture Notes in Mathematics, 452:176–179, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069556/.

McAvaney:1976:NSI

[McA76] K. L. McAvaney. The number and stability indices of C_{n}-trees. Lecture Notes in Mathematics, 560:142–148, 1976. CODEN LN-MAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097376/.

McAloon:1978:CCI

[McA78] Kenneth McAloon. A combinatorial characterization of inaccessible cardinals. Lecture Notes in Mathematics, 669:385–390, 1978.

CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103110/.

McConnell:1977:NMT

[McC77] James McConnell. Note on multiplication theorems for Schur functions. Lecture Notes in Mathematics, 579:252–257, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090020/.

McConnell:1978:GDR

[McC78a] J. C. McConnell. The global dimension of rings of differential operators. Lecture Notes in Mathematics, 641:189–197, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064847/.

McCormick:1978:RMR

[McC78b] S. F. McCormick. A revised mesh refinement strategy for Newton’s method applied to nonlinear two-point boundary value problems. Lecture Notes in Mathematics, 679:15–23, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067863/.

McCrory:1978:SGP

[McC78c] Clint McCrory. Stratified general position. Lecture Notes in Mathematics, 664:142–146, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061697/.

McCleary:1979:DSA

[McC79] John McCleary. modp decompositions of H-spaces; another approach. Lecture Notes in Mathematics, 763:70–87, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088078/.

McDonald:1978:SOA

[McD78] John N. McDonald. Some operators on L^{2}(dm) associated with finite Blaschke products. Lecture Notes in Mathematics, 693:125– 132, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064670/.

McFadden:1975:STI

[McF75] R. McFadden. Structure theorems for inverse semigroups. Lecture Notes in Mathematics, 450:197–208, 1975. CODEN LN-MAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062857/.

McGibbon:1978:WGS

[McG78] Charles A. McGibbon. Which group structures on S^{3} have a maximal torus? Lecture Notes in Mathematics, 657:353–360, 1978. CODEN LNMAA2. ISBN 3-540-08858-X (print), 3-540-35809-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069244/.

McKenna:1975:NFA

[McK75] Kenneth McKenna. New facts about Hilbert’s seventeenth problem. Lecture Notes in Mathematics, 498:220–230, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080982/.

McKean:1979:ISA

[McK79] H. P. McKean. Integrable systems and algebraic curves. Lecture Notes in Mathematics, 755:83–200, 1979. CODEN LN-MAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069806/.

McLeod:1975:SF

[McL75] J. B. McLeod. Swirling flow. Lecture Notes in Mathematics, 448: 242–255, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067089/.

McLeod:1979:KFE

[McL79] John McLeod. The Kunneth formula in equivariant K-theory. Lecture Notes in Mathematics, 741:316–333, 1979. CODEN LN-MAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062147/.

Mastrangelo-Dehen:1978:LFF

[MD78] Michèle Mastrangelo-Dehen. Laplacien fin de fonctions à moyenne convexe. (French) []. Lecture Notes in Mathematics, 681:157– 167, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065874/.

Mebkhout:1975:VPR

[Meb75] Zoghman Mebkhout. Valeur principale et residu simple des formes à singularités essentielles. (French) []. Lecture Notes in Mathematics, 482:190–215, 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080247/.

Mebkhout:1978:CLH

[Meb78] Zoghman Mebkhout. Cohomologie locale d’une hypersurface. (French) []. Lecture Notes in Mathematics, 670:89–119, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064396/.

Medeiros:1975:SWE

[Med75] L. A. Medeiros. Semilinear wave equations. Lecture Notes in Mathematics, 446:329–354, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070609/.

Mehri:1978:FNO

[Meh78a] B. Mehri. Forced nonlinear oscillation for certain third order differential equation. Lecture Notes in Mathematics, 679:73–83, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067868/.

Mehri:1978:ONT

[Meh78b] B. Mehri. Oscillation and nonoscillation theorems for a second order nonlinear functional differential equation. Lecture Notes in Mathematics, 631:97–106, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067466/.

Meinardus:1976:PSG

[Mei76] Günter Meinardus. Periodische Splinefunktionen. (German) [Periodic spline functions]. Lecture Notes in Mathematics, 501:177– 199, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079747/.

Meis:1978:SIE

[Mei78] Theodor Meis. Eine spezielle Integralgleichung erster Art. (German) []. Lecture Notes in Mathematics, 631:107–120, 1978. CO-DEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0067467/.

Mekler:1977:NFA

[Mek77] Alan Mekler. The number of κ-free abelian groups and the size of Ext. Lecture Notes in Mathematics, 616:323–331, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068203/.

Memin:1978:DMS

[Mém78] Jean Mémin. Decompositions multiplicatives de semimartingales exponentielles et applications. (French) []. Lecture Notes in Mathematics, 649:35–46, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064590/.

Metivier:1977:HSI

[Mét77a] Michel Métivier. Das Hilbertsche stochastische Integral. (German) []. Lecture Notes in Mathematics, 607:273–298, 1977. CODEN LNMAA2. ISBN 3-540-08434-7 (print), 3-540-37116-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069568/.

Metivier:1977:SIB

[Mét77b] Michel Métivier. Das Stochastische Integral bezüglich eines Semimartingals (reeller Fall). (German) []. Lecture Notes in Mathematics, 607:221–272, 1977. CODEN LNMAA2. ISBN 3-540-08434-7 (print), 3-540-37116-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069567/.

Metivier:1977:ETS

[Mét77c] Michel Métivier. Einführung in die Theorie der stochastischen Integration Der stetige Fall. (German) []. Lecture Notes in Mathematics, 607:1–90, 1977. CODEN LNMAA2. ISBN 3-540-08434-7 (print), 3-540-37116-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069564/.

Metivier:1977:GBP

[Mét77d] Michel Métivier. Grundlegende Begriffe für Prozesse. (German) []. Lecture Notes in Mathematics, 607:91–136, 1977. CODEN LNMAA2. ISBN 3-540-08434-7 (print), 3-540-37116-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069565/.

Metivier:1977:MQG

[Mét77e] Michel Métivier. Martingale und Quasimartingale. (German) []. Lecture Notes in Mathematics, 607:137–220, 1977. CODEN LN-MAA2. ISBN 3-540-08434-7 (print), 3-540-37116-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069566/.

Metivier:1977:RVQ

[Mét77f] Michel Métivier. Reelle und Vektorwertige Quasimartingale und die Theorie der Stochastischen Integration. (German) [], volume 607 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08434-7 (print), 3-540-37116-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 310 pp. LCCN QA3 .L28 no. 607. URL http:

//link.springer.com/book/10.1007/BFb0069563;   http://
www.springerlink.com/content/978-3-540-37116-8.

Metivier:1978:SIR

[Mét78a] Michel Métivier. Stochastic integration with respect to Hilbert valued martingales, representation theorems and infinite dimensional filtering. Lecture Notes in Mathematics, 695:13–25, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062651/.

Metraux:1978:QIP

[Mét78b] C. Métraux. Quelques inégalités pour martingales à paramètre bidimensionnel. (French) []. Lecture Notes in Mathematics, 649: 170–179, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064605/.

Meyer:1975:WMS

[Mey75a] Albert R. Meyer. Weak monadic second order theory of successor is not elementary-recursive. Lecture Notes in Mathematics, 453:132–154, 1975. CODEN LNMAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064872/.

Meyer:1975:CDE

[Mey75b] Prof. P.-A. Meyer. Complément sur la dualité entre H^{1} et BMO. (French) []. Lecture Notes in Mathematics, 465:237–238, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0102994.

Meyer:1975:QTF

[Mey75c] Prof. P.-A. Meyer. Questions de théorie des flots. (French) []. Lecture Notes in Mathematics, 465:1, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0102981.

Meyer:1975:SPI

[Mey75d] Prof. P.-A. Meyer. Séminaire de Probabilités IX Université de Strasbourg, volume 465 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 317 (est.) pp. LCCN QA3 .L28 no. 465. URL http:

//link.springer.com/book/10.1007/BFb0102980;   http://
www.springerlink.com/content/978-3-540-37518-0.

Meyer:1976:CIL

[Mey76a] P.-A. Meyer. Correction aux “Inégalités de Littlewood–Palet”. (French) []. Lecture Notes in Mathematics, 511:162–163, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101103.

Meyer:1976:GFS

[Mey76b] P.-A. Meyer. Generation of σ-fields by step processes. Lecture Notes in Mathematics, 511:118–124, 1976. CODEN LN-MAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101100/.

Meyer:1976:TPF

[Mey76c] P.-A. Meyer. La théorie de la prediction de F. Knight. (French) []. Lecture Notes in Mathematics, 511:86–103, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101098/.

Meyer:1976:ICF

[Mey76d] P.-A. Meyer. Les inégalités classiques. (French) []. Lecture Notes in Mathematics, 511:125–141, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101101/.

Meyer:1976:IGF

[Mey76e] P.-A. Meyer. Les inégalités générales. (French) []. Lecture Notes in Mathematics, 511:164–174, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101104/.

Meyer:1976:LCD

[Mey76f] P.-A. Meyer. L’opérateur carré du champ. (French) []. Lecture Notes in Mathematics, 511:142–161, 1976. CODEN LN-MAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101102/.

Meyer:1976:SGC

[Mey76g] P.-A. Meyer. Semi-groupes de convolution symétriques. (French) []. Lecture Notes in Mathematics, 511:175–183, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101105/.

Meyer:1976:SPX

[Mey76h] P.-A. Meyer. Séminaire de Probabilités X Université de Strasbourg, volume 511 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 593 pp. LCCN QA3 .L28 no. 511. URL http:

//link.springer.com/book/10.1007/BFb0101091;   http://
www.springerlink.com/content/978-3-540-38197-6.

Meyer:1976:RPG

[Mey76i] Paul André Meyer. Régularité des processus gaussiens. (French) []. Lecture Notes in Mathematics, 514:256–266, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080070/.

Meyer:1977:CFP

[Mey77a] P.-A. Meyer. Convergence faible de processus, d’après Mokobodzki. (French) []. Lecture Notes in Mathematics, 581:109– 119, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087191/.

Meyer:1977:DDP

[Mey77b] P.-A. Meyer. Le dual de H^{1}(R^{v}): Démonstrations probabilistes. (French) []. Lecture Notes in Mathematics, 581:132–195, 1977.

CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087193/.

Meyer:1977:RRB

[Mey77c] P.-A. Meyer. Résultats récents de Benveniste en théorie des flots. (French) []. Lecture Notes in Mathematics, 581:120–131, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087192/.

Meyer:1978:TJJ

[Mey78] P.-A. Meyer. Sur un théorème de J. Jacod. (French) []. Lecture Notes in Mathematics, 649:57–60, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064594/.

Meyer:1979:RCS

[Mey79] P.-A. Meyer. Une remarque sur le calcul stochastique dépendant d’un paramètre. (French) []. Lecture Notes in Mathematics, 721: 199–215, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070861/.

Mitchell:1978:GGM

[MG78] A. R. Mitchell and D. F. Griffiths. Generalised Galerkin methods for second order equations with significant first derivative terms. Lecture Notes in Mathematics, 630:90–104, 1978. CODEN LN-MAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067699/.

Muller:1979:OCP

[MG79] Hans-Georg Müller and Theo Gasser. Optimal convergence properties of kernel estimates of derivatives of a density function. Lecture Notes in Mathematics, 757:144–154, 1979. CODEN LN-MAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0098494/.

Maurer:1975:OSZ

[MH75] H. Maurer and U. Heidemann. Optimale Steuerprozesse mit Zustandsbeschränkungen. (German) []. Lecture Notes in Mathematics, 477:244–260, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079180/.

McAvaney:1977:GSG

[MHG77] K. L. McAvaney, D. A. Holton, and Douglas D. Grant. Groups with stable graphs. Lecture Notes in Mathematics, 622:148– 152, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069189/.

Michler:1975:GCB

[Mic75] Gerhard O. Michler. Green correspondence between blocks with cyclic defect groups II. Lecture Notes in Mathematics, 488:210– 235, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081226/.

Michel:1976:SEE

[Mic76a] P. Michel. Stricte ergodicité d’ensembles minimaux de substitution. (French) []. Lecture Notes in Mathematics, 532:189– 201, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080180/.

Michor:1976:NIM

[Mic76b] P. Michor. A note on the inverse mapping theorem of F. Berquier. Lecture Notes in Mathematics, 540:435–438, 1976. CODEN LN-MAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080869/.

Micula:1976:BNL

[Mic76c] Dr. G. Micula. Bemerkungen zur Numerischen Lösung von Anfangswertproblemen mit Hilfe Nichtlinearer Spline-Funktionen.

(German) []. Lecture Notes in Mathematics, 501:200–209, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079748/.

Michler:1977:SPM

[Mic77] Gerhard O. Michler. Small projective modules of finite groups. Lecture Notes in Mathematics, 586:34–42, 1977. CODEN LN-MAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087119/.

Michelsohn:1978:CBS

[Mic78a] M. L. Michelsohn. Clifford bundles, spinor bundles and Dirac operators. Lecture Notes in Mathematics, 664:147–162, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061698/.

Michelsohn:1978:COV

[Mic78b] M. L. Michelsohn. Cohomology operations and vector fields. Lecture Notes in Mathematics, 658:316–330, 1978. CODEN LN-MAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068723/.

Michor:1978:CFM

[Mic78c] Peter W. Michor. Complete functors and maximal operator ideals. Lecture Notes in Mathematics, 651:58–67, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067835/.

Michor:1978:CBM

[Mic78d] Peter W. Michor. Computable bifunctors and minimal operator ideals. Lecture Notes in Mathematics, 651:52–57, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067834/.

Michor:1978:CF

[Mic78e] Peter W. Michor. Computable functors. Lecture Notes in Mathematics, 651:17–26, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067831/.

Michor:1978:DF

[Mic78f] Peter W. Michor. Duality of functors. Lecture Notes in Mathematics, 651:27–38, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067832/.

Michor:1978:FCB

[Mic78g] Peter W. Michor. Functors and Categories of Banach Spaces: Tensor Products, Operator Ideals and Functors on Categories of Banach Spaces, volume 651 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 99 pp. LCCN QA3

.L28 no. 651. URL http://link.springer.com/book/10.1007/
BFb0067828;   http://www.springerlink.com/content/978-
3-540-35847-3.

Michor:1978:I

[Mic78h] Peter W. Michor. Introduction. Lecture Notes in Mathematics, 651:1–4, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067829/.

Michor:1978:P

[Mic78i] Peter W. Michor. Preliminaries. Lecture Notes in Mathematics, 651:5–16, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067830/.

Michor:1978:PTP

[Mic78j] Peter W. Michor. The projective (p, r, s)-tensor product. Lecture Notes in Mathematics, 651:68–97, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067836/.

Michor:1978:TPO

[Mic78k] Peter W. Michor. Tensor products and operator ideals. Lecture Notes in Mathematics, 651:39–51, 1978. CODEN LNMAA2. ISBN 3-540-08764-8 (print), 3-540-35847-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067833/.

Miller:1975:WFC

[Mil75a] Kenneth S. Miller. The Weyl fractional calculus. Lecture Notes in Mathematics, 457:80–89, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067098/.

Millett:1975:CF

[Mil75b] Kenneth C. Millett. Compact foliations. Lecture Notes in Mathematics, 484:277–287, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082159/.

Millar:1976:SSA

[Mil76] R. F. Millar. The singularities of solutions to analytic elliptic boundary value problems. Lecture Notes in Mathematics, 561: 73–87, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087627/.

Miller:1978:SCC

[Mil78a] John J. H. Miller. Sufficient conditions for the convergence, uniformly in ϵ, of a three point difference scheme for a singular perturbation problem. Lecture Notes in Mathematics, 679:85–91, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067869/.

Millett:1978:AGT

[Mil78b] Kenneth C. Millett, editor. Algebraic and Geometric Topology: Proceedings of a Symposium held at Santa Barbara in honor of Raymond L. Wilder, July 25–29, 1977, volume 664 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 664. URL http://link.springer.com/book/10.1007/
BFb0061684;   http://www.springerlink.com/content/978-
3-540-35758-2.

Miller:1979:WPA

[Mil79] R. K. Miller. Well posedness of abstract Volterra problems. Lecture Notes in Mathematics, 737:192–205, 1979. CODEN LN-MAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064507/.

Mislin:1978:CFD

[Mis78] Guido Mislin. Conditions for finite domination for certain complexes. Lecture Notes in Mathematics, 673:219–224, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064698/.

Mishchenko:1979:AT

[Mis79a] A. S. Mishchenko. C∗-algebras and K-theory. Lecture Notes in Mathematics, 763:262–274, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088090/.

Misiurewicz:1979:IMC

[Mis79b] Michał Misiurewicz. Invariant measures for continuous transformations of [0, 1] with zero topological entropy. Lecture Notes in Mathematics, 729:144–152, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063290/.

Miura:1976:BTI

[Miu76a] Robert M. Miura, editor. Bäcklund Transformations, the Inverse Scattering Method, Solitons, and Their Applications: NSF Research Workshop on Contact Transformations, volume 515 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 515. URL http://link.springer.com/book/10.1007/
BFb0081158;   http://www.springerlink.com/content/978-
3-540-38220-1.

Miura:1976:I

[Miu76b] Robert M. Miura. Introduction. Lecture Notes in Mathematics, 515:1–11, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081159/.

Miyoshi:1977:AMF

[Miy77] T. Miyoshi. Application of a mixed finite element method to a nonlinear problem of elasticity. Lecture Notes in Mathematics, 606:210–223, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064465/.

McLean:1975:AFS

[MJ75] Michael J. McLean and Daniel B. Johnston. An algorithm for finding the shortest terminal strings which can be produced from nonterminals in context-free grammars. Lecture Notes in Mathematics, 452:180–196, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069557/.

Mlak:1978:DRP

[Mla78] W. Mlak. Dissipative random processes. Lecture Notes in Mathematics, 656:139–169, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068817/.

Malliavin:1975:DDS

[MM75a] Marie Paule Malliavin and Paul Malliavin. Diagonalisation du système de de Rham–Hodge au dessus d’un espace Riemannien homogène. (French) []. Lecture Notes in Mathematics, 466:135– 146, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082203/.

Marek:1975:EMZ

[MM75b] W. Marek and A. Mostowski. On extendability of models of ZF set theory to the models of Kelley–Morse theory of classes. Lecture Notes in Mathematics, 499:460–542, 1975. CODEN LN-MAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079429/.

MacCamy:1979:NVV

[MM79a] R. C. MacCamy and Victor J. Mizel. Nonlinear vector-valued hereditary equations on the line. Lecture Notes in Mathematics, 737:206–219, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064508/.

Mohamed:1979:DDC

[MM79b] Saad Mohamed and Bruno J. Müller. Decomposition of dualcontinuous modules. Lecture Notes in Mathematics, 700:87– 94, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063462/.

Markley:1978:SAD

[MMP78] Nelson G. Markley, John C. Martin, and William Perrizo, editors. The Structure of Attractors in Dynamical Systems: Proceedings, North Dakota State University, June 20–24, 1977, volume 668 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 668. URL http://link.springer.com/book/10.1007/
BFb0101774;   http://www.springerlink.com/content/978-
3-540-35751-3.

McAsey:1978:NSA

[MMS78] M. McAsey, P. Muhly, and K.-S. Saito. Non-self-adjoint crossed products. Lecture Notes in Mathematics, 693:121–124, 1978. CO-DEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0064669/.

Meijer:1975:ETN

[MN75a] H. G. Meijer and H. Niederreiter. Équirépartition et théorie des nombres premiers. (French) []. Lecture Notes in Mathematics, 475:104–112, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074260/.

Metakides:1975:RTA

[MN75b] G. Metakides and A. Nerode. Recursion theory and algebra. Lecture Notes in Mathematics, 450:209–219, 1975. CODEN LN-MAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062858/.

Marcus:1979:MME

[MN79] Brian Marcus and Sheldon Newhouse. Measures of maximal entropy for a class of skew products. Lecture Notes in Mathematics, 729:105–125, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063288/.

Moet:1979:SPM

[Moe79] H. J. K. Moet. Singular perturbation methods in a onedimensional free boundary problem. Lecture Notes in Mathematics, 711:63–75, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062947/.

Mohamed:1979:CDC

[Moh79] Saad Mohamed. Continuous and dual-continuous modules. Lecture Notes in Mathematics, 700:226–227, 1979. CODEN LN-MAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063471.

Moishezon:1977:ASM

[Moi77a] Boris Moishezon. Algebraic surfaces and 4-manifolds. Lecture Notes in Mathematics, 586:43–49, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087120/.

Moishezon:1977:CSC

[Moi77b] Boris Moishezon. Complex Surfaces and Connected Sums of Complex Projective Planes, volume 603 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08355-3 (print), 3-540-37276-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 234 pp. LCCN QA3 .L28 no. 603. URL http:

//link.springer.com/book/10.1007/BFb0063355;   http://
www.springerlink.com/content/978-3-540-37276-9.

Moishezon:1977:ES

[Moi77c] Boris Moishezon. Elliptic surfaces. Lecture Notes in Mathematics, 603:116–222, 1977. CODEN LNMAA2. ISBN 3-540-08355-3 (print), 3-540-37276-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063359/.

Moishezon:1977:GPA

[Moi77d] Boris Moishezon. Generic projections of algebraic surfaces into CP^{3}. Lecture Notes in Mathematics, 603:72–115, 1977. CODEN LNMAA2. ISBN 3-540-08355-3 (print), 3-540-37276-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063358/.

Moishezon:1977:I

[Moi77e] Boris Moishezon. Introduction. Lecture Notes in Mathematics, 603:1–5, 1977. CODEN LNMAA2. ISBN 3-540-08355-3 (print), 3-540-37276-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063356/.

Moishezon:1977:RLT

[Moi77f] Boris Moishezon. R. Livne. A theorem about the modular group. Lecture Notes in Mathematics, 603:223–230, 1977. CODEN LN-MAA2. ISBN 3-540-08355-3 (print), 3-540-37276-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063360/.

Moishezon:1977:TSC

[Moi77g] Boris Moishezon. Topology of simply-connected algebraic surfaces of given degree n. Lecture Notes in Mathematics, 603:6–71, 1977. CODEN LNMAA2. ISBN 3-540-08355-3 (print), 3-540-37276-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063357/.

Moishezon:1978:TQT

[Moi78] Boris Moishezon. Topological questions in the theory of algebraic surfaces. Lecture Notes in Mathematics, 670:120–123, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064397/.

Mokobodzki:1976:LFS

[Mok76] Gabriel Mokobodzki. Sur la limite faible d’une suite de fonctions boreliennes. (French) []. Lecture Notes in Mathematics, 563:219– 259, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087581/.

Mokobodzki:1978:ERA

[Mok78a] Gabriel Mokobodzki. Espaces reticules et algèbre de fonctions stables par composition avec les fonctions C^{2}. (French) []. Lecture Notes in Mathematics, 681:209–214, 1978. CODEN LN-MAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065877/.

Mokobodzki:1978:ACD

[Mok78b] Gabriel Mokobodzki. Sur l’algèbre contenue dans le domaine etendu d’un générateur infinitesimal. (French) []. Lecture Notes in Mathematics, 681:168–187, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065875/.

Mokobodzki:1978:ORR

[Mok78c] Gabriel Mokobodzki. Sur l’opérateur de reduite. Remarques sur un travail de J. M. Bismut. (French) []. Lecture Notes in Mathematics, 681:188–208, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065876/.

Mokobodzki:1979:ECF

[Mok79] Gabriel Mokobodzki. Ensembles compacts de fonctions fortement surmédianes. (French) []. Lecture Notes in Mathematics, 713:178– 193, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071327/.

Molino:1975:AGL

[Mol75] Pierre Molino. Actions des groupes de Lie et presque-connexions. (French) []. Lecture Notes in Mathematics, 484:153–161, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082149/.

Moldestad:1977:A

[Mol77a] Johan Moldestad. Abstract. Lecture Notes in Mathematics, 574: 1–13, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087823/.

Moldestad:1977:AKT

[Mol77b] Johan Moldestad. Abstract Kleene theories. Lecture Notes in Mathematics, 574:60–63, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087830/.

Moldestad:1977:CLS

[Mol77c] Johan Moldestad. Calculation of the lengths of some computations. Lecture Notes in Mathematics, 574:109–119, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087833/.

Moldestad:1977:CD

[Mol77d] Johan Moldestad. The computation domain. Lecture Notes in Mathematics, 574:14–16, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087824/.

Moldestad:1977:CT

[Mol77e] Johan Moldestad. Computation theories on α. Lecture Notes in Mathematics, 574:54–59, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087829/.

Moldestad:1977:CHT

[Mol77f] Johan Moldestad. Computations in Higher Types, volume 574 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 194 (est.) pp. LCCN QA3 .L28 no. 574. URL http:

//link.springer.com/book/10.1007/BFb0087822;   http://
www.springerlink.com/content/978-3-540-37425-1.

Moldestad:1977:CKR

[Mol77g] Johan Moldestad. Connection with Kleene recursion in higher types. Lecture Notes in Mathematics, 574:26–33, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087826/.

Moldestad:1977:FCC

[Mol77h] Johan Moldestad. A final comment concerning the two types. Lecture Notes in Mathematics, 574:193–194, 1977. CODEN LN-MAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087836.

Moldestad:1977:G

[Mol77i] Johan Moldestad. Gaps. Lecture Notes in Mathematics, 574:120– 149, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087834/.

Moldestad:1977:KRN

[Mol77j] Johan Moldestad. Kleene recursion in normal objects of type n+2, n > 0. Lecture Notes in Mathematics, 574:50–53, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087828/.

Moldestad:1977:MAM

[Mol77k] Johan Moldestad. More about Mahloness. Lecture Notes in Mathematics, 574:99–108, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087832/.

Moldestad:1977:NCT

[Mol77l] Johan Moldestad. Normal computation theories on α. Lecture Notes in Mathematics, 574:64–98, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087831/.

Moldestad:1977:PFR

[Mol77m] Johan Moldestad. On Platek: “Foundations of recursion theory”. Lecture Notes in Mathematics, 574:150–192, 1977. CODEN LN-MAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087835/.

Moldestad:1977:RNL

[Mol77n] Johan Moldestad. Recursion in normal lists on α. Lecture Notes in Mathematics, 574:34–49, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087827/.

Moldestad:1977:R

[Mol77o] Johan Moldestad. Recursion on α. Lecture Notes in Mathematics, 574:17–25, 1977. CODEN LNMAA2. ISBN 3-540-08132-1 (print), 3-540-37425-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087825/.

Molino:1977:S

[Mol77p] Pierre Molino. G-structures. Lecture Notes in Mathematics, 588: 23–36, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091881/.

Molino:1977:GAF

[Mol77q] Pierre Molino. Généralisations et applications. (French) []. Lecture Notes in Mathematics, 588:149–156, 1977. CODEN LN-MAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091891/.

Molino:1977:IFI

[Mol77r] Pierre Molino. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 588:1–5, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091879/.

Molino:1977:PSP

[Mol77s] Pierre Molino. Presque-structures et problème d’Équivalence. (French) []. Lecture Notes in Mathematics, 588:65–76, 1977. CO-DEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0091884/.

Molino:1977:PSF

[Mol77t] Pierre Molino. Pseudogroupes et Γ-structures. (French) []. Lecture Notes in Mathematics, 588:49–64, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091883/.

Molino:1977:SOS

[Mol77u] Pierre Molino. Structures d’Ordre supérieur. (French) [Higher order structures]. Lecture Notes in Mathematics, 588:37–48, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091882/.

Molino:1977:SPM

[Mol77v] Pierre Molino. Structures plates; modèles standard. (French) [Flat structures; standard models]. Lecture Notes in Mathematics, 588:107–118, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091888/.

Molino:1977:TGFa

[Mol77w] Pierre Molino. Techniques générales 1. (French) []. Lecture Notes in Mathematics, 588:77–82, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091885/.

Molino:1977:TGFb

[Mol77x] Pierre Molino. Techniques générales 2. (French) []. Lecture Notes in Mathematics, 588:83–97, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091886/.

Molino:1977:TGFc

[Mol77y] Pierre Molino. Techniques générales 3. (French) []. Lecture Notes in Mathematics, 588:99–105, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091887/.

Molino:1977:TEFa

[Mol77z] Pierre Molino. Théorème d’Equivalence 1. (French) [Equivalence theorem 1]. Lecture Notes in Mathematics, 588:119–132, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091889/.

Molino:1977:TEFb

[Mol77-27] Pierre Molino. Théorème d’Equivalence 2. (French) [Equivalence theorem 2]. Lecture Notes in Mathematics, 588:133–148, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091890/.

Molino:1977:TSP

[Mol77-28] Pierre Molino. Théorie des G-Structures: Le Problème d’Equivalence: Notes rédigées avec la collaboration de F. Toupine. (French) [Theory of G-Structures: The Problem of Equivalence: Notes written with the collaboration of F. Toupine], volume 588 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LN-MAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 163 pp. LCCN QA3

.L28 no. 588. URL http://link.springer.com/book/10.1007/
BFb0091878;   http://www.springerlink.com/content/978-
3-540-37360-5.

Molino:1977:TJF

[Mol77-29] Pierre Molino. Théorie des jets. (French) []. Lecture Notes in Mathematics, 588:7–22, 1977. CODEN LNMAA2. ISBN 3-540-08246-8 (print), 3-540-37360-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091880/.

Moller:1977:HIS

[Möl77-30] Hans Michael Möller. Hermite interpolation in several variables using ideal-theoretic methods. Lecture Notes in Mathematics, 571: 155–163, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086571/.

Molchanov:1978:ENP

[Mol78] A. M. Molchanov. An estimate of the number of phases. Lecture Notes in Mathematics, 653:140–148, 1978. CODEN LN-MAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070090/.

Moore:1978:SF

[Moo78] John B. Moore. Statistical filtering. Lecture Notes in Mathematics, 680:152–169, 1978. CODEN LNMAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065316/.

Muller:1975:ILC

[MOP75] Prof. Gert H. Müller, Prof. Arnold Oberschelp, and Dr. Klaus Potthoff, editors. [vDash] ISILC Logic Conference: Proceedings of the International Summer Institute and Logic Colloquium, Kiel 1974, volume 499 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 499. URL http:

//link.springer.com/book/10.1007/BFb0079414;   http://
www.springerlink.com/content/978-3-540-38022-1.

Moreau:1976:ACA

[Mor76] J. J. Moreau. Application of convex analysis to the treatment of elastoplastic systems. Lecture Notes in Mathematics, 503:56– 89, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088746/.

Moreno:1977:APE

[Mor77a] C. J. Moreno. Analytic properties of Euler products of automorphic representations. Lecture Notes in Mathematics, 627:11– 26, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065295/.

Moreno:1977:EFT

[Mor77b] C. J. Moreno. Explicit formulas in the theory of automorphic forms. Lecture Notes in Mathematics, 626:73–216, 1977. CODEN LNMAA2. ISBN 3-540-08529-7 (print), 3-540-35985-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063065/.

Morgan:1977:RHT

[Mor77c] John W. Morgan. The rational homotopy theory of smooth, complex projective varieties (following Deligne, Griffiths, Morgan, and Sullivan [2]). Lecture Notes in Mathematics, 567:69–80, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096062/.

Morris:1977:SHL

[Mor77d] A. O. Morris. A survey on Hall–Littlewood functions and their applications to representation theory. Lecture Notes in Mathematics, 579:136–154, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090015/.

Morava:1978:TNS

[Mor78a] Jack Morava. The E_{2}-term of Novikov’s spectral sequence. Lecture Notes in Mathematics, 658:362–382, 1978. CODEN LN-MAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068726/.

Morava:1978:CCC

[Mor78b] Jack Morava. Completions of complex Cobordism. Lecture Notes in Mathematics, 658:349–361, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068725/.

More:1978:LMA

[Mor78c] Jorge J. Moré. The Levenberg–Marquardt algorithm: Implementation and theory. Lecture Notes in Mathematics, 630:105– 116, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067700/.

Moreno:1979:KFF

[Mor79a] Carlos Julio Moreno. Künneth formula for L-functions. Lecture Notes in Mathematics, 751:243–255, 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062712/.

Morgan:1979:CBD

[Mor79b] Elizabeth J. Morgan. Construction of balanced designs and related identities. Lecture Notes in Mathematics, 748:79–91, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0102686/.

Morton:1979:CES

[Mor79c] H. R. Morton. A criterion for an embedded surface in R^{3} to be unknotted. Lecture Notes in Mathematics, 722:93–98, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063193/.

Mostowski:1975:EF

[Mos75] A. Mostowski. An exposition of forcing. Lecture Notes in Mathematics, 450:220–282, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062859/.

Mosco:1976:IVP

[Mos76a] U. Mosco. Implicit variational problems and quasi variational inequalities. Lecture Notes in Mathematics, 543:83–156, 1976. CODEN LNMAA2. ISBN 3-540-07867-3 (print), 3-540-38075-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079943/.

Mostowski:1976:TRM

[Mos76b] A. Mostowski. Two remarks on the models of Morse’ set theory. Lecture Notes in Mathematics, 537:13–21, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096888/;   http:
//link.springer.com/chapter/10.1007/BFb0096889/.

Mosco:1977:EES

[Mos77a] U. Mosco. Error estimates for some variational inequalities. Lecture Notes in Mathematics, 606:224–236, 1977. CODEN LN-MAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064466/.

Moshinsky:1977:CTT

[Mos77b] M. Moshinsky. Canonical transformations and their representations in quantum mechanics. Lecture Notes in Mathematics, 570: 469–481, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087797/.

Moschovakis:1978:ISI

[Mos78a] Yiannis N. Moschovakis. Inductive scales on inductive sets. Lecture Notes in Mathematics, 689:185–192, 1978. CODEN LN-MAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069300/.

Mosco:1978:ESR

[Mos78b] U. Mosco. Sur l’existence de la solution regulière de l’inéquation quasi-variationnelle non linéaire du contrôle optimal impulsionnel et continu. (French) []. Lecture Notes in Mathematics, 665: 140–159, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061803/.

Moscovici:1979:VTC

[Mos79] Henri Moscovici. A vanishing theorem for L^{2}-cohomology in the nilpotent case. Lecture Notes in Mathematics, 728:201–210, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063344/.

Maruyama:1976:PTJ

[MP76] Gisiro Maruyama and Jurii V. Prokhorov, editors. Proceedings of the Third Japan–USSR Symposium on Probability Theory [Taskent 1975], volume 550 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 0-387-07995-5, 3-540-07995-5 (print), 3-540-37966-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 550; QA3 .L28 no. 550

QA273.A1. URL http://link.springer.com/book/10.1007/
BFb0077479;   http://www.springerlink.com/content/978-
3-540-37966-9.

Markley:1978:RMS

[MP78a] Nelson G. Markley and Michael E. Paul. Recipe minimal sets. Lecture Notes in Mathematics, 668:179–190, 1978. CODEN LN-MAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101791/.

Miller:1978:SAC

[MP78b] Haynes R. Miller and Stewart B. Priddy. On G and the stable Adams conjecture. Lecture Notes in Mathematics, 658:331– 348, 1978. CODEN LNMAA2. ISBN 3-540-08859-8 (print), 3-540-35808-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068724/.

Marcus:1979:BEA

[MP79a] Brian Marcus and Karl Petersen. Balancing ergodic averages. Lecture Notes in Mathematics, 729:126–143, 1979. CODEN LN-MAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063289/.

Marcus:1979:RFS

[MP79b] M. B. Marcus and G. Pisier. Random Fourier series on locally compact abelian groups. Lecture Notes in Mathematics, 721:72– 89, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070851/.

Makkai:1977:ARI

[MR77a] Michael Makkai and Gonzalo E. Reyes. Axioms and rules of inference valid in categories. Lecture Notes in Mathematics, 611: 96–131, 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066205/.

Makkai:1977:BHV

[MR77b] Michael Makkai and Gonzalo E. Reyes. Boolean and Heyting valued models. Lecture Notes in Mathematics, 611:132–142, 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066206/.

Makkai:1977:CT

[MR77c] Michael Makkai and Gonzalo E. Reyes. Classifying topoi. Lecture Notes in Mathematics, 611:272–297, 1977. CODEN LN-MAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066211/.

Makkai:1977:C

[MR77d] Michael Makkai and Gonzalo E. Reyes. Completeness. Lecture Notes in Mathematics, 611:143–164, 1977. CODEN LN-MAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066207/.

Makkai:1977:CC

[MR77e] Michael Makkai and Gonzalo E. Reyes. Conceptual completeness. Lecture Notes in Mathematics, 611:193–236, 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066209/.

Makkai:1977:CFO

[MR77f] Michael Makkai and Gonzalo E. Reyes. Corrections to “First order categorical logic”. Lecture Notes in Mathematics, 611:e1–e8, 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066212/.

Makkai:1977:ETG

[MR77g] Michael Makkai and Gonzalo E. Reyes. Existence theorems on geometric morphisms of topoi. Lecture Notes in Mathematics, 611: 165–192, 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066208/.

Makkai:1977:FOC

[MR77h] Michael Makkai and Gonzalo E. Reyes. First Order Categorical Logic: Model-Theoretical Methods in the Theory of Topoi and Related Categories, volume 611 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 301 pp. LCCN QA3

.L28 no. 611. URL http://link.springer.com/book/10.1007/
BFb0066201;   http://www.springerlink.com/content/978-
3-540-37100-7.

Makkai:1977:GT

[MR77i] Michael Makkai and Gonzalo E. Reyes. Grothendieck topoi. Lecture Notes in Mathematics, 611:11–67, 1977. CODEN LN-MAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066203/.

Makkai:1977:ILC

[MR77j] Michael Makkai and Gonzalo E. Reyes. Interpretation of the logic L_{∞_{ω}} in categories. Lecture Notes in Mathematics, 611:68–95, 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066204/.

Makkai:1977:I

[MR77k] Michael Makkai and Gonzalo E. Reyes. Introduction. Lecture Notes in Mathematics, 611:1–10, 1977. CODEN LNMAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066202/.

Makkai:1977:TC

[MR77l] Michael Makkai and Gonzalo E. Reyes. Theories as categories. Lecture Notes in Mathematics, 611:237–271, 1977. CODEN LN-MAA2. ISBN 3-540-08439-8 (print), 3-540-37100-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066210/.

Menaldi:1977:PVN

[MR77m] JJosé-Luis Menaldi and Edmundo Rofman. Sur les problèmes variationnels noncoercifs et l’équation du transport. (French) []. Lecture Notes in Mathematics, 606:204–209, 1977. CODEN LN-MAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064464/.

Mattila:1979:QMV

[MR79] P. Mattila and S. Rickman. Quasiregular mappings and value distribution. Lecture Notes in Mathematics, 743:204–211, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079493/.

Musial:1978:LVM

[MRN78] Kazimierz Musiał and Czesław Ryll-Nardzewski. Liftings of vector measures and their applications to RNP and WRNP. Lecture Notes in Mathematics, 645:162–171, 1978. CODEN LN-MAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069673/.

Melin:1975:FIO

[MS75] Anders Melin and Johannes Sjöstrand. Fourier integral operators with complex-valued phase functions. Lecture Notes in Mathematics, 459:120–223, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074195/.

Macdonald:1976:CCG

[MS76a] Sheila Oates Macdonald and Anne Penfold Street. On crystallographic colour groups. Lecture Notes in Mathematics, 560:149– 157, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097377/.

Marek:1976:UE

[MS76b] W. Marek and M. Srebrny. Urelements and extendability. Lecture Notes in Mathematics, 537:203–219, 1976. CODEN LN-MAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096902/.

Mazur:1976:PRC

[MS76c] Barry Mazur and Jean-Pierre Serre. Points rationnels des courbes modulaires x_{0}(N). (French) []. Lecture Notes in Mathematics, 514: 238–255, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080069/.

Morrel:1977:EOA

[MS77] Bernard B. Morrel and I. M. Singer, editors. K-Theory and Operator Algebras: Proceedings of a Conference held at the University of Georgia in Athens, Georgia, April 21–25, 1975, volume 575 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 575. URL http://link.springer.com/book/10.1007/
BFb0095696;   http://www.springerlink.com/content/978-
3-540-37423-7.

Menikoff:1978:EHO

[MS78a] A. Menikoff and J. Sjöstrand. The eigenvalues of hypoelliptic operators. Lecture Notes in Mathematics, 660:157–163, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082301/.

Miamee:1978:MID

[MS78b] A. G. Miamee and H. Salehi. On minimality of infinite dimensional stationary stochastic processes. Lecture Notes in Mathematics, 656:119–128, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068815/.

Miamee:1978:FNO

[MS78c] A. G. Miamee and H. Salehi. On the factorization of a nonnegative operator valued function. Lecture Notes in Mathematics, 656:129–137, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068816/.

Muller:1978:HST

[MS78d] Gert H. Müller and Dana S. Scott, editors. Higher Set Theory: Proceedings, Oberwolfach, Germany, April 13–23, 1977, volume 669 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 669. URL http://link.springer.com/book/10.1007/
BFb0103096;   http://www.springerlink.com/content/978-
3-540-35749-0.

Mullin:1978:NSC

[MS78e] R. C. Mullin and D. Stinson. Near-self-complementary designs and a method of mixed sums. Lecture Notes in Mathematics, 686: 59–67, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062517/.

McKay:1979:CSG

[MS79a] B. D. McKay and R. G. Stanton. The current status of the generalised Moore graph problem. Lecture Notes in Mathematics, 748:21–31, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102681/.

Memin:1979:CPP

[MS79b] J. Mémin and A. N. Shiryayev. Un critère prévisible pour l’uniforme intégrabilité des semimartingales exponentielles. (French) []. Lecture Notes in Mathematics, 721:147–161, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070858/.

Menikoff:1979:ECH

[MS79c] A. Menikoff and J. Sjöstrand. On the eigenvalues of a class of hypoelliptic operators II. Lecture Notes in Mathematics, 755:201– 247, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069807/.

Morgan:1976:DC

[MSW76] Elizabeth J. Morgan, Anne Penfold Street, and Jennifer Seberry Wallis. Designs from cyclotomy. Lecture Notes in Mathematics, 560:158–176, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097378/.

Marek:1976:STH

[MSZ76] Wiktor Marek, Marian Srebrny, and Andrzej Zarach, editors. Set Theory and Hierarchy Theory A Memorial Tribute to Andrzej Mostowski: Bierutowice, Poland 1975, volume 537 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 537. URL http://link.springer.com/book/10.1007/
BFb0096887;   http://www.springerlink.com/content/978-
3-540-38122-8.

Mukherjea:1976:MS

[MT76a] Prof. Arunava Mukherjea and Prof. Nicolas A. Tserpes. Measures on semigroups. Lecture Notes in Mathematics, 547:1–138, 1976. CODEN LNMAA2. ISBN 3-540-07987-4 (print), 3-540-37980-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079840/.

Mukherjea:1976:MTS

[MT76b] Prof. Arunava Mukherjea and Prof. Nicolas A. Tserpes. Measures on Topological Semigroups: Convolution Products and Random Walks, volume 547 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07987-4 (print), 3-540-37980-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 197 pp. LCCN QA3 .L28 no. 547. URL http:

//link.springer.com/book/10.1007/BFb0079839;   http://
www.springerlink.com/content/978-3-540-37980-5.

Mukherjea:1976:RRW

[MT76c] Prof. Arunava Mukherjea and Prof. Nicolas A. Tserpes. Recurrent random walks on topological groups and semigroups. Lecture Notes in Mathematics, 547:139–195, 1976. CODEN LN-MAA2. ISBN 3-540-07987-4 (print), 3-540-37980-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079841/.

May:1977:RPF

[MT77a] Warren May and Elias Toubassi. A result on problem 87 of L. Fuchs. Lecture Notes in Mathematics, 616:354–367, 1977. CO-DEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0068206/.

Miles:1977:IVM

[MT77b] J. Miles and D. Townsend. On the imaginary values of meromorphic functions. Lecture Notes in Mathematics, 599:93–95, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096828/.

Mossino:1977:CPN

[MT77c] J. Mossino and R. Temam. Certains problémes non linéaires de la physique des plasmas. (French) []. Lecture Notes in Mathematics, 606:237–260, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064467/.

Mujica:1978:RAF

[Muj78] Jorge Mujica. Representation of analytic functionals by vector measures. Lecture Notes in Mathematics, 645:147–161, 1978. CO-DEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069672/.

Mukherjea:1979:SRL

[Muk79] Arunava Mukherjea. Some remarks on limits of iterates of probability measures on groups and semigroups. Lecture Notes in Mathematics, 706:239–251, 1979. CODEN LNMAA2. ISBN 3-540-09124-6 (print), 3-540-35406-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063128/.

Muller:1975:IEP

[Mul75a] Iris Muller. Intégrales d’entrelacement pour GL(n, k) où k est un corps p-adique. (French) []. Lecture Notes in Mathematics, 497: 277–348, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078021/.

Muller:1975:ARF

[Mül75b] Wolfgang Müller. On Artin rings of finite representation type. Lecture Notes in Mathematics, 488:236–243, 1975. CODEN LN-MAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081227/.

Mullin:1975:NBW

[Mul75c] R. C. Mullin. A note on balanced weighing matrices. Lecture Notes in Mathematics, 452:28–41, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069541/.

Muller:1978:ENC

[Mül78a] W. Müller. Experiences on numerical calculation of fields. Lecture Notes in Mathematics, 679:93–103, 1978. CODEN LN-MAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067870/.

Mullin:1978:SES

[Mul78b] R. C. Mullin. A survey of extremal (r, λ)-systems and certain applications. Lecture Notes in Mathematics, 686:106–110, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062522/.

Muller:1979:IEP

[Mul79] I. Muller. Intégrales d’entrelacement pour un groupe de Chevalley sur un corps p-adique. (French) [Interleaving integrals for a group of Chevalley over a p-adic field]. Lecture Notes in Mathematics, 739:367–403, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062501/.

Munkholm:1979:WTP

[Mun79] Hans J. Munkholm. Whitehead torsion for PL fiber homotopy equivalences. Lecture Notes in Mathematics, 741:90–101, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062136/.

Musial:1976:ICP

[Mus76] Kazimierz Musiał. Inheritness of compactness and perfectness of measures by thick subsets. Lecture Notes in Mathematics, 541:

31–42, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081038/.

McKean:1977:SQO

[MvM77] Henry P. McKean and Pierre van Moerbeke. Sur le spectre de quelques opérateurs et les variétés de Jacobi. (French) []. Lecture Notes in Mathematics, 567:54–68, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096061/.

Marcus:1978:NCC

[MW78] Michael B. Marcus and Wojbor A. Woyczyński. A necessary condition for the central limit theorem on spaces of stable type. Lecture Notes in Mathematics, 644:327–339, 1978. CODEN LN-MAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066854/.

MacCamy:1979:NAV

[MW79] R. C. MacCamy and Philip Weiss. Numerical approximations for Volterra integral equations. Lecture Notes in Mathematics, 737: 173–191, 1979. CODEN LNMAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064506/.

Meister:1976:FTM

[MWW76] V. Erhard Meister, Wolfgang L. Wendland, and Norbert Weck, editors. Function Theoretic Methods for Partial Differential Equations: Proceedings of the International Symposium Held at Darmstadt, Germany, April 12–15, 1976, volume 561 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 561. URL http://link.springer.com/book/10.1007/
BFb0087619;   http://www.springerlink.com/content/978-
3-540-37536-4.

Marek:1978:CMO

[MZ78] W. Marek and P. Zbierski. On a class of models of the n-th order arithmetic. Lecture Notes in Mathematics, 669:361–374, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103108/.

Nachbin:1978:DSA

[Nac78] Leopoldo Nachbin. Sur la densité des sous-algèbres polynômiales d’applications continûment différentiables. (French) []. Lecture Notes in Mathematics, 694:196–202, 1978. CODEN LN-MAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063249/.

Nagasawa:1976:PAN

[Nag76] Masao Nagasawa. A probabilistic approach to non-linear Dirichlet problem. Lecture Notes in Mathematics, 511:184–193, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101106/.

Namba:1979:RV

[Nam79a] Makoto Namba. R_{n}(V) for n ≤ g. Lecture Notes in Mathematics, 767:58–123, 1979. CODEN LNMAA2. ISBN 3-540-09722-8 (print), 3-540-38518-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085982/.

Namba:1979:FED

[Nam79b] Makoto Namba. Families of effective divisors and linear systems on projective manifolds. Lecture Notes in Mathematics, 767:191– 225, 1979. CODEN LNMAA2. ISBN 3-540-09722-8 (print), 3-540-38518-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085984/.

Namba:1979:FHM

[Nam79c] Makoto Namba. Families of holomorphic maps of compact complex manifolds. Lecture Notes in Mathematics, 767:124–190, 1979. CODEN LNMAA2. ISBN 3-540-09722-8 (print), 3-540-38518-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085983/.

Namba:1979:FLS

[Nam79d] Makoto Namba. Families of linear systems on compact Riemann surfaces. Lecture Notes in Mathematics, 767:226–276, 1979. CO-DEN LNMAA2. ISBN 3-540-09722-8 (print), 3-540-38518-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0085985/.

Namba:1979:FMF

[Nam79e] Makoto Namba. Families of Meromorphic Functions on Compact Riemann Surfaces, volume 767 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09722-8 (print), 3-540-38518-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 284 pp. LCCN QA3

.L28 no. 767. URL http://link.springer.com/book/10.1007/
BFb0085979;   http://www.springerlink.com/content/978-
3-540-38518-9.

Namba:1979:RDT

[Nam79f] Makoto Namba. Review of the deformation theory of compact complex manifolds. Lecture Notes in Mathematics, 767:1–20, 1979. CODEN LNMAA2. ISBN 3-540-09722-8 (print), 3-540-38518-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085980/.

Namba:1979:S

[Nam79g] Makoto Namba. Structure of Hol(V, P^{1}). Lecture Notes in Mathematics, 767:21–57, 1979. CODEN LNMAA2. ISBN 3-540-09722-8 (print), 3-540-38518-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085981/.

Nanopoulos:1975:MIR

[Nan75] F. Nanopoulos. Mesures d’information et représentation de semigroupes associé. (French) []. Lecture Notes in Mathematics, 465: 154–205, 1975. CODEN LNMAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102990/.

Narayana:1978:RPU

[Nar78] T. V. Narayana. Recent progress and unsolved problems in dominance theory. Lecture Notes in Mathematics, 686:68–78, 1978.

CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062518/.

Nashed:1979:PTG

[Nas79a] M. Z. Nashed. On the perturbation theory for generalized inverse operators in Banach spaces. Lecture Notes in Mathematics, 701:180–195, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062081/.

Nashed:1979:FAM

[Nas79b] M. Zuhair Nashed, editor. Functional Analysis Methods in Numerical Analysis: Special Session, American Mathematical Society, St. Louis, Missouri 1977, volume 701 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 701. URL http:

//link.springer.com/book/10.1007/BFb0062071;   http://
www.springerlink.com/content/978-3-540-35530-4.

Nathanson:1977:NTD

[Nat77a] Melvyn B. Nathanson, editor. Number Theory Day: Proceedings of the Conference Held at Rockefeller University, New York 1976, volume 626 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08529-7 (print), 3-540-35985-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 626. URL http:

//link.springer.com/book/10.1007/BFb0063062;   http://
www.springerlink.com/content/978-3-540-35985-2.

Nathanson:1977:OBN

[Nat77b] Melvyn B. Nathanson. Oscillations of bases in number theory and combinatorics. Lecture Notes in Mathematics, 626:217–231, 1977. CODEN LNMAA2. ISBN 3-540-08529-7 (print), 3-540-35985-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063066/.

Nathanson:1979:NTC

[Nat79] Melvyn B. Nathanson, editor. Number Theory Carbondale 1979: Proceedings of the Southern Illinois Number Theory Conference Carbondale, March 30 and 31, 1979, volume 751 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 751. URL http://link.springer.com/book/10.1007/
BFb0062697;   http://www.springerlink.com/content/978-
3-540-34852-8.

Nazarova:1975:POS

[Naz75] L. A. Nazarova. Partially ordered sets with an infinite number of indecomposable representations. Lecture Notes in Mathematics, 488:244–252, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081228/.

Neeman:1977:AGL

[Ne’77] Y. Ne’eman. The application of graded Lie algebras to invariance considerations in particle physics. Lecture Notes in Mathematics, 570:109–144, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087786/.

Neeman:1978:GGT

[Ne’78] Yuval Ne’eman. Gravity is the gauge theory of the paralleltransport modification of the Poincaré group. Lecture Notes in Mathematics, 676:189–215, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063672/.

Nel:1976:CCT

[Nel76] L. D. Nel. Cartesian closed topological categories. Lecture Notes in Mathematics, 540:439–451, 1976. CODEN LNMAA2. ISBN 3-540-07859-2 (print), 3-540-38118-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080870/.

Nerode:1975:LF

[Ner75] A. Nerode. Logic and foundations. Lecture Notes in Mathematics, 450:283–290, 1975. CODEN LNMAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062860/.

Neuhaus:1976:WCU

[Neu76] G. Neuhaus. Weak convergence under contiguous alternatives of the empirical process when parameters are estimated: The D_{k} approach. Lecture Notes in Mathematics, 566:68–82, 1976. CODEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096880/.

Neuberger:1979:BVP

[Neu79] J. W. Neuberger. Boundary value problems for systems of nonlinear partial differential equations. Lecture Notes in Mathematics, 701:196–208, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062082/.

Neveu:1977:PPF

[Nev77] J. Neveu. Processus ponctuels. (French) []. Lecture Notes in Mathematics, 598:249–445, 1977. CODEN LNMAA2. ISBN 3-540-08340-5 (print), 3-540-37307-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097494/.

Newell:1976:IBB

[New76a] Alan C. Newell. The interrelation between Bäcklund transformations and the inverse scattering transform. Lecture Notes in Mathematics, 515:227–240, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081170/.

Newhouse:1976:CST

[New76b] Sheldon E. Newhouse. Conservative systems and two problems of Smale. Lecture Notes in Mathematics, 525:104–110, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077847/.

Newman:1977:DGP

[New77] M. F. Newman. Determination of groups of prime-power order. Lecture Notes in Mathematics, 573:73–84, 1977. CODEN LN-MAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087814/.

Nichanian:1975:TFD

[Nic75a] Marc Nichanian. Les transformées de Fourier des distributions de type positif sur SL(2, R) et la formule des traces. (French) []. Lecture Notes in Mathematics, 497:349–367, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078022/.

Nicolas:1975:PON

[Nic75b] J. L. Nicolas. Problèmes d’optimisation en nombres entiers et approximations diophantiennes. (French) []. Lecture Notes in Mathematics, 475:113–119, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074261/.

Nicolas:1977:CCA

[Nic77] Anne-Marie Nicolas. Sur les conditions de chaines ascendantes dans des groupes abeliens et des modules sans torsion. (French) []. Lecture Notes in Mathematics, 586:50–60, 1977. CODEN LN-MAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087121/.

Nickel:1978:ZZA

[Nic78] Karl Nickel. Ein Zusammenhang zwischen Aufgaben monotoner Art und Intervall-Mathematik. (German) []. Lecture Notes in Mathematics, 631:121–132, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067468/.

Niederreiter:1975:ISF

[Nie75a] H. Niederreiter. Indépendance de suites. (French) []. Lecture Notes in Mathematics, 475:120–131, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074262/.

Niederreiter:1975:RNT

[Nie75b] H. Niederreiter. Résultats nouveaux dans la théorie quantitative de l’équirépartition. (French) []. Lecture Notes in Mathematics, 475:132–154, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074263/.

Niethammer:1977:NAC

[Nie77] Wilhelm Niethammer. On the numerical analytic continuation of power series. Lecture Notes in Mathematics, 571:164–175, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086572/.

Niemi:1978:PTN

[Nie78] Hannu Niemi. Prediction theory for non-stationary sequences of random vectors. Lecture Notes in Mathematics, 656:183–205, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068819/.

Nirenberg:1975:PHL

[Nir75] L. Nirenberg. On a problem of Hans Lewy. Lecture Notes in Mathematics, 459:224–234, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074196/.

Nitsche:1977:CFE

[Nit77] Joachim Nitsche. L_{∞}-convergence of finite element approximations. Lecture Notes in Mathematics, 606:261–274, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064468/.

Nixdorff:1978:ADE

[Nix78] K. Nixdorff. An application of the differential equations of the sound ray. Lecture Notes in Mathematics, 679:105–118, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067871/.

Nohel:1979:NHV

[Noh79a] John A. Nohel. A nonlinear hyperbolic Volterra equation. Lecture Notes in Mathematics, 737:220–235, 1979. CODEN LN-MAA2. ISBN 3-540-09534-9 (print), 3-540-35035-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064509/.

Nohel:1979:NSP

[Noh79b] John A. Nohel. A nonlinear singularly perturbed Volterra functional differential equation. Lecture Notes in Mathematics, 730: 265–282, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064324/.

Norguet:1975:FPV

[Nor75] François Norguet, editor. Fonctions de Plusieurs Variables Complexes II: Séminaire François Norguet Janvier 1974–Juin 1975, volume 482 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 482. URL http:

//link.springer.com/book/10.1007/BFb0080242;   http://
www.springerlink.com/content/978-3-540-37601-9.

Norbury:1978:EPW

[Nor78a] John Norbury. The existence of periodic water waves. Lecture Notes in Mathematics, 648:117–128, 1978. CODEN LN-MAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066408/.

Nordgren:1978:COH

[Nor78b] Eric A. Nordgren. Composition operators on Hilbert spaces. Lecture Notes in Mathematics, 693:37–63, 1978. CODEN LN-MAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064659/.

Norguet:1978:FPV

[Nor78c] François Norguet, editor. Fonctions de Plusieurs Variables Complexes III: Séminaire François Norguet Octobre 1975–Juin 1977, volume 670 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 670. URL http:

//link.springer.com/book/10.1007/BFb0064390;   http://
www.springerlink.com/content/978-3-540-35745-2.

Norsett:1978:UDF

[Nør78d] Syvert P. Nørsett. On using the Du Fort Frankel scheme for determination of the velocity profile in turbulent boundary layer along an oscillating wall. Lecture Notes in Mathematics, 679:119– 126, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067872/.

Noverraz:1975:PCB

[Nov75] Philippe Noverraz. Pseudo-convexité et base de Schauder dans les ELC. (French) []. Lecture Notes in Mathematics, 474:63–82, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077399/.

Novak:1977:GTR

[Nov77a] Josef Novák, editor. General Topology and Its Relations to Modern Analysis and Algebra IV: Proceedings of the Fourth Prague Topological Symposium, 1976 Part A: Invited Papers, volume 609 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 609. URL http://link.springer.com/book/10.1007/
BFb0068667;   http://www.springerlink.com/content/978-
3-540-37108-3.

Noverraz:1977:MGE

[Nov77b] Philippe Noverraz. Sur la mesure gaussienne des ensembles polaires en dimension infinie. (French) []. Lecture Notes in Mathematics, 578:265–268, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091476/.

Noverraz:1978:MGE

[Nov78] Philippe Noverraz. Sur la mesure gaussienne des ensembles polaires en dimension infinie. (French) []. Lecture Notes in Mathematics, 694:203–213, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063250/.

Neumann:1978:SMP

[NR78a] Walter D. Neumann and Frank Raymond. Seifert manifolds, plumbing, μ-invariant and orientation reversing maps. Lecture Notes in Mathematics, 664:163–196, 1978. CODEN LN-MAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061699/.

Newman:1978:TAP

[NR78b] M. F. Newman and J. S. Richardson, editors. Topics in Algebra: Proceedings, 18th Summer Research Institute of the Australian Mathematical Society Australian National University Canberra, January 9–February 17, 1978, volume 697 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 697. URL http://link.springer.com/book/10.1007/
BFb0103118;   http://www.springerlink.com/content/978-
3-540-35549-6.

Nayak:1976:VVI

[NS76] M. K. Nayak and T. P. Srinivasan. Vector valued inner measures. Lecture Notes in Mathematics, 541:107–116, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081044/.

Nakagami:1979:ACA

[NT79a] Yoshiomi Nakagami and Masamichi Takesaki. Action, co-action and duality. Lecture Notes in Mathematics, 731:1–19, 1979. CO-DEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069743/.

Nakagami:1979:AGT

[NT79b] Yoshiomi Nakagami and Masamichi Takesaki. Applications to Galois theory. Lecture Notes in Mathematics, 731:111–128, 1979. CODEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069749/.

Nakagami:1979:DCP

[NT79c] Yoshiomi Nakagami and Masamichi Takesaki. Duality for Crossed Products of von Neumann Algebras, volume 731 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 139 pp. LCCN QA3

.L28 no. 731. URL http://link.springer.com/book/10.1007/
BFb0069742;   http://www.springerlink.com/content/978-
3-540-35125-2.

Nakagami:1979:EPC

[NT79d] Yoshiomi Nakagami and Masamichi Takesaki. Elementary properties of crossed products. Lecture Notes in Mathematics, 731: 20–34, 1979. CODEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069744/.

Nakagami:1979:ID

[NT79e] Yoshiomi Nakagami and Masamichi Takesaki. Integrability and dominance. Lecture Notes in Mathematics, 731:35–61, 1979. CO-DEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069745/.

Nakagami:1979:PAC

[NT79f] Yoshiomi Nakagami and Masamichi Takesaki. Perturbations of actions and co-actions. Lecture Notes in Mathematics, 731:88– 100, 1979. CODEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069747/.

Nakagami:1979:RCC

[NT79g] Yoshiomi Nakagami and Masamichi Takesaki. Relative commutant of crossed products. Lecture Notes in Mathematics, 731:101– 110, 1979. CODEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069748/.

Nakagami:1979:S

[NT79h] Yoshiomi Nakagami and Masamichi Takesaki. Spectrum. Lecture Notes in Mathematics, 731:62–87, 1979. CODEN LNMAA2. ISBN 3-540-09522-5 (print), 3-540-35125-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069746/.

Nunke:1977:WP

[Nun77] R. J. Nunke. Whitehead’s problem. Lecture Notes in Mathematics, 616:240–250, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068200/.

Nussbaum:1979:PSN

[Nus79] Roger D. Nussbaum. Periodic solutions of nonlinear autonomous functional differential equations. Lecture Notes in Mathematics, 730:283–325, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064325/.

Nastasescu:1979:FRM

[NV79a] C. Năstăsescu and F. Van Oystaeyen. Filtered rings and modules. Lecture Notes in Mathematics, 758:1–35, 1979. CODEN LNMAA2. ISBN 3-540-09708-2 (print), 3-540-38478-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067332/.

Nastasescu:1979:LCN

[NV79b] C. Năstăsescu and F. Van Oystaeyen. Local conditions for Noetherian graded rings. Lecture Notes in Mathematics, 758:127– 143, 1979. CODEN LNMAA2. ISBN 3-540-09708-2 (print), 3-540-38478-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067334/.

Nastasescu:1979:TGR

[NV79c] C. Năstăsescu and F. Van Oystaeyen. Topics in graded ring theory. Lecture Notes in Mathematics, 758:36–127, 1979. CODEN LNMAA2. ISBN 3-540-09708-2 (print), 3-540-38478-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067333/.

Nastasescu:1979:GFR

[NV79d] Constantin Năstăsescu and F. Van Oystaeyen. Graded and Filtered Rings and Modules, volume 758 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09708-2 (print), 3-540-38478-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 148 pp. LCCN QA3

.L28 no. 758. URL http://link.springer.com/book/10.1007/
BFb0067331;   http://www.springerlink.com/content/978-
3-540-38478-6.

Nedelec:1979:CEC

[NV79e] J. C. Nédélec and J. C. Verite. Computation of eddy currents on a surface in R^{3} by finite element methods. Lecture Notes in Mathematics, 704:111–126, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063617/.

Oberschelp:1977:MSN

[Obe77] Walter Oberschelp. Monotonicity for structure numbers in theories without identity. Lecture Notes in Mathematics, 579:297– 308, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090025/.

Oberst:1978:URI

[Obe78] Ulrich Oberst. The use of representations in the invariant theory of not necessarily reductive groups. Lecture Notes in Mathematics, 641:112–127, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064843/.

Oden:1975:CMI

[Ode75] Prof. J. Tinsley Oden, editor. Computational Mechanics: International Conference on Computational Methods in Nonlinear Mechanics, Austin, Texas, 1974, volume 461 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 461. URL http://link.springer.com/book/10.1007/
BFb0074145;   http://www.springerlink.com/content/978-
3-540-37503-6.

Oden:1976:TMH

[OL76] J. T. Oden and J. K. Lee. Theory of mixed and hybrid finiteelement approximations in linear elasticity. Lecture Notes in Mathematics, 503:90–109, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088747/.

Oden:1977:DMH

[OL77] J. T. Oden and J. K. Lee. Dual-mixed hybrid finite element method for second-order elliptic problems. Lecture Notes in Mathematics, 606:275–291, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064469/.

Oleinik:1978:NSE

[Ole78] O. A. Oleĭnik. On the Navier-Stokes equations in a domain with the moving boundary. Lecture Notes in Mathematics, 665:160– 169, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061804/.

Olive:1975:UDF

[Oli75] David Olive. Unitarity and discontinuity formulae. Lecture Notes in Mathematics, 449:133–142, 1975. CODEN LNMAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062920/.

Oliveira:1976:SAF

[Oli76] E. R. Arantes Oliveira. Some applications of functional analysis in the mathematical theory of structures. Lecture Notes in Mathematics, 503:166–176, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088752/.

Olsen:1978:CRI

[Ols78a] Catherine L. Olsen. A concrete representation of index theory in von Neumann algebras. Lecture Notes in Mathematics, 693: 133–143, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064671/.

Olson:1978:AGP

[Ols78b] Loren D. Olson, editor. Algebraic Geometry: Proceedings, Tromsø Symposium, Norway, June 27–July 8, 1977, volume 687 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 687. URL http://link.springer.com/book/10.1007/
BFb0062925;   http://www.springerlink.com/content/978-
3-540-35688-2.

Olver:1979:AHF

[Olv79] Peter J. Olver. Appendix: How to find the symmetry group of a differential equation. Lecture Notes in Mathematics, 762:200– 239, 1979. CODEN LNMAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087463/.

OMalley:1978:SPO

[O’M78] R. E. O’Malley, Jr. Singular perturbations and optimal control. Lecture Notes in Mathematics, 680:170–218, 1978. CODEN LN-MAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0065317/.

Oman:1975:NIW

[Oma75] John Oman. Norm identities which characterize inner product spaces. Lecture Notes in Mathematics, 490:116–133, 1975. CO-DEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081137/.

Onolri:1977:VFY

[Ono77] Enrico Onolri. V. Fock, 40 years later. Lecture Notes in Mathematics, 570:72–75, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087783/.

Oodaira:1978:NFW

[Ood78] Hiroshi Oodaira. Note on Freidlin–Wentzell type estimates for stochastic processes. Lecture Notes in Mathematics, 695:145– 153, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062662/.

Oort:1977:SCM

[Oor77] Frans Oort. Singularities of coarse moduli schemes. Lecture Notes in Mathematics, 586:61–76, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087122/.

Ollagnier:1979:DST

[OP79] Jean Moulin Ollagnier and Didier Pinchon. Dynamical systems of total orders. Lecture Notes in Mathematics, 729:153–162, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063291/.

Oppelstrup:1978:RMD

[Opp78] Jesper Oppelstrup. The RKFHB4 method for delay-differential equations. Lecture Notes in Mathematics, 631:133–146, 1978. CO-DEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0067469/.

Orno:1977:SRB

[Ørn77] Peter Ørno. A separable reflexive Banach space having no finite dimensional Čebyšev subspaces. Lecture Notes in Mathematics, 604:73–75, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069208/.

Ortiz:1978:NSN

[Ort78] E. L. Ortiz. On the numerical solution of nonlinear and functional differential equations with the tau method. Lecture Notes in Mathematics, 679:127–139, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067873/.

Orzech:1976:BGG

[Orz76] Morris Orzech. Brauer groups of graded algebras. Lecture Notes in Mathematics, 549:134–147, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077340/.

Osius:1975:LST

[Osi75a] Gerhard Osius. Logical and set theoretical tools in elementary topoi. Lecture Notes in Mathematics, 445:297–346, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061299/.

Osius:1975:NKJ

[Osi75b] Gerhard Osius. A note on Kripke–Joyal semantics for the internal language of Topoi. Lecture Notes in Mathematics, 445:349– 354, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061300/.

Osikawa:1978:CIN

[Osi78] Motosige Osikawa. Construction of ITPFI with non-trivial uncountable T -set. Lecture Notes in Mathematics, 650:186–188, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067402/.

Osofsky:1979:RPD

[Oso79] Barbara L. Osofsky. Remarks on the projective dimension of ℵunions. Lecture Notes in Mathematics, 734:223–235, 1979. CO-DEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0103161/.

Osswald:1975:SIL

[Oss75] H. Osswald. Über Skolemerweiterungen in der Intuitionistichen Logik mit Gleichheit. (German) []. Lecture Notes in Mathematics, 500:264–266, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079556/.

Ovrelid:1976:P

[Øvr76a] Nils Øvrelid. Problems. Lecture Notes in Mathematics, 512:193– 204, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080035/.

Ovrelid:1976:POE

[Øvr76b] Nils Øvrelid. Pseudodifferential operators and the ¯∂-equation. Lecture Notes in Mathematics, 512:185–192, 1976. CODEN LN-MAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080034/.

Oden:1975:CDF

[OW75] J. T. Oden and L. C. Wellford, Jr. Continuous and discontinuous finite element approximations of shock waves in nonlinear elastic solids. Lecture Notes in Mathematics, 461:149–168, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074153/.

Osborne:1978:NAP

[OW78] M. R. Osborne and G. A. Watson. Nonlinear approximation problems in vector norms. Lecture Notes in Mathematics, 630:117– 132, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067701/.

Oxley:1979:GCP

[Oxl79] James G. Oxley. A generalization of a covering problem of Mullin and Stanton for matroids. Lecture Notes in Mathematics, 748: 92–97, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102687/.

Pacholski:1976:LRP

[Pac76] L. Pacholski. On limit reduced powers, saturateness and universality. Lecture Notes in Mathematics, 537:221–239, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096903/.

Page:1979:AGF

[Pag79] Annie Page. Actions de groupes. (French) []. Lecture Notes in Mathematics, 740:9–24, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071050/.

Palamodov:1978:MVD

[Pal78a] V. P. Palamodov. Moduli in versal deformations of complex spaces. Lecture Notes in Mathematics, 683:74–115, 1978. CODEN LN-MAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063175/.

Palis:1978:RCD

[Pal78b] J. Palis. Rigidity of the centralizers of diffeomorphisms and structural stability of suspended foliations. Lecture Notes in Mathematics, 652:114–121, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063506/.

Papp:1979:LSL

[Pap79a] Z. Papp. Left stable left Noetherian rings. Lecture Notes in Mathematics, 700:228, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063472.

Papp:1979:STS

[Pap79b] Zoltán Papp. Spectrum, topologies and sheaves for left Noetherian rings. Lecture Notes in Mathematics, 700:204–214, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063467/.

Paquet:1979:SGG

[Paq79a] Luc Paquet. Semi-groupes généralisés et équations d’évolution. (French) []. Lecture Notes in Mathematics, 713:243–263, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071329/.

Paquet:1979:SGH

[Paq79b] Luc Paquet. Semi-groupes holomorphes en norme du sup. (French) []. Lecture Notes in Mathematics, 713:194–242, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071328/.

Parikh:1975:LCS

[Par75a] Prof. Dr. Rohit Parikh, editor. Logic Colloquium: Symposium on Logic Held at Boston, 1972–73, volume 453 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 453. URL http://link.springer.com/book/10.1007/
BFb0064868;   http://www.springerlink.com/content/978-
3-540-37483-1.

Parry:1975:STM

[Par75b] William Parry. Suspensions of topological Markov chains. Lecture Notes in Mathematics, 468:30–31, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082589.

Pareigis:1976:NAR

[Par76] Bodo Pareigis. Non-additive ring and module theory IV the Brauer group of a symmetric monoidal category. Lecture Notes in Mathematics, 549:112–133, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077339/.

Pardoux:1978:FDA

[Par78] E. Pardoux, CNRS. Filtrage de diffusions avec conditions frontières: caractérisation de la densité conditionnelle. (French) []. Lecture Notes in Mathematics, 636:163–188, 1978. CODEN LNMAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063264/.

Parry:1979:IOF

[Par79a] William Parry. An information obstruction to finite expected coding length. Lecture Notes in Mathematics, 729:163–168, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063292/.

Parry:1979:LAR

[Par79b] William Parry. The Lorenz attractor and a related population model. Lecture Notes in Mathematics, 729:169–187, 1979. CO-DEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063293/.

Parzen:1979:DQE

[Par79c] Emanuel Parzen. Density quantile estimation approach to statistical data modelling. Lecture Notes in Mathematics, 757:155– 180, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098495/.

Passi:1979:AGR

[Pas79a] Inder Bir S. Passi. The associated graded ring of a group ring. Lecture Notes in Mathematics, 715:106–129, 1979. CODEN LN-MAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067194/.

Passi:1979:DS

[Pas79b] Inder Bir S. Passi. Dimension subgroups. Lecture Notes in Mathematics, 715:13–19, 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067188/.

Passi:1979:DSF

[Pas79c] Inder Bir S. Passi. Dimension subgroups over fields. Lecture Notes in Mathematics, 715:36–49, 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067190/.

Passi:1979:GRT

[Pas79d] Inder Bir S. Passi. Group Rings and Their Augmentation Ideals, volume 715 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 137 pp. LCCN QA3 .L28 no. 715. URL http:

//link.springer.com/book/10.1007/BFb0067186;   http://
www.springerlink.com/content/978-3-540-35297-6.

Passi:1979:GRN

[Pas79e] Inder Bir S. Passi. Group rings of nilpotent groups. Lecture Notes in Mathematics, 715:20–35, 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067189/.

Passi:1979:NRA

[Pas79f] Inder Bir S. Passi. The nilpotent residue of an augmentation ideal. Lecture Notes in Mathematics, 715:95–105, 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067193/.

Passi:1979:PIG

[Pas79g] Inder Bir S. Passi. Polynomial ideals in group rings. Lecture Notes in Mathematics, 715:1–12, 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067187/.

Passi:1979:PMG

[Pas79h] Inder Bir S. Passi. Polynomial maps on groups and dimension subgroups. Lecture Notes in Mathematics, 715:50–75, 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067191/.

Passi:1979:RNA

[Pas79i] Inder Bir S. Passi. The residual nilpotence of the augmentation ideal. Lecture Notes in Mathematics, 715:76–94, 1979. CODEN LNMAA2. ISBN 3-540-09254-4 (print), 3-540-35297-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067192/.

Paugam:1978:IHA

[Pau78] Michel Paugam. Sur les invariants homologiques des anneaux locaux noetheriens: Un calcul de la cinquième deflection ϵ_{5}. (French) []. Lecture Notes in Mathematics, 641:298–338, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064854/.

Palis:1977:GTI

[PdC77] Jacob Palis and Manfredo do Carmo, editors. Geometry and Topology: III Latin American School of Mathematics Proceedings of the School held at the Instituto de Matemática Pura e Aplicada CNPg Rio de Janeiro July 1976, volume 597 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 597. URL http:

//link.springer.com/book/10.1007/BFb0085344;   http://
www.springerlink.com/content/978-3-540-37301-8.

Pixley:1975:CDY

[PE75] Carl Pixley and William Eaton. S^{1} cross a UV decomposition of S^{3} yields S^{1} × S^{3}. Lecture Notes in Mathematics, 438:166– 194, 1975. CODEN LNMAA2. ISBN 3-540-07137-7 (print), 3-540-37412-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066114/.

Pearce:1976:BSS

[Pea76] C. E. M. Pearce. Bordered symmetric square roots of the identity matrix. Lecture Notes in Mathematics, 560:177–184, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097379/.

Pelczynski:1975:UBR

[Peł75] A. Pełczyński. On unconditional bases and Rademacher averages. Lecture Notes in Mathematics, 472:119–129, 1975. CODEN LN-MAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081950/.

Pelczynski:1976:BSP

[Peł76] A. Pełczyński. On Banach space properties of uniform algebras. Lecture Notes in Mathematics, 512:109–116, 1976. CODEN LN-MAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080028/.

Pelant:1977:CPU

[Pel77] Jan Pelant. Combinatorial properties of uniformities. Lecture Notes in Mathematics, 609:154–165, 1977. CODEN LN-MAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068680/.

Peletier:1978:NEP

[Pel78a] L. A. Peletier. A nonlinear eigenvalue problem occurring in population genetics. Lecture Notes in Mathematics, 665:170–187, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061805/.

Pellaumail:1978:ISM

[Pel78b] J. Pellaumail. An inequality for the semi-martingales application to stochastic differential equations. Lecture Notes in Mathematics, 656:207–223, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068820/.

Perrin:1977:REC

[Per77] D. Perrin. La représentation ergodique des codes biprefixes. (French) []. Lecture Notes in Mathematics, 586:156–167, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087131/.

Pereyra:1978:FDS

[Per78a] V. Pereyra. Finite difference solution of two-point boundary value problems and symbolic manipulation. Lecture Notes in Mathematics, 630:133–143, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067702/.

Perrizo:1978:LPF

[Per78b] William Perrizo. A linearization process for flows. Lecture Notes in Mathematics, 668:211–224, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101793/.

Persson:1978:DCS

[Per78c] Ulf Persson. Double coverings and surfaces of general type. Lecture Notes in Mathematics, 687:168–195, 1978. CODEN LN-MAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062932/.

Perovic:1979:ISM

[Per79a] Miodrag Perović. Isolated singularity of the mean quasiconformal mappings. Lecture Notes in Mathematics, 743:212–214, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079494/.

Perronnet:1979:CM

[Per79b] A. Perronnet. The club MODULEF. Lecture Notes in Mathematics, 704:127–153, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063618/.

Pesenti:1976:PWF

[Pes76a] Danielle Pesenti. Produit de Wick des formes sesquilinéaires. (French) []. Lecture Notes in Mathematics, 518:120–143, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080409/.

Pesenti:1976:RCC

[Pes76b] Danielle Pesenti. Relations de commutation canoniques structurés de Fok et champs de Segal. (French) []. Lecture Notes in Mathematics, 518:144–183, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080410/.

Petty:1975:ICC

[Pet75] Clinton M. Petty. Intersectional configuration of convex sets. Lecture Notes in Mathematics, 490:194–205, 1975. CODEN LN-MAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081142/.

Petrich:1976:A

[Pet76a] Mario Petrich. Antiautomorphisms. Lecture Notes in Mathematics, 553:114–167, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094965/.

Petrich:1976:CAS

[Pet76b] Mario Petrich. Categories of Algebraic Systems: Vector and Projective Spaces, Semigroups, Rings and Lattices, volume 553 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CO-DEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 187 (est.) pp. LCCN QA3 .L28 no. 553. URL http:

//link.springer.com/book/10.1007/BFb0094957;   http://
www.springerlink.com/content/978-3-540-37954-6.

Petrich:1976:I

[Pet76c] Mario Petrich. Introduction. Lecture Notes in Mathematics, 553: 1–10, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094958/.

Petrich:1976:La

[Pet76d] Mario Petrich. Lattices. Lecture Notes in Mathematics, 553:50– 62, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094962/.

Petrich:1976:Lb

[Pet76e] Mario Petrich. The loop. Lecture Notes in Mathematics, 553: 99–113, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094964/.

Petrich:1976:PS

[Pet76f] Mario Petrich. Projective spaces. Lecture Notes in Mathematics, 553:63–98, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094963/.

Petrich:1976:R

[Pet76g] Mario Petrich. Rings. Lecture Notes in Mathematics, 553:39– 49, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094961/.

Petrich:1976:S

[Pet76h] Mario Petrich. Semigroups. Lecture Notes in Mathematics, 553: 19–38, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094960/.

Petrich:1976:SC

[Pet76i] Mario Petrich. Special cases. Lecture Notes in Mathematics, 553: 168–187, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094966/.

Petrich:1976:VS

[Pet76j] Mario Petrich. Vector spaces. Lecture Notes in Mathematics, 553: 11–18, 1976. CODEN LNMAA2. ISBN 3-540-07998-X (print), 3-540-37954-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094959/.

Petry:1977:CNQ

[Pét77] A. Pétry. On cardinal numbers in Quine’s New Foundations. Lecture Notes in Mathematics, 619:241–250, 1977. CODEN LN-MAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067655/.

Peters:1978:LTT

[Pet78a] Chris Peters. On the local Torelli theorem, a review of known results. Lecture Notes in Mathematics, 683:62–73, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063174/.

Petrie:1978:SS

[Pet78b] Ted Petrie. G surgery I — a survey. Lecture Notes in Mathematics, 664:197–233, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061700/.

Petry:1978:NUS

[Pet78c] Herbert-Rainer Petry. On the non-uniqueness of spin structure in superconductivity. Lecture Notes in Mathematics, 676:247– 254, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063674/.

Pettis:1978:RNT

[Pet78d] B. J. Pettis. On the Radon–Nikodým theorem. Lecture Notes in Mathematics, 644:340–355, 1978. CODEN LNMAA2. ISBN 3-540-08668-4 (print), 3-540-35906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066855/.

Petryshyn:1979:SNE

[Pet79] W. V. Petryshyn. On the solvability of nonlinear equations involving abstract and differential operators. Lecture Notes in Mathematics, 701:209–247, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062083/.

Pfeiffer:1975:VBO

[Pfe75] H. Pfeiffer. Eine Variante des Bezeichnungssystems W (X) für Ordinalzahlen. (German) []. Lecture Notes in Mathematics, 500: 267–270, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079557/.

Pflug:1975:LPG

[Pfl75a] Dr. Rolf Peter Pflug. Das Levi-Problem. (German) []. Lecture Notes in Mathematics, 432:150–200, 1975. CODEN LN-MAA2. ISBN 3-540-07027-3 (print), 3-540-37314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069145/.

Pflug:1975:HGH

[Pfl75b] Dr. Rolf Peter Pflug. Holomorphiegebiete. (German) [Holomorphy domains]. Lecture Notes in Mathematics, 432:6–43, 1975. CODEN LNMAA2. ISBN 3-540-07027-3 (print), 3-540-37314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069143/.

Pflug:1975:HPG

[Pfl75c] Dr. Rolf Peter Pflug. Holomorphiegebiete, pseudokonvexe Gebiete und das Levi-Problem. (German) [Domains of holomorphy, pseudoconvex regions and the Levi problem], volume 432 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07027-3 (print), 3-540-37314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 210 pp. LCCN QA3

.L28 no. 432. URL http://link.springer.com/book/10.1007/
BFb0069141;   http://www.springerlink.com/content/978-
3-540-37314-8.

Pflug:1975:PG

[Pfl75d] Dr. Rolf Peter Pflug. Pseudokonvexität. (German) []. Lecture Notes in Mathematics, 432:44–149, 1975. CODEN LN-MAA2. ISBN 3-540-07027-3 (print), 3-540-37314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069144/.

Pflug:1975:VGF

[Pfl75e] Dr. Rolf Peter Pflug. Vorwort. (German) [Foreword]. Lecture Notes in Mathematics, 432:1–5, 1975. CODEN LNMAA2. ISBN 3-540-07027-3 (print), 3-540-37314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069142/.

Pfluger:1976:CPS

[Pfl76] Albert Pfluger. On a coefficient problem for Schlicht functions. Lecture Notes in Mathematics, 505:79–91, 1976. CODEN LN-MAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081100/.

Pflug:1978:FPG

[Pfl78] Peter Pflug. Functions of polynomial growth and domains of holomorphy. Lecture Notes in Mathematics, 670:224–232, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064403/.

Pham:1975:MMF

[Pha75a] Frédéric Pham. Microanalyticité de la matrice S. (French) []. Lecture Notes in Mathematics, 449:83–101, 1975. CODEN LN-MAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062917/.

Pham:1975:HTP

[Pha75b] Prof. Frédéric Pham, editor. Hyperfunctions and Theoretical Physics: Rencontre de Nice, 21–30 Mai 1973, volume 449 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 449. URL http://link.springer.com/book/10.1007/
BFb0062913;   http://www.springerlink.com/content/978-
3-540-37454-1.

Philipp:1976:ASI

[Phi76] Walter Philipp. Almost sure invariance principles for empirical distribution functions of weakly dependent random variables. Lecture Notes in Mathematics, 566:83–105, 1976. CODEN LN-MAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096881/.

Philipp:1979:ASI

[Phi79] Walter Philipp. Almost sure invariance principles for sums of Bvalued random variables. Lecture Notes in Mathematics, 709:171– 193, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071957/.

Piene:1978:SFS

[Pie78] Ragni Piene. Some formulas for a surface in P^{3}. Lecture Notes in Mathematics, 687:196–235, 1978. CODEN LNMAA2. ISBN 3-540-08954-3 (print), 3-540-35688-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062933/.

Pinsky:1975:PMD

[Pin75a] Dr. Mark A. Pinsky, editor. Probabilistic Methods in Differential Equations: Proceedings of the Conference Held at the University of Victoria, August 19–20, 1974, volume 451 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 451. URL http://link.springer.com/book/10.1007/
BFb0068574;   http://www.springerlink.com/content/978-
3-540-37481-7.

Pinsky:1975:RE

[Pin75b] Mark A. Pinsky. Random evolutions. Lecture Notes in Mathematics, 451:89–99, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068581/.

Pindor:1979:PAR

[Pin79] Maciej Pindor. Padé approximants and rational functions as tools for finding poles and zeros of analytical functions measured experimentally. Lecture Notes in Mathematics, 765:338–351, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085588/.

Piranian:1977:PMM

[Pir77] George Piranian. The points of maximum modulus of a univalent function. Lecture Notes in Mathematics, 599:96–100, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096829/.

Piriou:1978:PRP

[Pir78] A. Piriou. Propagation et reflexion de la propriété de transmission des distributions de Fourier. (French) []. Lecture Notes in Mathematics, 660:172–180, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082303/.

Pisier:1976:LDL

[Pis76] Gilles Pisier. Sur la loi du logarithme itèrè dans les espaces de Banach. (French) []. Lecture Notes in Mathematics, 526:203– 210, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082354/.

Plante:1975:FMP

[Pla75] J. Plante. Foliations with measure preserving holonomy. Lecture Notes in Mathematics, 468:6–7, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082576.

Pleijel:1975:SST

[Ple75] Åke Pleijel. A survey of spectral theory for pairs of ordinary differential operators. Lecture Notes in Mathematics, 448:256– 272, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067090/.

Plymen:1979:AGR

[Ply79] R. J. Plymen. Automorphic group representations: The hyperfinite II_{1} factor and the Weyl algebra. Lecture Notes in Mathematics, 725:291–306, 1979. CODEN LNMAA2. ISBN 3-540-09512-8 (print), 3-540-35162-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062622/.

Poenaru:1976:DUV

[Poé76a] Valentin Poénaru. Déploiement (uni)-versel des fonctions Ginvariantes. (French) []. Lecture Notes in Mathematics, 510:60– 89, 1976. CODEN LNMAA2. ISBN 3-540-07630-1 (print), 3-540-38172-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079198/.

Poenaru:1976:SPS

[Poé76b] Valentin Poénaru. Singularités C^{∞} en Présence de Symétrie: en Particulier en Présence de la Symétrie d’un Groupe de Lie Compact. (French) [C^{∞} Singularities in the Presence of Symmetry: especially in the presence of Symmetry of a Compact Lie Group], volume 510 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07630-1 (print), 3-540-38172-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 174 pp. LCCN QA3 .L28 no. 510. URL http:

//link.springer.com/book/10.1007/BFb0079196;   http://
www.springerlink.com/content/978-3-540-38172-3.

Poenaru:1976:SSEa

[Poé76c] Valentin Poénaru. Stabilité structurelle équivariante. Lecture Notes in Mathematics, 510:92–136, 1976. CODEN LN-MAA2. ISBN 3-540-07630-1 (print), 3-540-38172-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079199/.

Poenaru:1976:SSEb

[Poé76d] Valentin Poénaru. Stabilité structurelle équivariante. Lecture Notes in Mathematics, 510:142–174, 1976. CODEN LN-MAA2. ISBN 3-540-07630-1 (print), 3-540-38172-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079200/.

Poenaru:1976:TIF

[Poé76e] Valentin Poénaru. Théorie des invariants C^{∞}. (French) []. Lecture Notes in Mathematics, 510:2–58, 1976. CODEN LNMAA2. ISBN 3-540-07630-1 (print), 3-540-38172-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079197/.

Pohlers:1975:UBP

[Poh75] W. Pohlers. An upper bound for the provability of transfinite induction in systems with N-times iterated inductive definitions.

Lecture Notes in Mathematics, 500:271–289, 1975. CODEN LN-MAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079558/.

Poitou:1977:MDD

[Poi77] Georges Poitou. Minorations de discriminants [d’après A. M. Odlyzko]. (French) []. Lecture Notes in Mathematics, 567:136– 153, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096066/.

Pollington:1979:HDS

[Pol79] A. D. Pollington. The Hausdorff dimension of a set of non-normal well approximable numbers. Lecture Notes in Mathematics, 751: 256–264, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062713/.

Pommaret:1975:TDS

[Pom75] J. P. Pommaret. Theory of deformation of structures. Lecture Notes in Mathematics, 484:57–78, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082146/.

Pommerenke:1976:IRS

[Pom76] Ch. Pommerenke. On inclusion relations for spaces of automorphic forms. Lecture Notes in Mathematics, 505:92–100, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081101/.

Popp:1977:AMT

[Pop77a] Herbert Popp. Applications of moduli theory to fibre spaces and the additivity formula for the Kodaira dimension of fibre spaces. Open problems. Lecture Notes in Mathematics, 620:144–158, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067447/.

Popp:1977:AQT

[Pop77b] Herbert Popp. Applications of the quotient theorems to moduli of algebraic varieties. Lecture Notes in Mathematics, 620:44– 54, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067440/.

Popp:1977:CMS

[Pop77c] Herbert Popp. Compactification of moduli spaces. Lecture Notes in Mathematics, 620:117–128, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067445/.

Popp:1977:FMS

[Pop77d] Herbert Popp. Fine moduli spaces. The universal families for stable curves with level n-structure. Lecture Notes in Mathematics, 620:129–143, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067446/.

Popp:1977:GQC

[Pop77e] Herbert Popp. Group quotients in the category of analytic spaces and the category of algebraic spaces. Lecture Notes in Mathematics, 620:32–43, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067439/.

Popp:1977:MSP

[Pop77f] Herbert Popp. Moduli spaces for polarized algebraic varieties. Lecture Notes in Mathematics, 620:18–31, 1977. CODEN LN-MAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067438/.

Popp:1977:MTC

[Pop77g] Herbert Popp. Moduli Theory and Classification Theory of Algebraic Varieties, volume 620 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 189 pp. LCCN QA3

.L28 no. 620. URL http://link.springer.com/book/10.1007/
BFb0067436;   http://www.springerlink.com/content/978-
3-540-37031-4.

Popp:1977:MTA

[Pop77h] Herbert Popp. Moduli theory of algebraic varieties and classification theory of compact complex spaces. Lecture Notes in Mathematics, 620:1–17, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067437/.

Popp:1977:MCM

[Pop77i] Herbert Popp. Mumford’s construction of the moduli variety for curves and polarized abelian varieties. Other applications of Mumford’s quotient theory. Lecture Notes in Mathematics, 620:79– 92, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067443/.

Popp:1977:OMT

[Pop77j] Herbert Popp. Other methods of treating moduli problems. Artin’s method of algebraic stacks. Griffiths’s method of period maps. Lecture Notes in Mathematics, 620:93–116, 1977. CODEN LN-MAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067444/.

Popp:1977:QAS

[Pop77k] Herbert Popp. Quotients for affine schemes by reductive algebraic groups. Lecture Notes in Mathematics, 620:55–67, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067441/.

Popp:1977:QCS

[Pop77l] Herbert Popp. Quotients in the category of schemes. Lecture Notes in Mathematics, 620:68–78, 1977. CODEN LNMAA2. ISBN 3-540-08522-X (print), 3-540-37031-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067442/.

Popp:1978:RDC

[Pop78] Herbert Popp. Recent developments in the classification theory of algebraic varieties. Lecture Notes in Mathematics, 641:281– 297, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064853/.

Porru:1979:QMN

[Por79] Giovanni Porru. Quasiconformal mappings in normed spaces. Lecture Notes in Mathematics, 743:215–222, 1979. CODEN LN-MAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079495/.

Poston:1976:VCM

[Pos76] Tim Poston. Various catastrophe machines. Lecture Notes in Mathematics, 525:111–126, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077848/.

Pourcin:1979:FHD

[Pou79] Geneviève Pourcin. Fibrés holomorphes dont la base et la fibre sont des espaces de stein. (French) []. Lecture Notes in Mathematics, 710:206–220, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069980/.

Powell:1978:FAN

[Pow78a] M. J. D. Powell. A fast algorithm for nonlinearly constrained optimization calculations. Lecture Notes in Mathematics, 630:144– 157, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067703/.

Powers:1978:RII

[Pow78b] R. T. Powers. Resistance inequalities for the isotropic Heisenberg model. Lecture Notes in Mathematics, 650:160–167, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067399/.

Powell:1979:VMM

[Pow79] M. J. D. Powell. Variable metric methods for constrained optimization. Lecture Notes in Mathematics, 704:62–72, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063615/.

Peitgen:1979:LSC

[PP79] Heinz-Otto Peitgen and Michael Prüfer. The Leray–Schauder continuation method is a constructive element in the numerical study of nonlinear eigenvalue and bifurcation problems. Lecture Notes in Mathematics, 730:326–409, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064326/.

Palmer:1978:LSE

[PPW78] Marion Rachel Palmer, William Parry, and Peter Walters. Large sets of endomorphisms and of g-measures. Lecture Notes in Mathematics, 668:191–210, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101792/.

Pastijn:1977:SM

[PR77a] F. Pastijn and H. Reynaerts. Semilattices of modules. Lecture Notes in Mathematics, 586:145–155, 1977. CODEN LN-MAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087130/.

Pelling:1977:PDE

[PR77b] M. J. Pelling and D. G. Rogers. A problem in the design of electrical circuits, a generalized subadditive inequality and the recurrence relation j(n, m) = j([n/2], m)+j([n+1/2], m)+j(n, m−1). Lecture Notes in Mathematics, 622:153–169, 1977. CODEN LN-MAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069190/.

Praeger:1977:PTP

[PR77c] Cheryl E. Praeger and Alan Rahilly. On partially transitive projective planes of certain Hughes types. Lecture Notes in Mathematics, 573:85–111, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087815/.

Prawitz:1975:CGT

[Pra75] Dag Prawitz. Comments on Gentzen-type procedures and the classical notion of truth. Lecture Notes in Mathematics, 500:290– 319, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079559/.

Pratelli:1978:VPT

[Pra78] Maurizio Pratelli. Une version probabiliste d’un théorème d’interpolation de G. Stampacchia. (French) []. Lecture Notes in Mathematics, 649:1–19, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064587/.

Pratelli:1979:SED

[Pra79] Maurizio Pratelli. Le support exact du temps local d’une martingale continue. (French) []. Lecture Notes in Mathematics, 721: 126–131, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070855/.

Preece:1976:NOG

[Pre76a] D. A. Preece. Non-orthogonal Graeco–Latin designs. Lecture Notes in Mathematics, 560:7–26, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097364/.

Preston:1976:AS

[Pre76b] Chris Preston. Attractive specifications. Lecture Notes in Mathematics, 534:160–191, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080572/.

Preston:1976:CMP

[Pre76c] Chris Preston. Continuous models: Point processes. Lecture Notes in Mathematics, 534:87–110, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080569/.

Preston:1976:EGS

[Pre76d] Chris Preston. Existence of Gibbs states. Lecture Notes in Mathematics, 534:33–45, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080566/.

Preston:1976:I

[Pre76e] Chris Preston. Introduction. Lecture Notes in Mathematics, 534: 1–10, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080564/.

Preston:1976:IS

[Pre76f] Chris Preston. Invariant specifications. Lecture Notes in Mathematics, 534:46–58, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080567/.

Preston:1976:LM

[Pre76g] Chris Preston. Lattice models. Lecture Notes in Mathematics, 534:59–86, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080568/.

Preston:1976:RF

[Pre76h] Chris Preston. Random Fields, volume 534 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2.

ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 200 pp. LCCN QA3

.L28 no. 534. URL http://link.springer.com/book/10.1007/
BFb0080563;   http://www.springerlink.com/content/978-
3-540-38193-8.

Preston:1976:RFS

[Pre76i] Chris Preston. Random fields and specifications. Lecture Notes in Mathematics, 534:11–32, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080565/.

Preston:1976:ST

[Pre76j] Chris Preston. Some thermodynamics. Lecture Notes in Mathematics, 534:137–159, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080571/.

Preston:1976:SIG

[Pre76k] Chris Preston. Specific information gain. Lecture Notes in Mathematics, 534:111–136, 1976. CODEN LNMAA2. ISBN 3-540-07852-5 (print), 3-540-38193-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080570/.

Procesi:1978:BHT

[Pro78] Claudio Procesi. Les bases de Hodge dans la théorie des invariants. (French) []. Lecture Notes in Mathematics, 641:128–144, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064844/.

Pearcy:1977:EAR

[PS77] Carl Pearcy and Norberto Salinas. Extensions of C∗-algebras and the reducing essential matricial spectra of an operator. Lecture Notes in Mathematics, 575:96–112, 1977. CODEN LN-MAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095705/.

Pare:1978:AFA

[PS78] Robert Paré and Dietmar Schumacher. Abstract families and the adjoint functor theorems. Lecture Notes in Mathematics, 661:1– 125, 1978. CODEN LNMAA2. ISBN 3-540-08914-4 (print), 3-540-35762-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061361/.

Prevot:1979:MAE

[PS79] D. Prevot and R. Schott. Marches aléatoires sur les espaces homogènes des groupes de Lie nilpotents simplement connexes. (French) []. Lecture Notes in Mathematics, 739:404–426, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062502/.

Padgett:1976:ASC

[PT76] W. J. Padgett and R. L. Taylor. Almost sure convergence of weighted sums of random elements in Banach spaces. Lecture Notes in Mathematics, 526:187–202, 1976. CODEN LN-MAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082353/.

Ptak:1977:NMI

[Ptá77] Vlastimil Pták. Nondiscrete mathematical induction. Lecture Notes in Mathematics, 609:166–178, 1977. CODEN LN-MAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068681/.

Puel:1978:PVP

[Pue78] J.-P. Puel. Un problème de valeur propre non linéaire et de frontière libre. (French) []. Lecture Notes in Mathematics, 665: 188–204, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061806/.

Peitgen:1979:FDE

[PW79] Heinz-Otto Peitgen and Hans-Otto Walther, editors. Functional Differential Equations and Approximation of Fixed Points: Proceedings, Bonn, July 1978, volume 730 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 730. URL http:

//link.springer.com/book/10.1007/BFb0064306;   http://
www.springerlink.com/content/978-3-540-35129-0.

Querre:1977:ACI

[Que77] Julien Querré. Sur les anneaux complètement intégralement clos. (French) []. Lecture Notes in Mathematics, 586:77–85, 1977. CO-DEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087123/.

Querre:1978:IAI

[Que78] Julien Querré. Intersections d’anneaux intégrés (II). (French) [Integrated intersections of rings (II)]. Lecture Notes in Mathematics, 641:358–367, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064856/.

Quine:1975:V

[Qui75] W. V. Quine. The variable. Lecture Notes in Mathematics, 453: 155–168, 1975. CODEN LNMAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064873/.

Quillen:1976:LQM

[Qui76a] Dan Quillen. Letter from Quillen to Milnor on (π_{i}0 → π_{i}^{S} → K_{i}Z). Lecture Notes in Mathematics, 551:182–188, 1976. CODEN LN-MAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080001/.

Quillen:1976:CCR

[Qui76b] Daniel Quillen. Characteristic classes of representations. Lecture Notes in Mathematics, 551:189–216, 1976. CODEN LN-MAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080002/.

Quinn:1979:FCB

[Qui79] Stephen J. Quinn. Factorisation of complete bipartite graphs into two isomorphic subgraphs. Lecture Notes in Mathematics, 748: 98–111, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0102688/.

Raboin:1977:REE

[Rab77] P. Raboin. Résolution de l’équation ¯∂f = F sur un espace de Hilbert. (French) []. Lecture Notes in Mathematics, 578:269– 279, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091477/.

Rabinowitz:1978:MPA

[Rab78a] Paul H. Rabinowitz. A minimax principle and applications to elliptic partial differential equations. Lecture Notes in Mathematics, 648:97–115, 1978. CODEN LNMAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066407/.

Raboin:1978:PDE

[Rab78b] P. Raboin. Le problème du ¯∂ sur un espace de Hilbert. (French) []. Lecture Notes in Mathematics, 694:214–227, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063251/.

Rahilly:1975:STP

[Rah75] Alan Rahilly. Some translation planes with relations which are not translations. Lecture Notes in Mathematics, 452:197–209, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069558/.

Rais:1975:ACG

[Raï75] M. Raïs. Action de certains groupes dans des espaces de fonctions C^{∞}. (French) []. Lecture Notes in Mathematics, 466:147–150, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082204/.

Rais:1978:ODB

[Raï78] Mustapha Raïs. Opérateurs différentiels bi-invariants [d’après M. Duflo]. (French) []. Lecture Notes in Mathematics, 677:125– 137, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070758/.

Rajagopalan:1977:CSS

[Raj77] M. Rajagopalan. Compact C-spaces and S-spaces. Lecture Notes in Mathematics, 609:179–189, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068682/.

Rall:1979:PMS

[Ral79a] L. B. Rall. Perturbation methods for the solution of linear problems. Lecture Notes in Mathematics, 701:248–281, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062084/.

Rallis:1979:ECR

[Ral79b] Stephen Rallis. The Eichler commutation relation and the continuous spectrum of the Weil representation. Lecture Notes in Mathematics, 728:211–244, 1979. CODEN LNMAA2. ISBN 3-540-09516-0 (print), 3-540-35131-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063345/.

Ramis:1978:AVT

[Ram78a] J. P. Ramis. Additif à “Variations sur le thème ‘gaga’ ”. (French) []. Lecture Notes in Mathematics, 694:278–279, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063253.

Ramis:1978:AIV

[Ram78b] J. P. Ramis. Additif II à “Variations sur le theme ‘gaga’ ”. (French) []. Lecture Notes in Mathematics, 694:280–289, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063254/.

Ramis:1978:GAG

[Ram78c] J.-P. Ramis. Géométrie analytique et géométrie algébrique (variations sur le thème ‘gaga’). (French) []. Lecture Notes in Mathematics, 694:228–277, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063252/.

Ramis:1979:PDT

[Ram79a] J.-P. Ramis. A propos du théorème de Borel–Ritt à plusieurs variables. (French) []. Lecture Notes in Mathematics, 712:289– 292, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062820/.

Ramis:1979:FAS

[Ram79b] Jean-Pierre Ramis. Frobenius avec singularités [d’après B. Malgrange, J. F. Mattei et R. Moussu]. (French) []. Lecture Notes in Mathematics, 710:290–299, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069986/.

Rankin:1977:RUW

[Ran77] R. A. Rankin. Ramanujan’s unpublished work on congruences. Lecture Notes in Mathematics, 601:3–15, 1977. CODEN LN-MAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063942/.

Ranicki:1979:LQT

[Ran79a] Andrew Ranicki. Localization in quadratic L-theory. Lecture Notes in Mathematics, 741:102–157, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062137/.

Ranicki:1979:TSO

[Ran79b] Andrew Ranicki. The total surgery obstruction. Lecture Notes in Mathematics, 763:275–316, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088091/.

Rapp:1976:LAD

[Rap76] A. Rapp. L’équation ‘∂’ avec décroissance au bord sur certains ouverts convexes d’un espace de Banach. (French) []. Lecture Notes in Mathematics, 524:21–29, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077995/.

Rasiowa:1975:MVA

[Ras75] H. Rasiowa. Many-valued algorithmic logic. Lecture Notes in Mathematics, 499:543–567, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079430/.

Ratiu:1977:ALB

[Raţ77] Tudor Raţiu. Appendix to Lecture I: Bifurcations, semiflows, and Navier-Stokes equations. Lecture Notes in Mathematics, 615:23– 35, 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068358/.

Ratliff:1978:CCC

[Rat78a] Louis J. Ratliff, Jr. The catenary chain conjecture. Lecture Notes in Mathematics, 647:79–85, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067936/.

Ratliff:1978:CCa

[Rat78b] Louis J. Ratliff, Jr. The chain conjecture. Lecture Notes in Mathematics, 647:45–52, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067929/.

Ratliff:1978:CCR

[Rat78c] Louis J. Ratliff, Jr. Chain Conjectures in Ring Theory: An Exposition of Conjectures on Catenary Chains, volume 647 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 133 pp. LCCN QA3

.L28 no. 647. URL http://link.springer.com/book/10.1007/
BFb0067924;   http://www.springerlink.com/content/978-
3-540-35874-9.

Ratliff:1978:CCb

[Rat78d] Louis J. Ratliff, Jr. Comments on (3.3.1) and conjecture (K). Lecture Notes in Mathematics, 647:88–89, 1978. CODEN LN-MAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067938.

Ratliff:1978:DBR

[Rat78e] Louis J. Ratliff, Jr. Definitions and basic results. Lecture Notes in Mathematics, 647:5–13, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067926/.

Ratliff:1978:DCW

[Rat78f] Louis J. Ratliff, Jr. The depth conjecture and the weak depth conjecture. Lecture Notes in Mathematics, 647:53–56, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067930/.

Ratliff:1978:DGC

[Rat78g] Louis J. Ratliff, Jr. The descended GB-conjecture and the GBconjecture. Lecture Notes in Mathematics, 647:63–66, 1978. CO-DEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0067932/.

Ratliff:1978:C

[Rat78h] Louis J. Ratliff, Jr. The H-conjecture. Lecture Notes in Mathematics, 647:57–62, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067931/.

Ratliff:1978:I

[Rat78i] Louis J. Ratliff, Jr. Introduction. Lecture Notes in Mathematics, 647:1–4, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067925/.

Ratliff:1978:NCC

[Rat78j] Louis J. Ratliff, Jr. The normal chain conjecture. Lecture Notes in Mathematics, 647:86–87, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067937.

Ratliff:1978:SCC

[Rat78k] Louis J. Ratliff, Jr. Some (catenary) chain conjectures. Lecture Notes in Mathematics, 647:20–44, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067928/.

Ratliff:1978:SE

[Rat78l] Louis J. Ratliff, Jr. Some examples. Lecture Notes in Mathematics, 647:90–92, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067939/.

Ratliff:1978:SRS

[Rat78m] Louis J. Ratliff, Jr. Some recently solved problems. Lecture Notes in Mathematics, 647:14–19, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067927/.

Ratliff:1978:SRQ

[Rat78n] Louis J. Ratliff, Jr. Some related questions. Lecture Notes in Mathematics, 647:93–98, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067940/.

Ratliff:1978:SAC

[Rat78o] Louis J. Ratliff, Jr. The strong avoidance conjecture and the avoidance conjecture. Lecture Notes in Mathematics, 647:67–69, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067933/.

Ratliff:1978:TLC

[Rat78p] Louis J. Ratliff, Jr. The taut-level conjecture. Lecture Notes in Mathematics, 647:75–78, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067935/.

Ratliff:1978:UC

[Rat78q] Louis J. Ratliff, Jr. The upper conjecture. Lecture Notes in Mathematics, 647:70–74, 1978. CODEN LNMAA2. ISBN 3-540-08758-3 (print), 3-540-35874-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067934/.

Ratiu:1979:STM

[Raţ79] Tudor Raţiu. On the smoothness of the time t-map of the KdV equation and the bifurcation of the eigenvalues of Hill’s operator. Lecture Notes in Mathematics, 755:248–294, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069808/.

Rauch:1975:LFP

[Rau75a] Jeffrey Rauch. Lecture #1. Five problems: an introduction to the qualitative theory of partial differential equations. Lecture Notes in Mathematics, 446:355–369, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070610/.

Rauch:1975:LMT

[Rau75b] Jeffrey Rauch. Lecture #2. The mathematical theory of crushed ice. Lecture Notes in Mathematics, 446:370–379, 1975. CODEN LNMAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070611/.

Rauch:1975:LSM

[Rau75c] Jeffrey Rauch. Lecture #3. Scattering by many tiny obstacles. Lecture Notes in Mathematics, 446:380–389, 1975. CODEN LN-MAA2. ISBN 3-540-07148-2 (print), 3-540-37440-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070612/.

Rauzy:1975:EEF

[Rau75d] Gérard Rauzy. Équirépartition et entropie. (French) []. Lecture Notes in Mathematics, 475:155–175, 1975. CODEN LN-MAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074264/.

Rauzy:1975:RMA

[Rau75e] Prof. Gérard Rauzy, editor. Répartition Modulo 1: Actes du Colloque de Marseille–Luminy, 4 au 7 Juin 1974, volume 475 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 475. URL http://link.springer.com/book/10.1007/
BFb0074253;   http://www.springerlink.com/content/978-
3-540-37579-1.

Rautmann:1978:USW

[Rau78] Reimund Rautmann. On the uniqueness and stability of weak solutions of a Fokker–Planck–Vlasov equation. Lecture Notes in Mathematics, 679:141–150, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067874/.

Rawnsley:1978:SPH

[Raw78] J. H. Rawnsley. Some properties of half-forms. Lecture Notes in Mathematics, 676:311–314, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063677/.

Rawnsley:1979:SDY

[Raw79] J. H. Rawnsley. Self-dual Yang–Mills fields. Lecture Notes in Mathematics, 755:295–312, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069809/.

Raymond:1978:RWW

[Ray78a] Frank Raymond. R. L. Wilder’s work on generalized manifolds — an appreciation. Lecture Notes in Mathematics, 664:7–32, 1978. CODEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061686/.

Raynaud:1978:FAT

[Ray78b] Michel Raynaud. Faisceaux amples et très amples [d’après T. Matsusaka]. (French) []. Lecture Notes in Mathematics, 677:46– 58, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070753/.

Razar:1977:VDS

[Raz77] Michael J. Razar. Values of Dirichlet series at integers in the critical strip. Lecture Notes in Mathematics, 627:1–10, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065294/.

Rofe-Beketov:1975:DIP

[RB75] F. S. Rofe-Beketov. Deficiency indices and properties of spectrum of some classes of differential operators. Lecture Notes in Mathematics, 448:273–293, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067091/.

Rebolledo:1978:ATM

[Reb78] R. Rebolledo. Sur les applications de la théorie des martingales à l’étude statistique d’une famille de processus ponctuels. (French) []. Lecture Notes in Mathematics, 636:27–70, 1978. CODEN LN-MAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063260/.

Rebolledo:1979:DML

[Reb79] Rolando Rebolledo. Decomposition de martingales locales et rarefaction des sauts. (French) []. Lecture Notes in Mathematics, 721: 138–146, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070857/.

Reddien:1979:DAB

[Red79] G. W. Reddien. Difference approximations to boundary value problems with deviating arguments. Lecture Notes in Mathematics, 701:282–297, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062085/.

Reed:1976:ANL

[Ree76a] Michael C. Reed. Abstract Non Linear Wave Equations, volume 507 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07617-4 (print), 3-540-38135-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 124 (est.) pp. LCCN QA3 .L28 no. 507. URL http:

//link.springer.com/book/10.1007/BFb0079271;   http://
www.springerlink.com/content/978-3-540-38135-8.

Reed:1976:EPS

[Ree76b] Michael C. Reed. Existence and properties of solutions. Lecture Notes in Mathematics, 507:5–66, 1976. CODEN LNMAA2. ISBN 3-540-07617-4 (print), 3-540-38135-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079273/.

Reed:1976:I

[Ree76c] Michael C. Reed. Introduction. Lecture Notes in Mathematics, 507:1–4, 1976. CODEN LNMAA2. ISBN 3-540-07617-4 (print), 3-540-38135-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079272/.

Reed:1976:ST

[Ree76d] Michael C. Reed. Scattering theory. Lecture Notes in Mathematics, 507:67–124, 1976. CODEN LNMAA2. ISBN 3-540-07617-4 (print), 3-540-38135-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079274/.

Reed:1978:SNL

[Ree78a] Michael C. Reed. Singularities in non-linear waves of Klein– Gordon type. Lecture Notes in Mathematics, 648:145–161, 1978. CODEN LNMAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066410/.

Rees:1978:SRT

[Ree78b] Elmer Rees. Some rank two bundles on P_{n}C, whose Chern classes vanish. Lecture Notes in Mathematics, 683:25–28, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063171/.

Reiner:1975:LFC

[Rei75] Irving Reiner. Locally free class groups of orders. Lecture Notes in Mathematics, 488:253–281, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081229/.

Reich:1976:QMD

[Rei76a] Edgar Reich. Quasiconformal mappings of the disk with given boundary values. Lecture Notes in Mathematics, 505:101–137, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081102/.

Reid:1976:SCL

[Rei76b] J. K. Reid. Sparse in-core linear programming. Lecture Notes in Mathematics, 506:176–189, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080124/.

Reid:1977:QPI

[Rei77a] J. D. Reid. Quasi-pure-injectivity and quasi-pure projectivity. Lecture Notes in Mathematics, 616:219–227, 1977. CODEN LN-MAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068198/.

Reid:1977:SLS

[Rei77b] John K. Reid. Solution of linear systems of equations: Direct methods (general). Lecture Notes in Mathematics, 572:102– 129, 1977. CODEN LNMAA2. ISBN 3-540-08130-5 (print), 3-540-37430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0116616/.

Reimer:1977:CSS

[Rei77c] Prof. Dr. M. Reimer. Clenshaw sums in several variables. Lecture Notes in Mathematics, 571:176–185, 1977. CODEN LN-MAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086573/.

Reiner:1978:IRG

[Rei78a] Irving Reiner. Integral representations: Genus, k-theory and class groups. Lecture Notes in Mathematics, 697:52–69, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103122/.

Reiner:1978:IRC

[Rei78b] Irving Reiner. Integral representations of cyclic p-groups. Lecture Notes in Mathematics, 697:70–87, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103123/.

Reiner:1978:IRF

[Rei78c] Irving Reiner. Integral representations of finite groups. Lecture Notes in Mathematics, 641:145–162, 1978. CODEN LN-MAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064845/.

Reiner:1979:TIR

[Rei79] Irving Reiner. Topics in integral representation theory. Lecture Notes in Mathematics, 744:1–143, 1979. CODEN LNMAA2. ISBN 3-540-09546-2 (print), 3-540-35007-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063059/.

Rentschler:1977:CAD

[Ren77] Rudolf Rentschler. Comportement de l’application de Dixmier par rapport à l’antiautomorphisme principal pour des algèbres de Lie complètement résolubles. (French) []. Lecture Notes in Mathematics, 586:93–100, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087125/.

Renault:1979:AGA

[Ren79a] Guy Renault. Actions de groupes et anneaux reguliers injectifs. (French) []. Lecture Notes in Mathematics, 734:236–248, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103162/.

Rentschler:1979:OSP

[Ren79b] Rudolf Rentschler. Orbites dans le spectre primitif de l’algèbre enveloppante d’une algèbre de Lie. (French) []. Lecture Notes in Mathematics, 740:88–98, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071055/.

Revesz:1976:TTM

[Rév76] P. Révész. Three theorems of multivariate empirical process. Lecture Notes in Mathematics, 566:106–126, 1976. CODEN LN-MAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096882/.

Rey:1978:I

[Rey78a] William J. J. Rey. Introduction. Lecture Notes in Mathematics, 690:1–7, 1978. CODEN LNMAA2. ISBN 3-540-09091-6 (print), 3-540-35598-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064678/.

Rey:1978:J

[Rey78b] William J. J. Rey. The jackknife. Lecture Notes in Mathematics, 690:17–29, 1978. CODEN LNMAA2. ISBN 3-540-09091-6 (print), 3-540-35598-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064680/.

Rey:1978:E

[Rey78c] William J. J. Rey. M-estimators. Lecture Notes in Mathematics, 690:30–76, 1978. CODEN LNMAA2. ISBN 3-540-09091-6 (print), 3-540-35598-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064681/.

Rey:1978:SD

[Rey78d] William J. J. Rey. On sampling distributions. Lecture Notes in Mathematics, 690:8–16, 1978. CODEN LNMAA2. ISBN 3-540-09091-6 (print), 3-540-35598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064679/.

Rey:1978:OA

[Rey78e] William J. J. Rey. Open avenues. Lecture Notes in Mathematics, 690:77–87, 1978. CODEN LNMAA2. ISBN 3-540-09091-6 (print), 3-540-35598-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064682/.

Rey:1978:RSM

[Rey78f] William J. J. Rey. Robust Statistical Methods, volume 690 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LN-MAA2. ISBN 3-540-09091-6 (print), 3-540-35598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 128 pp. LCCN QA3

.L28 no. 690. URL http://link.springer.com/book/10.1007/
BFb0064677;   http://www.springerlink.com/content/978-
3-540-35598-4.

Robinson:1979:CWL

[RF79] D. F. Robinson and L. R. Foulds. Comparison of weighted labelled trees. Lecture Notes in Mathematics, 748:119–126, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0102690/.

Rheinboldt:1975:SLS

[Rhe75] Werner C. Rheinboldt. On the solution of large, sparse sets of nonlinear equations. Lecture Notes in Mathematics, 461:169– 194, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074154/.

Rhin:1975:RMQ

[Rhi75] G. Rhin. Répartition modulo 1 de f(p_{n}) quand f est une série entière. (French) []. Lecture Notes in Mathematics, 475:176– 244, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074265/.

Ribet:1977:GRA

[Rib77] Kenneth A. Ribet. Galois representations attached to eigenforms with nebentypus. Lecture Notes in Mathematics, 601:18–52, 1977. CODEN LNMAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063943/.

Richter:1975:LRO

[Ric75] Wayne Richter. The least Σ_{2}^{1} and Π_{2}^{1} reflecting ordinals. Lecture Notes in Mathematics, 499:568–578, 1975. CODEN LN-MAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079431/.

Richman:1977:GVG

[Ric77] Fred Richman. A guide to valuated groups. Lecture Notes in Mathematics, 616:73–86, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068190/.

Riley:1979:EPP

[Ril79] Robert Riley. An elliptical path from parabolic representations to hyperbolic structures. Lecture Notes in Mathematics, 722:99– 133, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063194/.

Ringel:1975:RTL

[Rin75] Claus Michael Ringel. The representation type of local algebras. Lecture Notes in Mathematics, 488:282–305, 1975. CODEN LN-MAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081230/.

Rinzel:1976:NSS

[Rin76] John Rinzel. Nerve signaling and spatial stability of wave trains. Lecture Notes in Mathematics, 525:127–142, 1976. CODEN LN-MAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077849/.

Risler:1976:RRF

[Ris76] J. J. Risler. Résultats récents sur les fonctions de Nash. (French) []. Lecture Notes in Mathematics, 524:79–89, 1976. CODEN LN-MAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077999/.

Ritter:1976:FRA

[Rit76] Gunter Ritter. Famille résolvante associée à des opérateurs ne verifiant pas le principe du maximum. (French) []. Lecture Notes in Mathematics, 518:184–194, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080411/.

Ryll-Nardzewski:1975:TET

[RN75] C. Ryll-Nardzewski. Topics in ergodic theory. Lecture Notes in Mathematics, 472:131–156, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081951/.

Roache:1975:RDP

[Roa75] Patrick J. Roache. Recent developments and problem areas in computational fluid dynamics. Lecture Notes in Mathematics, 461: 195–256, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074155/.

Roberts:1975:AEC

[Rob75a] Leigh Roberts. All erections of a combinatorial geometry and their automorphism groups. Lecture Notes in Mathematics, 452: 210–213, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069559/.

Robinson:1975:SSD

[Rob75b] R. Clark Robinson. Structural stability for C^{1} diffeomorphisms. Lecture Notes in Mathematics, 468:21–23, 1975. CODEN LN-MAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082585/.

Roberts:1976:TSR

[Rob76a] Leslie G. Roberts. The K-theory of some reducible affine curves: a combinatorial approach. Lecture Notes in Mathematics, 551: 44–59, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079993/.

Robinson:1976:ISG

[Rob76b] D. F. Robinson. Integer sequences with given sum and restricted differences. Lecture Notes in Mathematics, 560:185–192, 1976. CODEN LNMAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097380/.

Robinson:1976:ODO

[Rob76c] Peter J. Robinson. Orthogonal designs in order sixteen. Lecture Notes in Mathematics, 560:193–197, 1976. CODEN LN-MAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097381/.

Robinson:1976:CAB

[Rob76d] Robert W. Robinson. Counting arrangements of bishops. Lecture Notes in Mathematics, 560:198–214, 1976. CODEN LN-MAA2. ISBN 3-540-08053-8 (print), 3-540-37537-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0097382/.

Robson:1976:CFO

[Rob76e] J. C. Robson. Cyclic and faithful objects in quotient categories with applications to Noetherian simple or Asano rings. Lecture Notes in Mathematics, 545:151–172, 1976. CODEN LN-MAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080310/.

Roberts:1977:NCS

[Rob77a] James W. Roberts. A nonlocally convex F -space with the Hahn– Banach approximation property. Lecture Notes in Mathematics, 604:76–81, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069209/.

Robinson:1977:ODO

[Rob77b] Peter J. Robinson. Orthogonal designs in order 24. Lecture Notes in Mathematics, 622:170–174, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069191/.

Robinson:1977:CUA

[Rob77c] R. W. Robinson. Counting unlabeled acyclic digraphs. Lecture Notes in Mathematics, 622:28–43, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069178/.

Robson:1977:QCW

[Rob77d] J. C. Robson. Quotient categories and Weyl algebras. Lecture Notes in Mathematics, 586:101–109, 1977. CODEN LN-MAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087126/.

Robinson:1978:ICL

[Rob78] Clark Robinson. Introduction to the closing lemma. Lecture Notes in Mathematics, 668:225–230, 1978. CODEN LNMAA2. ISBN 3-540-08925-X (print), 3-540-35751-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101794/.

Roberts:1979:STP

[Rob79a] Leslie G. Roberts. K_{2} of some truncated polynomial rings. Lecture Notes in Mathematics, 734:249–278, 1979. CODEN LN-MAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103163/.

Robinson:1979:DIP

[Rob79b] D. F. Robinson. Decomposition of integral pseudometrics. Lecture Notes in Mathematics, 748:112–118, 1979. CODEN LN-MAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0102689/.

Robinson:1979:IFV

[Rob79c] R. W. Robinson. Isomorphic factorisations VI: Automorphisms. Lecture Notes in Mathematics, 748:127–136, 1979. CODEN LN-MAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0102691/.

Rockland:1975:EPC

[Roc75a] Prof. Charles Rockland. Example: Poincaré complexes (¯∂_{b}). (French) []. Lecture Notes in Mathematics, 464:137–151, 1975. CODEN LNMAA2. ISBN 3-540-07175-X (print), 3-540-37511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096099/.

Rockland:1975:HAE

[Roc75b] Prof. Charles Rockland. Hypoellipticity and asymptotic eigenvalues in the abstract case. Lecture Notes in Mathematics, 464: 152–165, 1975. CODEN LNMAA2. ISBN 3-540-07175-X (print), 3-540-37511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096100/.

Rockland:1975:HEA

[Roc75c] Prof. Charles Rockland. Hypoellipticity and Eigenvalue Asymptotics, volume 464 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07175-X (print), 3-540-37511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 171 pp. LCCN QA3 .L28 no. 464. URL http:

//link.springer.com/book/10.1007/BFb0096096;   http://
www.springerlink.com/content/978-3-540-37511-1.

Rockland:1975:HLO

[Roc75d] Prof. Charles Rockland. Hypoellipticity with loss of one derivative. Lecture Notes in Mathematics, 464:12–136, 1975. CODEN LNMAA2. ISBN 3-540-07175-X (print), 3-540-37511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096098/.

Rockland:1975:I

[Roc75e] Prof. Charles Rockland. Introduction. Lecture Notes in Mathematics, 464:1–11, 1975. CODEN LNMAA2. ISBN 3-540-07175-X (print), 3-540-37511-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096097/.

Rockafellar:1976:IFN

[Roc76] R. Tyrrell Rockafellar. Integral functionals, normal integrands and measurable selections. Lecture Notes in Mathematics, 543: 157–207, 1976. CODEN LNMAA2. ISBN 3-540-07867-3 (print), 3-540-38075-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079944/.

Rochberg:1977:BMD

[Roc77] Richard Rochberg. The Banach–Mazur distance between function algebras on degenerating Riemann surfaces. Lecture Notes in Mathematics, 604:82–93, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069210/.

Rockafellar:1978:DOC

[Roc78] R. T. Rockafellar. Duality in optimal control. Lecture Notes in Mathematics, 680:219–257, 1978. CODEN LNMAA2. ISBN 3-540-08941-1 (print), 3-540-35714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065318/.

Roche:1979:CEM

[Roc79] J. R. Roche. Sur le calcul de l’exponentielle d’une matrice. (French) []. Lecture Notes in Mathematics, 765:246–256, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085584/.

Rodier:1975:MWC

[Rod75] F. Rodier. Modèle de Whittaker et caractères de représentations. (French) []. Lecture Notes in Mathematics, 466:151–171, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082205/.

Rodier:1977:DSR

[Rod77] François Rodier. Decomposition spectrale des Représentations lisses. (French) []. Lecture Notes in Mathematics, 587:177–195, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087921/.

Rodger:1979:FWS

[Rod79] Christopher A. Rodger. A family of weakly self-dual codes. Lecture Notes in Mathematics, 748:137–142, 1979. CODEN LN-MAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0102692/.

Roger:1975:CDF

[Rog75a] Cl. Roger. Caractères différentiels. (French) [Differential characters]. Lecture Notes in Mathematics, 484:162–178, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082150/.

Roggenkamp:1975:AIF

[Rog75b] Prof. Dr. K. W. Roggenkamp. The augmentation ideal of a finite group, an interesting module. Lecture Notes in Mathematics, 488:306–315, 1975. CODEN LNMAA2. ISBN 3-540-07406-6 (print), 3-540-37921-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081231/.

Rogers:1976:AGB

[Rog76a] Colin Rogers. On applications of generalized Bäcklund transformations to continuum mechanics. Lecture Notes in Mathematics, 515:106–135, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081165/.

Roguski:1976:EMZ

[Rog76b] S. Roguski. Extensions of models for ZFC to models for ZF + V = HOD with applications. Lecture Notes in Mathematics, 537:241–247, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096904/.

Rogers:1977:STT

[Rog77a] D. G. Rogers. A Schröder triangle: Three combinatorial problems. Lecture Notes in Mathematics, 622:175–196, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069192/.

Rogers:1977:STA

[Rog77b] Laurel Rogers. The structure of p-trees: Algebraic systems related to abelian groups. Lecture Notes in Mathematics, 616:57–72, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068189/.

Roguski:1977:TCH

[Rog77c] Stanisław Roguski. The theory of the class HOD. Lecture Notes in Mathematics, 619:251–255, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067656/.

Rogers:1979:ARS

[Rog79a] D. G. Rogers. An application of renewal sequences to the dimer problem. Lecture Notes in Mathematics, 748:143–153, 1979. CO-DEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0102693/.

Roggenkamp:1979:IRP

[Rog79b] K. W. Roggenkamp. Integral representations and presentations of finite groups. Lecture Notes in Mathematics, 744:145–275, 1979. CODEN LNMAA2. ISBN 3-540-09546-2 (print), 3-540-35007-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063060/.

Ronga:1975:SLA

[Ron75] F. Ronga. Stabilité locale des applications équivariantes. Lecture Notes in Mathematics, 484:23–35, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082142/.

Roos:1978:CNM

[Roo78a] Guy Roos. Cocycles de noyaux de Martinelli. (French) []. Lecture Notes in Mathematics, 670:365–369, 1978. CODEN LN-MAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064410/.

Roos:1978:FIP

[Roo78b] Guy Roos. Formules intégrales pour les formes différentielles sur C^{n}, II. (French) []. Lecture Notes in Mathematics, 670:31–52, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064393/.

Roos:1978:RUN

[Roo78c] Guy Roos. Recollement universel de noyaux. (French) []. Lecture Notes in Mathematics, 670:22–30, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064392/.

Roos:1979:RBP

[Roo79] Jan-Erik Roos. Relations between the Poincaré–Betti series of loop spaces and of local rings. Lecture Notes in Mathematics, 740: 285–322, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071068/.

Roquette:1975:NAH

[Roq75] Peter Roquette. Nonstandard aspects of Hilbert’s irreducibility theorem. Lecture Notes in Mathematics, 498:231–275, 1975. CO-DEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0080983/.

Ross:1975:BHE

[Ros75a] Bertram Ross. A brief history and exposition of the fundamental theory of fractional calculus. Lecture Notes in Mathematics, 457:1–36, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067096/.

Ross:1975:FCA

[Ros75b] Prof. Dr. Bertram Ross, editor. Fractional Calculus and Its Applications: Proceedings of the International Conference Held at the University of New Haven, June 1974, volume 457 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 457. URL http://link.springer.com/book/10.1007/
BFb0067095;   http://www.springerlink.com/content/978-
3-540-69975-0.

Rosinski:1976:IPB

[Ros76a] J. Rosiński. Invariance principle for Banach space valued random variables and under random partitions. Lecture Notes in Mathematics, 526:211–219, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082355/.

Rost:1976:SST

[Ros76b] H. Rost. Skorokhod stopping times of minimal variance. Lecture Notes in Mathematics, 511:194–208, 1976. CODEN LN-MAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101107/.

Rosenberg:1977:DDC

[Ros77] Harold Rosenberg. Les difféomorphismes du cercle [d’après M. R. Herman]. (French) []. Lecture Notes in Mathematics, 567:81– 98, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096063/.

Rosebrugh:1978:CAI

[Ros78a] Robert Rosebrugh. Coequalizers in algebras for an internal type. Lecture Notes in Mathematics, 661:243–260, 1978. CODEN LN-MAA2. ISBN 3-540-08914-4 (print), 3-540-35762-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061364/.

Rosinger:1978:ACA

[Ros78b] Elemer E. Rosinger. Associative, commutative algebras containing the distributions. Lecture Notes in Mathematics, 684:3–32, 1978. CODEN LNMAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067793/.

Rosinger:1978:DAC

[Ros78c] Elemer E. Rosinger. Dirac algebras containing the distributions. Lecture Notes in Mathematics, 684:33–59, 1978. CODEN LN-MAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067794/.

Rosinger:1978:DNP

[Ros78d] Elemer E. Rosinger. Distributions and Nonlinear Partial Differential Equations, volume 684 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 146 pp. LCCN QA3

.L28 no. 684. URL http://link.springer.com/book/10.1007/
BFb0067792;   http://www.springerlink.com/content/978-
3-540-35707-0.

Rosinger:1978:LIF

[Ros78e] Elemer E. Rosinger. Linear independent families of Dirac distributions. Lecture Notes in Mathematics, 684:111–119, 1978. CODEN LNMAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067798/.

Rosinger:1978:NSD

[Ros78f] Elemer E. Rosinger. Necessary structure of the distribution multiplications. Lecture Notes in Mathematics, 684:132–138, 1978. CODEN LNMAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067800/.

Rosinger:1978:PDD

[Ros78g] Elemer E. Rosinger. Products with Dirac distributions. Lecture Notes in Mathematics, 684:85–110, 1978. CODEN LN-MAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067797/.

Rosinger:1978:QPS

[Ros78h] Elemer E. Rosinger. Quantum particle scattering in potentials positive powers of the Dirac δ distribution. Lecture Notes in Mathematics, 684:70–84, 1978. CODEN LNMAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067796/.

Rosinger:1978:SNP

[Ros78i] Elemer E. Rosinger. Solutions of nonlinear partial differential equations application to nonlinear shock waves. Lecture Notes in Mathematics, 684:60–69, 1978. CODEN LNMAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067795/.

Rosinger:1978:SLP

[Ros78j] Elemer E. Rosinger. Support, local properties. Lecture Notes in Mathematics, 684:120–131, 1978. CODEN LNMAA2. ISBN 3-540-08951-9 (print), 3-540-35707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067799/.

Rosenblatt:1979:GMD

[Ros79] M. Rosenblatt. Global measures of deviation for kernel and nearest neighbor density estimates. Lecture Notes in Mathematics, 757: 181–190, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098496/.

Roth:1976:AOD

[Rot76a] Jean-Pierre Roth. Approximation des opérateurs dissipatifs. (French) []. Lecture Notes in Mathematics, 518:195–201, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080412/.

Roth:1976:FRT

[Rot76b] Jean-Pierre Roth. Formule de représentation et troncature des formes de Dirichlet sur R^{m}. (French) []. Lecture Notes in Mathematics, 563:260–274, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087582/.

Roth:1976:SGI

[Rot76c] Jean-Pierre Roth. Semi-groupes invariants sur un espace homogène. (French) []. Lecture Notes in Mathematics, 518:202– 213, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080413/.

Roth:1978:OEC

[Rot78] J.-P. Roth. Les opérateurs elliptiques comme générateurs infinitesimaux de semi-groupes de Feller. (French) []. Lecture Notes in Mathematics, 681:234–251, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065879/.

Roussarie:1978:CFD

[Rou78] Robert Roussarie. Constructions de feuilletages [d’après W. Thurston]. (French) []. Lecture Notes in Mathematics, 677:138– 154, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070759/.

Rourke:1979:PTG

[Rou79] C. P. Rourke. Presentations and the trivial group. Lecture Notes in Mathematics, 722:134–143, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063195/.

Roynette:1975:MAG

[Roy75] B. Roynette. Marches aléatoires sur les groupes d’Heisenberg. (French) []. Lecture Notes in Mathematics, 497:368–382, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078023/.

Royer:1976:PAD

[Roy76a] G. Royer. I. Propriété asymptotique des distributions de Wightman. (French) []. Lecture Notes in Mathematics, 518:214–227, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080414/.

Royer:1976:ITD

[Roy76b] G. Royer. II. Théorie de la diffusion de Haag–Ruelle. (French) []. Lecture Notes in Mathematics, 518:228–253, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080415/.

Roynette:1978:MAG

[Roy78] B. Roynette. Marches aléatoires sur les groupes de Lie. (French) []. Lecture Notes in Mathematics, 678:237–379, 1978. CODEN LNMAA2. ISBN 3-540-08938-1 (print), 3-540-35716-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062799/.

Reimann:1975:DG

[RR75a] Prof. Hans Martin Reimann and Prof. Thomas Rychener. Der Dualitaetssatz 0. (German) []. Lecture Notes in Mathematics, 487: 67–95, 1975. CODEN LNMAA2. ISBN 3-540-07404-X (print), 3-540-37917-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081829/.

Reimann:1975:SJN

[RR75b] Prof. Hans Martin Reimann and Prof. Thomas Rychener. Die Saetze von John–Nirenberg und Gehring. (German) []. Lecture Notes in Mathematics, 487:28–45, 1975. CODEN LNMAA2. ISBN 3-540-07404-X (print), 3-540-37917-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081827/.

Reimann:1975:FBMa

[RR75c] Prof. Hans Martin Reimann and Prof. Thomas Rychener. Funktionen Beschraenkter Mittlerer Oszillation. (German) []. Lecture Notes in Mathematics, 487:1–27, 1975. CODEN LNMAA2. ISBN 3-540-07404-X (print), 3-540-37917-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081826/.

Reimann:1975:FBMb

[RR75d] Prof. Hans Martin Reimann and Prof. Thomas Rychener. Funktionen beschränkter mittlerer Oszillation, volume 487 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07404-X (print), 3-540-37917-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 141 pp. LCCN QA3

.L28 no. 487. URL http://link.springer.com/book/10.1007/

BFb0081825;   http://www.springerlink.com/content/978-
3-540-37917-1.

Reimann:1975:MBG

[RR75e] Prof. Hans Martin Reimann and Prof. Thomas Rychener. Muckenhoupts A_{p}-Bedingung. (German) []. Lecture Notes in Mathematics, 487:46–66, 1975. CODEN LNMAA2. ISBN 3-540-07404-X (print), 3-540-37917-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081828/.

Reimann:1975:QAG

[RR75f] Prof. Hans Martin Reimann and Prof. Thomas Rychener. Quasikonforme Abbildungen. (German) [Quasi-conformal mapping]. Lecture Notes in Mathematics, 487:96–108, 1975. CODEN LNMAA2. ISBN 3-540-07404-X (print), 3-540-37917-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081830/.

Reimann:1975:RPB

[RR75g] Prof. Hans Martin Reimann and Prof. Thomas Rychener. Riesz-Potentiale von BMO-Funktionen 0. (German) []. Lecture Notes in Mathematics, 487:109–133, 1975. CODEN LNMAA2. ISBN 3-540-07404-X (print), 3-540-37917-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081831/.

Rahbar-Rochandel:1978:RES

[RR78] Hamid Rahbar-Rochandel. Relations entre la série de Betti d’un anneau local de Gorenstein R et celle de l’anneau R/socle R. (French) [Relations between the Betti series of a local Gorenstein ring R and that of the ring R /base R]. Lecture Notes in Mathematics, 641:339–357, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064855/.

Rahbar-Rochandel:1979:CDA

[RR79a] H. Rahbar-Rochandel. La cinquième déflection d’un anneau local noethérien. (French) []. Lecture Notes in Mathematics, 740: 205–212, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071062/.

Reiner:1979:IRT

[RR79b] Irving Reiner and Klaus W. Roggenkamp. Integral Representations: Topics in Integral Representation Theory by I. Reiner. Integral Representations and Presentations of Finite Groups by K. W. Roggenkamp, volume 744 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09546-2 (print), 3-540-35007-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 275 pp. LCCN QA3

.L28 no. 744. URL http://link.springer.com/book/10.1007/
BFb0063058;   http://www.springerlink.com/content/978-
3-540-35007-1.

Rallis:1975:DIP

[RS75] S. Rallis and G. Schiffmann. Distributions invariantes par le groupe orthogonal. (French) []. Lecture Notes in Mathematics, 497:494–642, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078025/.

Rehmann:1976:FPG

[RS76] U. Rehmann and C. Soulé. Finitely presented groups of matrices. Lecture Notes in Mathematics, 551:164–169, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079999/.

Reeb:1978:TTE

[RS78] G. Reeb and P. Schweitzer. Un théorème de Thurston etabli au moyen de l’analyse non standard. (French) []. Lecture Notes in Mathematics, 652:138, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063509.

Raviart:1977:MFE

[RT77] P.-A. Raviart and J. M. Thomas. A mixed finite element method for 2-nd order elliptic problems. Lecture Notes in Mathematics, 606:292–315, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064470/.

Rees:1978:CCI

[RT78] Elmer Rees and Emery Thomas. Complex Cobordism and intersections of projective varieties. Lecture Notes in Mathematics, 683:168–178, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063177/.

Rubenthaler:1975:DBI

[Rub75] Hubert Rubenthaler. Distributions bi-invariantes par SL_{n}(k). (French) [bi-invariant distributions SL_{n}(k)]. Lecture Notes in Mathematics, 497:383–493, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0078024/.

Rubenthaler:1979:SDR

[Rub79] Hubert Rubenthaler. Une série dégénérée de représentations de SL_{n}(R). (French) []. Lecture Notes in Mathematics, 739:427– 459, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062503/.

Ruckebusch:1978:TMR

[Ruc78a] G. Ruckebusch. On the theory of Markovian representation. Lecture Notes in Mathematics, 695:77–87, 1978. CODEN LN-MAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062656/.

Ruckebusch:1978:RMP

[Ruc78b] G. Ruckebusch. Représentations markoviennes de processus gaussiens stationnaires et applications statistiques. (French) []. Lecture Notes in Mathematics, 636:115–139, 1978. CODEN LN-MAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063262/.

Rudin:1977:NVS

[Rud77a] M. E. Rudin. A narrow view of set theoretic topology. Lecture Notes in Mathematics, 609:190–195, 1977. CODEN LN-MAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068683/.

Rudin:1977:HMD

[Rud77b] Walter Rudin. Holomorphic maps of discs into F -spaces. Lecture Notes in Mathematics, 599:104–108, 1977. CODEN LN-MAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096831/.

Ruelle:1976:LAP

[Rue76] David Ruelle. The Lorenz attractor and the problem of turbulence. Lecture Notes in Mathematics, 565:146–158, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091453/.

Ruelle:1977:FTF

[Rue77] David Ruelle. Formalisme thermodynamique. (French) [Thermodynamic formalism]. Lecture Notes in Mathematics, 567:154– 166, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096067/.

Ruhe:1977:CEE

[Ruh77] Axel Ruhe. Computation of eigenvalues and eigenvectors. Lecture Notes in Mathematics, 572:130–184, 1977. CODEN LN-MAA2. ISBN 3-540-08130-5 (print), 3-540-37430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0116617/.

Rumeur:1979:TIB

[Rum79] Huguette Rumeur. Théorème d’induction de Brauer et application au calcul des caractères modulaires. (French) []. Lecture Notes in Mathematics, 740:25–47, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071051/.

Rund:1976:VPB

[Run76] Hanno Rund. Variational problems and Bäcklund transformations associated with the sine-Gordon and Korteweg–de Vries equations and their extensions. Lecture Notes in Mathematics, 515:199– 226, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081169/.

Russo:1977:OTH

[Rus77] Bernard Russo. Operator theory in harmonic analysis. Lecture Notes in Mathematics, 604:94–102, 1977. CODEN LN-MAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069211/.

Rubel:1977:VSE

[RY77] L. A. Rubel and Chung-Chun Yang. Values shared by an entire function and its derivative. Lecture Notes in Mathematics, 599:101–103, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096830/.

Rychlik:1978:OAR

[Ryc78] Z. Rychlik. The order of approximation in the random central limit theorem. Lecture Notes in Mathematics, 656:225–236, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068821/.

Smolska-Adamowicz:1977:CSL

[SA77] Zofia Smólska-Adamowicz. Constructible semi-lattices of degrees of constructibility. Lecture Notes in Mathematics, 619:1–43, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067640/.

Sabbagh:1976:CAG

[Sab76] Gabriel Sabbagh. Caractérisation algébrique des groupes de type fini ayant un problème de mots résoluble. (French) []. Lecture Notes in Mathematics, 514:61–80, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080058/.

Sacerdote:1975:PMT

[Sac75] George S. Sacerdote. Projective model theory and coforcing. Lecture Notes in Mathematics, 498:276–306, 1975. CODEN LN-MAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080984/.

Saffman:1977:LIP

[Saf77] P. G. Saffman. Lecture IV a phenomenonological theory for the computation of turbulent shear flows. Lecture Notes in Mathematics, 615:71–82, 1977. CODEN LNMAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068361/.

Saito:1978:AAM

[Sai78] Kazuyuki Saitô. AW*-algebras with monotone convergence property and type III, non w*, AW*-factors. Lecture Notes in Mathematics, 650:131–134, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067394/.

Saito:1979:ABR

[Sai79a] Yoshimi Saitō. Asymptotic behavior of the radiative function. Lecture Notes in Mathematics, 727:53–99, 1979. CODEN LN-MAA2. ISBN 3-540-09514-4 (print), 3-540-35132-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0071000/.

Saito:1979:I

[Sai79b] Yoshimi Saitō. Introduction. Lecture Notes in Mathematics, 727: 1–9, 1979. CODEN LNMAA2. ISBN 3-540-09514-4 (print), 3-540-35132-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070998/.

Saito:1979:LAP

[Sai79c] Yoshimi Saitō. Limiting absorption principle. Lecture Notes in Mathematics, 727:10–52, 1979. CODEN LNMAA2. ISBN 3-540-09514-4 (print), 3-540-35132-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070999/.

Saito:1979:SR

[Sai79d] Yoshimi Saitō. Spectral representation. Lecture Notes in Mathematics, 727:100–134, 1979. CODEN LNMAA2. ISBN 3-540-09514-4 (print), 3-540-35132-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071001/.

Saito:1979:SRS

[Sai79e] Yoshimi Saitō. Spectral Representations for Schrödinger Operators With Long-Range Potentials, volume 727 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09514-4 (print), 3-540-35132-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 134 (est.) pp. LCCN QA3

.L28 no. 727. URL http://link.springer.com/book/10.1007/
BFb0070997;   http://www.springerlink.com/content/978-
3-540-35132-0.

Sakarovitch:1977:GIC

[Sak77] J. Sakarovitch. Sur les groupes infinis, considérés comme monoïdes syntaxiques de langages formels. (French) []. Lecture Notes in Mathematics, 586:168–179, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087132/.

Sakai:1978:RDT

[Sak78] Shôichirô Sakai. Recent developments in the theory of unbounded derivations in C∗-algebras. Lecture Notes in Mathematics, 650: 85–122, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067391/.

Salinas:1978:CPE

[Sal78] Norberto Salinas. A classification problem for essentially n-normal operators. Lecture Notes in Mathematics, 693:145–156, 1978. CO-DEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0064672/.

Samuelides:1976:AVC

[Sam76] M. Samuélidès. Analyticité vectorielle et champ scalaire neutre. (French) []. Lecture Notes in Mathematics, 518:254–265, 1976. CODEN LNMAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080416/.

Sanderson:1979:GMO

[San79] Brian J. Sanderson. The geometry of Mahowald orientations. Lecture Notes in Mathematics, 763:152–174, 1979. CODEN LN-MAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088085/.

Sarason:1976:AB

[Sar76a] Donald Sarason. Algebras between L^{∞} and H^{∞}. Lecture Notes in Mathematics, 512:117–130, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080029/.

Sard:1976:OA

[Sar76b] Arthur Sard. On optimal approximation. Lecture Notes in Mathematics, 501:220–225, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079750/.

Sard:1977:FSA

[Sar77] Arthur Sard. Function spaces for analysis. Lecture Notes in Mathematics, 571:186–195, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086574/.

Sargent:1978:DSP

[Sar78] R. W. H. Sargent. The decomposition of systems of procedures and algebraic equations. Lecture Notes in Mathematics, 630:158– 178, 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067704/.

Sato:1976:BSP

[Sat76a] Hiroshi Satô. Banach support of a probability measure in a locally convex space. Lecture Notes in Mathematics, 526:221–226, 1976. CODEN LNMAA2. ISBN 3-540-07793-6 (print), 3-540-38256-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082356/.

Sattinger:1976:PFC

[Sat76b] D. H. Sattinger. Pattern formation in convective phenomena. Lecture Notes in Mathematics, 565:159–173, 1976. CODEN LN-MAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091454/.

Sato:1979:HSP

[Sat79a] Hiroshi Satô. Hilbertian support of a probability measure on a Banach space. Lecture Notes in Mathematics, 709:195–205, 1979. CODEN LNMAA2. ISBN 3-540-09242-0 (print), 3-540-35341-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071958/.

Sattinger:1979:A

[Sat79b] D. H. Sattinger. Applications. Lecture Notes in Mathematics, 762: 151–199, 1979. CODEN LNMAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087462/.

Sattinger:1979:BME

[Sat79c] D. H. Sattinger. Bifurcation at multiple eigenvalues. Lecture Notes in Mathematics, 762:70–95, 1979. CODEN LNMAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087460/.

Sattinger:1979:EGR

[Sat79d] D. H. Sattinger. Elements of group representation theory. Lecture Notes in Mathematics, 762:96–150, 1979. CODEN LN-MAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087461/.

Sattinger:1979:MP

[Sat79e] D. H. Sattinger. Mathematical preliminaries. Lecture Notes in Mathematics, 762:18–36, 1979. CODEN LNMAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087458/.

Sattinger:1979:PEB

[Sat79f] D. H. Sattinger. Physical examples of bifurcation. Lecture Notes in Mathematics, 762:1–17, 1979. CODEN LNMAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087457/.

Sattinger:1979:SB

[Sat79g] D. H. Sattinger. Stability and bifurcation. Lecture Notes in Mathematics, 762:37–69, 1979. CODEN LNMAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087459/.

Sattinger:1979:GTM

[Sat79h] David H. Sattinger. Group Theoretic Methods in Bifurcation Theory, volume 762 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09715-5 (print), 3-540-38487-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 241 pp. LCCN QA3 .L28 no. 762. URL http:

//link.springer.com/book/10.1007/BFb0087456;   http://
www.springerlink.com/content/978-3-540-38487-8.

Sawyer:1975:ABR

[Saw75] Stanley Sawyer. An application of branching random fields to genetics. Lecture Notes in Mathematics, 451:100–112, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068582/.

Saxl:1977:RCG

[Sax77] Jan Saxl. Restrictions of characters, generosity, interchange and coloured graphs. Lecture Notes in Mathematics, 579:258– 266, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090021/.

Sander:1977:ICC

[SB77] G. Sander and P. Beckers. The influence of the choice of connectors in the finite element method. Lecture Notes in Mathematics, 606:316–342, 1977. CODEN LNMAA2. ISBN 3-540-08432-0 (print), 3-540-37158-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064471/.

Sbordone:1979:SCP

[Sbo79] Carlo Sbordone. On some convergence problems for quasiconformal mappings. Lecture Notes in Mathematics, 743:223–237, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079496/.

Chern:1977:CB

[sC77] Shiing shen Chern. Circle bundles. Lecture Notes in Mathematics, 597:114–131, 1977. CODEN LNMAA2. ISBN 3-540-08345-6 (print), 3-540-37301-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085351/.

Scarpellini:1975:BRS

[Sca75] B. Scarpellini. Bemerkungen zu Regel und Schema. (German) []. Lecture Notes in Mathematics, 500:320–340, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079560/.

Schmerl:1975:NEC

[Sch75a] James H. Schmerl. The number of equivalence classes of existentially complete structures. Lecture Notes in Mathematics, 498: 170–171, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080979.

Schmid:1975:SRA

[Sch75b] Wilfried Schmid. Some remarks about the discrete series characters of Sp(n, R). Lecture Notes in Mathematics, 466:172–194, 1975.

CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082206/.

Schmitt:1975:EPN

[Sch75c] Klaus Schmitt. Eigenvalue problems for nonlinear second order differential equations. Lecture Notes in Mathematics, 448:294– 306, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067092/.

Schmitt:1975:LUD

[Sch75d] Peter Schmitt. Linear uniform distribution. Lecture Notes in Mathematics, 475:245–252, 1975. CODEN LNMAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074266/.

Schober:1975:ASB

[Sch75e] Prof. Glenn Schober. Application of Schiffer’s boundary variation to linear problems. Lecture Notes in Mathematics, 478:92– 111, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077289/.

Schober:1975:AFC

[Sch75f] Prof. Glenn Schober. Application to families of conformal and q.c. mappings. Lecture Notes in Mathematics, 478:147–167, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0077293/.

Schober:1975:ASN

[Sch75g] Prof. Glenn Schober. Application to some nonlinear problems. Lecture Notes in Mathematics, 478:112–127, 1975. CODEN LN-MAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077290/.

Schober:1975:CFU

[Sch75h] Prof. Glenn Schober. Compact families F(D, ℓ_{∞}, ℓ_{∈}, P, Q) of univalent functions normalized by two linear functionals. Lecture Notes in Mathematics, 478:57–64, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077286/.

Schober:1975:EVM

[Sch75i] Prof. Glenn Schober. Elementary variational methods. Lecture Notes in Mathematics, 478:79–91, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077288/.

Schober:1975:ELE

[Sch75j] Prof. Glenn Schober. Extremal length and equicontinuity. Lecture Notes in Mathematics, 478:48–56, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077285/.

Schober:1975:FP

[Sch75k] Prof. Glenn Schober. Faber polynomials. Lecture Notes in Mathematics, 478:39–47, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077284/.

Schober:1975:FPR

[Sch75l] Prof. Glenn Schober. Functions with positive real part. Lecture Notes in Mathematics, 478:1–5, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077280/.

Schober:1975:PSC

[Sch75m] Prof. Glenn Schober. The Pólya–Schoenberg conjecture. Lecture Notes in Mathematics, 478:27–33, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077282/.

Schober:1975:PEP

[Sch75n] Prof. Glenn Schober. Properties of extreme points for some compact families F(D, ℓ_{∞}, ℓ_{∈}, P, Q). Lecture Notes in Mathematics, 478:65–78, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077287/.

Schober:1975:RCL

[Sch75o] Prof. Glenn Schober. Representation of continuous linear functionals. Lecture Notes in Mathematics, 478:34–38, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077283/.

Schober:1975:SPQ

[Sch75p] Prof. Glenn Schober. Some properties of quasiconformal mappings. Lecture Notes in Mathematics, 478:128–137, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077291/.

Schober:1975:SCC

[Sch75q] Prof. Glenn Schober. Special classes: convex, starlike, real, typically real, close-to-convex, bounded boundary rotation. Lecture Notes in Mathematics, 478:6–26, 1975. CODEN LNMAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077281/.

Schober:1975:SCQ

[Sch75r] Prof. Glenn Schober. Sufficient conditions for q.c. extensions. Lecture Notes in Mathematics, 478:168–171, 1975. CODEN LN-MAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077294/.

Schober:1975:UFS

[Sch75s] Prof. Glenn Schober. Univalent Functions — Selected Topics, volume 478 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 0-387-07391-4, 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 200 pp. LCCN QA3 .L28 no. 478. URL http:

//link.springer.com/book/10.1007/BFb0077279;   http://
www.springerlink.com/content/978-3-540-37587-6.

Schober:1975:VMQ

[Sch75t] Prof. Glenn Schober. A variational method for q.c. mappings. Lecture Notes in Mathematics, 478:138–146, 1975. CODEN LN-MAA2. ISBN 3-540-07391-4 (print), 3-540-37587-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077292/.

Schweitzer:1975:CLC

[Sch75u] Paul A. Schweitzer, S.J. Compact leaves of codimension one foliations. Lecture Notes in Mathematics, 484:273–276, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082158/.

Schweitzer:1975:CLF

[Sch75v] Paul A. Schweitzer, S.J. Compact leaves of foliations. Lecture Notes in Mathematics, 468:4–6, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082575/.

Schachermayer:1976:TGF

[Sch76a] M. Walter Schachermayer. Sur un théorème de Grothendieck. (French) []. Lecture Notes in Mathematics, 524:193–212, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0078006/.

Scheffer:1976:THD

[Sch76b] Vladimir Scheffer. Turbulence and Hausdorff dimension. Lecture Notes in Mathematics, 565:174–183, 1976. CODEN LN-MAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091455/.

Schlosser:1976:MSI

[Sch76c] Dr. Karl-Heinz Schloßer. Mehrdimensionale Spline-Interpolation mit Hilfe der Methode von Sard. (German) []. Lecture Notes in Mathematics, 501:247–264, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079752/.

Schmets:1976:AEF

[Sch76d] Jean Schmets. Application aux espaces de fonctions continues vectorielles. (French) []. Lecture Notes in Mathematics, 519:120– 142, 1976. CODEN LNMAA2. ISBN 3-540-07694-8 (print), 3-540-38226-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079251/.

Schmets:1976:CSC

[Sch76e] Jean Schmets. Conditions de séparabilité et de compacité faible. (French) []. Lecture Notes in Mathematics, 519:99–119, 1976. CO-DEN LNMAA2. ISBN 3-540-07694-8 (print), 3-540-38226-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079250/.

Schmets:1976:EAEa

[Sch76f] Jean Schmets. Espaces associés à un espace linaire à semi-normes. (French) []. Lecture Notes in Mathematics, 519:15–35, 1976. CO-DEN LNMAA2. ISBN 3-540-07694-8 (print), 3-540-38226-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0079247/.

Schmets:1976:EAEb

[Sch76g] Jean Schmets. Espaces associés aux espaces de fonctions continues. (French) []. Lecture Notes in Mathematics, 519:62–98, 1976. CODEN LNMAA2. ISBN 3-540-07694-8 (print), 3-540-38226-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079249/.

Schmets:1976:ECR

[Sch76h] Jean Schmets. Espaces complètement reguliers et séparés et espaces de fonctions continues. (French) []. Lecture Notes in Mathematics, 519:36–61, 1976. CODEN LNMAA2. ISBN 3-540-07694-8 (print), 3-540-38226-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079248/.

Schmets:1976:EFC

[Sch76i] Jean Schmets. Espaces de Fonctions Continues. (French) [Spaces of continuous functions], volume 519 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07694-8 (print), 3-540-38226-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 150 pp. LCCN QA3

.L28 no. 519. URL http://link.springer.com/book/10.1007/
BFb0079245;   http://www.springerlink.com/content/978-
3-540-38226-3.

Schmets:1976:PFP

[Sch76j] Jean Schmets. Préliminaires. (French) [Preliminaries]. Lecture Notes in Mathematics, 519:1–14, 1976. CODEN LNMAA2. ISBN 3-540-07694-8 (print), 3-540-38226-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079246/.

Schmidt:1976:AIEa

[Sch76k] Wolfgang M. Schmidt. Absolutely irreducible equations f(x, y) = 0. Lecture Notes in Mathematics, 536:92–133, 1976. CODEN LNMAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080441/.

Schmidt:1976:AIEb

[Sch76l] Wolfgang M. Schmidt. Absolutely irreducible equations f(x_{1}, …, x_{n}) = 0. Lecture Notes in Mathematics, 536:177–215, 1976. CODEN LNMAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080443/.

Schmidt:1976:CSE

[Sch76m] Wolfgang M. Schmidt. Character sums and exponential sums. Lecture Notes in Mathematics, 536:38–91, 1976. CODEN LN-MAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080440/.

Schmidt:1976:E

[Sch76n] Wolfgang M. Schmidt. Equations y^{d} = f(x) and y^{q} − y = f(x). Lecture Notes in Mathematics, 536:3–37, 1976. CODEN LN-MAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080439/.

Schmidt:1976:EMV

[Sch76o] Wolfgang M. Schmidt. Equations in many variables. Lecture Notes in Mathematics, 536:134–176, 1976. CODEN LNMAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080442/.

Schmidt:1976:EFF

[Sch76p] Wolfgang M. Schmidt. Equations over Finite Fields: An Elementary Approach, volume 536 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 276 pp. LCCN QA3

.L28 no. 536. URL http://link.springer.com/book/10.1007/
BFb0080437;   http://www.springerlink.com/content/978-
3-540-38123-5.

Schmidt:1976:I

[Sch76q] Wolfgang M. Schmidt. Introduction. Lecture Notes in Mathematics, 536:1–2, 1976. CODEN LNMAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080438.

Schmidt:1976:RAG

[Sch76r] Wolfgang M. Schmidt. Rudiments of algebraic geometry. The number of points in varieties over finite fields. Lecture Notes in Mathematics, 536:216–264, 1976. CODEN LNMAA2. ISBN 3-540-07855-X (print), 3-540-38123-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080444/.

Schumaker:1976:TCT

[Sch76s] Larry L. Schumaker. Toward a constructive theory of generalized spline functions. Lecture Notes in Mathematics, 501:265– 331, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079753/.

Scheinberg:1977:CSC

[Sch77a] Stephen Scheinberg. Cluster sets and corona theorems. Lecture Notes in Mathematics, 604:103–106, 1977. CODEN LN-MAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069212/.

Schempp:1977:AHB

[Sch77b] Prof. Dr. Walter Schempp. Approximation in G-homogeneous Banach spaces. Lecture Notes in Mathematics, 571:220–225, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086577/.

Schempp:1977:BPS

[Sch77c] Prof. Dr. Walter Schempp. Bernstein polynomials in several variables. (German) []. Lecture Notes in Mathematics, 571:212– 219, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086576/.

Schmid:1977:IHF

[Sch77d] Hans Joachim Schmid. Interpolation of harmonic functions. Lecture Notes in Mathematics, 571:226–249, 1977. CODEN LN-MAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086578/.

Schmid:1977:TCI

[Sch77e] Wilfried Schmid. Two character identities for semisimple Lie groups. Lecture Notes in Mathematics, 587:196–225, 1977. CO-DEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0087922/.

Schoenberg:1977:ZSD

[Sch77f] I. J. Schoenberg. On the zeros of the successive derivatives of integral functions II. Lecture Notes in Mathematics, 599:109– 116, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096832/.

Schutzenberger:1977:CRF

[Sch77g] M.-P. Schützenberger. La correspondance de Robinson. (French) []. Lecture Notes in Mathematics, 579:59–113, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090012/.

Schwartz:1977:SHE

[Sch77h] M.-H. Schwartz. Sections holomorphes d’espaces à fibres linéaires variables. (French) []. Lecture Notes in Mathematics, 578:280– 299, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091478/.

Schachermayer:1978:AMS

[Sch78a] W. Schachermayer. Addendum: Une modification standard de la démonstration non standard de Reeb et Schweitzer. (French) []. Lecture Notes in Mathematics, 652:139–140, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063510.

Schaeffer:1978:ANM

[Sch78b] David G. Schaeffer. An application of the Nash–Moser theorem to a free boundary problem. Lecture Notes in Mathematics, 648:129–143, 1978. CODEN LNMAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066409/.

Scharlemann:1978:SCM

[Sch78c] Martin Scharlemann. Smooth CE maps and smooth homeomorphisms. Lecture Notes in Mathematics, 664:234–240, 1978. CO-DEN LNMAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0061701/.

Scherer:1978:SSG

[Sch78d] Rudolf Scherer. Spiegelung von Stabilitätsbereichen. (German) []. Lecture Notes in Mathematics, 631:147–152, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067470/.

Schmidt:1978:BSV

[Sch78e] F. Schmidt. Banach-space valued stationary processes with absolutely continuous spectral function. Lecture Notes in Mathematics, 656:237–244, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068822/.

Schultz:1978:ACT

[Sch78f] Phillip Schultz. Annihilator classes of torsion-free abelian groups. Lecture Notes in Mathematics, 697:88–94, 1978. CODEN LN-MAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103124/.

Schweitzer:1978:DTF

[Sch78g] Paul A. Schweitzer, editor. Differential Topology, Foliations and Gelfand–Fuks Cohomology: Proceedings of the Symposium held at the Pontifica Universidade Católica do Rio de Janeiro, 5– 24 January, 1976, volume 652 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 652. URL http:

//link.springer.com/book/10.1007/BFb0063498;   http://
www.springerlink.com/content/978-3-540-38074-0.

Scheunert:1979:FC

[Sch79a] Manfred Scheunert. Formal constructions. Lecture Notes in Mathematics, 716:12–71, 1979. CODEN LNMAA2. ISBN 3-540-09256-0 (print), 3-540-35286-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070932/.

Scheunert:1979:I

[Sch79b] Manfred Scheunert. Introduction. Lecture Notes in Mathematics, 716:1–4, 1979. CODEN LNMAA2. ISBN 3-540-09256-0 (print), 3-540-35286-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070930/.

Scheunert:1979:PR

[Sch79c] Manfred Scheunert. Preparatory remarks. Lecture Notes in Mathematics, 716:5–11, 1979. CODEN LNMAA2. ISBN 3-540-09256-0 (print), 3-540-35286-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070931/.

Scheunert:1979:SLS

[Sch79d] Manfred Scheunert. Simple Lie superalgebras. Lecture Notes in Mathematics, 716:72–230, 1979. CODEN LNMAA2. ISBN 3-540-09256-0 (print), 3-540-35286-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070933/.

Scheunert:1979:SSF

[Sch79e] Manfred Scheunert. A survey of some further developments. Lecture Notes in Mathematics, 716:231–247, 1979. CODEN LN-MAA2. ISBN 3-540-09256-0 (print), 3-540-35286-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070934/.

Scheunert:1979:TLS

[Sch79f] Manfred Scheunert. The Theory of Lie Superalgebras: An Introduction, volume 716 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09256-0 (print), 3-540-35286-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 271 pp. LCCN QA3 .L28 no. 716. URL http:

//link.springer.com/book/10.1007/BFb0070929;   http://
www.springerlink.com/content/978-3-540-35286-0.

Schiffmann:1979:TKS

[Sch79g] G. Schiffmann. Travaux de Kostant sur la série principale. (French) []. Lecture Notes in Mathematics, 739:460–510, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062504/.

Schmidt:1979:UER

[Sch79h] Klaus Schmidt. Unique ergodicity and related problems. Lecture Notes in Mathematics, 729:188–198, 1979. CODEN LN-MAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063294/.

Schultz:1979:ICP

[Sch79i] Reinhard Schultz. Isotopy classes of periodic diffeomorphisms on spheres. Lecture Notes in Mathematics, 741:334–354, 1979. CO-DEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0062148/.

Schweiger:1979:MJP

[Sch79j] F. Schweiger. A modified Jacobi–Perron algorithm with explicitly given invariant measure. Lecture Notes in Mathematics, 729: 199–202, 1979. CODEN LNMAA2. ISBN 3-540-09517-9 (print), 3-540-35130-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063295/.

Scott:1975:DTL

[Sco75] Dana Scott. Data types as lattices. Lecture Notes in Mathematics, 499:579–651, 1975. CODEN LNMAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079432/.

Scott:1976:ABT

[Sco76] Alwyn C. Scott. The application of Bäcklund transforms to physical problems. Lecture Notes in Mathematics, 515:80–105, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081164/.

Scoble:1977:CAM

[Sco77] Jennifer A. Scoble. A combinatorial approach to map theory. Lecture Notes in Mathematics, 622:197–205, 1977. CODEN LN-MAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069193/.

Scott:1979:ABS

[Sco79] Ridgway Scott. Applications of Banach space interpolation to finite element theory. Lecture Notes in Mathematics, 701:298– 318, 1979. CODEN LNMAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062086/.

Swinnerton-Dyer:1977:ARC

[SD77] H. P. F. Swinnerton-Dyer. On ℓ-adic representations and congruences for coefficients of modular forms (II). Lecture Notes in Mathematics, 601:63–90, 1977. CODEN LNMAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063945/.

Swinnerton-Dyer:1975:ECM

[SDB75] H. P. F. Swinnerton-Dyer and B. J. Birch. Ellitpic curves and modular functions. Lecture Notes in Mathematics, 476:2–32, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097581/.

Seberry:1979:SRG

[Seb79] Jennifer Seberry. Some remarks on generalised Hadamard matrices and theorems of Rajkundalia on SBIBDS. Lecture Notes in Mathematics, 748:154–164, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0102694/.

Sec:1975:CEP

[Sec75] Antoinette Sec. Sur certaines équations de Pfaff complètement intégrables dans le champ complexe, (propriétés du feuilletage associé). (French) []. Lecture Notes in Mathematics, 484:224– 233, 1975. CODEN LNMAA2. ISBN 3-540-07405-8 (print), 3-540-37919-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082154/.

Segal:1977:HTA

[Seg77] Graeme Segal. K-homology theory and algebraic K-theory. Lecture Notes in Mathematics, 575:113–127, 1977. CODEN LN-MAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095706/.

Segal:1978:IST

[Seg78] Jack Segal. An introduction to shape theory. Lecture Notes in Mathematics, 673:225–242, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064699/.

Seidel:1975:MPE

[Sei75] Johan J. Seidel. Metric problems in elliptic geometry. Lecture Notes in Mathematics, 490:32–43, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081129/.

Sekiguchi:1976:KDC

[Sek76] Takeshi Sekiguchi. On the Krickeberg decomposition of continuous martingales. Lecture Notes in Mathematics, 511:209–215, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101108/.

Selberg:1977:RMF

[Sel77] Atle Selberg. Remarks on multiplicative functions. Lecture Notes in Mathematics, 626:232–241, 1977. CODEN LNMAA2. ISBN 3-540-08529-7 (print), 3-540-35985-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063067/.

Seneta:1976:FRV

[Sen76a] Eugene Seneta. Functions of regular variation. Lecture Notes in Mathematics, 508:1–52, 1976. CODEN LNMAA2. ISBN 3-540-07618-2 (print), 3-540-38137-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079659/.

Seneta:1976:RVF

[Sen76b] Eugene Seneta. Regularly Varying Functions, volume 508 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07618-2 (print), 3-540-38137-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 112 pp. LCCN QA3

.L28 no. 508. URL http://link.springer.com/book/10.1007/
BFb0079658;   http://www.springerlink.com/content/978-
3-540-38137-2.

Seneta:1976:SST

[Sen76c] Eugene Seneta. Some secondary theory of regularly varying functions. Lecture Notes in Mathematics, 508:53–91, 1976. CODEN LNMAA2. ISBN 3-540-07618-2 (print), 3-540-38137-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079660/.

Sen:1978:BRT

[Sen78] R. N. Sen. Bundle representations and their applications. Lecture Notes in Mathematics, 676:151–160, 1978. CODEN LN-MAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063668/.

Serre:1975:ALM

[Ser75a] Jean-Pierre Serre, editor. Algèbre locale, multiplicités; cours au Collège de France, 1957–1958. (French) [Local algebra, multiplicities; courses at the Collège of France, 1957–1958], volume 11 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CO-DEN LNMAA2. ISBN 0-387-07028-1, 3-540-07028-1 (paperback), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 11; QA3 .L28 no.

11, 1975 QA564. URL http://link.springer.com/book/10.
1007/978-3-540-37123-6;   http://www.springerlink.com/
content/978-3-540-37123-6.

Serre:1975:BMa

[Ser75b] Jean-Pierre Serre. Back matter. Lecture Notes in Mathematics, 11:157–160, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-37123-6/1.pdf.

Serre:1975:BMb

[Ser75c] Jean-Pierre Serre. Back matter. Lecture Notes in Mathematics, 11:157–160, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-37123-6/1.pdf.

Serre:1975:DCHa

[Ser75d] Jean-Pierre Serre. Dimension et codimension homologiques. (French) []. Lecture Notes in Mathematics, 11:67–123, 1975. CO-DEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-540-37123-6_
4/.

Serre:1975:DCHb

[Ser75e] Jean-Pierre Serre. Dimension et codimension homologiques. (French) []. Lecture Notes in Mathematics, 11:67–123, 1975. CO-DEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-540-37123-6_
4/.

Serre:1975:FMa

[Ser75f] Jean-Pierre Serre. Front matter. Lecture Notes in Mathematics, 11:I–X, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-37123-6/1.pdf.

Serre:1975:FMb

[Ser75g] Jean-Pierre Serre. Front matter. Lecture Notes in Mathematics, 11:I–X, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-37123-6/1.pdf.

Serre:1975:IPLa

[Ser75h] Jean-Pierre Serre. Idéaux premiers et localisation. (French) []. Lecture Notes in Mathematics, 11:1–7, 1975. CODEN LN-MAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-37123-6_1/.

Serre:1975:IPLb

[Ser75i] Jean-Pierre Serre. Idéaux premiers et localisation. (French) []. Lecture Notes in Mathematics, 11:1–7, 1975. CODEN LN-MAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-37123-6_1/.

Serre:1975:MFa

[Ser75j] Jean-Pierre Serre. Les multiplicités. (French) []. Lecture Notes in Mathematics, 11:124–156, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-37123-6_5/.

Serre:1975:MFb

[Ser75k] Jean-Pierre Serre. Les Multiplicités. (French) []. Lecture Notes in Mathematics, 11:124–156, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-37123-6_5/.

Serre:1975:OSFa

[Ser75l] Jean-Pierre Serre. Outils et sorites. (French) []. Lecture Notes in Mathematics, 11:8–37, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-37123-6_2/.

Serre:1975:OSFb

[Ser75m] Jean-Pierre Serre. Outils et sorites. (French) []. Lecture Notes in Mathematics, 11:8–37, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-37123-6_2/.

Serre:1975:TDFa

[Ser75n] Jean-Pierre Serre. Théorie de la dimension. (French) []. Lecture Notes in Mathematics, 11:38–66, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-37123-6_3/.

Serre:1975:TDFb

[Ser75o] Jean-Pierre Serre. Théorie de la dimension. (French) []. Lecture Notes in Mathematics, 11:38–66, 1975. CODEN LNMAA2. ISBN 3-540-07028-1 (print), 3-540-37123-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-37123-6_3/.

Serre:1975:VPE

[Ser75p] Jean-Pierre Serre. Valeurs propres des endomorphismes de Frobenius [d’après P. Deligne]. (French) []. Lecture Notes in Mathematics, 431:190–204, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066371/.

Servien:1976:TEF

[Ser76] Claude Servien. Sur la topologie d’un espace de fonctions entières avec poids. (French) []. Lecture Notes in Mathematics, 524:90– 95, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078000/.

Serre:1977:RLG

[Ser77] Jean-Pierre Serre. Représentations linéaires des groupes finis ‘algébriques’ [d’après Deligne–Lusztig]. (French) []. Lecture Notes in Mathematics, 567:256–273, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096074/.

Sergeraert:1979:DMT

[Ser79a] Francis Sergeraert. Bγ [d’après Mather et Thurston]. (French) []. Lecture Notes in Mathematics, 710:300–315, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069987/.

Serre:1979:PRC

[Ser79b] Jean-Pierre Serre. Points rationnels des courbes modulaires X_{o}(N) [d’après Barry Mazur [3], [4], [5]]. (French) []. Lecture Notes in Mathematics, 710:89–100, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069974/.

Shank:1975:TLR

[Sha75] H. Shank. The theory of left-right paths. Lecture Notes in Mathematics, 452:42–54, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069542/.

Shapiro:1976:MCA

[Sha76a] H. S. Shapiro. The modulus of continuity of an analytic function. Lecture Notes in Mathematics, 512:131–138, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080030/.

Shapiro:1976:HCO

[Sha76b] Jack M. Shapiro. On the homology and cohomology of the orthogonal and symplectic groups over a finite field of odd characteristic. Lecture Notes in Mathematics, 551:260–268, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080006/.

Shah:1977:EFB

[Sha77a] S. M. Shah. Entire functions of bounded index. Lecture Notes in Mathematics, 599:117–145, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096833/.

Shapiro:1977:CAE

[Sha77b] Harold S. Shapiro. Convergence almost everywhere of convolution integrals with a dilation parameter. Lecture Notes in Mathematics, 571:250–266, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086579/.

Shapiro:1977:RSA

[Sha77c] Joel H. Shapiro. Remarks on F -spaces of analytic functions. Lecture Notes in Mathematics, 604:107–124, 1977. CODEN LN-MAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069213/.

Shanahan:1978:AFP

[Sha78a] Patrick Shanahan. Applications of the fixed point theorem. Lecture Notes in Mathematics, 638:124–185, 1978. CODEN LN-MAA2. ISBN 3-540-08660-9 (print), 3-540-35921-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068269/.

Shanahan:1978:AIT

[Sha78b] Patrick Shanahan. Applications of the index theorem. Lecture Notes in Mathematics, 638:20–58, 1978. CODEN LNMAA2. ISBN 3-540-08660-9 (print), 3-540-35921-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068266/.

Shanahan:1978:ASF

[Sha78c] Patrick Shanahan. The Atiyah–Singer fixed point theorem. Lecture Notes in Mathematics, 638:110–123, 1978. CODEN LN-MAA2. ISBN 3-540-08660-9 (print), 3-540-35921-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068268/.

Shanahan:1978:ASI

[Sha78d] Patrick Shanahan. The Atiyah–Singer Index Theorem: An Introduction, volume 638 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08660-9 (print), 3-540-35921-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 224 pp. LCCN QA3 .L28 no. 638. URL http:

//link.springer.com/book/10.1007/BFb0068264;   http://
www.springerlink.com/content/978-3-540-35921-0.

Shanahan:1978:OP

[Sha78e] Patrick Shanahan. Outline of the proof. Lecture Notes in Mathematics, 638:59–109, 1978. CODEN LNMAA2. ISBN 3-540-08660-9 (print), 3-540-35921-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068267/.

Shanahan:1978:ST

[Sha78f] Patrick Shanahan. Statement of the theorem. Lecture Notes in Mathematics, 638:1–19, 1978. CODEN LNMAA2. ISBN 3-540-08660-9 (print), 3-540-35921-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068265/.

Shapiro:1979:CPA

[Sha79a] Harold N. Shapiro. A combinatorial problem in additive number theory. Lecture Notes in Mathematics, 751:265–282, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062714/.

Sharp:1979:NCE

[Sha79b] Rodney Y. Sharp. Necessary conditions for the existence of dualizing complexes in commutative algebra. Lecture Notes in Mathematics, 740:213–229, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071063/.

Sharpe:1979:GRA

[Sha79c] R. W. Sharpe. K_{2}(Z[Z/5]) is generated by relations among 2 × 2 matrices. Lecture Notes in Mathematics, 741:158–169, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062138/.

Shelah:1975:ERL

[She75] S. Shelah. Existence of rigid-like families of Abelian p-groups. Lecture Notes in Mathematics, 498:384–402, 1975. CODEN LN-MAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080986/.

Shirota:1975:SWP

[Shi75] Taira Shirota. On structures of L^{2}-well-posed mixed problems for hyperbolic operators. Lecture Notes in Mathematics, 459:235– 254, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074197/.

Shields:1978:SPO

[Shi78] Allen L. Shields. Some problems in operator theory. Lecture Notes in Mathematics, 693:157–167, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064673/.

Shnirman:1978:NUS

[Shn78] M. G. Shnirman. On non-uniqueness in some homogeneous networks. Lecture Notes in Mathematics, 653:31–36, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070082/.

Shub:1975:HTD

[Shu75] Mike Shub. Homology theory and dynamical systems. Lecture Notes in Mathematics, 468:36–39, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082592/.

Sibony:1976:MDA

[Sib76a] Nessim Sibony. Multi-dimensional analytic structure in the spectrum of a uniform algebra. Lecture Notes in Mathematics, 512: 139–165, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080031/.

Sibony:1976:NBA

[Sib76b] Nessim Sibony. Noyau de Bergman et applications biholomorphes dans des domaines strictement pseudo-convexes. (French) []. Lecture Notes in Mathematics, 514:145–158, 1976. CODEN LN-MAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080064/.

Sibony:1977:PDM

[Sib77] Nessim Sibony. Principe du maximum sur une variété C. R. et équations de Monge–Ampère complexes. (French) []. Lecture Notes in Mathematics, 578:14–27, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091460/.

Sidibe:1979:MLA

[Sid79] R. Sidibé. Martingales locales a accroissements independants. Lecture Notes in Mathematics, 721:132–137, 1979. CODEN LN-MAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070856/.

Sigmund:1975:PGE

[Sig75] Karl Sigmund. Propriétés générales et exceptionnelles des états statistiques de systèmes dynamiques stables. (French) []. Lecture Notes in Mathematics, 465:294–304, 1975. CODEN LN-MAA2. ISBN 3-540-07178-4 (print), 3-540-37518-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0102999/.

Sigmund:1976:NNT

[Sig76] Karl Sigmund. Nombres normaux et théorie ergodique. (French) []. Lecture Notes in Mathematics, 532:202–215, 1976. CODEN LNMAA2. ISBN 3-540-07803-7 (print), 3-540-38217-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080181/.

Sijbrand:1979:BAN

[Sij79] J. Sijbrand. Bifurcation analysis of a non linear free boundary problem from plasma physics. Lecture Notes in Mathematics, 711:76–93, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062948/.

Silverstein:1976:ADS

[Sil76a] Martin L. Silverstein. Associated Dirichlet spaces and boundary conditions. Lecture Notes in Mathematics, 516:60–114, 1976. CODEN LNMAA2. ISBN 3-540-07688-3 (print), 3-540-38222-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081338/.

Silverstein:1976:BTS

[Sil76b] Martin L. Silverstein. Boundary Theory for Symmetric Markov Processes, volume 516 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07688-3 (print), 3-540-38222-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 314 pp. LCCN QA3 .L28 no. 516. URL http:

//link.springer.com/book/10.1007/BFb0081336;   http://
www.springerlink.com/content/978-3-540-38222-5.

Silverstein:1976:ERB

[Sil76c] Martin L. Silverstein. Examples related to Brownian motion. Lecture Notes in Mathematics, 516:115–210, 1976. CODEN LN-MAA2. ISBN 3-540-07688-3 (print), 3-540-38222-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081339/.

Silverstein:1976:HHP

[Sil76d] Martin L. Silverstein. Hyperbolic half plane. Lecture Notes in Mathematics, 516:274–305, 1976. CODEN LNMAA2. ISBN 3-540-07688-3 (print), 3-540-38222-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081341/.

Silverstein:1976:SSC

[Sil76e] Martin L. Silverstein. Sample space construction. Lecture Notes in Mathematics, 516:1–59, 1976. CODEN LNMAA2. ISBN 3-540-07688-3 (print), 3-540-38222-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081337/.

Silverstein:1976:SP

[Sil76f] Martin L. Silverstein. Stable processes. Lecture Notes in Mathematics, 516:211–273, 1976. CODEN LNMAA2. ISBN 3-540-07688-3 (print), 3-540-38222-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081340/.

Simmons:1975:COT

[Sim75a] H. Simmons. The complexity of T^{f} and omitting types in F_{T}. Lecture Notes in Mathematics, 498:403–407, 1975. CODEN LN-MAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080987/.

Simms:1975:APH

[Sim75b] D. J. Simms. An application of polarisations and half-forms. Lecture Notes in Mathematics, 466:195–197, 1975. CODEN LN-MAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082207/.

Simms:1977:OGQ

[Sim77] D. J. Simms. An outline of geometric quantisation (d’après Kostant). Lecture Notes in Mathematics, 570:1–10, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087779/.

Simms:1978:SEG

[Sim78a] D. J. Simms. On the Schrödinger equation given by geometric quantisation. Lecture Notes in Mathematics, 676:351–356, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063680/.

Simon:1978:RSE

[Sim78b] Jacques Simon. Régularité de la solution d’une équation non linéaire dans R^{N}. (French) []. Lecture Notes in Mathematics, 665: 205–227, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061807/.

Sims:1978:MCG

[Sim78c] Charles C. Sims. A method for constructing a group from a subgroup. Lecture Notes in Mathematics, 697:125–136, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103127/.

Sims:1978:RAT

[Sim78d] Charles C. Sims. The role of algorithms in the teaching of algebra. Lecture Notes in Mathematics, 697:95–107, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103125/.

Sims:1978:SGT

[Sim78e] Charles C. Sims. Some group-theoretic algorithms. Lecture Notes in Mathematics, 697:108–124, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103126/.

Singer:1977:SRO

[Sin77] I. M. Singer. Some remarks on operator theory and index theory. Lecture Notes in Mathematics, 575:128–138, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095707/.

Sjerve:1978:GHR

[Sje78] Denis Sjerve. Generalized homological reduction theorems. Lecture Notes in Mathematics, 673:262–275, 1978. CODEN LN-MAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0064702/.

Sjostrand:1975:AFD

[Sjö75] Johannes Sjöstrand. Applications of Fourier distributions with complex phase functions. Lecture Notes in Mathematics, 459:255– 282, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074198/.

Sjogren:1976:AEB

[Sjö76a] Peter Sjögren. Approximation dans les espaces de Banach et mesures spectrales d’après Butzer, Nessel, Trebels. (French) []. Lecture Notes in Mathematics, 518:266–275, 1976. CODEN LN-MAA2. ISBN 3-540-07691-3 (print), 3-540-38225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080417/.

Sjogren:1976:PFH

[Sjö76b] Peter Sjögren. Une propriété des fonctions harmoniques positives d’après Dahlberg. (French) []. Lecture Notes in Mathematics, 563:275–282, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087583/.

Sjostrand:1978:CCO

[Sjö78] J. Sjöstrand. Sur certains complexes d’opérateurs pseudodifférentiels. (French) []. Lecture Notes in Mathematics, 660:181– 189, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082304/.

Skoda:1976:BVS

[Sko76] H. Skoda. Boundary values for the solutions of the ¯∂-equation and application to the Nevanlinna class. Lecture Notes in Mathematics, 512:166–179, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080032/.

Skoda:1978:MSF

[Sko78] Henri Skoda. Morphismes surjectifs et fibrés linéaires semi-positifs. (French) []. Lecture Notes in Mathematics, 694:290–324, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063255/.

Sleeman:1975:LDM

[Sle75] B. D. Sleeman. Left-definite multiparameter eigenvalue problems. Lecture Notes in Mathematics, 448:307–321, 1975. CODEN LN-MAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067093/.

Slomson:1976:DPG

[Slo76] Alan Slomson. Decision problems for generalized quantifiers — a survey. Lecture Notes in Mathematics, 537:249–258, 1976. CO-DEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096905/.

Slowikowski:1978:CWA

[Sło78] W. Słowikowski. Commutative Wick algebras II. Square integrable martingale algebras and Itô algebras. Lecture Notes in Mathematics, 695:179–191, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062665/.

Street:1979:BBA

[SM79a] Anne Penfold Street and Sheila Oates Macdonald. Balanced binary arrays I: The square grid. Lecture Notes in Mathematics, 748:165–198, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102695/.

Stuetzle:1979:SCA

[SM79b] Werner Stuetzle and Yashaswini Mittal. Some comments on the asymptotic behavior of robust smoothers. Lecture Notes in Mathematics, 757:191–195, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098497/.

Smale:1977:LID

[Sma77] Steve Smale. Lecture III: Dynamical systems and turbulence. Lecture Notes in Mathematics, 615:48–70, 1977. CODEN LN-MAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068360/.

Smith:1976:CIC

[Smi76a] Jonathan D. H. Smith. Central isotopy and cancellation. Lecture Notes in Mathematics, 554:69–92, 1976. CODEN LNMAA2. ISBN 3-540-07999-8 (print), 3-540-37939-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095451/.

Smith:1976:C

[Smi76b] Jonathan D. H. Smith. Centrality. Lecture Notes in Mathematics, 554:23–48, 1976. CODEN LNMAA2. ISBN 3-540-07999-8 (print), 3-540-37939-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095449/.

Smith:1976:DD

[Smi76c] Jonathan D. H. Smith. Direct decompositions. Lecture Notes in Mathematics, 554:49–68, 1976. CODEN LNMAA2. ISBN 3-540-07999-8 (print), 3-540-37939-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095450/.

Smith:1976:EO

[Smi76d] Jonathan D. H. Smith. Extensions and obstructions. Lecture Notes in Mathematics, 554:113–151, 1976. CODEN LNMAA2. ISBN 3-540-07999-8 (print), 3-540-37939-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095453/.

Smith:1976:MV

[Smi76e] Jonathan D. H. Smith. Mal’cev Varieties, volume 554 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07999-8 (print), 3-540-37939-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 158 pp. LCCN QA3

.L28 no. 554. URL http://link.springer.com/book/10.1007/
BFb0095447;   http://www.springerlink.com/content/978-
3-540-37939-3.

Smith:1976:PAE

[Smi76f] Jonathan D. H. Smith. Plain algebras and equational completeness. Lecture Notes in Mathematics, 554:93–112, 1976. CODEN LNMAA2. ISBN 3-540-07999-8 (print), 3-540-37939-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095452/.

Smith:1976:RN

[Smi76g] Jonathan D. H. Smith. Rudiments and notations. Lecture Notes in Mathematics, 554:1–22, 1976. CODEN LNMAA2. ISBN 3-540-07999-8 (print), 3-540-37939-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095448/.

Smirnov:1977:STA

[Smi77] Yu. M. Smirnov. Some topological aspects of the theory of topological transformation groups. Lecture Notes in Mathematics, 609: 196–204, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068684/.

Smith:1979:FGA

[Smi79] Martha Smith. Finitely generated algebras over a field. Lecture Notes in Mathematics, 700:229, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063473.

Smyrnelis:1978:PEE

[Smy78a] Emmanuel P. Smyrnélis. Polarité et effilement dans les espaces biharmoniques. (French) [Polarity and taper in the biharmonic spaces]. Lecture Notes in Mathematics, 681:252–276, 1978. CO-DEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0065880/.

Smyrnelis:1978:FHO

[Smy78b] Emmanuel P. Smyrnélis. Sur les fonctions hyperharmoniques d’ordre 2. (French) []. Lecture Notes in Mathematics, 681:277– 294, 1978. CODEN LNMAA2. ISBN 3-540-08947-0 (print), 3-540-35712-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065881/.

Snaith:1975:DLO

[Sna75] V. P. Snaith. Dyer–Lashof operations in K-theory. Lecture Notes in Mathematics, 496:103–294, 1975. CODEN LNMAA2. ISBN 3-540-07536-4 (print), 3-540-38026-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082286/.

Snaith:1978:SDC

[Sna78] Victor Snaith. Stable decompositions of classifying spaces with applications to algebraic Cobordism theories. Lecture Notes in Mathematics, 673:123–157, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064693/.

Sneddon:1975:UMP

[Sne75] Ian Naismith Sneddon. The use in mathematical physics of Erdélyi–Kober operators and of some of their generalizations. Lecture Notes in Mathematics, 457:37–79, 1975. CODEN LN-MAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067097/.

Sniatycki:1977:CGA

[Śni77] Jedrzej Śniatycki. On cohomology groups appearing in geometric quantization. Lecture Notes in Mathematics, 570:46–66, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087781/.

Sniatycki:1978:AGQ

[Śni78] Jedrzej Śniatycki. Application of geometric quantization in quantum mechanics. Lecture Notes in Mathematics, 676:357–367, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063681/.

Snider:1979:BGG

[Sni79] Robert L. Snider. Is the Brauer group generated by cyclic algebras? Lecture Notes in Mathematics, 734:279–301, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103164/.

Sario:1977:BGFb

[SNWC77a] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Biharmonic Green’s function β: Definition and existence. Lecture Notes in Mathematics, 605:340–382, 1977. CODEN LN-MAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064428/.

Sario:1977:BGFa

[SNWC77b] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Biharmonic Green’s function γ. Lecture Notes in Mathematics, 605: 298–339, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064427/.

Sario:1977:BBF

[SNWC77c] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Bounded biharmonic functions. Lecture Notes in Mathematics, 605:137–186, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064422/.

Sario:1977:BDF

[SNWC77d] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Bounded Dirichlet finite biharmonic functions. Lecture Notes in Mathematics, 605:216–236, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064424/.

Sario:1977:CTR

[SNWC77e] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Classification Theory of Riemannian Manifolds: Harmonic, quasiharmonic and biharmonic functions, volume 605 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 498 pp. LCCN QA3

.L28 no. 605. URL http://link.springer.com/book/10.1007/
BFb0064417;   http://www.springerlink.com/content/978-
3-540-37261-5.

Sario:1977:DFB

[SNWC77f] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Dirichlet finite biharmonic functions. Lecture Notes in Mathematics, 605:187–215, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064423/.

Sario:1977:HCG

[SNWC77g] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Hadamard’s conjecture on the Green’s function of a clamped plate. Lecture Notes in Mathematics, 605:444–472, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064430/.

Sario:1977:HF

[SNWC77h] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Harmonic functions. Lecture Notes in Mathematics, 605:26– 70, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064420/.

Sario:1977:HQB

[SNWC77i] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Harmonic, quasiharmonic, and biharmonic degeneracies. Lecture Notes in Mathematics, 605:237–244, 1977. CODEN LN-MAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064425/.

Sario:1977:LBO

[SNWC77j] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Laplace–Beltrami operator. Lecture Notes in Mathematics, 605: 12–25, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064419/.

Sario:1977:PHN

[SNWC77k] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Preface and historical note. Lecture Notes in Mathematics, 605: 1–11, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064418/.

Sario:1977:QF

[SNWC77l] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Quasiharmonic functions. Lecture Notes in Mathematics, 605: 71–136, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064421/.

Sario:1977:RON

[SNWC77m] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Relation of O_{β}^{N} to other null classes. Lecture Notes in Mathematics, 605:383–443, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064429/.

Sario:1977:RRB

[SNWC77n] Leo Sario, Mitsuru Nakai, Cecilia Wang, and Lung Ock Chung. Riesz representation of biharmonic functions. Lecture Notes in Mathematics, 605:245–297, 1977. CODEN LNMAA2. ISBN 3-540-08358-8 (print), 3-540-37261-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064426/.

Sochor:1976:AST

[Soc76] Antonin Sochor. The alternative set theory. Lecture Notes in Mathematics, 537:259–271, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096906/.

Sochor:1977:DCA

[Soc77] Antonín Sochor. Differential calculus in the alternative set theory. Lecture Notes in Mathematics, 619:273–284, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067658/.

Solovay:1978:CS

[Sol78a] Robert M. Solovay. A Δ_{3}^{1} coding of the subsets of ω_{ω}. Lecture Notes in Mathematics, 689:133–150, 1978. CODEN LN-MAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069297/.

Solovay:1978:IDA

[Sol78b] Robert M. Solovay. The independence of DC from AD. Lecture Notes in Mathematics, 689:171–183, 1978. CODEN LN-MAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069299/.

Sommer:1976:AGA

[Som76] Dr. Manfred Sommer. Alternanten Bei Gelchmässiger Approximation Mit Zweidimensionalen Splinefunktionen. (German) []. Lecture Notes in Mathematics, 501:339–370, 1976. CODEN LN-MAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079755/.

Souriau:1977:IGE

[Sou77] Jean-Marie Souriau. Interpretation géométrique des états quantiques. (French) []. Lecture Notes in Mathematics, 570:76–96, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087784/.

Souriau:1978:TGF

[Sou78] Jean-Marie Souriau. Thermodynamique et géométrie. (French) []. Lecture Notes in Mathematics, 676:369–397, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063682/.

Spector:1975:ATH

[Spe75a] René Spector. Apercu de la théorie des hypergroupes. (French) []. Lecture Notes in Mathematics, 497:643–673, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078026/.

Spellucci:1975:ARD

[Spe75b] Dr. Peter Spellucci. Algorithms for rational discrete least squares approximation. Lecture Notes in Mathematics, 477:261–275, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079181/.

Springer:1975:RBR

[Spr75] T. A. Springer. Relèvement de Brauer et représentations paraboliques de GL_{n}(F_{q}) [d’après G. Lusztig]. (French) []. Lecture Notes in Mathematics, 431:89–113, 1975. CODEN LN-MAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0066366/.

Springer:1977:SL

[Spr77a] T. A. Springer. SL_{2}(k). Lecture Notes in Mathematics, 585:43– 72, 1977. CODEN LNMAA2. ISBN 3-540-08242-5 (print), 3-540-37370-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095647/.

Springer:1977:AI

[Spr77b] T. A. Springer. The algebra of invariants. Lecture Notes in Mathematics, 585:15–42, 1977. CODEN LNMAA2. ISBN 3-540-08242-5 (print), 3-540-37370-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095646/.

Springer:1977:FG

[Spr77c] T. A. Springer. Finite groups. Lecture Notes in Mathematics, 585:73–111, 1977. CODEN LNMAA2. ISBN 3-540-08242-5 (print), 3-540-37370-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095648/.

Springer:1977:I

[Spr77d] T. A. Springer. Introduction. Lecture Notes in Mathematics, 585: 1–14, 1977. CODEN LNMAA2. ISBN 3-540-08242-5 (print), 3-540-37370-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095645/.

Springer:1977:IT

[Spr77e] T. A. Springer. Invariant Theory, volume 585 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08242-5 (print), 3-540-37370-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 112 pp. LCCN QA3

.L28 no. 585. URL http://link.springer.com/book/10.1007/
BFb0095644;   http://www.springerlink.com/content/978-
3-540-37370-4.

Springer:1977:RGW

[Spr77f] T. A. Springer. Représentations de groupes de Weyl et éléments nilpotents d’algèbres de Lie. (French) []. Lecture Notes in Mathematics, 586:86–92, 1977. CODEN LNMAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087124/.

Srebro:1976:QM

[Sre76] Uri Srebro. Quasiregular mappings. Lecture Notes in Mathematics, 505:148–163, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081104/.

Srebrny:1978:SCA

[Sre78] Marian Srebrny. Singular cardinals and analytic games. Lecture Notes in Mathematics, 669:391–422, 1978. CODEN LN-MAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103111/.

Srivastava:1978:LIS

[Sri78] J. N. Srivastava. On the linear independence of sets of 2^{q} columns of certain (1, −1) matrices with a group structure, and its connection with finite geometries. Lecture Notes in Mathematics, 686: 79–88, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062519/.

Srinivasan:1979:AC

[Sri79a] Bhama Srinivasan. The ℓ-adic cohomology. Lecture Notes in Mathematics, 764:31–59, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097144/.

Srinivasan:1979:C

[Sri79b] Bhama Srinivasan. Characters. Lecture Notes in Mathematics, 764:96–127, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097146/.

Srinivasan:1979:CR

[Sri79c] Bhama Srinivasan. Classification of representations. Lecture Notes in Mathematics, 764:128–167, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097147/.

Srinivasan:1979:CT

[Sri79d] Bhama Srinivasan. Classification of tori. Lecture Notes in Mathematics, 764:8–20, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097141/.

Srinivasan:1979:CLD

[Sri79e] Bhama Srinivasan. The construction of Lusztig–Deligne. Lecture Notes in Mathematics, 764:60–95, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097145/.

Srinivasan:1979:DSR

[Sri79f] Bhama Srinivasan. Discrete series representations and Harish-Chandra theory. Lecture Notes in Mathematics, 764:25–30, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097143/.

Srinivasan:1979:PSR

[Sri79g] Bhama Srinivasan. Principal series representations. Lecture Notes in Mathematics, 764:21–24, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097142/.

Srinivasan:1979:RFC

[Sri79h] Bhama Srinivasan. Representations of Finite Chevalley Groups: A Survey, volume 764 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 177 pp. LCCN QA3 .L28 no. 764. URL http:

//link.springer.com/book/10.1007/BFb0097139;   http://
www.springerlink.com/content/978-3-540-38488-5.

Srinivasan:1979:RRA

[Sri79i] Bhama Srinivasan. Review of results on algebraic groups. Lecture Notes in Mathematics, 764:1–7, 1979. CODEN LNMAA2. ISBN 3-540-09716-3 (print), 3-540-38488-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097140/.

Schaback:1976:ATP

[SS76a] Robert Schaback and Karl Scherer, editors. Approximation Theory: Proceedings of an International Colloquium Held at Bonn, Germany, June 8–11, 1976, volume 556 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 556. URL http:

//link.springer.com/book/10.1007/BFb0087392;   http://
www.springerlink.com/content/978-3-540-37552-4.

Schafer:1976:SIR

[SS76b] Werner Schäfer and Walter Schempp. Splineapproximation in intermediären Räumen. (German) []. Lecture Notes in Mathematics, 501:226–246, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079751/.

Schiffer:1976:DTQ

[SS76c] M. Schiffer and G. Schober. A distortion theorem for quasiconformal mappings. Lecture Notes in Mathematics, 505:138–147, 1976. CODEN LNMAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081103/.

Schurer:1976:ILT

[SS76d] F. Schurer and F. W. Steutel. On an inequality of Lorentz in the theory of Bernstein polynomials. Lecture Notes in Mathematics, 501:332–338, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079754/.

Simons:1976:WCS

[SS76e] Gordon Simons and William Stout. Weak convergence to stable laws by means of a weak invariance principle. Lecture Notes in Mathematics, 566:127–137, 1976. CODEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096883/.

Schafer:1977:EBB

[SS77a] Dr. Werner Schäfer and Prof. Dr. Walter Schempp. Error bounds for bivariate spline interpolation. Lecture Notes in Mathematics, 571:196–211, 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086575/.

Serre:1977:MFW

[SS77b] J.-P. Serre and H. M. Stark. Modular forms of weight 1/2. Lecture Notes in Mathematics, 627:27–67, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065296/.

Smith:1977:ESB

[SS77c] Mark A. Smith and Francis Sullivan. Extremely smooth Banach spaces. Lecture Notes in Mathematics, 604:125–137, 1977. CO-DEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0069214/.

Shulman:1978:RT

[SS78a] Herbert Shulman and James Stasheff. De Rham theory for bΓ. Lecture Notes in Mathematics, 652:62–74, 1978. CODEN LN-MAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063502/.

Sigrist:1978:EOG

[SS78b] François Sigrist and Ueli Suter. On the exponent and the order of the groups J̃(X). Lecture Notes in Mathematics, 673:116– 122, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064692/.

Seese:1977:CNT

[ST77a] D. G. Seese and H. P. Tuschik. Construction of nice trees. Lecture Notes in Mathematics, 619:257–271, 1977. CODEN LN-MAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067657/.

Shyr:1977:CBR

[ST77b] H. J. Shyr and G. Thierrin. Codes and binary relations. Lecture Notes in Mathematics, 586:180–188, 1977. CODEN LN-MAA2. ISBN 3-540-08243-3 (print), 3-540-37369-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087133/.

Staples:1975:CRT

[Sta75] John Staples. Church–Rosser theorems for replacement systems. Lecture Notes in Mathematics, 450:291–307, 1975. CODEN LN-MAA2. ISBN 3-540-07152-0 (print), 3-540-37480-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062861/.

Stallings:1977:CMF

[Sta77a] John Stallings. Coherence of 3-manifold fundamental groups. Lecture Notes in Mathematics, 567:167–173, 1977. CODEN LN-MAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096068/.

Stanley:1977:SCA

[Sta77b] Richard P. Stanley. Some combinatorial aspects of the Schubert calculus. Lecture Notes in Mathematics, 579:217–251, 1977. CO-DEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0090019/.

Stanton:1977:DBU

[Sta77c] Robert O. Stanton. Decomposition bases and Ulm’s theorem. Lecture Notes in Mathematics, 616:39–56, 1977. CODEN LN-MAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068188/.

Stark:1977:CFM

[Sta77d] H. M. Stark. Class fields and modular forms of weight one. Lecture Notes in Mathematics, 601:277–287, 1977. CODEN LN-MAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063951/.

Stampfli:1978:QD

[Sta78a] Joseph G. Stampfli. On a question of Deddens. Lecture Notes in Mathematics, 693:169–173, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064674/.

Stanton:1978:DKP

[Sta78b] Professor R. G. Stanton. The Doehlert–Klee problem: Part I, statistical background. Lecture Notes in Mathematics, 686:89– 100, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062520/.

Stasheff:1978:RHO

[Sta78c] James Stasheff. Rational homotopy-obstruction and perturbation theory. Lecture Notes in Mathematics, 673:7–31, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064687/.

Stavskaya:1978:MFI

[Sta78d] O. N. Stavskaya. Markov fields as invariant states for local processes. Lecture Notes in Mathematics, 653:113–121, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070088/.

Stafford:1979:CNM

[Sta79a] J. T. Stafford. Cancellation for nonprojective modules. Lecture Notes in Mathematics, 700:3–15, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063456/.

Stafford:1979:TNG

[Sta79b] J. T. Stafford. K-theory of Noetherian group rings. Lecture Notes in Mathematics, 734:302–322, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103165/.

Stafford:1979:SNR

[Sta79c] Toby Stafford. Simple Noetherian rings. Lecture Notes in Mathematics, 700:230–231, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063474.

Stefan:1975:AF

[Ste75a] Peter Stefan. Accessibility and foliations. Lecture Notes in Mathematics, 468:10–12, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082579/.

Stehle:1975:FPC

[Ste75b] Jean-Luc Stehlé. Fonctions plurisousharmoniques et convexité holomorphe de certains fibres analytiques. (French) []. Lecture Notes in Mathematics, 474:155–179, 1975. CODEN LN-MAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077406/.

Stehle:1975:RCE

[Ste75c] Jean-Luc Stehlé. Rectificatif concernant l’exposé: “Un théorème d’image directe propre”, publié dans le Séminaire Pierre Lelong, 1972/73, n^{o} 410. (French) []. Lecture Notes in Mathematics, 474: 180–182, 1975. CODEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077407/.

Stehle:1975:SEW

[Ste75d] Jean-Luc Stehlé. Sous-espaces de Weierstrass. (French) []. Lecture Notes in Mathematics, 482:337–350, 1975. CODEN LN-MAA2. ISBN 3-540-07397-3 (print), 3-540-37601-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080254/.

Stehle:1976:FQC

[Ste76a] Jean-Luc Stehlé. Faisceaux F -quasi-cohérents en géométrie analytique. (French) [Quasi-coherent F sheaves in analytic geometry]. Lecture Notes in Mathematics, 524:30–54, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077996/.

Stehle:1976:CPF

[Ste76b] Jean-Luc Stehlé. Un cas particulier de faisceau F -quasi-cohérent avec nullité des tôr. (French) []. Lecture Notes in Mathematics, 524:55–66, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077997/.

Stein:1976:AEP

[Ste76c] Michael R. Stein, editor. Algebraic K-Theory: Proceedings of the Conference Held at Northwestern University Evanston, January 12–16, 1976, volume 551 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 551. URL http:

//link.springer.com/book/10.1007/BFb0079990;   http://
www.springerlink.com/content/978-3-540-37964-5.

Stetter:1976:TTD

[Ste76d] Hans J. Stetter. Towards a theory for discretizations of stiff differential systems. Lecture Notes in Mathematics, 506:190–201, 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080125/.

Stegall:1977:PMC

[Ste77a] Charles Stegall. A proof of the martingale convergence theorem in Banach spaces. Lecture Notes in Mathematics, 604:138–141, 1977. CODEN LNMAA2. ISBN 3-540-08356-1 (print), 3-540-37262-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069215/.

Sternberg:1977:SRR

[Ste77b] Shlomo Sternberg. Some recent results on supersymmetry. Lecture Notes in Mathematics, 570:145–176, 1977. CODEN LN-MAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087787/.

Steel:1978:VC

[Ste78a] John R. Steel. On Vaught’s conjecture. Lecture Notes in Mathematics, 689:193–208, 1978. CODEN LNMAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069301/.

Steffens:1978:RFS

[Ste78b] Karsten Steffens. Regressive functions and stationary sets. Lecture Notes in Mathematics, 669:423–435, 1978. CODEN LN-MAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103112/.

Stein:1978:TPH

[Ste78c] Karl Stein. Topological properties of holomorphic and meromorphic mappings. Lecture Notes in Mathematics, 683:203–216, 1978. CODEN LNMAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063179/.

Stepanek:1978:CIM

[Ště78d] Petr Štěpánek. Cardinals in the inner model HOD. Lecture Notes in Mathematics, 669:437–454, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103113/.

Stern:1978:PCS

[Ste78e] Jacques Stern. Le problème des cardinaux singuliers [d’après R. B. Jensen et J. Silver]. (French) []. Lecture Notes in Mathematics, 677:59–72, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070754/.

Stern:1978:PRL

[Ste78f] Jacques Stern. Partitions of the real line into X_{1} closed sets. Lecture Notes in Mathematics, 669:455–460, 1978. CODEN LN-MAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103114/.

Sternberg:1978:RFT

[Ste78g] Shlomo Sternberg. On the role of field theories in our physical conception of geometry. Lecture Notes in Mathematics, 676:1– 80, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063665/.

Sternberg:1978:SPR

[Ste78h] Shlomo Sternberg. Some preliminary remarks on the formal variational calculus of Gel’fand and Dikii. Lecture Notes in Mathematics, 676:399–407, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063683/.

Stetter:1978:GEE

[Ste78i] H. J. Stetter. Global error estimation in ODE-solvers. Lecture Notes in Mathematics, 630:179–189, 1978. CODEN LN-MAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067705/.

Stetter:1978:CCT

[Ste78j] Hans J. Stetter. Considerations concerning a theory for ODEsolvers. Lecture Notes in Mathematics, 631:188–200, 1978. CO-DEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0067472/.

Steinberger:1979:ETD

[Ste79] Mark Steinberger. On the equivalence of the two definitions of the algebraic k-theory of a topological space. Lecture Notes in Mathematics, 763:317–331, 1979. CODEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088092/.

Stolarsky:1975:DSD

[Sto75a] Kenneth B. Stolarsky. Discrepancy and sums of distances between points of a metric space. Lecture Notes in Mathematics, 490:44–56, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081130/.

Stora:1975:SAJ

[Sto75b] R. Stora. Some applications of the Jost–Lehmann–Dyson theorem to the study of the global analytic structure of the N point function of quantum field theory. Lecture Notes in Mathematics, 449:163–184, 1975. CODEN LNMAA2. ISBN 3-540-07151-2 (print), 3-540-37454-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062922/.

Stone:1976:TMT

[Sto76] A. H. Stone. Topology and measure theory. Lecture Notes in Mathematics, 541:43–48, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081039/.

Stoll:1977:DR

[Sto77a] Wilhelm Stoll. The defect relation. Lecture Notes in Mathematics, 600:138–184, 1977. CODEN LNMAA2. ISBN 3-540-08341-3 (print), 3-540-37306-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062910/.

Stoll:1977:FMTa

[Sto77b] Wilhelm Stoll. The first main theorem. Lecture Notes in Mathematics, 600:55–72, 1977. CODEN LNMAA2. ISBN 3-540-08341-3 (print), 3-540-37306-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062907/.

Stoll:1977:FMTb

[Sto77c] Wilhelm Stoll. The first main theorem on parabolic spaces. Lecture Notes in Mathematics, 600:73–108, 1977. CODEN LN-MAA2. ISBN 3-540-08341-3 (print), 3-540-37306-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062908/.

Stoll:1977:I

[Sto77d] Wilhelm Stoll. Introduction. Lecture Notes in Mathematics, 600: 1–12, 1977. CODEN LNMAA2. ISBN 3-540-08341-3 (print), 3-540-37306-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062905/.

Stoll:1977:P

[Sto77e] Wilhelm Stoll. Preparations. Lecture Notes in Mathematics, 600: 13–54, 1977. CODEN LNMAA2. ISBN 3-540-08341-3 (print), 3-540-37306-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062906/.

Stoll:1977:SMT

[Sto77f] Wilhelm Stoll. The second main theorem. Lecture Notes in Mathematics, 600:109–137, 1977. CODEN LNMAA2. ISBN 3-540-08341-3 (print), 3-540-37306-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062909/.

Stoll:1977:VDP

[Sto77g] Wilhelm Stoll. Value Distribution on Parabolic Spaces, volume 600 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08341-3 (print), 3-540-37306-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 184 (est.) pp. LCCN QA3 .L28 no. 600. URL http:

//link.springer.com/book/10.1007/BFb0062904;   http://
www.springerlink.com/content/978-3-540-37306-3.

Stone:1977:MPM

[Sto77h] A. H. Stone. Measure-preserving maps. Lecture Notes in Mathematics, 609:205–210, 1977. CODEN LNMAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068685/.

Stoltzfus:1978:ACI

[Sto78a] Neal W. Stoltzfus. Algebraic computations of the integral concordance and double null concordance group of knots. Lecture Notes in Mathematics, 685:274–290, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062977/.

Stormer:1978:HSS

[Stø78b] Erling Størmer. Homogeneity of the state space of factors of type III_{1}. Lecture Notes in Mathematics, 650:168–172, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067400/.

Stolarsky:1979:NBP

[Sto79] Kenneth B. Stolarsky. The number of bits in a product of odd integers. Lecture Notes in Mathematics, 751:283–293, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062715/.

Struble:1975:FCO

[Str75] Raimond A. Struble. Fractional calculus in the operator field of generalized functions. Lecture Notes in Mathematics, 457:294– 297, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067114/.

Strauss:1976:ASQ

[Str76] Dr. Hans Strauß. Approximation mit splinefunktionen und quadraturformeln. (German) []. Lecture Notes in Mathematics, 501:371–419, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079756/.

Street:1977:QMD

[Str77] Anne Penfold Street. On quasi-multiple designs. Lecture Notes in Mathematics, 622:206–208, 1977. CODEN LNMAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069194/.

Stricker:1978:RCT

[Str78] C. Stricker. Une remarque sur les changements de temps et les martingales locales. (French) []. Lecture Notes in Mathematics, 649:20–21, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064588.

Strang:1979:FMP

[Str79a] Gilbert Strang. A family of model problems in plasticity. Lecture Notes in Mathematics, 704:292–305, 1979. CODEN LN-MAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063627/.

Strang:1979:MPP

[Str79b] Gilbert Strang. A minimax problem in plasticity theory. Lecture Notes in Mathematics, 701:319–333, 1979. CODEN LN-MAA2. ISBN 3-540-09110-6 (print), 3-540-35530-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062087/.

Stricker:1979:VSF

[Str79c] C. Stricker. Sur la p-variation des surmartingales. (French) []. Lecture Notes in Mathematics, 721:233–237, 1979. CODEN LN-MAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070864/.

Stricker:1979:REP

[Str79d] C. Stricker. Une remarque sur l’expose precedent. (French) []. Lecture Notes in Mathematics, 721:238–239, 1979. CODEN LN-MAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070865.

Stummel:1976:PDE

[Stu76a] Prof. Dr. Friedrich Stummel. Perturbation of domains in elliptic boundary value problems. Lecture Notes in Mathematics, 503:

110–136, 1976. CODEN LNMAA2. ISBN 3-540-07629-8 (print), 3-540-38165-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088748/.

Stute:1976:NCC

[Stu76b] W. Stute. A necessary condition for the convergence of the isotrope discrepancy. Lecture Notes in Mathematics, 566:138–140, 1976. CODEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096884/.

Schroder:1978:FPS

[STW78] Johann Schröder, Ulrich Trottenberg, and Kristian Witsch. On fast Poisson solvers and applications. Lecture Notes in Mathematics, 631:153–187, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067471/.

Subia:1975:FSF

[Sub75] Nelson Subia. Formule de Selberg et formes d’espaces hyperboliques compactes. (French) []. Lecture Notes in Mathematics, 497:674–700, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078027/.

Suffridge:1976:SFL

[Suf76] T. J. Suffridge. Starlike functions as limits of polynomials. Lecture Notes in Mathematics, 505:164–203, 1976. CODEN LN-MAA2. ISBN 3-540-07548-8 (print), 3-540-38088-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081105/.

Suffridge:1977:SCO

[Suf77] T. J. Suffridge. Starlikeness, convexity and other geometric properties of holomorphic maps in higher dimensions. Lecture Notes in Mathematics, 599:146–159, 1977. CODEN LNMAA2. ISBN 3-540-08343-X (print), 3-540-37303-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096834/.

Sundaresan:1975:SCR

[Sun75] Kondagunta Sundaresan. Smoothness classification of reflexive spaces I. Lecture Notes in Mathematics, 490:107–115, 1975. CO-DEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081136/.

Sunley:1979:PDR

[Sun79] Judith S. Sunley. Prime discriminants in real quadratic fields of narrow class number one. Lecture Notes in Mathematics, 751: 294–301, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062716/.

Suszko:1975:AFA

[Sus75] Roman Suszko. Abolition of the Fregean axiom. Lecture Notes in Mathematics, 453:169–239, 1975. CODEN LNMAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064874/.

Suslin:1979:CPP

[Sus79] A. A. Suslin. The cancellation problem for projective modules and related topics. Lecture Notes in Mathematics, 734:323–338, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103166/.

Suzuki:1978:IPC

[Suz78a] Masakazu Suzuki. Sur les intégrales premières de certains feuilletages analytiques complexes. (French) []. Lecture Notes in Mathematics, 670:53–79, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064394/.

Suzuki:1978:OHE

[Suz78b] Masakazu Suzuki. Sur les opérations holomorphes de C et de C∗ sur un espace de stein. (French) []. Lecture Notes in Mathematics, 670:80–88, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064395/.

Suzuki:1979:PRH

[Suz79] Osamu Suzuki. The problems of Riemann and Hilbert and the relations of Fuchs in several complex variables. Lecture Notes in Mathematics, 712:325–364, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062822/.

Stratila:1975:AAD

[SV75a] Dr. Şerban-Valentin Strâtilâ and Dr. Dan-Virgil Voiculescu. The C^{∗}-algebra associated to a direct limit of compact groups. Lecture Notes in Mathematics, 486:57–80, 1975. CODEN LNMAA2. ISBN 3-540-07403-1 (print), 3-540-37916-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082278/.

Stratila:1975:SAA

[SV75b] Dr. Şerban-Valentin Strâtilâ and Dr. Dan-Virgil Voiculescu. On the structure of af-algebras and their representations. Lecture Notes in Mathematics, 486:1–56, 1975. CODEN LNMAA2. ISBN 3-540-07403-1 (print), 3-540-37916-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082277/.

Stratila:1975:PI

[SV75c] Dr. Şerban-Valentin Strâtilâ and Dr. Dan-Virgil Voiculescu. The primitive ideals of (A(U(∞))). Lecture Notes in Mathematics, 486: 81–96, 1975. CODEN LNMAA2. ISBN 3-540-07403-1 (print), 3-540-37916-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082279/.

Stratila:1975:RAA

[SV75d] Dr. Şerban-Valentin Strâtilâ and Dr. Dan-Virgil Voiculescu. Representations of AF-Algebras and of the Group U(∞), volume 486 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07403-1 (print), 3-540-37916-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 154 (est.) pp. LCCN QA3 .L28 no. 486. URL http:

//link.springer.com/book/10.1007/BFb0082276;   http://
www.springerlink.com/content/978-3-540-37916-4.

Stratila:1975:STI

[SV75e] Dr. Şerban-Valentin Strâtilâ and Dr. Dan-Virgil Voiculescu. Some type II_{∞} factor representations of U(∞). Lecture Notes in Mathematics, 486:127–154, 1975. CODEN LNMAA2. ISBN 3-540-07403-1 (print), 3-540-37916-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082281/.

Stratila:1975:TIF

[SV75f] Dr. Şerban-Valentin Strâtilâ and Dr. Dan-Virgil Voiculescu. Type III factor representations of U(∞) in antisymmetric tensors. Lecture Notes in Mathematics, 486:97–126, 1975. CODEN LN-MAA2. ISBN 3-540-07403-1 (print), 3-540-37916-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0082280/.

Sanders:1979:AHO

[SV79] Jan A. Sanders and Ferdinand Verhulst. Approximations of higher order resonances with an application to Contopoulos’ model problem. Lecture Notes in Mathematics, 711:209–228, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062955/.

Sved:1977:GBC

[Sve77] Marta Sved. A generalisation of the binomial coefficients. Lecture Notes in Mathematics, 622:209–213, 1977. CODEN LN-MAA2. ISBN 3-540-08524-6 (print), 3-540-37020-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069195/.

Saracino:1975:ACC

[SW75a] D. Saracino and V. Weispfenning. On algebraic curves over commutative regular rings. Lecture Notes in Mathematics, 498:307– 383, 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080985/.

Saracino:1975:MTA

[SW75b] Dr. Daniel H. Saracino and Dr. Volker B. Weispfenning. Model Theory and Algebra: A Memorial Tribute to Abraham Robinson, volume 498 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 463 pp. LCCN QA3 .L28 no. 498. URL http:

//link.springer.com/book/10.1007/BFb0080972;   http://
www.springerlink.com/content/978-3-540-38057-3.

Schwichtenberg:1975:ITR

[SW75c] H. Schwichtenberg and S. S. Wainer. Infinite terms and recursion in higher types. Lecture Notes in Mathematics, 500:341–364, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079561/.

Street:1975:CMI

[SW75d] Dr. Anne Penfold Street and Dr. Walter Denis Wallis, editors. Combinatorial Mathematics III: Proceedings of the Third Australian Conference Held at the University of Queensland, 16–18 May, 1974, volume 452 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 452. URL http:

//link.springer.com/book/10.1007/BFb0069539;   http://
www.springerlink.com/content/978-3-540-37482-4.

Schweitzer:1978:PPR

[SW78a] Paul A. Schweitzer, S.J. and Andrew P. Whitman, S.J. Pontryagin polynomial residues of isolated foliation singularities. Lecture Notes in Mathematics, 652:95–103, 1978. CODEN LN-MAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063504/.

Schweitzer:1978:SFF

[SW78b] Paul A. Schweitzer, S.J. and Andrew P. Whitman, S.J. Structures feuilletées. (French) [Laminated structures]. Lecture Notes in Mathematics, 652:104–113, 1978. CODEN LNMAA2. ISBN 3-540-07868-1 (print), 3-540-38074-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063505/.

Smythe:1978:BP

[SW78c] R. T. Smythe and John C. Wierman. Bernoulli percolation. Lecture Notes in Mathematics, 671:28–62, 1978. CODEN LN-MAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063309/.

Smythe:1978:COP

[SW78d] R. T. Smythe and John C. Wierman. Conjectures and open problems. Lecture Notes in Mathematics, 671:170–176, 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063316/.

Smythe:1978:CFP

[SW78e] R. T. Smythe and John C. Wierman. Convergence of the first-passage processes. Lecture Notes in Mathematics, 671:78– 100, 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063311/.

Smythe:1978:DBP

[SW78f] R. T. Smythe and John C. Wierman. Definition of the basic processes and the existence of routes. Lecture Notes in Mathematics, 671:63–77, 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063310/.

Smythe:1978:FPP

[SW78g] R. T. Smythe and John C. Wierman. First-Passage Percolation on the Square Lattice, volume 671 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 196 pp. LCCN QA3 .L28 no. 671; QA3 .L28 no. 671 QA273.67. URL http:

//link.springer.com/book/10.1007/BFb0063306;   http://
www.springerlink.com/content/978-3-540-35744-5.

Smythe:1978:I

[SW78h] R. T. Smythe and John C. Wierman. Introduction. Lecture Notes in Mathematics, 671:1–7, 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063307/.

Smythe:1978:OPM

[SW78i] R. T. Smythe and John C. Wierman. Other percolation models on the square lattice. Lecture Notes in Mathematics, 671:158– 169, 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063315/.

Smythe:1978:P

[SW78j] R. T. Smythe and John C. Wierman. Preliminaries. Lecture Notes in Mathematics, 671:8–27, 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063308/.

Smythe:1978:RTP

[SW78k] R. T. Smythe and John C. Wierman. Renewal theory for percolation processes. Lecture Notes in Mathematics, 671:101–110, 1978. CODEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063312/.

Smythe:1978:RLH

[SW78l] R. T. Smythe and John C. Wierman. Route length and the height process. Lecture Notes in Mathematics, 671:128–157, 1978. CO-DEN LNMAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0063314/.

Smythe:1978:TC

[SW78m] R. T. Smythe and John C. Wierman. The time constant. Lecture Notes in Mathematics, 671:111–127, 1978. CODEN LN-MAA2. ISBN 3-540-08928-4 (print), 3-540-35744-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063313/.

Schempp:1977:CTF

[SZ77a] Prof. Dr. Walter Schempp and Prof. Dr. Karl Zeller, editors. Constructive Theory of Functions of Several Variables: Proceedings of a Conference Held at Oberwolfach April 25–May 1, 1976, volume 571 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08069-4 (print), 3-540-37496-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 571. URL http://link.springer.com/book/10.1007/
BFb0086559;   http://www.springerlink.com/content/978-
3-540-37496-1.

Serre:1977:MFOa

[SZ77b] Jean-Pierre Serre and Don Bernard Zagier, editors. Modular Functions of one Variable V: Proceedings International Conference, University of Bonn, Sonderforschungsbereich Theoretische Mathematik July 2–14, 1976, volume 601 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08348-0 (print), 3-540-37291-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 601. URL http:

//link.springer.com/book/10.1007/BFb0063941;   http://
www.springerlink.com/content/978-3-540-37291-2.

Serre:1977:MFOb

[SZ77c] Jean-Pierre Serre and Don Bernard Zagier, editors. Modular Functions of One Variable VI: Proceedings International Conference, University of Bonn, Sonderforschungsbereich Theoretische Mathematik July 2–14, 1976, volume 627 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 627. URL http:

//link.springer.com/book/10.1007/BFb0065293;   http://
www.springerlink.com/content/978-3-540-35984-5.

Szatzschneider:1975:MDV

[Sza75] W. Szatzschneider. A more deterministic version of Harris– Spitzer’s “Random constant velocity” model for infinite systems of particles. Lecture Notes in Mathematics, 472:157–168, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081952/.

Szafraniec:1978:APM

[Sza78] F. H. Szafraniec. Apropos of Professor Masani’s talk. Lecture Notes in Mathematics, 656:245–249, 1978. CODEN LN-MAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068823/.

Szczepaniak:1977:CT

[Szc77a] Zbigniew Szczepaniak. The consistency of the theory ZF + L^{1} ≠ HOD. Lecture Notes in Mathematics, 619:285–290, 1977. CODEN LNMAA2. ISBN 3-540-08521-1 (print), 3-540-37032-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067659/.

Szczyrba:1977:SSG

[Szc77b] Wiktor Szczyrba. On a symplectic structure of general relativity. Lecture Notes in Mathematics, 570:482–497, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087798/.

Szebehely:1975:RCM

[Sze75] V. Szebehely. Regularization in celestial mechanics. Lecture Notes in Mathematics, 461:257–263, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074156/.

Szeto:1976:PRA

[Sze76] George Szeto. The Pierce represencation of Azumaya algebras. Lecture Notes in Mathematics, 549:86–91, 1976. CODEN LN-MAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0077335/.

Szpiro:1976:COE

[Szp76] Lucien Szpiro. Cohomologie des ouvers de l’espace projectif sur un corps de caractéristique zéro. (French) []. Lecture Notes in Mathematics, 514:81–96, 1976. CODEN LNMAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080059/.

Szulga:1978:BCB

[Szu78] Jerzy Szulga. Boundedness and convergence of Banach lattice valued submartingales. Lecture Notes in Mathematics, 656:251– 256, 1978. CODEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068824/.

Tait:1975:RIT

[Tai75] William W. Tait. A realizability interpretation of the theory of species. Lecture Notes in Mathematics, 453:240–251, 1975. CO-DEN LNMAA2. ISBN 3-540-07155-5 (print), 3-540-37483-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0064875/.

Takahashi:1975:FJR

[Tak75a] Reiji Takahashi. Fonctions de Jacobi et représentations des groupes de Lie. (French) []. Lecture Notes in Mathematics, 497: 701–710, 1975. CODEN LNMAA2. ISBN 3-540-07537-2 (print), 3-540-38047-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0078028/.

Takeuti:1975:CPO

[Tak75b] Gaisi Takeuti. Consistency proofs and ordinals. Lecture Notes in Mathematics, 500:365–369, 1975. CODEN LNMAA2. ISBN 3-540-07533-X (print), 3-540-38020-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079562/.

Takens:1976:CES

[Tak76a] Floris Takens. Constrained equations; a study of implicit differential equations and their discontinuous solutions. Lecture Notes in Mathematics, 525:143–234, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077850/.

Takens:1976:IDE

[Tak76b] Floris Takens. Implicit differential equations: Some open problems. Lecture Notes in Mathematics, 535:237–253, 1976. CODEN LNMAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080503/.

Takahashi:1977:SF

[Tak77a] Reiji Takahashi. Spherical functions in spin_{0}(1, d)Spin(d − 1) for d = 2, 4 and 8. Lecture Notes in Mathematics, 587:226–240, 1977. CODEN LNMAA2. ISBN 3-540-08245-X (print), 3-540-37365-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087923/.

Takesaki:1977:FTI

[Tak77b] Masamichi Takesaki. Factors of type III. Lecture Notes in Mathematics, 575:139–156, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095708/.

Takemoto:1978:AAP

[Tak78a] Hideo Takemoto. C∗-algebras and applications to physics. Lecture Notes in Mathematics, 650:iii, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067404.

Takemoto:1978:ART

[Tak78b] Hideo Takemoto. On the algebraic reduction theory for countable direct summand C∗-algebras of separable C∗-algebras. Lecture Notes in Mathematics, 650:189–192, 1978. CODEN LN-MAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067403/.

Takenouchi:1978:NAF

[Tak78c] Osamu Takenouchi. A non-W ∗, AW*-factor. Lecture Notes in Mathematics, 650:135–139, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067395/.

Takeuti:1978:GNP

[Tak78d] Gaisi Takeuti. Gödel numbers of product spaces. Lecture Notes in Mathematics, 669:461–471, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103115/.

Takahashi:1979:QRA

[Tak79] Reiji Takahashi. Quelques résultats sur l’analyse harmonique dans l’espace symétrique non compact de rang I du type exceptionnel. (French) []. Lecture Notes in Mathematics, 739:511–567, 1979. CODEN LNMAA2. ISBN 3-540-09536-5 (print), 3-540-35023-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062505/.

Tammi:1978:AC

[Tam78a] Olli Tammi. Application to the coefficient a_{3} and a_{4}. Lecture Notes in Mathematics, 646:188–219, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070438/.

Tammi:1978:C

[Tam78b] Olli Tammi. The coefficient a_{3}. Lecture Notes in Mathematics, 646:148–162, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070436/.

Tammi:1978:EPB

[Tam78c] Olli Tammi. Extremum Problems for Bounded Univalent Functions, volume 646 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 313 pp. LCCN QA3 .L28 no. 646. URL http:

//link.springer.com/book/10.1007/BFb0070428;   http://
www.springerlink.com/content/978-3-540-35876-3.

Tammi:1978:FBB

[Tam78d] Olli Tammi. Functions with bounded boundary rotation. Lecture Notes in Mathematics, 646:99–116, 1978. CODEN LN-MAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070433/.

Tammi:1978:GC

[Tam78e] Olli Tammi. A generalization of the class S(b). Lecture Notes in Mathematics, 646:78–98, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070432/.

Tammi:1978:GLF

[Tam78f] Olli Tammi. Generalized Löwner-functions. Lecture Notes in Mathematics, 646:38–77, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070431/.

Tammi:1978:I

[Tam78g] Olli Tammi. Introduction. Lecture Notes in Mathematics, 646: 1–5, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070429/.

Tammi:1978:LF

[Tam78h] Olli Tammi. Löwner-functions. Lecture Notes in Mathematics, 646:6–37, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070430/.

Tammi:1978:GPI

[Tam78i] Olli Tammi. On generalizing the power inequality. Lecture Notes in Mathematics, 646:294–310, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070442/.

Tammi:1978:OC

[Tam78j] Olli Tammi. Optimization in the case N = 3. Lecture Notes in Mathematics, 646:266–286, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070440/.

Tammi:1978:OP

[Tam78k] Olli Tammi. Optimization of parameters for N = 1. Lecture Notes in Mathematics, 646:220–265, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070439/.

Tammi:1978:PIa

[Tam78l] Olli Tammi. The power inequality. Lecture Notes in Mathematics, 646:163–187, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070437/.

Tammi:1978:PIb

[Tam78m] Olli Tammi. The power inequality. Lecture Notes in Mathematics, 646:287–293, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070441/.

Tammi:1978:SCF

[Tam78n] Olli Tammi. The Schiffer-condition for S(b)-functions. Lecture Notes in Mathematics, 646:126–147, 1978. CODEN LN-MAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070435/.

Tammi:1978:VGF

[Tam78o] Olli Tammi. Variation of Green’s function. Lecture Notes in Mathematics, 646:117–125, 1978. CODEN LNMAA2. ISBN 3-540-08756-7 (print), 3-540-35876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070434/.

Tangora:1978:GCT

[Tan78] Martin C. Tangora. Generating Curtis tables. Lecture Notes in Mathematics, 673:243–253, 1978. CODEN LNMAA2. ISBN 3-540-08930-6 (print), 3-540-35737-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064700/.

Tartar:1978:NMR

[Tar78] L. Tartar. Une nouvelle méthode de résolution d’équations aux derivées partielles non linéaires. (French) []. Lecture Notes in Mathematics, 665:228–241, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061808/.

Tate:1975:ADT

[Tat75] J. Tate. Algorithm for determining the type of a singular fiber in an elliptic pencil. Lecture Notes in Mathematics, 476:33–52, 1975. CODEN LNMAA2. ISBN 3-540-07392-2 (print), 3-540-37588-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097582/.

Taylor:1977:GWM

[Tay77a] D. E. Taylor. Groups whose modular group rings have soluble unit groups. Lecture Notes in Mathematics, 573:112–117, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087816/.

Taylor:1977:TPB

[Tay77b] Joseph L. Taylor. Twisted products of Banach algebras and third čech cohomology. Lecture Notes in Mathematics, 575:157– 174, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095709/.

Taylor:1978:CWS

[Tay78a] Robert L. Taylor. Convergence of weighted sums in normed linear spaces. Lecture Notes in Mathematics, 672:108–145, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063211/.

Taylor:1978:GI

[Tay78b] Robert L. Taylor. General introduction. Lecture Notes in Mathematics, 672:1–5, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063206/.

Taylor:1978:LLNc

[Tay78c] Robert L. Taylor. Laws of large numbers in D[0, 1]. Lecture Notes in Mathematics, 672:153–184, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063213/.

Taylor:1978:LLNb

[Tay78d] Robert L. Taylor. Laws of large numbers in normed linear spaces. Lecture Notes in Mathematics, 672:71–107, 1978. CODEN LN-MAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063210/.

Taylor:1978:LLNa

[Tay78e] Robert L. Taylor. Laws of large numbers, uncorrelation, and convergence of weighted sums of random variables. Lecture Notes in Mathematics, 672:44–70, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063209/.

Taylor:1978:MP

[Tay78f] Robert L. Taylor. Mathematical preliminaries. Lecture Notes in Mathematics, 672:6–20, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063207/.

Taylor:1978:PA

[Tay78g] Robert L. Taylor. Possible applications. Lecture Notes in Mathematics, 672:185–201, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063214/.

Taylor:1978:REL

[Tay78h] Robert L. Taylor. Random elements in linear spaces. Lecture Notes in Mathematics, 672:21–43, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063208/.

Taylor:1978:RWS

[Tay78i] Robert L. Taylor. Randomly weighted sums. Lecture Notes in Mathematics, 672:146–152, 1978. CODEN LNMAA2. ISBN 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063212/.

Taylor:1978:SCW

[Tay78j] Robert L. Taylor. Stochastic Convergence of Weighted Sums of Random Elements in Linear Spaces, volume 672 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 0-387-08929-2, 3-540-08929-2 (print), 3-540-35740-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 216 pp. LCCN QA3 .L28 no. 672; QA3 .L28 no. 672 QA273.43. URL http:

//link.springer.com/book/10.1007/BFb0063205;   http://
www.springerlink.com/content/978-3-540-35740-7.

Taylor:1979:GK

[Tay79] Laurence R. Taylor. On the genera of knots. Lecture Notes in Mathematics, 722:144–154, 1979. CODEN LNMAA2. ISBN 3-540-09506-3 (print), 3-540-35186-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063196/.

tomDieck:1979:R

[tD79a] Tammo tom Dieck. λ-rings. Lecture Notes in Mathematics, 766: 27–69, 1979. CODEN LNMAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085968/.

tomDieck:1979:BRC

[tD79b] Tammo tom Dieck. The Burnside-ring of a compact Lie group. Lecture Notes in Mathematics, 766:82–158, 1979. CODEN LN-MAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085970/.

tomDieck:1979:BRF

[tD79c] Tammo tom Dieck. The Burnside ring of finite G-sets. Lecture Notes in Mathematics, 766:1–13, 1979. CODEN LNMAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085966/.

tomDieck:1979:EHC

[tD79d] Tammo tom Dieck. Equivariant homology and cohomology. Lecture Notes in Mathematics, 766:172–200, 1979. CODEN LN-MAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085972/.

tomDieck:1979:EHT

[tD79e] Tammo tom Dieck. Equivariant homotopy theory. Lecture Notes in Mathematics, 766:201–227, 1979. CODEN LNMAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085973/.

tomDieck:1979:GMB

[tD79f] Tammo tom Dieck. Geometric modules over the Burnside ring. Lecture Notes in Mathematics, 766:260–277, 1979. CODEN LN-MAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085975/.

tomDieck:1979:HEG

[tD79g] Tammo tom Dieck. Homotopy equivalent group representations. Lecture Notes in Mathematics, 766:228–259, 1979. CODEN LN-MAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085974/.

tomDieck:1979:HES

[tD79h] Tammo tom Dieck. Homotopy-equivalent stable G-vector bundles. Lecture Notes in Mathematics, 766:278–295, 1979. CODEN LNMAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085976/.

tomDieck:1979:IT

[tD79i] Tammo tom Dieck. Induction theory. Lecture Notes in Mathematics, 766:159–171, 1979. CODEN LNMAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085971/.

tomDieck:1979:HQF

[tD79j] Tammo tom Dieck. The J-homomorphism and quadratic forms. Lecture Notes in Mathematics, 766:14–26, 1979. CODEN LN-MAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085967/.

tomDieck:1979:PR

[tD79k] Tammo tom Dieck. Permutation representations. Lecture Notes in Mathematics, 766:70–81, 1979. CODEN LNMAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085969/.

tomDieck:1979:TGR

[tD79l] Tammo tom Dieck. Transformation Groups and Representation Theory, volume 766 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09720-1 (print), 3-540-38517-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 309 pp. LCCN QA3 .L28 no. 766. URL http:

//link.springer.com/book/10.1007/BFb0085965;   http://
www.springerlink.com/content/978-3-540-38517-2.

Topsoe:1976:TEC

[TDHJ76] Flemming Topsøe, Richard M. Dudley, and Jørgen Hoffmann-Jørgensen. Two examples concerning uniform convergence of measures w.r.t. balls in Banach spaces. Lecture Notes in Mathematics, 566:141–146, 1976. CODEN LNMAA2. ISBN 3-540-08061-9 (print), 3-540-37515-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096885/.

tomDieck:1979:HSF

[tDP79] T. tom Dieck and T. Petrie. The homotopy structure of finite group actions on spheres. Lecture Notes in Mathematics, 741: 222–243, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062142/.

Teissier:1975:TFG

[Tei75] Bernard Teissier. Théorèmes de finitude de géométrie analytique [d’après Heisuke Hironaka]. (French) []. Lecture Notes in Mathematics, 431:295–317, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066376/.

Temam:1976:AAC

[Tem76a] R. Temam. Applications de l’analyse convexe au calcul des variations. (French) []. Lecture Notes in Mathematics, 543:208– 237, 1976. CODEN LNMAA2. ISBN 3-540-07867-3 (print), 3-540-38075-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079945/.

Temam:1976:LES

[Tem76b] R. Temam. Local existence of ℓ^{∞} solutions of the Euler equations of incompressible perfect fluids. Lecture Notes in Mathematics, 565:184–194, 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091456/.

Temam:1976:TNS

[Tem76c] Roger Temam, editor. Turbulence and Navier Stokes Equations: Proceedings of the Conference Held at the University of Paris-Sud Orsay June 10–13 1975, volume 565 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08060-0 (print), 3-540-37516-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 565. URL http:

//link.springer.com/book/10.1007/BFb0091442;   http://
www.springerlink.com/content/978-3-540-37516-6.

Teply:1976:PSS

[Tep76] Mark L. Teply. Prime singular-splitting rings with finiteness conditions. Lecture Notes in Mathematics, 545:173–194, 1976. CODEN LNMAA2. ISBN 3-540-07985-8 (print), 3-540-37983-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080311/.

Teply:1979:S

[Tep79] Mark Teply. Subidealizers. Lecture Notes in Mathematics, 700: 232–234, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063475/.

Thompson:1975:TTT

[TH75] Gerald Thompson and William R. Hare. Tverberg-type theorems in convex product structures. Lecture Notes in Mathematics, 490:212–217, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0081144/.

Thimm:1977:DPF

[Thi77] W. Thimm. Sur les décompositions primaires des faisceaux analytiques cohérents. (French) []. Lecture Notes in Mathematics, 578:2–13, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091459/.

Thom:1976:TFW

[Tho76a] R. Thom. The two-fold way of catastrophe theory. Lecture Notes in Mathematics, 525:235–252, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077851/.

Thomas:1976:TSF

[Tho76b] G. Erik F. Thomas. Totally summable functions with values in locally convex spaces. Lecture Notes in Mathematics, 541:117– 131, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081045/.

Thomas:1977:FRB

[Tho77] Glânffrwd P. Thomas. Further results on Baxter sequences and generalized Schur functions. Lecture Notes in Mathematics, 579: 155–167, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090016/.

Thomas:1978:IRC

[Tho78] Erik Thomas. Integral representations in conuclear spaces. Lecture Notes in Mathematics, 645:172–179, 1978. CODEN LN-MAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069674/.

Thom:1979:CAQ

[Tho79a] R. Thom. Sur certains aspects qualitatifs de la théorie des équations aux derivées partielles. (French) []. Lecture Notes in Mathematics, 704:3–7, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063610/.

Thomason:1979:FQS

[Tho79b] R. W. Thomason. First quadrant spectral sequences in algebraic K-theory. Lecture Notes in Mathematics, 763:332–355, 1979. CO-DEN LNMAA2. ISBN 3-540-09721-X (print), 3-540-38520-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0088093/.

Thouvenot:1979:DFD

[Tho79c] Jean-Paul Thouvenot. La démonstration de Furstenberg du théorème de Szemerédi sur les progressions arithmétiques. (French) []. Lecture Notes in Mathematics, 710:221–232, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069981/.

Tikhonov:1979:PMP

[Tik79] A. N. Tikhonov. Les problèmes mal posés et les problèmes Mathématiques de traitement automatique de résultats d’Experiences. (French) []. Lecture Notes in Mathematics, 704:8–20, 1979. CO-DEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063611/.

Tits:1977:TMS

[Tit77] Jacques Tits. Travaux de Margulis sur les sous-groupes discrets de groupes de Lie. (French) []. Lecture Notes in Mathematics, 567: 174–190, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096069/.

Tits:1978:GWG

[Tit78] Jacques Tits. Groupes de Whitehead de groupes algébriques simples sur un corps [d’après V. P. Platonov et al.]. (French) []. Lecture Notes in Mathematics, 677:218–236, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070765/.

terMorsche:1976:RBF

[tM76] Hennie ter Morsche. On the relations between finite differences and derivatives of cardinal spline functions. Lecture Notes in Mathematics, 501:210–219, 1976. CODEN LNMAA2. ISBN 3-540-07543-7 (print), 3-540-38073-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079749/.

Tognoli:1976:DTR

[Tog76] A. Tognoli. A desingularisation theorem for real analytic varieties. Lecture Notes in Mathematics, 524:156–162, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0078004/.

Toledo:1975:FON

[Tol75a] Dr. Sue Toledo. First order number theory. Lecture Notes in Mathematics, 447:31–152, 1975. CODEN LNMAA2. ISBN 3-540-07149-0 (print), 3-540-37442-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062938/.

Toledo:1975:I

[Tol75b] Dr. Sue Toledo. Introduction. Lecture Notes in Mathematics, 447: 1–29, 1975. CODEN LNMAA2. ISBN 3-540-07149-0 (print), 3-540-37442-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062937/.

Toledo:1975:OHO

[Tol75c] Dr. Sue Toledo. Other higher order systems due to Schütte. Lecture Notes in Mathematics, 447:215–289, 1975. CODEN LN-MAA2. ISBN 3-540-07149-0 (print), 3-540-37442-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062940/.

Toledo:1975:SOL

[Tol75d] Dr. Sue Toledo. Second order logic. Lecture Notes in Mathematics, 447:153–214, 1975. CODEN LNMAA2. ISBN 3-540-07149-0 (print), 3-540-37442-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062939/.

Toledo:1975:TSF

[Tol75e] Dr. Sue Toledo. Tableau Systems for First Order Number Theory and Certain Higher Order Theories, volume 447 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07149-0 (print), 3-540-37442-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 339 pp. LCCN QA3

.L28 no. 447. URL http://link.springer.com/book/10.1007/
BFb0062936;   http://www.springerlink.com/content/978-
3-540-37442-8.

Toom:1978:MER

[Too78] A. L. Toom. Monotonic evolutions in real spaces. Lecture Notes in Mathematics, 653:1–14, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070080/.

Topsoe:1976:PCB

[Top76] Flemming Topsøe. Packings and coverings with balls in finite dimensional normed spaces. Lecture Notes in Mathematics, 541: 187–198, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081052/.

Tougeron:1976:MSA

[Tou76] J. C. Tougeron. Modules stratifiables sur l’anneau des germes de fonctions numériques c^{∞} à l’origine de R^{n}. (French) [Stratifiable modules on the ring of seeds of numerical functions c^{∞} at the origin of R^{n}]. Lecture Notes in Mathematics, 535:98–127, 1976. CODEN LNMAA2. ISBN 3-540-07853-3 (print), 3-540-38182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080498/.

Trautmann:1976:DCA

[Tra76] Günther Trautmann. Deformations of coherent analytic sheaves with finite singularities. Lecture Notes in Mathematics, 524:1– 20, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077994/.

Trautmann:1978:DMC

[Tra78a] Günther Trautmann. Deformations and moduli of coherent analytic sheaves with finite singularities. Lecture Notes in Mathematics, 670:233–302, 1978. CODEN LNMAA2. ISBN 3-540-08927-6 (print), 3-540-35745-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064404/.

Trautmann:1978:DSB

[Tra78b] Günther Trautmann. Deformations of sheaves and bundles. Lecture Notes in Mathematics, 683:29–41, 1978. CODEN LN-MAA2. ISBN 3-540-08949-7 (print), 3-540-35710-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0063172/.

Treves:1975:SOF

[Trè75] François Trèves. Second-order Fuchsian elliptic equations and eigenvalue asymptotics. Lecture Notes in Mathematics, 459:283– 340, 1975. CODEN LNMAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074199/.

Trnkova:1977:CAU

[Trn77] Věra Trnková. Categorial aspects are useful for topology. Lecture Notes in Mathematics, 609:211–225, 1977. CODEN LN-MAA2. ISBN 3-540-08437-1 (print), 3-540-37108-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068686/.

Trotter:1978:KMS

[Tro78] H. F. Trotter. Knot modules and Seifert matrices. Lecture Notes in Mathematics, 685:291–299, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062978/.

Truss:1976:SCK

[Tru76] J. Truss. Some cases of Künig’s lemma. Lecture Notes in Mathematics, 537:273–284, 1976. CODEN LNMAA2. ISBN 3-540-07856-8 (print), 3-540-38122-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096907/.

Truss:1978:NIS

[Tru78] J. Truss. A note on increasing sequences of constructibility degrees. Lecture Notes in Mathematics, 669:473–476, 1978. CODEN LNMAA2. ISBN 3-540-08926-8 (print), 3-540-35749-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103116/.

Tuan:1978:LDP

[Tua78] Pham Dinh Tuan. L’adéquation du processus multivariable gaussien continu stationnaire centre de densité spectrale rationnelle. (French) []. Lecture Notes in Mathematics, 636:141– 162, 1978. CODEN LNMAA2. ISBN 3-540-08658-7 (print), 3-540-35925-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063263/.

Tulczyjew:1977:SFF

[Tul77a] W. M. Tulczyjew. A symplectic formulation of field dynamics. Lecture Notes in Mathematics, 570:464–468, 1977. CODEN LN-MAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0087796/.

Tulczyjew:1977:SFP

[Tul77b] W. M. Tulczyjew. A symplectic formulation of particle dynamics. Lecture Notes in Mathematics, 570:457–463, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087795/.

Turpin:1975:EOE

[Tur75] Philippe Turpin. Espaces et opérateurs exponentiellement galbés. (French) []. Lecture Notes in Mathematics, 474:48–62, 1975. CO-DEN LNMAA2. ISBN 3-540-07189-X (print), 3-540-37554-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077398/.

Turgi:1977:STI

[Tur77] Mary Turgi. A sheaf-theoretic interpretation of the Kuroš theorem. Lecture Notes in Mathematics, 616:173–196, 1977. CODEN LNMAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068196/.

Turpin:1978:BPF

[Tur78] Philippe Turpin. Boundedness problems for finitely additive measures. Lecture Notes in Mathematics, 645:180–187, 1978. CODEN LNMAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069675/.

Tutte:1975:GCG

[Tut75] W. T. Tutte. The graph of the chromial of a graph. Lecture Notes in Mathematics, 452:55–61, 1975. CODEN LNMAA2. ISBN 3-540-07154-7 (print), 3-540-37482-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069543/.

Thirring:1978:UEF

[TW78a] W. Thirring and R. Wallner. The use of exterior forms in field theory. Lecture Notes in Mathematics, 676:171–178, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063670/.

Thu:1978:ENS

[TW78b] Nguyẽn Văn Thu and A. Weron. Examples of non-stationary Banach space valued stochastic processes of second order. Lecture Notes in Mathematics, 656:171–181, 1978. CODEN LN-MAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068818/.

Taylor:1979:SSF

[TW79] Laurence Taylor and Bruce Williams. Surgery spaces: Formulae and structure. Lecture Notes in Mathematics, 741:170–195, 1979. CODEN LNMAA2. ISBN 3-540-09545-4 (print), 3-540-35009-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062139/.

Tomiyama:1978:GSE

[TY78] Jun Tomiyama and Kôzô Yabuta. General short exact sequence theorem for Toeplitz operators of uniform algebras. Lecture Notes in Mathematics, 650:126–130, 1978. CODEN LNMAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067393/.

Ueno:1975:AFK

[Uen75a] Dr. Kenji Ueno. Addition formula for Kodaira dimensions of analytic fibre bundles. Lecture Notes in Mathematics, 439:172– 187, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070576/.

Ueno:1975:ARC

[Uen75b] Dr. Kenji Ueno. Algebraic reductions of complex varieties and complex manifolds of algebraic dimension zero. Lecture Notes in Mathematics, 439:141–171, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070575/.

Ueno:1975:ASA

[Uen75c] Dr. Kenji Ueno. Analytic spaces and algebraic varieties. Lecture Notes in Mathematics, 439:1–27, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070571/.

Ueno:1975:CV

[Uen75d] Dr. Kenji Ueno. Chapter VIII. Lecture Notes in Mathematics, 439:240–247, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070578/.

Ueno:1975:CAV

[Uen75e] Dr. Kenji Ueno. Classification of algebraic varieties and complex varieties. Lecture Notes in Mathematics, 439:94–140, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070574/.

Ueno:1975:CTA

[Uen75f] Dr. Kenji Ueno. Classification Theory of Algebraic Varieties and Compact Complex Spaces: Notes written in collaboration with P. Cherenack, volume 439 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xix + 278 pp. LCCN QA3 .L28 no. 439. URL http:

//link.springer.com/book/10.1007/BFb0070570;   http://
www.springerlink.com/content/978-3-540-37415-2.

Ueno:1975:DKD

[Uen75g] Dr. Kenji Ueno. D-dimensions and Kodaira dimensions. Lecture Notes in Mathematics, 439:28–75, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070572/.

Ueno:1975:ECM

[Uen75h] Dr. Kenji Ueno. Examples of complex manifolds. Lecture Notes in Mathematics, 439:188–239, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070577/.

Ueno:1975:FT

[Uen75i] Dr. Kenji Ueno. Fundamental theorems. Lecture Notes in Mathematics, 439:76–93, 1975. CODEN LNMAA2. ISBN 3-540-07138-5 (print), 3-540-37415-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070573/.

Urbanik:1975:EPM

[Urb75a] K. Urbanik. Extreme point method in probability theory. Lecture Notes in Mathematics, 472:169–194, 1975. CODEN LN-MAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081953/.

Urbanik:1975:SSP

[Urb75b] K. Urbanik. Stable symmetric probability laws in quantum mechanics. Lecture Notes in Mathematics, 472:195–206, 1975. CO-DEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0081954/.

UtrerasD:1979:CVT

[Utr79] F. Utreras D. Cross-validation techniques for smoothing spline functions in one or two dimensions. Lecture Notes in Mathematics, 757:196–232, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098498/.

Vaaler:1975:TTR

[Vaa75] J. Vaaler. A Tauberian theorem related to Weyl’s criterion. Lecture Notes in Mathematics, 475:253–258, 1975. CODEN LN-MAA2. ISBN 3-540-07388-4 (print), 3-540-37579-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0074267/.

Vaillant:1978:PSM

[Vai78] Jean Vaillant. Propriété de symétrie des matrices localisées d’une matrice fortement hyperbolique en un point multiple. (French) []. Lecture Notes in Mathematics, 660:190–205, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082305/.

Valentine:1975:AMS

[Val75] Joseph E. Valentine. Angles in metric spaces. Lecture Notes in Mathematics, 490:66–73, 1975. CODEN LNMAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081132/.

VanDaele:1978:FPC

[Van78a] A. Van Daele. Fixed points and commutation theorems. Lecture Notes in Mathematics, 650:140–144, 1978. CODEN LN-MAA2. ISBN 3-540-08762-1 (print), 3-540-35850-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067396/.

VanDeVen:1978:ECA

[Van78b] A. Van De Ven. On the Enriques classification of algebraic surfaces. Lecture Notes in Mathematics, 677:237–251, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070766/.

VanDeVen:1978:SRR

[Van78c] A. Van De Ven. Some recent results on surfaces of general type. Lecture Notes in Mathematics, 677:155–166, 1978. CODEN LN-MAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070760/.

VandenEssen:1979:RSD

[Van79] Arno Van den Essen. Reduction of singularities of the differential equation Ady = Bdx. Lecture Notes in Mathematics, 712: 44–59, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062814/.

Varopoulos:1976:CBA

[Var76] N. Th. Varopoulos. On a class of Banach algebras. Lecture Notes in Mathematics, 512:180–184, 1976. CODEN LNMAA2. ISBN 3-540-07682-4 (print), 3-540-38201-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080033/.

Varadarajan:1977:HAR

[Var77a] V. S. Varadarajan. Harmonic Analysis on Real Reductive Groups, volume 576 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08135-6 (print), 3-540-37420-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 521 pp. LCCN QA3 .L28 no. 576. URL http:

//link.springer.com/book/10.1007/BFb0097814;   http://
www.springerlink.com/content/978-3-540-37420-6.

Varadarajan:1977:IAR

[Var77b] V. S. Varadarajan. Invariant analysis on a real reductive Lie algebra. Lecture Notes in Mathematics, 576:1–174, 1977. CODEN LNMAA2. ISBN 3-540-08135-6 (print), 3-540-37420-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097815/.

Varadarajan:1977:IAD

[Var77c] V. S. Varadarajan. Invariant analysis on, and discrete series for real reductive groups. Lecture Notes in Mathematics, 576:175– 521, 1977. CODEN LNMAA2. ISBN 3-540-08135-6 (print), 3-540-37420-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097816/.

Varchenko:1978:NPE

[Var78] A. N. Varchenko. Newton polygons and estimates of oscillatory integrals. Lecture Notes in Mathematics, 677:29, 1978. CODEN LNMAA2. ISBN 3-540-08937-3 (print), 3-540-35719-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070751.

Vasilyev:1978:BMS

[Vas78] N. B. Vasilyev. Bernoulli and Markov stationary measures in discrete local interactions. Lecture Notes in Mathematics, 653:99– 112, 1978. CODEN LNMAA2. ISBN 3-540-08450-9 (print), 3-540-37044-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070087/.

vanderAa:1979:RPO

[vdAS79] Els van der Aa and Jan A. Sanders. The 1: 2: 1-resonance, its periodic orbits and integrals. Lecture Notes in Mathematics, 711:

187–208, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062954/.

vanderBurgh:1979:AVP

[vdB79] A. H. P. van der Burgh. On the asymptotic validity of perturbation methods for hyperbolic differential equations. Lecture Notes in Mathematics, 711:229–240, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062956/.

vandenEssen:1979:RSA

[vdE79] A. R. P. van den Essen. Regular singularities along normal crossings. Lecture Notes in Mathematics, 712:87–130, 1979. CODEN LNMAA2. ISBN 3-540-09250-1 (print), 3-540-35314-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062818/.

vanderKallen:1976:IS

[vdK76] Wilberd van der Kallen. Injective stability for K_{2}. Lecture Notes in Mathematics, 551:77–154, 1976. CODEN LNMAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079997/.

vanderLaan:1979:RAA

[vdLT79] G. van der Laan and A. J. J. Talman. A restart algorithm without an artificial level for computing fixed points on unbounded regions. Lecture Notes in Mathematics, 730:247–256, 1979. CODEN LNMAA2. ISBN 3-540-09518-7 (print), 3-540-35129-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064322/.

Vekua:1976:ENA

[Vek76] I. N. Vekua. Über einige neuere Anwendungen der verallgemeinerten Cauchy–Riemannschen Cleichungen in der Schalentheorie. (German) []. Lecture Notes in Mathematics, 561:88–99, 1976. CODEN LNMAA2. ISBN 3-540-08054-6 (print), 3-540-37536-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087628/.

Vergne:1975:RUG

[Ver75] Michèle Vergne. Représentations unitaires des groupes de Lie résolubles. (French) []. Lecture Notes in Mathematics, 431:205– 226, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066372/.

Verdier:1976:TRR

[Ver76] Jean-Louis Verdier. Le théorème de Riemann–Roch pour les variétés algébriques éventuellement singulières. (French) []. Lecture Notes in Mathematics, 514:159–175, 1976. CODEN LN-MAA2. ISBN 3-540-07686-7 (print), 3-540-38218-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080065/.

Verdier:1977:CDQ

[Ver77] Jean-Louis Verdier. Catégories derivées quelques résultats (État 0). (French) []. Lecture Notes in Mathematics, 569:262–311, 1977. CODEN LNMAA2. ISBN 3-540-08066-X (print), 3-540-37507-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091525/.

Verdier:1979:EDA

[Ver79a] Jean-Louis Verdier. Équations différentielles algébriques. (French) []. Lecture Notes in Mathematics, 710:101–122, 1979. CODEN LNMAA2. ISBN 3-540-09243-9 (print), 3-540-35340-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069975/.

Verhulst:1979:DJB

[Ver79b] F. Verhulst. The description of jumps between Kepler orbits by boundary layer methods. Lecture Notes in Mathematics, 711:176– 186, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062953/.

Verhulst:1979:AAT

[Ver79c] Ferdinand Verhulst. Asymptotic Analysis: From Theory to Application, volume 711 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 240 pp. LCCN QA3 .L28 no. 711. URL http:

//link.springer.com/book/10.1007/BFb0062943;   http://
www.springerlink.com/content/978-3-540-35332-4.

vonGolitschek:1976:ADP

[vG76] Manfred von Golitschek. Approximation durch Polynome mit ganzzahligen Koeffizienten. (German) []. Lecture Notes in Mathematics, 556:201–212, 1976. CODEN LNMAA2. ISBN 3-540-08001-5 (print), 3-540-37552-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087408/.

Vodopianov:1979:BVQ

[VGK79] S. K. Vodop′ianov, V. M. Gol′dšteĭn, and A. P. Kopylov. On boundary values of quasiconformal mappings in space. Lecture Notes in Mathematics, 743:238–261, 1979. CODEN LN-MAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079497/.

vanHarten:1979:AAM

[vH79a] Aart van Harten. Asymptotic approximations in magnetohydrodynamic singular perturbation problems. Lecture Notes in Mathematics, 711:94–124, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062949/.

vanHarten:1979:FBC

[vH79b] Aart van Harten. Feed-back control of singularly perturbed heating problems. Lecture Notes in Mathematics, 711:33–62, 1979. CODEN LNMAA2. ISBN 3-540-09245-5 (print), 3-540-35332-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062946/.

Viennot:1977:FGC

[Vie77] G. Viennot. Une forme géométrique de la correspondance de Robinson–Schensted. (French) []. Lecture Notes in Mathematics, 579:29–58, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090011/.

Viennot:1978:ALL

[Vie78a] Gérard Viennot. Algèbres de Lie Libres et Monoïdes Libres: Bases des Algèbres de Lie Libres et Factorisations des Monoïdes Libres, volume 691 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-09090-8 (print), 3-540-35601-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 124 pp. LCCN QA3 .L28 no. 691. URL http:

//link.springer.com/book/10.1007/BFb0067950;   http://
www.springerlink.com/content/978-3-540-35601-1.

Viennot:1978:BAL

[Vie78b] Gérard Viennot. Bases des algèbres de Lie libres. (French) []. Lecture Notes in Mathematics, 691:9–61, 1978. CODEN LN-MAA2. ISBN 3-540-09090-8 (print), 3-540-35601-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067952/.

Viennot:1978:BML

[Vie78c] Gérard Viennot. Bissections des monoïdes libres. (French) []. Lecture Notes in Mathematics, 691:62–93, 1978. CODEN LN-MAA2. ISBN 3-540-09090-8 (print), 3-540-35601-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067953/.

Viennot:1978:FML

[Vie78d] Gérard Viennot. Factorisations des monoïdes libres et familles basiques des algèbres de Lie libres. (French) []. Lecture Notes in Mathematics, 691:94–112, 1978. CODEN LNMAA2. ISBN 3-540-09090-8 (print), 3-540-35601-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067954/.

Viennot:1978:IFI

[Vie78e] Gérard Viennot. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 691:1–8, 1978. CODEN LNMAA2. ISBN 3-540-09090-8 (print), 3-540-35601-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067951/.

Vigneras:1977:FGE

[Vig77a] M.-F. Vignéras. Facteurs gamma et équations fonctionnelles. (French) []. Lecture Notes in Mathematics, 627:79–103, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065298/.

Vigneras:1977:STF

[Vig77b] Marie-France Vignéras. Séries thêta des formes quadratiques indéfinies. (French) []. Lecture Notes in Mathematics, 627:227– 239, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065303/.

vonKutschera:1975:ISN

[vK75] Franz von Kutschera. Intensional semantics for natural language. Lecture Notes in Mathematics, 499:445–459, 1975. CODEN LN-MAA2. ISBN 3-540-07534-8 (print), 3-540-38022-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0079428/.

vanMoerbeke:1979:AID

[vM79] Pierre van Moerbeke. About isospectral deformations of discrete Laplacians. Lecture Notes in Mathematics, 755:313–370, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069810/.

vanOystaeyen:1975:ASP

[vO75a] Prof. Freddy M. J. van Oystaeyen. Application: The symmetric part of the Brauer group. Lecture Notes in Mathematics, 444: 86–100, 1975. CODEN LNMAA2. ISBN 3-540-07146-6 (print), 3-540-37438-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068140/.

vanOystaeyen:1975:GL

[vO75b] Prof. Freddy M. J. van Oystaeyen. Generalities on localization. Lecture Notes in Mathematics, 444:4–15, 1975. CODEN LN-MAA2. ISBN 3-540-07146-6 (print), 3-540-37438-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068136/.

vanOystaeyen:1975:PSN

[vO75c] Prof. Freddy M. J. van Oystaeyen. Prime Spectra in Non-Commutative Algebra, volume 444 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07146-6 (print), 3-540-37438-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 100 (est.) pp. LCCN QA3

.L28 no. 444. URL http://link.springer.com/book/10.1007/
BFb0068134;   http://www.springerlink.com/content/978-
3-540-37438-1.

vanOystaeyen:1975:PAF

[vO75d] Prof. Freddy M. J. van Oystaeyen. Primes in algebras over fields. Lecture Notes in Mathematics, 444:58–85, 1975. CODEN LN-MAA2. ISBN 3-540-07146-6 (print), 3-540-37438-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068139/.

vanOystaeyen:1975:S

[vO75e] Prof. Freddy M. J. van Oystaeyen. Sheaves. Lecture Notes in Mathematics, 444:42–57, 1975. CODEN LNMAA2. ISBN 3-540-07146-6 (print), 3-540-37438-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068138/.

vanOystaeyen:1975:SKF

[vO75f] Prof. Freddy M. J. van Oystaeyen. Symmetric kernel functors. Lecture Notes in Mathematics, 444:16–41, 1975. CODEN LN-MAA2. ISBN 3-540-07146-6 (print), 3-540-37438-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068137/.

vanOystaeyen:1975:SLS

[vO75g] Prof. Freddy M. J. van Oystaeyen. Symmetric localization and sheaves. Lecture Notes in Mathematics, 444:1–3, 1975. CODEN LNMAA2. ISBN 3-540-07146-6 (print), 3-540-37438-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068135/.

Vogt:1978:CKG

[Vog78] Renate Vogt. Cobordismus von knoten. (German) []. Lecture Notes in Mathematics, 685:218–226, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062973/.

Voigt:1977:KBG

[Voi77a] Detlef Voigt. Das Kriterium von Blattner. (German) []. Lecture Notes in Mathematics, 592:75–77, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086130/.

Voigt:1977:RVP

[Voi77b] Detlef Voigt. Das Rechnen mit verschränkten Produkten. (German) []. Lecture Notes in Mathematics, 592:369–377, 1977. CO-DEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086146/.

Voigt:1977:SDA

[Voi77c] Detlef Voigt. Das Schema der Darstellungen einer Algebra. (German) []. Lecture Notes in Mathematics, 592:257–275, 1977. CO-DEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086138/.

Voigt:1977:SZM

[Voi77d] Detlef Voigt. Das Schema der Zerlegungen eines Moduls. (German) []. Lecture Notes in Mathematics, 592:275–280, 1977. CO-DEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086139/.

Voigt:1977:VTG

[Voi77e] Detlef Voigt. Das verallgemeinerte Taylorlemma. (German) []. Lecture Notes in Mathematics, 592:280–284, 1977. CODEN LN-MAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086140/.

Voigt:1977:STS

[Voi77f] Detlef Voigt. Der Satz von Taketa in der Situation endlicher, algebraischer Gruppen. (German) []. Lecture Notes in Mathematics, 592:187–201, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086135/.

Voigt:1977:ZMG

[Voi77g] Detlef Voigt. Der Zerlegungssatz von Mackey. (German) []. Lecture Notes in Mathematics, 592:306–326, 1977. CODEN LN-MAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086143/.

Voigt:1977:SSS

[Voi77h] Detlef Voigt. Die Sätze von Shoda in der Situation endlicher, algebraischer Gruppen. (German) []. Lecture Notes in Mathematics, 592:167–186, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086134/.

Voigt:1977:EDA

[Voi77i] Detlef Voigt. Einfache Darstellungen von auflösbaren, algebraischen Gruppen. (German) []. Lecture Notes in Mathematics, 592: 77–148, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086131/.

Voigt:1977:EID

[Voi77j] Detlef Voigt. Endomorphismenringe induzierter Darstellungen als verschränkte Produkte. (German) []. Lecture Notes in Mathematics, 592:377–407, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086147/.

Voigt:1977:EFL

[Voi77k] Detlef Voigt. Exkurs über die Frattinialgebra einer p-Liealgebra. (German) []. Lecture Notes in Mathematics, 592:243–256, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086137/.

Voigt:1977:FIM

[Voi77l] Detlef Voigt. Filtrierungen auf induzierten Moduln. (German) []. Lecture Notes in Mathematics, 592:327–348, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086144/.

Voigt:1977:FKB

[Voi77m] Detlef Voigt. Folgerungen aus dem Kriterium von Blattner. (German) []. Lecture Notes in Mathematics, 592:149–154, 1977. CO-DEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086132/.

Voigt:1977:IDT

[Voi77n] Detlef Voigt. Induzierte Darstellungen in der Theorie der endlichen, algebraischen Gruppen. (German) [], volume 592 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LN-MAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 413 pp. LCCN QA3

.L28 no. 592. URL http://link.springer.com/book/10.1007/
BFb0086128;   http://www.springerlink.com/content/978-
3-540-37355-1.

Voigt:1977:IMS

[Voi77o] Detlef Voigt. Induzierte Moduln, die von stabilen Moduln herrühren. (German) []. Lecture Notes in Mathematics, 592:287– 305, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086142/.

Voigt:1977:IKD

[Voi77p] Detlef Voigt. Induzierte und Koinduzierte Darstellungen. (German) []. Lecture Notes in Mathematics, 592:284–287, 1977. CO-DEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0086141/.

Voigt:1977:I

[Voi77q] Detlef Voigt. Introduction. Lecture Notes in Mathematics, 592: 1–74, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086129/.

Voigt:1977:MGG

[Voi77r] Detlef Voigt. Monomiale Gruppen. (German) [monomial groups]. Lecture Notes in Mathematics, 592:155–166, 1977. CODEN LN-MAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0086133/.

Voigt:1977:SLD

[Voi77s] Detlef Voigt. Satz und Lemma von Dade in der Situation endlicher, algebraischer Gruppen. (German) []. Lecture Notes in Mathematics, 592:202–242, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086136/.

Voigt:1977:VPG

[Voi77t] Detlef Voigt. Verschränkte Produkte. (German) []. Lecture Notes in Mathematics, 592:349–369, 1977. CODEN LNMAA2. ISBN 3-540-08251-4 (print), 3-540-37355-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086145/.

Voigt:1979:GAI

[Voi79] Detlef Voigt. Groupes algébriques infinitesimaux résolubles et leurs représentations linéaires irreductibles. (French) []. Lecture Notes in Mathematics, 740:48–68, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071052/.

Volger:1975:CTL

[Vol75a] Hugo Volger. Completeness theorem for logical categories. Lecture Notes in Mathematics, 445:51–86, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061293/.

Volger:1975:LCS

[Vol75b] Hugo Volger. Logical categories, semantical categories and topoi. Lecture Notes in Mathematics, 445:87–100, 1975. CODEN LN-MAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061294/.

Vergne:1975:CAS

[VR75] Michèle Vergne and Hugo Rossi. Continuation analytique de la série discréte holomorphe. (French) []. Lecture Notes in Mathematics, 466:198–225, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082208/.

vanRossum:1979:OEI

[vR79] H. van Rossum. Orthogonal expansions in indefinite inner product spaces. Lecture Notes in Mathematics, 765:172–183, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085580/.

Vuorinen:1979:AVA

[Vuo79] Matti Vuorinen. Asymptotic values and angular limits of quasiregular mappings. Lecture Notes in Mathematics, 743:262–267, 1979. CODEN LNMAA2. ISBN 3-540-09550-0 (print), 3-540-34861-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079498/.

Vo-Van-Tan:1977:CEC

[VVT77] Vo-Van-Tan. La classification des espaces 1-convexes. (French) []. Lecture Notes in Mathematics, 578:71–92, 1977. CODEN LN-MAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0091463/.

vonWeizsacker:1976:SNR

[vW76] Heinrich von Weizsäcker. Some negative results in theory of lifting. Lecture Notes in Mathematics, 541:159–172, 1976. CODEN LNMAA2. ISBN 3-540-07861-4 (print), 3-540-38107-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081049/.

vanWesep:1978:WDD

[vW78] Robert van Wesep. Wadge degrees and descriptive set theory. Lecture Notes in Mathematics, 689:151–170, 1978. CODEN LN-MAA2. ISBN 3-540-09086-X (print), 3-540-35626-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069298/.

Wachspress:1978:ICD

[Wac78] Eugene L. Wachspress. Isojacobic crosswind differencing. Lecture Notes in Mathematics, 630:190–199, 1978. CODEN LN-MAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067706/.

Wagoner:1976:CCA

[Wag76] J. B. Wagoner. Continuous cohomology and p-adic K-theory. Lecture Notes in Mathematics, 551:241–248, 1976. CODEN LN-MAA2. ISBN 3-540-07996-3 (print), 3-540-37964-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080004/.

Wagoner:1977:CPI

[Wag77] J. B. Wagoner. H-cobordisms, pseudo-isotopies, and analytic torsion. Lecture Notes in Mathematics, 575:175–191, 1977. CODEN LNMAA2. ISBN 3-540-08133-X (print), 3-540-37423-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095710/.

Wagschal:1978:PCC

[Wag78] Claude Wagschal. Problème de Cauchy à caractéristiques multiples dans les classes de Gevrey. (French) []. Lecture Notes in Mathematics, 660:206–216, 1978. CODEN LNMAA2. ISBN 3-540-08913-6 (print), 3-540-35764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0082306/.

Wagstaff:1979:AB

[Wag79] Samuel S. Wagstaff, Jr. Additive h-bases for n. Lecture Notes in Mathematics, 751:302–327, 1979. CODEN LNMAA2. ISBN 3-540-09559-4 (print), 3-540-34852-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062717/.

Wahlquist:1976:BTP

[Wah76] Hugo D. Wahlquist. Bäcklund transformation of potentials of the Korteweg–de Vries equation and the interaction of solitons with cnoidal waves. Lecture Notes in Mathematics, 515:162–183, 1976. CODEN LNMAA2. ISBN 3-540-07687-5 (print), 3-540-38220-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081167/.

Wahba:1979:CRT

[Wah79] Grace Wahba. Convergence rates of ‘thin plate’ smoothing splines when the data are noisy. Lecture Notes in Mathematics, 757:233– 245, 1979. CODEN LNMAA2. ISBN 3-540-09706-6 (print), 3-540-38475-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098499/.

Wainwright:1975:FR

[Wai75] W. L. Wainwright. A functional relation. Lecture Notes in Mathematics, 457:298–305, 1975. CODEN LNMAA2. ISBN 3-540-07161-X (print), 3-540-69975-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067115/.

Wallach:1975:URH

[Wal75a] Nolan R. Wallach. On the unitarizability of representations with highest weights. Lecture Notes in Mathematics, 466:226–231, 1975. CODEN LNMAA2. ISBN 3-540-07183-0 (print), 3-540-37524-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0082209/.

Walters:1975:E

[Wal75b] Peter Walters. Entropy. Lecture Notes in Mathematics, 458:70– 111, 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112506/.

Walters:1975:ES

[Wal75c] Peter Walters. Equilibrium states. Lecture Notes in Mathematics, 468:32–35, 1975. CODEN LNMAA2. ISBN 3-540-07184-9 (print), 3-540-37525-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082590/.

Walters:1975:ETI

[Wal75d] Peter Walters. Ergodic Theory — Introductory Lectures, volume 458 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 198 pp. LCCN QA3 .L28 no. 458. URL http:

//link.springer.com/book/10.1007/BFb0112501;   http://
www.springerlink.com/content/978-3-540-37494-7.

Walters:1975:ISI

[Wal75e] Peter Walters. Isomorphism and spectral invariants. Lecture Notes in Mathematics, 458:51–62, 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112504/.

Walters:1975:MPTa

[Wal75f] Peter Walters. Measure-preserving transformations. Lecture Notes in Mathematics, 458:16–50, 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112503/.

Walters:1975:MPTb

[Wal75g] Peter Walters. Measure-preserving transformations with pure point spectrum. Lecture Notes in Mathematics, 458:63–69, 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112505/.

Walters:1975:P

[Wal75h] Peter Walters. Preliminaries. Lecture Notes in Mathematics, 458: 1–15, 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112502/.

Walters:1975:TD

[Wal75i] Peter Walters. Topological dynamics. Lecture Notes in Mathematics, 458:112–139, 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112507/.

Walters:1975:TE

[Wal75j] Peter Walters. Topological entropy. Lecture Notes in Mathematics, 458:140–184, 1975. CODEN LNMAA2. ISBN 3-540-07163-6 (print), 3-540-37494-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112508/.

Waldschmidt:1976:PAF

[Wal76] Michel Waldschmidt. Propriétes arithmétiques de fonctions de plusieurs variables (i). (French) []. Lecture Notes in Mathematics, 524:106–129, 1976. CODEN LNMAA2. ISBN 3-540-07787-1 (print), 3-540-38251-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0078002/.

Waldschmidt:1977:TGV

[Wal77a] Michel Waldschmidt. Les travaux de G. V. Čudnovskiĭ sur les nombres transcendants. (French) []. Lecture Notes in Mathematics, 567:274–292, 1977. CODEN LNMAA2. ISBN 3-540-08064-3 (print), 3-540-37514-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096075/.

Waldschmidt:1977:PAF

[Wal77b] Michel Waldschmidt. Propriétés arithmétiques de fonctions de plusieurs variables (II). (French) []. Lecture Notes in Mathematics, 578:108–135, 1977. CODEN LNMAA2. ISBN 3-540-08256-5 (print), 3-540-37341-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091465/.

Waldvogel:1978:BVP

[Wal78a] J. Waldvogel. Boundary value problems in infinite intervals. Lecture Notes in Mathematics, 631:201–208, 1978. CODEN LN-MAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0067473/.

Wall:1978:CCM

[Wal78b] G. E. Wall. Commutator collection and module structure. Lecture Notes in Mathematics, 697:174–196, 1978. CODEN LN-MAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103129/.

Wall:1978:LMG

[Wal78c] G. E. Wall. Lie methods in group theory. Lecture Notes in Mathematics, 697:137–173, 1978. CODEN LNMAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103128/.

Walsh:1978:VMI

[Wal78d] John B. Walsh. Vector measures and the Itô integral. Lecture Notes in Mathematics, 645:188–197, 1978. CODEN LN-MAA2. ISBN 3-540-08669-2 (print), 3-540-35903-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0069676/.

Wamsley:1977:CSG

[Wam77] J. W. Wamsley. Computing soluble groups. Lecture Notes in Mathematics, 573:118–125, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087817/.

Wambecq:1979:SPR

[Wam79] A. Wambecq. Some properties of rational methods for solving ordinary differential equations. Lecture Notes in Mathematics, 765: 352–365, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085589/.

Wang:1975:RPM

[Wan75] Frank J. S. Wang. A random product of Markovian semi-groups of operators. Lecture Notes in Mathematics, 451:63–81, 1975. CODEN LNMAA2. ISBN 3-540-07153-9 (print), 3-540-37481-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068579/.

Wan:1979:BPO

[Wan79] Yieh-Hei Wan. Bifurcations of periodic orbits in autonomous systems. Lecture Notes in Mathematics, 755:371–377, 1979. CODEN LNMAA2. ISBN 3-540-09703-1 (print), 3-540-38462-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069811/.

Warfield:1976:CSC

[War76a] Robert B. Warfield, Jr. Completions and the structure of complete groups. Lecture Notes in Mathematics, 513:52–60, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080159/.

Warfield:1976:GAC

[War76b] Robert B. Warfield, Jr. Groups with abelian central quotient. Lecture Notes in Mathematics, 513:26–42, 1976. CODEN LN-MAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080157/.

Warfield:1976:HPF

[War76c] Robert B. Warfield, Jr. The Hall–Petresco formula and residual boundedness. Lecture Notes in Mathematics, 513:43–51, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080158/.

Warfield:1976:IRC

[War76d] Robert B. Warfield, Jr. Idempotent radicals on the category of nilpotent groups. Lecture Notes in Mathematics, 513:17–25, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080156/.

Warfield:1976:L

[War76e] Robert B. Warfield, Jr. Localization. Lecture Notes in Mathematics, 513:61–76, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080160/.

Warfield:1976:MC

[War76f] Robert B. Warfield, Jr. The Malcev correspondence. Lecture Notes in Mathematics, 513:104–111, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080164/.

Warfield:1976:NAK

[War76g] Robert B. Warfield, Jr. Nilpotent actions, Kolchin’s theorem, and Engel conditions. Lecture Notes in Mathematics, 513:77–82, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0080161/.

Warfield:1976:NG

[War76h] Robert B. Warfield, Jr. Nilpotent Groups, volume 513 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 111 (est.) pp. LCCN QA3

.L28 no. 513. URL http://link.springer.com/book/10.1007/
BFb0080152;   http://www.springerlink.com/content/978-
3-540-38205-8.

Warfield:1976:NGA

[War76i] Robert B. Warfield, Jr. Nilpotent groups admitting exponents in a ring. Lecture Notes in Mathematics, 513:83–95, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080162/.

Warfield:1976:R

[War76j] Robert B. Warfield, Jr. Rudiments. Lecture Notes in Mathematics, 513:1–5, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080153/.

Warfield:1976:TPL

[War76k] Robert B. Warfield, Jr. Tensor products and the lower central series. Lecture Notes in Mathematics, 513:9–16, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080155/.

Warfield:1976:URM

[War76l] Robert B. Warfield, Jr. Unipotent representations and Malcev completions of R-groups. Lecture Notes in Mathematics, 513:96– 103, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0080163/.

Warfield:1976:UCS

[War76m] Robert B. Warfield, Jr. The upper central series. Lecture Notes in Mathematics, 513:6–8, 1976. CODEN LNMAA2. ISBN 3-540-07683-2 (print), 3-540-38205-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080154/.

Ward:1977:NTC

[War77a] J. N. Ward. A note on the Todd–Coxeter algorithm. Lecture Notes in Mathematics, 573:126–129, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087818/.

Warfield:1977:SMA

[War77b] Robert B. Warfield, Jr. The structure of mixed abelian groups. Lecture Notes in Mathematics, 616:1–38, 1977. CODEN LN-MAA2. ISBN 3-540-08447-9 (print), 3-540-37069-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068187/.

Warfield:1979:SGM

[War79a] R. B. Warfield, Jr. Stable generation of modules. Lecture Notes in Mathematics, 700:16–33, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063457/.

Warfield:1979:MFB

[War79b] Robert B. Warfield, Jr. Modules over fully bounded Noetherian rings. Lecture Notes in Mathematics, 734:339–352, 1979. CODEN LNMAA2. ISBN 3-540-09529-2 (print), 3-540-35043-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103167/.

Wassermann:1976:SUC

[Was76] Gordon Wassermann. (r, s)-stable unfoldings and catastrophe theory. Lecture Notes in Mathematics, 525:253–262, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077852/.

Watson:1976:NAP

[Wat76] G. Alistair Watson, editor. Numerical Analysis: Proceedings of the Dundee Conference on Numerical Analysis, [July 1–4,] 1975, volume 506 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07610-7 (print), 3-540-38129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 506. URL http:

//link.springer.com/book/10.1007/BFb0080109;   http://
www.springerlink.com/content/978-3-540-38129-7.

Watkins:1978:CIG

[Wat78a] Mark E. Watkins. On the Cayley index of a group. Lecture Notes in Mathematics, 686:101–105, 1978. CODEN LNMAA2. ISBN 3-540-08953-5 (print), 3-540-35702-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062521/.

Watson:1978:NAP

[Wat78b] G. A. Watson, editor. Numerical Analysis: Proceedings of the Biennial Conference Held at Dundee, June 28–July 1, 1977, volume 630 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 630; QA1 .L471 v.630. URL http:

//link.springer.com/book/10.1007/BFb0067690;   http://
www.springerlink.com/content/978-3-540-35972-2.

Wirthmuller:1976:TSS

[WdPL76] Christopher G. Gibson Klaus Wirthmüller, Andrew A. du Plessis, and Eduard J. N. Looijenga. Topological Stability of Smooth Mappings, volume 552 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07997-1 (print), 3-540-37957-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 149 (est.) pp. LCCN QA3 .L28 no. 552. URL http:

//link.springer.com/book/10.1007/BFb0095244;   http://
www.springerlink.com/content/978-3-540-37957-7.

Weber:1977:CUP

[Web77] Michel Weber. Classes uniformes de processus gaussiens stationnaires. (French) []. Lecture Notes in Mathematics, 581:196– 256, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087194/.

Webb:1978:SNE

[Web78a] J. R. L. Webb. Strongly nonlinear elliptic equations. Lecture Notes in Mathematics, 665:242–256, 1978. CODEN LNMAA2. ISBN 3-540-08922-5 (print), 3-540-35756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061809/.

Weber:1978:A

[Web78b] Claude Weber. Appendix. Lecture Notes in Mathematics, 685: 270–273, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062976/.

Weber:1978:P

[Web78c] Claude Weber. Problems. Lecture Notes in Mathematics, 685: 309–311, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0062980/.

Weber:1978:TMA

[Web78d] Claude Weber. Torsion dans les modules d’Alexander. (French) []. Lecture Notes in Mathematics, 685:300–308, 1978. CODEN LNMAA2. ISBN 3-540-08952-7 (print), 3-540-35705-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0062979/.

Weck:1975:PEF

[Wec75] Norbert Weck. Über das Prinzip der Eindeutigen Fortsetzbarkeit in der Kontrolltheorie. (German) []. Lecture Notes in Mathematics, 477:276–284, 1975. CODEN LNMAA2. ISBN 3-540-07393-0 (print), 3-540-37591-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079182/.

Wedig:1978:ISM

[Wed78] W. Wedig. On the integration of sequences of moments’ equations in the stability theory of stochastic systems. Lecture Notes in Mathematics, 695:229–238, 1978. CODEN LNMAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0062669/.

Weidmann:1975:STP

[Wei75a] Joachim Weidmann. Spectral theory of partial differential operators. Lecture Notes in Mathematics, 448:71–111, 1975. CODEN LNMAA2. ISBN 3-540-07150-4 (print), 3-540-37444-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067081/.

Weil:1975:CJN

[Wei75b] André Weil. La cyclotomie jadis et naguère. (French) [The cyclotomy once and for all]. Lecture Notes in Mathematics, 431: 318–338, 1975. CODEN LNMAA2. ISBN 3-540-07023-0 (print), 3-540-37310-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066377/.

Weinstein:1975:MQC

[Wei75c] Alan Weinstein. On Maslov’s quantization condition. Lecture Notes in Mathematics, 459:341–372, 1975. CODEN LN-MAA2. ISBN 3-540-07180-6 (print), 3-540-37521-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074200/.

Weisfeiler:1976:APC

[Wei76a] Boris Weisfeiler. Algebraic properties of cellular algebras. Lecture Notes in Mathematics, 558:76–98, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089386/.

Weisfeiler:1976:ACS

[Wei76b] Boris Weisfeiler. An algorithm of construction of strongly regular graphs. Lecture Notes in Mathematics, 558:151–158, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089394/.

Weisfeiler:1976:AGC

[Wei76c] Boris Weisfeiler. An algorithm of graph canonization. Lecture Notes in Mathematics, 558:138–144, 1976. CODEN LN-MAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089392/.

Weisfeiler:1976:CAA

[Wei76d] Boris Weisfeiler. Cellular algebras arising in the theory of permutation groups. Lecture Notes in Mathematics, 558:38–40, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089380/.

Weisfeiler:1976:CSG

[Wei76e] Boris Weisfeiler. A construction of a stationary graph. Lecture Notes in Mathematics, 558:13–22, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089377/.

Weisfeiler:1976:CQC

[Wei76f] Boris Weisfeiler. Construction of the quotient in the case of cellular algebras of rank greater than one. Lecture Notes in Mathematics, 558:58–59, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089383.

Weisfeiler:1976:DS

[Wei76g] Boris Weisfeiler. Deep stabilization. Lecture Notes in Mathematics, 558:111–122, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089389/.

Weisfeiler:1976:ERU

[Wei76h] Boris Weisfeiler. Examples of results using the stability of depth 1. Lecture Notes in Mathematics, 558:123–128, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089390/.

Weisfeiler:1976:ICC

[Wei76i] Boris Weisfeiler. Imprimitive cells and construction of factorcells. Lecture Notes in Mathematics, 558:48–57, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089382/.

Weisfeiler:1976:KSR

[Wei76j] Boris Weisfeiler. Kernels and stability with respect to kernels. Lecture Notes in Mathematics, 558:104–110, 1976. CODEN LN-MAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089388/.

Weisfeiler:1976:M

[Wei76k] Boris Weisfeiler. Motivation. Lecture Notes in Mathematics, 558: 5–12, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089376/.

Weisfeiler:1976:CIG

[Wei76l] Boris Weisfeiler. On Construction and Identification of Graphs, volume 558 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 237 pp. LCCN QA3 .L28 no. 558. URL http:

//link.springer.com/book/10.1007/BFb0089374;   http://
www.springerlink.com/content/978-3-540-37539-5.

Weisfeiler:1976:SCS

[Wei76m] Boris Weisfeiler. On the structure of correct stationary graphs and cells having more than one normal subcell. Lecture Notes in Mathematics, 558:60–70, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089384/.

Weisfeiler:1976:PAG

[Wei76n] Boris Weisfeiler. A practical algorithm of graph canonization. Lecture Notes in Mathematics, 558:145–150, 1976. CODEN LN-MAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089393/.

Weisfeiler:1976:PC

[Wei76o] Boris Weisfeiler. Properties of cells. Lecture Notes in Mathematics, 558:23–28, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089378/.

Weisfeiler:1976:PCA

[Wei76p] Boris Weisfeiler. Properties of cellular algebras of rank greater than one. Lecture Notes in Mathematics, 558:29–37, 1976. CO-DEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0089379/.

Weisfeiler:1976:PPC

[Wei76q] Boris Weisfeiler. Properties of primitive cells. Lecture Notes in Mathematics, 558:71–75, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089385/.

Weisfeiler:1976:SCC

[Wei76r] Boris Weisfeiler. Some classes of cellular algebras. Lecture Notes in Mathematics, 558:41–47, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089381/.

Weisfeiler:1976:SDE

[Wei76s] Boris Weisfeiler. Some definitions and explanations about exhaustive search. Lecture Notes in Mathematics, 558:129–137, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089391/.

Weisfeiler:1976:SMS

[Wei76t] Boris Weisfeiler. Some modifications of stabilization. Lecture Notes in Mathematics, 558:99–103, 1976. CODEN LN-MAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089387/.

Weisfeiler:1976:SRA

[Wei76u] Boris Weisfeiler. Some remarks about the problem of graph identification. Lecture Notes in Mathematics, 558:1–4, 1976. CODEN LNMAA2. ISBN 3-540-08051-1 (print), 3-540-37539-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089375/.

Weinberger:1978:ABM

[Wei78a] H. F. Weinberger. Asymptotic behavior of a model in population genetics. Lecture Notes in Mathematics, 648:47–96, 1978. CODEN LNMAA2. ISBN 3-540-08759-1 (print), 3-540-35868-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066406/.

Weiss:1978:FCT

[Wei78b] Gary Weiss. The Fuglede commutativity theorem modulo the Hilbert–Schmidt class and generating functions for matrix operators. Lecture Notes in Mathematics, 693:175–184, 1978. CODEN LNMAA2. ISBN 3-540-09097-5 (print), 3-540-35557-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064675/.

Weron:1975:PTB

[Wer75] A. Weron. Prediction theory in Banach spaces. Lecture Notes in Mathematics, 472:207–228, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081955/.

Weron:1978:PTV

[Wer78] A. Weron, editor. Probability Theory on Vector Spaces: Proceedings, Trzebieszowice, Poland, September 1977, volume 656 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CO-DEN LNMAA2. ISBN 3-540-08846-6 (print), 3-540-35814-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 656; QA3 .L28 no. 656 QA273.43. URL http:

//link.springer.com/book/10.1007/BFb0068804;   http://
www.springerlink.com/content/978-3-540-35814-5.

Werner:1979:RMR

[Wer79] H. Werner. A reliable method for rational interpolation. Lecture Notes in Mathematics, 765:257–277, 1979. CODEN LN-MAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0085585/.

Whyburn:1978:RMF

[Why78] Lucille Whyburn. R. L. Moore’s first doctoral student at Texas. Lecture Notes in Mathematics, 664:33–37, 1978. CODEN LN-MAA2. ISBN 3-540-08920-9 (print), 3-540-35758-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0061687/.

Widlund:1978:CMM

[Wid78] Olof Widlund. Capacitance matrix methods for Helmholtz’s equation on general bounded regions. Lecture Notes in Mathematics, 631:209–219, 1978. CODEN LNMAA2. ISBN 3-540-08539-4 (print), 3-540-35970-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067474/.

Wiegand:1979:RBM

[Wie79] Roger Wiegand. Rings of bounded module type. Lecture Notes in Mathematics, 700:143–150, 1979. CODEN LNMAA2. ISBN 3-540-09107-6 (print), 3-540-35538-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063465/.

Wilcox:1975:STD

[Wil75a] Calvin H. Wilcox. Scattering Theory for the d’Alembert Equation in Exterior Domains, volume 442 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 0-387-07144-X, 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 184 pp. LCCN QA3

.L28 no. 442. URL http://link.springer.com/book/10.1007/
BFb0070581;   http://www.springerlink.com/content/978-
3-540-37429-9.

Wilcox:1975:AWF

[Wil75b] Dr. Calvin H. Wilcox. Asymptotic wave functions and energy distributions in exterior domains. Lecture Notes in Mathematics, 442:142–166, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070589/.

Wilcox:1975:I

[Wil75c] Dr. Calvin H. Wilcox. Introduction. Lecture Notes in Mathematics, 442:3–13, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070582/.

Wilcox:1975:SDEa

[Wil75d] Dr. Calvin H. Wilcox. Solutions of the d’Alembert equation in ∝^{n}. Lecture Notes in Mathematics, 442:14–38, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070583/.

Wilcox:1975:SDEb

[Wil75e] Dr. Calvin H. Wilcox. Solutions of the d’Alembert equation in arbitrary domains. Lecture Notes in Mathematics, 442:39–48, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070584/.

Wilcox:1975:SSSb

[Wil75f] Dr. Calvin H. Wilcox. Steady-state scattering theory and eigenfunction expansions for A. Lecture Notes in Mathematics, 442: 84–123, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070587/.

Wilcox:1975:SSSa

[Wil75g] Dr. Calvin H. Wilcox. Steady-state scattering theory in exterior domains and the limiting absorption principle. Lecture Notes in Mathematics, 442:49–75, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070585/.

Wilcox:1975:TDS

[Wil75h] Dr. Calvin H. Wilcox. Time-dependent scattering theory in exterior domains. Lecture Notes in Mathematics, 442:76–83, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070586/.

Wilcox:1975:WOA

[Wil75i] Dr. Calvin H. Wilcox. Wave operators and asymptotic solutions of the d’Alembert equation in exterior domains. Lecture Notes in Mathematics, 442:124–141, 1975. CODEN LNMAA2. ISBN 3-540-07144-X (print), 3-540-37429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070588/.

Williams:1976:SBM

[Wil76a] David Williams. On a stopped Brownian motion formula of H. M. Taylor. Lecture Notes in Mathematics, 511:235–239, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101110/.

Williams:1976:MP

[Wil76b] David Williams. The Q-matrix problem. Lecture Notes in Mathematics, 511:216–234, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101109/.

Williams:1977:LVS

[Wil77a] R. F. Williams. Lecture VII: The structure of Lorenz attractors. Lecture Notes in Mathematics, 615:94–112, 1977. CODEN LN-MAA2. ISBN 3-540-08445-2 (print), 3-540-37074-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068363/.

Williamson:1977:ORR

[Wil77b] S. G. Williamson. On the ordering, ranking, and random generation of basic combinatorial sets. Lecture Notes in Mathematics, 579:309–339, 1977. CODEN LNMAA2. ISBN 3-540-08143-7 (print), 3-540-37385-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090026/.

Wilson:1977:LSG

[Wil77c] John S. Wilson. Locally soluble groups satisfying the minimal condition for normal subgroups. Lecture Notes in Mathematics, 573:130–142, 1977. CODEN LNMAA2. ISBN 3-540-08131-3 (print), 3-540-37428-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087819/.

Wilson:1978:YGT

[Wil78a] Robin J. Wilson. 200 years of graph theory — a guided tour. Lecture Notes in Mathematics, 642:1–9, 1978. CODEN LN-MAA2. ISBN 3-540-08666-8 (print), 3-540-35912-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0070360/.

Wilson:1978:IRL

[Wil78b] William H. Wilson. Induced representations of life algebras. Lecture Notes in Mathematics, 697:197–204, 1978. CODEN LN-MAA2. ISBN 3-540-09103-3 (print), 3-540-35549-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0103130/.

Winkler:1975:MCS

[Win75] Peter M. Winkler. Model-completeness and Skolem expansions. Lecture Notes in Mathematics, 498:408–463, 1975. CODEN LN-MAA2. ISBN 3-540-07538-0 (print), 3-540-38057-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0080988/.

Wirthmuller:1976:SF

[Wir76] Klaus Wirthmüller. Stratifications and flows. Lecture Notes in Mathematics, 552:35–63, 1976. CODEN LNMAA2. ISBN 3-540-07997-1 (print), 3-540-37957-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095247/.

Wirz:1978:ISM

[Wir78] H. J. Wirz. On iterative solution methods for systems of partial differential equations. Lecture Notes in Mathematics, 679: 151–163, 1978. CODEN LNMAA2. ISBN 3-540-08940-3 (print), 3-540-35715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067875/.

Wexler-Kreindler:1978:PTE

[WK78] Elena Wexler-Kreindler. Propriétés de transfert des extensions d’Ore. (French) []. Lecture Notes in Mathematics, 641:235– 251, 1978. CODEN LNMAA2. ISBN 3-540-08665-X (print), 3-540-35913-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064851/.

Wexler-Kreindler:1979:SFT

[WK79] Elena Wexler-Kreindler. Series formelles tordues et conditions de chaines. (French) []. Lecture Notes in Mathematics, 740:99– 119, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071056/.

Wolfe:1975:MDS

[Wol75] Dorothy Wolfe. Metric dependence and a sum of distances. Lecture Notes in Mathematics, 490:206–211, 1975. CODEN LN-MAA2. ISBN 3-540-07417-1 (print), 3-540-37946-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081143/.

Wolf:1978:RAM

[Wol78] Joseph A. Wolf. Representations associated to minimal co-adjoint orrits. Lecture Notes in Mathematics, 676:329–349, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063679/.

Wong:1976:OTS

[Won76a] Yau-Chuen Wong. Orders and topologies on spaces consisting of families. Lecture Notes in Mathematics, 531:45–103, 1976. CO-DEN LNMAA2. ISBN 3-540-07800-2 (print), 3-540-38268-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077706/.

Wong:1976:SCT

[Won76b] Yau-Chuen Wong. Some characterizations of the topology of uniform convergence on order-bounded sets. Lecture Notes in Mathematics, 531:104–155, 1976. CODEN LNMAA2. ISBN 3-540-07800-2 (print), 3-540-38268-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077707/.

Wong:1976:SOV

[Won76c] Yau-Chuen Wong. A survey of ordered vector spaces. Lecture Notes in Mathematics, 531:1–44, 1976. CODEN LNMAA2. ISBN 3-540-07800-2 (print), 3-540-38268-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077705/.

Wong:1976:TUC

[Won76d] Yau-Chuen Wong. The Topology of Uniform Convergence on Order-Bounded Sets, volume 531 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNMAA2. ISBN 3-540-07800-2 (print), 3-540-38268-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 163 pp. LCCN QA3

.L28 no. 531. URL http://link.springer.com/book/10.1007/
BFb0077704;   http://www.springerlink.com/content/978-
3-540-38268-3.

Wong:1979:GN

[Won79a] Yau-Chuen Wong. General notations. Lecture Notes in Mathematics, 726:1–7, 1979. CODEN LNMAA2. ISBN 3-540-09513-6 (print), 3-540-35161-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062864/.

Wong:1979:NS

[Won79b] Yau-Chuen Wong. Nuclear spaces. Lecture Notes in Mathematics, 726:133–232, 1979. CODEN LNMAA2. ISBN 3-540-09513-6 (print), 3-540-35161-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062867/.

Wong:1979:SS

[Won79c] Yau-Chuen Wong. Schwartz spaces. Lecture Notes in Mathematics, 726:8–51, 1979. CODEN LNMAA2. ISBN 3-540-09513-6 (print), 3-540-35161-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062865/.

Wong:1979:SSN

[Won79d] Yau-Chuen Wong. Schwartz Spaces, Nuclear Spaces and Tensor Products, volume 726 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09513-6 (print), 3-540-35161-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 418 pp. LCCN QA3 .L28 no. 726. URL http:

//link.springer.com/book/10.1007/BFb0062863;   http://
www.springerlink.com/content/978-3-540-35161-0.

Wong:1979:TP

[Won79e] Yau-Chuen Wong. Tensor products. Lecture Notes in Mathematics, 726:233–342, 1979. CODEN LNMAA2. ISBN 3-540-09513-6 (print), 3-540-35161-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0062868/.

Wong:1979:TPO

[Won79f] Yau-Chuen Wong. Tensor products of ordered convex spaces. Lecture Notes in Mathematics, 726:343–403, 1979. CODEN LN-MAA2. ISBN 3-540-09513-6 (print), 3-540-35161-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062869/.

Wong:1979:VSS

[Won79g] Yau-Chuen Wong. Vector sequence spaces and absolutely summing mappings. Lecture Notes in Mathematics, 726:52–132, 1979. CODEN LNMAA2. ISBN 3-540-09513-6 (print), 3-540-35161-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0062866/.

Wood:1978:IC

[Woo78a] R. J. Wood. V -indexed categories. Lecture Notes in Mathematics, 661:126–140, 1978. CODEN LNMAA2. ISBN 3-540-08914-4 (print), 3-540-35762-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061362/.

Woodhouse:1978:GQW

[Woo78b] Nicholas Woodhouse. Geometric quantization and the WKB approximation. Lecture Notes in Mathematics, 676:295–309, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063676/.

Wormald:1979:CCC

[Wor79] Nicholas C. Wormald. Classifying K-connected cubic graphs. Lecture Notes in Mathematics, 748:199–206, 1979. CODEN LN-MAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0102696/.

Woyczynski:1975:GMB

[Woy75] W. A. Woyczyński. Geometry and martingales in Banach spaces. Lecture Notes in Mathematics, 472:229–275, 1975. CODEN LN-MAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0081956/.

Wraith:1975:LET

[Wra75] G. C. Wraith. Lectures on elementary topoi. Lecture Notes in Mathematics, 445:114–206, 1975. CODEN LNMAA2. ISBN 3-540-07164-4 (print), 3-540-37495-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061296/.

Wu:1978:IGT

[Wu78] Tai Tsun Wu. Introduction to gauge theory. Lecture Notes in Mathematics, 676:161–169, 1978. CODEN LNMAA2. ISBN 3-540-08935-7 (print), 3-540-35721-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063669/.

Wuytack:1979:CBT

[Wuy79a] L. Wuytack. Commented bibliography on techniques for computing Padé approximants. Lecture Notes in Mathematics, 765: 375–392, 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085591/.

Wuytack:1979:PAA

[Wuy79b] Luc Wuytack, editor. Padé Approximation and its Applications: Proceedings of a Conference held in Antwerp, Belgium, 1979, volume 765 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNMAA2. ISBN 3-540-09717-1 (print), 3-540-38511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 765. URL http:

//link.springer.com/book/10.1007/BFb0085571;   http://
www.springerlink.com/content/978-3-540-38511-0.

Yamada:1976:USS

[Yam76] Toshio Yamada. On the uniqueness of solutions of stochastic differential equations with reflecting barrier conditions. Lecture Notes in Mathematics, 511:240–244, 1976. CODEN LNMAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101111/.

Yamada:1978:CSE

[Yam78a] Toshio Yamada. Sur une construction des solutions d’équations différentielles stochastiques dans le cas non-lipschitzien. (French) []. Lecture Notes in Mathematics, 649:114–131, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064599/.

Yamaguchi:1978:FEP

[Yam78b] Hiroshi Yamaguchi. Fonctions entières paraboliques dans C^{2}. (French) []. Lecture Notes in Mathematics, 694:325–334, 1978. CODEN LNMAA2. ISBN 3-540-09101-7 (print), 3-540-35552-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063256/.

Yammine:1979:TCS

[Yam79] Sleiman Yammine. Les théorèmes de Cohen–Seidenberg en algèbre non commutative. (French) []. Lecture Notes in Mathematics, 740: 120–169, 1979. CODEN LNMAA2. ISBN 3-540-09537-3 (print), 3-540-35020-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071057/.

Yanagihara:1977:ARG

[Yan77a] Hiroshi Yanagihara. Adjoint representations of group schemes. Lecture Notes in Mathematics, 614:142–153, 1977. CODEN LN-MAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068387/.

Yanagihara:1977:AHS

[Yan77b] Hiroshi Yanagihara. Algebraic Hopf subalgebras. Lecture Notes in Mathematics, 614:97–112, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068384/.

Yanagihara:1977:ASW

[Yan77c] Hiroshi Yanagihara. Algebraic schemes on which a group scheme operates. Lecture Notes in Mathematics, 614:216–230, 1977. CO-DEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0068393/.

Yanagihara:1977:CHSa

[Yan77d] Hiroshi Yanagihara. Centralizers of Hopf subalgebras and group subschemes. Lecture Notes in Mathematics, 614:166–174, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068389/.

Yanagihara:1977:CF

[Yan77e] Hiroshi Yanagihara. Coalgebras over a field. Lecture Notes in Mathematics, 614:29–44, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068379/.

Yanagihara:1977:CHSb

[Yan77f] Hiroshi Yanagihara. Commutators of Hopf subalgebras and group subschemes. Lecture Notes in Mathematics, 614:175–190, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068390/.

Yanagihara:1977:FGF

[Yan77g] Hiroshi Yanagihara. Formal groups over a field. Lecture Notes in Mathematics, 614:68–81, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068382/.

Yanagihara:1977:GSF

[Yan77h] Hiroshi Yanagihara. Group schemes over a field. Lecture Notes in Mathematics, 614:1–14, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068377/.

Yanagihara:1977:GSG

[Yan77i] Hiroshi Yanagihara. Group subschemes of a group scheme. Lecture Notes in Mathematics, 614:15–28, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068378/.

Yanagihara:1977:HAF

[Yan77j] Hiroshi Yanagihara. Hopf algebras over a field. Lecture Notes in Mathematics, 614:45–56, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068380/.

Yanagihara:1977:IGV

[Yan77k] Hiroshi Yanagihara. Isogenies of group varieties. Lecture Notes in Mathematics, 614:231–242, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068394/.

Yanagihara:1977:JIG

[Yan77l] Hiroshi Yanagihara. Joins and intersections of group subschemes. Lecture Notes in Mathematics, 614:126–141, 1977. CODEN LN-MAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068386/.

Yanagihara:1977:KPI

[Yan77m] Hiroshi Yanagihara. Kernels of purely inseparable isogenies. Lecture Notes in Mathematics, 614:243–257, 1977. CODEN LN-MAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068395/.

Yanagihara:1977:LIS

[Yan77n] Hiroshi Yanagihara. Left invariant semi-derivations of group varieties. Lecture Notes in Mathematics, 614:203–215, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068392/.

Yanagihara:1977:LAA

[Yan77o] Hiroshi Yanagihara. Lie algebras attached to group schemes. Lecture Notes in Mathematics, 614:191–202, 1977. CODEN LN-MAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068391/.

Yanagihara:1977:NHS

[Yan77p] Hiroshi Yanagihara. Normalizers of Hopf subalgebras and group subschemes. Lecture Notes in Mathematics, 614:154–165, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068388/.

Yanagihara:1977:RRG

[Yan77q] Hiroshi Yanagihara. Rational representations of group schemes. Lecture Notes in Mathematics, 614:113–125, 1977. CODEN LN-MAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068385/.

Yanagihara:1977:SCB

[Yan77r] Hiroshi Yanagihara. Structures of colocal bigebras. Lecture Notes in Mathematics, 614:57–67, 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068381/.

Yanagihara:1977:SFG

[Yan77s] Hiroshi Yanagihara. Structures of formal groups as local rings. Lecture Notes in Mathematics, 614:82–96, 1977. CODEN LN-MAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0068383/.

Yanagihara:1977:THA

[Yan77t] Hiroshi Yanagihara. Theory of Hopf Algebras Attached to Group Schemes, volume 614 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNMAA2. ISBN 3-540-08444-4 (print), 3-540-37077-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 308 pp. LCCN QA3 .L28 no. 614. URL http:

//link.springer.com/book/10.1007/BFb0068376;   http://
www.springerlink.com/content/978-3-540-37077-2.

Yamaguti:1979:NDS

[YF79] Masaya Yamaguti and Hiroshi Fujii. On numerical deformation of singularities in nonlinear elasticity. Lecture Notes in Mathematics, 704:267–281, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063625/.

Yor:1976:TPP

[YM76] Marc Yor and P.-A. Meyer. Sur la théorie de la prediction, et le problème de decomposition des tribus F_{t+}^{o}. (French) []. Lecture Notes in Mathematics, 511:104–117, 1976. CODEN LN-MAA2. ISBN 3-540-07681-6 (print), 3-540-38197-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101099/.

Yoeurp:1979:SAS

[Yoe79] Chantha Yoeurp. Sauts additifs et sauts multiplicatifs des semimartingales. (French) []. Lecture Notes in Mathematics, 721:118– 125, 1979. CODEN LNMAA2. ISBN 3-540-09505-5 (print), 3-540-35189-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070854/.

Yor:1976:RFD

[Yor76] Marc Yor. Une remarque sur les formes de Dirichlet et les semimartingales. (French) []. Lecture Notes in Mathematics, 563:283– 292, 1976. CODEN LNMAA2. ISBN 3-540-08057-0 (print), 3-540-37526-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087584/.

Yor:1977:TDF

[Yor77] Marc Yor. Sur les théories du filtrage et de la prédiction. (French) []. Lecture Notes in Mathematics, 581:257–297, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087195/.

Yor:1978:GFS

[Yor78a] Marc Yor. Grossissement d’une filtration et semi-martingales: Théorèmes généraux. (French) []. Lecture Notes in Mathematics, 649:61–69, 1978. CODEN LNMAA2. ISBN 3-540-08761-3 (print), 3-540-35856-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064595/.

Yor:1978:QRC

[Yor78b] Marc Yor. Quelques résultats sur certaines mesures extremales. Applications à la représentation des martingales. (French) []. Lecture Notes in Mathematics, 695:27–36, 1978. CODEN LN-MAA2. ISBN 3-540-09098-3 (print), 3-540-35556-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0062652/.

Yor:1979:QIE

[Yor79] Marc Yor. Quelques interactions entre mesures vectorielles et intégrales stochastiques. (French) []. Lecture Notes in Mathematics, 713:264–281, 1979. CODEN LNMAA2. ISBN 3-540-09252-8 (print), 3-540-35308-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071330/.

Young:1975:ISL

[You75] David M. Young. Iterative solution of linear and nonlinear systems derived from elliptic partial differential equations. Lecture Notes in Mathematics, 461:265–296, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074157/.

Yamamoto:1979:FEA

[YS79] Yoshiyuki Yamamoto and Yoichi Sumi. Finite element analysis for stress intensity factors. Lecture Notes in Mathematics, 704: 154–168, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063619/.

Zabczyk:1975:NSS

[Zab75] J. Zabczyk. A note on semipolar sets for processes with independent increments. Lecture Notes in Mathematics, 472:277– 283, 1975. CODEN LNMAA2. ISBN 3-540-07190-3 (print), 3-540-37556-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0081957/.

Zagier:1977:CES

[Zag77a] D. Zagier. Correction to “The Eichler–Selberg trace formula on SL_{2}(Z)”. Lecture Notes in Mathematics, 627:171–173, 1977. CO-DEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0065300/.

Zagier:1977:MFW

[Zag77b] D. Zagier. Modular forms whose Fourier coefficients involve zetafunctions of quadratic fields. Lecture Notes in Mathematics, 627: 105–169, 1977. CODEN LNMAA2. ISBN 3-540-08530-0 (print), 3-540-35984-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065299/.

Zanzotto:1977:ENI

[Zan77] P. A. Zanzotto. Sur l’existence d’un noyau induisant un opérateur sous-markovien donné. (French) []. Lecture Notes in Mathematics, 581:298–302, 1977. CODEN LNMAA2. ISBN 3-540-08145-3 (print), 3-540-37383-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087196/.

Zeeman:1976:CEC

[Zee76a] Christopher Zeeman. The classification of elementary catastrophes of codimension ≤ 5. Lecture Notes in Mathematics, 525: 263–327, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0077853/.

Zeeman:1976:BM

[Zee76b] E. C. Zeeman. Brain modelling. Lecture Notes in Mathematics, 525:367–372, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077855/.

Zeeman:1976:EB

[Zee76c] E. C. Zeeman. Euler buckling. Lecture Notes in Mathematics, 525:373–395, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077856/.

Zeeman:1976:GFS

[Zee76d] E. C. Zeeman. Gastrulation and formation of somites in amphibia and birds. Lecture Notes in Mathematics, 525:396–401, 1976. CO-DEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0077857/.

Zeeman:1976:PD

[Zee76e] E. C. Zeeman. Prison disturbances. Lecture Notes in Mathematics, 525:402–406, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077858/.

Zeeman:1976:UBD

[Zee76f] E. C. Zeeman. The umbilic bracelet and the double-cusp catastrophe. Lecture Notes in Mathematics, 525:328–366, 1976. CODEN LNMAA2. ISBN 3-540-07791-X (print), 3-540-38254-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077854/.

Zelinsky:1976:LES

[Zel76a] D. Zelinsky. Long exact sequences and the Brauer group. Lecture Notes in Mathematics, 549:63–70, 1976. CODEN LNMAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0077333/.

Zelinsky:1976:BGP

[Zel76b] Daniel Zelinsky, editor. Brauer Groups: Proceedings of the Conference Held at Evanston, October 11–15, 1975, volume 549 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07989-0 (print), 3-540-37978-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 549. URL http://link.springer.com/book/10.1007/
BFb0077328;   http://www.springerlink.com/content/978-
3-540-37978-2.

Zienkiewicz:1975:VPP

[Zie75] O. C. Zienkiewicz. Visco-plasticity, plasticity, creep and viscoplastic flow. Lecture Notes in Mathematics, 461:297–328, 1975. CODEN LNMAA2. ISBN 3-540-07169-5 (print), 3-540-37503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074158/.

Zielke:1979:AF

[Zie79a] Roland Zielke. Adjoined functions. Lecture Notes in Mathematics, 707:61–78, 1979. CODEN LNMAA2. ISBN 3-540-09125-4 (print), 3-540-35405-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071037/.

Zielke:1979:AP

[Zie79b] Roland Zielke. Analytical properties. Lecture Notes in Mathematics, 707:50–60, 1979. CODEN LNMAA2. ISBN 3-540-09125-4 (print), 3-540-35405-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071036/.

Zielke:1979:DCS

[Zie79c] Roland Zielke. Discontinuous Čebyšev Systems, volume 707 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LN-MAA2. ISBN 3-540-09125-4 (print), 3-540-35405-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 111 pp. LCCN QA3

.L28 no. 707. URL http://link.springer.com/book/10.1007/
BFb0071032;   http://www.springerlink.com/content/978-
3-540-35405-5.

Zielke:1979:DD

[Zie79d] Roland Zielke. Domains of definition. Lecture Notes in Mathematics, 707:1–11, 1979. CODEN LNMAA2. ISBN 3-540-09125-4 (print), 3-540-35405-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071033/.

Zielke:1979:HSR

[Zie79e] Roland Zielke. Haar spaces and related concepts. Lecture Notes in Mathematics, 707:12–28, 1979. CODEN LNMAA2. ISBN 3-540-09125-4 (print), 3-540-35405-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071034/.

Zielke:1979:MP

[Zie79f] Roland Zielke. Miscellaneous properties. Lecture Notes in Mathematics, 707:79–104, 1979. CODEN LNMAA2. ISBN 3-540-09125-4 (print), 3-540-35405-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071038/.

Zielke:1979:S

[Zie79g] Roland Zielke. Subspaces. Lecture Notes in Mathematics, 707: 29–49, 1979. CODEN LNMAA2. ISBN 3-540-09125-4 (print), 3-540-35405-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071035/.

Zienkiewicz:1979:SRC

[ZKB79] O. C. Zienkiewicz, D. W. Kelly, and P. Bettess. The Sommerfeld (radiation) condition on infinite domains and its modelling in numerical procedures. Lecture Notes in Mathematics, 704:169– 203, 1979. CODEN LNMAA2. ISBN 3-540-09123-8 (print), 3-540-35411-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063620/.